A Complete Bibliography of Publications in Foundations of Physics

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

08 March 2024 Version 1.101

Title word cross-reference

$ [2899]. (1 + 1) [2200]. (1 + 2) [3620]. (δ) [3476]. 0 [1937]. 0 ≤ θ < ∞ [1983]. 1 [3093, 2341, 3484, 1937, 1960, 1220]. 1 + 1 [2334, 3641]. 1 − P N [3151]. 1/2 [2799, 2466, 1448, 2145, 843, 1072, 78, 448, 765, 1096]. 1/r [3340, 3187]. 11 [3839]. $110.00/$40.00 [2795]. 120 [3759]. $130.00 [2590]. $137.00 [2496]. $138.00 [2566]. $138/95 [2609]. $139.50 [2785]. $149.00 [2558]. $167.00 [2495]. $168 [2631]. 18 [2562]. $185 [2497]. $186.00 [2494]. $191.00 [2783]. $193.00 [2821]. 2 [2653]. $219.00 [2708]. $23.95 [2498]. $24 [2967]. $25.95 [2624]. $27.00 [2344]. $27.50 [2345]. $28.95 [2819]. 2S [2145]. 3 [2925, 2100, 4087]. $303.00/$99.00 [2903]. $35.00 [2499]. $35.00/$20.00 [2540]. $36.80 [2687]. 4 [1515, 2072, 3274, 3449, 3447, 2730, 4147, 2852]. $42.00 [2541]. $45.00 [2617]. $49.50 [2767]. $49.95 [2539]. 4π [3868]. 5 [838, 2730, 755, 3935]. $55.00 [2766]. 6 [2201]. $60.00 [2565]. 600 [3450]. $70 [2641, 2786]. $75 [2836]. 8 [1072]. $825.00 [2316]. $849.95 [2375]. $89.95 [2820]. $95.00 [2808]. $95.00/$40.00 [2677]. ^{(1)} [3049]. ^{(3)} [1925]. ^{−} [3318]. ^{2} [204]. _{1,3} [1584]. _{c} [2446]. A [2842]. AF C∗ [2272]. α_{s} [2568, 2552]. B [2731, 2925, 3265]. B^{(3)} [1734]. C∗ [4320]. C^{∗} [4267]. D [1889, 2418]. D = 2 [2044]. Δ^{++} [2693]. E [2731, 2925, 764, 2568, 2552]. E = mc^{2} [4442]. e^{+}e^{−} → ΛΛ̄ → π^{−}pπ^{+}p̄ [437]. F [1318, 2279]. f(R) [4738, 4200, 3902, 4226]. f(R, T) = R + αR^{2} + λT [4365]. f_{0}(980) − a_{0}(980) [2701]. G [556, 2197]. G_{2} [1947]. GL [1243]. GM = Rc^{2} [402]. H_{0} [4268]. j [1856]. K [4653, 2687, 412]. k = −1 [3869, 4240]. L [937]. Λ [3027, 2920, 4119]. M [2691, 1586]. M_{2} [1584]. k [4098]. CPT [3018]. PT [2330]. D [3223, 3439]. N [3063, 4347, 442, 3241, 4236]. N_{c} [1621]. Nc [2543]. O [1720]. O(3) [3126]. ω [412]. \overline{SA}(4, R) [1948]. P [1966, 35, 2672, 865, 1764, 1155]. P^{2} [1966]. ψ [1013, 4734, 4564, 4285, 4145, 4353]. Q [1383, 3430]. qr [3308]. R [714, 715, 802, 817, 878, 951, 1198, 1384, 1199, 1321]. R^{3} [2145]. R_{44} [194]. R × S^{3} [936, 1770]. S [3740, 2145, 830, 3600, 2125]. S^{3} [714, 715, 802, 817, 878, 951, 1198, 1384, 1199, 1321]. S^{3}R [2148]. S^{3} × R [1702]. S_{x} [1220]. S_{y} [1220]. S_{z} [1220]. SL(2, C) [2537]. Sp(2, R) [2013]. Sp(6, R) [2013]. SU(1, 1) [2125, 3229]. SU(2) [175, 481]. SU(2) × U(1) [1702]. SU(3) [2637]. SU(4) [2012, 2019]. su(n) [2019]. SU_{2} [936]. T [3817]. ΘΔ^{cs} [4281]. × [714, 715, 802, 817, 878, 951, 1198, 1384, 1199, 1321]. U(1) [1384, 1385, 3935, 2719]. U(3) [2011]. U_{q}(3) [2011]. V [3113, 2094, 3791, 2097]. x [1155].

-absoluteness [4733]. -adic [35]. -Algebraic [4320]. -Algebras [2418, 2272, 1889]. -Body [4347]. -branes [1764]. -Calculus [3430]. -Cell [3759, 3450]. -d [838]. -Differentiation [3223, 3439]. -Dimensional [2201, 4087, 755]. -Epistemic [4564, 4285, 4145]. -Extendibility [4098]. -fermion [1515]. -Field [3265]. -Fold [2562]. -Form [865]. -FRW [3027]. -Fuzzy [2418]. -Gauge [3113]. -Information [2279]. -invariant [1586]. -Kernel [937]. -level [3241, 442]. -Loop [2672]. -Loops [2687, 4653]. -Matrix [3600]. -Momentum [3447]. -numbers [3308]. -Ontic [4734, 4353, 4564, 4285]. -Potential [2852]. -Quantum [3229]. -representation [1584]. -simultaneity [1318]. -Space [2072]. -Spinors [1072]. -Symmetric [2330]. -Symmetry [3126]. -Tense [3740]. -Theory [2691]. -Type [2094, 3791, 2097]. -Wave [2125].

/119.00 [2600]. /89.00 [2600].

0 [2345, 2376, 2590, 2641, 2786, 2795, 2784, 2836, 2495, 2494, 2565, 2539, 2677, 2316, 2808, 2617, 2820, 2496, 2687, 2766, 2315, 2624, 2967]. 0-12-059877-9 [2820]. 0-12-219141-2 [2786]. 0-19-513006-5 [2315]. 0-19-513096-0 [2315]. 0-19-513202-5 [2617]. 0-19-850658-9 [2539]. 0-19-852063-8 [2766]. 0-262-19455-4 [2565]. 0-306-45660-5 [2495]. 0-393-02022-3 [2624]. 0-415-93858-9 [2641]. 0-465-09293-4 [2967]. 0-521-35281-9 [2590]. 0-521-48022-1 [2808]. 0-521-58056-0 [2795]. 0-521-58056-0/0-521-58958-4 [2795]. 0-521-58958-4 [2795].

0-521-59271-2 [2784]. 0-521-59284-4 [2376]. 0-521-82143-6 [2836]. 0-691-01019-6 [2767]. 0-7382-0234-7 [2345]. 0-7923-3565-1 [2496]. 0-7923-5737-X [2494]. 0-7923-6812-6 [2600]. 0-9683689-2-1 [2316]. 0/0 [2795].

1 [2783, 2708, 2609, 2903, 2821, 2785, 2316, 2808, 2496]. 1-4020-0087-1 [2609]. 1-4020-0149-5 [2785]. 1-4020-0932-1 [2708]. 1-4020-1120-2 [2783]. 1-4020-1805-3 [2821]. 1-4020-1850-9 [2903]. 1-419 [2375]. 1-541 [2497]. 10.1007 [3322]. 10.1007/s10701 [3322]. 10.1007/s10701-009-9351-4 [3322]. 19 [3113]. 1941 [2558]. 1990 [1184, 1192]. 1995 [2475]. 1997 [2075, 2474]. 1998 [2104, 2117, 2131, 2139, 2149, 2161, 2171, 2179, 2185, 2194, 2204, 2220, 2232]. 1999 [2246, 2256, 2262, 2268, 2277, 2284, 2293, 2306].

2 [2815, 2786, 2783, 1282, 2784, 2631, 127]. 2-Body [2804]. 2/ISBN [2677]. 2000 [2379, 2368, 2391, 2419, 2430, 2443]. 2003 [2669, 2678, 2689, 2703, 2709, 2716, 2725, 2735, 2743, 2751, 2759, 2769, 2776, 2787]. 2004 [2796, 2802, 2809, 2813, 2822, 2829, 2837]. 2005 [2891]. 2012 [3717]. 20th [2388]. 2nd [2899].

3 [250, 2821, 2687, 2624]. 3-540-43262-0 [2687]. 3/ [2694]. 32 [2688]. 38 [3206].

4 [2376, 2795, 2565, 3322, 2967]. 4D [4446].

5 [2495, 2785, 2617, 2315]. 50th [536].

6 [2836]. 60th [3228]. 65th [2206]. 69.95 [2899].

7 [2345].

8 [2766]. 80th [3347].

9 [2590, 2641, 2903, 2539, 2820]. 96 [2387]. 981 [2631]. 981-02-4039-2 [2631]. 9th [2367, 2378, 2390].

=0 [194].

A. [4462, 3814]. Abandon [2225]. Abelian [688, 4470]. Aberration [3009, 484, 1120]. ABL [2248]. Abnormal [3314]. Abraham [2388, 2499, 2566, 2103]. Absence [2120, 828]. Absolute [4213, 387, 1118, 1119, 242, 352, 339, 388, 432, 465, 351, 880]. absoluteness [4733]. Absorber [4403, 3344, 592]. Absorbers [3627]. Absorption [2441, 2754]. Abstract [2671, 4423]. Absurd [2227]. Academic [2786, 2783, 2708, 2609, 2600, 2474, 2497, 2494, 2903, 2785, 2820, 2496, 2566, 2429]. academy [4627]. Acausality [3730]. accelerate [358]. Accelerated [3032, 3135, 3439, 2905, 3059, 2572, 2722, 2134, 3623, 3909, 4396, 3941, 1219, 869, 4051]. Accelerating [3575, 774, 2995, 1733]. Acceleration [4006, 2661, 4601, 2906, 3160, 2854, 3890, 2280, 3705, 380, 660, 1740]. accept [518]. Accidental [535]. According [2373, 3264, 2561, 3282, 3178, 242, 432, 476]. Account [4572, 4358]. Achieve [4370]. achievements [4117]. Acting [3896]. Action [2826, 2783, 3099, 3990, 3956, 2897, 2331, 2066, 3523, 3436, 4420, 4294, 4413, 2902, 2363, 978, 156, 53, 981, 1462, 104, 4654, 4662, 278, 1186, 1258, 540, 839, 1542]. action-angle [981]. action-reaction [540]. Actions [4577]. Actors [4482]. Acts [2135]. Actual [3392]. adaptation [1591]. Addendum [3049, 299, 3132, 3322, 501]. Addition [2566, 1827, 302]. additive [1261]. Addressing [4268, 3526]. Adequate [3373]. Adiabatic [3118, 562]. Adiabatically [3479]. adic [35]. Adjoint [4028, 1922]. adjointness [624, 1526]. Adjustable [3503]. ADM [2464]. admissible [2018]. admit [1223]. ado [1716]. Adrien [551, 563, 727]. Advanced [1184, 1192]. Adventures [3518]. Adversaries [2948]. Aerts [1200]. Aesthetic [456, 47, 128, 184, 291, 407, 461]. Aestivation [4187, 4691]. Aether [201, 234, 550]. Affect [3649]. Affine [3328, 4608, 3215, 1158]. Afshar [3173, 4496, 3261, 3098]. After [4506, 4422]. Again [4734, 2656, 4499, 3358]. Against [2223, 3845, 3392, 4081, 3050, 4238, 4366, 4336, 3943, 2033, 803, 4476]. Agashe [3049]. Agent [4314]. Aharonov [2111, 2235, 2362, 2361, 2550, 2551, 3157, 3214, 1537, 1930, 3480, 1778, 1305, 399, 1438, 4224, 2713, 4094, 2680, 4363, 1531, 4321, 452, 1050, 4225]. Aims [4004, 4017]. Ain’t [3012]. Ajoy [2903]. al [3190, 3290]. al. [3113]. Alan [2624]. Albert [959, 591]. Alexander [2102]. Alfred [2819]. Algebra [2313, 4660, 4680, 3372, 3664, 2128, 3708, 2519, 2808, 4147, 4446, 2424, 3227, 3738, 416, 1573, 1578, 1094, 645, 1577, 1571, 2010, 1827, 764, 1579, 1586]. algebra-valued [1586]. Algebraic [4020, 2282, 1529, 3231, 338, 1543, 2178, 3525, 3571, 3675, 3783, 3377, 3539, 3023, 3367, 665, 4267, 4320, 438, 1768, 1299, 162, 213, 552, 875, 879, 4626, 2790]. Algebraically [2382]. Algebraization [2062]. Algebras [3564, 3634, 4259, 2620, 3740, 3461, 3580, 3672, 3683, 4172, 2412, 2534, 3211, 2191, 3373, 3897, 3521, 2418, 3204, 3490, 494, 2272, 3567, 4021, 2863, 3755, 1824, 1681, 3371, 1688, 393, 1889, 1975, 1844, 1572, 1881, 1135, 1938, 1833, 1589]. Algorithm [3983]. Algorithmic [4261, 1852]. Algorithms [4065]. Alignment [4567]. Alive [4366, 4336]. Alleged [4344]. alleges [1875]. Allows [2260]. Alloys [2437]. Almost [3665, 3476]. Alone [3646, 3225, 3226]. Along [3297]. already [389]. Also [4542, 1032]. Alternative [4598, 4723, 4051, 4468, 3662, 2630, 3487, 3374, 408, 1240, 1998, 1039]. Always [4157, 711]. Alwyn [2783, 2474]. Am [2952, 3822]. ambience [869]. Ambiguities [4395, 744]. ambiguous [4476]. among [503]. amount [266].

Amplification [1067]. amplifiers [977]. amplitude [1287, 1105]. Amplitudes [2835, 2892, 2463, 4471, 849]. Analog [2852, 661]. Analogies [3471, 754]. Analogs [2079]. analogue [2034]. Analogues [3483, 508]. Analogy [2078, 1691]. Analysis [3278, 4611, 3737, 4203, 3301, 329, 3869, 3962, 4312, 4725, 4730, 4490, 2608, 1023, 1118, 4090, 4142, 2605, 2081, 2870, 2955, 2956, 4431, 4260, 4360, 4702, 4630, 4638, 4643, 4656, 1907, 1826, 1765, 1140, 2033, 1778, 366, 408, 1109, 1119, 4185, 345, 942, 1588, 48, 1354, 761, 899]. Analytic [3027, 222]. Analytical [4124, 2457, 115, 153, 1346]. analyticity [1967]. Analyzers [3038]. Analyzing [4205, 3053]. Anandan [1778]. Anatolij [2497]. Anatomy [4405]. and/or [3065]. Andre [2316, 2269]. anecdotal [1871]. Angle [4523, 734, 981]. angled [713]. Angular [3340, 2287, 2513, 3934, 3051, 3055, 533, 2696, 3385, 2372, 4417, 3968, 380, 841, 268]. Angular-momentum [533]. Anisotropic [3959, 3116, 4334]. anisotropies [4659, 230, 249]. Anisotropy [3475, 4241]. Announcement [683, 872, 850, 834, 916, 931, 1057, 1484, 1607, 1609, 1782, 2180, 2186, 2181, 2294, 2305, 2257, 2347, 2366, 2377, 2367, 2378, 2390, 25, 57, 139, 140, 187, 195, 248, 259, 825, 1083, 1084, 1363, 1371, 1378, 1597, 1616, 1630]. ANNOUNCEMENTS [2195, 2205, 2221, 2233, 2247, 789, 1161, 1874, 1885, 1894, 1904, 1920]. Annular [2848]. anomalies [1211, 1188]. Anomalous [2332, 4447, 2259, 165, 876]. Anomaly [3052, 3077, 4111, 2854, 2973, 2021]. Ansatz [4468]. Answer [3077, 3256, 979]. answering [3625, 3617]. Anthony [2808]. Anthropic [4687, 2641]. Anti [3096, 3406, 2657, 1783]. Anti- [2657]. Anti-Gleason [3096]. antichronous [493]. Anticorrelations [3476]. Antimatter [2734, 1836]. Antimeson [2985]. antineutrinos [695]. Antiparticle [2568, 2552, 2892, 1502, 201, 505]. Antiparticles [493]. Antipodal [3940]. antisymmetric [461]. Any [4016, 3503, 4592]. Anyon [3844]. Anyone [2756]. Apeiron [2316]. apologies [1]. Apparatus [2126, 1921, 1355, 964]. Apparent [4604, 1603, 361]. Appearance [4349]. Applicability [4076, 2073]. Application [3903, 2720, 4231, 2622, 2175, 3743, 3684, 2997, 3374, 2188, 2013, 1552, 1562, 876, 713, 2019, 1354, 2454, 607, 3540]. Applications [3860, 2609, 3051, 2798, 2255, 2259, 3445, 4350, 2734, 3026, 2313, 1895, 2416, 4666, 2005, 1853, 1014, 2015, 4628, 4649, 1264, 2903]. Applied [3217, 3707, 3364, 3443, 82, 2772]. Appraisal [2890, 3359]. Appreciating [3632]. Appreciation [2325]. Approach [2789, 2164, 3863, 3661, 3178, 3014, 3690, 4144, 3621, 4419, 4444, 2198, 4478, 3777, 2545, 2701, 2818, 2996, 2915, 3034, 4160, 4224, 3163, 3444, 2236, 4711, 2437, 2800, 2817, 4317, 2882, 3046, 3574, 3142, 4530, 3694, 4333, 3435, 3326, 2729, 4617, 4456, 2597, 4715, 3333, 2627, 2630, 4589, 2508, 2496, 2874, 3753, 2384, 2705, 4452, 3474, 4375, 952, 433, 1379, 900, 912, 922, 950, 1295, 438, 630, 1347, 1686, 1768, 1234, 532, 552, 961, 12, 22, 36, 804, 857, 926, 1598, 1599, 1576, 1069, 2022, 583, 1419, 1847, 608, 445, 692, 879, 4626, 364, 1039, 911, 4652]. Approach [2708, 2245, 2178]. Approaches [3180, 3785, 3294, 2988, 3844, 813, 1725, 1062]. Approximate [3084, 3107, 2369, 2043]. Approximately [3231]. Approximating [4013, 823, 888]. Approximation [2098, 3752, 4530, 2730, 2907, 2382, 2099, 2359, 4084, 2023, 14, 295]. Approximations [2445]. approximative [583]. April [2139, 2149, 2256, 2262, 2268, 2716, 2725, 2735, 2802, 2891]. Apropos [2763]. Arageorgis [4267]. Arbitrarily [2421]. Arbitrariness [3582, 234]. Arbitrary [2479, 2991, 4044, 2017, 1856, 1983, 77]. arbitrary-order [77]. Archibald [835, 826]. Archimedean [2313, 4471, 4470]. Archimedes [2843]. Architecture [3854]. Area [2655, 1063, 3403]. Areal [4471]. Arena [2727]. arenas [1696]. Arfken [2820]. Argument [3648, 3043, 2223, 4233, 4351, 2673, 4236, 4197, 4235, 3392, 4726, 4346, 4238, 3256, 4230, 3943, 4267, 3692, 3082, 60, 2033, 497, 88, 4331, 1357, 4476, 1477, 1067]. Argumentation [2608]. Arguments [3243, 3548, 4431, 114, 607]. Arising [2282, 3445, 517]. Arithmetic [1228]. Armas [4448]. Arrow [3143, 4712, 2704, 2974, 2673, 4509, 3585, 3285, 2133, 592, 1516, 4206]. arrows [1079]. Art [4278, 3855, 3221]. Article [4277, 16]. Artificial [3908]. Asher [2961]. Asim [2077, 1499, 2076]. Asked [3688]. Aspect [2733, 1122, 159, 816]. Aspects [3351, 3951, 581, 1401, 4090, 4565, 4252, 2606, 3848, 2997, 4176, 3908, 2635, 92, 272, 1113, 1077, 974, 1023, 1016, 1617, 1258, 1300]. assembles [497]. assembly [1296]. Assessment [3611, 3127, 3128, 2824]. Assessments [4196]. assignments [642]. Assis [2316]. Associated [2162, 3836, 2719, 3464, 410, 537, 1298]. association [513]. Assuming [4477]. Assumption [3031, 521]. Assumptions [4467, 3811, 2067]. astronomical [1120]. Astronomy [2450]. Astrophysical [4148, 1176]. Astrophysics [4004, 2355]. Asymmetric [2333]. Asymmetrically [2383]. asymmetries [1343]. Asymmetry [3143, 2401, 4474, 359, 679, 87, 1357, 361]. Asymptotic [3674, 4128, 2351, 743, 758]. Asymptotically [4195, 1846, 1060]. Asynchronous [3896]. Atkinson [3806]. atmospheric [89]. Atom [2461, 3520, 3542, 4707, 2095, 3941, 1557, 828, 301, 408, 1536, 1825, 1519]. Atomic [2050, 2188, 3927, 3580, 3576, 2093, 3586, 3019, 2097, 1317, 287]. Atomism [4283]. Atoms [4050, 2385, 2092, 3306, 2828, 4722, 1150, 2645, 2969, 2619, 538]. Attempting [4404]. Attractive [3757, 2353]. attractors [1728]. attributes [95, 439]. August [1147, 1185, 1193, 2377, 1138, 2179, 2185, 2430, 2443, 2769, 2776, 2813]. Auletta [2631]. Author [2196, 2304, 2442, 2548, 2651, 2758]. Autistic [2842]. Autler [3791]. Automata [3847, 3852, 2322]. automatic [841]. Automaton [3641]. Automorphisms [1348]. Autonomous [2377, 2473]. Autosoliton [2363]. Available [2316, 209]. avenues [561]. Averaged [2461, 3051, 4481, 189]. averages [1743]. Averaging [4209]. avoid [1404]. Avoiding [3988]. Avoids [4229]. Away [4335]. Axial [4260, 498, 77, 101]. Axiom [3339, 199, 1619, 1992]. axiom-preserving [1992]. Axiomatic [3949, 1642, 2282, 3170, 2773, 2791, 583]. Axiomatics [86, 3282, 27]. Axioms [4162, 3529, 2807, 4142, 2738, 228, 617, 618, 2002].

B [1925, 3011, 2617, 2429, 67, 3937]. Ba [2013]. Back [2020, 4499, 1542]. Back-action [1542]. Back-reaction [2020]. Backflow [2088]. Background [2777, 3448, 3904, 2320, 489]. Background-Based [3904]. backgrounds [4121]. Backreaction [2654, 2653]. baker [1244]. Balanced [3404, 3201]. balancing [597]. Ball [2702]. Ballistic [2465]. Balls [3251]. Band [2433]. Bands [2100, 1063]. Bang [4143, 3280, 200, 4240, 4113]. Bang/Crunch [4143]. Barbieri [2114]. Barbour [2617]. Bare [4188]. Bargmann [3594, 3501]. Barrier [867, 2454, 1884]. Barriers [3733, 451, 70]. Barut [2077, 1499, 1502, 2076]. Baryon [2543]. Baryons [2697]. Based [3863, 4016, 4340, 4480, 2896, 4144, 3308, 3758, 3946, 4419, 4444, 2636, 3108, 3798, 3862, 3311, 3254, 3445, 2519, 3662, 3419, 4589, 4298, 4595, 4054, 2173, 4597, 2530, 4176, 4045, 3153, 3904, 3450, 3500, 3759, 567, 1304, 1714, 66, 526, 1263, 1264, 409, 2012, 394]. Bases [4604, 4167, 129]. Basic [2387, 2733, 2896, 2555, 2322, 2521, 4253, 3011, 3073, 2967]. Basil [3633]. Basis [2806, 3949, 2893, 2994, 3374, 370, 1739, 732, 382, 1120]. Bass [2932]. Bassi [3112]. Batrouni [2836]. Batteries [2753]. Bavaria [1184, 1192]. Bayes [1288, 801]. Bayesian [3898, 3415, 702]. BCS [340]. BDC [4323]. Be [3300, 2648, 3646, 2579, 2823, 4519, 3678, 3467, 3226, 711, 4506, 4193, 71, 1033, 1040, 3718, 3737, 1032, 4578, 3809, 3077, 3364, 3416, 4475, 373, 2953, 403, 3030, 1259, 3162, 4257, 3152, 1743, 3192]. Beables [3390, 1304, 1293]. Beam [3510, 2728, 2097, 987]. Beamsplitter [3896]. Bearing [891]. Becoming [4405, 3637, 191]. Been [3725, 389]. Before [3256, 3162, 3192, 1000]. Behavior [4483, 2308, 2584, 4325, 4086, 2120, 4497, 2619, 743, 1203, 767, 1970, 128, 4627]. Behaviors [3919, 4373]. Behind [4019]. Being [3637]. Beings [2772]. Bekenstein [2668]. Believe [2842, 2852, 24]. Belinfante [1067]. Bell [1681, 3476, 4052, 2955, 793, 2956, 2423, 1302, 572, 1300, 1412, 2394, 3217, 1275, 2935, 1678, 710, 1269, 1265, 3337, 4000, 2985, 60, 3544, 1450, 2942, 4376, 3739, 3592, 1768, 2414, 2190, 3561, 1982, 1714, 3804, 1383, 497, 1116, 88, 4516, 1930, 3424, 1288, 1214, 2700, 3897, 4508, 4482, 1443, 3667, 315, 2932, 3846, 1266, 1449, 2608, 3850, 4331, 3675, 1645, 3821, 3986, 4047, 4399, 3195, 212, 2950, 4024, 1069, 3434, 3985, 3213, 4096, 2228, 2560, 3835, 792, 1891, 4052, 1303, 2840, 3724, 3972, 910, 4145, 1325, 4607, 364, 1292, 3206, 3653, 4524, 4575]. Bell [3450, 3673, 3152, 2983, 439]. Bell-Statistical [4508]. Bell-Type [2955, 2956, 1681, 3835, 1678, 1443]. Bellac [2836]. Beller [2540]. Below [2341]. Ben [4031]. bending [362]. Benefits [3700]. Berezin [937]. Bergmann [760, 752, 742]. Berkeley [2252]. Berlin [2899, 2687]. Bernoulli [2482, 2612]. Berry [3763, 3326, 3193, 3516]. Best [4308]. Beta [2688, 2024, 2611, 2698]. Bethe [1960]. betting [642]. Between [2898, 4050, 3873, 4604, 3412, 3794, 3347, 2449, 3775, 2295, 4469, 3789, 2594, 2595, 3508, 3294, 2599, 4584, 4182, 2793, 4323, 2675, 4459, 2581, 2803, 4285, 4311, 4710, 3637, 3641, 989, 21, 1537, 1317, 642, 3341, 3031, 1006, 1277, 1828, 1691, 1153, 2954, 4263, 285, 754, 324, 3218, 1173, 1087, 1657, 898, 472, 77, 101]. Beyond [2571, 4101, 3947, 4610, 4097, 3254, 2480, 3743, 4530, 4199, 4554, 2111, 2043, 2566]. Bi [4143, 4557]. Bi-directional [4143]. Bi-Hamilton [4557]. Bianchi [1034, 841, 3902]. Bias [3240, 2641, 4484, 4687]. Bibliography [760, 1954, 2231]. Biedenharn [2011, 2007]. Bifurcation [2843, 2457, 2384]. Big [3280, 200, 4143, 4240, 4113]. Bilinear [3672]. Billiard [2479, 3249, 2451, 2469, 4206]. Billiards [2460, 2462]. bimetric [679, 600, 1086, 736, 558, 1107, 1175]. Binary [3895]. Binding [4139]. Biography [2539, 4640]. Biological [275, 1301]. Biology [3860, 4276, 508, 2734, 2753, 21, 266]. Biot [1605]. Bipartite [4219, 3448]. Bird’s [2516]. Bird’s-Eye [2516]. Birkhoffian [1510]. Birth [3814, 853, 965]. Birthday [2206, 2325, 2431, 2444, 2456, 2478, 3228, 3347, 2968, 2838, 529, 545, 502, 675, 1499, 2036, 1047, 1266, 1864, 958, 536, 591, 1944]. Bispinor [2308]. Bit [3886, 2489, 3298, 4370, 4055]. Bitbol [2184]. Bivariate [648, 3835]. BKS [3341]. BL [3371]. BL-algebras [3371]. Black [2918, 4369, 2658, 2646, 4109, 4124, 3310, 2672, 2812, 3754, 2356, 3136, 2664, 4116, 3998, 4730, 2656, 4157, 4043, 3338, 3726, 3168, 4357, 3090, 2645, 4082, 2663, 3743, 3803, 3973, 4069, 4701, 4414, 4520, 3064, 2654, 4126, 2667, 858, 2657, 2653, 3940, 4110, 492, 693, 2387]. black-body [693]. Black-Hole [2356, 2664, 2663]. Black-Swan [3726]. Blackbody [3083, 3342, 3547, 2421, 1113]. Blackett [745, 1176]. Blobs [3635]. Bloch [2578, 1786, 4488]. blocks [1991]. Blockworld [3637, 3148]. Bodies [2661, 4591, 785, 712]. Body [2724, 4380, 2774, 2775, 3351, 3691, 4573, 2494, 4530, 4338, 2804, 4347, 2328, 2140, 2151, 1961, 1539, 157, 1614, 573, 121, 225, 371, 1114, 693, 4636]. Bogolubov [2656, 822]. Bohm [4610, 4524, 4575, 1058, 1071, 1075, 2111, 3910, 4138, 2235, 2888, 2362, 2361, 2550, 2551, 3157, 1730, 2884, 3214, 469, 4572, 1537, 4543, 4478, 1930, 3480, 1305, 3922, 1438, 4224, 3294, 3521, 1467, 1486, 2713, 1047, 3658, 3297, 1598, 1599, 1725, 4094, 659, 3932, 1735, 2273, 2882, 4566, 2680, 4367, 4363, 1531, 1727, 2887, 1313, 4589, 4321, 452, 1050, 1999, 1809, 1600, 1801, 1049, 3147, 4225]. Bohmian [4531, 2298, 2297, 3200, 2314, 4435, 4133, 2057, 4578, 2880, 3980, 4718, 3293, 3303, 2236, 4500, 2881, 3249, 4505, 3179, 3260, 4507, 2397, 4263, 4486, 3257, 4512, 2873, 3693, 3930]. Bohr [2876, 1140, 1250, 3592, 3889, 166, 959, 3539, 717, 1063, 2824, 893]. Bohrian [3776]. Bohumianism [4279]. boil [2047]. boils [227]. Boltzmann [4068, 3994, 4162, 1792, 2673, 753, 1924, 4715, 4283, 2428, 4063]. Boltzmannian [4374]. Bondi [3403]. Bonus [1553]. Book [1626, 953, 2090, 808, 1748, 1905, 2344, 2345, 2375, 2376, 2387, 2388, 2389, 2590, 884, 1744, 622, 1377, 1483, 2254, 929, 1115, 3482, 923, 1915, 722, 1872, 1916, 915, 1038, 1664, 1429, 1816, 721, 1463, 1709, 1255, 938, 1942, 1649, 1493, 956, 1703, 1082, 925, 737, 824, 2641, 2786, 1722, 1830, 1941, 1956, 1330, 928, 1093, 1126, 1272, 1745, 1839, 1860, 2783, 1290, 883, 639, 603, 1561, 2708, 1862, 1902, 1652, 999, 707, 2609, 1388, 619, 769, 1673, 1567, 1795, 1796, 1797, 1798, 882, 2600, 2242, 1257, 1167, 1432, 1595, 1169, 1127, 650, 1037, 1665, 666, 2795, 1918, 723]. Book [1861, 2244, 2245, 1480, 1560, 1362, 1496, 1350, 771, 1566, 1569, 1497, 930, 778, 2784, 2836, 1817, 726, 1030, 1102, 1387, 2499, 2767, 1662, 2474, 2495, 1433, 1927, 1788, 672, 2178, 2283, 2497, 2541, 1495, 945, 738, 658, 1831, 1101, 1668, 1274, 1814, 739, 895, 1565, 1430, 1815, 807, 627, 809, 914, 1361, 1482, 628, 2494, 913, 2565, 955, 787, 2631, 1737, 1386, 1391, 805, 1029, 1331, 651, 2255, 657, 795, 2899, 3387, 652, 1370, 706, 1629, 1100, 2558, 1137, 1928, 1893, 2903, 2539, 788, 1182, 1628, 1168, 986, 632, 1045, 1390, 1917, 1221, 2128, 2821, 1672, 1559]. Book [602, 740, 1492, 1919, 2677, 777, 1926, 1256, 620, 1650, 1494, 2303, 1654, 1428, 1291, 2292, 2102, 1092, 2184, 1873, 673, 806, 1273, 903, 1046, 2785, 1822, 741, 1323, 1324, 1625, 1594, 1747, 1306, 998, 779, 944, 1098, 1779, 1943, 2316, 924, 3010, 2954, 984, 2540, 776, 1666, 631, 1002, 1376, 1674, 904, 1653, 1349, 1490, 4042, 939, 667, 1749, 1651, 1332, 2808, 1818, 638, 4089, 724, 885, 1431, 1806, 1207, 906, 719, 2193, 1663, 1840, 1596, 2291, 996, 796, 780, 1558, 2794, 1568, 1648, 1805, 1935, 997, 1181, 3011, 1841, 1859, 2498, 2617, 1819, 954, 2820, 985, 1479, 1460, 2496]. Book [1914, 1667, 674, 881, 1933, 1191, 775, 1375, 594, 2475, 907, 797, 1289, 644, 786, 1624, 2687, 1821, 2766, 1623, 1820, 1746, 957, 1212, 1389, 1934, 2566, 1481, 601, 770, 1099, 1339, 896, 1780, 725, 2103, 905, 1627, 1901, 1955, 2241, 2315, 2365, 2624, 2819, 2967, 720, 894, 1360, 621, 3221, 3814, 2243, 2101, 1459, 1491, 2429]. Books [2345, 3011, 2967]. Boolean [1681]. boosts [1827]. bootstrap [90]. Bootstrapping [548, 3849]. Bootstraps [2886]. Borders [4435]. Born [2389, 4255, 4308, 2901, 3577, 3947, 3187, 4598, 4716, 3925, 3254, 4090, 4142, 3542, 2869, 3558, 4554, 2445, 3760, 3968]. borne [899]. Bose [3929, 3703, 1910, 1369, 3456, 1018, 2439]. Boson [2436, 3260, 538, 4056]. Bosonic [3807, 3641, 1384, 1702, 1636, 1697]. Bosons [2511, 3936, 3402, 2656, 3325, 2953, 1369, 1937]. Boston [2708, 2609, 2600, 2496, 2429]. Boston/Dordrecht/London [2609]. Bostrom [2641]. Both [2870]. Bottle [3395]. bottles [1237]. Bounce [4701]. Bouncing [4014, 4007]. Bound [2936, 2622, 4301, 1509, 4512, 2114, 766, 2039, 30, 1638, 1613]. Bound-State [2114]. Boundary [4483, 2200, 1551, 4703, 3677, 839, 4620, 857, 831, 2102]. Bounded [3511]. Bounds [4100, 4018, 3788, 3820, 1515]. Box [3905, 2248, 155]. Boxes [3841]. Boyer [2680]. Bracket [4426]. Brackets [4271, 1582, 746]. Brain [2681, 324, 1396, 377]. brain-mind [1396]. Branch [3046]. Branching [1907, 542]. Brane [2667]. Branes [3091, 3650, 1764]. breakdown [412]. Breaking [3661, 3457, 2270, 2326, 4277, 3750, 851, 3748, 1160, 598, 2032, 174]. Breit [1961]. Bremsstrahlung [1546]. Bretthorst [2784]. Breuer [2766]. Brian [2819]. Bridges [3976]. Brief [4082, 4416, 3522]. Brightness [3055]. British [1905]. Broglian [1471]. Broglie [4138, 2884, 4572, 4610, 4543, 1467, 3658, 3297, 2882, 4566, 1999, 1600, 1801, 4524, 4575, 512, 519, 1470, 1271, 1466, 517, 3166, 992, 2853, 4568, 4699, 4550, 521, 4583, 947, 2574, 3514, 511, 1006, 991, 1489, 4697, 304, 104, 506, 510, 1298, 1485, 718, 513, 1477, 515]. Broken [2289, 2258, 2087, 2886, 2439, 1314]. Brouwer [2417, 1234]. Brownian [431, 2299, 3396, 3686, 2907, 2914, 2908, 3074]. BRST [3406, 3039]. Brukner [4489, 3924]. BTZ [3310]. Bub [469, 4343, 88, 659, 2101]. Bubble [2846, 4354]. Budapest [4685]. Building [4584]. Built [3020]. Bulk [2931]. Bullet [2873]. Bundle [4333, 4591, 1380, 1585]. Bundles [3966]. Bunge [27]. Burgers [3021]. Buridan [3578]. Burst [4126]. Bursts [2355]. Busch [4108, 4163, 2271]. Busch-Type [2271]. BZ [1784, 2416].

C [2431, 2444, 2456, 2477, 2494, 2316]. Cables [2373]. Calculability [2724]. Calculated [890]. Calculation [2189, 2385, 4713, 1580, 2024]. calculations [1445, 2023]. Calculus [3886, 3383, 3430, 2900, 4177, 2121, 3164, 98]. Caliber [3749]. Calibration [2534]. California [2786, 2820]. Call [1762, 1771, 1781, 1978, 2006, 2294, 2305, 2347, 850, 1616, 1630]. Cambridge [2345, 2375, 2376, 2389, 2590, 2795, 2784, 2836, 2541, 2565, 2808, 2498]. Came [2656]. Camel [3209, 4375]. Camilleri [3387]. Can [711, 71, 1411, 1033, 1040, 3718, 3300, 2648, 4578, 3809, 3280, 4718, 4559, 2579, 2823, 4258, 3077, 3364, 3416, 4475, 373, 3467, 3586, 1320, 3030, 2132, 4071, 4324, 327, 3226, 3162, 1809, 4257, 2873, 2667, 3152, 1743, 3192, 4034, 992, 1032, 1545, 4491, 358, 403, 3923, 3678]. Canada [2316]. Cancellations [4161]. Cannot [3646, 4242]. Canonical [3599, 2167, 3900, 909, 3446, 273, 836, 2765, 3604, 2153, 4430, 3714, 1606, 1072, 485, 482, 190, 1657]. Canonically [3837]. Capacities [4231]. Capacitors [2832]. Cappelli [3814]. Capture [3318]. Carathéodory [2582]. carbon [3748]. Cardinality [4609]. cardinals [3473]. Carmeli [2995, 911]. Carmi [399]. Carnot [969]. Carriers [508]. Carroll [3792]. Carry [2650]. Carrying [2237, 2532]. Cartan [3357, 2914, 3074, 2197, 3164, 4551]. Cartesian [2536, 3395]. cascade [1317]. Cascades [2457]. Case [4076, 2895, 4305, 2309, 3998, 3590, 3190, 2932, 4491, 4024, 4521, 3535, 3249, 2274, 4456, 4366, 4336, 2557, 3711, 3844, 4384, 3866, 3977, 4457, 4466, 1284, 1177, 1443, 1398, 1163, 1104, 2172, 845]. Case-Study [4366, 4336]. Casher [2713]. Casimir [1219, 2355, 1537, 2781]. Castell [2899]. Castellani [3814]. Cat [3630, 2952, 4258, 1011]. Categorial [3670]. Categorical [3565]. Categoricity [4587]. Categories [3636, 3566]. Categorizing [2988]. Category [2872]. Cats [4737]. Cauchy [4171, 4245]. Causal [52, 4454, 4072, 3995, 4535, 3636, 3271, 4305, 162, 2664, 3815, 4275, 3476, 3846, 1800, 390, 3443, 2605, 4453, 4252, 3282, 2977, 4320, 4474, 766, 1726, 647, 961, 942, 111, 1600, 1801]. Causality [3315, 4083, 3499, 3846, 4310, 1965, 3191, 2850, 3213, 2569, 3600, 2684, 4137, 99, 317, 942, 960, 979]. Causally [2823, 4502]. Causation [4207, 3375]. Cause [3163, 4289, 2991, 3525, 3541, 4044, 2569, 2251, 3346, 2026, 327]. Caused [2207, 4447]. Causes [4366, 4336, 4034]. causility [1061]. Cavalleri [613, 614, 612]. Cavities [2222, 2421, 1545, 1542]. Cavity [3953, 1598, 1599]. CDM [4119]. celebrated [4627]. Cell [3759, 3450, 541]. cells [70]. Cellular [3847, 3852, 3641]. censor [685]. Censors [2992]. Censorship [4357, 4503, 2029]. Center [2508, 1636, 540]. center-of-mass [1636]. Central [768, 1574, 4477, 2803]. Century [2388, 2855, 3745, 4572]. certain [1844]. Certainties [4502]. certainty [4627, 2241]. CFT [3274, 3449]. chain [2013, 136]. Chains [3621, 3232, 4098]. challenge [1865]. Chamber [2846, 3780, 4362, 4070]. Chance [2254, 2003, 4624]. Change [3836, 4424, 3955]. Changed [2579]. Changes [3785, 3545, 4570, 287]. Changing [2150, 4080, 3941]. Channel [3405]. Channeling [3166]. Channels [3573, 3572, 3590, 3798]. Chaos [2376, 1852, 2458, 3120, 2471, 2464, 4634, 4512, 1087, 1979, 820, 1967, 2022, 934, 4641, 4677, 973]. Chaotic [2479, 2470, 2459, 1970]. chaotic-like [1970]. Chaplygin [4148]. Character [4091, 4025, 1111, 708, 1424, 1136]. Characterizability [2360]. Characterization [3862, 2982, 3838, 1750]. Characterizations [3234]. Characterizing [3761, 2744]. Charge [2568, 2552, 3072, 3871, 2550, 3083, 4698, 4288, 2812, 3520, 4157, 2505, 2721, 3135, 2905, 3304, 3960, 3144, 3185, 2572, 2722, 278, 4390, 2424, 3623, 3909, 4396, 4417, 107, 78, 1720, 1552, 1562]. Charge-Symmetric [3960]. Charged [2646, 3396, 2506, 2630, 2788, 1769, 1106, 1592, 1384, 443, 759, 285, 1520]. Charges [2532, 3404, 3622, 3782, 774, 341]. Charles [1238, 1248, 1249, 1222]. Cheap [4237]. Chemical [2435]. Chemistry [2609, 3965, 2734, 4390, 3073, 4649]. Cheng [4603]. Chicago [2540]. Childhood [2239]. Chiral [677, 2326, 2543, 2044, 1963, 2021]. CHN [1948]. Choice [3278, 2358, 3671, 3885, 3190, 4179, 2081, 2940, 1042, 1603, 3290]. Choosing [3546]. Chris [2808]. Chronogenesis [3652]. chronons [1777]. chronous [493]. CHSH [3819, 3986, 4257]. Chuang [2541]. Circle [2775, 4470, 1711]. Circle-Valued [4470]. Circuit [3791, 861]. Circuits [3744, 2087]. Circulating [2728]. City [2239]. Cl [1584]. Claim [3611]. Claimed [3972]. Clarification [3070, 361]. Clarify [4562]. Clarifying [4332]. Class [4011, 3503, 582, 1854, 1995, 38]. Classes [3271]. Classic [3474, 691]. Classical [2898, 2398, 4483, 3199, 3022, 3661, 2556, 4385, 4322, 4337, 3699, 2308, 2074, 2187, 3867, 3301, 2876, 4514, 3995, 4593, 2888, 2362, 2550, 3083, 3342, 3547, 3663, 3927, 4698, 3412, 3343, 3794, 2209, 2115, 2449, 2999, 3155, 4511, 3113, 3484, 1205, 2295, 2421, 2485, 3224, 1134, 4083, 3789, 3458, 3774, 2116, 1507, 3707, 2079, 3096, 1706, 2966, 4364, 468, 3264, 1522, 2165, 4135, 2091, 2915, 2594, 2595, 2526, 4325, 4086, 4293, 3323, 3478, 2897, 3258, 3922, 1438, 3303, 2906, 4214, 3135, 3797, 4590, 2565, 1374, 4614, 4615, 4631, 4370, 3446, 4182, 2990, 4301, 3827, 2259, 2793, 3254]. Classical [3336, 3413, 3819, 4307, 4015, 3076, 3999, 3168, 3130, 1574, 2817, 2159, 941, 3932, 3385, 1997, 3414, 2120, 3129, 3249, 3958, 3212, 2506, 1691, 3266, 2278, 3820, 3426, 3952, 3471, 2630, 1767, 4589, 2153, 2154, 3639, 4540, 3623, 2930, 2078, 4176, 3579, 2004, 2958, 3938, 4533, 4408, 4700, 2140, 4636, 3742, 3429, 2619, 2182, 2810, 3626, 4105, 2965, 3477, 3016, 1689, 1041, 1343, 1271, 1582, 4605, 1106, 1592, 1236, 1108, 1178, 1538, 821, 820, 107, 328, 1545, 1693, 207, 633, 1203, 192, 1319, 1921, 830, 4026, 148, 206, 1527, 856, 1418, 1581, 1449, 214, 767, 1958, 981, 669]. classical [1634, 371, 447, 358, 704, 874, 1700, 1707, 537, 503, 1133, 1173, 14, 1354, 482, 1671, 4650, 1230, 3528, 1994, 1993, 1422, 625, 932, 1087, 1657, 173, 656, 2245, 2494, 2303, 2193]. Classical-like [1997]. Classical-Quantum [3958]. Classicality [3182, 4595, 3175]. Classically [3386]. Classification [1541, 4521, 4285, 321]. Clauser [3737, 2685]. Clausius [4374]. Clean [2609, 4649]. Clearwater [2344]. Click [2490]. Clicks [2615]. Clifford [2315, 3634, 2812, 2901, 3754, 645, 3521, 1572, 2519, 2727, 2926, 494, 1585, 1938, 1833, 1589, 2424, 1586, 2863, 2128]. Clifford-Algebra [2519]. Climbing [2633, 4640, 2539]. Clock [3166, 3685, 3874, 387, 355, 732, 277, 414, 455, 899, 454]. Clocks [4526, 3950, 4053, 3685, 3087, 2844, 2854, 3154, 4033, 660, 1114]. Clone [3298]. Cloning [2746, 3427, 2949, 3783]. Closed [2658, 3531, 3989, 2919, 2823, 3530, 838, 4453, 3287, 3358, 3194, 1056, 3551]. Closer [3252]. Closing [1711]. Closure [4275]. Cloud [3780, 4362, 4070]. Cloud-Chamber [4070]. Clue [4390]. Clumsiness [3526]. Cluster [2724, 2816, 3438]. CM [2471]. CM-Type [2471]. CMB [4114]. Co [4383]. Co-Existence [4383]. Coaccelerated [2921]. Coarse [2487, 262]. coarse-graining [262]. Coarsened [3816]. Coaxial [2373]. code [307]. coefficients [1734, 2019]. Coexistence [1677, 4397]. Cognition [2783, 3246, 4654, 4662]. Cognitive [3860, 2259]. Coherence [2463, 3926, 977, 1070, 858]. Coherent [2602, 2937, 3078, 3089, 3133, 3219, 2339, 4310, 1171, 3658, 2448, 3828, 3084, 3107, 1284, 1500]. Cohomology [2756]. coincidence [74, 1189]. coincidences [231]. Coincidental [1519]. coins [464]. Cold [2092, 4573, 2699]. Colin [2344, 2283]. Collapse [4695, 2848, 4131, 3486, 4002, 4127, 3927, 2740, 3721, 3656, 4714, 2056, 4156, 2831, 4377, 3896, 3834, 3262, 4539, 2216, 3700, 3803, 3504, 3684, 2217, 2380, 3506, 3652, 4724, 4596, 4366, 4336, 486, 3064, 4415, 4582, 2064, 2684, 4700, 4034, 1799, 803, 1790, 246, 1936, 194, 1544, 1419, 1851]. collapse-free [1936]. Collapse-induced [2056]. Collapses [3669]. Colleague [3632]. Collective [2274, 2281, 2041]. Collide [2715]. Colliding [3251, 3283]. collision [342, 534, 184]. Collisions [3614, 2342, 2337, 4347, 316, 2020, 2046]. Colomo [3814]. Color [2912]. Colored [4131]. Columbia [2240]. Combine [3861]. Combining [1013]. Come [2490, 1524]. Comic [4343]. Comment [2334, 4181, 3993, 4187, 3157, 3132, 4186, 4082, 3985, 4462, 4217, 1136, 4603, 3623, 3206, 4031, 3044, 3192, 421, 495, 3113]. Comments [596, 613, 3112, 3173, 97, 3220, 93, 352, 79, 385, 793, 1705, 2956, 4267, 2177, 568, 336, 386]. Commitment [2489, 4055]. Common [3163, 4289, 2991, 3476, 3525, 2949, 395, 3541, 4044, 2569, 4457, 4466, 3448, 3346, 2026]. Communication [2745, 2981, 3865, 4355, 4327, 3713, 3912, 2064, 3942, 353]. Communications [1616, 1630, 2283]. communicator [526]. Commutation [3446, 2791, 485]. Commutative [3233, 3764, 3234, 3567, 3603]. commutativity [2038, 623]. Commutator [4604]. commute [1966, 763]. comoving [484]. compacifications [1550]. Compact [1508, 3402, 4573, 3078, 3089, 3133, 3219, 1773, 1807]. compactified [1772]. Comparable [3982]. Comparative [2013, 2023]. compared [4411]. Comparison [3873, 3045]. COMPASS [3334]. Compatibility [2957, 3909, 462, 1677, 1317, 1998]. Compatible [1719, 4167, 4298, 763, 2585, 982, 522, 564]. Competing [2437]. Complement [2928]. Complementarity [2571, 2940, 11, 3565, 3889, 2929, 3365, 4737, 4025, 3776, 2879, 3098, 2965, 3384, 959]. Complementary [272, 4166, 2930, 1724, 988]. Complete [3886, 3751, 2946, 3212, 3697, 1554, 373, 1259, 475]. Completed [3484]. Completely [1285, 4717, 3130, 474, 428]. Completeness [2818, 3170, 3213, 4587, 2082, 1261]. Completing [4065, 4391]. Completion [2558, 1992]. Complex [4483, 3339, 4133, 137, 3645, 3658, 4498, 3708, 3997, 4386, 2454, 3084, 3107, 1785, 314, 1341, 1510]. Complex-Domain [2454]. Complex-Valued [4483]. Complexifier [3828]. Complexity [2745, 2981, 4037, 4472, 4474, 1852, 864, 291, 467]. component [677, 78]. Components [2814, 3032, 1096, 478, 29, 48]. Composite [3824, 2637, 1622, 3336, 2059, 2324, 530]. Composite-Velocity [2324]. Composition [4099, 1179, 911]. Compositions [3939]. Compound [2135, 2145, 3735, 2051, 4202, 3856]. Compoundation [2397]. Compressed [3008]. Compressibility [4037, 1883]. Computability [860]. Computation [3158, 4097, 2541, 3248, 861, 3686, 3153, 1838, 1932]. Computational [2918, 3531, 4065, 4205, 3231, 2971, 2128]. Compute [2132]. Computer [2633, 3720, 3285, 1548, 1132, 1131]. Computers [4207, 859]. Computing [2344, 2283, 2202, 4464, 3742]. concatenated [609]. Concatenation [2341]. Concealed [3786]. concentration [270]. Concept [3681, 3275, 2839, 3304, 3385, 2192, 4507, 4350, 1041, 966, 1097, 579, 410, 244, 1314, 33, 5, 18, 1998, 1044]. Conception [2276, 2146, 1987, 1489, 1829, 656]. Concepts [2387, 46, 1229, 4142, 3496, 207, 439, 2365]. Conceptual [3951, 2686, 4621, 3377, 3930, 1871, 1043, 2495]. conceptualization [1488]. Conceptually [4024]. Concerning [2683, 3637, 869, 399, 7, 784]. Conciliate [3772]. Conclusive [2580]. concrete [2027]. Condensates [3703, 2332, 3456, 3064]. Condensation [3929, 2439, 1369]. Condensed [2848, 3795, 3202, 1314, 1408]. condensed-state [1408]. Condition [2374, 3846, 560, 4408, 107, 1720, 562, 1477]. Conditional [963, 3224, 3819, 4711, 3001, 3142, 4294, 2986, 2080, 1243, 4413]. Conditions [4483, 2200, 4037, 4118, 3555, 2199, 3551, 3522, 1551]. Conducting [2421]. Conductor [2532]. Conductors [2753]. Cone [2334, 746, 1254, 608]. Cones [2771]. Conference [1338, 1003, 4569, 872, 1570, 897]. Confessions [4237]. Configuration [4561, 4026, 4086, 3937, 4422, 4389, 3333]. Configurations [4148, 2094, 1466]. confined [711]. Confinement [2207, 3751, 350]. confirmation [4510]. Confirming [2896]. Conflict [4506, 3362, 2599]. Conform [1458]. Conformal [1696, 1108, 1550, 3328, 4292, 2588, 3400, 2084, 3535, 4030, 4111, 2719, 3746, 3282, 3643, 1511, 447, 643, 1687, 1845].

Conformally [1772, 3524, 2657, 1525]. confrontation [847]. Confronts [3463]. Conjecture [4371, 3895, 547, 1443, 1176]. conjectures [446]. Conjugate [2856, 3837, 973, 2260, 909]. Connected [4376, 4422, 3221, 2834, 420]. Connecting [4131, 3083, 3321, 3322]. Connection [2937, 3799, 4165, 2803, 4263, 4710, 4049, 4176, 1396, 439]. Connections [1380, 2295, 1104, 1998]. Consciousness [4377, 3137, 2994, 2176, 2300, 3110, 4366, 4336, 701, 2944, 4034, 245, 876, 1188, 4476, 4637, 4642, 50, 2345]. Consciousness-Causes-Collapse [4034]. consciousness-related [1188]. Consequence [4477, 2831]. Consequences [3000, 3049, 185, 2860, 3384, 3811, 3713, 3060, 4425, 4707, 4402, 1772, 1996, 223, 1752, 379, 298]. Conservation [3468, 2287, 2943, 4293, 2949, 3304, 2271, 2380, 4208, 4415, 290, 1284, 530, 4398, 765, 841, 1851, 2028, 375, 495, 664, 4403]. Conservative [3679, 3517, 309]. Conserve [4157]. Conserved [3622, 498]. conserved-vector [498]. Consider [2947, 2260]. Considerable [3490]. Considerations [2550, 3219, 2058, 2680, 1263]. Considered [3300, 3971, 3162, 3192, 1259]. considering [1258]. Consistency [253, 3031, 4399, 2199, 3694, 3551, 1593, 252, 167]. Consistent [3418, 4451, 3046, 3467, 4273, 4427, 3694, 4035, 2382, 2065, 3462, 3766, 4210, 2653, 340, 1593, 3446, 1846, 3604, 445]. Consistent-Histories [2065]. Consistently [4242]. Consisting [2145]. Consists [4718]. constancy [336, 335]. Constant [3267, 2682, 2237, 2532, 4435, 2665, 2812, 3577, 4188, 3276, 4379, 4160, 4194, 4317, 2988, 3215, 4087, 2197, 4209, 412, 235, 126, 4331, 554, 4125, 1729, 556, 558, 256, 37, 3104]. Constant/ [3104]. Constantin [1675]. Constants [2457, 4461, 889, 287, 690, 495, 1417]. Constitution [3246]. Constitutive [4249]. Constraint [1799, 4430, 748]. Constraints [4455, 3544, 3739, 2744, 4426, 2211, 2543, 2155, 3724, 2692, 978, 338, 1976, 1764]. Construct [4464]. Constructibility [3945]. Constructing [4167]. Construction [2971, 2892, 2015, 2070, 3513, 4591, 1996, 937, 584, 1998, 1298]. Constructions [2777, 4096]. Constructive [2039, 3975, 926, 583]. constructive-axiomatic [583]. Constructivism [4283]. construed [660]. contact [1515, 214]. containing [1613]. Contemporary [2245, 4503, 2365, 23]. Content [4113, 371, 831]. Contents [2075, 2104, 2117, 2131, 2139, 2149, 2161, 2171, 2179, 2185, 2194, 2204, 2220, 2232, 2246, 2256, 2262, 2268, 2277, 2284, 2293, 2306, 2379, 2368, 2391, 2419, 2430, 2443, 2669, 2678, 2689, 2703, 2709, 2716, 2725, 2735, 2743, 2751, 2759, 2769, 2776, 2787, 2796, 2802, 2809, 2813, 2822, 2829, 2837, 2891]. Context [3661, 3208, 3829, 2280, 11]. Context-Dependent [3661]. Context-Invariant [3829]. Contexts [3888, 4275, 4174]. Contextual [4337, 3028, 4244, 4593, 4348, 1420, 2929, 3829, 980]. contextualism [1727]. Contextuality [4421, 3768, 3239, 4376, 3823, 3243, 3895, 4013, 4440, 3821, 4047, 3434, 3985, 4099, 3972, 4524, 4329]. Contextually [4391]. Continents [2103, 1877, 4623, 2499]. Contingency [3376]. continue [1710]. continued [1733]. Continuing [1675]. Continuity [3253, 4352, 946]. Continuous [2396, 4604, 1113, 2594, 2595, 4696, 1462, 4536, 4569, 2047, 3145, 2504, 3923, 587, 487, 2016, 2048, 1277]. Continuum [3180, 2807, 4696, 4464, 1417]. contract [411]. Contraction [3138, 3903]. Contractions [150]. contradicting [710]. Contrary [484, 3766]. contrast [1087]. Contrasting [3663]. Contribution [2801, 4039, 4462, 2134, 4319, 538, 515]. Contributions [3348, 101, 1287, 1393, 598, 826, 77]. Control [3728]. Controlled [4569, 2848, 1298]. Controlling [2718]. Controversial [372, 784]. controversy [1468]. Conundrum [3405]. convention [921]. Conventional [2364, 586]. Conventionalism [308]. Conventionality [2307, 2491, 108, 3356, 413, 4212]. Conventions [2576]. Convergence [702, 1006, 1216]. Convergent [2468]. Conversations [4448]. Converter [3907]. Convex [4591]. convexity [2030]. Convivial [3913]. Conway [3043, 3739, 3756]. Cooper [3929, 527]. Coordinate [296, 3331, 1940, 2671, 4240, 2536, 2153, 1263, 1264, 388]. Coordinate-free [1263, 1264]. Coordinates [2851, 3090, 1166, 2336, 624, 20, 206, 120, 981]. Copenhagen [3839, 4258, 1721, 1329, 2797]. Copernicus [2344]. Core [2529]. Coreflections [3571]. Cornelius [2897, 144]. Corporation [2495]. corpuscle [1741]. Corpuscular [3662]. Correct [3300, 3256]. Corrected [2899]. Correction [4385, 4444, 4355, 4366, 4413, 4356, 4245]. Corrections [2672, 3851, 2483, 4715, 2557, 832]. Correlata [2737, 2748]. Correlated [3897, 3675, 3354, 1059, 1268, 1444, 1855]. Correlation [4219, 3163, 2710, 3896, 450, 1315, 1027]. Correlations [3337, 2470, 4480, 3544, 109, 2135, 1116, 2435, 4248, 3798, 3850, 3912, 2123, 2227, 2748, 4596, 3696, 3218, 3332, 3020, 4145, 3877, 3626, 615, 1253, 1310, 1069, 1416, 784, 1061]. Correspondence [4483, 3959, 4215, 3566, 3446, 3249, 3808, 4272, 1657, 1514, 642, 97, 1449, 166]. corresponding [815, 2001, 1298]. Corrigenda [2642, 2688, 2591]. Corrigendum [1340, 1351, 1913]. Cosmic [1760, 4357, 2127, 2847, 3705, 1160, 685, 736, 709, 1670]. Cosmogenesis [3506, 3652]. Cosmological [3267, 2682, 4011, 2665, 1774, 1857, 3577, 4188, 4144, 4469, 4509, 4738, 4117, 230, 249, 4379, 3055, 3139, 4160, 3104, 2564, 165, 4194, 100, 3060, 4317, 3160, 2988, 287, 3215, 254, 3201, 4450, 2280, 4209, 966, 4268, 235, 4125, 4648, 4663, 4668, 256, 2677]. cosmologically [169]. Cosmologies [4007, 554, 2464, 838]. Cosmology [2313, 4337, 4, 4004, 4038, 2772, 3418, 3842, 3621, 4509, 4659, 3280, 2995, 4590, 4123, 4005, 2203, 3700, 3926, 3746, 2887, 81, 2973, 4728, 4704, 4122, 4334, 1848, 1696, 678, 1774, 1791, 4675, 4679, 1870, 1659, 643, 1335, 1442, 4628, 1912, 1458, 1670, 2786, 2255]. Cosmos [4670, 3064, 4703, 2819]. cost [1834]. Could [3592]. Coulomb [205, 2159, 3063, 2017, 2140]. Countable [3346]. Counterfactual [3243, 4355, 4327, 2248, 1705, 1704]. Counterfactuals [3552, 3465, 2249, 2238]. Couple [3991]. Coupled [2312, 2980, 1144, 3091, 2657, 3941]. Coupling [3340, 2552, 2458, 4056, 4428, 3253, 2694, 670, 690, 2019, 239, 1176]. Couplings [2568, 3750]. Covariance [258, 4342, 4234, 4173, 119, 4438, 2992, 1123, 553, 836, 866, 1166, 460, 1669]. Covariant [2724, 4318, 2085, 2069, 3394, 2502, 3235, 2925, 2911, 3245, 2507, 2543, 4271, 2913, 3201, 2426, 2427, 2140, 2863, 1455, 766, 1792, 3427, 600, 1086, 267, 286, 659, 146, 1162, 1957, 505]. cover [2967]. Covering [2420]. Covers [3307]. coworkers [979]. CP [2236, 2043]. CP-Violation [2236]. CPT [3327, 2043, 3330, 2791]. Cr [2437]. crack [838]. Created [2598, 1407]. creating [1793]. Creation [2135, 2267, 2564, 3111, 1000, 1005]. credibilities [642]. Criterion [3979, 3766, 1318, 1714]. Critical [4725, 3858, 336, 2680, 3607, 3389, 4630, 4638, 4643, 4656, 2033, 2055, 514, 1617, 1292, 1511]. critical-historical [4630, 4638, 4643, 4656]. Criticism [2951, 352, 351]. Critics [4040, 3625, 3617]. Critique [3889, 3137, 3110, 71, 425]. Croca [2794]. Cross [2287]. Cross-Term [2287]. Crossed [2461]. Crossing [4112, 2093, 3389]. crucial [1726]. Crull [4039, 3872]. Crunch [4143]. Cryptography [3302]. Crystallization [3966]. CSL [2217]. CTC [4408]. Curious [4076, 3844, 1218, 1344]. Current [2237, 3025, 2606, 2424, 1287, 1870]. Currents [4531, 2532, 443, 1056, 498]. Curvature [4056, 3090, 3489, 690, 2914, 4087, 2260, 1034]. curvature-torsion [1034]. Curved [3577, 3058, 3966, 3167, 654, 156]. Curves [3531, 3989, 2919, 4453, 3358, 3194, 3551]. Curvilinear [3150]. Cushing [2867]. Cut [3395, 1457]. cut-off [1457]. Cutoffs [3992]. Cybernetics [3395]. cycle [969]. Cyclic [3895, 3746, 1514]. cyclical [104]. Cylindrical [4045, 480].

D [3049, 2387, 2668, 3935, 838, 3484, 3274, 3449, 2925, 2730, 2100, 4147, 4408, 2653]. D-CTC [4408]. D. [3482]. D=2 [1878]. DABCO [4323]. D’Alembert [1853]. DAMA [3329]. DAMA/LIBRA [3329]. damped [2054]. Daniel [2206, 2231]. Dark [3140, 3263, 3052, 4484, 4738, 3948, 4713, 2995, 4200, 4590, 3104, 2598, 2699, 4051, 4443, 4113, 4373, 2071, 1425, 1659, 1760]. Dark-Matter [4113]. Darwinian [3420]. Darwinism [2938]. Data [2747, 2995, 4185, 1508]. David [2757, 4610, 3010, 2193, 2475, 1058, 1071, 1075, 2736, 3356, 1486, 1047, 2750, 1575]. day [1414]. days [514]. De-Rham [443]. Dead [4366, 4336, 4691]. deal [1336]. Debate [4039, 329, 1414]. debates [4627]. Debating [1701]. Decay [2720, 3590, 2611, 2688, 2698, 3388, 1971, 2024]. Decaying [3681, 3218, 2805, 1991]. Deceleration [4498]. December [2220, 2232, 2787, 2837]. decides [290]. Decision [3431]. Decisions [4560]. Decoding [3437]. Decoherence [2298, 2396, 3182, 3873, 2999, 3224, 3842, 4040, 3427, 3715, 3771, 4039, 4215, 2091, 3849, 3602, 3587, 3330, 2048, 4622, 2216, 3926, 3504, 4589, 1591, 3175, 3119, 3062, 4731, 4070, 3913, 2429, 1970, 2291]. Decoherence-Based [4589]. Decoherent [3628]. Decomposable [2724]. Decomposition [3580, 3372, 3664, 2272, 3515, 3755, 1034, 1844, 1172, 1226, 1233]. Decompositions [2806]. Decouple [4704]. decrease [499]. Dedicated [2477, 2750]. Deduction [2570, 3016]. Deep [4548, 3512, 3981]. Deeper [2422]. Defect [2973, 1543]. Defense [4270]. Defined [3422]. Defining [4101, 4254]. definiteness [1705, 1704]. Definition [3355, 2752, 376, 2954, 3546, 1418, 522, 564, 1413, 892]. Definitions [215, 82]. Deflection [3009]. deformable [118]. Deformation [3532, 4066]. deformations [1523, 2018]. Deformed [3699, 2289, 3824, 3864, 3325, 4271]. degeneracies [535]. Degeneracy [2586]. Degenerate [2416, 2436, 4226, 3064]. Degrees [3794, 3508, 3525, 3005, 3227, 2676]. Delay [4338, 2053]. Delayed [3278, 2358, 2940, 3671, 3885, 1603, 3190, 3290, 4179, 2081, 74]. Delayed-Choice [3278, 2358, 3671, 3190, 2940, 1603, 3290]. Delays [4179]. Deleting [2949]. Delimiting [4075]. Delocalized [2396]. delta [1534]. Demand [3954]. demanded [264]. democracy [80]. Democritean [203]. Demon [3533, 597, 3907, 3405, 4294, 2573, 2628, 1834, 4413]. demonstrated [360]. Demonstration [4223]. Demystifying [4003]. Dense [2274]. densities [648, 1105]. Density [3052, 4144, 2335, 2880, 3816, 3444, 3811, 3008, 2882, 3001, 4390, 4319, 2137, 2634, 1197, 4462, 305, 1159, 1718]. Departures [3825, 2359]. Dependence [4131, 3903, 3874, 4153, 3046, 4429, 316, 1103, 1512, 11, 1965, 690]. Dependent [3661, 3674, 2486, 3117, 4498, 3090, 2142, 2508, 2454, 3257, 451, 803, 3480, 1019]. Depth [4406]. Derivation [2374, 3994, 2486, 1836, 3529, 4541, 4293, 2897, 131, 302, 4445, 3434, 3985, 4273, 4427, 3643, 2869, 4028, 1801, 4488, 405, 235, 941, 276, 1922, 623]. Derivations [3052, 2770]. Derivative [3061, 2930, 98, 267, 326, 334, 1576]. Derived [3575, 2774, 3646, 3476, 3444, 3253, 3226, 71, 306]. Deriving [4308, 3047]. Descartes [1485]. Describable [4257]. Describe [2342, 4242]. Described [3386, 1785, 1298]. Describing [1711, 3914]. Description [525, 3028, 3300, 2289, 2878, 4501, 2554, 2555, 3870, 2170, 4156, 3902, 3336, 4015, 3414, 3212, 4321, 3094, 3866, 3977, 4220, 744, 1427, 1200, 451, 1693, 1685, 604, 1543, 380, 1909, 534, 1997, 586, 608, 1259, 1076, 1133]. Descriptions [4210, 479, 1395, 1409, 665]. Design [3960]. Designed [4006]. d’Espagnat [1593]. Destabilising [4454]. Destruction [3244]. detailed [597]. Details [3817]. detect [1411, 791]. Detectably [3649]. Detected [3030, 389]. Detection [2922, 3254, 2122, 3696, 3184, 4360, 1730, 1936]. Detector [3758, 3059, 1655]. Detector-Based [3758]. Determination [2301, 2124, 13, 31, 1128, 171, 19]. Determined [2995, 4389, 4372]. determiner [1063]. determines [1345]. Determinism [3251, 3283, 3739, 3402, 3804, 4250, 4052, 4070, 3762, 1020, 3974]. Deterministic [4151, 3758, 3839, 4204, 3485, 3130, 4500, 3641, 1076, 1281]. Deutsch [2874, 3079, 3086]. Development [2450, 2558, 3210, 3305, 32, 517, 1472, 1023, 33, 2053]. Developments [2474, 2387]. Deviations [3649]. Device [2124, 1033, 1165]. Devices [3695]. Dewar [4201]. diachronic [1021]. diagonal [1695]. Diagrammatic [4028]. Dialogue [2540, 3637, 1153]. Diathermal [3118]. Dichotomizations [4440].

Dickson [2254]. Did [1869, 2490, 2177, 656]. Diego [2786, 2820]. Diffeomorphisms [4041]. Difference [4057, 3640]. Different [3180, 3794, 3459, 3052, 2211, 2807, 4491, 2988, 3011, 3131, 350]. Differentiable [1939]. Differential [2352, 4177, 3643, 2146, 2280, 1987, 1254, 1829, 1586]. Differentiation [3223, 3439, 3855]. difficulties [342, 1043, 345]. Diffraction [3484, 3100, 2214, 4613, 1577]. Diffractions [4156]. Diffuse [4008]. Diffusion [3922, 365]. diffusions [1297]. Digital [4207]. Dilatation [3750, 177]. Dilation [3849, 106, 242]. Dilaton [2653]. Dilemma [3968]. Dimension [4016, 3038, 4299, 4356, 163, 2039]. Dimensional [3267, 2334, 3763, 2086, 3751, 3839, 3948, 2916, 3017, 2201, 2438, 4087, 4731, 4070, 176, 2027, 1042, 1785, 451, 773, 224, 274, 355, 320, 395, 1227, 1352, 731, 665, 755, 1589, 1104, 1298]. dimensionality [79]. Dimensions [2200, 2479, 2815, 2827, 3341, 4573, 3620, 3029, 3063, 4046, 4585, 4382, 3641, 3032, 460, 243, 3335, 297]. Dipankar [2495]. Dipole [4605, 2472, 4265, 2099, 1638, 1260]. Dirac [3340, 551, 563, 3189, 1539, 3674, 545, 4605, 4454, 2308, 365, 3852, 1557, 547, 802, 3751, 3813, 4287, 1564, 2814, 265, 1583, 301, 325, 2332, 1210, 1333, 4543, 4602, 3280, 1094, 2593, 2167, 746, 2136, 552, 727, 4428, 3521, 1467, 1425, 1760, 1876, 2068, 2267, 2564, 2598, 3111, 408, 3446, 3304, 3960, 1574, 941, 2088, 1502, 3167, 1563, 1740, 550, 3853, 494, 2914, 1520, 558, 4265, 3643, 4446, 2508, 1961, 1423, 1584]. Dirac-Type [3189]. Direct [1980, 2340, 1844, 2060, 523]. Direction [4539, 1308, 183, 217]. directional [4143]. directionalization [1327]. Directions [3475, 2805]. dirichlet [4701]. Disappearance [4540]. Discernibility [3183]. Discontinuous [1962, 126, 1277]. Discovery [2617, 4644, 360, 1406]. Discrepancy [4604]. Discrepant [2197]. Discrete [2646, 3083, 4604, 4609, 3621, 3956, 121, 3797, 2910, 3047, 4536, 4177, 4054, 491, 3915, 3714, 864, 1113, 73, 2043, 1618, 268, 1258, 1201, 664]. Discrete-Time [3047]. Discreteness [2485, 3369]. Discretized [4717]. discretum [1417]. discretum-Unified [1417]. Discrimination [3412]. Discussion [3473, 3075, 3790, 893]. Discussions [959]. Disjunction [2398]. Disk [3996, 2175, 380, 423, 422, 303]. dislocated [947]. Dislocations [3357]. Dismissed [4256]. disorder [967]. Disordered [2470, 2459, 1883, 1884]. Dispelling [2919, 3602]. Dispersion [3754, 3600, 77, 101, 948, 1223, 1067]. dispersion-dissipation [948]. dispersion-free [1223, 1067]. dispersions [241]. Displacement [3025]. disposition [4627]. Dispositions [1727]. Dispute [3175]. Dissipation [2275, 948]. Dissipative [578, 4386, 1685, 969, 971]. Dissolution [4558]. Distance [2898, 1578, 3900, 2995, 3055, 3139, 3713, 4013, 3523, 4420, 3016, 978, 1186]. distances [949]. Distant [4050, 4213, 378, 4248, 735, 108, 413]. Distinction [4178, 4570, 1077]. Distinctness [3008]. Distinguish [3971]. Distinguishability [2314, 3531, 3774]. Distinguishable [8, 596, 595]. Distinguishing [3692]. Distributed [2126]. Distribution [4162, 3149, 3117, 2984, 2411, 1763, 1783, 1448, 843, 536, 1923, 1924, 1601].

Distributions [4093, 109, 258, 1606, 457, 1856, 902, 367]. Disturbance [4165, 4447, 425, 2172]. disturbances [633]. Disturbing [1869]. Divergence [4425, 1006]. diversity [975]. Division [3564, 2620, 4021, 4700]. Do [4592, 710, 1125, 2650, 4725, 4561, 3715, 4404, 4044, 2227, 2763, 411, 1524, 3850, 1302, 285, 103, 763]. Doctrine [2997, 1899]. Dodelson [2786]. Does [4733, 827, 2361, 3644, 4549, 3362, 3667, 685, 3792, 3932, 3285, 1518, 3256, 4140, 3098, 518, 2936, 1302]. Dogmas [4429]. DOI [3322]. Domain [3572, 3782, 2454, 2073]. Domains [4202, 1906]. Dome [3845]. Dongen [3407]. Don’t [3612, 4077]. doom [838]. Doppler [1040, 4241, 660, 339, 432, 404, 880]. Doran [2808]. Dordrecht [2783, 2708, 2609, 2600, 2474, 2497, 2494, 2903, 2785, 2496, 2566, 2429]. Dordrecht/Boston/London [2708, 2496]. Dot [2458]. Dots [4131, 3034]. Double [3484, 4550, 4496, 3024, 2611, 2688, 2698, 3961, 2453, 3184, 1471, 302, 1117, 543, 565, 2024]. Double-Slit [3961, 3184, 3024, 1471, 1117]. double-valued [543, 565]. Doubly [3723, 2014]. Doubt [4627, 2241]. Down [3907, 4069]. Down-Converter [3907]. Downwards [4207]. Dramatically [3586]. Dressed [2336]. Dressing [4551]. drift [732]. drive [975]. Driven [4284, 2290, 1966]. Droplet [4362]. Droplets [4014]. Dual [261, 4517, 1964, 1632, 749]. dualism [1474, 1465]. Dualities [3029, 3891]. Duality [3048, 3863, 3687, 4393, 2485, 3951, 3864, 3614, 3889, 4302, 2909, 4459, 311, 4375, 3641, 1741, 1475, 1900, 521, 2000, 987, 983, 664]. Due [3832, 1557, 2093]. Duffin [3936, 2198, 3620]. Dumpty [1070]. During [3545, 1001, 2253]. Dvurecenskij [2497]. Dynamic [2093, 2781, 2070, 242, 1776]. Dynamical [2286, 3916, 4259, 1194, 2449, 4305, 2711, 3893, 2720, 2326, 3679, 3919, 3378, 2083, 3808, 2049, 4536, 3779, 3562, 3938, 3432, 2965, 4644, 1535, 1276, 1512, 1711, 2022, 1687, 1973, 693, 761, 210, 1754, 186]. Dynamically [2349, 3887]. Dynamics [3702, 4580, 1184, 1192, 2378, 4059, 3575, 2263, 2511, 4280, 2726, 3440, 2461, 4611, 2481, 3339, 2450, 2310, 4699, 2381, 2524, 3394, 2637, 2779, 2245, 3397, 3688, 3749, 4381, 4379, 3645, 3858, 4716, 3038, 4557, 3413, 2437, 3530, 3603, 2527, 3708, 2519, 2522, 474, 2630, 3944, 2557, 3570, 4057, 4361, 2958, 3268, 3930, 3073, 2045, 294, 1157, 4644, 1285, 785, 1620, 1456, 1786, 578, 1052, 2040, 2043, 1598, 1599, 1881, 1997, 700, 24, 415, 1847, 2055, 309, 2908, 1767, 1858, 428, 475, 377, 560, 1528, 1981, 1957, 4636, 908, 816, 424, 768, 2367, 2390, 2494, 2617].

E. [1022]. E.d.B.B. [1801]. E=mc [204]. Each [4491, 3692]. Early [829, 4704, 33, 1575, 513, 1206, 1014]. Earth [2557]. Easy [2315, 4632, 4647]. Echo [2792, 2141]. Economics [3862]. Economics-Based [3862]. Economy [3796]. ed [4448]. Eddington [764, 1961, 1670]. Edge [2755, 4551]. Edited [2600, 2821, 2784, 2283, 2243]. Edition [2090, 2899, 2292]. Editor [3312]. Editorial [1307, 1469, 3452, 3505, 2, 4073, 3292, 3840, 3035, 3115, 3125, 3881, 512, 3056]. Editors [2392]. Eds [3814, 2783, 2474]. Edwin [2784]. Effect [2111, 2358, 3861, 2235, 4330, 2361, 4432, 3214, 3740, 3520, 3306, 3067, 3987, 2309, 3669, 4241, 3461, 3580, 3672, 3683, 4172, 1688, 2412, 2534, 3372, 2465, 3480, 2707, 1975, 2212, 3232, 3373, 4224, 2905, 2507, 3262, 4094, 2781, 4363, 3490, 3734, 2272, 3953, 2546, 4321, 3844, 3562, 4551, 3755, 2842, 2852, 4225, 975, 1040, 2050, 1824, 1219, 1385, 890, 891, 1305, 130, 1438, 2056, 300, 660, 118, 339, 1531, 2049, 404, 327, 486, 2047, 452, 1050, 880]. Effective [2330, 2637, 2701, 2331, 2342, 4584, 4277, 3144, 3185, 4429, 2022, 288]. effective-photon [288]. Effectiveness [4499, 3497, 3426, 1548, 1949]. Effects [4050, 3240, 2641, 3295, 4545, 3101, 2713, 4051, 4111, 3586, 2278, 3712, 3335, 3650, 4687, 1946, 679, 750, 943]. Efficiency [3788, 1116]. Efforts [4704]. Efinger [952]. Ehrenfest [2707, 3472, 4066, 1439]. Eigenstate [3062]. Eigenvibrations [1504]. eight [78, 101]. eight-component [78]. eightieth [529, 545, 502, 1864]. Eikonal [2338, 2730, 2907]. Einstein [3049, 3070, 4184, 3010, 1130, 2566, 2177, 3000, 3867, 4127, 678, 3592, 3929, 4342, 978, 3149, 1032, 3703, 1734, 1252, 3280, 1910, 1369, 405, 480, 1251, 4428, 3281, 3456, 3841, 1051, 2868, 1869, 735, 1472, 804, 3649, 3914, 1827, 3772, 2422, 131, 302, 3185, 959, 457, 114, 1166, 1721, 717, 2883, 2035, 2927, 2729, 2771, 18, 783, 3696, 794, 446, 379, 3971, 2127, 2997, 298, 1076, 236, 933, 4587, 437, 155, 2143, 2197, 1485, 1600, 1801, 4735, 424, 4673, 3147, 784, 898, 524, 1661, 4683, 1899, 2967, 3407]. elaboration [1675]. elastic [480]. Electric [2237, 3072, 2461, 3871, 2505, 2721, 3071, 3472, 1260, 4260, 2852, 661, 341, 48]. electric-type [48]. electrically [1769]. electricity [89]. Electrodynamic [2295, 2804, 2603, 1205, 223, 669, 83]. Electrodynamics [2113, 676, 2538, 2373, 3404, 175, 481, 3301, 3083, 3978, 3214, 2115, 878, 3113, 4083, 2069, 3080, 3085, 3264, 2329, 2622, 2897, 3618, 3135, 2068, 2518, 4614, 4615, 4631, 4433, 2990, 4015, 2159, 2158, 2160, 2718, 3851, 4540, 3623, 1104, 3016, 2544, 2970, 1343, 546, 962, 1633, 538, 2033, 611, 1197, 1720, 606, 610, 605, 616, 865, 320, 581, 1958, 1853, 1613, 1617, 403, 1520, 1540, 1264, 318, 507, 1316, 1202, 306, 476, 2303, 2193]. Electrodynamics-2 [3851]. Electrodynamik [3049]. Electrogravitational [477]. Electromagnetic [2683, 2373, 2105, 2550, 3342, 4698, 4511, 2287, 2998, 3095, 2947, 3396, 3892, 4301, 4585, 4736, 2512, 2100, 4321, 559, 2424, 2859, 4417, 2694, 3941, 1477, 1033, 1040, 1466, 4613, 1108, 1178, 1205, 592, 833, 443, 1910, 192, 399, 459, 126, 199, 682, 1581, 280, 386, 471, 2000, 347, 636, 385, 361, 1328, 1298, 1176]. Electromagnetically [3791, 2097]. Electromagnetism [2362, 2909, 2910, 4030, 3215, 2112, 4159, 272, 1706, 1698, 1925, 1094, 600, 735, 1086, 1310, 289, 1162]. Electron [2848, 3342, 3166, 3605, 2435, 2526, 4708, 1577, 2005, 4182, 3318, 3304, 2201, 2120, 3981, 305, 3812, 4390, 4692, 4707, 2060, 538, 180, 265, 1786, 3034, 1518, 358, 293, 1776, 1133, 1196, 476]. Electronic [2753, 3257, 171]. Electrons [4059, 3320, 4605, 3409, 3444, 3265, 3273, 3352, 1883]. Electrostatic [2106, 323, 397]. Electrostatics [4433]. elegant [301]. Elektrodynamik [3000]. Element [2510, 1974]. elemental [348]. Elementary [4288, 3499, 2005, 2638, 2203, 4445, 1107, 1175, 444, 2321, 2108, 3503, 330, 1765, 1187, 1326, 1456, 1374, 131, 360, 1501, 712, 1417, 466]. Elements [2189, 2188, 2385, 1832, 3176, 2734, 3624, 2249, 1976, 1413, 1896]. Elevator [2922]. Elevator-Like [2922]. Eliminated [3460]. Eliminating [4557, 4307, 4692]. Elimination [2374, 2588]. Elise [3872]. Elliot [2590]. Ellis [4277]. Elsewhere [3926]. Elucidation [522, 564]. Elusive [3345]. Elusiveness [3494]. embeddable [1244]. Embedded [3269]. Embedding [2209, 1693, 2336, 2320, 1096, 584, 1697]. Embeddings [2356, 1095]. Emerge [3458, 3923]. Emergence [3452, 3457, 3028, 4244, 4535, 3453, 3459, 4276, 4227, 2091, 4086, 3456, 4310, 4378, 3029, 2817, 3603, 3378, 4464, 1444, 2487, 3801, 3268, 1998, 1985]. Emergent [3703, 4409, 4600, 3269, 4706]. Emerges [3839, 3443]. Emil [2389]. Emilio [2968, 2987]. Emission [2093, 1580, 1799, 1545]. Emitted [3306, 3893]. Emitting [2711, 2095]. Emmy [3482, 4684]. Empirical [2301, 1996, 2408, 1027, 1828, 4267, 4523, 4196, 960, 2409, 13, 31, 161, 1318, 323, 4510, 19, 1267]. Empiricist [2708, 4652]. Empty [4561, 3007, 2873, 480]. Emulsions [4131]. encountered [525]. Encounters [3137, 3110]. Endeavour [2968]. Energetic [4293, 1549]. Energies [2609, 316, 137, 4649]. Energy [3468, 3251, 3283, 3674, 2105, 391, 4563, 1557, 3140, 3263, 2355, 4398, 821, 3052, 3383, 316, 2287, 2485, 1454, 2513, 2711, 4067, 2814, 2148, 3701, 4484, 3276, 4079, 3561, 4201, 2917, 3275, 3948, 3990, 3051, 2554, 2555, 3034, 4200, 3786, 3311, 3907, 3104, 3892, 2337, 4335, 3791, 4051, 2192, 3957, 4403, 2722, 2902, 4540, 4373, 4319, 4415, 3915, 2667, 3522, 3941, 3016, 1295, 1194, 1195, 1753, 185, 222, 209, 841, 397, 1171, 4182, 382, 131, 733, 33, 4462, 1851, 1923, 485, 800, 1540, 368, 892, 466]. Energy-Dependent [3674]. Energy-Momentum [4319, 841, 4462]. Energy-Time [3957, 1194, 1195]. Engage [3496]. Engine [4284, 1787, 1132, 1131]. Engineered [3741]. Engines [3788]. English [362]. Enhanced [3960]. Enhancement [3559, 1970]. Enigma [3137, 3110, 1977]. Enough [4077, 3225]. Enrich [4491]. Ensemble [3687, 4611, 2441, 3175, 3589]. Ensembles [3689, 3008, 4617, 2870, 2496, 2050, 1252, 1251, 1218]. ensuing [1001]. Entangled [2301, 2222, 3413, 4298, 3977, 4436]. Entanglement [3041, 3468, 3203, 3244, 2571, 3093, 2940, 2746, 3310, 2224, 3351, 4343, 3370, 3417, 4528, 3671, 2226, 3508, 3190, 3290, 2948, 2979, 3559, 3040, 4078, 4307, 4711, 4098, 4530, 3684, 4025, 4420, 2659, 3953, 3332, 2764, 2078, 3236, 3448, 4571, 4474, 4402, 3507, 4702, 3940, 1009]. Entities [2135, 2145, 525, 1683, 1200, 1785]. Entity [2145]. Entropic [3584, 4220]. Entropies [3837, 3432]. Entropy [3761, 2665, 2138, 3120, 3310, 2672, 3754, 2319, 3824, 3681, 2664, 4730, 4438, 2761, 3258, 3922, 2979, 3422, 2666, 4323, 3928, 3122, 2638, 3836, 3414, 2655, 2663, 4701, 3883, 3642, 4416, 3403, 2530, 3546, 2667, 3856, 4374, 1787, 329, 1834, 1973, 2936]. Entropy/ [2936]. Environment [245, 3428]. Environmental [3244]. Environments [2938]. Eon [2772]. Epistemic [4203, 4490, 2857, 3281, 4564, 4481, 4570, 4285, 3494, 4145, 4136, 3551]. Epistemological [2890, 3854, 271, 1897]. epistemology [4654, 4662, 1752, 2783]. EPR [1410, 2358, 4480, 1140, 1250, 1468, 1714, 1930, 3937, 3897, 3190, 3290, 3163, 3476, 1253, 2608, 3675, 3539, 2976, 490, 3054, 1292, 2423, 1809, 2824, 979, 1061]. EPR-B [3937]. EPR-Like [1930]. EPR-paradox [490]. EPR-Type [3190, 3290]. EPRB [3038]. EPRL [3568]. Epstein [3382]. Equation [3876, 3340, 3994, 3674, 4605, 4454, 4012, 3867, 4022, 2486, 3813, 3946, 3703, 4148, 3807, 3095, 2520, 3397, 2136, 4477, 2721, 3620, 3651, 3870, 3521, 2574, 2166, 2525, 3253, 2264, 3960, 3447, 2352, 3445, 3997, 2719, 2907, 3853, 3712, 2142, 3501, 3643, 2508, 4735, 4488, 3291, 3152, 4171, 4245, 3016, 1455, 3736, 1539, 175, 365, 677, 1557, 1792, 1783, 715, 802, 1232, 1564, 1072, 1583, 78, 198, 325, 1210, 1333, 946, 1341, 1775, 1467, 753, 1641, 1800, 40, 941, 1853, 993, 1881, 1502, 1960, 1563, 1740, 347, 636, 421, 38, 1423, 1922, 932]. Equations [4228, 2374, 3189, 3575, 2074, 4014, 4158, 2482, 2612, 3078, 3089, 3133, 3219, 3021, 2085, 2943, 3815, 2588, 2520, 3398, 4428, 3925, 2925, 3649, 3914, 2835, 3445, 3185, 2695, 3167, 2730, 2729, 2771, 2914, 2133, 4479, 2930, 3753, 1359, 567, 714, 1032, 1637, 1775, 765, 480, 1382, 686, 993, 582, 831, 407, 1691, 87, 927, 537, 295, 693, 1354, 149, 1961, 1298, 1584, 1743, 1661, 716, 2760]. Equiareal [4471, 4470]. Equilibration [3119]. Equilibrium [3121, 1763, 2836, 2502, 4716, 2675, 4667, 4676, 4682, 4575, 1178, 1533, 965, 233, 56, 400, 3901, 453, 1973, 949, 2836]. Equivalence [2679, 4454, 4469, 4342, 3032, 2844, 4224, 4182, 3595, 2572, 2547, 2348, 4282, 3760, 92, 823, 888, 1621, 185, 1836, 119, 1124, 1611, 257, 1669, 562, 4232, 2591]. Equivalent [4014, 2079, 3397, 3398, 856, 4080, 3016, 14, 8]. Erased [3033]. Eraser [3885, 4179, 2081]. Erases [4179]. Erasure [3190, 3290, 3033]. Ergodic [3269]. Eriksen [3960]. Errata [901, 1995]. Erratum [132, 208, 449, 509, 571, 798, 799, 810, 811, 812, 886, 990, 2129, 2130, 2568, 3625, 2768, 94, 3305, 501, 3352]. Error [414, 975]. Errors [2561]. Ertel [1912, 2034]. Erwin [1004, 1008, 1024, 1001, 1000, 1005, 959, 1014, 1021]. Escape [2966]. Escuela [2377]. Esfeld [4039]. esoteric [4627]. Esposito [2102]. ESR [3424, 3654]. Essay [4343, 362]. Essence [3106, 3203]. Essential [2859, 4664, 4669, 442, 2820]. Establish [4061]. Estimated [3788]. estimates [829]. estimation [135, 1240]. Eternal [4691, 1160]. Eternalism [4211, 4290]. Ether [2855, 36, 732, 83, 186]. ether-drift [732]. Etherino [3065]. Ettore [2801, 3080, 3085]. Euclidean [4620, 2723, 4134, 665, 1985, 2102]. Eugene [2245, 2292, 544, 529]. Euler [3703]. EUR [2600]. eV [2341]. evade [847]. Evading [4719]. Evaluating [1637, 2331]. Evan [2345]. Evans [3126, 3127, 3128, 3265]. Evaporating [4157]. Evasion [4131]. Event [4733, 4437, 3662, 2811, 3670, 3147, 1783, 1421]. event-anti-event [1783]. Event-by-Event [3147]. Event-Time [4437]. Events [3682, 3726, 4453, 4434, 3558, 3181, 609, 1271, 1733, 1952, 1763, 1759, 1958, 188, 1154, 420]. Everett [4255, 2147, 344, 2884, 4516, 2401, 3416, 2626, 2670, 3005, 3466, 1049]. Everettian [4548, 4090, 4142]. Evert [2496]. Everything [4238, 1505].

Evidence [547, 803, 4528, 2611, 2688, 2698, 1188, 3562, 527, 1986, 1176]. Evolution [2848, 2918, 4287, 573, 3394, 2760, 2761, 2798, 3710, 4709, 4694, 3530, 3928, 3836, 4445, 2907, 2982, 3164, 270, 294, 1656, 1713, 1418, 511, 1364, 557, 512, 270]. Evolution-parameter [573]. evolutional [1042]. Evolutionary [975, 3494, 1591]. evolutions [1514, 2008, 935]. Evolve [2763]. Evolving [3615, 3479, 949]. Exact [4011, 2385, 3620, 3135, 3914, 2132, 3623, 1032, 845]. exactly [1541]. exactness [1909]. Example [3093, 2938, 1689, 1501]. examples [1090, 710, 310, 1276, 360, 2004]. Exchange [4328, 2586, 940, 505]. exchangeable [1606]. Excitation [3797, 1530]. Excitations [2108]. exciton [934]. exciton-phonon [934]. Exclusion [3316, 3320, 3409, 3576, 3331, 3678, 2706, 704]. Exclusive [3558, 211]. Exclusively [3390]. exhaustive [211]. Exhibit [2299]. Exhibits [4605, 30]. Exist [2361, 4725, 2991, 4044, 3972, 3346, 685, 103, 1809]. Existence [2408, 4383, 863, 1397, 2657, 992, 2039, 1906, 1983, 1322]. Exorcising [3405]. Exotic [2800, 911]. Exotica [4473]. expanded [1477]. Expanding [2655, 1504, 694, 381]. Expands [4460]. Expansion [3032, 1532]. expansions [743]. Expectation [2466, 963, 705]. Expectations [2968, 468]. Expected [2072, 4119]. Experiment [2358, 2922, 2937, 3737, 3306, 3409, 3173, 3100, 4496, 2916, 3937, 4482, 3290, 3038, 2214, 3261, 3556, 3818, 4304, 2296, 3009, 4366, 4336, 2614, 2589, 3184, 3098, 3162, 4001, 3192, 1733, 1471, 1658, 1761, 772, 574, 890, 1900, 399, 1051, 1472, 732, 471, 432, 457, 987, 1050, 7, 155, 1809, 899, 1201, 2831]. Experimental [3316, 3320, 2964, 3409, 3893, 3208, 3811, 3265, 3889, 2296, 3821, 501, 1062, 2851, 1119, 2680, 288, 3820, 2516, 452, 3706, 3562, 2603, 396, 1986]. experimentation [1605]. Experiments [4542, 3278, 4014, 3157, 2896, 3592, 3728, 4562, 3671, 3190, 3290, 2236, 2215, 3655, 2118, 4586, 2141, 4298, 2859, 3206, 3421, 3147, 710, 1317, 1383, 1315, 1603, 274, 1117, 303, 1487, 791, 783, 1867, 1076, 74, 437, 1189]. Explain [4559, 3724]. Explained [4484, 3885]. Explaining [3012, 3877]. Explanation [4421, 4308, 2235, 2551, 4065, 3478, 4214, 3476, 2605, 473, 1253, 418, 6]. Explanations [4309]. Explanatory [4406]. explicate [67]. Explicit [710, 971, 1298, 1790, 910]. Exploiting [3624]. Exponential [3072, 4286, 3590]. exponentially [1991]. Exponents [2563]. Exposed [3718]. Expressed [2816, 4350]. Expression [2587, 3135, 3623, 3016]. Expressions [2188, 1367]. Extended [2918, 2503, 2672, 2901, 3583, 163, 2276, 3444, 214, 3782, 2630, 2146, 2321, 4001, 4524, 1106, 1592, 265, 1187, 1326, 1456, 1987, 773, 698, 728, 1764, 445, 1439, 1829, 1202]. Extendibility [4098]. Extending [3139, 4314, 2172]. Extensible [2356]. Extension [3200, 3039, 3494, 607, 435, 1466, 269, 319]. Extensional [3498]. Extensions [3660, 3970, 3680, 1922]. extensive [892]. External [4454, 3361, 3960, 2120, 2321, 3474, 1151]. Extra [3032, 4344, 3335, 4196].

Extra-dimensions [3032, 3335]. Extra-Empirical [4196]. Extra-Structures [4344]. Extracting [4104]. Extravaganza [3416]. Extremal [4167, 2667]. Extremality [4170]. Extreme [2661, 3235, 3586, 1171]. extremely [412]. Extremum [2214]. Extrinsic [4249]. Eye [2516].

F [2766, 3814]. F. [1639, 1153]. Fabric [4670, 2819]. Fabrication [3744]. Fabrice [2836]. Face [3904]. Faces [4429]. Fact [4574, 50]. Fact-nets [4574]. Factorable [2323]. Factorizability [2685, 1924]. Factorization [2062, 677, 1588]. Factorizations [3021]. Factors [3211, 341]. Facts [2889, 4345, 4489, 4725, 3109, 4515]. factually [1451]. fails [1966]. Failure [3802, 2754, 2840, 368]. failures [791]. Faith [2968]. Fall [2213, 2886, 2742]. fallacious [1296]. Fallacy [4696]. Falling [3520]. False [4393]. Falsify [4034]. Familiar [2653]. Families [1018]. Family [3471, 664]. fancy [50]. Far [2648, 965]. Faraday [3067]. Farkas [615]. Fast [4378]. Faster [528, 1033, 1040, 389, 527, 41]. Faster-than-light [528, 389, 527]. Fatal [3434, 3985]. Fate [4458, 2772, 3935]. Fe [2437]. feasibility [1789]. feasible [890]. Feature [4713]. Features [3212, 2653, 604, 575]. February [2104, 2117, 2131, 2246, 2669, 2678, 2689, 2703, 2709, 2796]. feedback [1488]. Fellow [4312]. Femtosecond [3744]. Fermat [2352]. Fermi [4187, 3197, 4295, 2828, 3765, 733, 4736, 4691, 446, 426, 239]. Fermi-Gas [3765]. Fermion [2620, 2332, 3260, 4446, 3084, 3107, 1515]. Fermionic [3939, 2945, 1702]. Fermions [2335, 1804, 2148, 3325, 3782, 3202, 1770, 604, 2041]. Ferrero [2474]. Ferretti [426]. Ferris [2624]. Ferromagnetism [2433]. Ferromagnets [2432]. Festschrift [3347]. few [4627, 826]. Feynman [2164, 2147, 3214, 3602, 1725, 598, 4403, 1357, 3344, 406, 768]. Fiber [3966, 4591, 1456]. Fictitious [3439]. Fidelity [3410]. Field [2113, 2090, 2208, 3189, 2237, 3575, 2511, 4262, 2714, 2330, 3847, 4535, 714, 715, 802, 878, 936, 951, 1198, 4287, 2085, 3839, 3599, 2623, 3132, 4364, 4478, 2845, 2637, 2701, 4296, 3590, 4381, 2767, 2326, 3071, 2436, 3620, 2947, 3127, 3128, 968, 3525, 2971, 2213, 3311, 2342, 2925, 3265, 2905, 4112, 2696, 4301, 3413, 3557, 3675, 3726, 748, 3029, 3377, 2275, 3262, 4094, 3059, 4277, 2728, 3319, 3185, 4367, 731, 4389, 2156, 3532, 3820, 2773, 2791, 2569, 3472, 4429, 2626, 3466, 2099, 4321, 2110, 3289, 3201, 4326, 3988, 4520, 2372, 3935, 2930, 4582, 2350, 2782, 3753]. Field [2108, 3448, 2603, 4060, 2657, 2440, 3474, 3269, 3641, 854, 567, 1769, 1466, 1125, 176, 1946, 1551, 840, 679, 2009, 162, 213, 1032, 263, 1734, 606, 610, 605, 616, 670, 1632, 480, 1527, 1682, 1989, 841, 1421, 856, 255, 410, 1641, 1581, 736, 1500, 1574, 1598, 1599, 1576, 47, 128, 184, 291, 407, 461, 2055, 1691, 1687, 347, 636, 445, 704, 2026, 1585, 626, 927, 537, 1546, 665, 168, 2023, 755, 756, 819, 295, 48, 409, 1354, 149, 1528, 343, 757, 1417, 1328, 420, 1511, 4665, 1660, 1646, 1751]. Fields [2375, 3593, 2799, 4011, 4454, 3921, 2461, 2222, 4535, 4543, 3095, 2091, 3900, 2162, 2106, 3254, 3336, 3960, 3447, 3523, 2463, 3091, 2908, 3074, 2512, 4298, 4446, 3202, 3793, 2692, 2108, 3787, 4260, 2694, 4551, 2492, 310, 1939, 1072, 1384, 1702, 443, 479, 661, 1706, 1698, 1523, 1802, 222, 645, 765, 313, 2040, 575, 1151, 1846, 749, 488, 429, 456, 18, 2041, 1056, 1176, 1647]. Fierz [1588, 1938, 1661]. Fifth [2431, 2444, 2456, 2478, 3339, 675, 163, 1499, 1944, 1309]. Filed [2915]. filtering [1284]. Filters [29]. Final [3611, 1701, 1974]. Finch [902]. find [1986]. Fine [4421, 2812, 4197, 3548, 4720, 4346, 37, 1251, 2423]. Fine-Tuning [4421, 4197, 3548, 4720, 4346]. Finely [4406]. Finetuned [4161]. fingered [4318]. Finite [2789, 3244, 3473, 2815, 4717, 2335, 2656, 3006, 3756, 2991, 3525, 3396, 4405, 2892, 4044, 3023, 4170, 2959, 3668, 3755, 3171, 2027, 1785, 1134, 1606, 1909, 731, 537, 243]. finite-dimensional [2027, 1785, 731]. Finite-Time [3244]. Finitely [1261]. Finland [1138, 1147, 1185, 1193, 1608]. Finsler [567, 2714, 3650, 1104, 1998]. Finslerian [435, 3537, 4432]. Firewall [4043]. Firewalls [3883]. First [1363, 1371, 1378, 3745, 2941, 3898, 3317, 4149, 2698, 2899, 3479, 4362, 2426, 1455, 1878, 892]. first- [1878]. First-Order [3898, 1455]. Fishbowl [4603]. Fisher [4544, 1341, 2279, 3870, 2638]. Fisica [2377]. Fits [2995]. FitzGerald [3138]. Five [3267, 176, 4412, 3825, 1710, 1352, 297]. Five-Dimensional [3267, 176, 1352]. Fixed [2749, 3969, 146]. FLASH [526]. Flat [2923, 2302, 2514, 2729, 3524, 1525, 355, 1981]. flatness [758]. Flavor [3656]. Flawed [3392]. Flea [3630]. Flow [3534, 4317, 3570, 2874, 4084, 400, 1973]. Flows [3311, 4121]. FLRW [4240, 3935]. fluctuate [827]. fluctuation [575]. Fluctuations [4012, 3916, 2481, 3118, 4381, 4379, 2046, 4378, 3396, 4584, 4243, 3591, 2654, 4209, 2440, 3941, 1536, 2022, 447]. Fluid [4738, 3922, 4583, 3479, 4319, 1881, 4462, 2908, 4627]. fluids [1156]. Flux [3713, 2451, 2975, 4396, 2832, 2842]. Foam [2604, 3330, 3332]. Foaminess [2357]. Fock [3929, 2023, 2705, 1457]. Focusing [3177]. Fokker [3702, 1775, 1421]. Fold [2562]. Foliation [2514]. Foliations [4072]. Follow [2386, 4527]. Following [2801]. Follows [3547]. forays [4627]. Forbidden [3576]. Force [4494, 4325, 3932, 2218, 3954, 4520, 3159, 829, 1032, 1537, 774, 3058, 244, 2045]. Force-Free [2218, 4520]. Forces [4394, 3517, 3439, 4328, 2120, 2472, 1848, 30, 285, 1309, 1801]. Forcing [3707, 2777]. Forecast [2772]. forever [1875]. Foreword [3492, 2150]. Forewords [4610]. Forget [3481]. Forgotten [2967, 4673]. Form [3876, 3540, 4602, 3900, 865, 3772, 3986, 3479, 2302, 2426, 2427, 3736, 301, 349, 104]. Formal [4090, 2684, 90, 1146, 1263]. Formalism [4526, 2816, 3816, 4606, 2593, 4391, 4200, 2066, 4481, 3960, 3076, 2671, 3708, 3677, 2281, 4582, 2137, 2064, 1336, 3299, 3941, 573, 946, 1123, 633, 645, 1109, 748, 453, 1303, 2030, 1372, 482, 1680, 1344, 1334]. Formalisms [2892]. Formalist [4726]. formalization [244]. Formalized [2783, 4654, 4662]. Formation [4407]. Forms [3672, 560, 1587]. Formula [3909, 2012]. Formulae [2468]. Formulas [2483, 2462, 290]. Formulated [3449, 4389]. Formulation [4016, 3689, 3025, 2069, 4543, 4064, 2593, 3870, 2166, 3104, 501, 2765, 2527, 3677, 2155, 3848, 2428, 2424, 3787, 3079, 3086, 2260, 1633, 663, 1198, 1072, 325, 1210, 1333, 468, 192, 580, 408, 286, 1476, 396, 136, 1510, 1091, 2018, 933, 1657]. Formulations [3108, 3692, 1878, 1521, 216]. Fortuitousness [2490]. Forty [3606]. Foulis [618]. Found [3113, 3032, 2688, 3206]. Foundation [4330, 2702, 2288, 4723, 2989, 4036, 4434, 4581, 3975, 3778, 1533, 56, 1365, 1373, 125, 626]. Foundational [3728, 3123, 3659, 1442, 3629, 2230, 1617]. Foundations [1138, 1147, 1185, 1193, 1498, 1615, 1990, 2582, 4630, 4638, 4643, 4656, 4095, 2407, 3348, 3511, 3978, 3454, 4004, 3769, 3402, 3186, 2600, 2845, 28, 4306, 2495, 577, 2003, 3556, 4017, 113, 127, 250, 501, 2183, 2803, 3868, 2773, 2791, 3419, 2311, 4649, 4390, 3114, 4430, 4136, 3164, 4374, 141, 3606, 4652, 2970, 3609, 3719, 1608, 1640, 430, 330, 218, 1031, 1522, 579, 4621, 926, 396, 24, 846, 271, 4685, 4686, 324, 1381, 2368, 2391, 2419, 2430, 2443, 2669, 2678, 2689, 2703, 2709, 2716, 2725, 2735, 2743, 2751, 2759, 2769, 2776, 2787, 2796, 2802]. Foundations [2809, 2813, 2822, 2829, 2837, 2866, 2875, 2075, 2104, 2117, 2131, 2139, 2149, 2161, 2171, 2179, 2185, 2194, 2204, 2220, 2232, 2246, 2256, 2262, 2268, 2277, 2284, 2293, 2306, 2346, 2379, 2403, 2455, 2467, 2476, 2488, 2500, 2509, 2517, 2523, 2528, 2533, 2542, 2549, 2559, 2567, 2575, 2583, 2592, 2601, 2610, 2618, 2625, 2632, 2643, 2652, 2660, 2708, 2609, 2631]. Founding [2994]. Four [3763, 2815, 2827, 3271, 1210, 2916, 1367, 3017, 3791, 2738, 2750, 23, 1042, 1333, 224, 274, 355, 320, 395]. Four-Dimensional [2916, 3017, 1042, 224, 274, 355, 320, 395]. Four-Level [3791]. Four-space [1210, 1333]. Fourier [821]. Fourth [3061, 893]. FPR [1230]. Fractal [3383]. Fractal-Scale [3383]. Fractional [3919, 4585, 3844, 4299, 4356]. Fractional-Dimension [4299, 4356]. Fragmentalism [4518]. Fragmentation [2472]. Fragments [3218]. Frame [4526, 4131, 3969, 4260, 1758, 1219, 599, 869, 158, 480, 376, 81, 507]. Frames [3663, 3271, 4058, 3002, 2821, 2175, 2546, 3524, 3026, 102, 1810, 1823, 119, 4689, 452, 2821]. Framework [2816, 4178, 4205, 4293, 4391, 3897, 3150, 3764, 4574, 4365, 1042, 1276, 320, 395, 58, 178]. Framework/Interaction [4178]. Franco [2243, 2838]. Franklin [2733, 528]. Frauchiger [4351, 4360, 4203, 4242]. Frauchiger-Renner [4351, 4360]. Fraunhofer [4156]. Free [4131, 4050, 3203, 4262, 3847, 4535, 4284, 4419, 4444, 4154, 3444, 2213, 3265, 4247, 4273, 2122, 3532, 2218, 2840, 2360, 4520, 4719, 3044, 2674, 4360, 180, 1555, 1698, 1523, 1369, 1936, 1846, 1598, 1223, 863, 18, 346, 1263, 1264, 38, 1067, 3015, 3112]. Freedom [3794, 3508, 3525, 3005, 3227, 3079, 3086, 2676]. French [3080, 3085]. frequencies [412]. Frequency [3306, 3884, 2608, 3353, 2670, 1853, 1091, 489]. frequent [784]. Frequently [3688]. Fresnel [4156, 1063]. Freudenthal [27]. FRG [1184, 1192]. Friction [4494]. Friederich [4346]. Friedman [4121]. Friedrichs [2350]. Friend [4312, 4404, 4495, 4001, 2944, 4401]. Fritz [1610]. Frolov [2387]. Front [2095, 1620]. Frontier [2344]. Frontiers [1570]. Fronts [2508]. FRW [3027]. fuels [4649, 2609]. Fujita [2244, 2785]. Fulfilled [4408]. Fullerenes [2613]. fully [1423]. Function [3189, 2147, 3945, 4604, 4485, 4219, 3721, 2339, 3811, 4696, 3792, 3262, 2118, 3803, 3504, 2984, 4366, 4336, 3943, 4089, 2363, 3515, 950, 176, 1345, 246, 1203, 1009, 1760, 1419, 536, 1085, 1439, 1993, 1657]. Functional [2276, 3444, 2962, 4121, 4171, 4245, 1232, 949, 1354]. functionals [705]. Functioning [3436]. Functions [2479, 2433, 3862, 4046, 2418, 3237, 175, 481, 1437, 1310, 1027, 598, 324, 1601, 8, 1586, 1534, 1068, 607]. Functors [4577]. Fundamental [3876, 1842, 1849, 2387, 2074, 4178, 2553, 2825, 3428, 3018, 2474, 2947, 3451, 2494, 3111, 3826, 4120, 3009, 3097, 2821, 4216, 3964, 4339, 4736, 2785, 4199, 4429, 1209, 4710, 138, 2566, 1675, 1570, 1863, 1903, 570, 989, 342, 1157, 1886, 247, 889, 2043, 709, 1687, 147, 202, 219, 418, 178]. Furry [1930]. Further [319, 4414, 3070, 379, 3817, 517, 401]. Fused [4464]. Future [4609, 3436, 3496, 2624, 1128, 4688, 146, 1055]. Fuzziness [3205]. Fuzzy [3233, 2416, 1105, 2410, 2418, 96, 2400, 2976, 1979, 668, 1444, 188].

G [760, 2784, 2836, 742, 2821]. G. [613, 614, 1153, 612]. Galilean [4538, 3255, 3646, 2773, 3226, 3963, 2863]. Galilean-Covariant [2863]. Galilei [1987, 540]. Galileo [2063]. Gallery [2388]. Galois [4165]. Galoisian [3773]. Gambling [4016]. Game [3431, 2638]. Gamma [2355]. Gamma-Ray [2355]. Gamow [1991, 3790]. Gao [4089]. Garg [3526]. Gas [2310, 4148, 3765, 3116, 4715, 2383, 117]. Gases [3117, 2274, 115]. Gates [3231]. Gateway [4228]. Gauge [3648, 3784, 2374, 4322, 2693, 3775, 3113, 1702, 1770, 3951, 2637, 3614, 2326, 1976, 2900, 2162, 4433, 2990, 3253, 3357, 4363, 3750, 2719, 3091, 2087, 3698, 2153, 3202, 3487, 3935, 2328, 2327, 3787, 3079, 3086, 2290, 2440, 1199, 1758, 936, 1803, 1939, 1384, 1385, 3748, 1636, 918, 1611, 1158, 255, 600, 1086, 1760, 1314, 846, 1963, 537, 409, 1354, 1528, 343, 757, 1647, 1477, 4385]. gauge-covariant [600, 1086]. Gauge-Dilatation [3750]. Gauge-Independence [2328]. Gauge-Invariant [2374]. Gauge-Underdetermination [4363]. Gauge/Gravity [3951]. Gauges [2849, 968]. Gaunt [1961]. Gauss [4433, 3172, 269, 319, 2327]. Gaussian [4611, 313, 1453, 2982]. GBD [4553]. GBP [2600]. Ge [1799]. gedanken [1033, 1809, 1201, 2922, 2781]. Gedankenexperiment [4258]. General [2894, 4072, 3014, 2407, 4432, 2513, 2711, 4067, 4342, 2860, 4234, 4266, 4231, 3051, 1276, 2649, 3055, 3139, 3294, 355, 153, 4023, 4024, 4253, 4216, 4521, 3146, 3288, 866, 2677, 4352, 2924, 3677, 2152, 4080, 4424, 2748, 3256, 2597, 3837, 2512, 2153, 2515, 2426, 2427, 2073, 190, 435, 570, 31, 1454, 679, 696, 746, 841, 759, 855, 1374, 22, 349, 736, 61, 836, 212, 583, 19, 555, 1923, 54, 1669, 1951, 4648, 4663, 4668, 1107, 1175, 1320, 4627, 308, 444, 1091, 1858, 375, 239, 343, 757, 839, 524, 472, 1457, 1899]. General-Relativistic [4234, 2426, 2427, 841, 343, 757]. Generalised [4100, 4253, 4382]. Generality [3673]. Generalization [2038, 2876, 1299, 2700, 315, 3875, 3266, 2926, 3159, 1520, 1999, 1718]. generalizations [852]. Generalized [294, 3674, 4604, 1792, 2790, 2815, 2827, 3229, 3754, 2770, 3813, 3527, 2648, 3672, 78, 2198, 918, 94, 64, 3211, 3424, 85, 255, 3559, 2352, 3603, 1881, 2577, 697, 4068, 3767, 3808, 4729, 309, 3490, 3091, 1888, 474, 4066, 490, 4431, 3933, 3963, 1245, 1557, 438, 592, 624, 1714, 1094, 20, 206, 1240, 1223, 1510, 322, 428, 475, 364, 421, 2648]. Generally [3181]. Generate [3586]. Generated [2349]. generating [1033, 1017, 1358]. Generation [3427, 2270, 1500, 3828]. Generations [3381]. generators [1698, 494]. Generic [2297]. Generically [4408]. Genius [2388, 2699]. Gennaro [2631]. Geno [2734]. Geno- [2734]. Genuine [4537, 2490]. Geodesic [3534]. geodesics [1380]. Geometric [2901, 2414, 4660, 4680, 3039, 2900, 3017, 267, 326, 334, 3385, 1352, 3708, 2908, 4036, 3201, 3890, 2634, 4084, 3516, 4375, 1945, 1379, 1426, 1578, 1571, 1876, 1827, 61, 159, 3286, 1617, 626, 1579, 1423, 2808]. Geometrical [3784, 2704, 3219, 2465, 733, 3023, 4731, 4361, 3156, 704, 1172, 236, 1501, 1300]. Geometries [2356, 2807, 3074, 2653]. Geometrization [698, 728, 1422, 1423, 1316]. Geometrized [4352]. geometrizes [1875]. Geometrizing [3391]. Geometro [1326, 1297, 1829, 2146, 1987]. Geometro-Differential [2146, 1829, 1987]. Geometro-stochastic [1326, 1297]. Geometrodifferential [2276]. Geometrodynamic [4340, 831]. geometrodynamics [836]. geometrostatic [541]. Geometry [4059, 3272, 2714, 3634, 4104, 2620, 3978, 3052, 2537, 4428, 3830, 947, 4134, 2267, 2598, 2933, 4608, 2255, 3399, 3144, 3185, 3615, 3326, 3395, 2627, 4066, 4202, 3643, 3388, 4523, 2384, 816, 2959, 1503, 1731, 567, 234, 1550, 1426, 1578, 967, 1053, 1708, 1425, 533, 1405, 463, 4628, 539, 2002, 1647]. geon [833]. Geophysical [2997]. Georg [362]. George [2836, 2820, 2241]. Georges [4123]. Gerald [2565]. Gerlach [1733, 3264, 3386, 964, 2124]. Germany [2899]. Geroch [3522]. get [1545]. Ghatak [2903]. Ghins [2409]. Ghirardi [3112]. Ghost [3061, 4446, 1706]. Ghosts [2327]. GHZ [1907, 2223, 4599]. Giampiero [2102]. Giant [2441]. Gibbs [2183]. Giora [2177]. Giraffe [3259]. Giuseppe [2102]. given [648, 902]. giving [198]. Glass [4206]. Gleason [4016, 1281, 3230, 2770, 2414, 3947, 1282, 3096, 4171, 4245]. Gleason-Type [4016, 2770, 3947, 4171, 4245]. Glimmers [3286]. Global [1946, 1525, 2704, 4739, 4596, 10, 1533, 1953]. Glueball [2169]. Gluons [2637]. GNS [3566]. GNS-Correspondence [3566]. Go [4734, 1342, 4418, 4412, 4229, 3904, 3092]. goal [710]. god [1407, 1875, 464]. Gödel [2967, 2177, 1869, 199, 717, 4673]. Godoy [2785, 2244]. Goes [2252]. Going [2261]. Golden [4295]. Good [3998, 1283]. Gordon [4531, 714, 2588, 3651, 2574, 1800, 2275, 1700, 1298, 3291]. Gottfried [3152]. Governed [3702]. Grad [4063]. Graded [1602]. graining [262]. Grains [2441, 3723, 2487]. Gran [3329]. grand [122]. Graph [4084, 1968]. graphic [1052]. Graphical [3886, 3903]. Grassmann [994]. Gravitating [2662, 952, 3604, 3479]. Gravitation [2313, 851, 3072, 951, 3839, 4498, 143, 4352, 2923, 289, 3070, 2302, 2514, 3215, 499, 2557, 239, 1321, 2676, 272, 2031, 679, 1456, 1925, 670, 121, 600, 1086, 61, 100, 194, 818, 697, 1854, 1995, 18, 1767, 1163, 168, 232, 210]. Gravitation* [2927]. Gravitational [3106, 2208, 743, 3189, 567, 1769, 4011, 2511, 4109, 4127, 2650, 4469, 3067, 2085, 3276, 4079, 1032, 4528, 4201, 3058, 4709, 3071, 855, 3077, 3649, 120, 257, 4368, 2728, 2655, 242, 750, 2354, 4736, 4338, 3091, 3883, 3712, 3335, 2573, 694, 4361, 1732, 2603, 3159, 2810, 2260, 854, 1125, 678, 443, 661, 263, 670, 1001, 410, 575, 362, 749, 554, 1585, 558, 800, 756, 295, 48, 899, 1176]. Gravitationally [3520]. Gravitationally-Induced [3520]. Gravitator [2383]. Gravities [4051]. Gravitino [2697]. Gravitism [249]. Gravitization [3727, 3746]. gravitomagnetic [840]. Gravitomagnetism [2683, 4147]. Gravitons [3160, 3030, 695, 1217]. Gravity [4458, 4542, 3198, 3197, 2843, 3486, 3537, 3718, 4563, 3873, 3690, 4127, 4340, 3216, 3597, 2896, 3151, 3328, 2864, 3577, 4144, 4469, 3842, 3951, 3789, 3721, 4061, 4201, 4128, 4738, 3406, 4460, 2587, 3613, 4149, 3509, 3710, 4195, 4573, 3274, 3449, 2900, 2213, 2413, 3902, 4078, 3720, 4114, 3009, 3875, 2777, 4150, 4091, 3584, 4553, 3270, 2353, 3535, 3489, 4152, 2102, 2357, 3666, 3700, 3926, 2907, 2157, 1851, 3215, 3161, 2515, 4704, 3890, 4226, 4373, 4365, 3388, 2654, 4430, 4466, 4554, 4299, 4356, 3693, 2492, 2653, 3171, 3553, 2635, 1635, 1878, 574, 1507, 1112, 4620, 1158, 1846]. gravity [686, 1634, 1953, 1516, 1697, 1843, 4125, 1313, 1442, 1951, 489, 1661, 4448]. Gravity-Related [3721]. Gravity-Wave [2357]. greatest [514]. Green [1310, 2281]. Greenberger [2206, 2231, 2223, 1299, 4348, 1604, 4495, 4431, 2229]. Greene [2819]. Griffiths [2475]. Groblicki [902]. Groenewold [3315]. Grøn [422]. Grothendieck [3003, 2811]. Ground [2222, 2093, 3542, 3851, 4576]. grounding [348]. Group [4022, 4340, 2901, 2537, 2545, 1527, 1536, 3076, 3002, 4273, 4427, 4121, 4180, 458, 1033, 1040, 175, 481, 553, 1959, 1506, 409, 228]. Group-theoretic [228]. Group-Theoretical [3076]. Groups [2790, 2815, 3078, 3089, 3133, 3219, 3802, 2212, 3006, 3235, 1754, 2011, 1535, 172, 314, 1524, 1522, 634, 150, 535, 543, 565, 1017, 1657]. Grover [3983]. GRT [2896, 3151]. Gruber [93]. Grünbaum [423]. GRW [4506, 3043, 4358, 3923]. guide [4625, 4633, 318, 2315]. Günther [576, 585, 577]. Gutzwiller [2431, 2444, 2456, 2478, 2477, 2433, 2480]. Györgyi [591]. Gyrogravito [3812]. gyrogroup [2002]. Gyrogroups [2537, 1929]. Gyromagnetic [3812, 745]. gyroscope [1091]. Gyrovector [2537, 2578].

H [2090, 1238, 1248, 1249, 2344, 2766, 1584, 3318, 3542]. H-Atom [3542]. H.-P [2766]. Haag [4395, 3988]. Hadamard [2479]. Hadron [2337, 900, 912, 922, 80]. Hadronic [2609, 3335, 1783, 2046, 1530, 1992, 4649]. hadrons [45]. Hahn [1246, 1172, 1226]. Hair [2657]. Half [2192, 3868]. Half-Integral [3868]. Half-Quanta [2192]. Hall [2475, 1385, 3844, 4551]. Haller [2325]. Hamilton [1222, 568, 179, 441, 569, 2826, 2482, 2612, 2079, 3397, 3521, 4557, 4481, 3225, 1334]. Hamiltonian [823, 888, 294, 3301, 4101, 3689, 1620, 2814, 3962, 654, 1685, 946, 1123, 696, 4160, 3400, 3530, 496, 4491, 136, 1510, 3212, 4080, 4184, 4424, 2464, 3938, 2260]. Hamiltonians [2038, 1526]. Hamming [4499]. Handles [3914]. Hans [2376, 2820, 1590, 1912]. Hans-Jürgen [2376]. Happens [2132]. Hard [2135, 4063, 2967]. Hard-Sphere [4063]. Hardcover [3010, 3011, 2344, 2345, 2375, 2376, 2590, 2641, 2786, 2783, 2708, 2609, 2600, 2795, 2784, 2836, 2499, 2767, 2495, 2497, 2541, 2494, 2565, 2631, 2558, 2903, 2539, 2677, 2785, 2540, 2808, 2794, 2820, 2496, 2766, 2566, 2315, 2624, 2819, 2429]. hardcover/softcover [2795, 2903, 2677, 2540]. Hardy [3339, 4223, 3483, 4289, 3555, 4324, 2738]. Harmonic [3763, 4076, 2092, 2672, 3078, 3089, 3133, 3219, 3797, 3752, 3708, 4456, 3828, 3654, 1811, 1812, 606, 610, 616, 1073, 1699]. Harmony [3221]. Harris [2345]. Hartree [3929, 2023, 2705]. Harvard [2261]. Haunted [1129]. Having [3008, 1397, 101]. Hawking [2624, 2278, 3725]. Hayden [2874, 3079, 3086]. He-like [3318]. Healey [4181]. Heart [4163]. Heat [3788, 400, 821, 1132, 1131]. Heavy [3454, 3614, 2342]. Heavy-Ion [2342]. Hecht [2011]. Hegel [2599]. Height [3259]. Heiles [2457]. Heisenberg [290, 4555, 204, 4690, 4268, 2085, 2920, 3621, 4018, 1811, 1812, 3561, 4130, 3460, 3387, 959, 3773, 356, 2670, 4450, 4164, 2172, 1478, 3760]. Heisenberg-Picture [2670]. Helicity [3593, 1141, 1773, 1807]. Helium [3257, 3013]. Heliumlike [2096]. Helix [2345]. Helmholtz [3113, 2033, 2428, 2426, 2427, 2852]. Helmut [2558]. Helsinki [1608]. Hénon [2457]. Henry [1441, 1452, 1434, 2561, 142, 2285]. Heras [3113]. heresies [254]. hermeneutical [1347]. Hermitian [3211, 4028, 1526]. Hertz [269, 319, 2859]. Hestenes [1572, 2914, 1575]. Heuristic [4342, 869, 4521]. Heyting [2412]. Hidden [238, 3359, 2135, 4204, 3659, 3984, 4348, 1327, 2868, 2971, 3311, 3485, 3040, 1645, 2739, 3103, 3829, 3436, 3549, 2862, 3601, 2560, 3724, 4145, 2955, 2956, 4092, 3082, 2369, 1886, 469, 1785, 1996, 338, 801, 762, 212, 659, 217, 879, 2028, 2030, 910, 1397, 196, 1988, 3221]. Hidden-Variable [3659, 4348, 2868, 3311, 2739, 4145, 4092, 469, 212, 659, 910]. hidden-variables [1996]. Hide [2034]. Hiding [2747, 3743]. Hierachies [4195]. Hierarchical [3702]. Hierarchy [4194, 80]. Higgs [538, 4056, 4192, 1687, 4429, 3793]. High [3561, 4645, 2995, 2337, 3353, 2785, 2446, 4731, 316, 412, 1589, 489]. High-Dimensional [4731, 1589]. High-Energy [2337]. High-Frequency [3353]. High-redshift [2995]. Higher [504, 3139, 3029, 2718, 3970, 3793, 2692, 2315, 1991, 198, 457, 4632, 4647]. Higher-Order [2718, 3970, 504, 1991]. higher-spin [457]. Highlighting [3899]. Highlights [2691, 1394, 826]. Highly [3324, 3427, 181, 391]. Hilbert [1945, 2535, 1785, 4527, 3645, 1225, 4224, 3223, 3439, 1581, 1909, 2671, 2962, 1694, 1618, 1010, 1419, 1277, 3490, 331, 1692, 3793, 2024, 2182]. Hilbert-space [1909]. Hilbertian [1784]. Hiley [3633, 3632]. Himself [2952]. Historical [4191, 2845, 2558, 2557, 2997, 4630, 4638, 4643, 4656, 33, 1414]. Histories [3418, 4251, 3046, 2199, 3694, 2065, 3628, 3462, 3692, 3766, 1593, 1898]. historiography [1868]. History [1895, 2401, 2192, 1871, 32, 1023]. Hodology [4546]. Hole [2387, 2658, 4124, 4233, 3310, 2672, 3754, 2356, 4239, 3136, 2664, 4236, 3998, 3080, 3085, 4730, 2656, 3338, 4235, 3090, 4082, 4129, 4726, 2663, 3743, 4238, 3803, 4069, 4701, 4414, 4230, 4520, 2654, 2667, 2653, 3940, 492]. Holes [2918, 4369, 2646, 4109, 2812, 4116, 4157, 4043, 3168, 4357, 2645, 3973, 3064, 4126, 2657, 4110, 858]. Holism [4264, 3538]. holistic [772]. Holographic [3310, 4373, 3227, 3920]. Holography [3568, 3403]. Holomorphic [3489, 1135]. Holonomic [2692]. Holt [3737]. Home [2656, 2495]. Homer [4102]. Homogeneous [4340, 2482, 2612, 3806, 2464, 864, 1772, 756]. Homotopy [3480, 4235]. Hon [2177]. Hoops [3371]. Hopf [1938]. Horava [3597]. Horizon [4115, 2666, 3090]. Horizons [3716, 3443, 4136, 1755, 3717, 3747]. horizontal [1681]. Horne [2223, 3737, 1299, 4348, 1604, 4431, 2229, 2685, 4495]. Horwitz [4042, 1954, 1944]. Hot [3454, 1783]. Hove [4425]. Hubbard [2436, 2438, 2446]. Hubble [159, 4029]. Hubert [2687]. Human [4721, 324]. Humean [4279]. Humeanism [4152]. Humpty [1070]. HV [3829]. Hybrid [3689]. Hydrodynamic [4568, 4602, 3848, 3870]. hydrodynamical [1372]. Hydrodynamics [4158, 3928, 4263, 3911, 3944]. Hydrogen [2461, 2484, 3851, 4707, 1557, 301, 1612, 408, 1536, 1519]. Hydrogenic [2188, 2385]. Hyper [2734]. Hyper-Mathematics [2734]. Hyperbolic [2537, 3515, 2143, 2384, 2634, 2002, 607]. Hypercomplex [1892, 314]. hypergraphs [877, 995]. Hyperion [3175]. Hyperplane [1103]. Hyperspace [2315, 4632, 4647]. Hypothesis [4187, 4509, 2702, 3685, 4078, 4366, 4336, 4691, 3065, 4728, 4034, 1454, 165, 288]. Hypothetical [2235].

i-xvi [2497]. i-xxv [2541]. I. [3985]. Ia [2995]. IARD [3393]. Iceberg [3209]. Idea [4248, 2847]. Ideal [4715, 115]. Idealisations [4278]. Idealizations [3844]. ideas [1466, 511, 1006, 512]. Identical [2447, 3183, 4504]. Identification [4544, 3735, 158]. Identities [1034, 3836, 2009, 841, 255, 1588, 1938, 537, 1354]. Identity [3774, 4465, 4329, 2044]. Ideograms [3965]. ignoring [1949]. Igor [2387, 2624]. II [2375, 1842, 1849, 1863, 1903, 912, 3283, 31, 294, 1040, 481, 1592, 1035, 1250, 1195, 715, 2612, 3219, 891, 249, 2661, 2595, 606, 1632, 70, 4379, 2555, 2761, 3128, 3290, 3439, 22, 134, 334, 767, 3029, 1373, 1599, 62, 4142, 251, 1005, 2240, 1409, 456, 636, 2281, 202, 1175, 3944, 76, 1264, 4466, 1423, 1751]. II. [4470]. IIa [209]. IIb [233]. III [922, 802, 610, 2798, 241, 36, 225, 1014, 2252, 309, 219, 160]. Illinois [2540]. illuminated [831]. Illusion [3369, 122]. Illusory [3587]. Illustration [356]. Illustrations [2215, 186]. illustrative [1689]. Ilya [976, 958]. imagery [1347]. Imaginary [1571, 3436]. Imaging [2805]. immortal [1011]. Impact [4306]. Imperfect [3783, 4403, 1766, 845]. imperfection [1407]. Implementation [3662]. implemented [1976]. Implicate [3639, 3637, 970, 393, 67]. Implications [1787, 3430, 4043, 3332, 2446, 2030, 4396, 4196, 1525, 615, 1077, 165, 100, 1121, 414, 1897, 1176]. Implied [2136, 152, 974]. Implies [2684, 2958]. Imploding [1877]. Importance [2236, 2557, 3872]. Important [25, 57, 139, 140, 187, 195, 248, 259, 1083, 1084, 1597, 2181, 1844]. Impossibility [194, 3762]. impossible [516]. Imprint [4709]. Imprints [2650, 3477]. Improbable [4161, 1089]. improper [312]. Improved [3576]. IMT [3111]. inaccuracy [1213]. Inaugurated [2839]. Incertisque [2132]. Incerto [2132]. Including [3947, 3839, 3091, 832]. Inclusion [3851]. Incompatibility [1775, 2929, 2739, 2317]. Incompatible [4725, 3695, 3184, 3624, 96, 1155, 478, 1705, 1704]. Incomplete [2857, 4035, 2960]. Incompleteness [3281, 3841, 2859, 1066, 1220]. Incompressibility [3520]. Inconsistencies [765]. Inconsistency [1329, 2958]. inconstancy [1877]. Increased [291]. incredible [837]. indeed [1984]. Indefinite [2697]. Independence [2191, 4508, 921, 2777, 3367, 2328, 52, 162, 918, 1738]. Independent [3737, 2318]. Indeterminacy [4590, 3773, 425, 1194, 1195]. Indeterminism [3982, 3804, 4254]. Indeterministic [4552]. Index [2196, 2304, 2442, 2548, 2651, 2758, 3960, 1530]. indication [32, 67]. Indirect [2340]. indirectly [1025]. Indistinguishability [2307, 3774, 3075, 1723, 2364, 3640, 1319, 2695]. indistinguishable [1041, 1763]. Individual [4138, 1686, 2135, 2145, 3682, 1411]. Individuality [3772, 34, 70]. Individuals [3770]. Individuation [3363, 3411]. Indivisibility [3776]. Induced [4695, 3674, 3873, 3328, 2999, 3520, 4567, 2966, 2276, 2931, 4156, 2106, 3649, 3834, 3791, 4051, 2453, 2157, 3119, 3562, 3448, 2097, 642, 2056, 575, 1451, 169, 117, 1697, 1843]. Inducing [3267]. Induction [3845, 66, 229, 477]. ineffectiveness [1087]. Inelastic [2383]. Inequalities [2394, 2935, 3544, 2942, 3561, 1315, 3424, 3470, 3476, 3821, 3986, 2950, 3434, 3985, 4096, 2228, 2560, 3835, 3724, 2983, 1678, 710, 1450, 1982, 1681, 1116, 405, 1443, 1449, 1302, 572, 1300, 1412]. Inequality [3737, 2985, 3647, 2700, 3819, 3850, 4185, 4052, 2840, 212, 1891]. Inequivalence [1878]. Inequivalent [3992, 1906]. Inertia [2309, 3584, 2349, 2251, 2492, 2832, 1848]. Inertial [2134, 4600, 3663, 185, 158, 399, 1801]. Inevitability [4426]. inexactness [84]. Infeld [3542]. Inference [4308, 3788, 3530, 1979, 299, 1214, 702, 1312]. Inferences [3766]. Inferential [2865]. Inferred [3152]. Infinite [2682, 3806, 4405, 3129, 4077, 2290, 2653, 1227]. infinite-dimensional [1227]. Infinite-Time [4077]. infinitely [29, 3251]. Infinitesimal [2790]. Infinity [3392, 3502, 4396, 676, 854, 744]. Inflation [4006, 3832, 3999, 4118]. Inflationary [4056, 3842, 4240, 3926]. Influence [4561, 1766, 337, 730]. Influences [3244, 3724, 1898, 439]. Information [3863, 1184, 1192, 2367, 2378, 2390, 3860, 2894, 2865, 4095, 3761, 3745, 2936, 2301, 4544, 2938, 4203, 2940, 3238, 2885, 2942, 2864, 3824, 2744, 2031, 3573, 781, 3894, 3123, 2621, 4606, 2686, 3529, 2279, 4306, 3437, 4501, 3293, 3012, 4407, 3862, 3870, 2949, 2948, 1838, 2126, 4114, 2259, 3924, 3834, 3443, 2638, 4082, 4010, 3743, 3803, 3973, 3853, 4025, 4208, 3570, 3543, 2874, 4084, 1412, 3942, 3838, 1122, 218, 1033, 1040, 1787, 4658, 1834, 1341, 276, 266, 2032, 4082, 2541, 2899]. Information-Based [3863]. Information-Induced [3834]. Information-Theoretic [2744, 4606, 3529, 1412, 218, 276]. Informational [3976, 4017, 1220, 4452, 935, 949]. Informationally [2946]. Infrared [4601, 1846, 581]. inherently [1345]. Inhomogeneities [4659]. Inhomogeneous [3946]. Initial [4037, 4118, 3692, 867]. Initialization [2202]. inner [1261]. Inseparability [3979, 1286]. Insight [3290]. Insights [2480, 3059, 3116, 4414, 4700, 517]. insolubility [887]. Inspired [3272, 4014]. Instabilities [2658]. Instantaneous [3113, 4403]. Instantons [2333]. Instead [488]. Institute [1184, 1192]. instrumental [138, 467]. instrumentalism [1405]. Instruments [4311, 3237, 1766, 767]. Insufficient [2319]. Integrability [3629, 761, 3579]. Integrable [3621, 2267, 3890]. Integral [4026, 2897, 2581, 4445, 3868, 4350, 2109, 4571, 1742, 104, 1699]. Integrality [4470]. Integrals [2147, 3381, 4259, 2447, 3461, 4264, 3480, 4135, 2448, 4422, 4436, 574]. Integrated [3745, 3744, 3834]. Integration [2427, 1382, 1552, 1562, 1477]. intellectual [1394]. Intelligence [3908]. Intelligent [2772]. Intelligible [2878]. intensities [890]. Inter [4066]. Inter-model [4066]. Interacting [1883, 3636, 2656, 2154, 3218, 3793, 864, 1025, 1205, 62, 768]. Interaction [4050, 4380, 3203, 4262, 2774, 2775, 4178, 2550, 4698, 479, 4284, 4154, 1987, 3265, 2122, 4311, 232, 2674, 4260, 4360, 157, 161, 1555, 703, 682, 1599, 1502, 68, 76, 160, 1012, 1661, 664]. Interaction-Free [4050, 3203, 4284, 2674, 4360, 1555]. Interaction-line [479]. Interactions [2105, 2846, 2289, 3499, 4584, 4404, 3116, 4736, 2218, 2859, 2603, 1515, 275, 1614, 443, 1980, 459, 374, 1687, 168, 819, 354, 1422, 297, 77, 101]. interacts [2035]. Interchangeability [4154]. Interdisciplinary [2392]. Interfacing [3168]. Interference [1724, 2082, 3099, 3068, 2713, 3024, 2929, 3970, 4459, 3961, 3193, 3471, 3562, 2613, 1012, 3817, 4613, 890, 891, 1723, 1487]. interferences [1400]. Interfering [4010]. Interferometer [2619, 2095]. Interferometers [2357, 3515]. Interferometric [4097]. Interferometry [3781, 2210, 2856, 3508, 2213, 3514, 3518, 2605, 2218, 3516, 1730, 992, 803, 1201]. Interior [3136]. Intermediate [3527, 1689]. Internal [4526, 3166, 2321, 1717]. International [1842, 1849, 1863, 1903, 940, 1570, 1616, 1630]. Internet [2816]. Interpret [2830]. Interpretation [4542, 3255, 3860, 2396, 4280, 3699, 2584, 3627, 3224, 3842, 4568, 3898, 3242, 3885, 4606, 2686, 3715, 4039, 2629, 2780, 4258, 3651, 2119, 2868, 3311, 4564, 2565, 2631, 3581, 3924, 3387, 3003, 3122, 4094, 4491, 3906, 2762, 2273, 3493, 3773, 4029, 2887, 3891, 4229, 3776, 3943, 2626, 2670, 1303, 2469, 3175, 3202, 2173, 3589, 1372, 2530, 3872, 3291, 3152, 3299, 4272, 1344, 2797, 290, 4630, 4638, 4643, 4656, 71, 1013, 344, 1890, 1719, 662, 4690, 161, 1142, 1742, 1465, 1180, 323, 891, 3275, 765, 521, 1270, 55, 69, 134, 408, 1800, 61, 40, 941, 511, 1544, 1752]. interpretation [1721, 1364, 1986, 704, 63, 1007, 653, 111, 1172, 68, 76, 160, 1173, 1329, 4650, 74, 1718, 1600, 1801, 9, 240, 1061, 512, 1311]. Interpretations [1905, 2263, 2877, 4038, 3536, 2215, 4491, 2639, 4270, 2065, 3401, 2529, 1705, 3866, 3977, 4045, 1704, 4624, 1982, 23]. Interpretative [3203]. Interpreted [3416]. Interpreting [4526, 3178, 4619, 2236, 3193, 4442, 2101]. Interval [1235]. intervals [420]. intricacies [1117]. Intricate [4330]. intriguing [446]. Intrinsic [3361, 3222, 2948, 4249]. Introducing [2917]. Introduction [4095, 3631, 4616, 4671, 3248, 461, 3952, 3292, 4416, 4634, 3717, 640, 2184, 4641, 4677, 2376, 2475]. intuition [1949]. Intuitionist [4291]. Intuitionistic [3241, 3374]. Intuitive [4565, 152, 1074]. invalidation [446]. Invariance [4538, 4385, 4322, 3126, 3238, 3328, 2289, 2964, 2693, 4292, 2258, 2825, 4438, 2510, 2062, 3063, 1735, 2732, 2397, 3154, 2547, 537, 2515, 2404, 3379, 1304, 1832, 1939, 119, 1158, 1940, 1354, 562, 540, 1458, 2591]. Invariant [2374, 3763, 2806, 3025, 4204, 3394, 3303, 2925, 3829, 3595, 2018, 2606, 3567, 3181, 1521, 302, 1364, 277, 1586, 149]. Invariants [4130, 2280, 1985]. invention [1406]. Inverse [3757, 1883, 1230]. Inversion [2425, 280, 386, 385, 1588]. invertibility [1589]. Investigating [4347, 3848]. Investigation [4498]. Invisibility [4041]. invisible [549]. Invited [2477]. Involutive [3566]. Involving [2215, 3385, 1456]. Ion [3614, 2342]. Ionic [2753]. ionization [117]. Ions [3318, 2978, 2096]. Irrealism [4223]. Irreducible [4180, 241]. Irremediably [3362]. Irreversibility [3994, 962, 2974, 969, 270, 3801, 970, 230, 249, 431, 669, 700]. Irreversible [3143, 3686, 46, 62, 116, 112, 182, 656, 309]. Irsee [1184, 1192]. Irsee/Bavaria [1184, 1192]. Isaac [2541]. ISBN [2345, 2376, 2590, 2641, 2786, 2783, 2708, 2609, 2600, 2795, 2784, 2836, 2767, 2495, 2494, 2565, 2631, 2903, 2539, 2821, 2677, 2785, 2316, 2808, 2617, 2820, 2496, 2687, 2766, 2315, 2624, 2967]. Ischebeck [2899]. Ising [4313, 136]. Iso [2734]. Iso- [2734]. Isoduals [2734]. Isolated [2918, 2297, 4400, 3484, 4445, 1923]. Isolation [3570]. isometries [688]. Isomorphism [3341]. Isospin [762]. isotopes [2013]. isotropy [921]. Israel [4031]. Issue [2477, 2399, 4004, 3769, 4610, 3804, 4306, 3556, 3818, 4017, 4304, 4307, 3313, 4383, 3793, 4272]. Issues [4117, 2003, 2546, 3930]. Itô’s [3430]. Iteration [3674]. Itinerant [2433, 2432]. Itself [4242, 1407]. IV [878, 616, 2253]. Ives [3306]. ix [2600].

J [3113, 2245, 2565, 2794, 2475]. J. [3221]. Jack [2565]. Jackson [2116, 3017, 2193]. Jacob [2668]. Jacobi [2482, 2612, 3397, 3521, 4557, 4481, 1334]. Jácome [4448]. Jagdish [2558, 2539]. James [2867]. Jammer [1864, 2365]. Jan [2496]. Jang [3522]. Janis [423]. Jarrett [3213]. Jauch [154]. Jay [4448, 4725]. Jaynes [2784, 3802, 949]. Jean [1736, 1710]. Jean-Pierre [1710]. Jeans [1734].

Jeffrey [2101, 4343]. Jeroen [3407]. Jersey [2499]. Jet [2339]. Joe [3469]. Joensuu [1138, 1147, 1185, 1193]. Johann [362]. Johannes [3247, 3228]. John [2128, 2193, 835, 842, 873, 1275, 1265, 871, 1266, 4685, 4686, 1325, 826, 870, 2600]. Johnson [3806]. Joint [1066, 843, 3196, 457, 2411, 988, 1450, 109, 1097, 383, 522, 564, 1705, 1855, 367, 1704]. Jordan [3704, 1991, 3897, 1246, 598, 446, 1172, 1226, 3760, 3968, 1457]. Jorge [2245]. José [2245]. Josef [154]. Josef-Maria [154]. Joseph [1486]. Josephson [1756, 1922]. Josiah [2183]. Jovian [3009]. Julian [2539, 2303, 2617, 2239, 2240, 2252, 2253, 2261, 4640]. July [2377]. jumps [1759]. junctions [1756]. June [1184, 1192, 1608, 2161, 2171, 2277, 2284, 2379, 2368, 2391, 2419, 2743, 2751, 2759, 2809]. Jürgen [2376, 1590]. Just [3158, 4107, 2343, 2616]. Just-So [2343]. Justifying [4077].

K-Mesons [2720]. Kac [4278]. Kähler [2121, 3164]. Kalb [3091]. Kaluza [838, 745, 867, 3400, 698, 728, 1074, 4226, 1985]. KAM [972]. Kamenshchik [2102]. Kant [3246]. Kantian [4091]. Kaon [3327]. Kaons [3332]. Kapitsa [2702]. Karl [502, 1392, 1399, 1468, 1393, 1394]. Károlyházy [2587]. Kastner [4032, 2249]. KDP [1539]. KdV [1881]. Kelvin [2503]. Kemmer [3936, 2198, 3620]. Kepler [2461, 1352, 4271, 2025]. Kernel [937, 1994]. Kerr [2812, 1804, 4133, 3185]. Key [3066]. Keys [2316]. Kiechle [2687]. Kiehn [385]. Kill [2873]. Killing [2790, 2815, 2827]. Kimball [2539]. Kind [4724, 526]. Kinematical [3227, 899, 45]. Kinematics [2682, 3959, 2661, 2671, 4105, 1204, 223, 747, 942, 424]. Kinetic [4221, 582, 2718]. Kingdom [2376, 2389, 2590, 2795, 2784, 2836, 2541, 2498]. Kip [2624]. Kirchhoff [4158]. Kite [3683]. Klein [4531, 745, 867, 714, 1333, 3807, 448, 2588, 3651, 2574, 3400, 1800, 3304, 2275, 838, 698, 728, 3395, 1074, 1700, 4226, 1985, 1298, 3291]. KLOE [3327]. Kluwer [2783, 2708, 2609, 2600, 2474, 2497, 2494, 2903, 2785, 2496, 2566, 2429]. KMS [562]. KMS-condition [562]. Knopf [2819]. Knot [3962]. Knotted [2484]. Know [3612, 2227]. knowable [4627]. Knowledge [4091, 2748, 689]. Known [3592]. Kochen [3450, 3043, 1887, 4376, 3739, 1768, 3756, 566, 1888, 3519, 4382, 3500, 3759]. Kolmogorovian [3661, 2029, 3984, 3974]. Koopman [4719]. Korteweg [3021]. Kramers [677]. Krinsky [2038]. Kristian [3387]. Kruskal [754]. Kubo [2281]. Kurt [2325]. Kustaanheimo [1552, 1562].

L [2541, 2821]. L. [3971]. Lab [2261]. labels [1294]. Laboratory [3718, 4469, 1732, 3066]. Lagrangian [3736, 3301, 1198, 3815, 3827, 1576, 3866, 3977, 3933]. Lagrangian-Path [3933]. Lagrangians [3397, 856]. Lahti [3247, 3228]. Lamb [546, 2189, 2188, 2116, 2091, 1612]. Lanczos [144, 2897]. Landau [766, 1197, 3016]. Landscape [3453, 3608]. Lanford [3801]. Langevin [1775]. Language [2405, 3349, 3858, 3122, 1081, 1921, 1146, 1242]. Laplace [3256]. Laplacian [4081, 2914]. Larc [3491]. Large [3745, 4134, 4405, 2543, 3016, 1621, 593, 166, 556, 2024, 1670]. large- [1621]. Large-Scale [4134]. Larger [3923]. largest [712]. Larmor [3909]. Larry [2784]. Lasenby [2808]. Laser [3781, 3067, 3744, 2091, 274, 840, 117]. laser-induced [117]. laser-ranged [840]. Lasers [2096]. Lasing [2096]. lasso [1968]. Last [2855, 2352, 4126, 1001]. Late [4006, 4572]. latent [1110]. later [1130, 794]. Lattice [3740, 2335, 1231, 541]. Lattices [3835, 1247, 1225, 1235, 1348]. Laudisa [4401]. Laughlin [3011]. Law [3661, 2826, 4295, 3126, 4107, 3527, 3123, 4297, 4027, 3116, 3479, 4025, 2141, 3114, 3779, 2420, 2327, 4009, 569, 745, 67, 821, 862, 1112, 1734, 1910, 889, 681, 1827, 100, 194, 1605, 2648, 3117, 4433, 974, 4068, 4029, 2383, 3268, 2810, 2566]. Lawless [3469]. Lawrence [4725, 1954, 2007, 4042, 1944]. Laws [4455, 2286, 2556, 4467, 4037, 2893, 2825, 4477, 4250, 3376, 2271, 3502, 2616, 2380, 2028, 375, 495]. Lax [2170, 2350, 1680]. Lead [2805]. Leader [2590]. Leads [4544]. Leaf [2755]. Learning [4732]. Least [3796, 3956, 3772]. Lebesgue [1172]. Lectures [2912]. Lee [2043, 2350, 2359]. Legacy [2967, 4673, 446]. Leggett [3647, 3659, 3709, 3162, 3206, 3526, 3192]. Lehnert [2374]. Leibniz [2405, 4232]. Leibowitz [3221]. Lemaître [4121, 4123]. Lemma [615, 3430, 1281, 1282]. Lemmas [2414]. Length [2553, 4288, 2901, 3807, 709, 221]. lengths [1457]. Less [3459, 3842, 3926, 4119]. Lessons [3795, 4632, 4647, 2315]. Lett [3113]. Letter [3312]. Letters [2368, 2391, 2419, 2430, 2443, 2669, 2678, 2689, 2703, 2709, 2716, 2725, 2735, 2743, 2751, 2759, 2769, 2776, 2787, 2796, 2802, 2809, 2813, 2822, 2829, 2837, 2866, 2875, 2075, 2104, 2117, 2131, 2139, 2149, 2161, 2171, 2179, 2185, 2194, 2204, 2220, 2232, 2246, 2256, 2262, 2268, 2277, 2284, 2293, 2306, 2346, 2379, 2403, 2455, 2467, 2476, 2488, 2500, 2509, 2517, 2523, 2528, 2533, 2542, 2549, 2559, 2567, 2575, 2583, 2592, 2601, 2610, 2618, 2625, 2632, 2643, 2652, 2660]. Level [2301, 2481, 2093, 3791, 3952, 294, 782, 3241, 3036, 309, 442]. Levels [2485, 3639, 1557, 466]. lever [713]. Levine [899]. Lewis [2090]. Liar [3148]. Liboff [3405]. LIBRA [3329]. LiCN [2452]. Lie [4691, 314, 2016, 4535, 326, 334, 583, 2018]. Lie-admissible [2018]. Life [4116, 4720, 3124, 2103, 1089, 1765, 4623, 4637, 4642, 6, 2345]. lifelong [1394]. Lifetime [2918, 1765]. Lifetimes [3586]. Lifshitz [3597, 3016]. Liftings [3422]. Light [2334, 4008, 3959, 4038, 3306, 2853, 4241, 3685, 2856, 3598, 4588, 4185, 4243, 2728, 608, 2174, 2973, 3403, 2508, 3344, 4700, 4638, 4643, 4656, 1033, 1040, 4613, 1730, 1620, 389, 528, 1741, 323, 378, 527, 412, 1900, 41, 336, 362, 335, 921, 660, 732, 1449, 131, 339, 1015, 846, 414, 1880, 1189]. Light-Cone [2334, 608]. light-front [1620]. Light-Fronts [2508]. Light-slice [3403]. Lightman [2624]. Lightning [2702]. lights [4630]. Like [4506, 3431, 2922, 3149, 1930, 3622, 3870, 2166, 2525, 3212, 3825, 3094, 3890, 1466, 1070, 604, 3318, 1970, 1997, 3193, 437, 2264]. likelihood [1240, 643]. Limit [4483, 3199, 4337, 3320, 2888, 3151, 4107, 2999, 3751, 3224, 3409, 4083, 3576, 4364, 4477, 4214, 3249, 2803, 4589, 4450, 4063, 2603, 1799, 1574]. limitation [1732]. Limitations [1278, 4482, 18, 4537, 918]. Limited [1967, 3538, 783]. Limits [4181, 2964, 4132, 4186, 3303, 3163, 1078, 3818, 4304, 3958, 4077, 3169, 4122]. Lindblad [3447]. Lindbladian [4567]. Lindbladian-Induced [4567]. Line [1436, 3353, 4350, 2451, 4532, 560, 479]. Linear [1359, 1284, 2287, 1583, 2945, 2636, 3058, 2906, 3490, 1579, 2544, 2970, 711, 4579, 412, 686, 981, 543, 565, 1189, 1298]. Linearity [705]. Linearized [4527, 48]. lines [609, 496, 1534, 1519]. lingo [1296]. link [989]. Linked [4560]. Links [2664, 611]. Liouville [993, 421]. Liquid [3284, 3197, 3966, 2667]. Liquids [3354]. List [1639, 2075, 2104, 2117, 2131, 2139, 2149, 2161, 2171, 2179, 2185, 2194, 2204, 2220, 2232, 2246, 2256, 2262, 2268, 2277, 2284, 2293, 2306, 2379, 2368, 2391, 2419, 2430, 2443, 2669, 2678, 2689, 2703, 2709, 2716, 2725, 2735, 2743, 2751, 2759, 2769, 2776, 2787, 2796, 2802, 2809, 2813, 2822, 2829, 2837, 2891, 2961, 3350]. Little [4370]. liv [2609]. Live [2916]. Lives [4324, 4137]. Lobachevski [1769]. Local [3143, 3665, 3704, 2922, 2223, 3328, 2964, 2637, 2580, 2393, 3798, 4348, 3846, 3040, 3447, 4525, 3195, 3103, 3906, 3829, 1953, 2605, 4389, 3213, 3390, 3684, 3887, 2569, 2547, 2591, 4321, 4052, 2840, 4298, 3971, 4054, 1076, 4145, 2986, 1809, 4137, 1060, 1368, 3730, 2635, 102, 3659, 1317, 1996, 1936, 3879, 1158, 1604, 1253, 255, 4367, 4111, 457, 1444, 910, 270, 10, 1655, 1644, 623, 629, 2950]. Local-Realistic [4054]. Locality [3244, 2538, 3592, 2225, 2190, 2860, 2858, 2780, 4026, 3441, 3552, 980, 4146, 983, 4363, 3467, 3666, 4035, 4217, 1924, 792, 2626, 3466, 4052, 3971, 3624, 3368, 793, 3550, 60, 238, 4376, 4624, 891, 3638, 497, 88, 405, 680, 3555, 1414, 1028, 1297, 1951, 379, 298, 3562, 3079, 3086, 2254]. Localizable [4364]. Localization [4138, 4169, 3308, 3682, 2577, 2712, 800, 3629, 1454, 890, 891, 1908, 1759]. localizations [1203]. Localized [1271, 3782, 2459, 2100, 1461]. Locally [3758, 3525, 2950, 3171, 864]. locally-interacting [864]. Locating [3792]. Loci [2132]. Logamediate [3527]. Logarithmic [744, 2672, 758]. Logic [2398, 4203, 2399, 3241, 3774, 2784, 3731, 2178, 3571, 2570, 3395, 3825, 3859, 3991, 4457, 4466, 3153, 3374, 1406, 3657, 90, 1695, 1979, 814, 641, 1228, 1234, 11, 4661, 995, 1393, 125, 1451, 1152, 369, 1738, 4626, 1163, 1337, 919]. Logical [3771, 4399, 2989, 3435, 1211, 1738, 4431, 3778, 1140, 473, 15]. Logics [3371, 2560, 2400, 3395, 75, 462, 1143, 1679, 1688, 1556, 1602, 1246, 434, 1692, 1322, 1216, 1348]. Lokanathan [2903]. London [2641, 2786, 2708, 2609, 2600, 2496, 2624, 2429, 1905]. Long [2772, 374, 2281]. Long-range [374]. Long-Term [2772]. Long-Time [2281]. Longer [3586]. longitudinal [1706]. Look [2343, 3607, 3389, 3252, 179, 148, 475]. Looking [3955, 4206]. Loop [2672, 3621, 4091, 3270, 4152, 2515]. Loophole [3756, 3434, 3985, 2840, 3972, 3526]. Loophole-Free [2840]. loops [4653, 2687]. lore [254]. lore-and [254]. Lorentz [2591, 458, 3763, 3138, 4022, 4340, 2896, 3126, 2806, 2289, 2964, 3540, 3903, 1823, 1832, 1564, 4204, 2545, 4438, 3303, 2525, 2731, 2925, 3338, 349, 2911, 3063, 640, 1118, 1119, 1735, 1506, 1563, 1740, 493, 4536, 2547, 4271, 277, 4603, 3131, 4479, 3963, 1328, 1104, 607, 1743, 911]. Lorentz-Covariant [2911, 4271]. Lorentz-Invariant [4204, 3303, 277]. Lorentz-Scalar [2525]. Lorentz-Transformation-Based [2896]. Lorentzian [3711, 440, 498, 1985]. Lorenz [2374]. Loschmidt [103, 4206]. Loses [2884]. Losing [4069]. Loss [3251, 3283, 3123, 4082, 3973]. lost [500]. Louck [2011]. Louis [4610, 512, 519, 1470, 511, 991, 1489, 510, 513, 515]. Low [3701, 2340, 3038]. Low-Mass [2340]. Lower [4100, 2964, 2901]. Lowering [3915]. Ludwig [576, 585, 2415, 1642, 577]. Lukasiewicz [2400]. Luminal [3650]. Luminosity [3055]. Lunch [4247]. lutte [1710]. Luttinger [3284]. Lutz [2899]. LWI [2094]. Lyapunov [2563].

M [2708, 2609, 2242, 2494, 2821, 385, 3625, 3617]. M-Theory [3625, 3617]. M. [3126]. Maass [2471]. Mach [3355, 402, 247, 3622, 322, 2997, 37, 1458, 1988, 2071, 1899]. Machian [4644, 4463]. Machine [2996, 1356]. Machlup [384]. Mackey [1642, 1233]. macrocausality [623]. macroobservable [417]. macrophysical [1897]. Macrorealism [3526, 1342]. Macroscopic [4695, 2584, 4259, 4400, 3417, 3427, 2760, 2761, 2798, 3530, 3928, 3479, 3961, 2087, 4472, 359, 1711, 969, 377, 1743, 1542]. macroscopically [596, 595, 1202]. Macrosystems [2215]. Macroworld [2394]. Made [4565]. Madelung [3870, 3922, 4583, 4036]. Magic [4370]. Magnetic [4394, 2461, 3871, 2550, 547, 2332, 2340, 4708, 2326, 3620, 2107, 2437, 3812, 312, 2975, 4520, 2060, 4260, 3474, 2852, 878, 180, 661, 1698, 229, 397, 341, 2034]. Magnetized [2512]. Magneto [2908]. Magneto-fluid-dynamics [2908]. Magnetohydrodynamics [2503, 2001]. Magnetoresonances [1968]. Magnets [2842]. Magnifying [4206]. magnitudes [416]. Magnus [2045]. Mahood [2309]. major [278]. Majorana [2828, 3141, 2801, 2069, 3080, 3085, 3605, 3019, 3753, 716]. Make [3792, 4071, 2468]. makes [1554]. Making [3431, 4000, 4305, 4278, 3535, 2615, 2540, 2099, 975]. Malament [3356]. manifestation [1967]. Manifestations [553, 4461, 1209]. Manifesting [3732]. Manifestly [3394, 1455, 1792, 1162, 505]. Manifold [4498, 773, 994]. Manifolds [2962, 4253, 2102, 4087, 2834, 4620, 436]. manuals [1242, 1233]. Manuscript [3080, 3085]. Many [3251, 2938, 3351, 4573, 3248, 2494, 1010, 1419, 4530, 2400, 3471, 4263, 4429, 2065, 2913, 3977, 4137, 525, 71, 573, 225, 371, 637, 1986, 1358, 29, 4636]. Many-Body [3351, 4573, 2494, 4530, 573, 225, 371, 4636]. Many-Hilbert-spaces [1010, 1419]. Many-Particle [4263, 3977]. Many-Spin [2938]. many-universes [71]. Many-Valued [2400]. many-valuedness [637]. Many-Worlds [2065, 1986]. Map [3400, 4470]. mappings [1285, 199]. Maps [2250, 4098, 3835, 474, 4176, 428]. Mara [2540]. Marchildon [2763]. Marcin [4725]. Marek [4725, 3206]. Margenau [1441, 1452, 1434, 142, 2285]. margins [648, 902]. Maria [154]. Markers [4586]. market [4627]. Markiewicz [4725]. Markov [3232, 3846, 4098, 4389, 935]. Markov-informational [935]. Markovian [4274]. Marlan [2242]. Marriage [2261]. Martin [2431, 2444, 2456, 2478, 2477, 2480]. Masanes [4009]. Maser [732]. Mass [1295, 4288, 2812, 3701, 2258, 2270, 163, 2623, 3132, 2163, 3275, 3967, 2340, 2599, 2718, 3144, 3185, 4081, 3708, 3883, 2134, 2154, 4600, 2365, 1769, 1763, 1783, 1803, 1636, 207, 1523, 1369, 126, 419, 4182, 171, 499, 559, 37, 540, 1981, 2012, 718, 716]. Massachusetts [2345, 2565]. Masses [3591, 750]. Massive [4006, 3599, 3311, 2697, 2613, 1636]. Massless [4385, 4322, 2148, 4427, 1177, 1911]. Master [2044]. matched [743]. Matching [3707]. Material [3025, 3845, 3785]. Materiality [2858]. Mates [204]. Math [4527, 2993]. Mathematical [1972, 4705, 2893, 2702, 2593, 3220, 3445, 4574, 2982, 4582, 2820, 4206, 3134, 4664, 940, 1750, 1367, 104, 271, 1767, 178, 1298, 4669]. Mathematical-Billiard [4206]. Mathematically [3398]. Mathematics [2602, 2937, 2783, 4527, 4499, 3827, 3497, 2954, 3426, 3825, 2734, 3640, 4654, 4662]. Matrices [2335, 2880, 3232, 2038, 2014, 2041]. Matrix [2189, 2188, 2385, 3816, 192, 3008, 2882, 3600, 2634, 1197, 830, 1976, 494]. Matter [3284, 3348, 4371, 3263, 2085, 4568, 4549, 4128, 3948, 1708, 4713, 2995, 4590, 2993, 2342, 2267, 2564, 2598, 2931, 3811, 3795, 2058, 4112, 4078, 2699, 2675, 2752, 4459, 4443, 4113, 3853, 3560, 2734, 3202, 4326, 4139, 2071, 2834, 1295, 1783, 1835, 44, 670, 1877, 1425, 1659, 1760, 80, 1846, 1599, 1314, 2035, 755, 216, 470, 1262, 1646, 1988]. Matters [3454, 3377, 4627]. Matteucci [2551]. Maudlin [4082]. Maurice [551, 563, 727]. Max [2389, 2365, 1864]. MaxEnt [3530]. Maximal [1157, 3917]. Maximum [2649, 3749, 3258, 3446, 1240, 2638, 329]. Maxwell [4413, 2799, 3852, 3025, 3533, 1834, 1910, 2520, 3398, 1632, 597, 3907, 2925, 3914, 686, 2159, 2158, 2160, 3405, 1691, 1924, 4184, 2914, 4715, 4294, 2573, 2628, 4479, 2930, 1104, 1202, 1743]. Maxwellian [3117, 507]. Mayer [2565, 2816]. Mean [4733, 4544, 4032, 3833, 2436, 1765, 2023]. mean-field [2023]. Meaning [2345, 4219, 4525, 4089, 2404, 2249, 2674, 3156, 1837, 1564, 861, 212, 687, 1130, 1415, 757, 4637, 4642]. Meaningfulness [2825]. Means [4167, 3166, 3744, 2588, 3336, 2605, 3848, 1997]. Measurability [3196, 48]. Measure [4188, 2856, 4013, 4029, 2973, 3307, 4471, 85, 1241, 471, 1412, 2032]. Measured [3725]. Measurement [4592, 2683, 2208, 3203, 4280, 630, 2082, 3644, 2884, 1293, 3758, 4284, 4154, 3087, 3188, 3499, 4174, 3438, 4449, 4392, 3826, 4156, 2671, 3036, 4441, 4693, 4085, 4316, 3174, 3204, 3973, 4558, 3435, 3165, 188, 3419, 4366, 4336, 2132, 1085, 3145, 3491, 4261, 3567, 3923, 3543, 1680, 3153, 3738, 3626, 4210, 1049, 3913, 2318, 433, 296, 245, 1643, 164, 299, 887, 1048, 1285, 1895, 1142, 469, 1785, 353, 213, 918, 42, 1936, 532, 1756, 135, 526, 1603, 1062, 669, 1095, 1096, 659, 59, 700, 39, 321, 1435, 58, 96, 4476, 478, 729, 2030, 1355, 112, 1705, 848]. measurement [1896, 2172, 1542, 9, 960, 367, 623, 1704, 515, 1616, 1630]. Measurement-Based [3419, 3153]. measurement-disturbance [2172].

Measurement-Driven [4284]. Measurement-Induced [4156]. Measurements [4131, 3022, 4262, 3861, 2407, 2144, 4165, 2770, 4032, 3947, 4484, 574, 4567, 3096, 2946, 3279, 4559, 2780, 2393, 4274, 4003, 2793, 3245, 4323, 2271, 2881, 3957, 2216, 3882, 2862, 3465, 3697, 4341, 3023, 4423, 3462, 2080, 4537, 2674, 2197, 2959, 2960, 2429, 1837, 1427, 988, 1239, 1789, 1823, 1165, 1110, 1555, 1129, 1909, 1208, 1462, 2048, 1010, 4622, 321, 372, 943, 1897, 1655, 1644, 845, 2291]. Measures [3619, 2779, 3234, 1227, 4377, 3668, 3755, 734]. Measuring [1088, 2933, 3265, 767, 2357, 2118, 3734, 4311, 4126]. Mécanique [2184]. Mechanical [3993, 3300, 2295, 3823, 2323, 2066, 3053, 2817, 1506, 4316, 2266, 3465, 3287, 4366, 4336, 4728, 4479, 3676, 3880, 4031, 4045, 2317, 342, 157, 275, 596, 299, 1022, 1787, 1614, 1473, 1713, 1555, 859, 468, 97, 595, 244, 496, 1445, 1728, 730, 4476, 4627, 1259, 1258, 190, 623, 629, 173]. Mechanics [4542, 4483, 4228, 3199, 4506, 2200, 3255, 4548, 3860, 4629, 3888, 3178, 2556, 3420, 2086, 4605, 2307, 2314, 2890, 4016, 2584, 3799, 2222, 2877, 3634, 2778, 3627, 2876, 3995, 4143, 4330, 2082, 2885, 2168, 4485, 2981, 3511, 3454, 2449, 2864, 4398, 4717, 4115, 3949, 2999, 2974, 3899, 3383, 4511, 4318, 3224, 4343, 2783, 2057, 4489, 3327, 2708, 2878, 3242, 3186, 4493, 3839, 2801, 3815, 4579, 4725, 3646, 2880, 2621, 3707, 4527, 3096, 3391, 2917, 3885, 3969, 2686, 2857, 3529, 4437, 3777, 4064, 3980, 2593, 3108, 3087, 4451, 4541, 4718, 4714, 3796, 2402, 2629, 2780, 2818, 2915, 3428]. Mechanics [2594, 2595, 3590, 4517, 4463, 3293, 3879, 4293, 3323, 4616, 4671, 3006, 3373, 4163, 3012, 3258, 3170, 3841, 3177, 3277, 3565, 3294, 2119, 2868, 4378, 2170, 4564, 3150, 3223, 3439, 3764, 3322, 3321, 3363, 3411, 3826, 2003, 2058, 3827, 2259, 2929, 3818, 3924, 4304, 3297, 4166, 3387, 2480, 4015, 3810, 3960, 3076, 2558, 3772, 2962, 3399, 3928, 2903, 4500, 2881, 3443, 2215, 2739, 4094, 4491, 3008, 3906, 4058, 4570, 3385, 2059, 4090, 4142, 2762, 2830, 4146, 4574, 3249, 4332, 2183, 2577, 4518, 2581, 2561, 2176, 3046, 3743, 3918, 3142, 3109, 3179, 3260, 3467, 3666, 2924, 3694]. Mechanics [2803, 3868, 4333, 4397, 3266, 4077, 3601, 3727, 3746, 4035, 2397, 2953, 3887, 4217, 4412, 4411, 3854, 4270, 4492, 3662, 4617, 4456, 2804, 3395, 3367, 4263, 2112, 3776, 4344, 4486, 2670, 3005, 4710, 4071, 2070, 2110, 2311, 2913, 4054, 3226, 2173, 4719, 2504, 3975, 3589, 3917, 3558, 4607, 2530, 2529, 2647, 2585, 2782, 3382, 3866, 3977, 2420, 2576, 4529, 4136, 3528, 2364, 3933, 4457, 2219, 4533, 4106, 4282, 4537, 2873, 4303, 4220, 3164, 3904, 4524, 4575, 3429, 3079, 3086, 4374, 4401, 3042, 3152, 2230, 3778, 3770, 4034, 2493, 3477, 2318, 2797, 3269, 3714, 3719]. Mechanics [3102, 525, 1379, 609, 989, 4630, 4638, 4643, 4656, 228, 823, 888, 815, 587, 1013, 1066, 1656, 1464, 1871, 35, 1826, 164, 344, 1236, 504, 1890, 1347, 425, 649, 1048, 1293, 1440, 1025, 1792, 663, 662, 982, 1128, 1353, 1450, 1031, 4690, 109, 1142, 1931, 1426, 2027, 1054, 820, 963, 1065, 1823, 1939, 1400, 328, 1059, 1465, 4624, 1180, 1980, 1112, 66, 891, 357, 1204, 450, 946, 1123, 1317, 1996, 633, 1103, 588, 617, 192, 4666, 1937, 1234, 34, 70, 1984, 333, 73, 671, 148, 206, 1105, 1149, 521, 708, 618, 1097, 264, 1308, 197, 209]. mechanics [233, 241, 703, 1270, 961, 682, 1418, 1712, 1892, 2010, 1218, 150, 332, 876, 69, 134, 225, 300, 1078, 286, 767, 1023, 926, 1645, 1064, 427, 981, 511, 1006, 2014, 1117, 220, 251, 260, 322, 1000, 1005, 1014, 371, 453, 584, 1237, 1813, 522, 564, 1488, 4654, 4662, 345, 146, 1752, 5, 1721, 1364, 550, 147, 202, 219, 1986, 188, 692, 1154, 589, 792, 63, 271, 372, 2051, 1007, 653, 15, 111, 1162, 243, 268, 1992, 503, 729, 1242, 1173, 1501, 1372, 470, 793, 1705, 482, 75, 1671, 2004, 4650, 406, 973, 1012, 2001, 440]. mechanics [1718, 1600, 1801, 1060, 1644, 1068, 590, 215, 1201, 849, 394, 170, 240, 1061, 3, 367, 898, 1704, 4652, 306, 518, 512, 1044, 4725, 2565, 2631, 2292, 2316, 4042, 4089, 2496, 2475]. Mechanics-Matter [2058]. Mechanism [3899, 2446, 359]. Media [2100, 2504, 1794]. Mediated [4402]. medium [480, 1882]. Meet [4495, 4384]. Meeting [4183]. meetings [940]. Meets [4439, 4052]. Mehra [2558, 2539]. Meinhard [2565]. Membranes [2157, 1843]. Memoriam [2269, 2285]. Memory [3511, 3855, 377]. Mensky [2429, 3923]. Mercier [2269]. Mereology [3498]. Merging [4023]. Mermin [3010, 2757, 2749, 3188, 2736, 4348, 2750]. Merwe [2783, 2474]. Merzbacher [2292]. Meson [2985, 2543, 1977]. Mesons [2720, 2580, 285]. Mesoscopic [3789, 4550]. Meta [4185, 4510]. Meta-analysis [4185]. meta-empirical [4510]. Metallic [2441]. Metaphysical [3843]. Metaphysics [3939, 3671, 1868, 1294, 1416]. metastable [2052]. meters [1096]. Method [3729, 3445, 3828, 4471, 4470, 743, 823, 888, 31, 2054]. Methodological [3208]. methodology [86]. Methods [3674, 2820, 4655, 4664, 1529, 822, 4669, 2795]. Metric [3537, 3263, 2289, 3058, 3032, 4200, 2947, 3400, 4588, 4286, 2719, 2730, 2522, 4226, 2320, 2834, 480, 668, 553, 1158, 1854, 1995, 411, 1998]. metric-affine [1158]. metrical [575]. Metrically [2834]. Metrication [4030]. Metrics [3754, 2924]. Metzner [3403]. Mexico [2377]. Michael [2254, 2600, 2541, 2429]. Michel [2836, 2184, 3501, 2409]. Michell [3256]. Michelson [1658, 1761, 3515, 7, 880]. micro [220, 488]. Microcanonical [1606]. microcausality [671]. microgenometry [1985]. Micromaser [3190, 1500]. Micromorphic [2998]. Microphysical [1036]. Microphysics [2255, 370, 4628, 348]. microrealistic [251, 418]. Microscope [356]. Microscopic [359, 4027, 974, 2521, 2120, 1973, 975, 431, 1743]. Microwave [2462, 2469, 1542]. Midpoints [1929]. Might [3649, 464]. Miguel [2474]. Mikhailov [2852]. Mikio [2291, 2036]. Miklós [2600]. Milklós [2178]. Milky [3998]. Mills [2334, 4340, 698, 728, 1528]. Milton [2539]. Mind [4380, 2345, 44, 2993, 470, 2531, 157, 4627, 1396]. Mind-Body [4380, 157]. minds [4637, 4642]. Minimal [4280, 3807, 2946, 4064, 3764, 4013, 3901, 2694, 4056, 382, 1971]. Minimization [2530]. Minimum [3224, 2638, 126]. Minkowski [3582, 2289, 2790, 2815, 2827, 1854, 1995, 3105, 2771, 2146, 2585]. minor [278]. Mioara [2783]. Miracle [4076]. Miracles [3610]. Mirror [3610, 4326]. Mirrors [2856]. Misconceptions [4277, 465]. misinterpretations [1870]. Misleading [3297, 356]. miss [710]. Missed [3446]. Missing [4255, 2895]. misunderstanding [88]. misunderstandings [617]. Mittelstaedt [1148, 1139, 3347, 3350, 3348, 3805]. Mixed [3689, 4727, 2753, 377, 2634, 2182]. Mixed-system [377]. Mixing [2701]. Mixtures [3424, 51]. Möbius [1845]. Modal [3255, 2263, 2396, 4280, 2584, 2877, 3777, 2779, 4491, 4247, 2639, 4252, 3569, 2530, 2529, 919, 1718]. Modal-Hamiltonian [4491]. Modalities [3888]. Modality [4246]. Model [3027, 3431, 4011, 2396, 2187, 2312, 2399, 2896, 3151, 3794, 3758, 3568, 2853, 3946, 2828, 4738, 4117, 4148, 3948, 2629, 2780, 2996, 3424, 2091, 2329, 3258, 2436, 3765, 4348, 4560, 3303, 3444, 2342, 2564, 4519, 3053, 2831, 4404, 4182, 3254, 3557, 2337, 2800, 3029, 3003, 3210, 3305, 2201, 4588, 4005, 3919, 1735, 3129, 4367, 3330, 2438, 3882, 3750, 2217, 4724, 3887, 3662, 4715, 2154, 4298, 4425, 2446, 3971, 4054, 4145, 4479, 3935, 4415, 3654, 4070, 1196, 3762, 2350, 3382, 3579, 4502, 4045, 4576, 3171, 3102, 952, 1111, 340, 365, 45, 1621, 1538, 1895, 1803, 265]. model [604, 398, 1778, 1936, 1517, 520, 1756, 115, 753, 1174, 732, 30, 1697, 1687, 1973, 1136, 4066, 1825, 492, 327, 910, 210, 1999, 1993, 979, 1061]. Model-Theoretical [2800]. Modeling [3001, 4728, 1513]. Modelling [3829]. Models [4131, 4406, 3982, 3768, 3043, 2568, 2552, 4002, 3873, 3871, 4705, 3301, 4734, 3893, 3656, 4135, 3373, 4353, 3974, 4729, 3700, 4285, 2729, 4596, 3724, 4373, 3162, 4700, 4537, 2280, 3705, 2653, 3192, 3503, 2429, 530, 1621, 1799, 1473, 1878, 1790, 1276, 1512, 1446, 2013, 1759, 2044, 496, 2048, 1506, 1851, 2023, 256, 1754, 1528, 1655, 1245]. Moderately [2274]. Modern [4380, 1138, 1147, 1185, 1193, 1498, 1615, 1990, 4189, 3805, 4038, 4659, 4675, 4679, 2495, 2839, 4123, 1401, 3246, 3493, 1608, 1640, 4621, 1166, 49, 2004, 2786]. Modes [2148, 4583, 2274, 2281, 4551]. Modification [3400]. Modified [4458, 2896, 1187, 3658, 2271, 4553, 4051, 4586, 4373, 4218, 1800, 556, 406]. Modifying [3690]. Modular [3629]. Modulation [3318]. Modulus [2995]. Mohrhoff [2762]. Molecular [3860, 2452, 3120, 2437, 3019, 781, 732]. Molecular-Dynamics [2437]. Molecules [4722, 77, 101]. Moment [3871, 2550, 2332, 4708, 4265, 4260, 4470, 180, 229]. Momenta [3340, 2805, 909]. Moments [4605, 878, 1260]. Momentum [3468, 290, 2105, 4604, 1287, 2287, 2513, 3934, 3809, 3051, 3311, 2696, 3892, 3447, 4269, 3252, 4319, 2372, 4417, 832, 3968, 3092, 1284, 1525, 624, 841, 533, 1095, 4462, 305, 268, 1258, 43]. MOND [3890, 4299, 4356]. MOND-Like [3890]. MONDian [4144]. Monism [3639]. Monogamy [3912]. Monoids [3566]. monopole [1509]. Monopoles [2340, 547]. Montreal [2316]. Moon [2557]. moralizer [1265]. Morgan [2417]. Morley [1658, 1761, 7, 880]. Morse [2125]. Mortessagne [2836]. Mössbauer [3987, 74]. most [1549]. Motion [3306, 2943, 3815, 3691, 2299, 2760, 2721, 3786, 3396, 2166, 2525, 3925, 3502, 2695, 2907, 2133, 2134, 2360, 3016, 129, 574, 247, 633, 681, 759, 855, 1467, 40, 118, 390, 189, 993, 2054, 309, 87, 1729, 537, 346, 495, 505]. Motions [2914, 2908, 3074, 2516, 3026, 875, 1600]. Motivation [3380]. motivations [307]. Mott [4131, 2434]. Mount [2633]. Mountain [2539, 4640]. movement [431]. Moving [1758, 4319, 4462]. Moyal [3315, 3294]. MPT [505]. Much [3158, 3665, 3644, 4119, 1545]. Mugur [2783]. Multi [2433, 4318, 4046, 3515]. Multi-Band [2433]. Multi-fingered [4318]. Multi-Path [3515]. Multi-Time [4046]. Multicarrier [2753]. Multiconductor [3353]. multidimensional [688]. Multilocal [310]. Multinodal [4512]. Multinomial [3149]. Multipart [1437]. Multiparticle [2224, 2226, 4497]. Multipartite [4099]. Multiple [2222, 4046, 1427]. multiplicity [2011]. Multiply [4422]. Multiply-Connected [4422]. Multipole [1540]. Multivalued [3395]. multivector [1576]. Multiverse [4197, 4720, 4199]. Muon [3388]. Muonic [538]. Must [4157, 2953]. Mutual [3422]. Mutually [4167, 211, 3023, 3558]. Muynck [2708]. MV [2412]. my [513]. Mysteries [2748]. Mystery [3602]. Mysticism [2748, 4627]. Myth [2919]. Myths [3109].

N [3010]. N. [2757]. NaI [3324]. Naimark [1095, 1096]. name [4627]. Namiki [2036, 2291]. nanotubes [3748]. Narratability [3918]. Natanzon [1529]. Nathan [684, 699, 675, 1850]. NATO [940, 1184, 1192]. Natural [4193, 2144, 4229, 4177, 66]. Naturalness [4191, 4192, 4199, 4429, 4196, 4198]. Nature [4455, 3597, 4037, 4562, 2621, 4250, 3534, 3376, 2531, 4137, 4706, 864, 1269, 353, 1342, 1341, 615, 34, 70, 390, 1408, 1015, 373, 1721, 1585, 1983, 2042, 718, 23]. Naturgesetze [1022]. Nauenberg [3137]. Navier [1346]. Near [4134, 447]. Nearly [1694]. necessary [1345]. Necessity [3971, 263]. Need [3737, 4476]. needed [200]. Needs [4074]. Negations [4439]. Negative [2072, 3809, 2976, 1853]. negative-frequency [1853]. Negativity [4018]. Neglected [3653]. Neither [4179, 4211]. Nelson [4716, 1728]. Neo [4605, 2797, 1074]. Neo-classical [4605]. Neo-Copenhagen [2797]. neo-intuitive [1074]. Netherlands [2783, 2474, 2494, 2903, 2785, 2566]. nets [4574]. Networks [4409, 3628]. Neuenschwander [3482]. Neumann [2600, 4388, 3217, 3359, 918, 4102, 39, 312, 4685, 4686, 4719, 4488]. Neumark [1286]. Neural [4409, 377]. Neutral [3327, 3111, 3472, 3332, 4260]. Neutrality [4371, 3843]. Neutrino [3468, 4288, 3669, 2163, 3967, 2343, 2831, 3065]. Neutrinoless [2611, 2688, 2698, 2024]. Neutrinos [2699, 695, 716]. Neutron [2650, 992, 2210, 3508, 2214, 3518, 3560, 2512, 3516, 1730, 803, 1487, 1201]. Neutrons [3519]. never [227]. Newman [2812, 1804, 4133, 3185]. Newton [4012, 1908, 3502, 1691, 2035, 2197, 1485, 4551]. Newtonian [4563, 1635, 4144, 3271, 4079, 4292, 4201, 3749, 121, 3594, 3150, 4368, 4081, 4352, 4347, 1767, 2557, 295, 4299, 4356, 3908]. Newtonian- [4144]. Next [2772]. Nick [2641]. Niels [2876, 959]. Nielsen [2541]. Nieuwenhuizen [3972, 4047]. Nil [3967]. Nine [314]. Nine-vectors [314]. ninetieth [591]. ninety [1394]. NJ [2767, 3010]. No [3239, 3359, 4734, 4315, 4239, 3707, 1342, 4418, 4718, 2949, 4302, 4487, 4129, 4453, 4412, 4229, 3084, 3107, 4303, 3904, 3092, 484]. No-Cloning [2949]. No-Deleting [2949]. No-Forcing [3707]. No-Go [4734, 1342, 4418, 4412, 4229, 3904, 3092]. No-Matching [3707].

No-Splitting [4487]. Noble [3069]. Nobody [4343]. Nodal [4272]. Nodded [4102]. Noether [3482, 2009, 4088, 4684]. Noise [4131, 4165, 2483, 975, 1302, 848]. Noise-Disturbance [4165]. Non [4733, 3244, 2538, 3661, 3200, 3486, 3993, 4340, 3663, 2399, 3126, 4376, 4056, 4398, 3078, 3089, 3133, 3219, 4061, 4579, 4545, 3659, 2858, 3984, 3638, 3590, 3512, 4274, 4026, 3879, 2836, 1989, 4709, 2502, 4160, 4348, 3667, 3303, 3517, 3764, 3772, 4150, 3195, 4682, 2892, 3090, 3603, 3555, 2605, 3901, 4367, 3309, 4111, 2581, 3479, 4729, 3782, 4701, 2536, 4456, 1114, 4396, 2529, 3562, 3567, 3382, 2692, 3880, 3528, 3933, 4457, 4408, 4126, 4735, 2694, 2619, 3079, 3086, 3908, 3778, 2635, 4624, 244, 4667, 4676, 2016, 136, 2021, 1973, 1528, 2313, 294, 3324, 1811]. Non [1812, 688, 4031, 2254]. Non-Abelian [688]. Non-Archimedean [2313]. Non-Bohmian [3200]. Non-Cartesian [2536]. Non-Classical [2619, 3528]. Non-Commutative [3764, 3567, 3603]. Non-compact [3078, 3089, 3133, 3219]. Non-conservation [4398]. Non-conservative [3517]. Non-Contextual [4348]. Non-dirichlet [4701]. Non-Equilibrium [2836, 2502, 4682, 3901, 4667, 4676, 1973]. Non-exponential [3590]. Non-Geometrodynamic [4340]. Non-Gravity [3486]. non-Hamiltonian [136, 294]. Non-Heisenberg [1811, 1812]. Non-Holonomic [2692]. Non-Individuality [3772]. Non-inertial [3663]. Non-Kolmogorovian [3661, 3984]. non-lie [2016]. Non-linear [4579]. Non-Local [3195, 2605, 2635, 3659, 3879, 4367, 4111]. Non-Locality [3244, 2538, 2858, 4376, 3638, 3555, 3562, 3079, 3086, 4624, 2254]. Non-Lorentz-Invariance [3126]. Non-Markovian [4274]. Non-Materiality [2858]. Non-Minimal [2694, 4056]. Non-Newtonian [3908]. Non-Orthogonal [2529]. Non-Paulian [3324]. Non-Perturbative [2892, 4160]. Non-Quantum [2399, 3528, 4408]. non-quantum-mechanical [244]. Non-realism [3512]. Non-reflexive [3778]. Non-Relativistic [2581, 3382, 3933, 4735, 4026, 1989, 3303, 4456, 4457]. Non-renormalizability [4061]. Non-representative [3993, 3880, 4031]. non-Riemannian [2021]. Non-separability [3667]. Non-Singular [3090, 3479, 4126]. Non-smooth [4150]. Non-standard [1114]. Non-symmetric [4729]. Non-trivial [3782]. Non-unique [3309]. Non-vanishing [4709]. non-Yang [1528]. Non-zero [4396]. Nonadiabatic [1945]. Nonassociative [3204]. noncausal [327]. nonclassical [708, 637]. Noncollinearity [43]. Noncommutative [3272, 3383, 2800, 4269, 3388, 1090, 1365, 1373]. noncommuting [1066, 1828]. Noncomposite [3824]. nonconformal [968]. Nonconservation [3251, 3283]. Nonconservative [3815]. nonconserved [498]. nonconserved-axial [498]. Noncontextual [3984, 4013, 3762]. Nonconventional [3121]. Nondemolition [1643]. Nondispersive [2218, 304, 1794]. Nonequilibrium [3702, 948, 753]. nonet [1977]. Nonexistence [376, 879, 1067]. Nonideal [3929, 1208, 1315, 1443]. Noninertial [3002]. Noninvariant [1880]. Noninversion [2096]. Nonlinear [3106, 3871, 3383, 2450, 3946, 1112, 2441, 4709, 2276, 2425, 3445, 331, 3828, 2794, 952, 1359, 1583, 830, 1436, 62, 384, 1520, 1298, 656]. nonlinearities [1527]. nonlinearity [442]. Nonlocal [1848, 1111, 3946, 2393, 2739, 3829, 3549, 1136, 3971, 4582, 3162, 3942, 3705, 3192, 1823, 2056, 2044, 730]. Nonlocality [3768, 4000, 4480, 1281, 3239, 3592, 2098, 1282, 2524, 3483, 2323, 3243, 3362, 3470, 566, 3460, 4307, 4024, 3655, 3442, 4311, 2028, 2065, 2080, 4534, 2874, 3250, 4218, 152, 1299, 1301, 4331, 1619, 1026, 1060, 898, 2227]. nonlocalization [410]. Nonlocally [1059, 1793]. nonminimal [670, 1176]. Nonobjectivity [2402, 1154]. nonoccurrence [729]. nonparallel [1827]. Nonperturbative [2789, 546, 2835]. Nonquantum [3216]. Nonreality [4612]. nonrelativisitic [590]. Nonrelativistic [3428, 2953, 2788, 2099, 2110, 440, 507, 1535, 633, 273, 932]. Nonsense [3314]. nonseparability [1036]. Nonsignaling [4024]. nonsingular [256]. Nonsingularity [2729]. Nonspreading [332]. nonstandard [1517]. Nonsymmetric [3388]. Nontemporal [3014]. Nonuniqueness [1740]. Nonunitary [1992]. Nor [4179, 4211]. Normal [4583, 3314]. Normalized [4093, 1532]. Normed [4021]. Norton [2624, 3845]. Not-in [3825]. Notation [3554]. Note [4062, 4432, 3533, 2943, 4399, 902, 4523, 3432, 3056, 1307, 1469, 344, 1219, 1200, 617, 1305, 749, 514, 870, 3125]. Notes [3196]. nothing [1716]. Notion [2599, 4146, 4254, 4217, 4595, 3930, 3640, 862]. Notions [3794, 3615, 4198, 970]. Novel [3960, 3883, 3291, 2054, 1922]. November [2829]. Novikov [2387, 2624]. Nuclear [3324, 3701, 3141, 2342, 3019, 994, 829, 2046, 30, 287, 2023, 822, 2012, 436, 466]. nuclei [2023]. Nucleon [2846, 3334, 2339, 2372, 1622, 1509]. nucleon-monopole [1509]. nucleons [285]. Nucleus [2848, 3765]. Null [456, 3883, 2771, 3696, 3403, 696, 746, 1632, 1530, 461]. Null-Result [3696]. Null-surfaces [3403]. Number [2789, 2075, 2104, 2117, 2131, 2139, 2149, 2161, 2171, 2179, 2185, 2194, 2204, 2220, 2232, 2246, 2256, 2262, 2268, 2277, 2284, 2293, 2306, 2379, 2368, 2391, 2419, 2430, 2443, 2669, 2678, 2689, 2703, 2709, 2716, 2725, 2735, 2743, 2751, 2759, 2769, 2776, 2787, 2796, 2802, 2809, 2813, 2822, 2829, 2837, 2891, 4020, 4019, 4189, 3075, 2892, 3515, 556]. Numbers [4419, 4444, 2954, 35, 3308, 247, 314, 1571, 166, 1296, 764, 1670]. Numbers-Based [4419, 4444]. Numerical [3785, 2446]. nutshell [4665, 2767].

O [1553, 2242]. Object [4695, 3093, 4154, 4311]. objectification [1353, 1450]. Objection [4048]. objections [1721, 1949]. Objective [4695, 2878, 2629, 2780, 4391, 3976, 3205, 3700, 663, 662, 279, 1604]. Objectiveness [4393]. Objectivity [4181, 3155, 3242, 2402, 4132, 4186, 4085, 1401, 3368, 3379, 4515, 1080]. Objects [95, 3735, 4253, 3246, 4445, 2122, 2321, 1187, 1326, 1456, 1150, 1764, 1211, 2051]. Observability [3843, 49, 523]. Observable [3361, 341, 1772, 1145, 106]. Observables [3233, 4167, 3619, 4172, 2534, 3101, 3424, 2393, 877, 2250, 3196, 3235, 4711, 4166, 3365, 3697, 2741, 4080, 4184, 3837, 2411, 4170, 3181, 4105, 3092, 734, 988, 2027, 1828, 58, 96, 434, 1155, 763, 1220, 1705, 1216, 1060, 1368, 629, 1704]. Observation [2641, 2060, 2097, 3516, 1122, 4687, 2047, 65]. Observational [4093, 4416]. observations [1444]. Observatory [2831]. Observe [4323]. Observed [4733]. Observer [2715, 2921, 3827, 2614, 3954, 484, 1813]. Observers [3439, 3466, 261, 1461, 1750]. Observing [4441, 828]. obtained [1743]. obtaining [2027]. Obvious [837]. occasion [1047, 1266, 1610]. occupation [1296]. Occurrence [2623, 3132, 1603]. occurrences [1855]. Occurs [4207]. October [2075, 2194, 2204, 2293, 2306, 2822]. Octonion [1947, 314]. octonion/quaternion [314]. Octonionic [1833]. Oehme [2043, 2359]. Off [1802, 2718, 2909, 2910, 2506, 1162, 1981, 1310, 1958, 1457]. Off-mass-shell [1981]. Off-Shell [2506, 1802, 1162, 1310, 1958]. Old [2338]. Once [4734]. One [3158, 2344, 2063, 3751, 3298, 4312, 2856, 2520, 4094, 2067, 4070, 2260, 1411, 451, 2039, 378, 921, 471, 1723, 1320, 1880, 1263, 1264]. One-Dimensional [3751, 4070, 451]. One-Particle [4094]. One-Photon [2520, 1723]. One-Way [3158, 2856, 2067, 378, 921, 471, 1880]. Ones [2209]. Ongoing [2861, 3600]. Only [4718, 3012, 2947, 4242, 3346, 1397]. Onsager [948, 384]. Ontic [4734, 4353, 4564, 4570, 4285]. Ontological [3982, 3768, 4313, 3811, 4392, 3377, 4285, 3494, 3887, 3854, 4537, 1890]. Ontologists [4296]. Ontology [3665, 3888, 3498, 4315, 4493, 4718, 4519, 4547, 4497, 3776, 4089, 3693, 4137, 3770, 4610]. opacity [949]. opaqueness [316]. Open [2367, 2378, 2390, 3283, 4678, 4681, 4552, 3816, 3836, 1634, 2680, 2766, 4651, 4657, 4674, 2243]. Operates [2781]. Operation [587, 404]. Operation-valued [587]. Operational [3278, 3380, 4266, 125, 1246, 3367, 2627, 4595, 2081, 433, 261, 1234, 1224, 589]. operationalism [588]. Operationalist [4153]. Operations [2295, 3783, 4300, 2570, 2400, 629]. Operator [2374, 4219, 3934, 2163, 2779, 2191, 3447, 3490, 3395, 2110, 3225, 2008, 1232, 633, 419, 427, 1777, 1657]. Operators [3740, 2079, 3777, 3449, 4487, 3800, 3001, 3490, 2914, 4607, 3915, 1066, 624, 1573, 97, 487, 731, 1263, 1264, 1586, 1718]. Oppenheim [4009]. Oppenheimer [3753, 2445]. Opposite [2805]. Optical [3889, 3172, 2805, 890, 1383]. Optically [3741]. Optics [2945, 3172, 3560, 4613, 1015, 1794, 1018, 2389, 2242]. optimally [1985]. Optimization [2536]. Option [3512]. Oracles [4100]. Orbit [2479, 2458, 2468, 829]. Orbital [3934, 2432, 2436, 3444, 3318, 2372, 1536]. Orbital- [2436]. Orbital-Free [3444]. Orbiting [4417]. Orbits [2457, 2485, 3981, 1574]. Order [3861, 3898, 2825, 2432, 3785, 2718, 3970, 3061, 3129, 3174, 2487, 3639, 3637, 3640, 1455, 504, 32, 67, 970, 1991, 1072, 1878, 393, 1235, 278, 77, 101]. Order-Invariance [2825]. Order-Unit [3174]. Ordered [2436, 181, 391, 377, 420]. Ordering [2570, 3314, 458]. orders [32]. Ordinary [4350]. Organization [4597, 962]. Organizers [3716]. Orhan [1499, 2076]. Orientability [4556]. Origin [3420, 4410, 2584, 3276, 2258, 2702, 2949, 3077, 3795, 3584, 3932, 2663, 2804, 3629, 2420, 4600, 2832, 864, 966, 230, 249, 669, 700, 1516, 709, 2049, 89]. Original [2983]. Origins [959, 1438]. Ortho [3013, 493]. Ortho- [3013]. ortho-chronous [493]. orthoalgebras [1247, 1233]. Orthocomplementation [3380]. orthodox [663, 662]. Orthogonal [2529]. Orthogonality [3229, 4702]. ortholattices [1694]. Orthomodular [3835, 1247, 1227, 1348]. Orthospaces [814]. orthosymmetric [1694]. Oscillating [3287, 492]. Oscillations [3468, 3656, 3669, 2163, 3280, 3967, 2343, 2453]. Oscillator [3763, 4076, 3699, 2672, 3078, 3089, 3133, 3219, 3797, 3752, 3919, 3708, 4456, 3828, 3654, 1811, 1812, 606, 610, 616, 1073, 1502]. oscillator-zero-point [606, 610, 616]. oscillators [1699]. Otfried [2899]. Other [4239, 2248, 4491, 3610, 2613, 3692]. ‘Oumuamua [4510]. outcome [1512]. Outcomes [3279, 4404, 3567]. outer [2011]. Overlapping [4202]. Overlaps [3289]. Oversight [4102]. Overview [2495, 4621]. Oxford [2388, 2539, 2677, 2617, 2766, 2315].

P [1954, 2344, 2387, 2283, 2561, 4042, 2766, 3814, 3113]. packet [51, 1719, 803, 1453, 304, 2053]. Packets [3258, 1033, 1040, 332]. Page [4526]. pages [2600, 2429]. Pair [2115, 973]. Pairing [2446]. Pairs [2568, 2552, 2985, 3929, 508, 201]. Pais [2388, 2499, 2103]. Palle [2967]. Panorama [4373]. Paper [2824, 1022, 613, 93, 4462, 902]. Papers [2294, 2305, 2347, 2477, 850, 901, 1616, 1630, 1762, 1771, 1781, 1978, 2006, 465]. Para [3013, 440, 507, 498]. Para-helium [3013]. para-Lorentzian [440, 498]. para-Maxwellian [507]. Paraconsistent [1143, 3657]. Paradigm [2894, 4128, 3650]. Paradox [3048, 1826, 4514, 3214, 4223, 3807, 3483, 3865, 2707, 2248, 4302, 4082, 3803, 1130, 2977, 2383, 3706, 4360, 4706, 978, 1333, 448, 566, 804, 1240, 1062, 1237, 81, 490, 933, 1897, 1179, 898, 4187, 2358, 4480, 3117, 3384, 3017, 3069, 4242, 3304, 2989, 3555, 2976, 3472, 794, 1190, 4324, 4691, 2067, 4206, 3148, 2810]. Paradoxes [3423, 4410, 4732, 4315, 4494, 525, 1404, 1930, 647, 691, 1121, 942, 103, 729]. Parallel [4324, 4137]. parallelism [735]. paramagnetic [2005]. Parameter [4337, 3361, 3394, 1512, 4498, 1042, 573, 1364, 1851]. Parameter-Dependent [4498]. Parameterized [3988]. Parameters [2693, 4192, 4429, 3503, 323]. Parametric [3154]. Parametrization [3990, 4112]. Parametrized [2156, 2152, 2302, 295, 1204, 1637]. Parity [2833, 280, 386, 4150, 385, 3450, 3500, 3759, 530]. Part [3436, 31, 32, 1281, 1194, 1195, 2815, 1282, 197, 209, 233, 241, 3439, 153, 113, 127, 250, 1365, 1373, 1598, 1599, 4152, 220, 251, 1395, 19, 68, 76, 160, 4457, 4466]. Partial [2941, 469, 1679, 3715, 2352, 3465]. partially [420]. Particle [3048, 2789, 3863, 2590, 2297, 3665, 4555, 4138, 1466, 3088, 4393, 3343, 2168, 3166, 3781, 3962, 3075, 2720, 4419, 4444, 4251, 3905, 4325, 3937, 2136, 2554, 2555, 2211, 41, 3508, 3396, 4328, 3889, 4302, 3896, 3399, 4094, 2835, 2892, 4459, 2088, 2577, 3708, 128, 4273, 4427, 2118, 2506, 3037, 3591, 4263, 2630, 2788, 4265, 2586, 3389, 2764, 4180, 2080, 2363, 3866, 3977, 2109, 4260, 4060, 2290, 2320, 3503, 1106, 1592, 1474, 1400, 1059, 1465, 443, 1475, 1900, 521, 1382, 1876, 1073, 4646, 4672, 659, 124, 2000, 987, 1502, 983, 184, 146, 1613, 1699, 1729, 1294]. particle [1891, 1700, 1707, 537, 444, 201, 346, 38, 1600, 625, 505, 664, 2522]. particle-antiparticle [1502, 201, 505]. particle-family [664]. Particle-like [1466]. Particle-Type [2363]. particle-wave [987]. Particles [4531, 609, 3575, 4385, 4322, 226, 3936, 3140, 2466, 2871, 2693, 3583, 3751, 2310, 3701, 4133, 2447, 3183, 3458, 3992, 4364, 2198, 4718, 4504, 2299, 2554, 2211, 3900, 2276, 3303, 3111, 3254, 1958, 2969, 2203, 4445, 2695, 3179, 3868, 2926, 2953, 1294, 2110, 2913, 2146, 2613, 952, 1041, 530, 1765, 1177, 1730, 1025, 1448, 843, 389, 528, 1614, 1713, 357, 448, 527, 1319, 1987, 759, 855, 1374, 1424, 483, 488, 415, 1613, 1960, 360, 445, 1856, 285, 2051, 1520, 874, 1107, 1175, 14, 8, 1829, 712, 1202, 43, 768, 466]. Particular [3357, 4627, 755]. Particulate [4473]. Partitions [4527, 3877]. Parton [350, 2337]. Parts [3066, 3867, 3735, 1853, 1922]. Passing [4383]. Past [3278, 4552, 4509, 2579, 4728, 4360, 4372, 1128, 146, 1154]. Patching [3946]. Path [2147, 3381, 4259, 1742, 2447, 4264, 3480, 4026, 2066, 2448, 4286, 4422, 4445, 1699, 4350, 4586, 3515, 3933, 4436, 2109, 4571, 1382, 1552, 1562, 1725]. Pathologies [2574]. Paths [3062, 1577]. Pattern [3961]. Patterns [4407]. Paul [529, 551, 563, 545, 4108, 4163, 727]. Pauli [3316, 3320, 1508, 3813, 3409, 198, 3576, 1802, 3331, 3678, 4048, 598, 494, 704, 1776, 4488, 1661]. Pauli-Exclusion-Principle [3576]. Paulian [3324]. Paulions [3304]. PBR [4612, 4718, 4397, 4140, 3692]. Peaceful [4397]. peculiarities [260]. Peierls [766]. Pekka [3247, 3228]. Pendulum [2063]. penetration [867]. Penrose [3341, 4357, 4523]. Pentagrams [3423]. Percepts [3496]. Peres [2961, 3341, 4348, 2805, 3500]. Perfect [4480, 3531, 4172, 3476, 3344, 4319, 1407, 4462]. Perfectly [2421]. Perform [2840]. performing [1743]. Perihelion [2151]. Periodic [2457, 2485, 2468, 578]. Permittivity [4243]. Permutation [2706]. Perseus [2345]. Persistence [2871]. Personal [4572, 2077, 871, 870, 142]. Perspectival [2584, 4522, 4290]. Perspectivalism [4173]. Perspective [4191, 3334, 3242, 3771, 4153, 4017, 4507, 3286, 4503, 3389, 4319, 3608, 4361, 3551, 2811, 90, 1631, 350, 1414, 4462]. Perspectives [3324, 3785, 1989, 4230, 3840, 4621, 2495]. Perturbations [2125, 4124, 4144, 4313, 3058, 4141, 2354, 1962]. Perturbative [3406, 2339, 2892, 4160]. Perturbed [3763, 2461, 3274, 4520, 610, 616, 1453]. Peter [760, 752, 1148, 1139, 3347, 3350, 3348, 3805, 742]. Petiau [3936, 2198, 3620]. Petruccione [2766]. Phase [2207, 3763, 2362, 2550, 2551, 3157, 2901, 3383, 3577, 3775, 3898, 3427, 2198, 4739, 2856, 3879, 2596, 1382, 3830, 3752, 4323, 4405, 3060, 2680, 4367, 2463, 4093, 3601, 2218, 2382, 2536, 3193, 4715, 2973, 2573, 1999, 2863, 3738, 3516, 2440, 4471, 4470, 1945, 438, 60, 668, 708, 875, 961, 341, 534, 1715, 1063, 1133, 1600]. Phase-covariant [3427]. Phase-like [3193]. Phase-Shift [3752]. Phase-Space [4093, 2863, 1382, 60, 708, 961, 534]. Phase-Spaces [3577]. Phases [2465, 3517, 1514, 1501]. Phenomena [2187, 3149, 3701, 2259, 3172, 3519, 4706, 876, 418, 375, 1056, 1039, 1511, 1660, 1751]. Phenomenological [3077, 3161]. Phenomenology [3665, 2169, 4573, 4704, 1971, 2048, 203, 2429]. Phenomenon [3537, 3096, 4490, 3264, 3386, 4731, 159]. Phi [1824]. Phi-symmetric [1824]. Phillips [2170, 2350, 1680]. Philosopher [1139, 4690, 1434, 3387]. Philosopher-physicist [1139]. Philosophical [2407, 4191, 3805, 2003, 26]. philosophique [2184]. Philosophy [1905, 3347, 3377, 3395, 3161, 4074, 2498, 4384, 3840, 4687, 772, 66, 1392, 1023, 1021, 1393, 1173, 4635, 4074, 2641, 2365]. Phonon [2978, 934]. Phonons [3273, 3352, 483]. Photo [519, 544, 576, 585, 551, 563, 684, 699, 752, 1004, 1008, 976, 1058, 1071, 1075, 1024, 1148, 1238, 1249, 1275, 1399, 1441, 1452, 1470, 1553]. Photoeffect [2189, 2188, 2385]. Photoionization [2099]. Photon [4006, 4169, 2520, 3713, 2264, 2718, 2000, 2453, 4298, 1993, 50, 548, 1164, 1475, 1936, 126, 1908, 1723, 987, 288, 455, 977, 231, 454]. Photonic [3744]. Photonics [3745]. Photons [4705, 2710, 3273, 3352, 1693, 285]. Photos [835, 842, 873]. Phys [3113, 2688, 3206]. Physical [129, 2841, 4178, 4244, 2633, 3950, 3300, 4467, 4065, 3219, 2893, 2860, 4721, 3777, 2916, 3937, 4477, 2326, 4250, 4508, 382, 4091, 396, 501, 2765, 3007, 3574, 4599, 460, 4423, 27, 4595, 3911, 3944, 1172, 4145, 4581, 4319, 4267, 590, 3838, 3016, 525, 1675, 1683, 102, 940, 67, 1128, 1142, 1564, 416, 549, 1750, 1793, 918, 1144, 604, 860, 1870, 521, 579, 69, 134, 225, 941, 981, 237, 586, 4462, 1563, 1740, 278, 1120, 1188, 1949, 145, 1259, 1415, 757, 284, 1855, 1044, 1217, 1905]. physically [1554, 1844, 106]. Physicist [1265, 1434, 3387, 4366, 4336, 510, 2103, 1139, 4690, 95, 4623, 4476]. Physicists [4660, 4680, 4664, 44, 1865, 4669, 2388, 2128, 2808, 2820]. Physics [4380, 1138, 1147, 1185, 1193, 1498, 1615, 1990, 2075, 2104, 2117, 2131, 2139, 2149, 2161, 2171, 2179, 2185, 2194, 2204, 2220, 2232, 2246, 2256, 2262, 2268, 2277, 2284, 2293, 2306, 2377, 2346, 2379, 2403, 2387, 2368, 2391, 2419, 2430, 2443, 2455, 2467, 2476, 2488, 2500, 2509, 2517, 2523, 2528, 2533, 2542, 2549, 2559, 2567, 2575, 2583, 2592, 2601, 2610, 2618, 2625, 2632, 2643, 2652, 2660, 2669, 2678, 2689, 2703, 2709, 2716, 2725, 2735, 2743, 2751, 2759, 2769, 2776, 2787, 2796, 2802, 2809, 2813, 2822]. Physics [2829, 2837, 2866, 2875, 3452, 3716, 2582, 2571, 3272, 3761, 2074, 2602, 2937, 4291, 3140, 3995, 4593, 3547, 4189, 3361, 4259, 4712, 2620, 3347, 3348, 3805, 4467, 2405, 2282, 2895, 2421, 3781, 3947, 4552, 3561, 3141, 3774, 3769, 2963, 4275, 4276, 3375, 3785, 3101, 2845, 28, 3087, 2495, 1124, 3451, 2807, 4246, 2494, 4328, 2839, 577, 2003, 4120, 3097, 3137, 3392, 4405, 3357, 2094, 3445, 4553, 2821, 4475, 3246, 3493, 2616, 3110, 4468, 2785, 3212, 4199, 2727, 2954, 3426, 4311, 3855, 3019, 3395, 3519, 3965, 4208, 3825, 4074, 4177, 2734, 2127, 4159, 3145, 539, 3368]. Physics [4384, 4655, 3840, 3872, 2566, 3374, 6, 3640, 2365, 3717, 3495, 3221, 1570, 1608, 1640, 1041, 1631, 31, 1696, 32, 67, 1886, 1686, 1784, 21, 772, 862, 1548, 247, 1757, 26, 1921, 969, 830, 508, 1001, 4621, 1392, 1838, 1023, 99, 861, 4646, 4672, 1314, 1408, 1489, 1229, 59, 86, 124, 373, 1335, 49, 19, 1435, 863, 360, 54, 4685, 4686, 4689, 324, 1283, 557, 463, 1209, 444, 160, 138, 1196, 1808, 826, 4635, 2002, 178, 1422, 1423, 4637, 4642, 1087, 141, 656, 1842, 1849, 2366, 1863, 1903, 2590, 2600, 2795, 2474, 2794, 2498, 2243]. physics-Axiomatics [86]. Pickover [2315]. Picture [4433, 4036, 2670, 1478, 3460, 2754]. Pictures [2846, 3760, 355]. Pienaar [4489]. Pierre [1736, 1710]. Pilot [4538, 3910, 4572, 4579, 3956, 2574, 4530, 3289, 4534, 4092, 516, 517, 4610]. Pilot-Wave [4538, 3910, 4579, 3956, 2574, 4530, 4534, 4092, 4610]. piloted [1298]. Pion [2846, 2339]. Pioneer [3052, 4123, 2854, 2973, 3077]. Piron [1675, 1642, 1677, 1686, 2414, 617, 618]. Pistons [3118]. Pitowsky [1174, 3974]. place [1868]. Planary [1911]. Planck [3702, 4337, 4288, 1734, 1910, 1775, 235, 126, 4331, 2192]. Planckian [4503, 4209]. planckions [712, 1206]. Plane [3315]. planes [1318]. Planetary [3094]. Plasma [3472]. Plates [2532]. Platforms [2174, 2627]. Plato [1875]. Players [3431]. Plenum [2495]. plus [1675]. Podolsky [978, 3149, 1051, 804, 2422, 457, 783, 3696, 1076, 933, 437, 3147, 784, 898, 114, 2883]. Poems [2750]. Poincaré [1810, 398, 1959, 4180, 4384]. Point [2550, 3342, 3547, 3927, 3672, 2749, 3797, 3827, 4547, 2881, 2192, 2926, 3883, 3448, 2492, 4644, 1108, 2966, 95, 604, 606, 610, 605, 616, 1027, 761, 190, 768, 3269]. point-like [604]. Pointers [4693]. Pointillisme [4591]. Points [3464]. Poisson [1582, 4426, 4271]. Poland [2378]. Polar [4375]. Polarisation [4157]. polarity [278]. Polarizable [2597]. Polarizable-Vacuum [2597]. Polarization [2113, 4243]. Polarized [2604]. polaron [813]. pole [1532, 107]. Pollifrone [2102]. Polology [2846]. Polydimensional [2519]. polynomial [1527]. Polyparticles [2812]. Pontryagin [1530]. Popov [3929]. Popper [1399, 502, 1468, 1392, 2296, 1393, 2589, 1394]. Popperian [1395, 1409]. population [1284]. Portrait [2388]. Portraits [2573]. posed [151, 91]. Poset [2417]. Posets [2416]. Position [4604, 4156, 3252, 3092, 292, 427, 1095, 1655, 1644, 394, 768]. Positive [2665, 2779, 1853, 3677, 474, 1285, 428]. Positive- [1853]. positivism [1405]. positivity [475]. possessing [77]. Possibilist [3581]. Possibilistic [3555]. Possibility [3861, 4250, 4323, 3288, 3734, 635, 1035, 60, 1160, 1986, 63]. Possible [2208, 3420, 1142, 3380, 4251, 4714, 3660, 2413, 3772, 3146, 4092, 1600, 1464, 611, 165, 791, 446, 1189, 89, 1772, 412]. Possibly [3366]. Post [2298, 2557, 2870, 2496, 1218, 295, 1344, 386]. Post-Decoherence [2298]. Post-Newtonian [2557, 295]. Post-Selected [2870, 1218, 1344]. posterior [702]. Posters [3747]. Postponing [3278]. postscript [512]. Postselected [3531]. Postulate [4255, 4387, 4603, 433, 71, 1278, 1440, 1931, 1180, 918, 4331, 384, 131, 892]. Postulates [1656, 252, 167, 406, 197]. postulational [58]. pot [227]. Potential [3757, 3674, 2125, 3072, 2330, 4544, 4594, 3187, 3832, 2381, 2588, 3620, 3862, 3922, 3786, 2525, 3063, 4367, 2914, 4540, 3335, 2140, 4402, 2842, 2852, 711, 175, 1541, 630, 451, 1884, 1051, 53, 110, 1453, 2022]. potentialities [920]. Potentiality [4175, 80]. Potentials [3340, 3200, 2457, 102, 1883, 1529, 205, 399, 1313]. POV [3234]. POVM [4019, 3464]. POVMs [3233]. Power [4521, 1867]. Poynting [3892, 4396]. Pozzi [2551]. pp [2344, 2345, 2376, 2590, 316, 2641, 2786, 2783, 2708, 2609, 2795, 2784, 2836, 2767, 2495, 2541, 2494, 2565, 2631, 2899, 2558, 2903, 2539, 2821, 2677, 2316, 3010, 2540, 2808, 2794, 3011, 2498, 2617, 2820, 2496, 2687, 2766, 2566, 2315, 2624, 2819, 2967]. pp. [2388, 2389]. Practical [3662, 4578, 2005]. Practice [4191]. Pragmatic [1195]. Pragmatist [3178, 4559, 3602]. Praise [3188, 4374]. Pre [2416, 2159, 3286, 2870, 1218, 1344, 2158, 2160]. Pre- [2870, 1218, 1344]. Pre-BZ [2416]. Pre-geometric [3286]. Pre-Maxwell [2159, 2158, 2160]. precede [327]. Precession [2713, 3501, 2151, 1690, 2002]. precessional [1091]. Precise [3781, 2954, 441, 568]. Precisely [2814]. Precision [3756]. Precocious [2351]. Predator [4611]. predicament [278]. Predictable [2814]. predicted [196]. Prediction [2337, 3146, 1312, 2842]. Predictions [4298, 1383, 783]. Predictive [3928]. Preface [2690, 2392, 2501, 3716, 3408, 3563, 3769, 2934, 4306, 2644, 2717, 2904, 3057, 3393, 3556, 3818, 4304, 4190, 2]. Preferred [2514, 1823, 1739, 158, 507]. Preliminaries [307]. Preliminary [825, 1616, 1630, 216]. prematter [1424]. Prentice [2475]. Preparation [3768, 3921, 1435, 4486, 783, 2318, 586, 1644]. Preparation-limited [783]. preparations [1686]. Prequantum [3413, 3557]. Presence [3807, 2736, 3620, 2020, 1531]. Present [4023, 4254, 196, 1414]. present-day [1414]. presentation [505]. Presented [3747]. Presentism [4383, 4211]. Presentist [4518]. Preservation [2979]. Preserve [3650]. preserved [1936]. preserving [1992]. prespace [1053]. Press [2375, 2376, 2388, 2389, 2590, 2786, 2795, 2784, 2836, 2499, 2767, 2541, 2565, 2539, 2677, 3010, 2540, 2808, 2498, 2617, 2820, 2766, 2315]. Pressure [4319, 4462]. Prevent [3280]. previous [901]. Prey [4611]. Price [2624]. Prigogine [976, 970, 958]. Primacy [3494]. Primary [2564]. Primas [475]. Prime [3515]. Primitive [4547, 2569]. Primitivism [4539]. Princeton [2499, 2767, 3010]. principal [2044]. Principle [2679, 4555, 2074, 2778, 3959, 2981, 2405, 2386, 4342, 2920, 3802, 3123, 3576, 4215, 3331, 3796, 2649, 3956, 2288, 3032, 3622, 2844, 3841, 4224, 3525, 2214, 3889, 2929, 2296, 2962, 3578, 3901, 3773, 2572, 3853, 3855, 2569, 2547, 2591, 2348, 4429, 4087, 2902, 3917, 2530, 3963, 4282, 2324, 4272, 2230, 3209, 92, 570, 179, 441, 569, 823, 888, 1157, 1633, 1643, 329, 1836, 1124, 156, 857, 257, 166, 1213, 583, 1462, 831, 269, 319, 49, 312, 704, 2026, 949, 2042, 368, 72, 74, 540, 790, 1201, 568, 3316, 3355, 3320, 402, 3409, 247, 3749, 3678, 322]. Principle [2695, 37, 1458, 1988, 2071]. Principles [4613, 3317, 2825, 2949, 3960, 3853, 4270, 2849, 4595, 2706, 2992, 97, 1476, 1669, 819, 10, 839, 959, 2389]. Printing [2899]. Prior [4477, 710, 660]. Priori [4249, 1461]. prism [1900]. privileged [599]. Probabilistic [3687, 4100, 3014, 4219, 2310, 4266, 2929, 3882, 3596, 1979, 1931, 86, 124, 429, 522, 564, 345, 949]. probabilistic-informational [949]. Probabilistics [500]. Probabilities [2395, 3238, 1268, 4135, 2737, 2793, 2929, 3819, 3001, 3142, 2199, 2862, 4093, 2976, 4502, 1066, 1471, 1450, 1939, 702, 1398, 383, 1395, 1409, 1355, 1397, 196]. Probability [3022, 4151, 3661, 3931, 4435, 2399, 3339, 2770, 3224, 3884, 3707, 2072, 3108, 2410, 2715, 3559, 4661, 2058, 4711, 3414, 2088, 3205, 4729, 3179, 4311, 3464, 2670, 902, 4597, 2986, 4106, 3410, 4471, 1683, 1090, 1448, 258, 1065, 1473, 4624, 1461, 1243, 1757, 468, 648, 889, 680, 1224, 1097, 579, 86, 522, 564, 586, 1847, 1016, 1856, 1063, 1412, 849, 1855, 367, 2784]. Probe [4154]. Probing [3781, 3818, 4304]. Problem [4592, 2208, 4062, 2125, 2461, 2312, 3927, 2884, 2806, 2999, 3113, 3484, 3605, 2338, 3087, 3279, 2818, 2526, 2596, 2326, 3438, 3679, 4160, 3667, 4392, 4194, 3053, 2202, 4112, 2422, 2817, 3036, 2159, 4118, 4317, 4216, 3061, 4152, 4332, 3493, 3743, 3574, 2988, 4530, 3973, 4338, 2804, 4347, 4271, 3608, 2328, 2151, 4303, 2114, 3742, 2280, 4452, 3913, 2944, 2011, 1508, 887, 1048, 1353, 107, 781, 1123, 1766, 151, 1936, 121, 1160, 225, 91, 581, 1725, 1407, 1676, 1352, 2017, 1028, 1843, 863, 2030, 426, 2025, 1230, 1957, 625, 960, 898, 36, 2423]. Problematic [4134, 2606]. Problems [1842, 1849, 4007, 3805, 611, 613, 2686, 2474, 3451, 3104, 4005, 614, 4558, 2454, 612, 1863, 1903, 4630, 4638, 4643, 4656, 1541, 26, 1225, 846, 1442, 372]. Proca [2374, 1707, 3753]. Procedure [4209, 2027]. Procedures [4261, 164]. Proceedings [3716, 3393, 3717]. Process [3431, 2358, 3874, 2393, 3413, 4475, 4316, 4389, 2089, 4336, 1284, 1122, 1966, 1420, 188, 4476, 270, 4366]. Processes [3143, 2111, 3324, 2144, 3636, 4478, 3826, 3218, 4320, 587, 1025, 1426, 1134, 1606, 1279, 580, 767, 1549, 62, 384, 534, 669, 1277, 169, 1728, 116, 112, 182, 123]. Processualism [4475]. Produce [3066, 3740, 1125]. Produced [3067]. Product [2399, 4433, 1832, 1947, 1261, 1844, 1223]. Production [2115, 2339, 3922, 3928, 1122]. Products [4535, 1556, 1602]. Professor [838]. Program [2311, 1726, 1996, 1711, 1017]. Programme [4526, 940, 2886]. Progress [4007, 1158]. progressive [1983]. Projection [4219, 4694, 1278, 1440, 1180, 832]. projections [1095, 1096]. Projective [3947, 1053, 3838, 1177, 533, 755]. Projectors [2529]. Proof [2029, 3359, 3751, 3470, 1443, 2731, 2925, 3895, 4090, 4142, 4254, 4507, 2953, 337, 730, 3831, 4049, 2229, 2038, 887, 1299, 2039]. Proofs [3341, 3762, 3450, 3500, 3759]. propagating [1882]. Propagation [2373, 3959, 3649, 4301, 4585, 2859, 4613, 471]. Propagator [4445, 4350, 538]. Propagators [4454, 598]. Propensities [1416, 1395, 1409, 1396]. Propensity [3931, 1757, 4358]. Proper [2164, 4062, 4419, 4444, 2167, 2765, 3604, 2527, 1366, 3282]. Proper-Time [2167, 2527]. Properties [3433, 2396, 3028, 4244, 3301, 4239, 3816, 3777, 3790, 2322, 3222, 4112, 4301, 3510, 3207, 4129, 4155, 3684, 4497, 589, 3354, 3184, 3624, 2321, 4180, 3755, 1525, 1793, 1146, 580, 995, 981, 733, 496, 47, 731, 1019, 278, 977, 182, 1646]. Property [2704, 2417, 3580, 2272, 3389, 3269, 1246, 1172, 1226, 1322, 892]. Proportionality [3403]. Proposal [2712, 4724, 3693, 1074]. Proposals [2783, 4654, 4662]. Proposed [4562, 501, 4094, 3601, 396, 1897]. proposition [1684, 1231]. Propositions [3295, 589]. Prospects [3514, 2926]. Protective [4559, 1837]. Protocols [3821]. Proton [4576]. prove [992]. Proved [3678]. Provide [3256, 4324]. Provides [4065]. Proving [4438]. provoking [1050]. Pseudo [2939, 3371, 3461, 3683, 1133]. Pseudo-classical [1133]. Pseudo-Telepathy [2939]. Pseudoharmonic [3620]. Pseudoscalar [2580]. psychokinetic [486]. Psychological [2259]. Psychology [3860, 1406]. PT [4030]. Publications [1248, 2757, 2668, 2961, 3247, 3633, 3350, 2987, 1639, 1736]. published [362]. Publishers [2708, 2609, 2600, 2474, 2497, 2494, 2903, 2785, 2496, 2429]. Publishing [2495]. Pulmannová. [2497]. pulsar [492]. Pulse [177]. Pulses [3744]. Pure [51, 2620, 2946, 3401, 3236, 574, 347, 636]. Purely [2310, 3815, 3215, 1318]. Purification [4315]. purposes [4578]. Pursuit [3553, 1675]. Puzzle [531, 4730, 4296]. Puzzling [2895, 3848, 1009]. PV [2597]. Pythagoras [2143]. Pythagorean [2522, 2143].

Q [3186]. Q-spaces [3186]. QBism [4256, 4418, 4314, 4411]. QBism’s [4312]. QBist [3433, 3822]. QCD [2568, 2552, 4545, 1948]. QED [2164, 2774, 1532, 2355, 2072, 2326, 1611, 2425, 3953, 1546, 2328, 2114]. QFT [3468, 4507, 2926]. qHV [3829]. QM [3424, 2803, 4140, 4229]. QRPA [2024]. quadratic [570, 53, 1231]. Quadrature [3365]. Quadrupole [4260]. Qualia [4103]. Qualitative [604, 2337]. Quandaries [4412]. Quanta [2192, 3442, 131, 1294, 2889, 4103]. quantal [1684, 241, 1087]. quantique [2184]. Quantising [2872]. Quantitative [4065]. Quantitatively [4024]. Quantities [3017, 4058, 4081, 283, 1142, 284]. Quantity [3193]. Quantization [815, 3081, 3301, 1773, 1807, 4528, 4598, 3599, 2473, 1966, 3990, 3039, 4381, 4379, 593, 3478, 20, 3034, 2162, 2066, 3446, 2264, 1574, 3262, 2468, 2203, 3532, 2975, 4066, 4430, 2109, 1922, 2842, 3760, 3968, 4375, 107, 654, 1326, 1685, 1720, 94, 64, 1160, 1552, 1562, 909, 273, 2037, 322, 1352, 1847, 1415, 489, 481]. quantize [2054]. Quantized [1456, 3058, 3227, 193]. Quantizing [4216, 263]. Quantum [2898, 2113, 4542, 4526, 1689, 1683, 2398, 92, 102, 4483, 4228, 3198, 3197, 4580, 2207, 3199, 4506, 2200, 2208, 2535, 3022, 1616, 1630, 1842, 1849, 2344, 2345, 2375, 2392, 3716, 3255, 4548, 3431, 3860, 3200, 2571, 2111, 3888, 2841, 2894, 2865, 4369, 2479, 3910, 3178, 3564, 3420, 3931, 3982, 4280, 3486, 3704, 4095, 4262, 4337, 1271, 1733, 3861, 3699, 4181, 3761, 2935, 3718, 2307, 2890, 35, 3993, 4016, 4410, 2187, 2330, 2584, 3799, 3921, 948, 2458, 2222, 2877, 4611, 2484, 3634, 3867, 3847, 3852, 4022, 1020, 4544, 4594, 4448, 2633, 2938, 3916, 4203, 32]. Quantum [67, 2778, 3627, 4127, 4104, 1347, 3995, 4143, 4330, 4014, 4340, 4593, 4000, 3663, 2082, 2063, 4101, 2745, 2939, 3300, 4678, 4681, 3412, 2399, 649, 3238, 3361, 2746, 1440, 2489, 2885, 4259, 4535, 4712, 3343, 4400, 982, 3689, 3794, 3240, 3348, 3360, 4485, 2981, 3511, 3498, 2209, 4359, 2449, 4734, 3229, 4167, 4165, 2864, 4398, 4717, 4115, 3241, 3949, 2999, 2974, 4038, 3899, 4065, 4205, 3383, 2319, 3155, 2282, 2416, 2770, 3728, 4511, 4158, 3520, 3484, 2895, 3681, 3351, 2744, 2135, 2295, 2485, 3308, 3224, 3758, 3418, 3744, 2853, 4562, 3842, 4343, 1726, 353, 3894, 4315, 3380]. Quantum [3231, 3621, 3869, 3149, 213, 2265, 3417, 3427, 3947, 4083, 4489, 3789, 3327, 2747, 1110, 4624, 3857, 1180, 2708, 2878, 3183, 3242, 3458, 3774, 4173, 4522, 574, 3075, 3186, 4493, 3839, 2801, 4610, 4579, 4725, 1713, 2621, 4116, 3371, 3619, 3707, 3823, 4128, 3703, 3832, 1555, 4527, 4284, 4313, 2600, 4223, 4567, 3080, 3085, 3406, 3391, 2563, 357, 2381, 2917, 3394, 3885, 3302, 4364, 4606, 1383, 2686, 2857, 859, 4154, 3280, 3948, 4478, 3529, 1523, 3100, 26, 3771, 4437, 2845, 2587, 3874, 3777, 4064, 3483, 2465, 3415, 4306, 4618, 42, 2435, 2593, 2165, 3108, 3087, 3279]. Quantum [4451, 4541, 4559, 4714, 3796, 4097, 2402, 2629, 2780, 2818, 2996, 4504, 2323, 4221, 4135, 2520, 3428, 4666, 3362, 3437, 3660, 2594, 2595, 3590, 3613, 4149, 4517, 4463, 4274, 3956, 4325, 3293, 3879, 2299, 2596, 3323, 4381, 4379, 4391, 3833, 3645, 333, 2329, 2474, 2495, 4616, 2737, 4671, 3441, 3552, 3731, 3858, 2622, 3588, 3710, 3034, 4195, 1052, 1224, 1279, 1446, 2283, 2322, 2497, 2541, 3006, 3232, 3373, 4174, 4163, 3798, 3425, 3012, 3222, 3258, 3438, 1097, 1611, 3274, 3449, 4355, 4327, 2715, 4716, 264, 3538, 3281, 4449, 3455, 1308, 3430, 3602, 3976, 4132, 4186, 3196, 3695, 2471]. Quantum [3170, 3841, 3177, 4560, 4723, 3190, 4353, 3594, 703, 3277, 3565, 3248, 3830, 2906, 3033, 3294, 3521, 1421, 3525, 4310, 2119, 2868, 3786, 4557, 4378, 3517, 2170, 2162, 3068, 3598, 320, 4564, 2979, 3223, 3439, 3764, 2066, 3925, 2631, 3422, 2413, 3322, 3321, 3571, 3363, 3411, 3559, 3735, 3826, 4175, 2084, 3658, 4156, 2905, 2126, 2202, 1241, 3889, 2058, 4141, 1109, 4584, 3460, 4179, 4242, 4370, 156, 4433, 3720, 4114, 3536, 3649, 4301, 2259, 2793, 2929, 3254, 3336, 3556, 3713, 3819, 3818, 3850, 3924, 4017, 4023, 4304, 4307, 4331, 4711, 3364, 3297, 2697, 3675]. Quantum [3783, 4439, 4166, 2899, 995, 3387, 3726, 2480, 4015, 3810, 3960, 3076, 3999, 3168, 3896, 4323, 2558, 3772, 3834, 2123, 2777, 2671, 2817, 2962, 3036, 3399, 3377, 3137, 4091, 3510, 2160, 2401, 427, 3207, 4441, 4693, 501, 3443, 3970, 2215, 2739, 3791, 4085, 4098, 4094, 993, 3059, 2835, 3604, 3603, 4491, 3008, 3906, 2950, 4058, 4168, 4570, 3932, 3366, 4737, 3385, 2059, 2781, 4185, 3270, 920, 4090, 4142, 3414, 3722, 2994, 3319, 2762, 2830, 2273, 4146, 4574, 700, 4152, 4010, 2022, 3249, 4155, 4367, 4332, 2570, 2577, 584, 4518, 2581, 3957, 1516, 2561, 2763, 3205, 3732, 2102]. Quantum [2463, 2176, 2300, 2681, 3800, 4654, 4662, 3046, 3743, 3001, 3918, 3961, 2216, 447, 3110, 4316, 4389, 4468, 3142, 3174, 3204, 3655, 3767, 3808, 3983, 3109, 1847, 2199, 3467, 3666, 2266, 3586, 2924, 3677, 3741, 3700, 3694, 3926, 2803, 3868, 4333, 4397, 4558, 3958, 3212, 2907, 3435, 3539, 3266, 4093, 2453, 3465, 1296, 3494, 2157, 3165, 3601, 3697, 3727, 3746, 3532, 4035, 4341, 2953, 1728, 3887, 4217, 4412, 4411, 2887, 3023, 2748, 2954, 3426, 3853, 3854, 4025, 4311, 4270, 2536, 4492, 3662, 1619, 4617, 4456, 3952, 2773, 2791, 1692, 2560, 2400, 474, 2804, 2087, 2914, 3074]. Quantum [3395, 3419, 3560, 3519, 3193, 3471, 4715, 4036, 4626, 2569, 3367, 4042, 4263, 3354, 2788, 2112, 3776, 1337, 4344, 4208, 3464, 4366, 4336, 874, 3161, 2614, 2132, 4089, 2626, 2670, 3005, 3466, 4710, 2586, 4071, 2070, 4321, 2110, 2982, 1186, 2154, 2515, 4052, 243, 268, 4298, 3848, 4569, 3911, 4147, 4446, 3094, 2738, 1322, 4434, 3401, 4294, 4413, 4704, 2291, 3570, 3569, 2764, 4054, 2081, 4159, 4390, 4540, 3020, 3988, 2794, 4581, 3844, 2173, 819, 4719, 3637, 3064, 3491, 3975, 3589, 3917, 4582, 3558, 4597, 4607, 3654, 2530, 2529, 3762, 3562, 4055, 2496, 3368, 3859]. Quantum [1026, 1396, 2531, 3550, 2647, 934, 2585, 3513, 4634, 4641, 4677, 2350, 2782, 3382, 3148, 4176, 2004, 3307, 2137, 3866, 3977, 3676, 3880, 3991, 4031, 919, 1230, 2064, 2420, 2576, 4431, 3878, 4523, 4529, 4136, 1897, 4361, 2475, 2364, 2958, 3933, 4457, 4466, 4139, 2219, 4700, 4106, 2992, 4282, 4537, 3421, 4257, 2766, 2249, 2238, 2873, 4303, 4554, 4045, 2603, 3742, 4084, 4220, 3153, 3164, 3693, 3872, 3345, 4576, 3904, 4060, 4524, 4575, 3429, 2667, 3079, 3086, 3942, 3551, 2182, 1068, 4452, 3432, 3856, 3738, 2290, 4472, 4401, 3042, 3082, 3152, 3299, 4474, 4551]. Quantum [2492, 849, 2959, 3507, 3626, 3668, 3680, 3755, 4210, 3237, 394, 2811, 3670, 4471, 4470, 3474, 4105, 2653, 4665, 3369, 2230, 3717, 4702, 2317, 2983, 3495, 3657, 3778, 3770, 4034, 2493, 3920, 3477, 3635, 4375, 4591, 2318, 2797, 306, 1660, 1751, 2544, 3269, 3553, 3641, 3714, 3719, 524, 2635, 3102, 2101, 1044, 1670, 1311, 952, 525, 433, 1379, 1170, 1183, 1863, 1903, 1695, 1455, 609, 766, 1284, 4630, 4638, 4643, 4656, 228, 1041, 71, 815, 13, 31, 218, 294, 342, 587, 546, 962, 1013, 1066, 1404, 1582, 157, 275, 538, 1656, 245, 596, 2050]. quantum [1643, 814, 1464, 1871, 1514, 82, 164, 344, 1236, 299, 504, 1022, 1787, 1952, 152, 438, 630, 1304, 462, 1508, 4651, 4657, 4674, 641, 425, 887, 1946, 60, 1048, 1140, 1293, 4619, 1025, 663, 662, 1551, 1128, 1285, 1239, 1353, 1450, 1789, 1895, 1031, 4690, 782, 109, 1142, 1931, 4658, 1426, 1642, 1686, 1784, 2027, 1054, 820, 21, 1785, 963, 1065, 1823, 1547, 1939, 1134, 1614, 978, 1400, 1473, 1228, 1143, 1679, 2009, 1385, 328, 162, 451, 1059, 1465, 1197, 3748, 1507, 1533, 4639, 137, 193, 1967, 66, 1693, 891, 4620, 1204, 1521, 946, 1123, 1317, 1996]. quantum [1757, 468, 633, 1524, 1103, 588, 617, 1688, 279, 1522, 615, 1921, 1214, 1937, 1543, 578, 1234, 1984, 1129, 671, 97, 148, 206, 130, 1105, 1149, 1682, 1989, 521, 595, 708, 618, 197, 209, 233, 241, 4515, 11, 1756, 580, 135, 526, 1603, 1270, 2040, 1420, 875, 961, 1051, 4621, 2056, 199, 682, 1418, 1712, 1892, 2010, 1218, 150, 1602, 332, 244, 1641, 1759, 876, 1078, 1301, 286, 2043, 1909, 1151, 1846, 581, 767, 113, 127, 250, 926, 1365, 1373, 1645, 1476, 99, 1598, 1599, 17, 1064, 292, 981, 487, 396, 1062, 1715, 166, 496, 2014, 1881]. quantum [687, 1544, 1997, 1117, 59, 1965, 1208, 866, 220, 251, 488, 260, 371, 453, 1462, 1237, 125, 1813, 522, 564, 1395, 1409, 1451, 1488, 1018, 1010, 4622, 643, 731, 846, 345, 146, 203, 1752, 169, 1613, 608, 5, 19, 211, 276, 309, 1435, 2049, 1721, 1445, 1152, 1364, 1843, 1080, 4125, 691, 1286, 1444, 1211, 550, 87, 54, 1313, 1231, 1986, 514, 188, 445, 485, 692, 1297, 1442, 1405, 1617, 1669, 1951, 434, 1154, 1729, 589, 792, 63, 271, 372, 1738, 2026, 4685, 4686, 972, 418, 730, 2051, 403, 1156, 1007, 4476, 1381, 4627]. quantum [1085, 15, 111, 1259, 1162, 1540, 2018, 1992, 1303, 160, 503, 729, 1242, 2030, 1355, 2047, 523, 491, 1173, 14, 428, 442, 475, 1829, 1501, 1372, 364, 470, 793, 701, 1055, 1705, 75, 1671, 3528, 406, 4408, 1216, 973, 1348, 848, 1356, 1718, 1600, 1801, 858, 4637, 4642, 83, 1060, 1644, 590, 1039, 832, 190, 1201, 1280, 1657, 170, 1344, 1542, 9, 65, 1049, 845, 240, 1061, 1334, 367, 623, 629, 898, 1704, 4652, 173, 518, 515, 472, 1661, 1457, 1449, 3148, 2090, 2376, 2254, 2783, 2242, 2244, 2767, 2178, 2283, 2541, 2903, 2292, 2540, 2496]. Quantum [2429]. Quantum-Bayesian [3415]. Quantum-Describable [4257]. quantum-electromagnetic [199]. Quantum-Events [4434]. Quantum-First [4149]. Quantum-Induced [3649]. Quantum-Like [3431, 3094]. Quantum-Matter-Spacetime [3348]. Quantum-Mechanical [3300, 3823, 2817, 3465, 3676, 2317, 1787, 1614, 1473]. Quantum-Mechanics-Free [4719]. Quantum-Particle [2835]. Quantum-realistic [1311]. Quantum-related [102]. Quantum-State [3415]. quantum-statistical [276]. Quantum-to-Classical [3076]. Quantum/ [1449]. Quantumness [3217, 4409, 4300]. Quaplectic [3002]. Quark [2336, 1621, 1803, 520, 446]. quarks [711, 226, 45, 323, 473]. Quasi [3473, 1448, 3141, 4722, 3829, 2016, 1856, 3472, 2311, 258, 432, 1601, 932]. Quasi-cardinals [3473]. quasi-classical [932]. Quasi-continuous [2016]. quasi-distribution [1601]. quasi-Doppler [432]. Quasi-Neutral [3472]. Quasi-probability [1448, 1856, 258]. Quasi-Separability [4722]. Quasi-set [3473]. Quasi-Set-Theoretical [2311]. Quasi-Stationary [3141]. Quasiclassical [3455, 2382, 2440, 2445]. Quasicrystals [2752]. quasideterministic [1906]. Quasielectrostatic [2295]. quasilinear [273]. Quasiparticle [2754]. Quasiparticles [3813]. Quasiprobability [4018, 3980]. Quaternion [3270, 4147, 4446, 284, 193, 198, 314]. Quaternion-Loop [3270]. Quaternionic [3905, 3645, 1945, 1946, 2010, 1091]. Qubit [3573, 4487, 4729]. Qubits [3947, 2980, 2634]. Query [4100]. Quest [2861, 4610, 2968, 4199, 1675, 1394]. Question [3987, 3843, 2842, 2852]. questioned [732]. Questions [3688, 2243, 164, 1677, 1634, 1647]. Quintessence [4011, 2598]. Quintet [4412]. Quo [4566].

R [1468, 2609, 385, 2794, 3011, 3221]. Rôle [4135]. Radial [3081, 1530, 3472]. radiance [3586]. Radiating [2354, 358]. Radiation [2511, 2312, 3083, 3342, 3547, 3927, 2138, 3120, 2421, 2966, 2091, 2661, 4601, 3618, 2505, 2906, 3135, 3797, 3053, 2905, 2675, 4368, 2506, 2572, 2217, 2722, 2788, 3623, 3909, 4396, 3725, 3448, 3016, 3269, 743, 1972, 1108, 1178, 1113, 1205, 1910, 855, 120, 471, 1638, 1357, 1540, 48, 693, 1660, 1646, 1751, 2261]. radiative [116]. Radiopure [3324]. Radius [3256]. Radix [3796]. Raggio [475]. Rainbow [3754]. Raising [3915]. Ramond [3091]. Randall [1238, 1248, 1249, 1222, 618]. Random [2740, 3118, 4343, 4325, 3290, 2040, 3038, 3254, 3336, 4440, 4096, 3820, 2908, 2100, 2041, 1280, 4329, 1426, 1205, 1134, 1606, 313, 1882, 1618, 1188]. Randomness [4222, 3429, 1533, 1447]. range [374]. ranged [840]. rank [461]. rapidly [785]. Rarefied [3117]. Rarita [3532]. Rastall [1111, 295]. Rate [2966, 3318, 3941, 3016, 642]. Rates [4383]. Rather [2976]. ratio [745, 302]. rational [1587]. rationale [145]. Rationality [2879]. Ratios [3812]. Ray [2460, 2355, 2469, 1799, 2754]. Ray-Splitting [2460]. Rayleigh [1734]. Rays [3500, 3893, 1018, 709]. Re [4442, 3359, 2824]. Re-Appraisal [3359]. Re-Assessment [2824]. Re-interpretation [3275]. Reaction [2312, 3618, 2906, 3053, 2905, 2506, 2788, 2134, 3909, 4396, 2020, 855, 540]. Reactions [3335, 437]. Read [4603]. Reading [3874]. reaffirmation [1015]. Real [3138, 4562, 2963, 4275, 3087, 4739, 3538, 3290, 4353, 3536, 2851, 1135, 4532, 3589, 3507, 3503, 314, 1571, 1574, 1150]. Real-Vector-Space [3538, 3507]. Realism [4548, 2890, 4178, 2223, 2225, 4522, 2406, 588, 4559, 2996, 2580, 1149, 3772, 4525, 2950, 4443, 3050, 3442, 3854, 1405, 4283, 2840, 4290, 2986, 4218, 982, 1930, 1921, 1936, 3512, 980, 457, 1444, 1865]. realist [1317]. Realistic [4053, 1402, 4449, 3195, 2950, 2118, 2122, 3943, 4054, 3162, 3192, 1783, 1180, 1253, 220, 488, 63, 1311]. Reality [2841, 3300, 3360, 4734, 2853, 2878, 4173, 3349, 3437, 3879, 4670, 2737, 680, 4310, 4175, 3557, 3130, 4091, 1401, 3176, 2227, 4464, 3435, 3105, 3854, 4311, 3004, 1487, 4052, 4298, 3624, 4055, 3513, 4212, 2249, 2873, 4472, 3495, 1950, 1731, 1170, 1183, 1269, 1141, 1832, 1714, 615, 1989, 225, 1413, 4627, 1259, 281, 1897, 1809, 1896, 1600, 960, 1267, 1478, 1898, 1403, 1044, 1217]. realizable [988]. Realization [2276, 2296, 3465, 4719, 328]. Really [2607, 4032, 1342, 412]. Realms [3455]. Reappearance [4540]. Rearrangement [4407]. Reason [2319]. Reasonable [3426]. Reasoning [2865, 4514]. Reasons [2248]. Reassessing [4374]. Reassessment [3647]. Rechenberg [2558]. Reciprocal [3982, 3577, 3002]. Reciprocity [3187, 2324, 948]. Recoil [3987]. Reconciled [3459]. reconciliation [1335]. Reconciling [4221, 3148]. Reconsideration [2877]. Reconsidered [2792, 1741, 1260]. Reconsidering [3188, 4459]. Reconstructing [3539]. Reconstruction [3857, 4738, 3810, 3859, 170, 2960, 60, 417]. reconstructions [2027]. Records [4360]. recover [317]. Recovery [4098, 1355]. Rectangular [2451]. recurrence [3748]. Recursive [1413]. Rédei [2600, 2178]. Redshift [3139, 2995, 257]. Redshift-Distance [3139]. Redshifts [4469, 3139, 165]. Reduced [3099, 3715]. Reduction [3452, 3440, 3453, 3459, 3893, 2393, 3679, 3456, 4081, 3727, 4066, 4202, 4170, 3491, 3227, 4533, 51, 42, 1276, 1512, 1711, 1618, 1163, 1283, 231, 848, 174]. Redundancy [4233]. Redundant [2938]. Reexamination [114]. Refer [3715]. Reference [4526, 3031, 4058, 2821, 2175, 3026, 102, 1219, 376, 81, 3395, 4689, 452, 507]. Refined [4358]. Refinement [1233]. Reflecting [3606]. Reflection [2890, 983]. Reflections [2833, 3924]. reflexive [3778]. Reformulation [4463, 4570]. refutation [60]. Refutations [3889]. refute [992]. Refuted [3162, 3192]. Regained [4082]. Regarding [4232, 131]. Regge [2169, 831]. Regime [4550, 3858, 4735]. Regions [4435]. Registration [3545, 586]. Regular [2653]. Regularising [4135]. Regularization [1963, 4269, 1802]. Regularizing [4335]. Reichenbach [3525, 2026, 1336]. Reichenbachian [2991, 3541, 4044, 3346]. Reinterpretation [3306, 2173, 39, 3]. reinterpreted [1439]. Reinterpreting [2673, 1357]. Reinvention [3857]. Reject [2248]. Rejoinder [93]. Rekindling [4572]. Related [2550, 3157, 3721, 2546, 3237, 102, 3748, 1117, 1188]. Relating [3846, 3714]. Relation [2301, 4544, 4165, 2812, 3775, 3561, 3934, 3275, 3979, 2211, 3139, 3889, 4459, 3595, 2581, 3252, 3227, 4702, 766, 1194, 1195, 402, 412, 1367, 382, 1764, 2172, 893, 898, 472]. Relational [4629, 3899, 4205, 4489, 4493, 4725, 3969, 4437, 3087, 3588, 4491, 4058, 4697, 4146, 4339, 4574, 4333, 4397, 4558, 4217, 4412, 4411, 4071, 3637, 3054, 3148, 4106, 3930, 4360, 4401, 2316]. Relationism [3242, 2406]. Relations [3794, 2449, 3229, 2386, 3754, 3824, 4130, 4289, 4310, 2793, 3912, 4323, 3385, 3957, 2791, 3837, 4066, 3600, 2234, 204, 948, 847, 734, 1034, 1828, 1016, 485, 790, 1457]. Relationship [3294, 1277, 324, 1173, 236]. Relationships [21, 2287, 2954]. Relative [3306, 4345, 3884, 3051, 98, 4168, 262, 3601, 2670, 4137, 247, 869, 4515, 1715, 1591, 1412, 4725]. Relatives [3410]. relativisitic [198]. Relativistic [283, 4228, 2724, 2789, 4531, 4506, 2200, 4059, 3575, 3763, 2511, 4138, 3043, 3088, 4605, 3440, 2714, 2503, 2733, 3072, 3799, 1952, 4594, 3301, 1535, 2466, 2189, 2188, 2385, 3342, 3547, 1783, 2168, 3214, 3078, 3089, 3133, 3219, 4568, 4234, 3391, 573, 2381, 2524, 2720, 2917, 3394, 3275, 2636, 3397, 1276, 2393, 2707, 1937, 3905, 4517, 4426, 4381, 380, 1959, 3765, 2807, 875, 2971, 3311, 2170, 2162, 2494, 2599, 2166, 2525, 660, 3892, 3045, 2264, 3447, 818, 3523, 3919, 2821, 3176, 2581, 1794, 2527, 415, 2712, 3851, 3179, 4507, 2506, 2907, 1764, 2155, 2519, 2522]. Relativistic [2729, 2791, 4715, 2112, 3282, 2586, 2154, 4049, 1992, 2913, 2023, 2426, 2427, 2557, 2764, 2902, 3064, 2504, 4707, 3013, 482, 2350, 2782, 3382, 3933, 2151, 4282, 2324, 3159, 2705, 4735, 3073, 3680, 768, 2493, 2243, 609, 766, 1040, 1343, 1635, 1295, 1304, 1580, 1763, 1792, 180, 1939, 1614, 1713, 1980, 137, 207, 1706, 1036, 357, 1204, 1521, 1637, 1103, 1369, 4026, 671, 645, 1989, 1382, 841, 3303, 961, 682, 1418, 1712, 366, 150, 400, 1574, 1638, 496, 124, 322, 2017, 1699, 1364, 550, 4456, 404, 485, 4689, 1162, 1186, 1858, 1133, 933, 221, 282]. relativistic [560, 4457, 343, 757, 1957, 4636, 1179, 2002, 1298, 424, 4042]. Relativistically [3636, 659, 1455]. Relativity [3198, 4506, 1905, 4213, 2877, 3014, 2407, 3083, 3927, 3959, 4432, 2491, 3475, 2386, 2672, 2901, 3583, 3754, 2513, 2711, 4067, 3646, 4231, 4451, 3051, 4297, 2723, 3362, 2649, 4293, 2288, 3031, 3055, 3139, 4224, 3150, 2518, 2084, 3723, 4608, 3581, 3827, 4023, 2962, 3399, 2521, 3603, 3002, 4253, 4216, 2989, 4521, 3964, 3146, 3288, 4352, 4254, 2639, 4383, 2924, 3981, 2519, 4414, 2152, 4080, 4424, 3256, 2597, 493, 2512, 4663, 4668, 2849, 4710, 2360, 3226, 3711, 4461, 3169, 3637, 3026, 2370, 2576, 2073, 4033, 3650, 261, 430, 435, 570, 1013, 161, 817, 1774, 1857, 611, 613, 253]. relativity [1454, 1164, 679, 163, 192, 387, 696, 746, 252, 773, 264, 759, 855, 199, 1374, 12, 22, 36, 921, 736, 167, 61, 131, 614, 345, 555, 1923, 1845, 1121, 81, 713, 692, 1669, 1951, 317, 791, 1130, 1107, 1175, 1320, 308, 145, 444, 1091, 68, 76, 160, 2042, 465, 1732, 10, 239, 1336, 612, 655, 899, 908, 917, 1104, 839, 911, 524, 472, 1457, 1899, 1950, 191, 3691, 3463, 4648, 4689, 3010, 2821, 2677]. Relativity-Proof [4254]. Relativized [1409, 1395]. Relativizing [2370]. Relaxation [4716, 4141, 2440]. relevance [74]. Relieve [3667]. Remark [3076, 3522]. Remarks [2658, 2395, 1564, 2107, 3104, 4091, 2762, 4726, 2763, 3883, 474, 3333, 282, 297, 472, 1217, 764, 1671, 3792]. Remiddi [2114]. Reminiscences [513]. Remote [3921, 3907]. Removal [205]. Renner [4203, 4351, 4242, 4360]. Renormalizability [3597, 4061]. Renormalizable [3274, 1846]. Renormalization [3757, 2990, 4395, 4121]. Renormalized [4192, 690]. Repeal [2603]. Repeatable [1789]. Replacing [3937]. replicating [542]. replication [275]. Reply [3182, 3113, 4032, 3112, 4040, 4489, 2309, 3173, 423, 422, 3356, 4186, 386, 3709, 4332, 3985, 3539, 475, 3206, 2249, 3872, 4401, 528, 1593, 618, 614, 455, 612]. Report [897, 1338, 1003]. Represent [4251]. Representation [4065, 2135, 2145, 2310, 3100, 2276, 3254, 2577, 3174, 3204, 2862, 3464, 2350, 3933, 2863, 3670, 340, 1785, 1383, 1938, 935, 1584, 607]. Representational [4233, 4231]. Representations [3784, 3903, 3573, 4018, 3992, 2545, 2536, 2400, 4180, 4607, 708, 1506, 543, 565, 2019, 1833, 1671, 973, 542]. representative [3993, 3880, 4031]. Reproducibility [3861, 583]. Reprogramming [4617, 2496]. repulsion [4692]. Repulsive [4051, 2353, 1032]. Requirements [2633, 1355]. Research [2845, 3208, 2886, 2311, 3600, 1001, 1408]. Resilient [3427]. Resistance [3901]. Resistive [2237]. Resolution [4604, 448, 2596, 2599, 3017, 2606, 2977, 2423, 2810, 2011, 1977, 1240, 1843, 1121, 625]. Resolutions [4172, 2448]. resolved [531, 1936]. Resolving [4187, 3489, 4691, 3175]. Resonance [2693, 2350, 2005, 1971]. Resonator [2122]. Resource [2633, 3370]. Respected [3632]. Respecting [4312]. Response [4201, 1252, 3137, 1132, 4277, 1190, 4031, 2824, 569, 1140, 1250, 1014]. Responses [4726]. rest [126, 171, 716]. Restricted [2706]. Restrictions [4266, 3447]. Result [4418, 4709, 2680, 3696, 660, 880]. Results [3329, 4124, 2964, 2446, 4630, 4638, 4643, 4656, 1142, 1784, 1225, 2004, 1344]. Retaining [2860]. Retardation [2188, 2385, 387]. retarded [1735]. retirement [1610]. Retroactive [943, 1603, 337]. Retrocausal [4204, 4451, 4145]. Retrocausality [3165]. Retrodiction [1823]. Revealed [3508]. Reversal [4002, 2974, 4541, 4557, 3925, 530, 2008, 1079, 55, 369, 1717, 505]. Reverse [4467]. Reversible [3788, 3268]. Review [2090, 1748, 1905, 2344, 2345, 2375, 2376, 2387, 2388, 2389, 2590, 1744, 2254, 3482, 4124, 4448, 2063, 3328, 2641, 2786, 4343, 1745, 2783, 2708, 2609, 2600, 2242, 1521, 2795, 2244, 2245, 1362, 2784, 2836, 2499, 2767, 2474, 2495, 4195, 2178, 2283, 2497, 2541, 1361, 2494, 2565, 2631, 16, 2255, 3407, 2899, 3387, 2558, 2903, 2539, 2128, 2821, 2677, 2303, 2292, 2102, 2184, 2785, 1747, 2316, 3010, 2540, 4042, 1749, 2808, 4089, 2193, 2997, 2291, 2794, 3011, 2498, 2617, 2820, 2496, 2475, 2687, 2766, 1746, 2566, 2103, 2241, 2365, 2624, 2819, 2967, 1360, 3221, 3814]. Review [2243, 2101, 2429, 1626, 953, 808, 884, 622, 1377, 1483, 929, 1115, 923, 1915, 722, 1872, 1916, 915, 1038, 1664, 1429, 1816, 721, 1463, 1709, 1255, 938, 1942, 1649, 1493, 956, 1703, 1082, 925, 737, 824, 1722, 1830, 1941, 1956, 1330, 928, 1093, 1126, 1272, 1839, 1860, 1290, 883, 639, 603, 1561, 1862, 1652, 999, 707, 1388, 619, 769, 1673, 1567, 1795, 1796, 1797, 1798, 882, 1257, 1167, 1432, 1595, 1169, 1127, 650, 1037, 1665, 666, 1918, 723, 1861, 1480, 1560, 1496, 1350, 771, 1566, 1569, 1497, 930, 778, 1817, 726, 1030, 1102, 1387, 1662, 1433, 1927, 1788, 672, 1495]. review [945, 738, 658, 1831, 1101, 1668, 399, 1274, 1814, 739, 895, 1565, 1430, 1815, 807, 627, 809, 914, 1482, 628, 913, 955, 787, 1737, 1386, 1391, 805, 1029, 1331, 651, 657, 795, 652, 1370, 706, 1629, 1100, 1137, 1928, 1893, 788, 1182, 1628, 1168, 632, 1045, 1390, 1917, 1221, 1672, 1559, 602, 740, 1492, 1919, 777, 1926, 1256, 620, 1650, 1494, 1654, 1428, 1291, 1092, 1873, 673, 806, 1273, 903, 1046, 1822, 741, 1323, 1324, 1625, 1594, 1306, 998, 779, 944, 1098, 1779, 1943, 924, 984, 776, 1666, 631, 1002, 1376, 1674, 904, 1653, 1349, 1490, 939, 667, 1651, 1332, 1818, 638]. review [724, 885, 1431, 1806, 1207, 906, 719, 1663, 1840, 1596, 996, 796, 780, 1558, 1568, 1648, 1805, 1935, 997, 1181, 1841, 1859, 1819, 954, 985, 1479, 1460, 1292, 1914, 1667, 674, 881, 1933, 1191, 775, 1375, 594, 907, 797, 1289, 644, 786, 1624, 1821, 1623, 1820, 957, 1212, 1389, 1934, 1481, 601, 770, 1099, 1339, 896, 1780, 725, 905, 1627, 1955, 720, 894, 621, 1459, 1491]. Reviews [2315, 1902, 986, 1901]. Revised [4433, 655, 908, 917, 899]. revisitation [640]. Revisited [4388, 3000, 3049, 2679, 2826, 3927, 2415, 2482, 2612, 2057, 2945, 3136, 3069, 4194, 3541, 3497, 3037, 2348, 2589, 4513, 531, 1690, 453, 1337, 559, 318, 506, 893, 1477]. Revisiting [3551]. Revolution [4306, 2540]. revolutionize [656]. Rham [443]. RHIC [2502]. Riccati [4228]. Ricci [4317, 236]. Richard [2624]. Riemannian [419, 110, 143, 2021]. Riesz [3580, 2272, 3755]. right [713]. right-angled [713]. Rigid [3987, 3691, 3037, 3026]. Rigorous [3737, 276, 1682, 2043]. Rigour [3377]. Rindler [2677]. Ring [3067, 4278, 3318]. Rings [4052]. ripples [1063]. Rise [2886, 1000, 1005, 1014]. Rizzi [2821]. Road [3198, 3297]. Robbins [3326]. Roberto [2498]. Robertson [4121, 3869, 2729, 655, 899, 908, 917]. Robustness [3208]. rocket [899]. rocket-borne [899]. Rodrigues [1190]. Rods [3087]. Roger [4357]. Rohrlich [1639, 1610]. Role [2208, 3203, 3511, 2999, 4511, 2880, 3588, 3922, 3003, 4338, 3494, 3600, 2531, 4105, 1886, 1789, 3748, 1734, 1081, 1866, 4331, 292]. Roles [4154, 2413]. Romanticism [4139]. room [4476]. Root [2280]. Roots [2845, 3723, 2278]. Rosa [1190]. Rosen [3147, 684, 699, 3469, 678, 675, 978, 3149, 1051, 804, 2422, 457, 114, 1850, 2883, 783, 3696, 1076, 933, 437, 784, 898]. Rotating [3081, 4115, 2098, 2265, 3996, 2821, 303, 2175, 2512, 2174, 2627, 2546, 4260, 2177, 785, 380, 423, 480, 422, 1869, 750, 4689, 1114, 452]. Rotation [2733, 2922, 2790, 2815, 3540, 3903, 4469, 3045, 1698, 1179, 477]. Rotational [2732]. Rotations [3868, 607]. Rothman [2241]. rough [1445]. Route [3415]. Routes [3131]. Routledge [2641]. Rovelli [3296]. Rovellian [4578]. Roy [2316, 2255]. RQRPA [2024]. Ruggiero [2821]. Rule [4255, 4308, 2770, 3947, 4598, 3034, 4716, 3594, 3925, 2248, 3446, 3254, 4090, 4142, 4140, 2869, 3558, 4554, 1832, 801, 4295]. Rules [2479, 1065, 3529, 99, 1638, 376, 360, 542, 845]. Runaway [3053]. Russellian [3639]. RW [4588]. Rydberg [3520]. Ryder [2090].

S [3049, 1976, 2903, 2624]. S-matrix [1976]. s10701-009-9351-4 [3322]. S60.00 [2784]. Sachs [3403]. Sacrifice [4406]. Safe [4195]. Safety [4128]. Sagnac [4432, 2707, 2546]. Sagredo [1153]. Salecker [3874]. Saletan [2245]. Salpeter [1960]. Salvador [2244, 2785]. Salviati [1153]. Same [3008]. sample [1240, 1223]. samples [1686]. San [2786, 2820]. Sans [4396]. Santilli [2609]. Santos [2968, 2987]. Sasso [3329]. satellites [840]. satisfying [710]. Savart [1605]. Say [2667]. Says [4718]. Scalable [2633, 2633]. Scalar [4458, 2525, 2713, 2914, 3201, 3935, 2108, 2657, 1977, 1911, 1598, 1599, 3911]. Scale [4228, 3745, 2672, 3383, 2510, 4134, 4027, 3060, 3201, 1969, 1158, 598]. scale-breaking [598]. Scale-Relativistic [4228]. Scaled [4505]. Scales [2901, 2341, 4195, 4450, 3923]. Scaling [1178, 3342, 2457, 2351]. Scarani [3206]. Scars [2452]. Scattering [2724, 2789, 3484, 3545, 2835, 2892, 2543, 2459, 2350, 2140, 451, 203]. Scenario [3121, 4144, 4200]. Scenarios [3527, 4099]. Schächter [2783]. Scheme [2434, 4429, 2096, 584, 729, 768]. Schemes [2982]. Schieve [2494]. Schlegel [455, 454]. Schlosshauer [3182]. Schmelzer [3985]. School [2366, 2377]. Schopenhauer [2599]. Schott [774]. Schrödinger [1022, 4480, 94, 1359, 3876, 1004, 1008, 1024, 3736, 1013, 1271, 3088, 4012, 2486, 714, 1232, 3946, 3839, 1112, 2588, 3095, 64, 1341, 1009, 1775, 4477, 1001, 4258, 3870, 2213, 2166, 2525, 1641, 3413, 3253, 2264, 40, 3630, 1006, 4737, 1000, 1005, 959, 1014, 1021, 3997, 1019, 2719, 1016, 1011, 3712, 2142, 1007, 2902, 4707, 3013, 38, 3938, 1012, 2705, 1298, 4488, 3152, 932, 3760]. Schrödinger-like [2264, 2166, 2525]. Schuetz [2673]. Schuster [745, 1176]. Schütte [1887]. Schwarzschild [3136, 1043, 3642, 3256, 751, 4520, 2320]. Schwinger [2010, 2303, 2239, 2240, 2252, 2253, 2261, 4640, 3532, 3738, 2539]. Science [1905, 2388, 2641, 3376, 1401, 4281, 1868, 4, 4687, 133, 1081, 1866, 4661, 1023, 500, 1393, 1153, 324, 4627, 23, 2784]. sciences [940]. Scientific [3729, 2825, 2631, 2539, 1865, 2879, 2794, 940, 1996, 1000, 4640, 360]. Scope [4017, 3673, 3195]. Scopes [4004]. Scott [2344, 2786]. screen [1676]. Scully [2242, 2091, 3190, 3290]. Search [3166, 2699, 4089, 1170, 1183, 1538, 972]. Searches [2340]. Sebens [3792]. Second [2090, 1072, 2648, 193, 2162, 2264, 4265, 1878, 461, 4295, 3527, 3117, 3123, 4027, 974, 2253, 3116, 2141, 2383, 3114, 3268, 2810]. Second-Law [3117, 2810]. Second-order [1072, 1878]. second-rank [461]. Sector [3911]. Sectors [4701]. Secure [3865]. Securing [4515]. seeing [649]. Seen [2213, 439]. Selected [2231, 2798, 2870, 1218, 1344]. Selection [2641, 2715, 99, 1638, 3766, 340, 4687]. Selective [4178]. Self [2113, 2848, 3687, 340, 2457, 2999, 624, 3058, 1632, 2106, 2662, 3446, 3792, 4182, 1846, 3604, 4051, 3479, 2382, 3395, 1358, 1540, 2902, 4692, 346, 2653, 962, 398, 703, 749, 445, 1546, 1922, 1526, 1661]. Self-accelerated [4051]. self-adjoint [1922]. Self-adjointness [624, 1526]. Self-Consistent [2382, 2653, 340, 3446, 1846, 3604, 445]. Self-controlled [2848]. Self-dual [1632, 749]. Self-Duality [3687]. Self-Energy [2902, 4182, 1540]. Self-Field [2113, 1546]. Self-force [3058]. Self-generating [1358]. Self-Gravitating [2662, 3604, 3479]. Self-Induced [2999, 2106]. self-interaction [703, 1661]. Self-Locating [3792]. Self-mass [4182]. self-organization [962]. Self-reference [3395]. Self-repulsion [4692]. Self-Similarity [2457]. self-stresses [398]. Self-turbulence [346]. Selleri [2838, 2243]. Semantic [3370, 1930, 2818, 2996, 3670, 1921]. Semantics [2895, 59, 1296]. Semi [3224, 4026]. Semi-Classical [3224, 4026]. Semiclassical [2568, 2552, 2466, 2551, 2468, 2463, 2469, 2454, 4260, 1514, 653]. Semiconductors [2753]. Semigroup [2146, 2008, 1987]. semiorder [1235]. Sense [4000, 3792, 2615, 4071]. sensitivity [1967]. Separability [3438, 3772, 4722, 3367, 2983, 3667, 560]. Separable [3887, 1060, 1267]. Separate [3476]. separated [525, 1683, 1200]. Separation [3520, 3545, 3841]. September [1905, 2378]. Sequential [3022, 3861, 4407, 1239, 1110]. Series [2647, 2351]. Serious [4343]. Set [3283, 4527, 4730, 2400, 2311, 1808, 3766, 3473, 1223, 406, 420]. Set-Theoretic [4730]. Sets [4072, 2416, 3490, 4252, 2978, 668, 96]. Setting [4153, 3487, 3793, 4267, 4320]. Seventieth [2325, 2968, 2838, 2036, 1047, 958]. Seventy [2431, 2444, 2456, 2478, 675, 1710, 742]. Seventy-Fifth [2431, 2444, 2456, 2478, 675]. seventy-five [1710]. several [58]. Shades [4363]. Shadows [4470]. Shalt [3298]. Shan [4089]. Shannon [2638, 2883]. Shape [3688, 2062, 2755, 4347, 4533]. Shaped [2421]. Shapes [3999, 1436]. Sharing [3507]. sharp [2027]. Sharpness [3240, 3682]. Shell [2718, 2909, 2910, 2506, 1802, 1310, 1958, 1162, 1981]. Shift [2207, 2189, 2188, 2362, 2550, 3157, 2116, 3752, 2680, 546, 1612, 155]. shifting [1877]. Shifts [2551, 2218]. Shigeji [2244, 2785]. Shimony [3737, 2956]. Short [2063, 2055, 293, 2096, 633, 1305]. Short-time [2055, 293]. Should [2225, 4519, 2830, 1259]. Show [3982, 3098]. shown [32]. SIC [4019, 3464]. SICs [4020, 4309, 4021]. sideband [508]. Sierpinski [2061]. Signac [4591]. Signal [3592, 2850, 1109, 426]. Signaling [3239, 4714]. Signalling [3942, 1051]. Signals [2373, 3113, 2977, 1036, 1174]. Signature [3893]. Signatures [3388]. Significance [4342, 3659, 3377, 102, 267, 326, 334, 687]. Signposts [3297]. Similarity [3674, 2457, 3640]. Simple [4076, 3674, 2223, 2312, 2938, 2607, 2629, 4560, 4214, 3266, 4386, 301, 398, 30, 1993]. Simplest [4019]. Simplex [3464]. Simplicio [1153]. Simplified [1070]. Simply [4256]. Simpson [1240]. Simulating [1729]. Simulation [4496, 3851, 3147, 1932]. Simulations [3989]. Simultaneity [4213, 3582, 2491, 3356, 3996, 3595, 1318, 108, 413]. Simultaneous [3096, 3624, 367, 1095, 1096, 58, 135]. Sinai [2469]. since [547]. Singal [4462]. Singapore [2631, 2794]. Single [3910, 3093, 3343, 1400, 3937, 3508, 3399, 3510, 1398, 2118, 3442, 4341, 4617, 2496, 2080, 4001, 3866, 1271, 1633, 1733, 2050, 1475, 659, 1117, 146, 1891, 1655]. Single-case [1398]. Single-Particle [3937, 2080, 1400]. single-photon [1475]. single-slit [1117]. single-time [146]. Single-World [4001]. Singlet [3937]. singletons [1523]. Singular [3340, 3090, 273, 3479, 761, 4126]. singular-point [761]. Singularities [3457, 4109, 4357, 3489, 3883, 1695, 205, 1879, 751, 1755, 1215]. Singularity [3757, 3280, 4112, 4240, 447, 643]. Sinusoidal [407]. Sir [502, 1392]. Sisir [2255]. Sites [1247]. Sitter [3198, 3916, 754, 1829, 409, 2657]. Situation [4023, 2516]. Six [2086, 1464, 2315, 773, 4632, 4647, 101]. Six-Dimensional [2086, 773]. sixtieth [1266]. sixty [1675, 1499, 1944]. sixty-fifth [1499, 1944]. sixty-plus [1675]. Size [4549, 3874, 3055, 2991, 4044]. Skepticism [4196]. Sky [4120]. Skyrme [1621]. Slave [2436]. Slave-Boson [2436]. slice [3403]. Slightly [3446]. Slit [2395, 3484, 4496, 3261, 3961, 3655, 4586, 3184, 3421, 2095, 1471, 1472, 3024, 1117, 2172]. Slowly [2512]. SM [3202]. Small [4313, 2299, 4530]. smallest [712]. Smearing [3619]. smearon [488]. Smooth [4063, 4150]. Smoothness [2800]. Snygg [2128]. Social [2259]. Society [1905]. Sodium [2097]. Soft [3512]. softcover [2795, 2903, 2677, 2316, 2540, 2498, 2617, 2687, 2315]. Solar [2343, 2699]. Soldner [362]. Solenoid [4698]. solenoids [1537]. Solid [3533, 4276, 2628]. Solid-State [3533, 2628]. Solipsism [3913]. Solitodynamics [4579]. Soliton [4699, 1825]. Solitons [3066, 1298, 832]. Solomon [3082]. Solution [4592, 4550, 3679, 4392, 3806, 3061, 3185, 4701, 4324, 2114, 4272, 1346, 1043, 2976]. Solutions [3027, 4011, 3021, 3605, 2502, 3620, 3053, 3914, 3604, 2142, 3653, 1032, 480, 1527, 753, 407, 1563, 751, 38, 1298]. solvable [1541]. Solvay [3592]. solve [1048]. Solved [3384, 647]. Solvents [4369]. Solving [2884]. Some [3000, 1736, 342, 4309, 3939, 813, 988, 310, 1784, 4239, 2545, 3051, 2323, 3430, 4043, 575, 2413, 3220, 2058, 3003, 2215, 1401, 4129, 4726, 47, 1043, 543, 565, 751, 764, 216, 1671, 852, 3049, 323, 1001, 580, 260, 2015, 254, 784, 4091]. Sometimes [4549]. Sommerfeld [531, 1063]. Sonic [4033]. Sophisticated [4237]. Sound [4387, 4603, 4033]. Source [3867, 3713, 4243, 3981, 3955, 3345, 743, 176, 484, 354]. Sources [4115, 3883, 2318, 1911, 1697]. Space [3267, 4062, 4109, 4169, 2861, 3081, 3916, 4359, 2289, 2815, 3754, 3946, 4133, 4561, 2072, 3864, 2198, 4460, 2587, 3415, 3331, 3058, 2916, 3956, 4501, 4026, 4086, 3937, 4379, 4670, 2510, 4709, 3682, 4713, 3538, 3966, 459, 4224, 3830, 634, 2574, 3223, 3439, 3363, 3723, 2671, 2817, 2851, 4697, 3330, 2923, 4422, 3708, 3997, 4585, 2357, 2730, 4093, 2727, 2926, 3105, 2382, 2536, 3642, 3074, 4715, 4347, 3333, 4536, 2627, 4202, 4271, 3332, 4147, 4446, 4434, 3793, 2487, 4145, 2573, 2146, 3567, 2585, 2420, 4529, 3227, 4139, 3963, 4533, 2992, 4045, 908, 2863, 2182, 3738, 2657, 3507]. Space [4471, 4470, 3838, 4110, 1503, 1945, 1731, 827, 1769, 815, 1696, 1907, 438, 1177, 60, 1525, 599, 1054, 1702, 1804, 1165, 654, 967, 1210, 1333, 1698, 387, 222, 333, 1382, 708, 747, 875, 961, 150, 575, 1581, 156, 1909, 1827, 1151, 110, 143, 534, 1118, 1119, 242, 352, 388, 432, 1451, 1618, 447, 643, 1277, 1166, 608, 331, 1692, 754, 927, 665, 268, 1063, 1133, 2024, 1829, 1999, 465, 1104, 1600, 351, 590, 880, 1215]. Space-Time [4359, 4446]. Spacecraft [2781]. spacelike [1310]. Spaces [2535, 3433, 2806, 2790, 2815, 2827, 2901, 3383, 3577, 2537, 4527, 3645, 2212, 2259, 2800, 3174, 3490, 2536, 3668, 1724, 1772, 1785, 3186, 1261, 1243, 1975, 668, 1241, 1227, 1240, 1223, 1694, 1010, 1419, 1231, 188, 1985]. Spacetime [2682, 2086, 4012, 3582, 4556, 4594, 4104, 3547, 3348, 3978, 2790, 2827, 4609, 2148, 3789, 4528, 2406, 3839, 2604, 4419, 2636, 3594, 3400, 4498, 4584, 3029, 3003, 2201, 4129, 4152, 2175, 2354, 307, 322, 2203, 3007, 1395, 1409, 4464, 3666, 3167, 2926, 2302, 2514, 2771, 4054, 3062, 3462, 1055, 3148, 4436, 2121, 3181, 330, 1380, 1773, 1807, 1770, 1573, 1879, 1577, 1571, 4688, 339, 583, 1254, 1854, 1995, 321, 485, 1297, 1767, 409, 1981, 1321, 4444, 2624]. Spacetimes [2665, 4239, 4240, 3524, 3282, 3201, 419, 2021, 1755]. Spacing [2470]. Spatial [2207, 3475, 3308, 4134, 3494, 3005, 3227, 676, 854, 744, 1318, 1368]. Spatially [3151, 2459]. Spatially-VSL [3151]. Spatio [4254]. Spatio-temporal [4254]. Spatiotemporal [4081]. Speak [3792]. Special [2477, 4213, 1635, 3014, 3959, 2144, 4004, 2491, 3475, 3769, 4610, 3646, 4306, 4451, 4297, 2723, 3362, 3031, 2932, 3723, 3827, 3556, 3818, 4017, 4304, 3603, 2677, 3313, 4383, 3981, 2729, 2849, 2360, 3954, 3226, 3169, 2151, 817, 1774, 1857, 611, 613, 192, 387, 252, 264, 199, 167, 937, 131, 614, 1121, 713, 317, 493, 4648, 4663, 4668, 145, 68, 76, 160, 38, 465, 1732, 1336, 612, 655, 899, 908, 917, 1104, 424, 911, 3691, 3463, 1355]. special-relativistic [424]. Species [2211]. Specker [3450, 1887, 4376, 1768, 566, 1888, 3519, 4382, 3500, 3759]. Spectra [2646, 2452, 2462, 2469, 487]. Spectral [2470, 2481, 4598, 3619, 4172, 2451, 3567, 1365, 1373, 731, 1519]. Spectroscopy [2169, 2168, 1436]. Spectrum [3547, 2136, 3797, 3447, 3791, 3915, 900, 912, 922, 1017]. spectrum-generating [1017]. Speed [3117, 4241, 3598, 1932, 3009, 4588, 4243, 2174, 2603, 41, 921]. Speed-Dependent [3117]. speeds [1983]. Speedup [3899, 4065, 4205, 4370, 3345]. Speedups [3531]. Spekkens [3886, 3554]. Sphere [2733, 4063]. spherical [120]. Spherically [4115, 1923, 1755]. sphinx [4627]. Spin [4531, 3593, 2799, 4385, 4322, 4605, 3799, 2458, 3093, 2938, 3337, 2466, 3381, 3334, 3328, 2812, 4219, 180, 2145, 3621, 3317, 450, 2432, 3211, 2165, 2436, 2807, 3322, 3321, 2792, 4646, 4672, 3319, 1690, 3313, 3708, 2695, 3882, 2732, 3868, 3326, 3037, 2953, 2791, 3047, 3333, 2141, 3831, 4049, 3793, 3628, 1601, 3335, 2124, 2692, 4523, 3257, 4524, 1455, 829, 1538, 1141, 1448, 1803, 843, 1939, 1072, 78, 198, 1070, 448, 1937, 765, 1402, 1959, 1096, 659, 457, 2017, 1960, 1856, 478, 1220, 1196, 1644, 476, 2694, 2590]. Spin- [2799, 3093, 2145, 2436, 2145, 1448, 843, 1072, 78, 448, 1937, 765, 1096, 1960, 1856, 1220]. Spin-0 [4531]. Spin-2 [3593]. Spin-3 [2694]. Spin-3/ [2694]. Spin-Dependent [3257]. Spin-Geometry [4523]. Spin-Orbit [2458, 829]. Spin-Statistics [3799, 3319, 2732, 3326, 2791, 3333]. Spin-Zero [2953]. Spinelli [613, 614, 612]. Spinless [3303, 4445, 357, 1613]. Spinning [2110, 2109, 1614, 1520, 43]. Spinor [2085, 3331, 3937, 2971, 3519, 755, 2108, 1550, 393, 1053, 267, 326, 334, 347, 636, 927]. Spinorial [4471]. Spinors [2620, 1072, 2545, 3937, 555, 4424, 3787, 677, 875, 1572, 2015]. spinors-Construction [2015]. Spins [4313, 2697]. Spiral [3357]. Split [2814, 987]. split-beam [987]. Splitter [3510]. Splitting [3867, 2460, 4487, 3791, 2469]. Spontaneity [3004]. Spontaneous [3669, 1545, 1203, 2093, 4539, 4277, 2712, 2217, 4724, 4415, 4582, 498, 851, 1580, 1799, 1160, 1759, 375]. Spontaneously [3893, 2095, 1314]. Sporadic [4021]. spread [1287]. Spreads [2587]. Springer [2344, 2899, 2558, 2687]. Springer-Verlag [2344]. Square [3757, 3249, 2290, 451]. Squashed [4098]. Squeezed [4185, 848]. SQUID [1342]. SR [2780]. SRT [2546]. Stability [1262, 2317, 46, 582]. Stabilization [3064]. Stabilizer [3554]. Stable [4345, 2363, 233]. stage [557, 1206]. Staircase [3357]. Standard [2417, 4117, 2731, 2925, 4519, 4005, 3750, 2182, 3714, 1895, 1114, 1355]. Standards [2933, 1088, 411]. Standing [2911]. Standpoint [1791, 135]. Stapp [2561, 2955]. Star [2650, 2512]. Stark [2093, 2507]. Stars [3969, 3064, 1877]. State [3284, 3433, 2935, 3440, 2301, 3921, 4143, 1177, 3531, 2806, 4734, 3533, 4509, 4276, 4148, 497, 3415, 4727, 2323, 2448, 2093, 2273, 3542, 3851, 3727, 2142, 3074, 4202, 2383, 2628, 2124, 3491, 3779, 4055, 2064, 2684, 231, 2634, 2114, 4576, 3692, 3838, 1503, 1411, 31, 294, 1427, 1783, 2039, 42, 1512, 1500, 1638, 1408, 19, 486, 201, 256, 377, 318, 401, 1644, 590, 174]. State-Collapse [2064]. State-to-State [4509]. State-Vectors [3692]. Statement [2805, 441, 568]. statements [784]. States [2484, 3663, 3412, 4400, 3689, 2957, 3078, 3089, 3133, 3219, 4489, 3141, 1573, 3461, 2946, 3715, 3101, 3790, 2622, 3425, 3897, 3281, 2436, 3976, 4353, 3658, 2061, 2697, 3675, 4300, 4570, 3129, 2763, 2681, 3741, 4093, 3828, 3464, 2586, 3954, 3236, 4436, 3084, 3107, 4512, 3551, 2983, 51, 13, 596, 1991, 1686, 1811, 1812, 574, 1261, 1475, 595, 233, 1604, 1418, 1171, 1530, 2052, 30, 1073, 1229, 687, 920, 1544, 1613, 211, 1728, 1085, 1159, 1591, 1509, 1355, 467, 848, 1280, 123, 1067]. Statevector [1618]. Static [833, 3960, 1911, 480, 1424, 1755]. Stationary [2237, 3813, 3141, 753, 2512, 661]. Statistical [2582, 3761, 1826, 1890, 4330, 2168, 3689, 3454, 4318, 2485, 2828, 4130, 2795, 2244, 615, 2836, 2250, 2823, 4508, 4564, 2003, 3413, 4015, 3928, 4682, 3603, 2183, 4077, 3287, 3354, 4140, 4628, 2311, 2913, 2173, 4176, 4655, 4408, 4084, 4374, 3152, 989, 823, 888, 71, 299, 1020, 1719, 1792, 1980, 4618, 668, 99, 4667, 4676, 700, 453, 584, 2055, 276, 4497, 2255]. Statistics [3315, 2816, 587, 2307, 2314, 3799, 3149, 2447, 3317, 4504, 2211, 3322, 3321, 1882, 3319, 700, 3313, 2732, 3326, 2791, 2451, 3333, 3831, 4049, 3844, 218, 60, 1142, 1134, 1910, 1224, 14, 610]. Status [3324, 2144, 4485, 4128, 1612, 3545, 2611, 2688, 4368, 3114, 3762, 191]. Steady [2383, 318, 401]. Steady-State [2383]. Steering [3687, 868]. Stefan [4068]. Stellar [3486]. Step [3294, 1534]. Stephen [2624]. Steps [2596, 2791]. Stern [1733, 3264, 3386, 964, 2124]. Steuernagel [3173]. Steven [2375]. Stewart [1275, 1265, 1266, 1325]. Stiefel [1552, 1562]. Still [4506, 4091, 2680, 1050]. Stilwell [3306]. Stochastic [4012, 1025, 606, 610, 605, 616, 671, 668, 3679, 4716, 3413, 396, 501, 4088, 1015, 4505, 3997, 3851, 2483, 2089, 2982, 1242, 2654, 4220, 4575, 173, 2544, 2970, 1346, 370, 900, 912, 922, 587, 611, 1326, 450, 1105, 1279, 1446, 521, 580, 1476, 384, 2014, 373, 415, 1277, 188, 692, 1297, 293, 478, 503, 763, 729, 910, 1196, 318, 1600, 1801, 306, 518, 476, 1282]. stock [4627]. Stöckmann [2376].

Stokes [1346]. Stöltzner [2600]. Storage [2938, 3318]. Storage-Ring [3318]. Story [2839]. Strange [2839, 4384]. Strangeness [1835, 4492]. Strategy [3208, 702]. Stratification [3855]. stratifying [297]. Strength [3912, 829]. Stress [2943]. Stress-Tensors [2943]. stresses [398]. Strictly [4503, 61]. String [2658, 3729, 2386, 3611, 3625, 3617, 3613, 3614, 3615, 3891, 2606, 3616, 2371, 3606, 3814, 1636, 1017]. String-Theoretic [3891]. Strings [3029, 3179, 2926, 3607, 2371, 968]. Stringy [3330]. Stroboscopic [2473]. Strong [4544, 679, 2435, 3960, 2099, 3724, 123, 1669, 168, 664]. Strong-Field [2099, 679]. Stronger [3446, 2123]. Stronger-Than-Quantum [2123]. Strongly [3757, 2479, 3695, 3354]. Structural [2406, 1355, 4180, 575]. Structure [3687, 3704, 3547, 4233, 3334, 2746, 2790, 2815, 2812, 1614, 3231, 3371, 2702, 2587, 4426, 3830, 3223, 3439, 80, 2437, 2363, 4650, 3938, 2992, 3421, 2634, 4110, 609, 4644, 1886, 1684, 1473, 781, 1750, 604, 459, 1622, 1151, 583, 522, 564, 1518, 1080, 1973, 1156, 1767, 15, 2023, 37, 908, 1215, 994, 2565]. Structures [2392, 2282, 4609, 3380, 3619, 3798, 3378, 4339, 3023, 2908, 4386, 4344, 2811, 3670, 1683, 1582, 1784, 1620, 4639, 969, 1488, 1254, 971, 2497]. Students [2917]. Studies [2392, 7]. Study [1184, 1192, 3420, 2329, 383, 4366, 4336, 2557, 1284, 4644, 2013, 2023]. Stueckelberg [4006, 2162, 2719]. Stuff [4069]. Suarez [3206]. Sub [3586]. Sub-radiance [3586]. Subject [4325, 2680]. suboptic [1040]. Subsidiary [2947]. Subspaces [2980]. Substantivalism [2406]. Substantivalist [4237]. Subsystem [2957, 4719]. Subsystems [2059]. Success [1949]. Successive [3540, 3903, 1328]. such [4627]. Sudarshan [4719, 2241]. Sudbury [2831]. Suggestion [437]. Suhail [2242]. Sum [2479, 1556, 3172]. summary [1474]. Summer [2366, 2377]. sums [1681]. Super [2334, 3485, 4370, 3974, 3650, 3668, 115, 757]. Super- [4370]. Super-determinism [3974]. super-gauge [757]. super-ideal [115]. Super-Luminal [3650]. Superconducting [3520, 2980, 3791, 3782, 2462, 340]. Superconductivity [4645, 2438, 363, 359, 3748, 4618, 1964, 1922, 2244, 2785]. Superconductivity* [2912]. Superconductors [2060, 1531]. Superdeterminism [4565]. Superfluid [2650, 4147, 4446, 201, 256, 354, 2045]. Superfluids [2978]. supergravity [698, 728, 561]. Supergroups [917]. superheavy [2023]. Superloop [3864]. Superluminal [3583, 3113, 4714, 640, 2850, 2516, 3724, 2977, 2064, 3942, 635, 1035, 172, 526, 1174, 1600]. Supermassive [3998]. Supermeasured [4508]. Supermetrics [3323]. Supermultiplicity [2017]. Supernovae [2995]. Superoperators [2960]. superoptic [1033, 1040]. Superoscillations [4594, 4287]. Superposition [1236, 4400, 3417, 160, 596, 595, 1201]. Superpositions [3427, 4310, 3598, 3657, 1542]. Superrelativity [1974]. Superselection [3594, 376]. supersonics [105]. superspace [917]. Superstring [3610, 3609]. Superstrings [2691, 3260, 3719]. Supersymmetric [3864, 2914]. Supersymmetry [2691, 3782, 552, 561, 2375]. Supertask [2266]. Supertasks [3288].

Supervenience [3453]. Supplementarity [2929]. Supplementary [2647]. support [518]. Supports [2840, 919]. Surface [2532, 3055, 3116, 696, 1632]. Surfaces [1534, 3403]. Surfing [4632, 4647, 2315]. Surmounting [3395]. Surplus [4233]. Surprise [2177, 1869]. Surprising [3384, 4102, 1869]. Surrounding [3930]. Survey [520, 4096, 54, 3023, 1211]. Survive [4561, 4258]. Sussman [2565]. SuSy [1135]. Sutherland [978, 979]. Swampland [4513]. Swan [3726]. Swapping [2940, 4420]. Sylvia [2497]. Symmetric [3995, 4115, 3899, 4205, 1790, 4699, 3960, 4596, 4502, 3042, 1824, 1911, 4729, 1923, 1755, 2330]. Symmetrical [3651]. Symmetries [1343, 2790, 2815, 2827, 3775, 3406, 3101, 2910, 2083, 2439, 2153, 2706, 274, 2043, 280, 386, 2044, 2016, 385, 1348, 664]. Symmetrized [2238, 8]. Symmetry [3457, 4002, 2307, 2314, 3342, 3126, 3328, 782, 4144, 3327, 2270, 2636, 4135, 73, 3710, 2326, 3177, 3235, 3925, 3338, 2084, 3445, 4058, 4277, 4030, 3750, 2087, 3519, 3610, 1381, 2030, 3403, 3131, 3579, 2404, 4045, 174, 851, 1108, 1178, 133, 1042, 3748, 1160, 634, 255, 224, 341, 320, 395, 46, 1314, 535, 312, 1209, 1717, 2012, 2032, 240, 77, 101]. Symmetry-Based [4144]. Symmetry-breaking [174]. Symplectic [3634, 4381, 4379, 4533, 4471, 4470, 3209, 4375, 1582, 1997]. Symposium [1138, 1498, 1615, 1640, 1842, 1849, 1990, 1863, 1903, 1147, 1185, 1193, 1608]. Synaptic [3664]. Synchronization [2849, 4710, 4471, 4470, 611, 613, 614, 1366, 1114, 1957, 612]. Synchronizations [2175]. synchrony [378]. synthesis [4630, 4638, 4643, 4656]. Syntopic [321]. System [4014, 3794, 3241, 2145, 2711, 3231, 3327, 3816, 3031, 3508, 3303, 3036, 4013, 2792, 2059, 3836, 3961, 4497, 2142, 2557, 3938, 2108, 1346, 989, 1411, 1783, 1128, 782, 3417, 606, 610, 616, 1712, 273, 1502, 1019, 1923, 1776, 1509, 949, 934, 377]. Systematic [405, 942]. Systems [3702, 2898, 2683, 2367, 2378, 2390, 3661, 3888, 3088, 2301, 3337, 2470, 4678, 4681, 4400, 3689, 2449, 3824, 3520, 4318, 3484, 3681, 3351, 2135, 2295, 2513, 3823, 2473, 2555, 2760, 2761, 2798, 4573, 3222, 4348, 2991, 2170, 2979, 2662, 3336, 3413, 3530, 3895, 3806, 4405, 2765, 3604, 3378, 3414, 3541, 4044, 3741, 4617, 2914, 2459, 3094, 2764, 3462, 2496, 2805, 4063, 4408, 4257, 4084, 4092, 3551, 2182, 3432, 3856, 3346, 2965, 3269, 3714, 1695, 2050, 864, 1508, 4651, 4657, 4674, 1684, 821, 1205, 661, 1685, 573, 1766, 1144, 1268, 578, 397, 355, 280, 386, 1151, 920, 237]. systems [586, 2054, 2055, 1444, 1231, 385, 1188, 1220, 761, 1717, 2004, 1344, 173, 2766]. Szent [591]. Szilard [1787].

T [2867, 2784, 2316, 3488, 2043, 2446]. Tabletop [4542]. Tac [3509]. Tachyon [1204, 942, 647]. Tachyons [2658, 172, 105, 381, 1125, 317]. Tail [2281]. Takahashi [2009, 255]. Take [3644, 4236]. Tale [2499, 2103, 4623, 1961]. Taming [3757]. Tanaka [2795]. tangent [1380]. Tanya [4343]. Targeting [2935]. Targets [2848]. tautology [1887]. Taxonomy [4178]. Technologies [3745]. technology [324, 138]. Telegdi [3501]. Telegraph [2493]. Teleological [4421]. Teleparallel [2121, 4334, 1998]. Teleparallelism [2197, 1422, 1423]. Telepathy [2939]. Teleportation [2358, 3040, 2273, 3983]. Tell [4485, 1342]. Temperature [4645, 2656, 3396, 4696, 2785, 966, 1729]. Temporal [4072, 3638, 3024, 4433, 4539, 3378, 4596, 4254]. Tempore [2132]. ten [1104]. ten-dimensional [1104]. Tense [3740]. Tension [2649, 4268]. Tensor [4458, 2105, 4292, 3397, 3892, 98, 1556, 1602, 736, 1223, 461]. Tensor-Vector-Scalar [4458]. Tensorial [2200]. Tensors [2943, 3051, 504, 236]. Tentative [2587]. Term [3203, 3867, 2772, 2287, 2718, 4735, 376, 1189]. Terms [2744, 3837, 3787, 175, 481, 1534, 839]. Test [2724, 2399, 840, 2916, 3265, 501, 3601, 3662, 2462, 2840, 3706, 4257, 3526, 3316, 161, 1726, 1975, 759, 396, 288, 712, 655, 899, 908, 917, 1600]. Testable [3656, 521]. tested [403]. Testing [2222, 4556, 4550, 1116, 3485, 4185, 3558, 1201, 433, 538, 1604, 1104]. Tests [4542, 3217, 2985, 3327, 2580, 4005, 504, 2043]. tetrad [1086]. Tetrode [1421]. Tetron [3210, 3305]. TeV [3060]. TeVeS [3400]. Texture [4670]. TF [2699]. Their [3301, 2620, 1789, 4065, 3051, 4445, 1121, 2606, 2734, 2557, 1675, 710, 1730, 1525, 95, 508, 731, 1764, 278, 3074, 751, 2469, 375, 1829, 2002]. Them [2545]. themata [1866]. Theme [4478, 2229, 1964, 646]. Theological [1870]. Theology [3376]. Theorem [4388, 4131, 2503, 4016, 4072, 2489, 3230, 2942, 4734, 2415, 2414, 3113, 2190, 3947, 3804, 4718, 3341, 2932, 3892, 4047, 3195, 2739, 4088, 2352, 2950, 4395, 3964, 4599, 2732, 4140, 4229, 3831, 2428, 2426, 2427, 3988, 2955, 2956, 4523, 3044, 3801, 2143, 3653, 4382, 3500, 3759, 3092, 1907, 1269, 1887, 1768, 2009, 1714, 1342, 1288, 1214, 315, 1645, 302, 383, 4684, 1286, 1888, 792, 972, 1439, 1303, 2034, 910, 364, 793, 1292, 852, 540, 439, 4410, 4203, 4000, 4612, 3015, 3112, 4477, 3667, 3960, 4399, 2803, 4397, 3972, 3450, 3522, 3673, 3482]. Theorems [3217, 4107, 3707, 4487, 4412, 2791, 3904, 4171, 4245, 1588, 170]. Theoretic [2744, 3573, 4730, 4606, 3529, 3891, 3487, 3838, 218, 1989, 748, 276, 1412, 228]. Theoretical [2366, 2377, 2408, 2839, 4498, 3076, 4123, 2800, 1117, 4273, 4427, 3855, 2311, 2350, 1605, 2409, 635, 1035, 1165, 1921, 171, 1258, 178]. Theories [2387, 3687, 4100, 2330, 4178, 3690, 3239, 3339, 3583, 2416, 3701, 4079, 4545, 3659, 3984, 4266, 2165, 4714, 670, 3588, 4573, 2823, 2948, 2494, 3485, 4112, 3045, 3195, 3709, 2739, 4539, 2821, 4277, 3061, 4339, 3743, 4273, 4427, 2785, 3820, 4080, 2773, 2606, 3289, 2557, 2930, 4267, 3650, 3904, 4092, 2566, 4161, 3181, 4474, 3838, 1675, 4644, 745, 82, 710, 867, 1281, 1551, 213, 1282, 1521, 1317, 1996, 860, 1527, 1976, 841, 1158, 634, 1604, 980, 255, 244, 1088, 212, 1118, 1119, 1854, 1995, 846, 1963, 58, 2028, 1505, 465, 343, 757, 1511]. Theory [3702, 2724, 2789, 4538, 4506, 3784, 2535, 2375, 2334, 4020, 3473, 2843, 3200, 2571, 3910, 3886, 3564, 3931, 3982, 4280, 3704, 4095, 4262, 4181, 3761, 3088, 4605, 2714, 2890, 2301, 3729, 2602, 2937, 3847, 4340, 2888, 3342, 2082, 4101, 4678, 4681, 4189, 4535, 4359, 2449, 4038, 2901, 2319, 3929, 2895, 2744, 3308, 772, 3894, 3611, 2860, 4572, 3857, 3839, 4579, 3625, 3617, 2998, 2623, 3080, 3085, 3132, 4602, 2720, 3394, 4364, 4478, 4645, 2637, 2701, 4516, 2244, 4097, 4221, 3660, 3590, 3613, 2167, 3956, 4381, 2784, 3833, 2767, 3509, 3858, 2622, 2410, 4174, 2502, 3274, 3538]. Theory [4449, 2436, 4132, 4186, 3127, 3128, 4723, 2900, 4224, 3294, 3525, 3444, 4557, 2574, 2170, 2162, 2413, 2342, 2267, 2931, 2691, 4242, 4114, 3536, 3413, 3556, 3818, 4023, 4304, 4653, 3364, 3675, 3783, 4439, 3726, 3304, 2558, 3377, 4235, 2903, 4693, 3357, 501, 3443, 3970, 3103, 4094, 2521, 4358, 4553, 3366, 2989, 4521, 3964, 2994, 3319, 3615, 4566, 3542, 3997, 4389, 3767, 3808, 3353, 3390, 3677, 4333, 3539, 117, 4403, 4199, 2155, 2156, 3091, 4414, 3287, 2274, 2281, 4184, 2954, 3853, 3952, 2791, 3554, 4386, 3047, 4036, 2569, 4423, 3610, 4589, 4429, 2626, 3466, 4049, 3848]. Theory [2738, 3389, 3569, 2902, 4226, 4326, 3988, 2424, 4581, 3344, 3637, 3491, 4582, 4430, 4597, 3762, 281, 2531, 3013, 2350, 2782, 4001, 3307, 3616, 2137, 2576, 2454, 4534, 3421, 2766, 2238, 2280, 2705, 4060, 4196, 3082, 3299, 2440, 3507, 3606, 2260, 2544, 3641, 3609, 4706, 952, 1950, 433, 129, 296, 1455, 766, 989, 1972, 71, 342, 1271, 1701, 1733, 1643, 363, 813, 32, 67, 152, 517, 1304, 4613, 4651, 4657, 4674, 425, 176, 1946, 1285, 1239, 1789, 1895, 161, 782, 714, 715, 802, 817, 878, 936, 951, 1198, 1642, 599, 469, 549, 592, 781, 679, 2009, 1384]. theory [1385, 1702, 1770, 162, 1187, 1326, 137, 193, 1554, 1706, 1910, 279, 192, 4618, 42, 4666, 830, 1129, 1612, 1527, 85, 1682, 773, 553, 1708, 264, 197, 209, 233, 241, 135, 968, 1603, 1420, 1421, 856, 1467, 199, 575, 1641, 600, 1086, 12, 22, 36, 153, 4661, 533, 2005, 2043, 2052, 1536, 1846, 113, 127, 250, 1574, 1365, 1373, 1476, 17, 1576, 292, 396, 1974, 818, 659, 242, 352, 339, 388, 432, 866, 838, 488, 86, 1510, 697, 39, 373, 47, 128, 184, 291, 456, 461, 1010, 731, 345, 586, 203, 2055, 1691, 689, 1727]. theory [19, 1080, 691, 1231, 514, 278, 96, 445, 692, 704, 18, 556, 2026, 791, 1130, 1283, 558, 626, 1381, 168, 68, 76, 160, 455, 755, 2030, 964, 375, 491, 216, 1829, 409, 822, 1354, 364, 1055, 361, 1808, 1680, 149, 1417, 1336, 655, 899, 908, 917, 351, 186, 932, 1202, 454, 4665, 9, 65, 1049, 1199, 1321, 623, 629, 173, 515, 524, 472, 3029, 2090, 1905, 3814, 2785, 2687]. Theory/ [3029]. There [2286, 3789, 3879, 2737, 4475, 3494, 1283, 4303, 2676, 107, 484, 1403]. Thermal [3121, 3950, 2675, 4405, 4414, 4349, 869]. thermalizing [2020]. Thermo [3870]. Thermo-hydrodynamic [3870]. Thermodynamic [4077, 1016, 4386, 4283, 3706, 2234, 283, 823, 888, 1178]. Thermodynamical [2673, 2676]. Thermodynamically [3686]. Thermodynamics [3158, 2582, 133, 2295, 2421, 4297, 4438, 3338, 2662, 62, 189, 4682, 3603, 3901, 3479, 4414, 4334, 2653, 2640, 948, 1447, 197, 209, 233, 241, 56, 153, 400, 46, 4667, 4676, 276, 442, 282, 72, 892, 656, 4295, 3527, 3123, 4027, 3116, 2141, 3114, 4009, 2836]. Thermometry [3722]. thermophysics [116]. Thermostatics [153]. Thermostatistics [3325]. Thermosynthetic [3124]. These [4506, 2227, 2741]. Thing [3007]. Things [2742]. thinker [510]. Thinking [4599]. Third [1160, 2292, 4009]. Thomas [3540, 3903, 3501, 1179, 2002]. Thorne [2624]. thorough [942]. Thou [3298]. Thought [3512, 2215, 399, 1867, 1050]. thought-provoking [1050].

Thoughts [2338, 24, 1021]. Threat [4152]. Three [2724, 3088, 3180, 1054, 1982, 3823, 3948, 3341, 3798, 3017, 2248, 3502, 2835, 3655, 3405, 2764, 3558, 4529, 3421, 4382, 121, 665, 1961, 1298]. Three-Body [2724, 121]. Three-Box [2248]. Three-Channel [3405]. Three-Dimensional [3948, 3017, 665, 1298]. Three-Particle [3088, 2835, 2764]. Three-Slit [3655]. Three-Space [4529, 1054]. Threshold [2472, 557]. Thrust [2843]. thus [3825]. thus-Whether [3825]. Tic [3509]. Tic-Tac-Toe [3509]. Tier [4617, 2496]. Tilted [2463]. Time [458, 3733, 3143, 4006, 2164, 3244, 4062, 3612, 4002, 2775, 4109, 2553, 2861, 3995, 3950, 3644, 2486, 3361, 4053, 4712, 4359, 2815, 2704, 2974, 3899, 4205, 4658, 4318, 2673, 1547, 2919, 3946, 4133, 4509, 803, 3402, 4699, 122, 4227, 4721, 3816, 4490, 2210, 3990, 4437, 2587, 3874, 4541, 3480, 2916, 2167, 3956, 4501, 3849, 3323, 4379, 2760, 2761, 2798, 4670, 3588, 2510, 3682, 4713, 4557, 2574, 3068, 2710, 3150, 3925, 3363, 3795, 2425, 3318, 2817, 3530, 3928, 3585, 2401, 3207, 2851, 4046, 2765, 3090, 4539, 4168, 4697, 3385, 3836, 4475, 3285, 3330, 3957, 2527, 3800, 4585, 2053]. Time [4383, 2357, 3549, 4338, 2730, 4077, 1080, 2281, 3358, 87, 4625, 4633, 3010, 3642, 485, 3004, 2142, 3154, 3074, 3395, 3047, 4536, 4532, 4596, 2133, 2132, 4710, 3332, 4147, 4446, 4434, 2508, 4465, 2146, 2930, 2585, 1717, 3923, 2420, 2454, 3227, 4349, 4139, 2992, 3488, 3817, 2238, 4045, 4452, 2847, 3596, 3042, 2967, 3680, 3156, 2177, 4673, 4471, 2640, 240, 4110, 1731, 815, 530, 1656, 1907, 2008, 966, 1194, 1195, 599, 592, 1079, 679, 1702, 1804, 1165, 451, 1969, 967, 1521, 230, 249, 853, 578, 98, 222, 333, 1308, 747, 459, 634, 1869, 1077, 366, 395, 150]. time [55, 575, 3723, 106, 280, 386, 382, 1827, 1151, 1725, 1118, 1119, 242, 352, 388, 432, 1965, 84, 79, 183, 1777, 1516, 1451, 1618, 447, 643, 177, 1166, 146, 1019, 2055, 608, 1843, 1366, 385, 369, 293, 1130, 1320, 3481, 2042, 256, 368, 74, 1356, 908, 351, 420, 880, 656, 4206, 2899, 2315]. time-delayed [74]. Time-Dependent [2486, 2142, 2508, 2454, 803, 3480, 451, 1019]. Time-energy [485, 222, 382, 368]. Time-Evolution [3836]. Time-Reading [3874]. Time-Reversal [2974, 3925]. time-scale [1969]. Time-Symmetric [3899, 4205, 4596, 3042]. Time-Symmetrized [2238]. time-varying [256]. Timeless [3785, 4086, 4475]. Timelike [3531, 3989, 2919, 4046, 3358, 3194, 3551]. Timelines [3800]. Times [3081, 2150, 3966, 3245, 2463, 3282, 1696, 1726, 1215]. Timothy [2624]. Tip [3209]. Tired [2853]. Tl [3324]. Today [4164]. Toe [3509]. Tokamak [3472]. Told [3592]. Tolerance [1879]. toll [1294]. tomographic [950]. Tomography [3704, 3229, 2083, 1997]. Tomoyasu [2795]. Tony [2241]. Too [4239]. Tool [2128]. Tools [2971, 1228]. Top [2107, 1803]. Topics [3272]. Topoi [3003, 2811]. Topological [3517, 1552, 1562, 3782, 2975, 2973, 4735]. Topology [1631, 4735, 827, 714, 715, 802, 878, 936, 951, 1198, 1384, 1199, 436]. Topos [3777, 2413, 3024, 4439]. Toroid [4698]. Torretti [2498]. Torsion [2908, 3074, 3395, 2260, 1034, 747, 555, 1215]. Torsional [2068, 4269]. Torun [2378]. Torus [3962]. toss [464]. Total [3892, 4347, 2930]. Totally [4343]. tours [1247]. Townes [3791]. Toy [3886, 3554, 3102]. Trace [2468, 2483, 2462, 468]. Traces [3715]. Track [4362]. Tracks [3780]. Tradeoff [4465]. Tradition [3855]. Trajectories [4531, 2298, 2297, 4555, 4435, 4133, 3099, 961, 4557, 3658, 1073, 4010, 4505, 2146, 4707, 3257, 4060, 587, 1271, 1400, 1059, 1987, 1453]. Trajectory [3910, 3100, 3108, 2119, 3382, 4272, 1725]. Trajectory-Based [3108]. Trans [4497, 4503]. Trans-Planckian [4503]. Trans-statistical [4497]. Transactional [3627, 3885, 4478, 3581, 1742]. transactions [1214]. Transcendental [4091]. Transfer [2335, 3942, 2038]. transferred [1033, 1040]. Transform [2032, 1940]. Transformation [3674, 2896, 3802, 4043, 4112, 3892, 2581, 4536, 4603, 3237, 1552, 1562, 496, 1845, 1244, 1657, 911]. Transformations [3255, 3540, 3903, 2588, 2636, 2518, 2731, 2925, 2851, 4479, 3963, 2290, 283, 296, 1758, 349, 214, 640, 388, 237, 331, 493, 537, 1579, 1328, 1368, 190, 607]. Transient [820, 130, 123]. Transition [2434, 3343, 4511, 2594, 2595, 3232, 3076, 4094, 3060, 4729, 2862, 3410, 4471, 148, 206, 5, 1063]. Transitions [2188, 3180, 2355, 3898, 3576, 4323, 4405, 503, 616]. transitive [1715]. Translation [3207, 362]. Translations [2827]. Transmission [3353]. transmit [285]. Transmutation [3319]. Transparency [3791, 2097]. Transport [2465, 2502, 2274, 2281, 2753, 1758, 1810, 821]. Transpose [2941]. Transposition [149]. Transposition-invariant [149]. Transquantum [2726]. Trap [2092, 3752]. Trapped [2978]. trapping [1284]. Travel [2919, 3004, 2177, 635, 1035, 1869]. Traveler [2315, 4625, 4633]. Traveling [3021]. Traveling-Wave [3021]. Traversable [4553]. Traversable-Wormhole [4553]. Traversal [3733]. treated [339]. Treatment [2435, 4047, 228, 1787, 1070, 429]. Treder [1590]. Trends [2497, 4639]. Trial [3445]. triality [1833]. Trials [3737]. Triangle [2061]. Triangulations [4305]. Tribute [2867, 2077, 2968, 4357, 871, 142]. triggering [1655]. trigonometric [607]. Trigonometry [3425]. Triple [4286]. trivial [3782]. Troubles [3939]. Trouton [3069]. True [4393, 3597, 4052, 2518]. Trump [2494]. Truth [2408]. Trying [4485, 4468]. Tsirelson [3574]. Tumulka [3112]. Tuned [4406]. Tuning [4421, 4197, 3548, 4720, 4346]. Tunnel [118, 300]. Tunneling [2966, 1884, 2710, 3245, 1725, 3463, 2453, 2089, 2454, 3680, 1726, 1969, 1965]. tunnelling [1577]. turbulence [346]. Turbulent [2103, 4623]. Turn [3809]. Turning [3845]. Twentieth [4572]. Twenty [3745]. Twenty-First [3745]. Twin [3865, 1130, 1121, 2172, 2989, 794, 1190, 2067]. twin-slit [2172]. Twins [1826]. twistors [686]. Two [4483, 4151, 3431, 1539, 1090, 3088, 2774, 2775, 2395, 2301, 4143, 2092, 4167, 3813, 3605, 2825, 279, 3590, 3398, 3937, 2499, 3034, 3261, 3896, 3036, 4487, 2438, 4338, 4617, 4586, 2142, 2586, 2508, 3844, 2912, 4180, 2930, 2496, 2328, 7, 2140, 2151, 1961, 3653, 2103, 4198, 2095, 1647, 2318, 294, 677, 782, 1614, 1059, 1537, 1900, 1077, 1472, 1073, 1027, 1723, 1237, 184, 1613, 4623, 309, 729]. Two- [3088]. Two-Body [2774, 2775, 4338, 2328, 2140, 2151, 1961, 1539, 1614]. Two-Center [2508]. two-component [677]. Two-Dimensional [2438]. Two-Electron [3605, 3034]. Two-Level [2301, 3036, 294, 782, 309]. Two-Particle [3896, 2586, 4180, 1059, 1073, 184, 1613]. two-photon [1723]. two-point [1027]. two-prism [1900]. Two-Qubit [4487]. Two-Slit [2395, 3261, 4586, 2095, 1472]. Two-State [2142]. Two-State-Vector [4143]. Two-Tier [4617, 2496]. Twofold [4105]. Type [3189, 4016, 2457, 2770, 3947, 3372, 3664, 2471, 3190, 3235, 2271, 4235, 2094, 3791, 3835, 2955, 2956, 2363, 2097, 4171, 4245, 1678, 1681, 1383, 1443, 3290, 1088, 2016, 48, 1104, 1457, 2995, 3902]. Type-Decomposition [3372, 3664]. Type-I [3902]. Types [3823, 3798, 2930]. Typicality [3108].

U.K [2808]. U.S.A [2820, 2819]. Ubiquitous [3093]. Ultimate [2344]. Ultra [4573]. Ultra-Cold [4573]. Ultracold [3741]. Ultrafast [3741]. ultrametric [1909]. ultraviolet [1846]. Ulvi [569]. Umkehrung [1022]. unacceptable [1125]. unacknowledged [1294]. Unambiguous [3446]. Unavoidable [4311]. Unbiased [4167, 3023]. Unbounded [2665]. uncertain [146]. Uncertainty [4580, 4555, 4544, 734, 2386, 3754, 4067, 3561, 3934, 4130, 2587, 3979, 3879, 1312, 4481, 3792, 2296, 959, 3957, 3837, 3252, 4087, 3963, 790, 2234, 4164, 2317, 3209, 2318, 290, 204, 847, 1194, 1195, 4268, 1367, 382, 1462, 485, 368, 74, 893, 1457]. Unconceived [4075]. Unconditional [2043, 3850]. Unconnected [2834]. Unconstrained [2157, 1843]. Unconventional [3444]. Undecidability [3279]. Undecidable [3153]. Underdetermination [2890, 4363, 4267]. Underlying [2778, 3023, 570, 2002]. Underpinning [3040]. Understanding [4351, 4215, 4541, 2037, 2422, 3010, 4586, 4140, 1501, 2315, 1675, 152, 4632, 4647]. Understands [4343]. Understood [2273, 630]. Undesirable [3191]. undistorted [1983]. Undivided [3637]. Uneasy [3991]. unexcited [77, 101]. Unfamiliar [2653]. Unfinished [2968]. Ungar [2566]. Unidirectionality [3488, 3817]. Unification [4593, 2901, 1925, 3451, 3399, 3260, 2908, 3637, 3779, 2197, 681, 692, 1074, 4683, 3407]. Unified [4011, 1803, 4738, 2593, 2996, 222, 3127, 3128, 3150, 4156, 3902, 1152, 3091, 3837, 4326, 1417, 176, 1708, 197, 209, 233, 241, 255, 857, 1847, 1687, 626, 168, 755, 149, 83]. uniformities [590]. Uniformly [2905, 3059, 2572, 4396, 4260, 1219]. Unifying [3784, 2556, 645, 345]. Unimodular [4125]. Unique [4708, 366, 3309, 1226, 1233]. Uniqueness [3582, 2972, 3878, 537]. Unit [2568, 2552, 3174]. Unital [3572, 1223]. Unitarily [2079]. Unitarity [1515, 2979, 3940, 99]. Unitary [2724, 4242, 4404, 2835, 2892, 4300, 1655, 1644, 1846, 1579]. Unitary-Only [4242]. United [2376, 2389, 2590, 2795, 2784, 2836, 2541, 2498]. Unitless [2279]. Units [2553, 4736, 215]. Unity [2448, 17]. Universal [2147, 3945, 3231, 2212, 3792, 3960, 2271, 3585, 4461, 3915, 1488, 1320, 239]. Universality [2481, 1078, 2041]. Universe [4008, 4053, 4056, 3869, 4727, 4134, 3220, 3902, 4051, 4704, 4457, 4466, 2177, 181, 391, 1504, 1547, 965, 1160, 1869, 868, 1549, 637, 701, 318, 401, 694, 507, 381, 1206, 3143, 3066, 3469, 4143, 2704, 2607, 1753, 3527, 4659, 3437, 3032, 3455, 2995, 3364, 3309, 3287, 1358, 3062, 3637, 3134, 3011]. Universes [4151, 2315, 71, 688, 4632, 4647, 2209]. University [2377, 2375, 2376, 2388, 2389, 2590, 2795, 2784, 2836, 2499, 2767, 2541, 2539, 2677, 3010, 2540, 2808, 2498, 2617, 2766, 2315, 1001, 2240]. unknowable [4627]. Unknown [3080, 3085]. Unmeasurable [3544]. Unnormalizable [4093]. Unobserved [3012]. Unphysical [1563]. Unpredictability [3592]. Unreasonable [4499, 3497, 3426, 2219, 1548, 1949]. Unremarkable [1727]. Unresolved [2573]. Unruh [2905, 2278, 3734]. Unsharp [3360, 2282, 2416, 987, 1642, 1784, 2027, 1679, 1688, 1146]. Unsharpness [4175]. unsolved [1225]. Unspeakable [2942]. Unstable [2693, 4239, 2170, 4129, 1765, 1427, 1712]. Untangling [3041]. until [3247, 3350]. Unwrapping [3194]. Update [2217]. Upgraded [4258]. upon [526, 2925]. Upper [2901, 1397]. UPWs [1983]. urn [1245]. Ursell [2816]. USA [3010, 2967]. Usage [2248]. USD [2600]. Use [4242, 302, 2002, 911]. Used [2189, 2188, 2385, 4257]. Uses [3802]. Using [126, 2781, 4586, 3193, 2124, 3562, 2060, 3032, 1604, 1827, 1096, 437, 899, 1201, 623]. Usual [2897, 3017]. utilization [1301].

V [2245, 936, 2261]. V. [3437]. Vacua [2333]. Vacuum [2113, 2334, 2855, 4454, 3663, 2138, 3120, 2355, 3052, 3383, 4511, 1620, 3781, 2265, 3276, 2309, 2332, 2966, 3095, 3948, 2587, 2299, 2554, 2555, 4157, 4378, 4335, 2781, 4243, 3591, 2597, 2659, 2134, 4298, 3911, 3944, 2667, 2492, 3941, 1720, 412, 1160, 256, 174]. Vacuum-Destabilising [4454]. Vadis} [4566]. Vaidman [1778, 4332, 4031]. Valeri [2387]. Validity [3527, 2860, 2685, 82, 1637, 384]. Value [3980, 3676, 3092, 709, 1445, 266]. Valued [4483, 2400, 3237, 4470, 587, 543, 565, 1587, 1586]. valuedness [637]. Values [2889, 3993, 2466, 4032, 2072, 4447, 4155, 4341, 3562, 3880, 3991, 867, 1142, 963, 1793, 1851, 701, 4031]. vanishing [4709, 643]. Variable [3359, 3701, 3659, 3984, 4348, 3038, 2868, 3311, 2066, 2739, 3103, 3829, 3601, 2154, 4569, 4145, 2956, 4092, 469, 762, 212, 659, 554, 217, 2028, 910]. Variable-Mass [2154]. Variables [4204, 2971, 3485, 3040, 2929, 3895, 4440, 3436, 3549, 2862, 2954, 2560, 3074, 3145, 2955, 3082, 2369, 2260, 4329, 1886, 238, 1996, 338, 801, 1327, 1645, 879, 2030, 1397]. Variance [3809, 3833]. Variant [2279]. variants [1947]. variation [198]. Variational [2434, 2074, 182, 72, 570, 941]. Variations [646, 2025, 2229, 4478, 1964]. Varied [3453]. Variety [3917, 1157]. Various [2215, 274]. Varying [2826, 4588, 256]. Vassallo [4039]. Vecchia [3814]. Vector [4458, 3340, 3189, 2511, 4143, 2806, 4527, 3538, 3311, 53, 110, 3708, 3997, 1587, 927, 3507, 175, 1512, 399, 486, 1263, 1264, 231, 1056, 498, 174]. Vector-spinor [927]. Vector-valued [1587]. Vectorial [3540, 607]. Vectors [2323, 3692, 314]. Vedral [3437]. velocities [1040, 41]. Velocity [1079, 2856, 2850, 2110, 2067, 2324, 2566, 1033, 378, 412, 336, 335, 471, 302, 305, 1880, 426, 1179, 43, 911]. verification [452, 1732]. Verlag [2344]. Version [2584, 2223, 3111, 4088, 4225, 1022, 1421, 220, 251]. Versions [3398, 4294, 4413, 1790, 698, 728]. Versus [3022, 4203, 3739, 4039, 2408, 4307, 4046, 3829, 4218, 2409, 1539, 1195, 1677, 2895, 1507, 1930, 2402, 3017, 3246, 2035, 81, 3474, 1049, 505, 2071]. Vertex [3449]. Vervoort [3971]. Very [4248]. Vessot [899]. VI [951]. Via [2421, 3674, 1232, 840, 3427, 3656, 1936, 4293, 3323, 3870, 1552, 1562, 4694, 1109, 4254, 3142, 3532, 105, 4334, 4702]. Viable [4592, 4588, 3401]. Vibrations [2452]. Vienna [1001]. View [4262, 4259, 4354, 4509, 3672, 2857, 3281, 3602, 3827, 2881, 2516, 4644, 95, 214, 1021, 1166, 1130, 794]. viewed [1242]. Viewpoint [2929, 1939, 633, 869]. Viewpoints [4301]. views [279, 1001]. Vigier [1736, 1710, 979]. vii [2375, 2499]. viii [2967]. Villars [1802]. Violate [3850]. Violating [4508, 3912]. Violation [2394, 3320, 3409, 2236, 2296, 4150, 3330, 3724, 3098, 3488, 3817, 1982, 891, 375]. Violations [3191, 3131]. VIP [3320, 3409]. virial [852]. Virtual [2568, 2552, 2554, 3160, 3273, 3352, 4110, 2000, 285]. Viscous [4494, 2908]. Visibility [3261]. Visible [549]. Vision [4478]. visit [401]. Visited [2656]. Visualizability [3843]. visualized [2004]. Voigt [223, 880]. Vol [2387, 2785, 3113]. Volume [2075, 2104, 2117, 2131, 2139, 2149, 2161, 2171, 2179, 2185, 2196, 2304, 2194, 2204, 2220, 2232, 2246, 2256, 2262, 2268, 2277, 2284, 2293, 2306, 2442, 2379, 2375, 2368, 2391, 2419, 2430, 2443, 2548, 2651, 2758, 2669, 2678, 2689, 2703, 2709, 2716, 2725, 2735, 2743, 2751, 2759, 2769, 2776, 2787, 2796, 2802, 2809, 2813, 2822, 2829, 2837, 2891, 2510, 3818, 4304, 2558, 2862]. Volumes [2327]. Vortex [4576, 1858, 2045]. vortex-dynamics [1858]. Vortices [2060]. Vorticity [2428, 2426, 2034, 3944]. Vries [3021]. vs [609, 772, 3108, 4331, 4570, 3213, 3494, 27, 3368, 2983]. VSL [3151].

W [3126, 2494, 2624]. Walk [2740, 1618]. Walker [2345, 3869, 2729, 4121]. Walks [4022, 2952]. Wallace [4090, 4142]. Walls [2421, 3782]. Want [4119]. War [2253]. Ward [2009, 255, 2044]. Warped [3935]. Was [3386, 2941]. watched [227]. Watching [4526, 2047]. Wave [3048, 4538, 3863, 3189, 2479, 3910, 2147, 4605, 2125, 3945, 4393, 2433, 4485, 3078, 3089, 3133, 3219, 2098, 2711, 1741, 3021, 1465, 4572, 3721, 4579, 2339, 1475, 3956, 2574, 2166, 2525, 3811, 225, 300, 3889, 2058, 3649, 3792, 4302, 3262, 4046, 4459, 2357, 4530, 2118, 3803, 3504, 3212, 2730, 3560, 4366, 4336, 3943, 4089, 3289, 2424, 4707, 4534, 2873, 4092, 664, 51, 1033, 1040, 516, 517, 1719, 1437, 1474, 1072, 803, 574, 1345, 172, 78, 198, 1637, 246, 1203, 1341, 1009, 1900, 521, 332, 69, 134, 1023, 1453, 511, 1006, 304, 1015, 1000, 1005, 1014]. wave [2000, 987, 983, 2053, 1419, 653, 1258, 8, 506, 1012, 2001, 1993, 1298, 1068, 625, 3, 1646, 512, 4610]. Wave-corpuscle [1741]. Wave-Emitting [2711]. Wave-Function [4366, 4336, 1419]. Wave-Like [3212]. Wave-Particle [4459, 1465, 1475, 1474, 1900, 2000, 983, 625]. wave-theoretical [1258]. Waveforms [2471]. Wavefunction [2740, 4557, 2831, 3896, 2216, 3983, 2380, 1731, 1790]. Wavefunctions [4605, 2469, 4145]. Wavelength [2096]. Wavelet [2605]. wavelets [1271, 1733, 40, 118, 390]. Waves [3106, 2841, 4109, 2650, 2853, 4562, 4568, 4561, 2998, 4709, 2911, 4697, 2969, 4585, 2100, 2060, 2320, 1271, 678, 1835, 992, 947, 126, 1882, 2000, 1983, 523, 694, 1298, 1600, 718]. Way [3158, 3200, 2846, 2856, 4010, 2067, 2562, 3650, 378, 921, 471, 1880, 3998]. Weak [2571, 3687, 3993, 1450, 4032, 3183, 3839, 3809, 3980, 4447, 4003, 2881, 4155, 4736, 1159, 3562, 3676, 3880, 3991, 4031, 3543, 1896, 459, 1453, 498]. Weak-measurement [1896]. Weatherall [4201]. Weber [2820, 2373, 2033]. Websites [2816]. Wedges [844]. Weg [3100]. Wegner [3534]. Weighting [3117]. weights [1223]. Weinberg [2375]. Weirdness [4410]. Welcher [3100]. Well [2336, 3825, 1966, 91, 2453, 2290]. Well-Dressed [2336]. well-posed [91]. wells [451, 2463]. Were [2846]. Werner [959]. Wesson [2931, 3111]. Weyl [2914, 3648, 3267, 4062, 3852, 4022, 715, 1426, 3052, 4018, 4292, 2801, 1187, 2258, 2270, 3599, 1088, 1425, 1876, 2068, 2267, 2564, 2598, 2933, 3111, 3074, 3524, 463, 558, 3201, 3890]. Weyl-type [1088]. Weylian [1659, 2931]. Wheeler [835, 842, 873, 2358, 862, 871, 861, 838, 4403, 1357, 3344, 826, 870, 768]. Whence [4548, 4113]. Where [1966, 1524, 2490]. Wherefore [4113]. Whether [2916, 3825, 1345]. Which [2144, 2225, 4251, 3548, 4010, 2132, 4691, 2839, 30, 1223]. Which-Way [4010]. while [1949]. Whiteheadian [1069, 364]. Whither [4113]. Who [2952]. Whole [3364]. widths [177]. wiggly [1764]. Wightman [1908]. Wigner [544, 4183, 529, 950, 60, 4354, 843, 4312, 497, 3874, 3294, 1908, 4499, 4404, 534, 1453, 4495, 383, 536, 3497, 2984, 1085, 1439, 1063, 4001, 4401, 1657, 2944]. Wignerian [646]. Wilczyński [484]. Will [4131, 3015, 3112, 2299, 3044, 863]. Willard [2183]. Willem [2708]. Williams [2344, 2283]. WIMP [3208]. Winding [2261]. Wire [2237]. Wires [2373]. Wisdom [2565]. Within [2780, 4475, 4365, 4707, 4502, 2280, 4143, 1042, 1711, 320, 2953, 372, 3047, 1173]. Without [2535, 1, 3665, 3916, 4053, 4107, 3470, 4477, 4508, 2995, 3841, 2448, 3447, 3416, 4475, 4010, 3983, 2748, 3854, 2439, 4534, 4139, 3942, 3250, 2371, 2967, 4673, 3770, 3503, 525, 3182, 226, 1440, 3230, 2942, 109, 862, 1180, 1342, 1369, 1124, 1976, 909, 427, 2175, 1764, 2155, 1687, 358, 2099, 2067, 2955, 2956, 1068, 424, 849]. Witnessing [3559]. Witten [1621]. WKB [1019]. Wolf [2389]. Wolfgang [2677, 392]. Wonderful [3482, 4684]. Woodhouse [3282]. Woodward [2309]. Wootters [4526, 1472, 3889]. Work [4394, 2936, 397, 2480, 2383, 1047, 1266, 1610, 1000]. working [165]. Works [3546, 1736]. Workshop [1616, 1630, 2367, 2378, 2390, 3747]. World [2972, 3458, 2963, 4275, 2631, 3536, 2911, 2401, 2253, 2615, 2561, 3732, 2015, 4350, 4077, 3020, 2794, 560, 4001, 4673, 3477, 3296, 609, 4619, 314, 1711, 4515, 355, 496, 1407, 4623, 327, 2103, 2101, 2967]. World-Line [4350]. Worldlines [4609, 4251]. Worlds [2715, 3248, 3471, 2065, 4137, 3221, 1464, 1911, 1986, 1358]. Wormald [455]. Wormhole [4553, 4402]. Wormhole-Mediated [4402]. Wormholes [4157, 3914, 4365]. Written [2941]. Wrong [2116, 2868, 3366, 2951, 2741].

X [1947, 2687, 1799, 3893, 2494, 2754]. X-ray [2754, 1799]. X-rays [3893]. xi [2376, 2794]. xii [2539, 2819]. xiii [2641, 2786, 2899, 2677]. xiv [2344, 2808, 2496]. xix [2785]. xlii [2566]. xlviii [2903]. xv [2795, 2767, 3010]. xvi [2836, 2499, 2497, 2498, 2429]. xvii [2495, 2565]. xviii [2783, 3011, 2315]. xxi [2590, 2766]. xxii [2375, 2820]. xxiv [2708, 2617, 2315]. xxix [2784]. xxv [2541]. xxxii [2631]. xxxvi [2558].

Y-product [1947]. Yamaleev [3919]. Yang [2334, 4340, 2043, 698, 728, 3227, 1528, 2359]. Years [4506, 4422, 3389, 3606, 1001, 1575]. Yes [4040]. Yield [4404, 4341]. York [2344, 2375, 2388, 2641, 2495, 2558, 2539, 2677, 2239, 3011, 2617, 2475, 2766, 2315, 2624, 2819, 2967]. Young [2239, 2240, 2252, 2253, 2261, 3100]. Yourgrau [2967, 392]. Yu [2102]. Yukawa [205]. Yurtsever [569].

Zacatecas [2377]. Zadeh [2417, 1234]. Zee [2767]. Zeilinger [2223, 1299, 4348, 1604, 3924, 4495, 4431, 2229]. Zeno [2358, 2050, 2980, 4494, 2049, 2047]. Zermelo [103]. Zero [2344, 3342, 3547, 3927, 2148, 2966, 2554, 2555, 3797, 2192, 2953, 3448, 2492, 3269, 1108, 1523, 606, 610, 605, 616, 4396]. Zero-Energy [2554, 2555]. zero-mass [1523]. Zero-Point [3342, 3547, 3927, 2192, 3448, 2966, 3269, 1108, 605]. zeropoint [1660, 1646, 1751]. Zeros [2484]. Zeroth [4297]. zigzagging [979, 1061]. Zitterbewegung [1582, 4605, 1270, 1513, 3277, 217, 3579, 1994, 1993, 3474]. Zn [4323]. zones [1063]. Zubairy [2242]. Zukowski [3206, 4725]. Zur [3000, 3049]. Zurek [1472, 3889, 2869].

References

Anonymous:1970:A

[1] Anonymous. Without apologies. Foundations of Physics, 1(1):1, ???? 1970. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

      URL http://link.springer.com/article/10.1007/BF00708648.

Margenau:1970:EP

[2] Henry Margenau and Wolfgang Yourgrau. Editorial preface. Foundations of Physics, 1(1):1–3, ???? 1970. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

      article/10.1007/BF00708649.

deBroglie:1970:RWM

[3] Louis de Broglie. The reinterpretation of wave mechanics. Foundations of Physics, 1(1):5–15, ???? 1970. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00708650.

Bergmann:1970:CS

[4] Peter G. Bergmann. Cosmology as a science. Foundations of Physics, 1(1):17–22, ???? 1970. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00708651.

Park:1970:CTQ

[5] James L. Park. The concept of transition in quantum mechanics. Foundations of Physics, 1(1):23–33, ???? 1970. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00708652.

Wigner:1970:PEL

[6] Eugene P. Wigner. Physics and the explanation of life. Foundations of Physics, 1(1):35–45, March 1970. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://

link.springer.com/article/10.1007/BF00708653;   http://www.
springerlink.com/content/p0453n1p7835xt66.

Tornebohm:1970:TSC

[7] Håkan Törnebohm. Two studies concerning the Michelson–Morley experiment. Foundations of Physics, 1(1):47–56, ???? 1970. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708654.

Siegel:1970:DEP

[8] Armand Siegel. Distinguishable equivalent particles with symmetrized wave functions. Foundations of Physics, 1(1):57–68, ???? 1970. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708655.

Zeh:1970:IMQ

[9] H. D. Zeh. On the interpretation of measurement in quantum theory. Foundations of Physics, 1(1):69–76, ???? 1970. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00708656.

Treder:1970:GLP

[10] Hans-J. Treder. Global and local principles of relativity. Foundations of Physics, 1(1):77–94, ???? 1970. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00708657.

Heelan:1970:CCD

[11] Patrick A. Heelan. Complementarity, context dependence, and quantum logic. Foundations of Physics, 1(2):95–110, ???? 1970. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708721.

Janossy:1970:NAT

[12] L. Jánossy. A new approach to the theory of relativity. Foundations of Physics, 1(2):111–131, ???? 1970. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00708722.

Band:1970:EDQ

[13] William Band and James L. Park. The empirical determination of quantum states. Foundations of Physics, 1(2):133–144, ???? 1970. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708723.

Siegel:1970:CAQ

[14] Armand Siegel. On the classical approximation in the quantum statistics of equivalent particles. Foundations of Physics, 1(2):145–171, ???? 1970. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00708724.

Ruttimann:1970:LSQ

[15] G. T. Rüttimann. On the logical structure of quantum mechanics. Foundations of Physics, 1(2):173–182, ???? 1970. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00708725.

Jauch:1970:RA

[16] J. M. Jauch. Review article. Foundations of Physics, 1(2):183–189, ???? 1970. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00708726.

Lande:1971:UQT

[17] Alfred Landé. Unity in quantum theory. Foundations of Physics, 1 (3):191–202, ???? 1971. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00708606.

Rayski:1971:LCF

[18] Jerzy Rayski. Limitations of the concept of free fields in Einstein’s theory of gravitation. Foundations of Physics, 1(3):203–209, ???? 1971. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708607.

Park:1971:GTE

[19] James L. Park and William Band. A general theory of empirical state determination in quantum physics: Part I. Foundations of Physics, 1 (3):211–226, ???? 1971. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00708608.

Gruber:1971:QGC

[20] Gary R. Gruber. Quantization in generalized coordinates. Foundations of Physics, 1(3):227–234, ???? 1971. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00708609. See erratum [93].

Cochran:1971:RBQ

[21] Andrew A. Cochran. Relationships between quantum physics and biology. Foundations of Physics, 1(3):235–250, ???? 1971. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708610.

Janossy:1971:NAT

[22] L. Jánossy. A new approach to the theory of relativity. II. The general theory of relativity. Foundations of Physics, 1(3):251–267, ???? 1971. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00708611.

Wisdom:1971:FCI

[23] J. O. Wisdom. Four contemporary interpretations of the nature of science. Foundations of Physics, 1(3):269–284, ???? 1971. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708612.

Mercier:1971:TDF

[24] André Mercier. Thoughts on the dynamics of foundations, or what I believe. Foundations of Physics, 1(3):285–292, ???? 1971. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708613.

Anonymous:1971:IA

[25] Anonymous. Important announcement. Foundations of Physics, 1 (3):i, ???? 1971. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00708605.

Fock:1971:QPP

[26] Vladimir Fock. Quantum physics and philosophical problems. Foundations of Physics, 1(4):293–306, ???? 1971. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00708579.

Salt:1971:PAF

[27] David Salt. Physical axiomatics: Freudenthal vs. Bunge. Foundations of Physics, 1(4):307–313, ???? 1971. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00708580.

Freudenthal:1971:MAF

[28] Hans Freudenthal. More about foundations of physics. Foundations of Physics, 1(4):315–323, ???? 1971. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00708581.

Shimony:1971:FIM

[29] Abner Shimony. Filters with infinitely many components. Foundations of Physics, 1(4):325–328, ???? 1971. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00708582.

Kobus:1971:SMN

[30] J. P. Kobus and M. Z. Nashed. A simple model for nuclear forces which exhibits bound states. Foundations of Physics, 1(4):329–337, ???? 1971. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00708583.

Band:1971:GME

[31] William Band and James L. Park. A general method of empirical state determination in quantum physics: Part II. Foundations of Physics, 1(4):339–357, ???? 1971. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00708584.

Bohm:1971:QTI

[32] D. Bohm. Quantum theory as an indication of a new order in physics. Part A. The development of new orders as shown through the history of physics. Foundations of Physics, 1(4):359–381, ???? 1971. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708585.

Lindsay:1971:CEE

[33] R. B. Lindsay. The concept of energy and its early historical development. Foundations of Physics, 1(4):383–393, ???? 1971. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708586.

Goldman:1971:MIN

[34] Stanford Goldman. The mechanics of individuality in nature. Foundations of Physics, 1(4):395–408, ???? 1971. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00708587.

Beltrametti:1972:QMA

[35] E. G. Beltrametti and G. Cassinelli. Quantum mechanics and p-adic numbers. Foundations of Physics, 2(1):1–7, January 1972. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708614.

Janossy:1972:NAT

[36] L. Jánossy. A new approach to the theory of relativity. III. Problem of the ether. Foundations of Physics, 2(1):9–25, January 1972. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708615.

Tomozawa:1972:MPM

[37] Yukio Tomozawa. Mach’s Principle, mass, and the fine structure constant. Foundations of Physics, 2(1):27–34, January 1972. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708616.

Synge:1972:SCS

[38] J. L. Synge. A special class of solutions of the Schrödinger equation for a free particle. Foundations of Physics, 2(1):35–40, January 1972. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708617.

Moldauer:1972:RNT

[39] P. A. Moldauer. A reinterpretation of von Neumann’s theory of measurement. Foundations of Physics, 2(1):41–47, January 1972. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708618.

Koga:1972:MWI

[40] Toyoki Koga. The motion of wavelets — an interpretation of the Schrödinger equation. Foundations of Physics, 2(1):49–78, January 1972. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00708619.

Gruber:1972:PVF

[41] Gary R. Gruber. Particle velocities faster than the speed of light. Foundations of Physics, 2(1):79–82, January 1972. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00708620.

Fujiwara:1972:QTS

[42] Izuru Fujiwara. Quantum theory of state reduction and measurement. Foundations of Physics, 2(2–3):83–110, April 1972. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00708495.

deBeauregard:1972:NVM

[43] Olivier Costa de Beauregard. Noncollinearity of velocity and momentum of spinning particles. Foundations of Physics, 2(2–3):111–127, April 1972. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00708496.

Freundlich:1972:MMP

[44] Yehudah Freundlich. Mind, matter, and physicists. Foundations of Physics, 2(2–3):129–148, April 1972. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00708497.

Biedenharn:1972:KMQ

[45] L. C. Biedenharn, R. Y. Cusson, M. Y. Han, and J. D. Louck. A kinematical model for quarks and hadrons. Foundations of Physics, 2(2– 3):149–159, April 1972. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00708498.

Lavenda:1972:CSS

[46] B. H. Lavenda. Concepts of stability and symmetry in irreversible thermodynamics. I. Foundations of Physics, 2(2–3):161–179, April 1972. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00708499.

Muraskin:1972:SPA

[47] M. Muraskin. Some properties of an “aesthetic” field theory. Foundations of Physics, 2(2–3):181–188, April 1972. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00708500.

Smith:1972:MAE

[48] Gerrit J. Smith. Measurability analysis of the electric-type components of the linearized gravitational radiation field. Foundations of Physics, 2 (2–3):189–222, April 1972. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00708501.

Omelianovsky:1972:POM

[49] M. E. Omelianovsky. On the principle of observability in modern physics. Foundations of Physics, 2(2–3):223–239, April 1972. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00708502.

Woodward:1972:PCF

[50] James F. Woodward, André de Klerk, Gail Kahler, Kathrine Leber, Peter Pompei, Daniel Schultz, and Sharon Stern. Photon consciousness: Fact or fancy? Foundations of Physics, 2(2–3):241–244, April 1972. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708503.

AndradeeSilva:1972:PSM

[51] João Andrade e Silva. Pure states, mixtures, and reduction of the wave packet. Foundations of Physics, 2(4):245–256, October 1972. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708410.

Avishai:1972:CI

[52] Y. Avishai and H. Ekstein. Causal independence. Foundations of Physics, 2(4):257–270, October 1972. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00708411.

Lanczos:1972:VPQ

[53] C. Lanczos. Vector potential and quadratic action. Foundations of Physics, 2(4):271–285, October 1972. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00708412.

Piron:1972:SGQ

[54] C. Piron. Survey of general quantum physics. Foundations of Physics, 2 (4):287–314, October 1972. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00708413.

Hwang:1972:NIT

[55] Sun-Tak Hwang. A new interpretation of time reversal. Foundations of Physics, 2(4):315–326, October 1972. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00708414.

Jauch:1972:NFE

[56] J. M. Jauch. On a new foundation of equilibrium thermodynamics. Foundations of Physics, 2(4):327–332, October 1972. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00708415.

Anonymous:1973:IA

[57] Anonymous. Important announcement. Foundations of Physics, 3 (1):1, March 1973. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00708595.

Prugovecki:1973:PFT

[58] Eduard Prugovecki. A postulational framework for theories of simultaneous measurement of several observables. Foundations of Physics, 3(1):3–18, March 1973. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00708596.

Margenau:1973:PSQ

[59] Henry Margenau and James L. Park. The physics and the semantics of quantum measurement. Foundations of Physics, 3(1):19–28, March 1973. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00708597.

Bub:1973:PPS

[60] Jeffrey Bub. On the possibility of a phase-space reconstruction of quantum statistics: a refutation of the Bell–Wigner locality argument. Foundations of Physics, 3(1):29–44, March 1973. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00708598.

Kobus:1973:SGI

[61] J. P. Kobus. A strictly geometric interpretation of gravitation in general relativity. Foundations of Physics, 3(1):45–51, March 1973. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708599.

Lavenda:1973:TNI

[62] B. H. Lavenda. Thermodynamics of nonlinear, interacting irreversible processes. II. Foundations of Physics, 3(1):53–88, March 1973. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708600.

Rayski:1973:PMR

[63] Jerzy Rayski. The possibility of a more realistic interpretation of quantum mechanics. Foundations of Physics, 3(1):89–100, March 1973. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00708601.

Finkel:1973:GSQ

[64] Robert Warren Finkel. Generalized Schrödinger quantization. Foundations of Physics, 3(1):101–108, March 1973. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00708602.

Zeh:1973:TQT

[65] H. D. Zeh. Toward a quantum theory of observation. Foundations of Physics, 3(1):109–116, March 1973. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00708603.

Elsasser:1973:NPQ

[66] Walter M. Elsasser. A natural philosophy of quantum mechanics based on induction. Foundations of Physics, 3(1):117–137, March 1973. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708604.

Bohm:1973:QTI

[67] David Bohm. Quantum theory as an indication of a new order in physics. B. Implicate and explicate order in physical law. Foundations of Physics, 3(2):139–168, June 1973. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00708436.

Schlegel:1973:IISa

[68] Richard Schlegel. An interaction interpretation of special relativity theory. Part I. Foundations of Physics, 3(2):169–184, June 1973. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708437.

Janossy:1973:PIW

[69] L. Jánossy. The physical interpretation of wave mechanics. I. Foundations of Physics, 3(2):185–202, June 1973. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00708438.

Goldman:1973:MIN

[70] Stanford Goldman. The mechanics of individuality in nature. II. Barriers, cells, and individuality. Foundations of Physics, 3(2):203–228, June 1973. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00708439.

Ballentine:1973:CSP

[71] L. E. Ballentine. Can the statistical postulate of quantum theory be derived? — a critique of the many-universes interpretation. Foundations of Physics, 3(2):229–240, June 1973. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00708440.

Szamosi:1973:VPT

[72] G. Szamosi. Variational principle in thermodynamics. Foundations of Physics, 3(2):241–246, June 1973. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00708441.

Greenspan:1973:SDM

[73] D. Greenspan. Symmetry in discrete mechanics. Foundations of Physics, 3(2):247–253, June 1973. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00708442.

Tefft:1973:RTD

[74] Wayne E. Tefft. The relevance of time-delayed coincidence Mössbauer experiments for the interpretation of the uncertainty principle. Foundations of Physics, 3(2):255–263, June 1973. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00708443.

Strauss:1973:LQM

[75] Martin Strauss. Logics for quantum mechanics. Foundations of Physics, 3(2):265–276, June 1973. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00708444.

Schlegel:1973:IISb

[76] Richard Schlegel. An interaction interpretation of special relativity theory. Part II. Foundations of Physics, 3(3):277–295, September 1973. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00708672.

vanderMerwe:1973:DIB

[77] Alwyn J. van der Merwe and Johannes H. van der Merwe. Dispersion interactions between unexcited molecules possessing axial symmetry: arbitrary-order contributions. Foundations of Physics, 3(3):297– 311, September 1973. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00708673.

Edmonds:1973:GCE

[78] James D. Edmonds, Jr. Generalized charge in the eight-component spin-1/2 wave equation. Foundations of Physics, 3(3):313–319, September 1973. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00708674.

Mirman:1973:CDT

[79] R. Mirman. Comments on the dimensionality of time. Foundations of Physics, 3(3):321–333, September 1973. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00708675.

Kantorovich:1973:SHM

[80] Aharon Kantorovich. Structure of hadron matter: Hierarchy, democracy, or potentiality? Foundations of Physics, 3(3):335–349, September 1973.

CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00708676.

Prokhovnik:1973:CVR

[81] S. J. Prokhovnik. Cosmology versus relativity — the reference frame paradox. Foundations of Physics, 3(3):351–358, September 1973. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00708677.

Benioff:1973:DVA

[82] Paul Benioff. On definitions of validity applied to quantum theories. Foundations of Physics, 3(3):359–379, September 1973. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708678.

Wallace:1973:UQE

[83] Bryan G. Wallace. The unified quantum electrodynamic ether. Foundations of Physics, 3(3):381–388, September 1973. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00708679.

Miller:1973:IT

[84] G. Miller. The inexactness of time. Foundations of Physics, 3(3): 389–398, September 1973. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00708680.

Gudder:1973:GMT

[85] Stanley Gudder. Generalized measure theory. Foundations of Physics, 3(3):399–411, September 1973. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00708681.

Mayants:1973:PTP

[86] L. S. Mayants. On probability theory and probabilistic physicsaxiomatics and methodology. Foundations of Physics, 3(4):413–433, December 1973. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00709111.

Phipps:1973:TAQ

[87] T. E. Phipps, Jr. Time asymmetry and quantum equations of motion. Foundations of Physics, 3(4):435–455, December 1973. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00709112.

Freedman:1973:BMB

[88] S. Freedman and E. Wigner. On Bub’s misunderstanding of Bell’s locality argument. Foundations of Physics, 3(4):457–458, December 1973. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00709113;
http://www.springerlink.com/content/r17r1242148r3870.

Waahlin:1973:POA

[89] L. Wåhlin. A possible origin of atmospheric electricity. Foundations of Physics, 3(4):459–472, December 1973. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00709114.

Allen:1973:BPF

[90] A. D. Allen. The bootstrap from the perspective of formal logic. Foundations of Physics, 3(4):473–475, December 1973. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00709115.

Jaynes:1973:WPP

[91] E. T. Jaynes. The well-posed problem. Foundations of Physics, 3(4): 477–492, December 1973. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00709116.

Aharonov:1973:QAE

[92] Y. Aharonov and G. Carmi. Quantum aspects of the equivalence principle. Foundations of Physics, 3(4):493–498, December 1973. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00709117.

Landau:1973:RCP

[93] B. V. Landau. Rejoinder: Comments on a paper by Gruber. Foundations of Physics, 3(4):499–500, December 1973. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00709118. See [20].

Finkel:1973:EGS

[94] R. W. Finkel. Erratum: Generalized schrödinger quantization. Foundations of Physics, 3(4):501, December 1973. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00709119.

Freundlich:1974:OTA

[95] Yehudah Freundlich. Objects and their attributes: a physicist’s point of view. Foundations of Physics, 4(1):1–8, March 1974. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00708549.

Prugovecki:1974:FST

[96] E. Prugovecki. Fuzzy sets in the theory of measurement of incompatible observables. Foundations of Physics, 4(1):9–18, March 1974. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708550.

Gruber:1974:CCP

[97] Gary R. Gruber. Comments on the correspondence principles of quantum mechanical operators. Foundations of Physics, 4(1):19–22, March 1974. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00708551.

Gleyzal:1974:RTC

[98] André Gleyzal. Relative tensor calculus and the tensor time derivative. Foundations of Physics, 4(1):23–30, March 1974. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00708552.

Kyrala:1974:SRC

[99] A. Kyrala. Selection rules, causality, and unitarity in statistical and quantum physics. Foundations of Physics, 4(1):31–51, March 1974. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00708553.

Kobus:1974:CIN

[100] J. P. Kobus. Cosmological implications of a new law of gravitation. Foundations of Physics, 4(1):53–64, March 1974. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00708554.

vanderMerwe:1974:DIB

[101] Alwyn J. van der Merwe and Johannes H. van der Merwe. Dispersion interactions between unexcited molecules having axial symmetry: Contributions of order six and eight. Foundations of Physics, 4(1):65–74, March 1974. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00708555.

Aharonov:1974:QRR

[102] Yakir Aharonov and Gideon Carmi. Quantum-related reference frames and the local physical significance of potentials. Foundations of Physics, 4(1):75–81, March 1974. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00708556.

Rothstein:1974:LZP

[103] Jerome Rothstein. Loschmidt’s and Zermelo’s paradoxes do not exist. Foundations of Physics, 4(1):83–89, March 1974. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00708557.

Mucci:1974:MFB

[104] Joseph F. Mucci. On the mathematical form of de Broglie’s cyclical action integral. Foundations of Physics, 4(1):91–95, March 1974. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708558.

Shankara:1974:TS

[105] T. S. Shankara. Tachyons via supersonics. Foundations of Physics, 4 (1):97–104, March 1974. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00708559.

Kantor:1974:TDP

[106] W. Kantor. Is time dilation physically observable? Foundations of Physics, 4(1):105–113, March 1974. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00708560.

Cohen:1974:TQC

[107] H. A. Cohen. Is there a quantization condition for the classical problem of charge and pole? Foundations of Physics, 4(1):115–120, March 1974. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00708561.

Feenberg:1974:CDS

[108] Eugene Feenberg. Conventionality in distant simultaneity. Foundations of Physics, 4(1):121–126, March 1974. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00708562.

Cartwright:1974:CJD

[109] Nancy Cartwright. Correlations without joint distributions in quantum mechanics. Foundations of Physics, 4(1):127–136, March 1974. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708563.

Lanczos:1974:VPR

[110] C. Lanczos. Vector potential and Riemannian space. Foundations of Physics, 4(1):137–147, March 1974. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00708564.

Rybakov:1974:CIQ

[111] Yu. P. Rybakov. On the causal interpretation of quantum mechanics. Foundations of Physics, 4(2):149–161, June 1974. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00712683.

Sperber:1974:MIP

[112] Gunnar Sperber. On measurement and irreversible processes. Foundations of Physics, 4(2):163–179, June 1974. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00712684.

Krips:1974:FQTa

[113] H. Krips. Foundations of quantum theory. Part I. Foundations of Physics, 4(2):181–193, June 1974. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00712685.

Moldauer:1974:RAE

[114] P. A. Moldauer. Reexamination of the arguments of Einstein, Podolsky, and Rosen. Foundations of Physics, 4(2):195–205, June 1974. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00712686.

Huang:1974:AMS

[115] Y. K. Huang. Analytical model for super-ideal gases. Foundations of Physics, 4(2):207–213, June 1974. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00712687.

Rueda:1974:ITR

[116] A. Rueda. On the irreversible thermophysics of radiative processes. Foundations of Physics, 4(2):215–226, June 1974. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00712688.

Panarella:1974:TLI

[117] E. Panarella. Theory of laser-induced gas ionization. Foundations of Physics, 4(2):227–259, June 1974. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00712689.

Koga:1974:TEM

[118] Toyoki Koga. Tunnel effect in the motion of deformable wavelets. Foundations of Physics, 4(2):261–266, June 1974. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00712690.

Earman:1974:CIE

[119] J. Earman. Covariance, invariance, and the equivalence of frames. Foundations of Physics, 4(2):267–289, June 1974. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00712691.

Kobus:1974:GRS

[120] J. P. Kobus. Gravitational radiation in spherical coordinates. Foundations of Physics, 4(2):291–297, June 1974. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00712692.

Greenspan:1974:DNG

[121] Donald Greenspan. Discrete Newtonian gravitation and the three-body problem. Foundations of Physics, 4(2):299–310, June 1974. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00712693.

Ellis:1974:TGI

[122] Homer G. Ellis. Time, the grand illusion. Foundations of Physics, 4 (2):311–319, June 1974. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00712694.

deBroglie:1974:SPT

[123] Louis de Broglie. Strong processes and transient states. Foundations of Physics, 4(3):321–333, September 1974. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00708539.

Mayants:1974:RPP

[124] L. S. Mayants. On a relativistic particle in probabilistic physics. Foundations of Physics, 4(3):335–353, September 1974. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00708540.

Mittelstaedt:1974:OFQ

[125] P. Mittelstaedt and E. W. Stachow. Operational foundation of quantum logic. Foundations of Physics, 4(3):355–365, September 1974. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708541.

Honig:1974:MPR

[126] William M. Honig. A minimum photon “rest mass”- using Planck’s constant and discontinuous electromagnetic waves. Foundations of Physics, 4(3):367–380, September 1974. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00708542.

Krips:1974:FQTb

[127] H. Krips. Foundations of quantum theory. Part 2. Foundations of Physics, 4(3):381–394, September 1974. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00708543.

Muraskin:1974:PBA

[128] Murray Muraskin and Beatrice Ring. Particle behavior in aesthetic field theory. Foundations of Physics, 4(3):395–405, September 1974. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708544.

Allen:1974:PBN

[129] A. D. Allen. Physical bases for a new theory of motion. Foundations of Physics, 4(3):407–412, September 1974. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00708545.

Gudder:1974:TQE

[130] S. Gudder. A transient quantum effect. Foundations of Physics, 4(3): 413–416, September 1974. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00708546.

Lessen:1974:DEP

[131] Martin Lessen. Derivation of Einstein’s postulate regarding the energy of light quanta from a more elementary postulate and special relativity. Foundations of Physics, 4(3):417–419, September 1974. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708547.

Anonymous:1974:E

[132] Anonymous. Erratum. Foundations of Physics, 4(3):421, September 1974. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00708548.

Callen:1974:TSS

[133] Herbert Callen. Thermodynamics as a science of symmetry. Foundations of Physics, 4(4):423–443, December 1974. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00708519.

Janossy:1974:PIW

[134] L. Jánossy. The physical interpretation of wave mechanics. II. Foundations of Physics, 4(4):445–452, December 1974. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00708520.

Helstrom:1974:SMS

[135] Carl W. Helstrom. “Simultaneous measurement” from the standpoint of quantum estimation theory. Foundations of Physics, 4(4): 453–463, December 1974. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00708521.

Marchand:1974:NHF

[136] J.-P. Marchand and P. A. Martin. A non-Hamiltonian formulation of the Ising chain. Foundations of Physics, 4(4):465–472, December 1974. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00708522.

Edmonds:1974:CER

[137] James D. Edmonds, Jr. Complex energies in relativistic quantum theory. Foundations of Physics, 4(4):473–479, December 1974. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708523.

Schumacher:1974:FPI

[138] D. L. Schumacher. Fundamental physics and instrumental technology. Foundations of Physics, 4(4):481–497, December 1974. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708524.

Anonymous:1974:IA

[139] Anonymous. Important announcement. Foundations of Physics, 4 (4):i, December 1974. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00708518.

Anonymous:1975:IAa

[140] Anonymous. Important announcement. Foundations of Physics, 5 (1):1, March 1975. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01100307.

Yourgrau:1975:FP

[141] W. Yourgrau and A. van der Merwe. Foundations of physics. Foundations of Physics, 5(1):3–4, March 1975. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01100308.

Yourgrau:1975:HMP

[142] Wolfgang Yourgrau. Henry Margenau: a personal tribute. Foundations of Physics, 5(1):5–7, March 1975. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01100309.

Lanczos:1975:GRS

[143] C. Lanczos. Gravitation and Riemannian space. Foundations of Physics, 5(1):9–18, March 1975. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01100310.

Yourgrau:1975:CL

[144] Wolfgang Yourgrau. Cornelius Lanczos (1893–1974). Foundations of Physics, 5(1):19–20, March 1975. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01100311.

Ruark:1975:PRS

[145] Arthur E. Ruark. The physical rationale for special relativity. Foundations of Physics, 5(1):21–36, March 1975. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01100312.

Noyes:1975:FPU

[146] H. Pierre Noyes. Fixed past and uncertain future: a single-time covariant quantum particle mechanics. Foundations of Physics, 5(1):37–43, March 1975. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF01100313.

Phipps:1975:TFM

[147] T. E. Phipps, Jr. Toward a fundamental mechanics. I. Foundations of Physics, 5(1):45–58, March 1975. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01100314.

Gruber:1975:NLT

[148] Gary R. Gruber. A new look at the transition of classical to quantum mechanics. Foundations of Physics, 5(1):59–61, March 1975. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01100315.

Tautz:1975:TIE

[149] M. F. Tautz. Transposition-invariant equations for the unified field theory. Foundations of Physics, 5(1):63–74, March 1975. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01100316.

Huddleston:1975:CST

[150] P. L. Huddleston, M. Lorente, and P. Roman. Contractions of space–time groups and relativistic quantum mechanics. Foundations of Physics, 5 (1):75–87, March 1975. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01100317.

Friedman:1975:PP

[151] Kenneth S. Friedman. A problem posed. Foundations of Physics, 5 (1):89–91, March 1975. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01100318.

Bohm:1975:IUN

[152] D. J. Bohm and B. J. Hiley. On the intuitive understanding of nonlocality as implied by quantum theory. Foundations of Physics, 5(1): 93–109, March 1975. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01100319.

Jauch:1975:ATP

[153] Josef-Maria Jauch. Analytical thermodynamics. Part I. Thermostatics– general theory. Foundations of Physics, 5(1):111–132, March 1975. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01100320.

Yourgrau:1975:JMJ

[154] Wolfgang Yourgrau. Josef-Maria Jauch (1914–1974). Foundations of Physics, 5(1):133–134, March 1975. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01100321.

Treder:1975:ESE

[155] H.-J. Treder. The Einstein shift in Einstein’s box experiment. Foundations of Physics, 5(1):135–142, March 1975. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01100322.

Kawai:1975:QAP

[156] T. Kawai. Quantum action principle in curved space. Foundations of Physics, 5(1):143–158, March 1975. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01100323.

Bass:1975:QMM

[157] L. Bass. A quantum mechanical mind-body interaction. Foundations of Physics, 5(1):159–172, March 1975. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01100324.

Gordon:1975:IPI

[158] C. N. Gordon. The identification of a preferred inertial frame. Foundations of Physics, 5(1):173–183, March 1975. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01100325.

Konitz:1975:GAH

[159] Herbert Konitz. A geometric aspect of Hubble’s phenomenon. Foundations of Physics, 5(1):185–191, March 1975. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01100326.

Schlegel:1975:SQR

[160] Richard Schlegel. Superposition in quantum and relativity physics — an interaction interpretation of special relativity theory: Part III. Foundations of Physics, 5(2):197–215, June 1975. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00717437.

Cannon:1975:ETI

[161] W. H. Cannon and O. G. Jensen. An empirical test of the interaction interpretation of the theory of relativity. Foundations of Physics, 5(2):217–227, June 1975. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00717438.

DeFacio:1975:CIA

[162] B. DeFacio. Causal independence in algebraic quantum field theory. Foundations of Physics, 5(2):229–237, June 1975. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00717439.

Edmonds:1975:ERM

[163] James D. Edmonds, Jr. Extended relativity: Mass and the fifth dimension. Foundations of Physics, 5(2):239–249, June 1975. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00717440.

Benioff:1975:PMQ

[164] Paul Benioff. On procedures for the measurement of questions in quantum mechanics. Foundations of Physics, 5(2):251–255, June 1975. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00717441.

Jaakkola:1975:CIA

[165] T. Jaakkola, M. Moles, J. P. Vigier, J. C. Pecker, and W. Yourgrau. Cosmological implications of anomalous redshifts — a possible working hypothesis. Foundations of Physics, 5(2):257–269, June 1975. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00717442.

Liboff:1975:BCP

[166] Richard L. Liboff. Bohr correspondence principle for large quantum numbers. Foundations of Physics, 5(2):271–293, June 1975. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00717443.

Kingsley:1975:CPS

[167] James M. Kingsley. On the consistency of the postulates of special relativity. Foundations of Physics, 5(2):295–300, June 1975. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00717444.

Sarfatti:1975:TUF

[168] Jack Sarfatti. Toward a unified field theory of gravitation and strong interactions. Foundations of Physics, 5(2):301–307, June 1975. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00717445.

Opher:1975:QPC

[169] Reuven Opher. Are quantum processes cosmologically induced? Foundations of Physics, 5(2):309–321, June 1975. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00717446.

Zabey:1975:RTQ

[170] P. C. Zabey. Reconstruction theorems in quantum mechanics. Foundations of Physics, 5(2):323–342, June 1975. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00717447.

Kobus:1975:TDE

[171] J. P. Kobus. A theoretical determination of the electronic rest mass. Foundations of Physics, 5(2):343–347, June 1975. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00717448.

Duffey:1975:TSW

[172] George H. Duffey. Tachyons and superluminal wave groups. Foundations of Physics, 5(2):349–354, June 1975. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00717449.

delaPena:1975:STC

[173] L. de la Peña and A. M. Cetto. Stochastic theory for classical and quantum mechanical systems. Foundations of Physics, 5(2):355–370, June 1975. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00717450.

Zeh:1975:SBV

[174] H. D. Zeh. Symmetry-breaking vacuum and state vector reduction. Foundations of Physics, 5(2):371–373, June 1975. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00717451.

Barut:1975:ETF

[175] A. O. Barut and S. Malin. Electrodynamics in terms of functions over the group SU(2). I. The equation of the vector potential. Foundations of Physics, 5(3):375–386, September 1975. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00708884.

Browne:1975:FDU

[176] P. F. Browne. Five-dimensional unified field theory: The source function. Foundations of Physics, 5(3):387–398, September 1975. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708885.

Newburgh:1975:PWT

[177] Ronald G. Newburgh. Pulse widths and time dilatation. Foundations of Physics, 5(3):399–405, September 1975. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00708886.

Var:1975:NMF

[178] Robert E. Var. On a new mathematical framework for fundamental theoretical physics. Foundations of Physics, 5(3):407–431, September 1975. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00708887.

Bailey:1975:NLH

[179] Cecil D. Bailey. A new look at Hamilton’s principle. Foundations of Physics, 5(3):433–451, September 1975. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00708888.

Chow:1975:SMM

[180] T. L. Chow. Spin magnetic moment of a free relativistic electron. Foundations of Physics, 5(3):453–454, September 1975. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00708889.

Bell:1975:HOU

[181] A. B. Bell and D. M. Bell. A highly ordered universe. Foundations of Physics, 5(3):455–480, September 1975. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00708890.

Stark:1975:VPI

[182] A. Stark. Variational properties of irreversible processes. Foundations of Physics, 5(3):481–490, September 1975. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00708891.

Mirman:1975:DT

[183] R. Mirman. The direction of time. Foundations of Physics, 5(3):491– 511, September 1975. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00708892.

Muraskin:1975:TPC

[184] M. Muraskin and B. Ring. A two-particle collision in aesthetic field theory. Foundations of Physics, 5(3):513–523, September 1975. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708893.

Davidon:1975:CIE

[185] William C. Davidon. Consequences of the inertial equivalence of energy. Foundations of Physics, 5(3):525–542, September 1975. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708894.

Whealton:1975:IDT

[186] J. H. Whealton. Illustrations of a dynamical theory of the ether. Foundations of Physics, 5(3):543–553, September 1975. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00708895.

Anonymous:1975:IAb

[187] Anonymous. Important announcement. Foundations of Physics, 5(4): 555, December 1975. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00708428.

Prugovecki:1975:MQM

[188] Eduard Prugovecki. Measurement in quantum mechanics as a stochastic process on spaces of fuzzy events. Foundations of Physics, 5(4): 557–571, December 1975. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00708429.

Lavenda:1975:TAM

[189] B. H. Lavenda. Thermodynamics of averaged motion. Foundations of Physics, 5(4):573–589, December 1975. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00708430.

Witriol:1975:GQM

[190] Norman M. Witriol. General quantum mechanical canonical point transformations. Foundations of Physics, 5(4):591–605, December 1975. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00708431.

Capek:1975:RSB

[191] Milic Capek. Relativity and the status of becoming. Foundations of Physics, 5(4):607–617, December 1975. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00708432.

Frost:1975:MFS

[192] Authur A. Frost. Matrix formulation of special relativity in classical mechanics and electromagnetic theory. Foundations of Physics, 5(4): 619–641, December 1975. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00708433.

Edmonds:1975:SQQ

[193] James D. Edmonds, Jr. Second quantized quaternion quantum theory. Foundations of Physics, 5(4):643–648, December 1975. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708434.

Kobus:1975:ICU

[194] J. P. Kobus. Impossibility of collapse under the law of gravitation R_{44} =0. Foundations of Physics, 5(4):649–654, December 1975. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708435.

Anonymous:1976:IAa

[195] Anonymous. Important announcement. Foundations of Physics, 6 (1):1, February 1976. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00708658.

deBroglie:1976:PPH

[196] Louis de Broglie, Georges Lochak, Juan Alberto Beswick, and José Vassalo-Pereira. Present, predicted, and hidden probabilities. Foundations of Physics, 6(1):3–14, February 1976. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00708659.

Hatsopoulos:1976:UQTa

[197] George N. Hatsopoulos and Elias P. Gyftopoulos. A unified quantum theory of mechanics and thermodynamics. Part I. Postulates. Foundations of Physics, 6(1):15–31, February 1976. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00708660.

Edmonds:1976:RHS

[198] J. D. Edmonds, Jr. A relativisitic “higher spin” quaternion wave equation giving a variation on the Pauli equation. Foundations of Physics, 6 (1):33–36, February 1976. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00708661.

Honig:1976:GAM

[199] William M. Honig. Gödel axiom mappings in special relativity and quantum-electromagnetic theory. Foundations of Physics, 6(1):37– 57, February 1976. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00708662.

Allen:1976:BBN

[200] Allen D. Allen. The big bang is not needed. Foundations of Physics, 6 (1):59–63, February 1976. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00708663.

Sinha:1976:ASS

[201] K. P. Sinha, C. Sivaram, and E. C. G. Sudarshan. Aether as a superfluid state of particle-antiparticle pairs. Foundations of Physics, 6 (1):65–70, February 1976. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00708664.

Phipps:1976:TFMa

[202] T. E. Phipps, Jr. Toward a fundamental mechanics. II. Foundations of Physics, 6(1):71–82, February 1976. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00708665.

Noyes:1976:DPQ

[203] H. Pierre Noyes. A Democritean phenomenology for quantum scattering theory. Foundations of Physics, 6(1):83–100, February 1976. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708666.

Bell:1976:MEH

[204] A. B. Bell and D. M. Bell. Mates to e=mc^{2} and to the Heisenberg uncertainty relations. Foundations of Physics, 6(1):101–106, February 1976. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00708667.

Corsiglia:1976:RSC

[205] L. Corsiglia. Removal of singularities in the Coulomb and Yukawa potentials. Foundations of Physics, 6(1):107–109, February 1976. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708668.

Gruber:1976:TCQ

[206] Gary R. Gruber. On the transition from classical to quantum mechanics in generalized coordinates. Foundations of Physics, 6(1):111– 113, February 1976. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00708669.

Eriksen:1976:CRC

[207] Erik Eriksen and Kjell Vøyenli. The classical and relativistic concepts of mass. Foundations of Physics, 6(1):115–124, February 1976. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708670.

Anonymous:1976:E

[208] Anonymous. Erratum. Foundations of Physics, 6(1):125, February 1976. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00708671.

Hatsopoulos:1976:UQTb

[209] George N. Hatsopoulos and Elias P. Gyftopoulos. A unified quantum theory of mechanics and thermodynamics. Part IIa. Available energy. Foundations of Physics, 6(2):127–141, April 1976. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00708955.

Stephenson:1976:DMG

[210] L. M. Stephenson. A dynamical model for gravitation. Foundations of Physics, 6(2):143–155, April 1976. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00708956.

Park:1976:MEE

[211] James L. Park and William Band. Mutually exclusive and exhaustive quantum states. Foundations of Physics, 6(2):157–172, April 1976. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00708957.

Lochak:1976:BIG

[212] Georges Lochak. Has Bell’s inequality a general meaning for hiddenvariable theories? Foundations of Physics, 6(2):173–184, April 1976. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00708958.

DeFacio:1976:QMA

[213] B. DeFacio. Quantum measurement and algebraic quantum field theories. Foundations of Physics, 6(2):185–192, April 1976. CODEN FNDPA4.

ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00708959.

Kohler:1976:EVC

[214] Robert H. Kohler. Extended view of classical contact transformations. Foundations of Physics, 6(2):193–208, April 1976. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00708960.

Winans:1976:DUM

[215] J. Gibson Winans. Definitions and units in mechanics. Foundations of Physics, 6(2):209–219, April 1976. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00708961.

Shekhawat:1976:SPF

[216] V. Shekhawat. Some preliminary formulations toward a new theory of matter. Foundations of Physics, 6(2):221–235, April 1976. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708962.

Perkins:1976:DZH

[217] Granville A. Perkins. The direction of the Zitterbewegung: a hidden variable. Foundations of Physics, 6(2):237–248, April 1976. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708963.

Band:1976:NIT

[218] William Band and James L. Park. New information-theoretic foundations for quantum statistics. Foundations of Physics, 6(3):249–262, June 1976. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00708800.

Phipps:1976:TFMb

[219] T. E. Phipps, Jr. Toward a fundamental mechanics. III. Foundations of Physics, 6(3):263–273, June 1976. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00708801.

Maxwell:1976:TMR

[220] Nicholas Maxwell. Toward a micro realistic version of quantum mechanics. Part I. Foundations of Physics, 6(3):275–292, June 1976. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708802.

Streltsov:1976:RL

[221] V. N. Strel’tsov. On the relativistic length. Foundations of Physics, 6 (3):293–298, June 1976. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00708803.

Gleyzal:1976:UFA

[222] André Gleyzal. Unified fields of analytic space–time-energy. Foundations of Physics, 6(3):299–303, June 1976. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00708804.

Gluckman:1976:VKE

[223] Albert G. Gluckman. Voigt kinematics and electrodynamic consequences. Foundations of Physics, 6(3):305–316, June 1976. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00708805.

Hsu:1976:NFD

[224] J. P. Hsu. New four-dimensional symmetry. Foundations of Physics, 6(3):317–339, June 1976. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00708806.

Janossy:1976:WMP

[225] L. Jánossy. Wave mechanics and physical reality. III. The many-body problem. Foundations of Physics, 6(3):341–350, June 1976. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708807.

Bell:1976:PQ

[226] A. B. Bell and D. M. Bell. Particles without quarks. Foundations of Physics, 6(3):351–366, June 1976. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00708808.

Wolsky:1976:WWP

[227] A. M. Wolsky. Why “a watched pot never boils”. Foundations of Physics, 6(3):367–369, June 1976. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00708809.

Ax:1976:GTT

[228] James Ax. Group-theoretic treatment of the axioms of quantum mechanics. Foundations of Physics, 6(4):371–399, August 1976. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00715028.

Gupta:1976:MMM

[229] N. D. Sen Gupta. On magnetic moment and magnetic induction. Foundations of Physics, 6(4):401–405, August 1976. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00715029.

Gal-Or:1976:COIa

[230] Benjamin Gal-Or. Cosmological origin of irreversibility, time, and time anisotropies. I. Foundations of Physics, 6(4):407–426, August 1976. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00715030.

Szczepanski:1976:SVR

[231] A. Szczepański. State vector reduction and photon coincidences. Foundations of Physics, 6(4):427–433, August 1976. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00715031.

Schlegel:1976:IG

[232] Richard Schlegel. Interaction, not gravitation. Foundations of Physics, 6(4):435–438, August 1976. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00715032.

Hatsopoulos:1976:UQTc

[233] George N. Hatsopoulos and Elias P. Gyftopoulos. A unified quantum theory of mechanics and thermodynamics. Part IIb. Stable equilibrium states. Foundations of Physics, 6(4):439–455, August 1976. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00715033.

Browne:1976:AGA

[234] P. F. Browne. Arbitrariness of geometry and the aether. Foundations of Physics, 6(4):457–471, August 1976. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00715034.

Hobart:1976:CDP

[235] R. H. Hobart. A cosmological derivation of Planck’s constant. Foundations of Physics, 6(4):473–476, August 1976. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00715035.

Sida:1976:GRE

[236] D. W. Sida. A geometrical relationship for the Einstein and Ricci tensors. Foundations of Physics, 6(4):477–483, August 1976. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00715036.

Mayants:1976:TPS

[237] L. S. Mayants. On transformations of physical systems. Foundations of Physics, 6(5):485–510, October 1976. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00715104.

Bub:1976:HVL

[238] Jeffrey Bub. Hidden variables and locality. Foundations of Physics, 6 (5):511–525, October 1976. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00715105.

Treder:1976:GUF

[239] Hans-Jürgen Treder. Gravitation and universal Fermi coupling in general relativity. Foundations of Physics, 6(5):527–538, October 1976. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00715106.

deBeauregard:1976:TSI

[240] O. Costa de Beauregard. Time symmetry and interpretation of quantum mechanics. Foundations of Physics, 6(5):539–559, October 1976. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00715107.

Hatsopoulos:1976:UQTd

[241] George N. Hatsopoulos and Elias P. Gyftopoulos. A unified quantum theory of mechanics and thermodynamics. Part III. Irreducible quantal dispersions. Foundations of Physics, 6(5):561–570, October 1976. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00715108.

Marinov:1976:GDT

[242] Stefan Marinov. Gravitational (dynamic) time dilation according to absolute space–time theory. Foundations of Physics, 6(5):571–581, October 1976. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00715109.

Santhanam:1976:QMF

[243] T. S. Santhanam and A. R. Tekumalla. Quantum mechanics in finite dimensions. Foundations of Physics, 6(5):583–587, October 1976. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00715110.

Ionescu-Pallas:1976:CFF

[244] Nicholas Ionescu-Pallas and Liviu Sofonea. The concept of force and the formalization of non-quantum-mechanical theories. Foundations of Physics, 6(5):589–597, October 1976. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00715111.

Bedford:1976:ECQ

[245] Donald Bedford and Derek Wang. Environment, consciousness, and quantum measurement. Foundations of Physics, 6(5):599–605, October 1976. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00715112.

Frederick:1976:CWF

[246] Carl Frederick. The collapse of the wave function. Foundations of Physics, 6(5):607–611, October 1976. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00715113.

Eaves:1976:MPR

[247] R. E. Eaves. Mach’s Principle, relative motion, and fundamental numbers of physics. Foundations of Physics, 6(5):613–620, October 1976. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00715114.

Anonymous:1976:IAb

[248] Anonymous. Important announcement. Foundations of Physics, 6(6): 621, December 1976. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00708942.

Gal-Or:1976:COIb

[249] Benjamin Gal-Or. Cosmological origin of irreversibility, time, and time anisotropies. II. Gravitism. Foundations of Physics, 6(6):623– 637, December 1976. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00708943.

Krips:1976:FQT

[250] H. Krips. Foundations of quantum theory. Part 3. Foundations of Physics, 6(6):639–659, December 1976. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00708944.

Maxwell:1976:TMV

[251] Nicholas Maxwell. Toward a microrealistic version of quantum mechanics. Part II. Foundations of Physics, 6(6):661–676, December 1976. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00708945.

Gron:1976:CPS

[252] Ø. Grøn and M. Nicola. The consistency of the postulates of special relativity. Foundations of Physics, 6(6):677–680, December 1976. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00708946.

Chatham:1976:CR

[253] R. E. Chatham. Consistency in relativity. Foundations of Physics, 6(6): 681–685, December 1976. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00708947.

Prokhovnik:1976:CLS

[254] S. J. Prokhovnik. Cosmological lore-and some heresies. Foundations of Physics, 6(6):687–705, December 1976. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00708948.

Hsu:1976:GLG

[255] J. P. Hsu. Generalized local gauge symmetry and the Ward–Takahashi identities in unified field theories. Foundations of Physics, 6(6):707– 716, December 1976. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00708949.

Sinha:1976:SVS

[256] K. P. Sinha, C. Sivaram, and E. C. G. Sudarshan. The superfluid vacuum state, time-varying cosmological constant, and nonsingular cosmological models. Foundations of Physics, 6(6):717–726, December 1976. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708950.

Landsberg:1976:GRE

[257] P. T. Landsberg and N. T. Bishop. Gravitational redshift and the equivalence principle. Foundations of Physics, 6(6):727–737, December 1976. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00708951.

Cohen:1976:CQP

[258] Leon Cohen. Covariance for quasi-probability distributions. Foundations of Physics, 6(6):739–741, December 1976. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00708952.

Anonymous:1977:IA

[259] Anonymous. Important announcement. Foundations of Physics, 7 (1–2):1, February 1977. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00715237.

Mayants:1977:SPQ

[260] L. S. Mayants. On some peculiarities of quantum mechanics. Foundations of Physics, 7(1–2):3–28, February 1977. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00715238.

Anderson:1977:DOO

[261] R. Anderson and G. E. Stedman. Dual observers in operational relativity. Foundations of Physics, 7(1–2):29–33, February 1977. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00715239.

Marchand:1977:RCG

[262] Jean-Paul Marchand. Relative coarse-graining. Foundations of Physics, 7 (1–2):35–49, February 1977. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00715240.

Eppley:1977:NQG

[263] Kenneth Eppley and Eric Hannah. The necessity of quantizing the gravitational field. Foundations of Physics, 7(1–2):51–68, February 1977. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00715241.

Harding:1977:QMD

[264] Charles Harding. Quantum mechanics as demanded by the special theory of relativity. Foundations of Physics, 7(1–2):69–76, February 1977. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00715242.

Cvijanovich:1977:NEM

[265] G. B. Cvijanovich and J. P. Vigier. New extended model of the Dirac electron. Foundations of Physics, 7(1–2):77–96, February 1977. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00715243.

Volkenstein:1977:AVI

[266] M. V. Volkenstein. The amount and value of information in biology. Foundations of Physics, 7(1–2):97–109, February 1977. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00715244.

Jhangiani:1977:GSS

[267] V. Jhangiani. Geometric significance of the spinor covariant derivative. Foundations of Physics, 7(1–2):111–120, February 1977. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00715245.

Santhanam:1977:QMD

[268] T. S. Santhanam. Quantum mechanics in discrete space and angular momentum. Foundations of Physics, 7(1–2):121–127, February 1977. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00715246.

Moore:1977:EGH

[269] Richard L. Moore. An extension of the Gauss–Hertz principle. Foundations of Physics, 7(1–2):129–136, February 1977. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00715247.

Thomas:1977:IVV

[270] Richard N. Thomas. Irreversibility, evolution, Evolution, and the process of local concentration. Foundations of Physics, 7(1–2):137–150, February 1977. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00715248.

Rayski:1977:EMF

[271] Jerzy Rayski. Epistemological and mathematical foundations of quantum mechanics. Foundations of Physics, 7(3–4):151–164, April 1977. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00709004.

Browne:1977:CAG

[272] P. F. Browne. Complementary aspects of gravitation and electromagnetism. Foundations of Physics, 7(3–4):165–183, April 1977. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00709005.

Kawai:1977:CQN

[273] T. Kawai. Canonical quantization of a nonrelativistic singular quasilinear system. Foundations of Physics, 7(3–4):185–204, April 1977. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00709006.

Hsu:1977:LEV

[274] J. P. Hsu. Laser experiments and various four-dimensional symmetries. Foundations of Physics, 7(3–4):205–220, April 1977. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00709007.

Bass:1977:BRQ

[275] L. Bass. Biological replication by quantum mechanical interactions. Foundations of Physics, 7(3–4):221–231, April 1977. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00709008.

Park:1977:RIT

[276] James L. Park and William Band. Rigorous information-theoretic derivation of quantum-statistical thermodynamics. I. Foundations of Physics, 7(3–4):233–244, April 1977. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00709009.

Schlegel:1977:LIC

[277] Richard Schlegel. A Lorentz-invariant clock. Foundations of Physics, 7 (3–4):245–253, April 1977. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00709010.

Post:1977:MMP

[278] E. J. Post. A minor or a major predicament of physical theory? (charge and action polarity and their order properties). Foundations of Physics, 7(3–4):255–277, April 1977. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00709011.

Freundlich:1977:TVO

[279] Yehudah Freundlich. Two views of an objective quantum theory. Foundations of Physics, 7(3–4):279–300, April 1977. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00709012.

Kiehn:1977:PTI

[280] R. M. Kiehn. Parity and time inversion symmetries of electromagnetic systems. Foundations of Physics, 7(5–6):301–311, June 1977. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:

//link.springer.com/article/10.1007/BF00711483. See comments

[385, 386].

Stapp:1977:TR

[281] Henry Pierce Stapp. Theory of reality. Foundations of Physics, 7(5– 6):313–323, June 1977. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00711484.

Streltsov:1977:RRT

[282] V. N. Strel’tsov. Remarks on relativistic thermodynamics. Foundations of Physics, 7(5–6):325–330, June 1977. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00711485.

Agmon:1977:RTT

[283] Noam Agmon. Relativistic transformations of thermodynamic quantities. Foundations of Physics, 7(5–6):331–339, June 1977. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00711486.

Winans:1977:QPQ

[284] J. Gibson Winans. Quaternion physical quantities. Foundations of Physics, 7(5–6):341–349, June 1977. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00711487.

Rietdijk:1977:HDV

[285] C. W. Rietdijk. How do “virtual” photons and mesons transmit forces between charged particles and nucleons? Foundations of Physics, 7 (5–6):351–374, June 1977. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00711488.

Kasper:1977:CFQ

[286] U. Kasper, E. Kreisel, and H. J. Treder. On the covariant formulation of quantum mechanics. Foundations of Physics, 7(5–6):375–389, June 1977. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00711489.

Ohanian:1977:CCA

[287] Hans C. Ohanian. Cosmological changes in atomic and nuclear constants. Foundations of Physics, 7(5–6):391–404, June 1977. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00711490.

Panarella:1977:ETE

[288] E. Panarella. Experimental test of the effective-photon hypothesis. Foundations of Physics, 7(5–6):405–419, June 1977. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00711491.

Pandres:1977:GE

[289] D. Pandres, Jr. Gravitation and electromagnetism. Foundations of Physics, 7(5–6):421–430, June 1977. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00711492.

Angelidis:1977:MCD

[290] Th. D. Angelidis. Momentum conservation decides Heisenberg’s interpretation of the uncertainty formulas. Foundations of Physics, 7(5– 6):431–449, June 1977. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00711493.

Muraskin:1977:ICA

[291] Murray Muraskin and Beatrice Ring. Increased complexity in aesthetic field theory. Foundations of Physics, 7(5–6):451–458, June 1977. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00711494.

Laue:1978:RPQ

[292] Hans Laue. The role of position in quantum theory. Foundations of Physics, 8(1–2):1–30, February 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00708483.

Raskin:1978:STS

[293] Paul D. Raskin. Short-time stochastic electron. Foundations of Physics, 8 (1–2):31–44, February 1978. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00708484.

Band:1978:GTL

[294] William Band and James L. Park. Generalized two-level quantum dynamics. II. Non-Hamiltonian state evolution. Foundations of Physics, 8 (1–2):45–58, February 1978. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00708485.

Smalley:1978:PPN

[295] Larry L. Smalley. Parametrized post-Newtonian approximation and Rastall’s gravitational field equations. Foundations of Physics, 8(1– 2):59–68, February 1978. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00708486.

Altschul:1978:CTT

[296] Martin S. Altschul. Coordinate transformations and the theory of measurement. Foundations of Physics, 8(1–2):69–92, February 1978. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00708487.

Vertosick:1978:RSI

[297] Frank T. Vertosick, Jr. Remarks on stratifying interactions in five dimensions. Foundations of Physics, 8(1–2):93–102, February 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708488.

Selleri:1978:CEL

[298] F. Selleri. On the consequences of Einstein locality. Foundations of Physics, 8(1–2):103–116, February 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00708489.

Benoist:1978:ASI

[299] Rodney W. Benoist, Jean-Paul Marchand, and Wolfgang Yourgrau. Addendum to statistical inference and quantum mechanical measurement. Foundations of Physics, 8(1–2):117–118, February 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708490.

Janossy:1978:WMT

[300] L. Jánossy. Wave mechanics and the tunnel effect. Foundations of Physics, 8(1–2):119–122, February 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00708491.

Edmonds:1978:ESF

[301] James D. Edmonds, Jr. An elegant but “simple” form for the Dirac hydrogen atom. Foundations of Physics, 8(1–2):123–129, February 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00708492.

Liebscher:1978:DEV

[302] Dierck-Ekkehard Liebscher. Derivation of Einstein’s velocity addition theorem through use of the invariant double ratio. Foundations of Physics, 8(1–2):131–135, February 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00708493.

Marinov:1978:RDE

[303] Stefan Marinov. Rotating disk experiments. Foundations of Physics, 8(1– 2):137–156, February 1978. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00708494.

Mackinnon:1978:NBW

[304] L. Mackinnon. A nondispersive de Broglie wave packet. Foundations of Physics, 8(3–4):157–176, April 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00715205.

Perkins:1978:EVM

[305] Granville A. Perkins. Electron velocity and momentum density. Foundations of Physics, 8(3–4):177–189, April 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00715206.

delaPena-Auerbach:1978:QMD

[306] L. de la Peña-Auerbach and A. M. Cetto. Quantum mechanics derived from stochastic electrodynamics. Foundations of Physics, 8(3– 4):191–210, April 1978. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00715207.

McCollum:1978:SCP

[307] Gin McCollum. Spacetime code: Preliminaries and motivations. Foundations of Physics, 8(3–4):211–228, April 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00715208.

Roxburgh:1978:CGR

[308] I. W. Roxburgh and R. K. Tavakol. Conventionalism and general relativity. Foundations of Physics, 8(3–4):229–237, April 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00715209.

Park:1978:GTL

[309] James L. Park and William Band. Generalized two-level quantum dynamics. III. Irreversible conservative motion. Foundations of Physics, 8 (3–4):239–254, April 1978. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00715210.

Capri:1978:MFS

[310] A. Z. Capri. Multilocal fields — some examples. Foundations of Physics, 8(3–4):255–269, April 1978. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00715211.

Wyss:1978:D

[311] Walter Wyss. Duality. Foundations of Physics, 8(3–4):271–275, April 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00715212.

Post:1978:MSI

[312] E. J. Post. Magnetic symmetry, improper symmetry, and Neumann’s principle. Foundations of Physics, 8(3–4):277–294, April 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00715213.

Gudder:1978:GRF

[313] S. P. Gudder. Gaussian random fields. Foundations of Physics, 8(3– 4):295–302, April 1978. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00715214.

Edmonds:1978:NVC

[314] James D. Edmonds, Jr. Nine-vectors, complex octonion/quaternion hypercomplex numbers, lie groups, and the ‘real’ world. Foundations of Physics, 8(3–4):303–311, April 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00715215.

Herbert:1978:GBT

[315] Nick Herbert and Jack Karush. Generalization of Bell’s theorem. Foundations of Physics, 8(3–4):313–317, April 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00715216.

Chou:1978:EDO

[316] T. T. Chou. Energy dependence of opaqueness for pp collisions at high energies. Foundations of Physics, 8(5–6):319–328, June 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708565.

Recami:1978:HRC

[317] Erasmo Recami. How to recover causality in special relativity for tachyons. Foundations of Physics, 8(5–6):329–340, June 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708566.

Surdin:1978:SSU

[318] M. Surdin. The steady state universe revisited, with stochastic electrodynamics as a guide. Foundations of Physics, 8(5–6):341–357, June 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00708567.

Moore:1978:FEG

[319] Richard L. Moore. Further extension of the Gauss–Hertz principle. Foundations of Physics, 8(5–6):359–370, June 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00708568.

Hsu:1978:QEW

[320] J. P. Hsu. Quantum electrodynamics within the framework of a new fourdimensional symmetry. Foundations of Physics, 8(5–6):371–391, June 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00708569.

Newburgh:1978:SMN

[321] Ronald G. Newburgh. Syntopic measurements: a new classification of spacetime measurement. Foundations of Physics, 8(5–6):393–398, June 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00708570.

Meessen:1978:SQG

[322] A. Meessen. Spacetime quantization, generalized relativistic mechanics, and Mach’s Principle. Foundations of Physics, 8(5–6):399–415, June 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00708571.

Eagles:1978:EIS

[323] D. M. Eagles. An electrostatic interpretation of some empirical parameters of light quarks. Foundations of Physics, 8(5–6):417–421, June 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00708572.

Roederer:1978:RBH

[324] Juan G. Roederer. On the relationship between human brain functions and the foundations of physics, science, and technology. Foundations of Physics, 8(5–6):423–438, June 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00708573.

Edmonds:1978:YAF

[325] James D. Edmonds, Jr. Yet another formulation of the Dirac equation. Foundations of Physics, 8(5–6):439–444, June 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00708574.

Jhangiani:1978:GSSa

[326] V. Jhangiani. Geometric significance of the spinor Lie derivative. I. Foundations of Physics, 8(5–6):445–462, June 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00708575.

Schmidt:1978:CEP

[327] Helmut Schmidt. Can an effect precede its cause? A model of a noncausal world. Foundations of Physics, 8(5–6):463–480, June 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708576.

Davidson:1978:CRQ

[328] Mark Davidson. A classical realization of quantum mechanics. Foundations of Physics, 8(5–6):481–492, June 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00708577.

Cyranski:1978:AME

[329] John F. Cyranski. Analysis of the maximum entropy principle “debate”. Foundations of Physics, 8(5–6):493–506, June 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708578.

Ax:1978:EFS

[330] James Ax. The elementary foundations of spacetime. Foundations of Physics, 8(7–8):507–546, August 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00717578.

Phillips:1978:NTH

[331] Peter R. Phillips. Nonlinear transformations in Hilbert space. Foundations of Physics, 8(7–8):547–564, August 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00717579.

Ignatovich:1978:NWP

[332] V. K. Ignatovich. Nonspreading wave packets in quantum mechanics. Foundations of Physics, 8(7–8):565–571, August 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00717580.

Green:1978:QMS

[333] H. S. Green. Quantum mechanics of space and time. Foundations of Physics, 8(7–8):573–591, August 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00717581.

Jhangiani:1978:GSSb

[334] V. Jhangiani. Geometric significance of the spinor Lie derivative. II. Foundations of Physics, 8(7–8):593–601, August 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00717582.

Kalotas:1978:CVL

[335] T. M. Kalotas and A. R. Lee. On the constancy of the velocity of light. Foundations of Physics, 8(7–8):603–607, August 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00717583.

Hsu:1978:CCC

[336] J. P. Hsu and T. N. Sherry. Critical comments on “On the constancy of the velocity of light”. Foundations of Physics, 8(7–8):609–614, August 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00717584.

Rietdijk:1978:PRI

[337] C. W. Rietdijk. Proof of a retroactive influence. Foundations of Physics, 8 (7–8):615–628, August 1978. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00717585.

Fine:1978:ACH

[338] Arthur Fine and Paul Teller. Algebraic constraints on hidden variables. Foundations of Physics, 8(7–8):629–636, August 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00717586.

Marinov:1978:LDE

[339] Stefan Marinov. The light Doppler effect treated by absolute spacetime theory. Foundations of Physics, 8(7–8):637–652, August 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00717587.

Benson:1978:SCS

[340] Alvin K. Benson. Self-consistent selection of a superconducting representation for the BCS model. Foundations of Physics, 8(9–10): 653–666, October 1978. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00717497.

Hsu:1978:OPF

[341] J. P. Hsu. Observable phase factors and symmetry of electric and magnetic charges. Foundations of Physics, 8(9–10):667–675, October 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00717498.

Band:1978:SFD

[342] William Band and James L. Park. Some fundamental difficulties with quantum mechanical collision theory. Foundations of Physics, 8(9–10): 677–694, October 1978. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00717499.

Treder:1978:GRG

[343] Hans-Jürgen Treder and Wolfgang Yourgrau. On general-relativistic and gauge field theories. Foundations of Physics, 8(9–10):695–708, October 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00717500.

Benioff:1978:NEI

[344] Paul Benioff. A note on the Everett interpretation of quantum mechanics. Foundations of Physics, 8(9–10):709–720, October 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00717501.

Neumann:1978:PAD

[345] Manfred Neumann. A probabilistic analysis of the difficulties of unifying quantum mechanics with the theory of relativity. Foundations of Physics, 8(9–10):721–733, October 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00717502.

Sivashinsky:1978:STM

[346] G. Sivashinsky. Self-turbulence in the motion of a free particle. Foundations of Physics, 8(9–10):735–744, October 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00717503.

Perkins:1978:SEP

[347] Granville A. Perkins. A spinor equation of the pure electromagnetic field. Foundations of Physics, 8(9–10):745–757, October 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00717504.

Spigelman:1978:TEG

[348] Joseph H. Spigelman. Toward an elemental grounding of microphysics. Foundations of Physics, 8(9–10):759–785, October 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00717505.

Kalotas:1978:GFL

[349] T. M. Kalotas and A. R. Lee. On the general form of Lorentz transformations. Foundations of Physics, 8(9–10):787–792, October 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00717506.

Edmonds:1978:PCD

[350] James D. Edmonds, Jr. Parton confinement: a different perspective. Foundations of Physics, 8(9–10):793–795, October 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00717507.

Vrcelj:1978:CAS

[351] Zoran Vrcelj. A criticism of the “absolute space–time theory”. Foundations of Physics, 8(9–10):797–800, October 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00717508. See comments [352].

Marinov:1978:CCA

[352] S. Marinov. Comments on: “A criticism of the ‘absolute space–time theory’ ”. Foundations of Physics, 8(9–10):801–804, October 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00717509. See [351].

Cyranski:1978:QMC

[353] John F. Cyranski. Quantum measurement as a communication with nature. Foundations of Physics, 8(11–12):805–822, December 1978. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00715055.

Sinha:1978:SSA

[354] K. P. Sinha and E. C. G. Sudarshan. The superfluid as a source of all interactions. Foundations of Physics, 8(11–12):823–831, December 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00715056.

Hsu:1978:GFF

[355] J. P. Hsu and J. A. Underwood. General flat four-dimensional world pictures and clock systems. Foundations of Physics, 8(11–12):833– 843, December 1978. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00715057.

Roychoudhuri:1978:HMM

[356] Chandrasekhar Roychoudhuri. Heisenberg microscope — misleading illustration. Foundations of Physics, 8(11–12):845–849, December 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://www.springerlink.com/content/t58j8k8m73666lu7/.

Fanchi:1978:QMR

[357] John R. Fanchi and R. Eugene Collins. Quantum mechanics of relativistic spinless particles. Foundations of Physics, 8(11–12):851–877, December 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL https://link.springer.com/article/10.1007/BF00715059.

Pearle:1978:WCC

[358] Philip Pearle. When can a classical electron accelerate without radiating? Foundations of Physics, 8(11–12):879–891, December 1978. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

https://link.springer.com/article/10.1007/BF00715060.

Benson:1978:MMM

[359] Alvin K. Benson. Microscopic mechanism for the macroscopic asymmetry of superconductivity. Foundations of Physics, 8(11–12):893– 904, December 1978. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/BF00715061.

Pietschmann:1978:RSD

[360] Herbert Pietschmann. The rules of scientific discovery demonstrated from examples of the physics of elementary particles. Foundations of Physics, 8(11–12):905–919, December 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/BF00715062.

Stephenson:1978:CAA

[361] L. M. Stephenson. Clarification of an apparent asymmetry in electromagnetic theory. Foundations of Physics, 8(11–12):921–926, December 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL https://link.springer.com/article/10.1007/BF00715063.

Jaki:1978:JGS

[362] Stanley L. Jaki. Johann Georg von Soldner and the gravitational bending of light, with an English translation of his essay on it published in 1801. Foundations of Physics, 8(11–12):927–950, December 1978. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

https://link.springer.com/article/10.1007/BF00715064.

Bell:1978:NTS

[363] A. B. Bell and D. M. Bell. New theory of superconductivity. Foundations of Physics, 8(11–12):951–957, December 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/BF00715065.

Stapp:1979:WAQ

[364] Henry P. Stapp. Whiteheadian approach to quantum theory and the generalized Bell’s theorem. Foundations of Physics, 9(1–2):1–25, February 1979. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00715049.

Bess:1979:DMD

[365] Leon Bess. A diffusion model for the Dirac equation. Foundations of Physics, 9(1–2):27–54, February 1979. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00715050.

Hsu:1979:ATR

[366] J. P. Hsu. The analysis of time: Is the relativistic time unique? Foundations of Physics, 9(1–2):55–69, February 1979. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00715051.

deMuynck:1979:SMJ

[367] Willem M. de Muynck, Peter A. E. M. Janssen, and Alexander Santman. Simultaneous measurement and joint probability distributions in quantum mechanics. Foundations of Physics, 9(1–2):71–122, February 1979. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00715052.

Sorkin:1979:FTE

[368] Rafael Sorkin. On the failure of the time-energy uncertainty principle. Foundations of Physics, 9(1–2):123–128, February 1979. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00715053.

Post:1979:LTR

[369] E. J. Post. The logic of time reversal. Foundations of Physics, 9(1– 2):129–161, February 1979. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00715054.

Aron:1979:SBM

[370] J. C. Aron. A stochastic basis for microphysics. Foundations of Physics, 9(3–4):163–191, April 1979. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00715178.

Melsheimer:1979:CCM

[371] Olaf Melsheimer. On the classical content of many-body quantum mechanics. Foundations of Physics, 9(3–4):193–215, April 1979. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00715179.

Rayski:1979:CPM

[372] J. Rayski. Controversial problems of measurements within quantum mechanics. Foundations of Physics, 9(3–4):217–236, April 1979. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00715180.

Moore:1979:CSP

[373] Steven M. Moore. Can stochastic physics be a complete theory of nature? Foundations of Physics, 9(3–4):237–259, April 1979. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00715181.

Mickens:1979:LRI

[374] Ronald E. Mickens. Long-range interactions. Foundations of Physics, 9 (3–4):261–269, April 1979. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00715182.

Shah:1979:GTC

[375] K. Tahir Shah. A general theory of conservation laws, their violation, and spontaneous phenomena. Foundations of Physics, 9(3–4):271–282, April 1979. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00715183.

Mirman:1979:NSR

[376] R. Mirman. Nonexistence of superselection rules: Definition of term frame of reference. Foundations of Physics, 9(3–4):283–299, April 1979. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00715184.

Stuart:1979:MSB

[377] C. I. J. M. Stuart, Y. Takahashi, and H. Umezawa. Mixed-system brain dynamics: Neural memory as a macroscopic ordered state. Foundations of Physics, 9(3–4):301–327, April 1979. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00715185.

Feenberg:1979:DSO

[378] Eugene Feenberg. Distant synchrony and the one-way velocity of light. Foundations of Physics, 9(5–6):329–337, June 1979. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00708525.

Schiavulli:1979:FCE

[379] L. Schiavulli and F. Selleri. Further consequences of Einstein locality. Foundations of Physics, 9(5–6):339–352, June 1979. CODEN FNDPA4.

ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00708526.

Gron:1979:RDR

[380] Ø. Grøn. Relativistic description of a rotating disk with angular acceleration. Foundations of Physics, 9(5–6):353–369, June 1979. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:

//link.springer.com/article/10.1007/BF00708527. See comments

[422, 423].

Walstad:1979:TEU

[381] Allan Walstad. Tachyons in an expanding universe. Foundations of Physics, 9(5–6):371–374, June 1979. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00708528.

Kim:1979:PBM

[382] Y. S. Kim and Marilyn E. Noz. Physical basis for minimal time-energy uncertainty relation. Foundations of Physics, 9(5–6):375–387, June 1979. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00708529.

Mugur-Schachter:1979:SWT

[383] M. Mugur-Schächter. Study of Wigner’s theorem on joint probabilities. Foundations of Physics, 9(5–6):389–404, June 1979. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00708530.

Lavenda:1979:VOM

[384] B. H. Lavenda. On the validity of the Onsager–Machlup postulate for nonlinear stochastic processes. Foundations of Physics, 9(5–6): 405–420, June 1979. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00708531.

Post:1979:CPT

[385] E. J. Post. Comments on “Parity and time inversion symmetries of electromagnetic systems” by R. M. Kiehn. Foundations of Physics, 9 (5–6):421–424, June 1979. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00708532. See [280, 386].

Kiehn:1979:RPC

[386] R. M. Kiehn. Reply to Post’s comments on “Parity and time inversion symmetries of electromagnetic systems”. Foundations of Physics, 9(5– 6):425–426, June 1979. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00708533. See [280, 385].

Giannoni:1979:CRA

[387] Carlo Giannoni. Clock retardation, absolute space, and special relativity. Foundations of Physics, 9(5–6):427–444, June 1979. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708534.

Marinov:1979:CTA

[388] Stefan Marinov. The coordinate transformations of the absolute space– time theory. Foundations of Physics, 9(5–6):445–460, June 1979. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00708535.

Cooper:1979:FTL

[389] J. C. Cooper. Have faster-than-light particles already been detected? Foundations of Physics, 9(5–6):461–466, June 1979. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00708536.

Koga:1979:CNM

[390] Toyoki Koga. Causal nature of the motion of wavelets. Foundations of Physics, 9(5–6):467–470, June 1979. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00708537.

Bell:1979:EHO

[391] A. B. Bell and D. M. Bell. Energy in a highly ordered universe. Foundations of Physics, 9(5–6):471–477, June 1979. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00708538.

Margenau:1980:WY

[392] Henry Margenau. Wolfgang Yourgrau (1908–1979). Foundations of Physics, 10(1–2):3–6, February 1980. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00709013.

Frescura:1980:IOA

[393] F. A. M. Frescura and B. J. Hiley. The implicate order, algebras, and the spinor. Foundations of Physics, 10(1–2):7–31, February 1980. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00709014.

Young:1980:QMB

[394] Ralph H. Young. Quantum mechanics based on position. Foundations of Physics, 10(1–2):33–56, February 1980. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00709015.

Hsu:1980:CTF

[395] J. P. Hsu and T. N. Sherry. Common time in a four-dimensional symmetry framework. Foundations of Physics, 10(1–2):57–76, February 1980. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00709016.

Lee:1980:PFQ

[396] V. J. Lee. Physical foundations of quantum theory: Stochastic formulation and proposed experimental test. Foundations of Physics, 10(1– 2):77–107, February 1980. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00709017.

Howard:1980:WEE

[397] Roger Howard. Work and energy in electrostatic and magnetic systems. Foundations of Physics, 10(1–2):109–136, February 1980. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00709018.

Garcia:1980:SMP

[398] Luis García and Juan M. Usón. A simple model for Poincaré selfstresses. Foundations of Physics, 10(1–2):137–149, February 1980. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00709019.

Harris:1980:RAC

[399] Jonathan H. Harris and Mark D. Semon. A review of the Aharonov– Carmi thought experiment concerning the inertial and electromagnetic vector potentials. Foundations of Physics, 10(1–2):151–162, February 1980. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00709020.

Kellerman:1980:HFR

[400] Peter L. Kellerman. Heat flow in relativistic equilibrium thermodynamics. Foundations of Physics, 10(1–2):163–173, February 1980. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00709021.

Surdin:1980:FVS

[401] M. Surdin. A further visit to the steady state universe. Foundations of Physics, 10(1–2):175–179, February 1980. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00709022.

Corben:1980:MPR

[402] H. C. Corben. Mach’s Principle and the relation GM = Rc^{2}. Foundations of Physics, 10(1–2):181–184, February 1980. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00709023.

Ringhofer:1980:WCT

[403] K. Ringhofer and H. Salecker. What can be tested in quantum electrodynamics? Foundations of Physics, 10(3–4):185–196, April 1980. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00715066.

Prokhovnik:1980:ORD

[404] S. J. Prokhovnik. The operation of the relativistic Doppler effect. Foundations of Physics, 10(3–4):197–208, April 1980. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00715067.

Garuccio:1980:SDA

[405] A. Garuccio and F. Selleri. Systematic derivation of all the inequalities of Einstein locality. Foundations of Physics, 10(3–4):209–216, April 1980. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00715068.

Thankappan:1980:MSF

[406] V. K. Thankappan and P. Gopalakrishna Nambi. A modified set of Feynman postulates in quantum mechanics. Foundations of Physics, 10 (3–4):217–236, April 1980. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00715069.

Muraskin:1980:SSA

[407] M. Muraskin. Sinusoidal solutions to the aesthetic field equations. Foundations of Physics, 10(3–4):237–242, April 1980. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00715070.

Josephson:1980:AFA

[408] J. Josephson. An alternative formulation for the analysis and interpretation of the Dirac hydrogen atom. Foundations of Physics, 10(3– 4):243–266, April 1980. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00715071.

Smrz:1980:GFT

[409] P. K. Smrz. A gauge field theory of spacetime based on the de Sitter group. Foundations of Physics, 10(3–4):267–280, April 1980. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00715072.

Ikeda:1980:CNA

[410] Satoshi Ikeda. On the concept of “nonlocalization” associated with the gravitational field. Foundations of Physics, 10(3–4):281–287, April 1980. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00715073.

Phipps:1980:DMS

[411] T. E. Phipps, Jr. Do metric standards contract? Foundations of Physics, 10(3–4):289–307, April 1980. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00715074.

Fujiwara:1980:LVV

[412] Kunio Fujiwara. Is the light velocity in vacuum really a constant? Possible breakdown of the linear ω–k relation at extremely high frequencies. Foundations of Physics, 10(3–4):309–331, April 1980. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00715075.

Ohrstrom:1980:CDS

[413] Peter Øhrstrøm. Conventionality in distant simultaneity. Foundations of Physics, 10(3–4):333–343, April 1980. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00715076.

Schlegel:1980:LCE

[414] Richard Schlegel. The light clock: Error and implications. Foundations of Physics, 10(3–4):345–351, April 1980. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00715077.

Namsrai:1980:RDS

[415] Khavtgain Namsrai. Relativistic dynamics of stochastic particles. Foundations of Physics, 10(3–4):353–361, April 1980. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00715078.

Cooke:1980:APM

[416] R. M. Cooke and J. Hilgevoord. The algebra of physical magnitudes. Foundations of Physics, 10(5–6):363–373, June 1980. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00708739.

Melsheimer:1980:RM

[417] Olaf Melsheimer. On the reconstruction of the macroobservable. Foundations of Physics, 10(5–6):375–401, June 1980. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00708740.

Rietdijk:1980:MEF

[418] C. W. Rietdijk. A microrealistic explanation of fundamental quantum phenomena. Foundations of Physics, 10(5–6):403–457, June 1980. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00708741.

Huleihil:1980:MOR

[419] Kh. Huleihil and S. Malin. The mass operator in Riemannian spacetimes. Foundations of Physics, 10(5–6):459–467, June 1980. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00708742.

Vroegindewey:1980:CPO

[420] P. G. Vroegindewey, V. Ja. Kreinovic, and O. M. Kosheleva. From a connected, partially ordered set of events to a partially ordered field of time intervals. Foundations of Physics, 10(5–6):469–484, June 1980. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00708743.

Steeb:1980:CGL

[421] W.-H. Steeb. A comment on the generalized Liouville equation. Foundations of Physics, 10(5–6):485–493, June 1980. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00708744.

Grunbaum:1980:RDR

[422] Adolf Grünbaum and Allen I. Janis. The rotating disk: Reply to Grøn. Foundations of Physics, 10(5–6):495–498, June 1980. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00708745. See [380, 423].

Gron:1980:RDR

[423] Øyvind Grøn. The rotating disk: Reply to Grünbaum and Janis. Foundations of Physics, 10(5–6):499–501, June 1980. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00708746. See [380, 422].

Wesley:1980:EDS

[424] J. P. Wesley. Einstein dynamics without special-relativistic kinematics. Foundations of Physics, 10(5–6):503–511, June 1980. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00708747.

Brown:1981:CDT

[425] Harvey R. Brown and Michael L. G. Redhead. A critique of the disturbance theory of indeterminacy in quantum mechanics. Foundations of Physics, 11(1–2):1–20, February 1981. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00715192.

Shirokov:1981:FFP

[426] M. Shirokov. Fermi–Ferretti problem and signal velocity. Foundations of Physics, 11(1–2):21–36, February 1981. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00715193.

Laue:1981:QMP

[427] Hans Laue. Quantum mechanics without the position operator. Foundations of Physics, 11(1–2):37–46, February 1981. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00715194.

Simmons:1981:CPM

[428] Ralph F. Simmons, Jr. and James L. Park. On completely positive maps in generalized quantum dynamics. Foundations of Physics, 11(1– 2):47–55, February 1981. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.

1007/BF00715195. See remarks [474].

Mayants:1981:PTF

[429] L. S. Mayants. On the probabilistic treatment of fields. Foundations of Physics, 11(1–2):57–76, February 1981. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00715196.

Aron:1981:FR

[430] J. C. Aron. The foundations of relativity. Foundations of Physics, 11(1– 2):77–101, February 1981. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00715197.

Gordon:1981:BMM

[431] L. G. M. Gordon. Brownian movement and microscopic irreversibility. Foundations of Physics, 11(1–2):103–113, February 1981. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00715198.

Marinov:1981:QDE

[432] Stefan Marinov. The quasi-Doppler experiment according to absolute space–time theory. Foundations of Physics, 11(1–2):115–120, February 1981. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00715199.

Aharonov:1981:OAT

[433] Y. Aharonov and M. Vardi. An operational approach for testing the postulate of measurement in quantum theory. Foundations of Physics, 11(1– 2):121–125, February 1981. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00715200.

Pulmannova:1981:OQL

[434] S. Pulmannová. On the observables on quantum logics. Foundations of Physics, 11(1–2):127–136, February 1981. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00715201.

Asanov:1981:FEG

[435] G. S. Asanov. A Finslerian extension of general relativity. Foundations of Physics, 11(1–2):137–154, February 1981. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00715202.

deWet:1981:TNM

[436] J. A. de Wet. On the topology of nuclear manifolds. Foundations of Physics, 11(1–2):155–169, February 1981. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00715203.

Tornqvist:1981:SEP

[437] Nils A. Törnqvist. Suggestion for Einstein–Podolsky–Rosen experiments using reactions like e^{+}e^{−} → ΛΛ̄ → π^{−}pπ^{+}p̄. Foundations of Physics, 11(1–2):171–177, February 1981. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00715204.

Bohm:1981:QAA

[438] D. Bohm and B. J. Hiley. On a quantum algebraic approach to a generalized phase space. Foundations of Physics, 11(3–4):179–203, April 1981. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00726266.

dEspagnat:1981:CIA

[439] B. d’Espagnat. The concepts of influences and of attributes as seen in connection with Bell’s theorem. Foundations of Physics, 11(3–4): 205–234, April 1981. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00726267.

Vargas:1981:NPL

[440] J. G. Vargas. Nonrelativistic para-Lorentzian mechanics. Foundations of Physics, 11(3–4):235–278, April 1981. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00726268.

Bailey:1981:MPS

[441] Cecil D. Bailey. On a more precise statement of Hamilton’s principle. Foundations of Physics, 11(3–4):279–296, April 1981. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:

//link.springer.com/article/10.1007/BF00726269. See comments

[569, 568].

Simmons:1981:ENL

[442] Ralph F. Simmons, Jr. and James L. Park. The essential nonlinearity of N-level quantum thermodynamics. Foundations of Physics, 11(3– 4):297–305, April 1981. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00726270.

Dodson:1981:RCC

[443] C. T. J. Dodson and R. W. Tucker. De-Rham currents and charged particle interactions in electromagnetic and gravitational fields. Foundations of Physics, 11(3–4):307–328, April 1981. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00726271.

Sachs:1981:EPP

[444] Mendel Sachs. Elementary particle physics from general relativity. Foundations of Physics, 11(3–4):329–354, April 1981. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00726272.

Prugovecki:1981:SCA

[445] Eduard Prugovecki. A self-consistent approach to quantum field theory for extended particles. Foundations of Physics, 11(5–6):355–382, June 1981. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00727074.

Santilli:1981:ILE

[446] Ruggero Maria Santilli. An intriguing legacy of Einstein, Fermi, Jordan, and others: The possible invalidation of quark conjectures. Foundations of Physics, 11(5–6):383–472, June 1981. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00727075.

Narlikar:1981:QCF

[447] J. V. Narlikar. Quantum conformal fluctuations near the classical space– time singularity. Foundations of Physics, 11(5–6):473–492, June 1981. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00727076.

Fanchi:1981:RKP

[448] John R. Fanchi. Resolution of the Klein paradox for spin-1/2 particles. Foundations of Physics, 11(5–6):493–498, June 1981. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00727077.

Anonymous:1981:E

[449] Anonymous. Erratum. Foundations of Physics, 11(5–6):499, June 1981. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00727078.

Faris:1982:SCS

[450] William G. Faris. Spin correlation in stochastic mechanics. Foundations of Physics, 12(1):1–26, January 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00726872.

Dewdney:1982:QPD

[451] C. Dewdney and B. J. Hiley. A quantum potential description of onedimensional time-dependent scattering from square barriers and square wells. Foundations of Physics, 12(1):27–48, January 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00726873.

Semon:1982:EVA

[452] Mark D. Semon. Experimental verification of an Aharonov–Bohm effect in rotating reference frames. Foundations of Physics, 12(1): 49–57, January 1982. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00726874.

Melsheimer:1982:FEQ

[453] Olaf Melsheimer. The formalism of equilibrium quantum statistical mechanics revisited. Foundations of Physics, 12(1):59–84, January 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00726875.

Wormald:1982:SPC

[454] Laurence I. Wormald. Schlegel’s photon clock theory. Foundations of Physics, 12(1):85–88, January 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00726876. See reply [455].

Schlegel:1982:SPC

[455] Richard Schlegel. Schlegel’s photon clock theory: a reply to Wormald. Foundations of Physics, 12(1):89–92, January 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00726877. See [454].

Muraskin:1982:AFN

[456] M. Muraskin. Aesthetic fields — null theory II. Foundations of Physics, 12(1):93–100, January 1982. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00726878.

Mermin:1982:JDL

[457] N. D. Mermin and Gina M. Schwarz. Joint distributions and local realism in the higher-spin Einstein–Podolsky–Rosen experiment. Foundations of Physics, 12(2):101–135, February 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00736844.

Agodi:1982:TOL

[458] A. Agodi and M. A. Cassarino. Time ordering and the Lorentz group. Foundations of Physics, 12(2):137–152, February 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00736845.

Hestenes:1982:STS

[459] David Hestenes. Space–time structure of weak and electromagnetic interactions. Foundations of Physics, 12(2):153–168, February 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00736846.

Post:1982:PDC

[460] E. J. Post. Physical dimensions and covariance. Foundations of Physics, 12(2):169–195, February 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00736847.

Muraskin:1982:ISR

[461] M. Muraskin. Introduction of a second-rank antisymmetric tensor into null aesthetic field theory. Foundations of Physics, 12(2):197– 205, February 1982. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00736848.

Brabec:1982:CQL

[462] J. Brabec and P. Pták. On compatibility in quantum logics. Foundations of Physics, 12(2):207–212, February 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00736849.

Rosen:1982:WGP

[463] Nathan Rosen. Weyl’s geometry and physics. Foundations of Physics, 12(3):213–248, March 1982. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00726849.

Pearle:1982:MGT

[464] Philip Pearle. Might God toss coins? Foundations of Physics, 12(3): 249–263, March 1982. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00726850.

Torr:1982:MRP

[465] D. G. Torr and P. Kolen. Misconceptions in recent papers on special relativity and absolute space theories. Foundations of Physics, 12 (3):265–284, March 1982. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00726851.

deWet:1982:NEL

[466] J. A. de Wet. On nuclear energy levels and elementary particles. Foundations of Physics, 12(3):285–300, March 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00726852.

Svetlichny:1982:ICS

[467] George Svetlichny. The instrumental complexity of states. Foundations of Physics, 12(3):301–326, March 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00726853.

Finch:1982:CPQ

[468] Peter D. Finch. Classical probability and the quantum mechanical trace formulation for expectations. Foundations of Physics, 12(4): 327–345, April 1982. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00726781.

Christensen:1982:PMB

[469] Jens Peter Christensen and Richard D. Mattuck. Partial measurement in the Bohm–Bub hidden-variable theory. Foundations of Physics, 12 (4):347–361, April 1982. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00726782.

Stapp:1982:MMQ

[470] Henry P. Stapp. Mind, matter, and quantum mechanics. Foundations of Physics, 12(4):363–399, April 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00726783.

Kolen:1982:EMO

[471] P. Kolen and D. G. Torr. An experiment to measure the one-way velocity of propagation of electromagnetic radiation. Foundations of Physics, 12 (4):401–411, April 1982. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00726784.

vBorzeszkowski:1982:RRB

[472] H.-H. v. Borzeszkowski and H.-J. Treder. Remarks on the relation between general relativity and quantum theory. Foundations of Physics, 12(4):413–418, April 1982. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00726785.

Gudder:1982:LEQ

[473] Stanley P. Gudder. A logical explanation for quarks. Foundations of Physics, 12(4):419–431, April 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00726786.

Raggio:1982:RCP

[474] G. A. Raggio and H. Primas. Remarks on “On Completely Positive Maps in Generalized Quantum Dynamics”. Foundations of Physics, 12 (4):433–435, April 1982. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00726787. See [428].

Simmons:1982:ALC

[475] Ralph F. Simmons, Jr. and James L. Park. Another look at complete positivity in generalized quantum dynamics: Reply to Raggio and Primas.

Foundations of Physics, 12(4):437–439, April 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00726788. See [474].

delaPena:1982:SEA

[476] L. de la Peña and A. Jáuregui. The spin of the electron according to stochastic electrodynamics. Foundations of Physics, 12(5):441–465, May 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00729994.

Woodward:1982:EIR

[477] James F. Woodward. Electrogravitational induction and rotation. Foundations of Physics, 12(5):467–478, May 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00729995.

Schroeck:1982:SMI

[478] Franklin E. Schroeck, Jr. On the stochastic measurement of incompatible spin components. Foundations of Physics, 12(5):479–497, May 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00729996.

Duffey:1982:ILD

[479] George H. Duffey. Interaction-line descriptions of fields. Foundations of Physics, 12(5):499–508, May 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00729997.

Gron:1982:MSC

[480] Ø. Grøn. The metric in a static cylindrical elastic medium and in an empty rotating frame as solutions of Einstein’s field equations. Foundations of Physics, 12(5):509–520, May 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00729998.

Barut:1982:ETF

[481] A. O. Barut, S. Malin, and M. Semon. Electrodynamics in terms of functions over the group SU(2): II. Quantization. Foundations of Physics, 12(5):521–530, May 1982. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00729999.

Steeb:1982:RCM

[482] W.-H. Steeb and David E. Miller. Relativistic classical mechanics and canonical formalism. Foundations of Physics, 12(5):531–542, May 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00730000.

Marinov:1982:PP

[483] Stefan Marinov. Are phonons particles? Foundations of Physics, 12 (5):543–550, May 1982. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00730001.

Marinov:1982:CWT

[484] S. Marinov. Contrary to Wilczyński, there is no aberration for comoving source and observer. Foundations of Physics, 12(5):551–553, May 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00730002.

Prugovecki:1982:TEU

[485] Eduard Prugovecki. Time-energy uncertainty and relativistic canonical commutation relations in quantum spacetime. Foundations of Physics, 12(6):555–564, June 1982. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00731928.

Schmidt:1982:CSV

[486] Helmut Schmidt. Collapse of the state vector and psychokinetic effect. Foundations of Physics, 12(6):565–581, June 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00731929.

Leaf:1982:CSQ

[487] Boris Leaf. The continuous spectra of quantum operators. Foundations of Physics, 12(6):583–606, June 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00731930.

Maxwell:1982:IPF

[488] Nicholas Maxwell. Instead of particles and fields: a micro realistic quantum “smearon” theory. Foundations of Physics, 12(6):607–631, June 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00731931.

vBorzeszkowski:1982:HFB

[489] H-H v. Borzeszkowski. On high frequency background quantization of gravity. Foundations of Physics, 12(6):633–643, June 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00731932.

Selleri:1982:GEP

[490] F. Selleri. Generalized EPR-paradox. Foundations of Physics, 12 (7):645–659, July 1982. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00729804.

Shale:1982:DQT

[491] David Shale. Discrete quantum theory. Foundations of Physics, 12 (7):661–687, July 1982. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00729805.

Sachs:1982:PMO

[492] Mendel Sachs. A pulsar model from an oscillating black hole. Foundations of Physics, 12(7):689–708, July 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00729806.

Recami:1982:ASR

[493] Erasmo Recami and Waldyr A. Rodrigues, Jr. Antiparticles from special relativity with ortho-chronous and antichronous Lorentz transformations. Foundations of Physics, 12(7):709–718, July 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00729807.

Poole:1982:PDM

[494] Charles P. Poole, Jr. and Horacio A. Farach. Pauli–Dirac matrix generators of Clifford algebras. Foundations of Physics, 12(7):719–738, July 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00729808.

Steeb:1982:CCL

[495] W.-H. Steeb, J. Schröter, and W. Erig. A comment on conservation laws and constants of motion. Foundations of Physics, 12(7):739–742, July 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00729809.

Lock:1982:TPW

[496] James A. Lock. The transformation properties of world lines in relativistic quantum mechanical Hamiltonian models. Foundations of Physics, 12 (8):743–757, August 1982. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00731687.

Finch:1982:SAB

[497] Peter D. Finch. State assembles and the Bell–Wigner locality argument. Foundations of Physics, 12(8):759–764, August 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00731688.

Vargas:1982:SPL

[498] J. G. Vargas. Spontaneous para-Lorentzian conserved-vector and nonconserved-axial weak currents. Foundations of Physics, 12(8): 765–779, August 1982. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00731689.

Schlegel:1982:GMD

[499] Richard Schlegel. Gravitation and mass decrease. Foundations of Physics, 12(8):781–795, August 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00731690.

Mayants:1982:PLS

[500] L. S. Mayants. Probabilistics: a lost science. Foundations of Physics, 12 (8):797–811, August 1982. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00731691.

Lee:1982:PFQ

[501] V. J. Lee. Physical foundations of quantum theory: Stochastic formulation and proposed experimental test — addendum and erratum. Foundations of Physics, 12(8):813–819, August 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00731692.

Bondi:1982:SKP

[502] Hermann Bondi. Sir Karl Popper — on his eightieth birthday. Foundations of Physics, 12(9):821–823, September 1982. CODEN FNDPA4.

ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01884994.

Schroeck:1982:TAC

[503] Franklin E. Schroeck, Jr. The transitions among classical mechanics, quantum mechanics, and stochastic quantum mechanics. Foundations of Physics, 12(9):825–841, September 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01884995.

Bergia:1982:HOT

[504] S. Bergia and F. Cannata. Higher-order tensors and tests of quantum mechanics. Foundations of Physics, 12(9):843–859, September 1982. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01884996.

deBeauregard:1982:MVM

[505] O. Costa de Beauregard. Mpt versus: a manifestly covariant presentation of motion reversal and particle-antiparticle exchange. Foundations of Physics, 12(9):861–871, September 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01884997.

Surdin:1982:BWR

[506] M. Surdin. De Broglie’s wave revisited. Foundations of Physics, 12(9): 873–888, September 1982. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01884998.

Vargas:1982:NPM

[507] Jose G. Vargas. Nonrelativistic para-Maxwellian electrodynamics with preferred reference frame in the universe. Foundations of Physics, 12(9): 889–905, September 1982. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01884999.

Goldman:1982:CSP

[508] Stanford Goldman. Carriers and sideband pairs and their analogues in physics and biology. Foundations of Physics, 12(9):907–917, September 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF01885000.

Anonymous:1982:E

[509] Anonymous. Erratum. Foundations of Physics, 12(9):919, September 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF01885001.

Vigier:1982:LBP

[510] Jean-Pierre Vigier. Louis de Broglie — physicist and thinker. Foundations of Physics, 12(10):923–930, October 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01889266.

Lochak:1982:EIL

[511] Georges Lochak. The evolution of the ideas of Louis de Broglie on the interpretation of wave mechanics. Foundations of Physics, 12(10): 931–953, October 1982. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.

1007/BF01889267. See comment [512].

vanderMerwe:1982:EPE

[512] Alwyn van der Merwe. Editorial postscript to “The evolution of the ideas of Louis de Broglie on the interpretation of wave mechanics”. Foundations of Physics, 12(10):955–962, October 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01889268. See [511].

deBeauregard:1982:RME

[513] O. Costa de Beauregard. Reminiscences on my early association with Louis de Broglie. Foundations of Physics, 12(10):963–969, October 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01889269.

Popper:1982:CNG

[514] Karl Popper. A critical note on the greatest days of quantum theory. Foundations of Physics, 12(10):971–976, October 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1007/

BF01889270. 10.1007/BF01889270.

eSilva:1982:CLB

[515] João Andrade e Silva. On the contribution of Louis de Broglie to the quantum theory of measurement. Foundations of Physics, 12(10): 977–987, October 1982. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889271.

Bell:1982:IPW

[516] J. S. Bell. On the impossible pilot wave. Foundations of Physics, 12 (10):989–999, October 1982. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889272.

Bohm:1982:BPW

[517] D. J. Bohm and B. J. Hiley. The de Broglie pilot wave theory and the further development of new insights arising out of it. Foundations of Physics, 12(10):1001–1016, October 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01889273.

delaPena:1982:DQM

[518] L. de la Peña and A. M. Cetto. Does quantum mechanics accept a stochastic support? Foundations of Physics, 12(10):1017–1037, October 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF01889274.

Anonymous:1982:PLB

[519] Anonymous. Photo of Louis de Broglie. Foundations of Physics, 12 (10):ii, October 1982. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889265.

Gudder:1982:SQM

[520] Stanley P. Gudder. Survey of a quark model. Foundations of Physics, 12(11):1041–1055, November 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01300545.

Gueret:1982:BWP

[521] Ph. Gueret and J.-P. Vigier. De Broglie’s wave particle duality in the stochastic interpretation of quantum mechanics: a testable physical assumption. Foundations of Physics, 12(11):1057–1083, November 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01300546.

Mugur-Schachter:1982:EPS

[522] Mioara Mugur-Schächter. Elucidation of the probabilistic structure of quantum mechanics and definition of a compatible joint probability.

Foundations of Physics, 12(11):1085, November 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01300547.

Selleri:1982:DOQ

[523] F. Selleri. On the direct observability of quantum waves. Foundations of Physics, 12(11):1087–1112, November 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01300548.

vBorzeszkowski:1982:QTE

[524] H.-H. v. Borzeszkowski and H.-J. Treder. Quantum theory and Einstein’s general relativity. Foundations of Physics, 12(11):1113–1129, November 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF01300549.

Aerts:1982:DMS

[525] Dirk Aerts. Description of many separated physical entities without the paradoxes encountered in quantum mechanics. Foundations of Physics, 12(12):1131–1170, December 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00729621.

Herbert:1982:FSC

[526] Nick Herbert. FLASH — a superluminal communicator based upon a new kind of quantum measurement. Foundations of Physics, 12(12): 1171–1179, December 1982. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00729622.

Franklin:1982:CEF

[527] Allan Franklin. Cooper’s evidence for faster-than-light particles. Foundations of Physics, 12(12):1181–1182, December 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00729623.

Cooper:1982:FTL

[528] J. C. Cooper. Faster-than-light particles: a reply to Franklin. Foundations of Physics, 12(12):1183–1193, December 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00729624.

Barut:1983:EPW

[529] Asim O. Barut and Alwyn van der Merwe. To Eugene Paul Wigner on his eightieth birthday. Foundations of Physics, 13(1):3–5, January 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01889406.

Barut:1983:CPT

[530] A. O. Barut. On conservation of parity and time reversal and composite models of particles. Foundations of Physics, 13(1):7–12, January 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01889407.

Biedenharn:1983:SPR

[531] L. C. Biedenharn. The “Sommerfeld Puzzle” revisited and resolved. Foundations of Physics, 13(1):13–34, January 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01889408.

Gudder:1983:AM

[532] Stanley P. Gudder. An approach to measurement. Foundations of Physics, 13(1):35–49, January 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01889409.

Judd:1983:AMT

[533] B. R. Judd. Angular-momentum theory and projective geometry. Foundations of Physics, 13(1):51–59, January 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01889410.

Lee:1983:WPS

[534] Hai-Woong Lee and Marlan O. Scully. The Wigner phase-space description of collision processes. Foundations of Physics, 13(1):61–72, January 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF01889411.

Moshinsky:1983:ADS

[535] Marcos Moshinsky. Accidental degeneracies and symmetry groups. Foundations of Physics, 13(1):73–82, January 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01889412.

OConnell:1983:WDF

[536] R. F. O’Connell. The Wigner distribution function — 50th birthday. Foundations of Physics, 13(1):83–92, January 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01889413.

Rund:1983:IIA

[537] Hanno Rund. Invariance identities associated with finite gauge transformations and the uniqueness of the equations of motion of a particle in a classical gauge field. Foundations of Physics, 13(1):93–114, January 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF01889414.

Bauer:1983:MAT

[538] W. G. Bauer and H. Salecker. Muonic atoms testing the electron propagator of quantum electrodynamics and the Higgs boson contribution. Foundations of Physics, 13(1):115–132, January 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01889415.

Souriau:1983:PG

[539] Jean-Marie Souriau. Physics and geometry. Foundations of Physics, 13 (1):133–151, January 1983. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889416.

Treder:1983:GIA

[540] Hans-Jürgen Treder. Galilei invariance, action-reaction principle, and center of mass theorem. Foundations of Physics, 13(1):153–160, January 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF01889417.

Wheeler:1983:GLC

[541] John Archibald Wheeler. The geometrostatic lattice cell. Foundations of Physics, 13(1):161–173, January 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01889418.

Wybourne:1983:BRR

[542] B. G. Wybourne. Branching rules and replicating representations. Foundations of Physics, 13(1):175–182, January 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01889419.

Neeman:1983:SDVa

[543] Yuval Ne’eman. Some double-valued representations of the linear groups. Foundations of Physics, 13(1):183, January 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01889420.

Anonymous:1983:PEW

[544] Anonymous. Photo of Eugene Wigner. Foundations of Physics, 13 (1):ii, January 1983. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889405.

Barut:1983:PDH

[545] Asim O. Barut and Alwyn van der Merwe. Paul Dirac on his eightieth birthday. Foundations of Physics, 13(2):187–188, February 1983. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01889479.

Barut:1983:NQE

[546] A. O. Barut and J. Kraus. Nonperturbative quantum electrodynamics: The Lamb shift. Foundations of Physics, 13(2):189–194, February 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01889480.

Cabrera:1983:MME

[547] Blas Cabrera and W. Peter Trower. Magnetic monopoles: Evidence since the Dirac conjecture. Foundations of Physics, 13(2):195–215, February 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF01889481.

Chew:1983:BP

[548] Geoffrey F. Chew. Bootstrapping the photon. Foundations of Physics, 13 (2):217–246, February 1983. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889482.

Cooper:1983:VIP

[549] Leon N. Cooper. Visible and invisible in physical theory. Foundations of Physics, 13(2):247–251, February 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01889483.

Petroni:1983:DAR

[550] Nicola Cufaro Petroni and Jean Pierre Vigier. Dirac’s aether in relativistic quantum mechanics. Foundations of Physics, 13(2):253–286, February 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF01889484.

Anonymous:1983:PPAa

[551] Anonymous. Photo of Paul Adrien Maurice Dirac. Foundations of Physics, 13(2):ii, February 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01889478.

Gursey:1983:DAA

[552] Feza Gürsey. A Dirac algebraic approach to supersymmetry. Foundations of Physics, 13(3):289–296, March 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01906178.

Halpern:1983:MGC

[553] Leopold Halpern. Manifestations of group covariance in a metric theory. Foundations of Physics, 13(3):297–310, March 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01906179.

Narlikar:1983:CVG

[554] J. V. Narlikar. Cosmologies with variable gravitational constant. Foundations of Physics, 13(3):311–323, March 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01906180.

Penrose:1983:STG

[555] Roger Penrose. Spinors and torsion in general relativity. Foundations of Physics, 13(3):325–339, March 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01906181.

Recami:1983:MLN

[556] Erasmo Recami. A modified large number theory with constant G. Foundations of Physics, 13(3):341–346, March 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01906182.

Rompe:1983:PTN

[557] R. Rompe and H.-J. Treder. Is physics at the threshold of a new stage of evolution? Foundations of Physics, 13(3):347–361, March 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01906183.

Rosen:1983:BWD

[558] Nathan Rosen. The bimetric Weyl–Dirac theory and the gravitational constant. Foundations of Physics, 13(3):363–372, March 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01906184.

Schwinger:1983:EMR

[559] Julian Schwinger. Electromagnetic mass revisited. Foundations of Physics, 13(3):373–383, March 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01906185.

Sudarshan:1983:FRD

[560] E. C. G. Sudarshan and N. Mukunda. Forms of relativistic dynamics with world line condition and separability. Foundations of Physics, 13 (3):385–393, March 1983. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01906186.

Taylor:1983:NAS

[561] J. G. Taylor. New avenues in supersymmetry and supergravity. Foundations of Physics, 13(3):395–407, March 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01906187.

Thirring:1983:EAI

[562] W. Thirring. On the equivalence of adiabatic invariance and the KMScondition. Foundations of Physics, 13(3):409–415, March 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01906188.

Anonymous:1983:PPAb

[563] Anonymous. Photo of Paul Adrien Maurice Dirac. Foundations of Physics, 13(3):i, March 1983. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01906177.

Mugur-Schachter:1983:EPS

[564] Mioara Mugur-Schächter. Elucidation of the probabilistic structure of quantum mechanics and definition of a compatible joint probability. Foundations of Physics, 13(4):419–465, April 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00730892.

Neeman:1983:SDVb

[565] Yuval Ne’eman. Some double-valued representations of the linear groups. Foundations of Physics, 13(4):467–480, April 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00730893.

Heywood:1983:NKS

[566] Peter Heywood and Michael L. G. Redhead. Nonlocality and the Kochen– Specker paradox. Foundations of Physics, 13(5):481–499, May 1983. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00729511.

Asanov:1983:GFE

[567] G. S. Asanov. Gravitational field equations based on Finsler geometry. Foundations of Physics, 13(5):501–527, May 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00729512.

Yurtsever:1983:CMP

[568] Ulvi Yurtsever. Comments on “A more precise statement of Hamilton’s principle”. Foundations of Physics, 13(5):529–537, May 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00729513. See [441, 569].

Bailey:1983:HLH

[569] Cecil D. Bailey. Hamilton’s law or Hamilton’s principle: a response to Ulvi Yurtsever. Foundations of Physics, 13(5):539–544, May 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00729514. See [441,

568].

Atkins:1983:FQV

[570] William K. Atkins. A fundamental quadratic variational principle underlying general relativity. Foundations of Physics, 13(5):545–552, May 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00729515.

Anonymous:1983:E

[571] Anonymous. Erratum. Foundations of Physics, 13(5):553, May 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00729516.

Rastall:1983:BI

[572] Peter Rastall. The bell inequalities. Foundations of Physics, 13(6): 555–570, June 1983. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00730098.

Fanchi:1983:RMB

[573] John R. Fanchi and Weldon J. Wilson. Relativistic many-body systems: Evolution-parameter formalism. Foundations of Physics, 13(6): 571–605, June 1983. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00730099.

Dodonov:1983:QIM

[574] V. V. Dodonov, V. I. Man’ko, and V. N. Rudenko. Quantum integrals of motion and gravity wave experiment: Measurements in pure quantum states. Foundations of Physics, 13(6):607–628, June 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00730100.

Ikeda:1983:SSF

[575] Satoshi Ikeda. Some structural features induced by the space–time metrical fluctuation in the theory of gravitational fields. Foundations of Physics, 13(6):629–636, June 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00730101.

Anonymous:1983:PGLa

[576] Anonymous. Photo of Günther Ludwig. Foundations of Physics, 13(7):638, July 1983. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889344.

Jenc:1983:GLF

[577] F. Jenc, W. Maass, O. Melsheimer, H. Neumann, and A. van der Merwe. Günther Ludwig and the foundations of physics. Foundations of Physics, 13(7):639–641, July 1983. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889345.

Gisin:1983:DQD

[578] N. Gisin. Dissipative quantum dynamics for systems periodic in time. Foundations of Physics, 13(7):643–654, July 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01889346.

Hartkamper:1983:FPP

[579] Alois Hartkämper and Heinz-Jürgen Schmidt. On the foundations of the physical probability concept. Foundations of Physics, 13(7): 655–672, July 1983. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889347.

Hellwig:1983:FQS

[580] K.-E. Hellwig and W. Stulpe. A formulation of quantum stochastic processes and some of its properties. Foundations of Physics, 13 (7):673–699, July 1983. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889348.

Kraus:1983:AIP

[581] K. Kraus. Aspects of the infrared problem in quantum electrodynamics. Foundations of Physics, 13(7):701–713, July 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01889349.

Maass:1983:SCK

[582] W. Maass. On the stability of a class of kinetic equations. Foundations of Physics, 13(7):715–729, July 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01889350.

Mayr:1983:CAA

[583] Dieter Mayr. A constructive-axiomatic approach to the Lie structure in general spacetime by the principle of approximative reproducibility. Foundations of Physics, 13(7):731–743, July 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01889351.

Melsheimer:1983:QSM

[584] Olaf Melsheimer. Quantum statistical mechanics as a construction of an embedding scheme. Foundations of Physics, 13(7):745–758, July 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01889352.

Anonymous:1983:PGLb

[585] Anonymous. Photo of Günther Ludwig. Foundations of Physics, 13 (8):760, August 1983. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01906268.

Neumann:1983:DPR

[586] Holger Neumann. The description of preparation and registration of physical systems and conventional probability theory. Foundations of Physics, 13(8):761–778, August 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01906269.

Barchielli:1983:SCT

[587] A. Barchielli, L. Lanz, and G. M. Prosperi. Statistics of continuous trajectories in quantum mechanics: Operation-valued stochastic processes. Foundations of Physics, 13(8):779–812, August 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01906270.

Foulis:1983:ROQ

[588] D. Foulis, C. Piron, and C. Randall. Realism, operationalism, and quantum mechanics. Foundations of Physics, 13(8):813–841, August 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01906271.

Randall:1983:POP

[589] C. H. Randall and D. J. Foulis. Properties and operational propositions in quantum mechanics. Foundations of Physics, 13(8):843–857, August 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF01906272.

Werner:1983:PUS

[590] Reinhard Werner. Physical uniformities on the state space of nonrelativisitic quantum mechanics. Foundations of Physics, 13(8):859– 881, August 1983. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01906273.

Pauling:1983:ASG

[591] Linus Pauling. Albert Szent–Györgyi on his ninetieth birthday. Foundations of Physics, 13(9):883–886, September 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00732063.

Cramer:1983:AET

[592] John G. Cramer. The arrow of electromagnetic time and the generalized absorber theory. Foundations of Physics, 13(9):887–902, September 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00732064.

Greenberger:1983:QL

[593] Daniel M. Greenberger. Quantization in the large. Foundations of Physics, 13(9):903–951, September 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00732065.

Taylor:1983:BR

[594] J. G. Taylor, Banesh Hoffmann, T. E. Phipps, Jr., and Marek Demianski. Book review. Foundations of Physics, 13(9):953–962, September 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00732066.

Gutkowski:1983:QMS

[595] D. Gutkowski and M. V. Valdes Franco. On the quantum mechanical superposition of macroscopically distinguishable states. Foundations of Physics, 13(10):963–986, October 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00729517. See comments [596].

Bedford:1983:CQM

[596] D. Bedford and D. Wang. Comments on “On the quantum mechanical superposition of macroscopically distinguishable states”. Foundations of Physics, 13(10):987–988, October 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00729518. See [595].

Gordon:1983:MDD

[597] L. G. M. Gordon. Maxwell’s demon and detailed balancing. Foundations of Physics, 13(10):989–997, October 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00729519.

Papp:1983:SBC

[598] E. Papp. The scale-breaking contributions of the Pauli–Jordan functions and of the Feynman propagators. Foundations of Physics, 13(10): 999–1011, October 1983. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00729520.

Chang:1983:STT

[599] T. Chang. A space–time theory with a privileged frame. Foundations of Physics, 13(10):1013–1022, October 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00729521.

Israelit:1983:GCB

[600] Mark Israelit and Nathan Rosen. A gauge-covariant bimetric theory of gravitation and electromagnetism. Foundations of Physics, 13(10): 1023–1045, October 1983. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00729522.

Wehrl:1983:BR

[601] Alfred Wehrl. Book review. Foundations of Physics, 13(10):1047– 1049, October 1983. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00729523.

Malin:1983:BR

[602] Shimon Malin. Book review. Foundations of Physics, 13(10):1051– 1053, October 1983. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00729524.

Desloge:1983:BR

[603] Edward A. Desloge. Book review. Foundations of Physics, 13(10): 1055–1057, October 1983. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00729525.

Fryberger:1983:MSP

[604] David Fryberger. A model for the structure of point-like fermions: Qualitative features and physical description. Foundations of Physics, 13(11): 1059–1100, November 1983. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00728138.

Goedecke:1983:SES

[605] G. H. Goedecke. Stochastic electrodynamics. I. On the stochastic zeropoint field. Foundations of Physics, 13(11):1101–1119, November 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00728139.

Goedecke:1983:SEIa

[606] G. H. Goedecke. Stochastic electrodynamics. II. The harmonic oscillatorzero-point field system. Foundations of Physics, 13(11):1121–1138, November 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00728140.

Yamasaki:1983:ETH

[607] H. Yamasaki. Extension of trigonometric and hyperbolic functions to vectorial arguments and its application to the representation of rotations and Lorentz transformations. Foundations of Physics, 13(11): 1139–1154, November 1983. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00728141.

Papp:1983:LCA

[608] E. Papp. Light-cone approach to the quantum space–time description. Foundations of Physics, 13(11):1155–1165, November 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00728142.

Arshansky:1983:PVE

[609] R. Arshansky, L. P. Horwitz, and Y. Lavie. Particles vs. events: The concatenated structure of world lines in relativistic quantum mechanics. Foundations of Physics, 13(12):1167–1194, December 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00727992.

Goedecke:1983:SEIb

[610] G. H. Goedecke. Stochastic electrodynamics. III. Statistics of the perturbed harmonic oscillator-zero-point field system. Foundations of Physics, 13(12):1195–1220, December 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00727993.

Cavalleri:1983:PSS

[611] G. Cavalleri and G. Spinelli. Problems of synchronization in special relativity and possible links with stochastic electrodynamics. Foundations of Physics, 13(12):1221–1229, December 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00727994. See comments [612, 613, 614].

Vargas:1983:PSS

[612] Jose G. Vargas. Problems of synchronization in special relativity: a reply to G. Cavalleri and G. Spinelli. Foundations of Physics, 13(12): 1231–1235, December 1983. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.

1007/BF00727995. See [611, 613, 614].

Chang:1983:PSS

[613] Tsao Chang. Problems of synchronization in special relativity: Comments on a paper by G. Cavalleri and G. Spinelli. Foundations of Physics, 13(12):1237–1240, December 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00727996. See [611, 612, 614].

Marinov:1983:PSS

[614] Stefan Marinov. Problems of synchronization in special relativity: a reply to G. Cavalleri and G. Spinelli. Foundations of Physics, 13(12): 1241–1244, December 1983. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.

1007/BF00727997. See [611, 612, 613].

Garg:1984:FLN

[615] Anupam Garg and N. D. Mermin. Farkas’s lemma and the nature of reality: Statistical implications of quantum correlations. Foundations of Physics, 14(1):1–39, January 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00741645.

Goedecke:1984:SEI

[616] G. H. Goedecke. Stochastic electrodynamics. IV. Transitions in the perturbed harmonic oscillator-zero-point field system. Foundations of Physics, 14(1):41–63, January 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00741646.

Foulis:1984:NMP

[617] D. J. Foulis and C. H. Randall. A note on misunderstandings of Piron’s axioms for quantum mechanics. Foundations of Physics, 14 (1):65–81, January 1984. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00741647. See reply [618].

Hadjisavvas:1984:PAQ

[618] N. Hadjisavvas and F. Thieffine. Piron’s axioms for quantum mechanics: a reply to Foulis and Randall. Foundations of Physics, 14(1): 83–88, January 1984. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00741648. See [617].

Eastman:1984:BR

[619] Timothy E. Eastman and Evan Fales. Book review. Foundations of Physics, 14(1):89–99, January 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00741649.

McCormack:1984:BR

[620] Francis J. McCormack. Book review. Foundations of Physics, 14(1): 101–102, January 1984. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00741650.

deBeauregard:1984:BR

[621] Olivier Costa de Beauregard. Book review. Foundations of Physics, 14 (1):103–105, January 1984. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00741651.

Arms:1984:BR

[622] Judith M. Arms. Book review. Foundations of Physics, 14(1):107– 109, January 1984. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00741652.

deMuynck:1984:DLC

[623] W. M. de Muynck and J. P. H. W. van den Eijnde. A derivation of local commutativity from macrocausality using a quantum mechanical theory of measurement. Foundations of Physics, 14(2):111–146, February 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00729970.

Domingos:1984:SAM

[624] J. M. Domingos and M. H. Caldeira. Self-adjointness of momentum operators in generalized coordinates. Foundations of Physics, 14(2): 147–154, February 1984. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00729971.

Wesley:1984:RCW

[625] J. P. Wesley. A resolution of the classical wave-particle problem. Foundations of Physics, 14(2):155–170, February 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00729972.

Rosen:1984:GFU

[626] Nathan Rosen and Gerald E. Tauber. A geometric foundation for a unified field theory. Foundations of Physics, 14(2):171–186, February 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00729973.

Hestenes:1984:BR

[627] David Hestenes. Book review. Foundations of Physics, 14(2):187– 191, February 1984. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00729974.

Horwitz:1984:BR

[628] L. P. Horwitz. Book review. Foundations of Physics, 14(2):193– 198, February 1984. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00729975.

deMuynck:1984:QMT

[629] Willem M. de Muynck. A quantum mechanical theory of local observables and local operations. Foundations of Physics, 14(3):199–253, March 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00730210.

Bohm:1984:MUT

[630] D. Bohm and B. J. Hiley. Measurement understood through the quantum potential approach. Foundations of Physics, 14(3):255–274, March 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00730211.

Prigogine:1984:BR

[631] Ilya Prigogine. Book review. Foundations of Physics, 14(3):275– 277, March 1984. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00730212.

Leipholz:1984:BR

[632] H. H. E. Leipholz. Book review. Foundations of Physics, 14(3): 279–280, March 1984. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00730213.

Finch:1984:OFQ

[633] Peter D. Finch. The operator formalism of quantum mechanics from the viewpoint of short disturbances in nonrelativistic classical motion. Foundations of Physics, 14(4):281–306, April 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00738920.

Hiskes:1984:STT

[634] Anne L. D. Hiskes. Space–time theories and symmetry groups. Foundations of Physics, 14(4):307–332, April 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00738921.

Bishop:1984:STT

[635] N. T. Bishop. Is superluminal travel a theoretical possibility? Foundations of Physics, 14(4):333–340, April 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00738922.

Perkins:1984:SEP

[636] Granville A. Perkins. A spinor equation of the pure electromagnetic field. II. Foundations of Physics, 14(4):341–349, April 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00738923.

Murphy:1984:NMV

[637] George L. Murphy. On the “nonclassical many-valuedness” of the universe. Foundations of Physics, 14(4):351–357, April 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00738924.

Roskies:1984:BR

[638] Ralph Roskies. Book review. Foundations of Physics, 14(4):359– 361, April 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00738925.

Derrick:1984:BR

[639] Malcolm Derrick and Pierre Ramond. Book review. Foundations of Physics, 14(4):363–365, April 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00738926.

Maccarrone:1984:ISL

[640] G. D. Maccarrone and Erasmo Recami. The introduction of superluminal Lorentz transformations: a revisitation. Foundations of Physics, 14(5):367–407, May 1984. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00738808.

Brody:1984:QL

[641] T. A. Brody. On quantum logic. Foundations of Physics, 14(5): 409–430, May 1984. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00738809.

Gaudard:1984:CBC

[642] Marie Gaudard. The correspondence between credibilities and induced betting rate assignments. Foundations of Physics, 14(5):431–441, May 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00738810.

Narlikar:1984:VLS

[643] J. V. Narlikar. The vanishing likelihood of space–time singularity in quantum conformal cosmology. Foundations of Physics, 14(5):

443–456, May 1984. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00738811.

Toms:1984:BR

[644] David J. Toms and Michael Conrad. Book review. Foundations of Physics, 14(5):457–465, May 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00738812.

Greider:1984:UCA

[645] K. R. Greider. A unifying Clifford algebra formalism for relativistic fields. Foundations of Physics, 14(6):467–506, June 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00736596.

Landau:1984:VWT

[646] B. V. Landau. Variations on a Wignerian theme. Foundations of Physics, 14(6):507–534, June 1984. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00736597.

Girard:1984:TCP

[647] Réjean Girard and Louis Marchildon. Are tachyon causal paradoxes solved? Foundations of Physics, 14(6):535–548, June 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00736598.

Finch:1984:BPD

[648] Peter D. Finch and Roman Groblicki. Bivariate probability densities with given margins. Foundations of Physics, 14(6):549–552, June 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00736599. See note

[902].

Broyles:1984:QMS

[649] A. A. Broyles. Quantum mechanics of seeing. Foundations of Physics, 14(6):553–560, June 1984. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00736600.

Franklin:1984:BR

[650] Allan Franklin. Book review. Foundations of Physics, 14(6):561– 563, June 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00736601.

Just:1984:BR

[651] Kurt Just. Book review. Foundations of Physics, 14(6):565–567, June 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00736602.

Klein:1984:BR

[652] Abraham Klein, L. E. Reichl, and Paul J. Wiita. Book review. Foundations of Physics, 14(6):569–577, June 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00736603.

Rosen:1984:SIW

[653] Nathan Rosen. A semiclassical interpretation of wave mechanics. Foundations of Physics, 14(7):579–605, July 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00738743.

Domingos:1984:QHC

[654] J. M. Domingos and M. H. Caldeira. The quantization of the Hamiltonian in curved space. Foundations of Physics, 14(7):607–623, July 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00738744.

Vargas:1984:RRT

[655] José G. Vargas. Revised Robertson’s test theory of special relativity. Foundations of Physics, 14(7):625–651, July 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00738745.

vonBorzeszkowski:1984:DNI

[656] Horst-Heino von Borzeszkowski and Renate Wahsner. Did nonlinear irreversible thermodynamics revolutionize the classical time conception of physics? Foundations of Physics, 14(7):653–670, July 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00738746.

Kilmister:1984:BR

[657] C. W. Kilmister, F. S. Levin, and Burt Ovrut. Book review. Foundations of Physics, 14(7):671–681, July 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00738747.

Hamermesh:1984:BR

[658] Morton Hamermesh. Book review. Foundations of Physics, 14 (7):683, July 1984. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00738748.

Longtin:1984:RCB

[659] Luc Longtin and Richard D. Mattuck. Relativistically covariant Bohm– Bub hidden-variable theory for spin measurement of a single particle. Foundations of Physics, 14(8):685–703, August 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00736616.

Kelly:1984:RDE

[660] Edward M. Kelly. Relativistic Doppler effect in light clocks construed as a result of prior acceleration. Foundations of Physics, 14 (8):705–720, August 1984. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00736617.

Embacher:1984:AEM

[661] Franz Embacher. The analog of electric and magnetic fields in stationary gravitational systems. Foundations of Physics, 14(8):721–738, August 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00736618.

Burgos:1984:OIO

[662] M. E. Burgos. An objective interpretation of orthodox quantum mechanics. Foundations of Physics, 14(8):739–752, August 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00736619.

Burgos:1984:OFO

[663] M. E. Burgos. An objective formulation of orthodox quantum mechanics. Foundations of Physics, 14(8):753–766, August 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00736620.

vanderSpuy:1984:WPF

[664] E. van der Spuy. Wave, particle-family duality and the conservation of discrete symmetries in strong interaction. Foundations of Physics, 14 (8):767–775, August 1984. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00736621.

Salingaros:1984:AFD

[665] Nikos Salingaros and Yehiel Ilamed. Algebraic field descriptions in three-dimensional Euclidean space. Foundations of Physics, 14(8): 777–797, August 1984. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00736622.

Fraser:1984:BR

[666] J. T. Fraser and Joseph Macek. Book review. Foundations of Physics, 14 (8):799–807, August 1984. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00736623.

Rindler:1984:BR

[667] Wolfgang Rindler, R. Bruce Lindsay, O. Costa de Beauregard, and Barry Barish. Book review. Foundations of Physics, 14(8):809–819, August 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00736624.

Guz:1984:SPS

[668] W. Guz. Stochastic phase spaces, fuzzy sets, and statistical metric spaces. Foundations of Physics, 14(9):821–848, September 1984. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00737552.

Leiter:1984:OIC

[669] Darryl Leiter. On the origin of irreversibility in classical electrodynamic measurement processes. Foundations of Physics, 14(9):849– 863, September 1984. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00737553.

Goenner:1984:TGN

[670] H. F. M. Goenner. Theories of gravitation with nonminimal coupling of matter and the gravitational field. Foundations of Physics, 14(9): 865–881, September 1984. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00737554.

Greenwood:1984:SMR

[671] D. P. Greenwood and E. Prugovecki. Stochastic microcausality in relativistic quantum mechanics. Foundations of Physics, 14(9):883– 906, September 1984. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00737555.

Gudder:1984:BR

[672] Stanley Gudder and Vladimir Naroditsky. Book review. Foundations of Physics, 14(9):907–913, September 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00737556.

Narducci:1984:BR

[673] L. M. Narducci. Book review. Foundations of Physics, 14(9):915– 917, September 1984. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00737557.

Stelle:1984:BR

[674] Kellogg S. Stelle and Holger Neumann. Book review. Foundations of Physics, 14(9):919–922, September 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00737558.

Carmeli:1984:NRH

[675] Moshe Carmeli and Alwyn van der Merwe. For Nathan Rosen on his seventy-fifth birthday. Foundations of Physics, 14(10):923–924, October 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF01889245.

Alexander:1984:ESI

[676] Matthew Alexander and Peter G. Bergmann. Electrodynamics at spatial infinity. Foundations of Physics, 14(10):925–951, October 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01889246.

Biedenharn:1984:CTC

[677] L. C. Biedenharn and L. P. Horwitz. Chiral two-component spinors and the factorization of Kramers’s equation. Foundations of Physics, 14(10): 953–961, October 1984. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889247.

Carmeli:1984:ERG

[678] M. Carmeli and Ch. Charach. The Einstein–Rosen gravitational waves and cosmology. Foundations of Physics, 14(10):963–986, October 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01889248.

Damour:1984:SFE

[679] Thibault Damour. Strong-field effects and time asymmetry in general relativity and in bimetric gravitation theory. Foundations of Physics, 14 (10):987–995, October 1984. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889249.

Gudder:1984:RLP

[680] Stanley P. Gudder. Reality, locality, and probability. Foundations of Physics, 14(10):997–1010, October 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01889250.

Halpern:1984:ULM

[681] Leopold Halpern. On the unification of the law of motion. Foundations of Physics, 14(10):1011–1026, October 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01889251.

Horwitz:1984:EIR

[682] L. P. Horwitz. On the electromagnetic interaction in relativistic quantum mechanics. Foundations of Physics, 14(10):1027–1046, October 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01889252.

Anonymous:1984:A

[683] Anonymous. Announcement. Foundations of Physics, 14(10):1047, October 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889253.

Anonymous:1984:PNRa

[684] Anonymous. Photo of Nathan Rosen. Foundations of Physics, 14 (10):i, October 1984. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889244.

Israel:1984:DCC

[685] W. Israel. Does a cosmic censor exist? Foundations of Physics, 14(11): 1049–1059, November 1984. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01882488.

Kozameh:1984:MEL

[686] Carlos N. Kozameh, Ezra T. Newman, and John R. Porter. Maxwell’s equations, linear gravity, and twistors. Foundations of Physics, 14(11): 1061–1081, November 1984. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01882489.

Malin:1984:MSQ

[687] S. Malin. The meaning and significance of quantum states. Foundations of Physics, 14(11):1083–1094, November 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01882490.

Mansouri:1984:NAI

[688] Freydoon Mansouri and Louis Witten. Non-Abelian isometries of multidimensional universes. Foundations of Physics, 14(11):1095– 1106, November 1984. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01882491.

Pachner:1984:TK

[689] J. Pachner. A theory of knowledge. Foundations of Physics, 14(11): 1107–1120, November 1984. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01882492.

Parker:1984:CDR

[690] Leonard Parker. Curvature dependence of renormalized coupling constants. Foundations of Physics, 14(11):1121–1129, November 1984. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01882493.

Peres:1984:CPQ

[691] Asher Peres. The classic paradoxes of quantum theory. Foundations of Physics, 14(11):1131–1145, November 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01882494.

Prugovecki:1984:SQM

[692] E. Prugovecki. The stochastic quantum mechanics approach to the unification of relativity and quantum theory. Foundations of Physics, 14(12): 1147–1162, December 1984. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889316.

Stachel:1984:DEB

[693] John Stachel. The dynamical equations of black-body radiation. Foundations of Physics, 14(12):1163–1168, December 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01889317.

Tauber:1984:GWE

[694] Gerald E. Tauber. Gravitational waves in an expanding universe. Foundations of Physics, 14(12):1169–1183, December 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01889318.

Weber:1984:GNA

[695] J. Weber. Gravitons, neutrinos, and antineutrinos. Foundations of Physics, 14(12):1185–1209, December 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01889319.

Goldberg:1984:HGR

[696] J. N. Goldberg. The Hamiltonian of general relativity on a null surface. Foundations of Physics, 14(12):1211–1216, December 1984. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01889320.

Moffat:1984:GTG

[697] John W. Moffat. Generalized theory of gravitation. Foundations of Physics, 14(12):1217–1252, December 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01889321.

Neeman:1984:GYM

[698] Yuval Ne’eman. Geometrization in the Yang–Mills, extended supergravity, and Klein–Kaluza versions. Foundations of Physics, 14(12): 1253, December 1984. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889322.

Anonymous:1984:PNRb

[699] Anonymous. Photo of Nathan Rosen. Foundations of Physics, 14 (12):i, December 1984. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889315.

Mashkevich:1985:QSD

[700] V. S. Mashkevich. Quantum statistical dynamics: Statistics origin, measurement, and irreversibility. Foundations of Physics, 15(1): 1–33, January 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00738735.

Stapp:1985:CVQ

[701] Henry P. Stapp. Consciousness and values in the quantum universe. Foundations of Physics, 15(1):35–47, January 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00738736.

Gaudard:1985:CPP

[702] Marie Gaudard. Convergence of posterior probabilities in the Bayesian inference strategy. Foundations of Physics, 15(1):49–62, January 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00738737.

Hestenes:1985:QMS

[703] David Hestenes. Quantum mechanics from self-interaction. Foundations of Physics, 15(1):63–87, January 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00738738.

Ranada:1985:GIP

[704] Antonio F. Rañada. A geometrical interpretation of the Pauli exclusion principle in classical field theory. Foundations of Physics, 15(1): 89–100, January 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00738739.

Gudder:1985:LEF

[705] Stanley P. Gudder. Linearity of expectation functionals. Foundations of Physics, 15(1):101–111, January 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00738740.

Knight:1985:BR

[706] Peter Knight. Book review. Foundations of Physics, 15(1):113– 114, January 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00738741.

Duffey:1985:BR

[707] George H. Duffey. Book review. Foundations of Physics, 15(1): 115–116, January 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00738742.

Guz:1985:NCP

[708] W. Guz. On the nonclassical character of the phase-space representations of quantum mechanics. Foundations of Physics, 15(2):121–128, February 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00735283.

Namsrai:1985:OCR

[709] Kh. Namsrai. On the origin of cosmic rays and the value of fundamental length. Foundations of Physics, 15(2):129–143, February 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00735284.

Bergia:1985:EET

[710] S. Bergia, F. Cannata, and V. Monzoni. Explicit examples of theories satisfying Bell’s inequalities: Do they miss their goal prior to contradicting experiments? Foundations of Physics, 15(2):145–154, February 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00735285.

Ai:1985:CQA

[711] H. B. Ai and J. P. Hsu. Can quarks always be confined by a linear potential? Foundations of Physics, 15(2):155–159, February 1985. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00735286.

Treder:1985:PLE

[712] H. J. Treder. The planckions as largest elementary particles and as smallest test bodies. Foundations of Physics, 15(2):161–166, February 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00735287.

Prokhovnik:1985:ASR

[713] S. J. Prokhovnik and K. P. Kovács. The application of special relativity to the right-angled lever. Foundations of Physics, 15(2):167–173, February 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00735288.

Carmeli:1985:FTTa

[714] M. Carmeli. Field theory on R × S^{3} topology. I: The Klein–Gordon and Schrödinger equations. Foundations of Physics, 15(2):175–184, February 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00735289.

Carmeli:1985:FTTb

[715] M. Carmeli and S. Malin. Field theory on R × S^{3} topology. II: The Weyl equation. Foundations of Physics, 15(2):185–191, February 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00735290.

vonBorzeszkowski:1985:MER

[716] H.-H. von Borzeszkowski and H. J. Treder. On the Majorana equations and the rest mass of neutrinos. Foundations of Physics, 15(2): 193–199, February 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00735291.

Peres:1985:EGB

[717] Asher Peres. Einstein, Gödel, Bohr. Foundations of Physics, 15(2): 201–205, February 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00735292.

Wignall:1985:BWN

[718] J. W. G. Wignall. De Broglie waves and the nature of mass. Foundations of Physics, 15(2):207–227, February 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00735293.

Santos:1985:BR

[719] Emilio Santos. Book review. Foundations of Physics, 15(2):229– 231, February 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00735294.

dEspagnat:1985:BR

[720] Bernard d’Espagnat. Book review. Foundations of Physics, 15(2): 233–235, February 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00735295.

Bracken:1985:BR

[721] A. J. Bracken. Book review. Foundations of Physics, 15(2):237– 239, February 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00735296.

Bergmann:1985:BR

[722] Peter G. Bergmann. Book review. Foundations of Physics, 15(2): 241–242, February 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00735297.

Gaillard:1985:BR

[723] Mary K. Gaillard. Book review. Foundations of Physics, 15(2):243– 244, February 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00735298.

Rowlinson:1985:BR

[724] J. S. Rowlinson. Book review. Foundations of Physics, 15(2):245– 246, February 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00735299.

West:1985:BR

[725] Geoffrey B. West. Book review. Foundations of Physics, 15(2):247– 249, February 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00735300.

Greechie:1985:BR

[726] Richard J. Greechie. Book review. Foundations of Physics, 15(2): 251–256, February 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00735301.

Halpern:1985:PAM

[727] Leopold Halpern. Paul Adrien Maurice Dirac (1902–1984). Foundations of Physics, 15(3):257–259, March 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00737316.

Neeman:1985:GYM

[728] Yuval Ne’eman. Geometrization in the Yang–Mills extended supergravity and Klein–Kaluza versions. Foundations of Physics, 15(3): 261–277, March 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00737317.

Schroeck:1985:NTP

[729] F. E. Schroeck, Jr. On the nonoccurrence of two paradoxes in the measurement scheme of stochastic quantum mechanics. Foundations of Physics, 15(3):279–302, March 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00737318.

Rietdijk:1985:PQM

[730] C. W. Rietdijk and F. Selleri. Proof of a quantum mechanical nonlocal influence. Foundations of Physics, 15(3):303–317, March 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00737319.

Naroditsky:1985:FOT

[731] Vladimir Naroditsky. Field operators and their spectral properties in finite-dimensional quantum field theory. Foundations of Physics, 15 (3):319–331, March 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00737320.

Kelly:1985:MED

[732] Edward M. Kelly. Maser ether-drift experiment questioned on basis of molecular light clock model. Foundations of Physics, 15(3): 333–337, March 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00737321.

Liboff:1985:GPF

[733] Richard L. Liboff. Geometrical properties of the Fermi energy. Foundations of Physics, 15(3):339–352, March 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00737322.

Breitenberger:1985:UMU

[734] Ernst Breitenberger. Uncertainty measures and uncertainty relations for angle observables. Foundations of Physics, 15(3):353–364, March 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00737323.

Israelit:1985:EDP

[735] Mark Israelit and Nathan Rosen. Einstein: Distant parallelism and electromagnetism. Foundations of Physics, 15(3):365–377, March 1985. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00737324.

Kerrighan:1985:CFT

[736] D. B. Kerrighan. The cosmic field tensor in bimetric general relativity. Foundations of Physics, 15(3):379–386, March 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00737325.

Chirikov:1985:BR

[737] Boris V. Chirikov. Book review. Foundations of Physics, 15(3): 387–389, March 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00737326.

Halzen:1985:BR

[738] Francis Halzen and Lawrence W. Fagg. Book review. Foundations of Physics, 15(3):391–397, March 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00737327.

Heller:1985:BR

[739] Michael Heller. Book review. Foundations of Physics, 15(3):399– 401, March 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00737328.

March:1985:BR

[740] Norman H. March. Book review. Foundations of Physics, 15(3): 403–405, March 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00737329.

Oplatka:1985:BR

[741] Avraham Oplatka. Book review. Foundations of Physics, 15(3): 407–408, March 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00737330.

Komar:1985:PGB

[742] Arthur Komar. For Peter G. Bergmann at seventy. Foundations of Physics, 15(4):409–410, April 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01889276.

Anderson:1985:GRS

[743] James L. Anderson. Gravitational radiation, source behavior, and the method of matched asymptotic expansions. Foundations of Physics, 15 (4):411–418, April 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889277.

Ashtekar:1985:LAD

[744] Abhay Ashtekar. Logarithmic ambiguities in the description of spatial infinity. Foundations of Physics, 15(4):419–431, April 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01889278.

Barut:1985:GRK

[745] A. O. Barut and Thomas Gornitz. On the gyromagnetic ratio in the Kaluza–Klein theories and the Schuster–Blackett law. Foundations of Physics, 15(4):433–437, April 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01889279.

Goldberg:1985:DBG

[746] Joshua N. Goldberg. Dirac brackets for general relativity on a null cone. Foundations of Physics, 15(4):439–450, April 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01889280.

Hehl:1985:KTS

[747] Friedrich W. Hehl. On the kinematics of the torsion of space–time. Foundations of Physics, 15(4):451–471, April 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01889281.

Komar:1985:FTC

[748] Arthur Komar. Field theoretic constraint formalism. Foundations of Physics, 15(4):473–485, April 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01889282.

Kozameh:1985:NSD

[749] Carlos N. Kozameh and Ezra T. Newman. A note on self-dual gravitational fields. Foundations of Physics, 15(4):487–495, April 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01889283.

Mashhoon:1985:GER

[750] Bahram Mashhoon. Gravitational effects of rotating masses. Foundations of Physics, 15(4):497–515, April 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01889284.

Rosen:1985:SSS

[751] Nathan Rosen. Some Schwarzschild solutions and their singularities. Foundations of Physics, 15(4):517–529, April 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01889285.

Anonymous:1985:PPB

[752] Anonymous. Photo of Peter Bergmann. Foundations of Physics, 15(4):ii, April 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889275.

Ianiro:1985:SNS

[753] N. Ianiro and J. L. Lebowitz. Stationary nonequilibrium solutions of model Boltzmann equation. Foundations of Physics, 15(5):531–544, May 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF01882480.

Rindler:1985:ABK

[754] Wolfgang Rindler. Analogies between Kruskal space and de Sitter space. Foundations of Physics, 15(5):545–551, May 1985. CODEN FNDPA4.

ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01882481.

Schmutzer:1985:SMP

[755] Ernst Schmutzer. Spinor matter in a particular 5-dimensional projective unified field theory. Foundations of Physics, 15(5):553–569, May 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01882482.

Schucking:1985:HGF

[756] E. L. Schucking. The homogeneous gravitational field. Foundations of Physics, 15(5):571–577, May 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01882483.

Treder:1985:PMG

[757] Hans-Jürgen Treder. On the physical meaning of gauge and supergauge in general-relativistic field theories. Foundations of Physics, 15 (5):579–604, May 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01882484.

Winicour:1985:LAF

[758] Jeffrey Winicour. Logarithmic asymptotic flatness. Foundations of Physics, 15(5):605–616, May 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01882485.

Hogan:1985:MCT

[759] Peter A. Hogan and Ivor Robinson. The motion of charged test particles in general relativity. Foundations of Physics, 15(5):617–627, May 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01882486.

Anonymous:1985:BPG

[760] Anonymous. Bibliography: Peter G. Bergmann. Foundations of Physics, 15(5):629–635, May 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01882487.

Steeb:1985:IDS

[761] W.-H. Steeb, M. Kloke, B. M. Spieker, and A. Kunick. Integrability of dynamical systems and the singular-point analysis. Foundations of Physics, 15(6):637–666, June 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00738293.

Levine:1985:IHV

[762] R. Y. Levine. Isospin as a hidden variable. Foundations of Physics, 15(6):667–676, June 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00738294.

Schroeck:1985:CSO

[763] Franklin E. Schroeck, Jr. Compatible stochastic observables that do not commute. Foundations of Physics, 15(6):677–681, June 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00738295.

Salingaros:1985:SRA

[764] Nikos Salingaros. Some remarks on the algebra of Eddington’s E numbers. Foundations of Physics, 15(6):683–691, June 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00738296.

Greider:1985:IIC

[765] Kenneth R. Greider. Inconsistencies in the interpretation of the conservation equations for spin-1/2 fields. Foundations of Physics, 15 (6):693–700, June 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00738297.

Arshansky:1985:LPR

[766] R. Arshansky and L. P. Horwitz. The Landau–Peierls relation and a causal bound in covariant relativistic quantum theory. Foundations of Physics, 15(6):701–715, June 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00738298.

Kraus:1985:MPQ

[767] K. Kraus. Measuring processes in quantum mechanics. II. The classical behavior of measuring instruments. Foundations of Physics, 15 (6):717–730, June 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00738299.

deBeauregard:1985:RDI

[768] O. Costa de Beauregard. Relativistic dynamics of interacting point particles: Central position of the Wheeler–Feynman scheme. Foundations of Physics, 15(6):731–736, June 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00738300.

Eccles:1985:BR

[769] John C. Eccles and C. J. Isham. Book review. Foundations of Physics, 15(6):737–743, June 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00738301.

Wehrl:1985:BR

[770] Alfred Wehrl and H. B. G. Casimir. Book review. Foundations of Physics, 15(6):745–749, June 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00738302.

Gisin:1985:BR

[771] N. Gisin. Book review. Foundations of Physics, 15(6):751–752, June 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00738303.

Cyranski:1985:TVE

[772] John F. Cyranski. Theory vs. experiment: a holistic philosophy of physics. Foundations of Physics, 15(7):753–771, July 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00739845.

Hagston:1985:ETR

[773] W. E. Hagston and I. D. Cox. An extended theory of relativity in a sixdimensional manifold. Foundations of Physics, 15(7):773–805, July 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00739846.

Evans:1985:ACS

[774] A. B. Evans. Accelerating charges and the Schott force. Foundations of Physics, 15(7):807–816, July 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00739847.

Stuewer:1985:BR

[775] Roger H. Stuewer. Book review. Foundations of Physics, 15(7): 817–818, July 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00739848.

Power:1985:BR

[776] E. A. Power. Book review. Foundations of Physics, 15(7):819–820, July 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00739849.

Matzner:1985:BR

[777] Richard A. Matzner. Book review. Foundations of Physics, 15(7): 821–822, July 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00739850.

Graham:1985:BR

[778] W. R. M. Graham. Book review. Foundations of Physics, 15(7): 823–824, July 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00739851.

Phipps:1985:BR

[779] Thomas E. Phipps, Jr. Book review. Foundations of Physics, 15 (7):825–828, July 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00739852.

Seitz:1985:BR

[780] Frederick Seitz. Book review. Foundations of Physics, 15(7):829– 831, July 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00739853.

Cyranski:1985:ITP

[781] John F. Cyranski. Information theory and the problem of molecular structure. Foundations of Physics, 15(8):833–849, August 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00738317.

Carinena:1985:STT

[782] José F. Cariñena and Mariano Santander. Symmetry theory in a twolevel quantum system. Foundations of Physics, 15(8):851–859, August 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00738318.

Russell:1985:PLP

[783] B. R. Russell. Preparation-limited predictions in Einstein–Podolsky– Rosen experiments. Foundations of Physics, 15(8):861–869, August 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00738319.

deBeauregard:1985:SFC

[784] O. Costa de Beauregard. On some frequent but controversial statements concerning the Einstein–Podolsky–Rosen correlations. Foundations of Physics, 15(8):871–887, August 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00738320.

Carmeli:1985:DRR

[785] M. Carmeli. The dynamics of rapidly rotating bodies. Foundations of Physics, 15(8):889–904, August 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00738321.

Toms:1985:BR

[786] David J. Toms and Jayant Narlikar. Book review. Foundations of Physics, 15(8):905–913, August 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00738322.

Inkson:1985:BR

[787] John C. Inkson. Book review. Foundations of Physics, 15(8):915– 917, August 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00738323.

Landsberg:1985:BR

[788] Peter T. Landsberg. Book review. Foundations of Physics, 15(8): 919–920, August 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00738324.

Anonymous:1985:A

[789] Anonymous. Announcements. Foundations of Physics, 15(8):921– 923, August 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00738325.

Uffink:1985:UPU

[790] J. B. M. Uffink and J. Hilgevoord. Uncertainty principle and uncertainty relations. Foundations of Physics, 15(9):925–944, September 1985. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00739034.

Rodrigues:1985:EDP

[791] Waldyr Alves Rodrigues, Jr. and Jayme Tiomno. On experiments to detect possible failures of relativity theory. Foundations of Physics, 15(9): 945–961, September 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00739035.

Rastall:1985:LBT

[792] Peter Rastall. Locality, Bell’s theorem, and quantum mechanics. Foundations of Physics, 15(9):963–972, September 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00739036. See comments [793].

Stapp:1985:CLB

[793] Henry P. Stapp. Comments on “Locality, Bell’s theorem, and quantum mechanics”. Foundations of Physics, 15(9):973–976, September 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00739037. See [792].

Sachs:1985:ELV

[794] Mendel Sachs. On Einstein’s later view of the Twin Paradox. Foundations of Physics, 15(9):977–980, September 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00739038. See comments [1130].

Kilmister:1985:BR

[795] C. W. Kilmister. Book review. Foundations of Physics, 15(9):981– 984, September 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00739039.

Seeger:1985:BR

[796] Raymond J. Seeger. Book review. Foundations of Physics, 15(9):985– 988, September 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00739040.

Tipler:1985:BR

[797] Frank J. Tipler. Book review. Foundations of Physics, 15(9):989– 992, September 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00739041.

Anonymous:1985:Ea

[798] Anonymous. Erratum. Foundations of Physics, 15(9):993, September 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00739042.

Anonymous:1985:Eb

[799] Anonymous. Erratum. Foundations of Physics, 15(9):995, September 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00739043.

Rosen:1985:LGE

[800] Nathan Rosen. Localization of gravitational energy. Foundations of Physics, 15(10):997–1008, October 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00732842.

Gudder:1985:BRH

[801] Stanley Gudder and Thomas Armstrong. Bayes’ rule and hidden variables. Foundations of Physics, 15(10):1009–1017, October 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00732843.

Carmeli:1985:FTTc

[802] M. Carmeli and S. Malin. Field theory on R × S^{3} topology. III: The Dirac equation. Foundations of Physics, 15(10):1019–1029, October 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00732844.

Dewdney:1985:TDN

[803] C. Dewdney, A. Garuccio, Ph. Gueret, A. Kyprianidis, and J. P. Vigier. Time-dependent neutron interferometry: Evidence against wave packet collapse? Foundations of Physics, 15(10):1031–1042, October 1985. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00732845.

Kaufherr:1985:NAE

[804] Tirzah Kaufherr. A new approach to the Einstein–Podolsky–Rosen paradox. Foundations of Physics, 15(10):1043–1051, October 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00732846.

Johnson:1985:BR

[805] G. W. Johnson. Book review. Foundations of Physics, 15(10):1053– 1055, October 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00732847.

Newman:1985:BR

[806] E. T. Newman. Book review. Foundations of Physics, 15(10):1057– 1059, October 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00732848.

Herbert:1985:BR

[807] Nick Herbert. Book review. Foundations of Physics, 15(10):1061– 1063, October 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00732849.

Angelelli:1985:BR

[808] Ignacio Angelelli. Book review. Foundations of Physics, 15(10):1065– 1069, October 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00732850.

Hiscock:1985:BR

[809] William A. Hiscock. Book review. Foundations of Physics, 15(10): 1071–1072, October 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00732851.

Anonymous:1985:Ec

[810] Anonymous. Erratum. Foundations of Physics, 15(10):1073, October 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00732852.

Anonymous:1985:Ed

[811] Anonymous. Erratum. Foundations of Physics, 15(10):1075, October 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00732853.

Anonymous:1985:Ee

[812] Anonymous. Erratum. Foundations of Physics, 15(10):1077, October 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00732854.

Bogolubov:1985:SAP

[813] N. N. Bogolubov and N. N. Bogolubov, Jr. Some approaches to polaron theory. Foundations of Physics, 15(11):1079–1177, November 1985. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00733225.

Bell:1985:OQL

[814] J. L. Bell. Orthospaces and quantum logic. Foundations of Physics, 15(12):1179–1202, December 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00735530.

Banai:1985:QST

[815] M. Banai. Quantization of space–time and the corresponding quantum mechanics. Foundations of Physics, 15(12):1203–1245, December 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00735531.

Videira:1985:GAD

[816] A. L. L. Videira, A. L. Rocha Barros, and N. C. Fernandes. Geometry as an aspect of dynamics. Foundations of Physics, 15(12):1247– 1262, December 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00735532.

Carmeli:1985:STR

[817] M. Carmeli. R × S^{3} special theory of relativity. Foundations of Physics, 15(12):1263–1273, December 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00735533.

Logunov:1986:RTG

[818] A. A. Logunov and M. A. Mestvirishvili. Relativistic theory of gravitation. Foundations of Physics, 16(1):1–26, January 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00735177.

Shirkov:1986:QPF

[819] D. V. Shirkov. Quantum principles in field interactions. Foundations of Physics, 16(1):27–38, January 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00735178.

Chirikov:1986:TCQ

[820] Boris V. Chirikov. Transient chaos in quantum and classical mechanics. Foundations of Physics, 16(1):39–49, January 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00735179.

Casati:1986:ETF

[821] Giulio Casati. Energy transport and the Fourier heat law in classical systems. Foundations of Physics, 16(1):51–61, January 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00735180.

Soloviev:1986:BMN

[822] V. G. Soloviev. Bogolubov’s methods in nuclear theory. Foundations of Physics, 16(1):63–70, January 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00735181.

Bakulev:1986:PTEa

[823] A. P. Bakulev, N. N. Bogolubov, Jr., and A. M. Kurbatov. The principle of thermodynamic equivalence in statistical mechanics: The method of approximating Hamiltonian. Foundations of Physics, 16(1):71, January 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00735182.

Chisolm:1986:BR

[824] J. S. R. Chisolm. Book review. Foundations of Physics, 16(1): 73–75, January 1986. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00735183.

Anonymous:1986:PA

[825] Anonymous. Preliminary announcement. Foundations of Physics, 16 (1):77, January 1986. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00735184.

Thorne:1986:JAW

[826] Kip S. Thorne and Wojciech H. Zurek. John Archibald Wheeler: a few highlights of his contributions to physics. Foundations of Physics, 16 (2):79–89, February 1986. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889373.

Anderson:1986:DTS

[827] Arlen Anderson and Bryce DeWitt. Does the topology of space fluctuate? Foundations of Physics, 16(2):91–105, February 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01889374.

Dicke:1986:OAA

[828] R. H. Dicke. On observing the absence of an atom. Foundations of Physics, 16(2):107–113, February 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01889375.

Barschall:1986:EES

[829] H. H. Barschall and Louis Brown. Early estimates of the strength of the nuclear spin-orbit force. Foundations of Physics, 16(2):115–124, February 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF01889376.

Glimm:1986:MTC

[830] J. Glimm and D. H. Sharp. An S matrix theory for classical nonlinear physics. Foundations of Physics, 16(2):125–141, February 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01889377.

Miller:1986:GCR

[831] Warner Allen Miller. The geometrodynamic content of the Regge equations as illuminated by the boundary of a boundary principle. Foundations of Physics, 16(2):143–169, February 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01889378.

Wilets:1986:MPS

[832] Lawrence Wilets. Momentum projection of solitons including quantum corrections. Foundations of Physics, 16(2):171–185, February 1986. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01889379.

Demianski:1986:SEG

[833] Marek Demianski. Static electromagnetic geon. Foundations of Physics, 16(2):187–190, February 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01889380.

Anonymous:1986:Aa

[834] Anonymous. Announcement. Foundations of Physics, 16(2):191– 192, February 1986. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889381.

Anonymous:1986:PJA

[835] Anonymous. Photos of John Archibald Wheeler. Foundations of Physics, 16(2):i–ii, February 1986. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889372.

Kuchar:1986:CGG

[836] Karel V. Kuchar. Canonical geometrodynamics and general covariance. Foundations of Physics, 16(3):193–208, March 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01889472.

Lukash:1986:OIL

[837] V. N. Lukash and I. D. Novikov. Obvious, incredible …. Foundations of Physics, 16(3):209–225, March 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01889473.

Matzner:1986:PWC

[838] Richard A. Matzner and Anthony Mezzacappa. Professor Wheeler and the crack of doom: Closed cosmologies in the 5-d Kaluza–Klein theory. Foundations of Physics, 16(3):227–248, March 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01889474.

York:1986:BTA

[839] James W. York, Jr. Boundary terms in the action principles of general relativity. Foundations of Physics, 16(3):249–257, March 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01889475.

Ciufolini:1986:TGF

[840] Ignazio Ciufolini. Test of the gravitomagnetic field via laser-ranged satellites. Foundations of Physics, 16(3):259–265, March 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01889476.

Hehl:1986:BIA

[841] Friedrich W. Hehl and J. Dermott McCrea. Bianchi identities and the automatic conservation of energy-momentum and angular momentum in general-relativistic field theories. Foundations of Physics, 16 (3):267–293, March 1986. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889477.

Anonymous:1986:PJWa

[842] Anonymous. Photos of John Wheeler. Foundations of Physics, 16 (3):i–ii, March 1986. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889471.

Cohen:1986:JWD

[843] Leon Cohen and Marlan O. Scully. Joint Wigner distribution for spin-1/2 particles. Foundations of Physics, 16(4):295–310, April 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01882690.

DeWitt-Morette:1986:W

[844] Cécile DeWitt-Morette, Stephen G. Low, Lawrence S. Schulman, and Anwar Y. Shiekh. Wedges I. Foundations of Physics, 16(4):311–349, April 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF01882691.

dEspagnat:1986:QRE

[845] Bernard d’Espagnat. Are the quantum rules exact? The case of the imperfect measurements. Foundations of Physics, 16(4):351–360, April 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF01882692.

Neeman:1986:PQF

[846] Yuval Ne’eman. The problems in quantum foundations in the light of gauge theories. Foundations of Physics, 16(4):361–377, April 1986. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01882693.

Braginsky:1986:HEC

[847] V. B. Braginsky and F. Ya. Khalili. How to evade the confrontation with the uncertainty relations. Foundations of Physics, 16(4):379–382, April 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF01882694.

Unruh:1986:MQN

[848] W. G. Unruh. The measurement of quantum noise reduction in squeezed states. Foundations of Physics, 16(4):383–389, April 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01882695.

Wootters:1986:QMP

[849] William K. Wootters. Quantum mechanics without probability amplitudes. Foundations of Physics, 16(4):391–405, April 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01882696.

Anonymous:1986:ACP

[850] Anonymous. Announcement and call for papers. Foundations of Physics, 16(4):407–408, April 1986. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01882697.

Bekenstein:1986:GSS

[851] Jacob D. Bekenstein. Gravitation and spontaneous symmetry breaking. Foundations of Physics, 16(5):409–422, May 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01882725.

Teller:1986:SGV

[852] Edward Teller. Some generalizations of the virial theorem. Foundations of Physics, 16(5):423–435, May 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01882726;   http://www.osti.
gov/energycitations/product.biblio.jsp?osti_id=5408332;   http:

//www.springerlink.com/content/m543kn8473665585/. Part IV. In-

vited Papers Dedicated to John Archibald Wheeler.

Geheniau:1986:BT

[853] J. Géhéniau and I. Prigogine. The birth of time. Foundations of Physics, 16(5):437–443, May 1986. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01882727.

Alexander:1986:GFS

[854] Matthew Alexander and Peter G. Bergmann. The gravitational field at spatial infinity. Foundations of Physics, 16(5):445–454, May 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01882728.

Hogan:1986:GRR

[855] P. A. Hogan and I. Robinson. Gravitational radiation reaction on the motion of particles in general relativity. Foundations of Physics, 16 (5):455–464, May 1986. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01882729.

Hojman:1986:ELC

[856] Sergio Hojman and L. C. Shepley. Equivalent Lagrangians in classical field theory. Foundations of Physics, 16(5):465–481, May 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01882730.

Kheyfets:1986:BBP

[857] Arkady Kheyfets. The boundary of a boundary principle: a unified approach. Foundations of Physics, 16(5):483–497, May 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01882731.

Wald:1986:BHQ

[858] Robert M. Wald. Black holes and quantum coherence. Foundations of Physics, 16(5):499–506, May 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01882732.

Feynman:1986:QMC

[859] Richard P. Feynman. Quantum mechanical computers. Foundations of Physics, 16(6):507–531, June 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01886518.

Geroch:1986:CPT

[860] Robert Geroch and James B. Hartle. Computability and physical theories. Foundations of Physics, 16(6):533–550, June 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01886519.

Landauer:1986:CPW

[861] Rolf Landauer. Computation and physics: Wheeler’s meaning circuit? Foundations of Physics, 16(6):551–564, June 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01886520.

Deutsch:1986:WNL

[862] David Deutsch. On Wheeler’s notion of “law without law” in physics. Foundations of Physics, 16(6):565–572, June 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01886521.

Peres:1986:EFW

[863] Asher Peres. Existence of “free will” as a problem of physics. Foundations of Physics, 16(6):573–584, June 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01886522.

Bennett:1986:NOC

[864] Charles H. Bennett. On the nature and origin of complexity in discrete, homogeneous, locally-interacting systems. Foundations of Physics, 16 (6):585–592, June 1986. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01886523.

Henneaux:1986:FE

[865] Marc Henneaux and Claudio Teitelboim. p-form electrodynamics. Foundations of Physics, 16(7):593–617, July 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01889624.

Mashhoon:1986:GCQ

[866] Bahram Mashhoon. General covariance and quantum theory. Foundations of Physics, 16(7):619–635, July 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01889625.

Brill:1986:BPI

[867] Dieter R. Brill. Barrier penetration and initial values in Kaluza–Klein theories. Foundations of Physics, 16(7):637–649, July 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01889626.

Isenberg:1986:SU

[868] James Isenberg. Steering the universe. Foundations of Physics, 16 (7):651–665, July 1986. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889627.

Gerlach:1986:HVC

[869] Ulrich H. Gerlach. Heuristic viewpoint concerning the thermal ambience relative to an accelerated frame. Foundations of Physics, 16 (7):667–677, July 1986. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889628.

Treder:1986:JWP

[870] Hans-Jürgen Treder. To John Wheeler: a personal note. Foundations of Physics, 16(7):679, July 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01889629.

Drell:1986:JWP

[871] Sidney Drell. John Wheeler: a personal tribute. Foundations of Physics, 16(7):681–683, July 1986. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889630.

Anonymous:1986:AC

[872] Anonymous. Announcement of conference. Foundations of Physics, 16(7):685, July 1986. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889631.

Anonymous:1986:PJWb

[873] Anonymous. Photos of John Wheeler. Foundations of Physics, 16 (7):i–ii, July 1986. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889623.

Rosen:1986:QPC

[874] Nathan Rosen. Quantum particles and classical particles. Foundations of Physics, 16(8):687–700, August 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00735376.

Holland:1986:RAS

[875] P. R. Holland. Relativistic algebraic spinors and quantum motions in phase space. Foundations of Physics, 16(8):701–719, August 1986. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00735377.

Jahn:1986:QMC

[876] Robert G. Jahn and Brenda J. Dunne. On the quantum mechanics of consciousness, with application to anomalous phenomena. Foundations of Physics, 16(8):721–772, August 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00735378.

Gudder:1986:OH

[877] S. P. Gudder and G. T. Rüttimann. Observables on hypergraphs. Foundations of Physics, 16(8):773–790, August 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00735379.

Carmeli:1986:FTT

[878] M. Carmeli and S. Malin. Field theory on R × S^{3} topology. IV: Electrodynamics of magnetic moments. Foundations of Physics, 16(8): 791–806, August 1986. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00735380.

Redei:1986:NHV

[879] Miklós Rédei. Nonexistence of hidden variables in the algebraic approach. Foundations of Physics, 16(8):807–815, August 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00735381.

Wesley:1986:MMR

[880] J. P. Wesley. Michelson–Morley result, a Voigt–Doppler effect in absolute space–time. Foundations of Physics, 16(8):817–824, August 1986. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00735382.

Stengers:1986:BR

[881] Isabelle Stengers and Max Jammer. Book review. Foundations of Physics, 16(8):825–833, August 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00735383.

Ellis:1986:BR

[882] George Ellis. Book review. Foundations of Physics, 16(8):835– 837, August 1986. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00735384.

Demianski:1986:BR

[883] Marek Demianski, S. Neil Rasband, and Herbert C. Corben. Book review. Foundations of Physics, 16(8):839–845, August 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00735385.

Antman:1986:BR

[884] Stuart S. Antman. Book review. Foundations of Physics, 16(8): 847–849, August 1986. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00735386.

Rudaz:1986:BR

[885] Serge Rudaz. Book review. Foundations of Physics, 16(8):851– 853, August 1986. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00735387.

Anonymous:1986:E

[886] Anonymous. Erratum. Foundations of Physics, 16(8):855, August 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00735388.

Brown:1986:IPQ

[887] Harvey R. Brown. The insolubility proof of the quantum measurement problem. Foundations of Physics, 16(9):857–870, September 1986. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00765334.

Bakulev:1986:PTEb

[888] A. P. Bakulev, N. N. Bogoljubov, Jr., and A. M. Kurbatov. The principle of thermodynamic equivalence in statistical mechanics: The method of approximating Hamiltonian. Foundations of Physics, 16(9): 871–882, September 1986. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00765335.

Frieden:1986:PLF

[889] B. Roy Frieden. A probability law for the fundamental constants. Foundations of Physics, 16(9):883–903, September 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00765336.

Engelke:1986:ELIa

[890] C. E. Engelke and C. W. Engelke. The effect of localization on interference. I. Calculated intensities for a feasible optical experiment. Foundations of Physics, 16(9):905–916, September 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00765337.

Engelke:1986:ELIb

[891] Charles E. Engelke. The effect of localization on interference. II. Bearing on locality violation and the interpretation of quantum mechanics. Foundations of Physics, 16(9):917–921, September 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00765338.

Zanchini:1986:DEP

[892] Enzo Zanchini. On the definition of extensive property energy by the first postulate of thermodynamics. Foundations of Physics, 16(9):923– 935, September 1986. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00765339.

deBeauregard:1986:BDF

[893] O. Costa de Beauregard. Bohr’s discussion of the fourth uncertainty relation revisited. Foundations of Physics, 16(9):937–939, September 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00765340.

dEspagnat:1986:BR

[894] Bernard d’Espagnat and Lochlainn O’Raifeartaigh. Book review. Foundations of Physics, 16(9):941–949, September 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00765341.

Heller:1986:BR

[895] Michael Heller and W. T. Morris. Book review. Foundations of Physics, 16(9):951–955, September 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00765342.

Welford:1986:BR

[896] Walter Welford, Geraint Rosser, S. Fujita, and Donald Richards. Book review. Foundations of Physics, 16(9):957–965, September 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00765343.

Lounesto:1986:RC

[897] Pertti Lounesto. Report on conference. Foundations of Physics, 16(9): 967–971, September 1986. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00765344.

deMuynck:1986:RBE

[898] Willem M. de Muynck. On the relation between the Einstein–Podolsky– Rosen paradox and the problem of nonlocality in quantum mechanics. Foundations of Physics, 16(10):973–1002, October 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01892937.

Vargas:1986:KGA

[899] José G. Vargas. Kinematical and gravitational analysis of the rocketborne clock experiment by Vessot and Levine using the revised Robertson’s test theory of special relativity. Foundations of Physics, 16(10): 1003–1020, October 1986. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01892938.

Aron:1986:SAHa

[900] J. C. Aron. A stochastic approach to the hadron spectrum. I. Foundations of Physics, 16(10):1021–1060, October 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01892939.

Anonymous:1986:EPP

[901] Anonymous. Errata to previous papers. Foundations of Physics, 16 (10):1060, October 1986. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01892940.

Schweizer:1986:PDG

[902] B. Schweizer and A. Sklar. Probability distributions with given margins: Note on a paper by Finch and Groblicki. Foundations of Physics, 16(10): 1061–1064, October 1986. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01892941. See [648].

Norton:1986:BR

[903] John Norton. Book review. Foundations of Physics, 16(10):1065– 1067, October 1986. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01892942.

Pyenson:1986:BR

[904] Lewis Pyenson. Book review. Foundations of Physics, 16(10):1069– 1071, October 1986. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01892943.

Williams:1986:BR

[905] L. Pearce Williams. Book review. Foundations of Physics, 16(10): 1073–1075, October 1986. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01892944.

Saleh:1986:BR

[906] Bahaa E. A. Saleh, Donna K. Bandy, Larry Smalley, and David Hawkins. Book review. Foundations of Physics, 16(10):1077–1085, October 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01892945.

Temmer:1986:BR

[907] Georges M. Temmer. Book review. Foundations of Physics, 16(10): 1087–1088, October 1986. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01892946.

Vargas:1986:RRTa

[908] José G. Vargas and Douglas G. Torr. Revised Robertson’s test theory of special relativity: Space–time structure and dynamics. Foundations of Physics, 16(11):1089–1126, November 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00761283.

Just:1986:CQC

[909] K. Just and L. S. The. Canonical quantization without conjugate momenta. Foundations of Physics, 16(11):1127–1141, November 1986. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00761284.

Seipp:1986:BTE

[910] H. P. Seipp. Bell’s theorem and an explicit stochastic local hiddenvariable model. Foundations of Physics, 16(11):1143–1152, November 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00761285.

deBeauregard:1986:CEU

[911] O. Costa de Beauregard. On Carmeli’s exotic use of the Lorentz transformation and on the velocity composition approach to special relativity. Foundations of Physics, 16(11):1153–1157, November 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00761286.

Aron:1986:SAHb

[912] J. C. Aron. A stochastic approach to the hadron spectrum. II. Foundations of Physics, 16(11):1159–1210, November 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00761287.

Humphreys:1986:BR

[913] Paul Humphreys and J. S. R. Chisholm. Book review. Foundations of Physics, 16(11):1211–1219, November 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00761288.

Hiscock:1986:BR

[914] William A. Hiscock. Book review. Foundations of Physics, 16(11): 1221–1223, November 1986. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00761289.

Berry:1986:BR

[915] Michael Berry and O. W. Greenberg. Book review. Foundations of Physics, 16(11):1225–1228, November 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00761290.

Anonymous:1986:Ab

[916] Anonymous. Announcement. Foundations of Physics, 16(11):1229, November 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00761291.

Vargas:1986:RRTb

[917] José G. Vargas. Revised Robertson’s test theory of special relativity: Supergroups and superspace. Foundations of Physics, 16(12):1231– 1261, December 1986. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00732119.

Feuchtwang:1986:GGI

[918] T. E. Feuchtwang, E. Kazes, and P. H. Cutler. Generalized gauge independence and the physical limitations on the von Neumann measurement postulate. Foundations of Physics, 16(12):1263–1284, December 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00732120.

Svetlichny:1986:QSM

[919] George Svetlichny. Quantum supports and modal logic. Foundations of Physics, 16(12):1285–1295, December 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00732121.

Malin:1986:QSP

[920] Shimon Malin. Quantum states and potentialities of quantum systems. Foundations of Physics, 16(12):1297–1305, December 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00732122.

Kannenberg:1986:IIC

[921] L. Kannenberg. Independence of the isotropy convention for the oneway light speed in relativity. Foundations of Physics, 16(12):1307– 1313, December 1986. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00732123.

Aron:1986:SAHc

[922] J. C. Aron. A stochastic approach to the hadron spectrum. III. Foundations of Physics, 16(12):1315–1328, December 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00732124.

Bennett:1986:BR

[923] M. K. Bennett. Book review. Foundations of Physics, 16(12):1329– 1331, December 1986. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00732125.

Pierce:1986:BR

[924] Allan D. Pierce. Book review. Foundations of Physics, 16(12):1333– 1335, December 1986. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00732126.

Chen:1986:BR

[925] Bang-Yen Chen. Book review. Foundations of Physics, 16(12):1337– 1338, December 1986. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00732127.

Kummer:1987:CAF

[926] Hans Kummer. A constructive approach to the foundations of quantum mechanics. Foundations of Physics, 17(1):1–62, January 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00751152.

Rosen:1987:VSS

[927] Nathan Rosen and Gerald E. Tauber. Vector-spinor space and field equations. Foundations of Physics, 17(1):63–99, January 1987. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00751153.

Cushing:1987:BR

[928] James T. Cushing and John L. Friedman. Book review. Foundations of Physics, 17(1):101–109, January 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00751154.

Barut:1987:BR

[929] A. O. Barut. Book review. Foundations of Physics, 17(1):111– 113, January 1987. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00751155.

Gordon:1987:BR

[930] William E. Gordon. Book review. Foundations of Physics, 17(1): 115–116, January 1987. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00751156.

Anonymous:1987:A

[931] Anonymous. Announcement. Foundations of Physics, 17(1):117, January 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00751157.

Wignall:1987:NSE

[932] J. W. G. Wignall. The nonrelativistic Schrödinger equation in “quasiclassical” theory. Foundations of Physics, 17(2):123–147, February 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00733204.

Smith:1987:RFE

[933] Gerrit J. Smith and Robert Weingard. A relativistic formulation of the Einstein–Podolsky–Rosen paradox. Foundations of Physics, 17(2): 149–171, February 1987. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00733205.

Steeb:1987:QCE

[934] W.-H. Steeb, J. A. Louw, and A. Kunick. Quantum chaos of an excitonphonon system. Foundations of Physics, 17(2):173–181, February 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00733206.

Schachter:1987:MIR

[935] François Schächter. A Markov-informational representation of evolutions. Foundations of Physics, 17(2):183–192, February 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00733207.

Carmeli:1987:FTTa

[936] M. Carmeli and S. Malin. Field theory on R×S^{3} topology. V: SU_{2} gauge theory. Foundations of Physics, 17(2):193–200, February 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00733208.

Ktorides:1987:SCB

[937] C. N. Ktorides and L. C. Papaloucas. A special construction of Berezin’s L-kernel. Foundations of Physics, 17(2):201–207, February 1987. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00733209.

Brooke:1987:BR

[938] James A. Brooke. Book review. Foundations of Physics, 17(2):209– 213, February 1987. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00733210.

Rieckers:1987:BR

[939] Alfred Rieckers and John C. Inkson. Book review. Foundations of Physics, 17(2):215–220, February 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00733211.

Anonymous:1987:PMN

[940] Anonymous. Programme of meetings NATO international scientific exchange programme mathematical and physical sciences. Foundations of Physics, 17(2):I–VI, February 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00733212.

Lavenda:1987:CVD

[941] B. H. Lavenda. Classical variational derivation and physical interpretation of Dirac’s equation. Foundations of Physics, 17(3):221–237, March 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00733091.

Recami:1987:TKC

[942] Erasmo Recami. Tachyon kinematics and causality: a systematic thorough analysis of the tachyon causal paradoxes. Foundations of Physics, 17(3):239–296, March 1987. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00733092.

Rietdijk:1987:REM

[943] C. W. Rietdijk. Retroactive effects from measurements. Foundations of Physics, 17(3):297–311, March 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00733093.

Phipps:1987:BR

[944] Thomas E. Phipps, Jr. Book review. Foundations of Physics, 17(3): 313–319, March 1987. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00733094.

Halpern:1987:BR

[945] Leopold Halpern. Book review. Foundations of Physics, 17(3): 321–327, March 1987. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00733095.

Ferrari:1987:CEH

[946] L. Ferrari. The continuity equation and the Hamiltonian formalism in quantum mechanics. Foundations of Physics, 17(4):329–343, April 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00733372.

Holland:1987:GDB

[947] P. R. Holland. Geometry of dislocated de Broglie waves. Foundations of Physics, 17(4):345–363, April 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00733373.

Beretta:1987:QTN

[948] Gian Paolo Beretta. Quantum thermodynamics of nonequilibrium. Onsager reciprocity and dispersion-dissipation relations. Foundations of Physics, 17(4):365–381, April 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00733374.

Schachter:1987:PIO

[949] François Schächter. The probabilistic-informational opacity functional, Jaynes’s principle, and distances to equilibrium of an evolving system.

Foundations of Physics, 17(4):383–396, April 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00733375.

Bertrand:1987:TAW

[950] J. Bertrand and P. Bertrand. A tomographic approach to Wigner’s function. Foundations of Physics, 17(4):397–405, April 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00733376.

Carmeli:1987:FTTb

[951] M. Carmeli and S. Malin. Field theory on R × S^{3} topology. VI: Gravitation. Foundations of Physics, 17(4):407–417, April 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00733377.

Adomian:1987:NAE

[952] G. Adomian. A new approach to the Efinger model for a nonlinear quantum theory for gravitating particles. Foundations of Physics, 17 (4):419–423, April 1987. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00733378.

Achinstein:1987:BR

[953] Peter Achinstein. Book review. Foundations of Physics, 17(4): 425–433, April 1987. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00733379.

Smalley:1987:BR

[954] Larry L. Smalley. Book review. Foundations of Physics, 17(4): 435–439, April 1987. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00733380.

Inagaki:1987:BR

[955] Shogo Inagaki. Book review. Foundations of Physics, 17(4):441– 443, April 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00733381.

Capri:1987:BR

[956] Anton Z. Capri, Frank Weinhold, Mahinder S. Uberoi, and A. P. Balachandran. Book review. Foundations of Physics, 17(4):445–451, April 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00733382.

Volkenstein:1987:BR

[957] Mikhail Volkenstein. Book review. Foundations of Physics, 17(4): 453–457, April 1987. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00733383.

Nicolis:1987:IPH

[958] G. Nicolis, L. E. Reichl, and A. van der Merwe. For Ilya Prigogine on his seventieth birthday. Foundations of Physics, 17(5):459–460, May 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01559697.

Mehra:1987:NBD

[959] Jagdish Mehra. Niels Bohr discussions with Albert Einstein, Werner Heisenberg, and Erwin Schrödinger — the origins of the Principles of Uncertainty and Complementarity. Foundations of Physics, 17(5): 461–506, May 1987. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01559698.

dEspagnat:1987:ERE

[960] B. d’Espagnat. Empirical reality, empirical causality, and the measurement problem. Foundations of Physics, 17(5):507–529, May 1987. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01559699.

Holland:1987:TCP

[961] P. R. Holland, A. Kyprianidis, and J. P. Vigier. Trajectories and causal phase-space approach to relativistic quantum mechanics. Foundations of Physics, 17(5):531–548, May 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01559700.

Barut:1987:IOS

[962] A. O. Barut. Irreversibility, organization, and self-organization in quantum electrodynamics. Foundations of Physics, 17(6):549–559, June 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01882786.

Cohen:1987:CEV

[963] Leon Cohen and Chongmoon Lee. Conditional expectation values in quantum mechanics. Foundations of Physics, 17(6):561–574, June 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01882787.

Scully:1987:TSG

[964] Marlan O. Scully, Willis E. Lamb, Jr., and Asim Barut. On the theory of the Stern–Gerlach apparatus. Foundations of Physics, 17 (6):575–583, June 1987. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01882788.

Geheniau:1987:UBF

[965] Jules Géheniau, Edgard Gunzig, and Isabelle Stengers. The universe: a birth far from equilibrium? Foundations of Physics, 17(6):585–601, June 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF01882789.

Brout:1987:CTO

[966] R. Brout. On the concept of time and the origin of the cosmological temperature. Foundations of Physics, 17(6):603–619, June 1987. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01882790.

Englert:1987:DST

[967] F. Englert. From disorder to space–time geometry. Foundations of Physics, 17(6):621–635, June 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01882791.

Henneaux:1987:FTS

[968] Marc Henneaux. Field theory of strings in nonconformal gauges. Foundations of Physics, 17(6):637–652, June 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01882792.

Glansdorff:1987:IMP

[969] P. Glansdorff. Irreversibility in macroscopic physics: From Carnot cycle to dissipative structures. Foundations of Physics, 17(7):653–666, July 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF01889539.

Bohm:1987:IOP

[970] David Bohm. The implicate order and Prigogine’s notions of irreversibility. Foundations of Physics, 17(7):667–677, July 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01889540.

Rossler:1987:EDS

[971] Otto E. Rössler. Explicit dissipative structures. Foundations of Physics, 17(7):679–688, July 1987. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889541.

Reichl:1987:SQK

[972] L. E. Reichl and W. A. Lin. The search for a quantum KAM theorem. Foundations of Physics, 17(7):689–697, July 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01889542.

Tomita:1987:CPR

[973] Kazuhisa Tomita. Conjugate pair of representations in chaos and quantum mechanics. Foundations of Physics, 17(7):699–711, July 1987. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01889543.

Kondepudi:1987:MAI

[974] D. K. Kondepudi. Microscopic aspects implied by the Second Law. Foundations of Physics, 17(7):713–722, July 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01889544.

Allen:1987:EDE

[975] P. M. Allen and J. M. McGlade. Evolutionary drive: The effect of microscopic diversity, error making, and noise. Foundations of Physics, 17(7):723–738, July 1987. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889545.

Anonymous:1987:PIP

[976] Anonymous. Photo of Ilya Prigogine. Foundations of Physics, 17(7):ii, July 1987. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889538.

Selleri:1987:CPP

[977] F. Selleri. Coherence properties of photon amplifiers. Foundations of Physics, 17(8):739–757, August 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00733264.

Cufaro-Petroni:1987:EPR

[978] N. Cufaro-Petroni, C. Dewdney, P. R. Holland, A. Kyprianidis, and J. P. Vigier. Einstein–Podolsky–Rosen constraints on quantum action at a distance: The Sutherland paradox. Foundations of Physics, 17 (8):759–773, August 1987. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00733265.

deBeauregard:1987:ZCE

[979] O. Costa de Beauregard. On the zigzagging causality EPR model: answer to Vigier and coworkers and to Sutherland. Foundations of Physics, 17 (8):775–785, August 1987. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00733266.

Hoekzema:1987:LRC

[980] Dick Hoekzema. Locality and realism in contextual theories. Foundations of Physics, 17(8):787–797, August 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00733267.

Leacock:1987:PPL

[981] Robert A. Leacock. The physical properties of linear and action-angle coordinates in classical and quantum mechanics. Foundations of Physics, 17(8):799–807, August 1987. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00733268.

Burgos:1987:QMC

[982] M. E. Burgos. Quantum mechanics is compatible with realism. Foundations of Physics, 17(8):809–812, August 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00733269.

Mugur-Schachter:1987:LRW

[983] Mioara Mugur-Schächter. Locality, reflection, and wave-particle duality. Foundations of Physics, 17(8):813–857, August 1987. CODEN FNDPA4.

ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00733270.

Poole:1987:BR

[984] Charles P. Poole, Jr. and Gilles Labonté. Book review. Foundations of Physics, 17(8):859–867, August 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00733271.

Squires:1987:BR

[985] E. J. Squires and Francesco Guerra. Book review. Foundations of Physics, 17(8):869–873, August 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00733272.

Leggett:1987:BR

[986] Anthony J. Leggett, Peter Graneau, and Mitchell C. Begelman. Book reviews. Foundations of Physics, 17(8):875–889, August 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00733273.

Mittelstaedt:1987:UPW

[987] P. Mittelstaedt, A. Prieur, and R. Schieder. Unsharp particle-wave duality in a photon split-beam experiment. Foundations of Physics, 17(9): 891–903, September 1987. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00734319.

Busch:1987:SRJ

[988] Paul Busch. Some realizable joint measurements of complementary observables. Foundations of Physics, 17(9):905–937, September 1987. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00734320.

Atmanspacher:1987:FLB

[989] H. Atmanspacher and H. Scheingraber. A fundamental link between system theory and statistical mechanics. Foundations of Physics, 17(9): 939–963, September 1987. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00734321.

Anonymous:1987:E

[990] Anonymous. Erratum. Foundations of Physics, 17(9):965, September 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00734322.

Lochak:1987:LB

[991] Georges Lochak. Louis de Broglie (1892–1987). Foundations of Physics, 17(10):967–970, October 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00938006.

Croca:1987:NIC

[992] J. R. Croca. Neutron interferometry can prove (or refute) the existence of de Broglie’s waves. Foundations of Physics, 17(10):971–980, October 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00938007.

Liboff:1987:QEM

[993] Richard L. Liboff. Quantum equations of motion and the Liouville equation. Foundations of Physics, 17(10):981–991, October 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00938008.

deWet:1987:NSG

[994] J. A. de Wet. Nuclear structure on a Grassmann manifold. Foundations of Physics, 17(10):993–1018, October 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00938009.

Klay:1987:QLP

[995] Matthias P. Kläy. Quantum logic properties of hypergraphs. Foundations of Physics, 17(10):1019–1036, October 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00938010.

Schweber:1987:BR

[996] Silvan S. Schweber. Book review. Foundations of Physics, 17(10): 1037–1039, October 1987. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00938011.

Shimony:1987:BR

[997] Abner Shimony. Book review. Foundations of Physics, 17(10):1041– 1043, October 1987. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00938012.

Peters:1987:BR

[998] Philip C. Peters. Book review. Foundations of Physics, 17(10):1045– 1047, October 1987. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00938013.

Dudley:1987:BR

[999] Donald G. Dudley. Book review. Foundations of Physics, 17(10): 1049–1050, October 1987. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00938014.

Mehra:1987:ESRa

[1000] Jagdish Mehra. Erwin Schrödinger and the rise of wave mechanics. I. Schrödinger’s scientific work before the creation of wave mechanics. Foundations of Physics, 17(11):1051–1112, November 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01889803.

Halpern:1987:ESV

[1001] Leopold Halpern. Erwin Schrödinger’s views on gravitational physics during his last years at the University of Vienna and some research ensuing from it. Foundations of Physics, 17(11):1113–1130, November 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01889804.

Prigogine:1987:BR

[1002] Ilya Prigogine. Book review. Foundations of Physics, 17(11):1131– 1136, November 1987. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889805.

Kilmister:1987:CR

[1003] C. W. Kilmister. Conference report. Foundations of Physics, 17(11): 1137–1140, November 1987. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889806.

Anonymous:1987:PESa

[1004] Anonymous. Photo of Erwin Schrödinger. Foundations of Physics, 17 (11):ii, November 1987. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889802.

Mehra:1987:ESRb

[1005] Jagdish Mehra. Erwin Schrödinger and the rise of wave mechanics. II. The creation of wave mechanics. Foundations of Physics, 17(12): 1141–1188, December 1987. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889592.

Lochak:1987:CDB

[1006] Georges Lochak. Convergence and divergence between the ideas of de Broglie and Schrödinger in wave mechanics. Foundations of Physics, 17(12):1189–1203, December 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01889593.

Rohrlich:1987:SIQ

[1007] Fritz Rohrlich. Schrödinger and the interpretation of quantum mechanics. Foundations of Physics, 17(12):1205–1220, December 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01889594.

Anonymous:1987:PESb

[1008] Anonymous. Photo of Erwin Schrödinger. Foundations of Physics, 17 (12):ii, December 1987. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889591.

Ghirardi:1988:PES

[1009] G. C. Ghirardi, A. Rimini, and T. Weber. The puzzling entanglement of Schrödinger’s wave function. Foundations of Physics, 18(1): 1–27, January 1988. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01882871.

Namiki:1988:MHS

[1010] Mikio Namiki. Many-Hilbert-spaces theory of quantum measurements. Foundations of Physics, 18(1):29–55, January 1988. CODEN FNDPA4.

ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01882872.

Peres:1988:SIC

[1011] Asher Peres. Schrödinger’s immortal cat. Foundations of Physics, 18 (1):57–76, January 1988. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01882873.

Treder:1988:IIS

[1012] Hans-Jürgen Treder and Horst-Heino von Borzeszkowski. Interference and interaction in Schrödinger’s wave mechanics. Foundations of Physics, 18(1):77–93, January 1988. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01882874.

Barut:1988:CRQ

[1013] A. O. Barut. Combining relativity and quantum mechanics: Schrödinger’s interpretation of ψ. Foundations of Physics, 18(1):95–105, January 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01882875.

Mehra:1988:ESR

[1014] Jagdish Mehra. Erwin Schrödinger and the rise of wave mechanics. III. Early response and applications. Foundations of Physics, 18(2):107–184, February 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1007/
BF01882930;   http://www.springerlink.com/content/p144515610k37j02/

. Part IV. Invited Papers Commemorating the Centenary of the Birth of Erwin Schrödinger, 12 August 1887.

Marshall:1988:SOR

[1015] Trevor Marshall and Emilio Santos. Stochastic optics: a reaffirmation of the wave nature of light. Foundations of Physics, 18(2):185–223, February 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1007/
BF01882931;   http://www.springerlink.com/content/r342562317734r75/

. Part IV. Invited Papers Commemorating the Centenary of the Birth of Erwin Schrödinger, 12 August 1887.

Park:1988:TAS

[1016] James L. Park. Thermodynamic aspects of Schrödinger’s probability relations. Foundations of Physics, 18(2):225–244, February 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://

link.springer.com/article/10.1007/BF01882932;   http://www.
springerlink.com/content/v2666233mgk4t133/. Part IV. Invited Pa-

pers Commemorating the Centenary of the Birth of Erwin Schrödinger, 12 August 1887.

Neeman:1988:SGG

[1017] Yuval Ne’eman. The spectrum-generating groups program and the string. Foundations of Physics, 18(3):245–275, March 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01891492.

Mukunda:1988:FBR

[1018] N. Mukunda, E. C. G. Sudarshan, and R. Simon. Families of Bose rays in quantum optics. Foundations of Physics, 18(3):277–306, March 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01891493.

ORaifeartaigh:1988:WPT

[1019] L. O’Raifeartaigh and Andreas Wipf. WKB properties of the timedependent Schrödinger system. Foundations of Physics, 18(3):307– 329, March 1988. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01891494.

Bitsakis:1988:QSD

[1020] Eftichios Bitsakis. Quantum statistical determinism. Foundations of Physics, 18(3):331–355, March 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01891495.

Melgar:1988:PES

[1021] Miguel Ferrero Melgar. The philosophy of Erwin Schrödinger: a diachronic view of Schrödinger’s thoughts. Foundations of Physics, 18 (3):357–371, March 1988. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01891496.

Bergmann:1988:QMV

[1022] Otto Bergmann. A quantum mechanical version of the paper by E. Schrödinger “Über die Umkehrung der Naturgesetze”. Foundations of Physics, 18(3):373–378, March 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01891497.

Kuhn:1988:ADW

[1023] Wilfried Kuhn. Analysis of the development of wave mechanics: aspects from the history of physics and the philosophy of science. Foundations of Physics, 18(3):379–399, March 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01891498.

Anonymous:1988:PES

[1024] Anonymous. Photo of Erwin Schrödinger. Foundations of Physics, 18 (3):iii–iv, March 1988. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01891491.

Buonomano:1988:SPI

[1025] V. Buonomano and A. F. Prado de Andrade. Stochastic processes for indirectly interacting particles and stochastic quantum mechanics. Foundations of Physics, 18(4):401–426, April 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00732547.

Stapp:1988:QN

[1026] Henry P. Stapp. Quantum nonlocality. Foundations of Physics, 18 (4):427–448, April 1988. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00732548.

Landau:1988:ETP

[1027] Lawrence J. Landau. Empirical two-point correlation functions. Foundations of Physics, 18(4):449–460, April 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00732549.

Mugur-Schachter:1988:WPL

[1028] Mioara Mugur-Schächter. What is the problem in the locality problem? Foundations of Physics, 18(4):461–470, April 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00732550.

Jones:1988:BR

[1029] Roger Jones, Myron Tribus, Don N. Page, and Bernard H. Lavenda. Book review. Foundations of Physics, 18(4):471–487, April 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00732551.

Greenberg:1988:BR

[1030] O. W. Greenberg. Book review. Foundations of Physics, 18(4): 489–490, April 1988. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00732552.

Bussey:1988:FQM

[1031] P. J. Bussey. The foundations of quantum mechanics. Foundations of Physics, 18(5):491–528, May 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00732743.

Dietz:1988:NES

[1032] Werner Dietz. New exact solutions of Einstein’s field equations: Gravitational force can also be repulsive! Foundations of Physics, 18 (5):529–547, May 1988. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00732744.

Band:1988:CITa

[1033] William Band. Can information be transferred faster than light? I. A gedanken device for generating electromagnetic wave packets with superoptic group velocity. Foundations of Physics, 18(5):549–562, May 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00732745.

Davies:1988:NCT

[1034] John B. Davies. New curvature-torsion relations through decomposition of the Bianchi identities. Foundations of Physics, 18(5):563–569, May 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00732746.

Bishop:1988:STT

[1035] N. T. Bishop. Is superluminal travel a theoretical possibility? II. Foundations of Physics, 18(5):571–574, May 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00732747.

Evrard:1988:MNR

[1036] Daniel Evrard and Francçois Thieffine. Microphysical nonseparability and relativistic signals. Foundations of Physics, 18(5):575–583, May 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00732748.

Franklin:1988:BR

[1037] Allan Franklin. Book review. Foundations of Physics, 18(5):585– 588, May 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00732749.

Bertolotti:1988:BR

[1038] Mario Bertolotti. Book review. Foundations of Physics, 18(5): 589–590, May 1988. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00732750.

Wignall:1988:AAQ

[1039] J. W. G. Wignall. An alternative approach to quantum phenomena. Foundations of Physics, 18(6):591–624, June 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00734564.

Band:1988:CITb

[1040] William Band. Can information be transferred faster than light? II. The relativistic Doppler effect on electromagnetic wave packets with suboptic and superoptic group velocities. Foundations of Physics, 18 (6):625–638, June 1988. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00734565.

Bach:1988:CIP

[1041] Alexander Bach. The concept of indistinguishable particles in classical and quantum physics. Foundations of Physics, 18(6):639–649, June 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00734566.

Chang:1988:CEP

[1042] T. Chang. On the choice of evolutional parameter within a framework of four-dimensional symmetry. Foundations of Physics, 18(6):651–658, June 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00734567.

Narlikar:1988:SSS

[1043] J. V. Narlikar and T. Padmanabhan. The Schwarzschild solution: Some conceptual difficulties. Foundations of Physics, 18(6):659–668, June 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00734568.

vonBorzeszkowski:1988:QMP

[1044] Horst-Heino von Borzeszkowski and Renate Wahsner. Quantum mechanics and the physical reality concept. Foundations of Physics, 18 (6):669–681, June 1988. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00734569.

Leschke:1988:BR

[1045] Hajo Leschke. Book review. Foundations of Physics, 18(6):683–685, June 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00734570.

Norton:1988:BR

[1046] John Norton. Book review. Foundations of Physics, 18(6):687–689, June 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00734571.

Jammer:1988:DBH

[1047] Max Jammer. David Bohm and his work — on the occasion of his seventieth birthday. Foundations of Physics, 18(7):691–699, July 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00734150.

Bub:1988:HSM

[1048] Jeffrey Bub. How to solve the measurement problem of quantum mechanics. Foundations of Physics, 18(7):701–722, July 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00734151.

Zeh:1988:MBV

[1049] H.-D. Zeh. Measurement in Bohm’s versus Everett’s quantum theory. Foundations of Physics, 18(7):723–730, July 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00734152.

Semon:1988:ABE

[1050] Mark D. Semon and John R. Taylor. The Aharonov–Bohm effect: Still a thought-provoking experiment. Foundations of Physics, 18(7): 731–740, July 1988. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00734153.

Holland:1988:QPS

[1051] P. R. Holland and J. P. Vigier. The quantum potential and signalling in the Einstein–Podolsky–Rosen experiment. Foundations of Physics, 18(7):741–750, July 1988. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00734154.

Gudder:1988:QGD

[1052] Stanley P. Gudder. Quantum graphic dynamics. Foundations of Physics, 18(7):751–776, July 1988. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00734155.

Frescura:1988:PSG

[1053] F. A. M. Frescura. Projective spinor geometry and prespace. Foundations of Physics, 18(8):777–808, August 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01889310.

Chew:1988:TSQ

[1054] G. F. Chew and H. P. Stapp. Three-space from quantum mechanics. Foundations of Physics, 18(8):809–831, August 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01889311.

Stapp:1988:SFQ

[1055] Henry P. Stapp. Spacetime and future quantum theory. Foundations of Physics, 18(8):833–849, August 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01889312.

Tucker:1988:CCV

[1056] R. W. Tucker. Closed currents, vector fields, and phenomena. Foundations of Physics, 18(8):851–864, August 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01889313.

Anonymous:1988:A

[1057] Anonymous. Announcement. Foundations of Physics, 18(8):865, August 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF01889314.

Anonymous:1988:PDBa

[1058] Anonymous. Photo of David Bohm. Foundations of Physics, 18 (8):ii, August 1988. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889309.

Dewdney:1988:NCT

[1059] C. Dewdney. Nonlocally correlated trajectories in two-particle quantum mechanics. Foundations of Physics, 18(9):867–886, September 1988. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01855940.

Wan:1988:LON

[1060] K. Kong Wan. Local observables, nonlocality, and asymptotically separable quantum mechanics. Foundations of Physics, 18(9):887– 911, September 1988. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01855941.

deBeauregard:1988:ZCM

[1061] O. Costa de Beauregard. On the zigzagging causility model of EPR correlations and on the interpretation of quantum mechanics. Foundations of Physics, 18(9):913–938, September 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01855942.

Leggett:1988:EAQ

[1062] A. J. Leggett. Experimental approaches to the quantum measurement paradox. Foundations of Physics, 18(9):939–952, September 1988. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01855943.

Schleich:1988:APS

[1063] W. Schleich, H. Walther, and J. A. Wheeler. Area in phase space as determiner of transition probability: Bohr–Sommerfeld bands, Wigner ripples, and Fresnel zones. Foundations of Physics, 18(10):953–968, October 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF01909932.

Landsberg:1988:WQM

[1064] P. T. Landsberg. Why quantum mechanics? Foundations of Physics, 18 (10):969–982, October 1988. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01909933.

Cohen:1988:RPQ

[1065] Leon Cohen. Rules of probability in quantum mechanics. Foundations of Physics, 18(10):983–998, October 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01909934.

Barut:1988:JPN

[1066] A. O. Barut, M. Bozić, and Z. Marić. Joint probabilities of noncommuting operators and incompleteness of quantum mechanics. Foundations of Physics, 18(10):999–1012, October 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01909935.

deObaldia:1988:ABA

[1067] Elida de Obaldia, Abner Shimony, and Frederick Wittel. Amplification of Belinfante’s argument for the nonexistence of dispersion-free states. Foundations of Physics, 18(10):1013–1021, October 1988. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01909936.

Wang:1988:QMW

[1068] Lipo Wang and R. F. O’Connell. Quantum mechanics without wave functions. Foundations of Physics, 18(10):1023–1033, October 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01909937.

Malin:1988:WAB

[1069] Shimon Malin. A Whiteheadian approach to Bell’s correlations. Foundations of Physics, 18(10):1035–1044, October 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01909938.

Englert:1988:SCL

[1070] Berthold-Georg Englert, Julian Schwinger, and Marlan O. Scully. Is spin coherence like Humpty–Dumpty? I. Simplified treatment. Foundations of Physics, 18(10):1045–1056, October 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01909939.

Anonymous:1988:PDBb

[1071] Anonymous. Photo of David Bohm. Foundations of Physics, 18 (10):i, October 1988. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01909931.

Cufaro-Petroni:1988:SOW

[1072] Nicola Cufaro-Petroni, Philippe Gueret, and Jean-Pierre Vigier. Secondorder wave equation for spin-1/2 fields: 8-spinors and canonical formulation. Foundations of Physics, 18(11):1057–1075, November 1988. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01892948.

Kyprianidis:1988:TTP

[1073] A. Kyprianidis. Trajectories of two-particle states for the harmonic oscillator. Foundations of Physics, 18(11):1077–1091, November 1988. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01892949.

Rosen:1988:NIP

[1074] Steven M. Rosen. A neo-intuitive proposal for Kaluza–Klein unification. Foundations of Physics, 18(11):1093–1139, November 1988. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01892950.

Anonymous:1988:PDBc

[1075] Anonymous. Photo of David Bohm. Foundations of Physics, 18 (11):i, November 1988. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01892947.

Selleri:1988:LDD

[1076] F. Selleri and A. Zeilinger. Local deterministic description of Einstein– Podolsky–Rosen experiments. Foundations of Physics, 18(12):1141– 1158, December 1988. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889429.

Horwitz:1988:TAT

[1077] L. P. Horwitz, R. I. Arshansky, and A. C. Elitzur. On the two aspects of time: The distinction and its implications. Foundations of Physics, 18(12):1159–1193, December 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01889430.

Josephson:1988:LUQ

[1078] Brian D. Josephson. Limits to the universality of quantum mechanics. Foundations of Physics, 18(12):1195–1204, December 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01889431.

Cramer:1988:VRA

[1079] John G. Cramer. Velocity reversal and the arrows of time. Foundations of Physics, 18(12):1205–1212, December 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01889432.

Peat:1988:TSO

[1080] F. David Peat. Time, structure, and objectivity in quantum theory. Foundations of Physics, 18(12):1213–1231, December 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01889433.

Ford:1988:RLS

[1081] Alan Ford and F. David Peat. The role of language in science. Foundations of Physics, 18(12):1233–1242, December 1988. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01889434.

Chang:1988:BR

[1082] Shau-Jin Chang. Book review. Foundations of Physics, 18(12):1243– 1245, December 1988. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889435.

Anonymous:1988:IA

[1083] Anonymous. Important announcement. Foundations of Physics, 18 (12):i, December 1988. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889428.

Anonymous:1989:IA

[1084] Anonymous. Important announcement. Foundations of Physics, 19 (1):1, January 1989. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00737763.

Royer:1989:MQS

[1085] Antoine Royer. Measurement of quantum states and the Wigner function. Foundations of Physics, 19(1):3–32, January 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00737764.

Israelit:1989:GCB

[1086] Mark Israelit. A gauge-covariant bimetric tetrad theory of gravitation and electromagnetism. Foundations of Physics, 19(1):33–55, January 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00737765.

Woo:1989:CIC

[1087] C. H. Woo. Chaos, ineffectiveness, and the contrast between classical and quantal physics. Foundations of Physics, 19(1):57–76, January 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00737766.

Israelit:1989:MSW

[1088] Mark Israelit. Measuring standards in Weyl-type theories. Foundations of Physics, 19(1):77–90, January 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00737767.

Baez:1989:LI

[1089] John C. Baez. Is life improbable? Foundations of Physics, 19(1): 91–95, January 1989. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00737768.

Bar-Natan:1989:TEN

[1090] Dror Bar-Natan. Two examples in noncommutative probability. Foundations of Physics, 19(1):97–104, January 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00737769.

Sachs:1989:PFG

[1091] Mendel Sachs. The precessional frequency of a gyroscope in the quaternionic formulation of general relativity. Foundations of Physics, 19(1): 105–108, January 1989. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00737770.

Moshinsky:1989:BR

[1092] Marcos Moshinsky and Brian Skyrms. Book review. Foundations of Physics, 19(1):109–114, January 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00737771.

Cushing:1989:BRa

[1093] James T. Cushing, Grégoire Nicolis, and F. David Peat. Book review. Foundations of Physics, 19(1):115–123, January 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00737772.

Fryberger:1989:GED

[1094] David Fryberger. On generalized electromagnetism and Dirac algebra. Foundations of Physics, 19(2):125–159, February 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00734522.

Levine:1989:SMP

[1095] Robert Y. Levine and Robert R. Tucci. On the simultaneous measurement of position and momentum: Naimark embeddings and projections. Foundations of Physics, 19(2):161–173, February 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00734523.

Levine:1989:SMS

[1096] Robert Y. Levine and Robert R. Tucci. On the simultaneous measurement of spin components using spin-1/2 meters: Naimark embedding and projections. Foundations of Physics, 19(2):175–187, February 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00734524.

Hall:1989:QMC

[1097] Michael J. W. Hall. Quantum mechanics and the concept of joint probability. Foundations of Physics, 19(2):189–207, February 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00734525.

Phipps:1989:BR

[1098] Thomas E. Phipps, Jr. Book review. Foundations of Physics, 19(2): 209–219, February 1989. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00734526.

Weisskopf:1989:BR

[1099] Victor F. Weisskopf. Book review. Foundations of Physics, 19(2): 221–223, February 1989. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00734527.

Korringa:1989:BR

[1100] J. Korringa. Book review. Foundations of Physics, 19(2):225– 226, February 1989. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00734528.

Happer:1989:BR

[1101] William Happer. Book review. Foundations of Physics, 19(2):227– 228, February 1989. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00734529.

Greenberg:1989:BR

[1102] O. W. Greenberg. Book review. Foundations of Physics, 19(2):229– 230, February 1989. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00734530.

Fleming:1989:HDR

[1103] Gordon N. Fleming and Harry Bennett. Hyperplane dependence in relativistic quantum mechanics. Foundations of Physics, 19(3): 231–267, March 1989. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00734559.

Vargas:1989:EML

[1104] Jose G. Vargas and Douglas G. Torr. Electrodynamics of the Maxwell– Lorentz type in the ten-dimensional space of the testing of special relativity: a case for Finsler type connections. Foundations of Physics, 19 (3):269–291, March 1989. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00734560.

Gudder:1989:FAD

[1105] Stanley Gudder. Fuzzy amplitude densities and stochastic quantum mechanics. Foundations of Physics, 19(3):293–317, March 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00734561.

Beil:1989:ECC

[1106] R. G. Beil. The extended classical charged particle. Foundations of Physics, 19(3):319–338, March 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00734562.

Rosen:1989:EPBa

[1107] Nathan Rosen. Elementary particles in bimetric general relativity. Foundations of Physics, 19(3):339–348, March 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00734563.

Boyer:1989:CSC

[1108] Timothy H. Boyer. Conformal symmetry of classical electromagnetic zero-point radiation. Foundations of Physics, 19(4):349–365, April 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00731830.

Kannenberg:1989:QFS

[1109] L. Kannenberg. Quantum formalism via signal analysis. Foundations of Physics, 19(4):367–383, April 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00731831.

Dicke:1989:QMS

[1110] Robert H. Dicke. Quantum measurements, sequential and latent. Foundations of Physics, 19(4):385–395, April 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00731832.

Bedford:1989:NCR

[1111] Donald Bedford and Henry P. Stapp. Nonlocal character of the Rastall model. Foundations of Physics, 19(4):397–406, April 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:

//link.springer.com/article/10.1007/BF00731833. See comment

[1136].

Efinger:1989:NSM

[1112] H. J. Efinger. Nonlinear Schrödinger mechanics and the law of gravity. Foundations of Physics, 19(4):407–418, April 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00731834.

Cetto:1989:CDA

[1113] A. M. Cetto and L. de la Penã. Continuous and discrete aspects of blackbody radiation. Foundations of Physics, 19(4):419–437, April 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00731835.

Rosenblum:1989:NSS

[1114] Arnold Rosenblum. Non-standard synchronization of clocks outside a rotating body. Foundations of Physics, 19(4):439–442, April 1989. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00731836.

Bellomo:1989:BR

[1115] Nicola Bellomo and Jayant Narlikar. Book review. Foundations of Physics, 19(4):443–451, April 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00731837.

Fine:1989:CET

[1116] Arthur Fine. Correlations and efficiency: Testing the Bell inequalities. Foundations of Physics, 19(5):453–478, May 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00734654.

Manoukian:1989:TIS

[1117] Edward B. Manoukian. Theoretical intricacies of the single-slit, the double-slit, and related experiments in quantum mechanics. Foundations of Physics, 19(5):479–504, May 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00734655.

Maciel:1989:AAS

[1118] A. K. A. Maciel and J. Tiomno. Analysis of absolute space–time Lorentz theories. Foundations of Physics, 19(5):505–519, May 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00734656.

Maciel:1989:EAA

[1119] A. K. A. Maciel and J. Tiomno. Experimental analysis of absolute space– time Lorentz theories. Foundations of Physics, 19(5):521–530, May 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00734657.

Prokhovnik:1989:PBA

[1120] S. J. Prokhovnik and W. T. Morris. The physical basis of astronomical aberration. Foundations of Physics, 19(5):531–539, May 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00734658.

Prkhovnik:1989:TPS

[1121] S. J. Prkhovnik. The twin paradoxes of special relativity: Their resolution and implications. Foundations of Physics, 19(5):541–552, May 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00734659.

Atmanspacher:1989:AIP

[1122] Harald Atmanspacher. The aspect of information production in the process of observation. Foundations of Physics, 19(5):553–577, May 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00734660.

Ferrari:1989:CPH

[1123] L. Ferrari. The covariance problem and the Hamiltonian formalism in quantum mechanics. Foundations of Physics, 19(5):579–605, May 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00734661.

Gruszczak:1989:PEP

[1124] J. Gruszczak, M. Heller, and P. Multarzyński. Physics with and without the equivalence principle. Foundations of Physics, 19(5):607–618, May 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00734662.

Bishop:1989:DTP

[1125] N. T. Bishop. Do tachyons produce an unacceptable gravitational field? Foundations of Physics, 19(5):619–624, May 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00734663.

Cushing:1989:BRb

[1126] James T. Cushing. Book review. Foundations of Physics, 19(5): 625–627, May 1989. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00734664.

Fox:1989:BR

[1127] Ronald F. Fox. Book review. Foundations of Physics, 19(5):629–632, May 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00734665.

Busch:1989:DPF

[1128] Paul Busch and Pekka J. Lahti. The determination of the past and the future of a physical system in quantum mechanics. Foundations of Physics, 19(6):633–678, June 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00731904.

Greenberger:1989:HMQ

[1129] Daniel M. Greenberger and Alaine YaSin. “Haunted” measurements in quantum theory. Foundations of Physics, 19(6):679–704, June 1989. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00731905.

Rodrigues:1989:MTT

[1130] Waldyr A. Rodrigues, Jr. and Marcio A. F. Rosa. The meaning of time in the theory of relativity and “Einstein’s later view of the Twin Paradox”. Foundations of Physics, 19(6):705–724, June 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00731906. See [794].

deBeauregard:1989:CHE

[1131] O. Costa de Beauregard. The computer and the heat engine. Foundations of Physics, 19(6):725–727, June 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00731907. See response [1132].

Landauer:1989:RCH

[1132] Rolf Landauer. Response to “The computer and the heat engine”. Foundations of Physics, 19(6):729–732, June 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00731908. See [1131].

Sherry:1989:PCP

[1133] G. C. Sherry. Pseudo-classical phase space description of the relativistic electron. Foundations of Physics, 19(6):733–741, June 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00731909.

Costantini:1989:CQS

[1134] D. Costantini and U. Garibaldi. Classical and quantum statistics as finite random processes. Foundations of Physics, 19(6):743–754, June 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00731910.

Pereira:1989:RHS

[1135] R. A. Marques Pereira. Real and holomorphic SuSy algebras. Foundations of Physics, 19(6):755–782, June 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00731911.

Rastall:1989:CNC

[1136] Peter Rastall. Comment on “Nonlocal character of the Rastall model”. Foundations of Physics, 19(6):783, June 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00731912. See [1111].

Kreuzer:1989:BR

[1137] H. J. Kreuzer and O. W. Greenberg. Book review. Foundations of Physics, 19(6):785–786, June 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00731913.

Anonymous:1989:SFMa

[1138] Anonymous. Symposium on the foundations of modern physics 1990 13– 17 August 1990, joensuu, Finland. Foundations of Physics, 19(6):i, June 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00731914.

Busch:1989:PMP

[1139] Paul Busch and Alwyn van der Merwe. Peter Mittelstaedt: Philosopherphysicist. Foundations of Physics, 19(7):789–791, July 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01889300.

Bub:1989:BRE

[1140] Jeffrey Bub. On Bohr’s response to EPR: a quantum logical analysis. Foundations of Physics, 19(7):793–805, July 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01889301.

Busch:1989:RSH

[1141] Paul Busch and Franklin E. Schroeck, Jr. On the reality of spin and helicity. Foundations of Physics, 19(7):807–872, July 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01889302.

Cassinelli:1989:MSI

[1142] Gianni Cassinelli and Pekka J. Lahti. The measurement statistics interpretation of quantum mechanics: Possible values and possible measurement results of physical quantities. Foundations of Physics, 19 (7):873–890, July 1989. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889303.

DallaChiara:1989:PQL

[1143] Maria Luisa Dalla Chiara and Roberto Giuntini. Paraconsistent quantum logics. Foundations of Physics, 19(7):891–904, July 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01889304.

Foulis:1989:CPS

[1144] David J. Foulis. Coupled physical systems. Foundations of Physics, 19(7):905–922, July 1989. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889305.

Grabowski:1989:WO

[1145] Marian Grabowski. What is an observable? Foundations of Physics, 19(7):923–930, July 1989. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889306.

Giuntini:1989:TFL

[1146] Roberto Giuntini and Heinz Greuling. Toward a formal language for unsharp properties. Foundations of Physics, 19(7):931–945, July 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01889307.

Anonymous:1989:SFMb

[1147] Anonymous. Symposium on the Foundations of Modern Physics 1990, 13–17 August 1990, Joensuu, Finland. Foundations of Physics, 19 (7):947, July 1989. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889308.

Anonymous:1989:PPM

[1148] Anonymous. Photo of Peter Mittelstaedt. Foundations of Physics, 19(7):i, July 1989. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889299.

Gudder:1989:RQM

[1149] Stanley Gudder. Realism in quantum mechanics. Foundations of Physics, 19(8):949–970, August 1989. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883150.

Ludwig:1989:ATR

[1150] G. Ludwig. Atoms: are they real or are they objects? Foundations of Physics, 19(8):971–983, August 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01883151.

Kluppel:1989:STS

[1151] M. Klüppel and H. Neumann. The space–time structure of quantum systems in external fields. Foundations of Physics, 19(8):985–998, August 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF01883152.

Pavicic:1989:UQL

[1152] Mladen Pavicić. Unified quantum logic. Foundations of Physics, 19 (8):999–1016, August 1989. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883153.

Piron:1989:NDN

[1153] Constantin Piron. New dialogue on a new science between F. Salviati, G. Sagredo, and ?. Simplicio. Foundations of Physics, 19(8):1017– 1025, August 1989. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883154.

Quadt:1989:NPE

[1154] Ralf Quadt. The nonobjectivity of past events in quantum mechanics. Foundations of Physics, 19(8):1027–1035, August 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01883155.

Reiter:1989:IO

[1155] H. Reiter and W. Thirring. Are x and p incompatible observables? Foundations of Physics, 19(8):1037–1039, August 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01883156.

Ristig:1989:SQF

[1156] M. L. Ristig. The structure of quantum fluids. Foundations of Physics, 19 (8):1041–1050, August 1989. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883157.

Barbour:1989:MVN

[1157] Julian B. Barbour. Maximal variety as a new fundamental principle of dynamics. Foundations of Physics, 19(9):1051–1073, September 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01883158.

Hehl:1989:PMA

[1158] Friedrich W. Hehl, J. Dermott McCrea, Eckehard W. Mielke, and Yuval Ne’eman. Progress in metric-affine gauge theories of gravity with local scale invariance. Foundations of Physics, 19(9):1075–1100, September 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF01883159.

Ruttimann:1989:WDS

[1159] Gottfried T. Rüttimann. Weak density of states. Foundations of Physics, 19(9):1101–1112, September 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01883160.

Heintzmann:1989:TQP

[1160] H. Heintzmann. Third quantization: The problem of the cosmic vacuum, spontaneous symmetry breaking, and the possibility of an eternal universe. Foundations of Physics, 19(9):1113–1120, September 1989. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01883161.

Anonymous:1989:A

[1161] Anonymous. Announcements. Foundations of Physics, 19(9):1121– 1123, September 1989. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883162.

Saad:1989:SEM

[1162] David Saad, L. P. Horwitz, and R. I. Arshansky. Off-shell electromagnetism in manifestly covariant relativistic quantum mechanics. Foundations of Physics, 19(10):1125–1149, October 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00731876.

Rohrlich:1989:LRC

[1163] Fritz Rohrlich. The logic of reduction: The case of gravitation. Foundations of Physics, 19(10):1151–1170, October 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00731877.

Cormier-Delanoue:1989:RP

[1164] Christian Cormier-Delanoue. From relativity to the photon. Foundations of Physics, 19(10):1171–1190, October 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00731878.

Desloge:1989:TDS

[1165] Edward A. Desloge. A theoretical device for space and time measurements. Foundations of Physics, 19(10):1191–1213, October 1989. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00731879.

Norton:1989:CCE

[1166] John Norton. Coordinates and covariance: Einstein’s view of space– time and the modern view. Foundations of Physics, 19(10):1215– 1263, October 1989. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00731880.

Evenson:1989:BR

[1167] William E. Evenson. Book review. Foundations of Physics, 19(10): 1265–1269, October 1989. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00731881.

Legget:1989:BR

[1168] Anthony J. Legget. Book review. Foundations of Physics, 19(10): 1271–1273, October 1989. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00731882.

Ford:1989:BR

[1169] Joseph Ford and Gottfried T. Rüttimann. Book review. Foundations of Physics, 19(10):1275–1281, October 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00731883.

Anonymous:1989:SQRa

[1170] Anonymous. In search of quantum reality. Foundations of Physics, 19 (10):1283, October 1989. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00731884.

Howard:1989:CSE

[1171] Stephen D. Howard and Sanat K. Roy. Coherent states as extreme energy states. Foundations of Physics, 19(11):1285–1298, November 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00732751.

Schindler:1989:PGI

[1172] Christian Schindler. Physical and geometrical interpretation of the Jordan–Hahn and the Lebesgue decomposition property. Foundations of Physics, 19(11):1299–1314, November 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00732752.

Sibelius:1989:IWP

[1173] Patrick Sibelius. An interpretation within philosophy of the relationship between classical mechanics and quantum mechanics. Foundations of Physics, 19(11):1315–1326, November 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00732753.

Jozsa:1989:PMS

[1174] Richard Jozsa. The Pitowsky model and superluminal signals. Foundations of Physics, 19(11):1327–1335, November 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00732754.

Rosen:1989:EPBb

[1175] Nathan Rosen. Elementary particles in bimetric general relativity. II. Foundations of Physics, 19(11):1337–1344, November 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00732755.

Woodward:1989:NCE

[1176] James F. Woodward. On nonminimal coupling of the electromagnetic and gravitational fields: The astrophysical evidence for the Schuster– Blackett conjecture and its implications. Foundations of Physics, 19(11): 1345–1361, November 1989. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00732756.

Boya:1989:SSP

[1177] Luis J. Boya. State space as projective space. The case of massless particles. Foundations of Physics, 19(11):1363–1370, November 1989. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00732757.

Boyer:1989:SST

[1178] Timothy H. Boyer. Scaling symmetry and thermodynamic equilibrium for classical electromagnetic radiation. Foundations of Physics, 19(11): 1371–1383, November 1989. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00732758.

Ungar:1989:RVC

[1179] Abraham A. Ungar. The relativistic velocity composition paradox and the Thomas rotation. Foundations of Physics, 19(11):1385– 1396, November 1989. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00732759.

Dieks:1989:QMP

[1180] D. Dieks. Quantum mechanics without the projection postulate and its realistic interpretation. Foundations of Physics, 19(11):1397– 1423, November 1989. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00732760.

Shimony:1989:BR

[1181] Abner Shimony. Book review. Foundations of Physics, 19(11):1425– 1429, November 1989. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00732761.

Landsberg:1989:BR

[1182] P. T. Landsberg and James W. Dufty. Book review. Foundations of Physics, 19(11):1431–1434, November 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00732762.

Anonymous:1989:SQRb

[1183] Anonymous. In search of quantum reality. Foundations of Physics, 19 (11):1435, November 1989. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00732763.

Anonymous:1989:NASa

[1184] Anonymous. NATO Advanced Study Institute: Information Dynamics in Irsee/Bavaria (FRG) June 15–26, 1990. Foundations of Physics, 19(11): 1437–1438, November 1989. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00732764.

Anonymous:1989:SFMc

[1185] Anonymous. Symposium on the Foundations of Modern Physics 1990 August 13–17, 1990, Joensuu, Finland. Foundations of Physics, 19(11): 1439, November 1989. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00732765.

Salisbury:1989:QRA

[1186] Donald C. Salisbury and Michael Pollot. Quantum relativistic action at a distance. Foundations of Physics, 19(12):1441–1477, December 1989.

CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00732507.

Drechsler:1989:MWT

[1187] W. Drechsler. Modified Weyl theory and extended elementary objects. Foundations of Physics, 19(12):1479–1497, December 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00732508.

Radin:1989:ECR

[1188] Dean I. Radin and Roger D. Nelson. Evidence for consciousness-related anomalies in random physical systems. Foundations of Physics, 19(12): 1499–1514, December 1989. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00732509.

Varga:1989:PLT

[1189] Péter Varga. On the possible linear term in light coincidence experiments. Foundations of Physics, 19(12):1515–1523, December 1989. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00732510.

Sachs:1989:RRR

[1190] Mendel Sachs. Response to Rodrigues and Rosa on the Twin Paradox. Foundations of Physics, 19(12):1525–1528, December 1989. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00732511.

Stott:1989:BR

[1191] P. E. Stott. Book review. Foundations of Physics, 19(12):1529– 1530, December 1989. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00732512.

Anonymous:1989:NASb

[1192] Anonymous. NATO Advanced Study Institute: Information Dynamics in Irsee/Bavaria (FRG) June 15–26, 1990. Foundations of Physics, 19(12): 1531–1532, December 1989. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00732513.

Anonymous:1989:SFMd

[1193] Anonymous. Symposium on the Foundations of Modern Physics 1990 August 13–17, 1990, Joensuu, Finland. Foundations of Physics, 19 (12):1533, December 1989. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00732514.

Busch:1990:ETUa

[1194] Paul Busch. On the energy-time uncertainty relation. Part I: Dynamical time and time indeterminacy. Foundations of Physics, 20 (1):1–32, January 1990. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00732932.

Busch:1990:ETUb

[1195] Paul Busch. On the energy-time uncertainty relation. Part II: Pragmatic time versus energy indeterminacy. Foundations of Physics, 20 (1):33–43, January 1990. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00732933.

Spavieri:1990:MES

[1196] Gianfranco Spavieri. Model of the electron spin in stochastic physics. Foundations of Physics, 20(1):45–61, January 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00732934.

Diosi:1990:LDM

[1197] L. Diósi. Landau’s density matrix in quantum electrodynamics. Foundations of Physics, 20(1):63–70, January 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00732935.

Carmeli:1990:FTT

[1198] M. Carmeli and A. Malka. Field theory on R × S^{3} topology: Lagrangian formulation. Foundations of Physics, 20(1):71–110, January 1990. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00732936.

Zet:1990:GTT

[1199] G. Zet. Gauge theory on R × S^{3} topology. Foundations of Physics, 20 (1):111–117, January 1990. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00732937.

Cattaneo:1990:NAD

[1200] Gianpiero Cattaneo and Giuseppe Nisticò. A note on Aerts’ description of separated entities. Foundations of Physics, 20(1):119–132, January 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00732938.

Wolf:1990:TDQ

[1201] C. Wolf. Testing discrete quantum mechanics using neutron interferometry and the superposition principle — a gedanken experiment. Foundations of Physics, 20(1):133–137, January 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00732939.

Wignall:1990:MET

[1202] J. W. G. Wignall. Maxwell electrodynamics from a theory of macroscopically extended particles. Foundations of Physics, 20(2):139– 158, February 1990. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00731644.

Frenkel:1990:SLW

[1203] Andor Frenkel. Spontaneous localizations of the wave function and classical behavior. Foundations of Physics, 20(2):159–188, February 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00731645.

Fanchi:1990:TKP

[1204] John R. Fanchi. Tachyon kinematics in parametrized relativistic quantum mechanics. Foundations of Physics, 20(2):189–224, February 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00731646.

Cole:1990:CES

[1205] Daniel C. Cole. Classical electrodynamic systems interacting with classical electromagnetic random radiation. Foundations of Physics, 20(2): 225–240, February 1990. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00731647.

vBorzeszkowski:1990:PES

[1206] H. H. v. Borzeszkowski and H. J. Treder. Planckions and the early stage of the universe. Foundations of Physics, 20(2):241–250, February 1990.

CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00731648.

Ryan:1990:BR

[1207] Michael P. Ryan, Jr. Book review. Foundations of Physics, 20(2): 251–253, February 1990. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00731649.

Martens:1990:NQM

[1208] Hans Martens and Willem M. de Muynck. Nonideal quantum measurements. Foundations of Physics, 20(3):255–281, March 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00731693.

Rosen:1990:FMS

[1209] Joe Rosen. Fundamental manifestations of symmetry in physics. Foundations of Physics, 20(3):283–307, March 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00731694.

Evans:1990:FSF

[1210] A. B. Evans. Four-space formulation of Dirac’s equation. Foundations of Physics, 20(3):309–335, March 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00731695.

Peruzzi:1990:LAQ

[1211] G. Peruzzi. Logical anomalies of quantum objects. A survey. Foundations of Physics, 20(3):337–352, March 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00731696.

Wald:1990:BR

[1212] Robert Wald. Book review. Foundations of Physics, 20(3):353– 356, March 1990. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00731697.

Martens:1990:IP

[1213] Hans Martens and Willem M. de Muynck. The inaccuracy principle. Foundations of Physics, 20(4):357–380, April 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00731707.

Garrett:1990:BTI

[1214] A. J. M. Garrett. Bell’s theorem, inference, and quantum transactions. Foundations of Physics, 20(4):381–402, April 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00731708.

deAndrade:1990:SSS

[1215] L. C. Garcia de Andrade. The structure of singularities in space–times with torsion. Foundations of Physics, 20(4):403–416, April 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00731709.

Tome:1990:COQ

[1216] W. Tomé and S. Gudder. Convergence of observables on quantum logics. Foundations of Physics, 20(4):417–434, April 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00731710.

vonBorzeszkowski:1990:RPR

[1217] Horst-Heino von Borzeszkowski. Remarks on the physical reality of gravitons. Foundations of Physics, 20(4):435–445, April 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00731711.

Hu:1990:CQM

[1218] Wayne Hu. The curious quantum mechanics of pre- and post-selected ensembles. Foundations of Physics, 20(4):447–458, April 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00731712.

Beyer:1990:NCE

[1219] Horst Beyer and Jürgen Nitsch. A note on a Casimir effect in a uniformly accelerated reference frame. Foundations of Physics, 20(4):459–469, April 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00731713.

Singer:1990:IIO

[1220] Matthias Singer and Werner Stulpe. Informational incompleteness of the observables S_{x}, S_{y}, S_{z} for spin-1 systems. Foundations of Physics, 20(4):471–472, April 1990. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00731714.

Lindenberg:1990:BR

[1221] Katja Lindenberg. Book review. Foundations of Physics, 20(4): 473–475, April 1990. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00731715.

Foulis:1990:CHR

[1222] D. J. Foulis and M. K. Bennett. Charles Hamilton Randall: 1928–1987. Foundations of Physics, 20(5):473–476, May 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01883235.

Lock:1990:TPG

[1223] Robin H. Lock. The tensor product of generalized sample spaces which admit a unital set of dispersion-free weights. Foundations of Physics, 20(5):477–498, May 1990. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883236.

Gudder:1990:QPO

[1224] Stanley Gudder. Quantum probability and operational statistics. Foundations of Physics, 20(5):499–527, May 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01883237.

Gross:1990:HLN

[1225] Herbert Gross. Hilbert lattices: New results and unsolved problems. Foundations of Physics, 20(5):529–559, May 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01883238.

Schindler:1990:UJH

[1226] Christian Schindler. The unique Jordan–Hahn decomposition property. Foundations of Physics, 20(5):561–573, May 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01883239.

Keller:1990:MID

[1227] Hans A. Keller. Measures on infinite-dimensional orthomodular spaces. Foundations of Physics, 20(5):575–604, May 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01883240.

Dacey:1990:ATQ

[1228] J. C. Dacey. Arithmetic tools for quantum logic. Foundations of Physics, 20(5):605–619, May 1990. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883241.

Ludwig:1990:CSP

[1229] G. Ludwig. Concepts of states in physics. Foundations of Physics, 20 (6):621–633, June 1990. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889451.

Svetlichny:1990:IFP

[1230] George Svetlichny. On the inverse FPR problem: Quantum is classical. Foundations of Physics, 20(6):635–650, June 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01889452.

Piziak:1990:LTQ

[1231] Robert Piziak. Lattice theory, quadratic spaces, and quantum proposition systems. Foundations of Physics, 20(6):651–665, June 1990. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01889453.

Catlin:1990:SEO

[1232] Donald E. Catlin. The Schrödinger equation via an operator functional equation. Foundations of Physics, 20(6):667–690, June 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01889454.

Younce:1990:RUM

[1233] Matthew B. Younce. Refinement and unique Mackey decomposition for manuals and orthoalgebras. Foundations of Physics, 20(6):691–700, June 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF01889455.

Giuntini:1990:BZL

[1234] Roberto Giuntini. Brouwer–Zadeh logic and the operational approach to quantum mechanics. Foundations of Physics, 20(6):701–714, June 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01889456.

Janowitz:1990:IOS

[1235] M. F. Janowitz. Interval order and semiorder lattices. Foundations of Physics, 20(6):715–732, June 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01889457.

Bennett:1990:SQC

[1236] M. K. Bennett and D. J. Foulis. Superposition in quantum and classical mechanics. Foundations of Physics, 20(6):733–744, June 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01889458.

Mielnik:1990:PTB

[1237] Bogdan Mielnik. The paradox of two bottles in quantum mechanics. Foundations of Physics, 20(6):745–755, June 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01889459.

Anonymous:1990:PCHa

[1238] Anonymous. Photo of Charles H. Randall. Foundations of Physics, 20(6):i, June 1990. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889450.

Busch:1990:QTS

[1239] Paul Busch, Gianni Cassinelli, and Pekka J. Lahti. On the quantum theory of sequential measurements. Foundations of Physics, 20 (7):757–775, July 1990. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889690.

Klay:1990:MLE

[1240] Matthias P. Kläy and David J. Foulis. Maximum likelihood estimation on generalized sample spaces: an alternative resolution of Simpson’s paradox. Foundations of Physics, 20(7):777–799, July 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01889691.

Kalmbach:1990:QMS

[1241] G. Kalmbach. Quantum measure spaces. Foundations of Physics, 20 (7):801–821, July 1990. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889692.

Schroeck:1990:SQM

[1242] F. E. Schroeck, Jr. and D. J. Foulis. Stochastic quantum mechanics viewed from the language of manuals. Foundations of Physics, 20 (7):823–858, July 1990. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889693.

Edwards:1990:CPS

[1243] C. Martin Edwards and Gottfried T. Rüttimann. On conditional probability in GL spaces. Foundations of Physics, 20(7):859–872, July 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01889694.

Schweizer:1990:BTE

[1244] B. Schweizer and A. Sklar. The baker’s transformation is not embeddable. Foundations of Physics, 20(7):873–879, July 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01889695.

Wright:1990:GUM

[1245] Ron Wright. Generalized urn models. Foundations of Physics, 20 (7):881–903, July 1990. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889696.

Olubummo:1990:OLH

[1246] Yewande Olubummo and Thurlow A. Cook. Operational logics and the Hahn–Jordan property. Foundations of Physics, 20(7):905–913, July 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF01889697.

Greechie:1990:STO

[1247] Richard J. Greechie. Sites and tours in orthoalgebras and orthomodular lattices. Foundations of Physics, 20(7):915–923, July 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01889698.

Anonymous:1990:PCHb

[1248] Anonymous. Publications of Charles H. Randall. Foundations of Physics, 20(7):925–927, July 1990. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889699.

Anonymous:1990:PCHc

[1249] Anonymous. Photo of Charles H. Randall. Foundations of Physics, 20(7):i, July 1990. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889689.

Bub:1990:BRE

[1250] Jeffrey Bub. On Bohr’s response to EPR: II. Foundations of Physics, 20 (8):929–941, August 1990. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00738373.

Guy:1990:FEE

[1251] Reed Guy and Robert Deltete. Fine, Einstein, and ensembles. Foundations of Physics, 20(8):943–965, August 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00738374.

Fine:1990:EER

[1252] A. Fine. Einstein and ensembles: Response. Foundations of Physics, 20 (8):967–989, August 1990. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00738375.

Hoffmann:1990:LRE

[1253] M. Hoffmann. A local realistic explanation of EPR correlations. Foundations of Physics, 20(8):991–1003, August 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00738376.

Multarzynski:1990:DCS

[1254] Piotr Multarzyński and Michael Heller. The differential and cone structures of spacetime. Foundations of Physics, 20(8):1005–1015, August 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00738377.

Brody:1990:BR

[1255] Burt Brody. Book review. Foundations of Physics, 20(8):1017– 1019, August 1990. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00738378.

Mazo:1990:BR

[1256] Robert M. Mazo. Book review. Foundations of Physics, 20(8):1021– 1023, August 1990. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00738379.

Enz:1990:BR

[1257] Charles P. Enz and A. Lodder. Book review. Foundations of Physics, 20 (8):1025–1031, August 1990. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00738380.

Sibelius:1990:MWT

[1258] Patrick Sibelius. The mechanical and the wave-theoretical aspects of momentum considering discrete action. Foundations of Physics, 20(9): 1033–1059, September 1990. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00731851.

Ryff:1990:SQM

[1259] L. C. B. Ryff. Should quantum mechanical description of physical reality be considered complete? Foundations of Physics, 20(9):1061– 1078, September 1990. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00731852.

Rupertsberger:1990:EDM

[1260] H. Rupertsberger. Electric dipole moments reconsidered. Foundations of Physics, 20(9):1079–1089, September 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00731853.

Dvurecenskij:1990:FAS

[1261] Anatolij Dvurecenskij, Tibor Neubrunn, and Sylvia Pulmannová. Finitely additive states and completeness of inner product spaces. Foundations of Physics, 20(9):1091–1102, September 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00731854.

Thirring:1990:SM

[1262] W. Thirring. Stability of matter. Foundations of Physics, 20(9):1103– 1110, September 1990. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00731855.

Smith:1990:CFOa

[1263] C. Ray Smith, Steven R. Rolf, and Ramarao Inguva. Coordinate-free operators based on one vector. I. Formal considerations. Foundations of Physics, 20(9):1111–1122, September 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00731856.

Smith:1990:CFOb

[1264] C. Ray Smith, Steven R. Rolf, and Ramarao Inguva. Coordinate-free operators based on one vector. II. Applications in electrodynamics. Foundations of Physics, 20(9):1123–1133, September 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00731857.

Bertlmann:1990:JSB

[1265] Reinhold A. Bertlmann. John Stewart Bell — physicist and moralizer. Foundations of Physics, 20(10):1135–1138, October 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01889461.

Jammer:1990:JSB

[1266] Max Jammer. John Stewart Bell and his work — on the occasion of his sixtieth birthday. Foundations of Physics, 20(10):1139–1145, October 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF01889462.

dEspagnat:1990:TSE

[1267] Bernard d’Espagnat. Towards a separable “empirical reality”? Foundations of Physics, 20(10):1147–1172, October 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01889463.

Garuccio:1990:PCS

[1268] A. Garuccio, V. L. Lepore, and F. Selleri. Probabilities for correlated systems. Foundations of Physics, 20(10):1173–1189, October 1990. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01889464.

Bertlmann:1990:BTN

[1269] R. A. Bertlmann. Bell’s theorem and the nature of reality. Foundations of Physics, 20(10):1191–1212, October 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01889465.

Hestenes:1990:ZIQ

[1270] David Hestenes. The Zitterbewegung interpretation of quantum mechanics. Foundations of Physics, 20(10):1213–1232, October 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01889466.

Barut:1990:QTS

[1271] A. O. Barut. Quantum theory of single events: Localized de Broglie wavelets, Schrödinger waves, and classical trajectories. Foundations of Physics, 20(10):1233–1240, October 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01889467.

Cushing:1990:BR

[1272] James T. Cushing, Roberto Peccei, Paul Teller, and Leopold Halpern. Book review. Foundations of Physics, 20(10):1241–1263, October 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01889468.

Newman:1990:BR

[1273] Edward H. Newman. Book review. Foundations of Physics, 20(10): 1265–1267, October 1990. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889469.

Harrison:1990:BR

[1274] James F. Harrison. Book review. Foundations of Physics, 20(10): 1269–1270, October 1990. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889470.

Anonymous:1990:PJS

[1275] Anonymous. Photo of John Stewart Bell. Foundations of Physics, 20 (10):iii, October 1990. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889460.

Ghirardi:1990:RDR

[1276] G. C. Ghirardi, R. Grassi, and P. Pearle. Relativistic dynamical reduction models: General framework and examples. Foundations of Physics, 20(11):1271–1316, November 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01883487.

Nicrosini:1990:RBC

[1277] Oreste Nicrosini and Alberto Rimini. On the relationship between continuous and discontinuous stochastic processes in Hilbert space. Foundations of Physics, 20(11):1317–1327, November 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01883488.

Ballentine:1990:LPP

[1278] L. E. Ballentine. Limitations of the projection postulate. Foundations of Physics, 20(11):1329–1343, November 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01883489.

Gudder:1990:QSP

[1279] Stanley Gudder. Quantum stochastic processes. Foundations of Physics, 20(11):1345–1363, November 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01883490.

Wootters:1990:RQS

[1280] William K. Wootters. Random quantum states. Foundations of Physics, 20(11):1365–1378, November 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01883491.

Brown:1990:NGL

[1281] H. R. Brown and G. Svetlichny. Nonlocality and Gleason’s lemma. Part I. Deterministic theories. Foundations of Physics, 20(11):1379– 1387, November 1990. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883492.

Elby:1990:NGL

[1282] Andrew Elby. Nonlocality and Gleason’s lemma. Part 2. Stochastic theories. Foundations of Physics, 20(11):1389–1397, November 1990. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01883493.

Rohrlich:1990:TGP

[1283] Fritz Rohrlich. There is good physics in theory reduction. Foundations of Physics, 20(11):1399–1412, November 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01883494.

Aspect:1990:LMC

[1284] Alain Aspect and Robin Kaiser. Linear momentum conservation in coherent population trapping: a case study for a quantum filtering process. Foundations of Physics, 20(12):1413–1428, December 1990. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01883515.

Busch:1990:CPM

[1285] Paul Busch and Pekka J. Lahti. Completely positive mappings in quantum dynamics and measurement theory. Foundations of Physics, 20(12): 1429–1439, December 1990. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883516.

Peres:1990:NTQ

[1286] Asher Peres. Neumark’s theorem and quantum inseparability. Foundations of Physics, 20(12):1441–1453, December 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01883517.

Cohen:1990:MSA

[1287] Leon Cohen. Momentum spread: amplitude and current contributions. Foundations of Physics, 20(12):1455–1473, December 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01883518.

Garrett:1990:BTB

[1288] A. J. M. Garrett. Bell’s theorem and Bayes’ theorem. Foundations of Physics, 20(12):1475–1512, December 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01883519.

Tisza:1990:BR

[1289] Laszlo Tisza. Book review. Foundations of Physics, 20(12):1513– 1517, December 1990. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883520.

DeFacio:1990:BR

[1290] Brian DeFacio. Book review. Foundations of Physics, 20(12):1519– 1521, December 1990. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883521.

Mitchell:1990:BR

[1291] J. W. Mitchell. Book review. Foundations of Physics, 20(12):1523– 1524, December 1990. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883522.

Stapp:1991:EBT

[1292] Henry P. Stapp. EPR and Bell’s theorem: a critical review. Foundations of Physics, 21(1):1–23, January 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01883560.

Bub:1991:MBQ

[1293] Jeffrey Bub. Measurement and “beables” in quantum mechanics. Foundations of Physics, 21(1):25–42, January 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01883561.

Redhead:1991:PPL

[1294] Michael Redhead and Paul Teller. Particles, particle labels, and quanta: The toll of unacknowledged metaphysics. Foundations of Physics, 21 (1):43–62, January 1991. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883562.

Bitsakis:1991:MME

[1295] Eftichios Bitsakis. Mass, matter, and energy. A relativistic approach. Foundations of Physics, 21(1):63–81, January 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01883563.

Park:1991:QAS

[1296] James L. Park. Quantum assembly semantics: The fallacious lingo of occupation numbers. Foundations of Physics, 21(1):83–92, January 1991.

CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01883564.

Prugovecki:1991:GSL

[1297] Eduard Prugovecki. Geometro-stochastic locality in quantum spacetime and quantum diffusions. Foundations of Physics, 21(1):93–124, January 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF01883565.

Vigier:1991:EMC

[1298] Jean-Pierre Vigier. Explicit mathematical construction of relativistic nonlinear de Broglie waves described by three-dimensional (wave and electromagnetic) solitons “piloted” (controlled) by corresponding solutions of associated linear Klein–Gordon and Schrödinger equations. Foundations of Physics, 21(2):125–148, February 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01889529.

Clifton:1991:GGH

[1299] Robert K. Clifton, Michael L. G. Redhead, and Jeremy N. Butterfield. Generalization of the Greenberger–Horne–Zeilinger algebraic proof of nonlocality. Foundations of Physics, 21(2):149–184, February 1991. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01889530.

Tyapkin:1991:GAB

[1300] Alexei A. Tyapkin and Milan Vindushka. The geometrical aspects of the bell inequalities. Foundations of Physics, 21(2):185–195, February 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01889531.

Josephson:1991:BUQ

[1301] Brian D. Josephson and Fotini Pallikari-Viras. Biological utilization of quantum nonlocality. Foundations of Physics, 21(2):197–207, February 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF01889532.

Marshall:1991:WDN

[1302] Trevor W. Marshall. What does noise do to the bell inequalities? Foundations of Physics, 21(2):209–219, February 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01889533.

Santos:1991:IQF

[1303] Emilio Santos. Interpretation of the quantum formalism and Bell’s theorem. Foundations of Physics, 21(2):221–241, February 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01889534.

Bohm:1991:RIQ

[1304] D. Bohm and B. J. Hiley. On the relativistic invariance of a quantum theory based on beables. Foundations of Physics, 21(2):243–250, February 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF01889535.

Greenberger:1991:SNA

[1305] Daniel Greenberger. A short note on the Aharonov–Bohm effect. Foundations of Physics, 21(2):251–255, February 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01889536.

Percus:1991:BR

[1306] Jerome K. Percus and Michael F. Shlesinger. Book review. Foundations of Physics, 21(2):257–261, February 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01889537.

Anonymous:1991:EN

[1307] Anonymous. Editorial note. Foundations of Physics, 21(2):i, February 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF01889528.

Hasegawa:1991:QMD

[1308] H. Hasegawa, T. Petrosky, I. Prigogine, and S. Tasaki. Quantum mechanics and the direction of time. Foundations of Physics, 21(3): 263–281, March 1991. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883634.

Treder:1991:FF

[1309] Hans-Jürgen Treder. On the fifth forces. Foundations of Physics, 21 (3):283–298, March 1991. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883635.

Land:1991:GFS

[1310] M. C. Land and L. P. Horwitz. Green’s functions for off-shell electromagnetism and spacelike correlations. Foundations of Physics, 21 (3):299–310, March 1991. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883636.

vonWeizsacker:1991:QRI

[1311] C. F. von Weizsäcker and Th. Görnitz. Quantum-realistic interpretation. Foundations of Physics, 21(3):311–321, March 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01883637.

Hilgevoord:1991:UPI

[1312] Jan Hilgevoord and Jos Uffink. Uncertainty in prediction and in inference. Foundations of Physics, 21(3):323–341, March 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01883638.

Pitowsky:1991:BQP

[1313] Itamar Pitowsky. Bohm’s quantum potentials and quantum gravity. Foundations of Physics, 21(3):343–352, March 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01883639.

Leggett:1991:CSB

[1314] Anthony J. Leggett and Fernando Sols. On the concept of spontaneously broken gauge symmetry in condensed matter physics. Foundations of Physics, 21(3):353–364, March 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01883640.

Fine:1991:INC

[1315] Arthur Fine. Inequalities for nonideal correlation experiments. Foundations of Physics, 21(3):365–378, March 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01883641.

Vargas:1991:GE

[1316] Jose G. Vargas. On the geometrization of electrodynamics. Foundations of Physics, 21(4):379–401, April 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00733354.

Ferrero:1991:CBQ

[1317] M. Ferrero and T. W. Marshall. The compatibility between quantum mechanics and local realist theories in atomic cascade experiments. Foundations of Physics, 21(4):403–415, April 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00733355.

Coleman:1991:EPS

[1318] Robert Alan Coleman and Herbert Korte. An empirical, purely spatial criterion for the planes of F -simultaneity. Foundations of Physics, 21 (4):417–437, April 1991. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00733356.

Fujita:1991:ICP

[1319] S. Fujita. On the indistinguishability of classical particles. Foundations of Physics, 21(4):439–457, April 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00733357.

Rosen:1991:COU

[1320] Nathan Rosen. Can one have a universal time in general relativity? Foundations of Physics, 21(4):459–472, April 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00733358.

Zet:1991:GTS

[1321] G. Zet, C. Pasnicu, and M. Agop. Gravitation theory in the spacetime R × S^{3}. Foundations of Physics, 21(4):473–481, April 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00733359.

Schindler:1991:QLE

[1322] Christian Schindler. Quantum logics with the existence property. Foundations of Physics, 21(4):483–498, April 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00733360.

Padmanabhan:1991:BRa

[1323] T. Padmanabhan. Book review. Foundations of Physics, 21(4): 499–501, April 1991. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00733361.

Padmanabhan:1991:BRb

[1324] T. Padmanabhan and Bruce Hevly. Book review. Foundations of Physics, 21(4):503–507, April 1991. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00733362.

Shimony:1991:JSB

[1325] Abner Shimony. John Stewart Bell: 1928–1990. Foundations of Physics, 21(5):509–512, May 1991. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00733256.

Drechsler:1991:GSQ

[1326] Wolfgang Drechsler and Eduard Prugovecki. Geometro-stochastic quantization of a theory for extended elementary objects. Foundations of Physics, 21(5):513–546, May 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00733257.

Hnilo:1991:HVD

[1327] Alejandro A. Hnilo. Hidden variables with directionalization. Foundations of Physics, 21(5):547–567, May 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00733258.

Ungar:1991:SLT

[1328] Abraham A. Ungar. Successive Lorentz transformations of the electromagnetic field. Foundations of Physics, 21(5):569–589, May 1991. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00733259.

Stuart:1991:ICI

[1329] C. I. J. M. Stuart. Inconsistency of the Copenhagen interpretation. Foundations of Physics, 21(5):591–622, May 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00733260.

Coon:1991:BR

[1330] Sidney A. Coon. Book review. Foundations of Physics, 21(5): 623–625, May 1991. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00733261.

Judd:1991:BR

[1331] Brian R. Judd. Book review. Foundations of Physics, 21(5):627–629, May 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00733262.

Rizzo:1991:BR

[1332] Thomas G. Rizzo. Book review. Foundations of Physics, 21(5): 631–632, May 1991. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00733263.

Evans:1991:KPF

[1333] A. B. Evans. Klein’s paradox in a four-space formulation of Dirac’s equation. Foundations of Physics, 21(6):633–647, June 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00733274.

deCastro:1991:QHJ

[1334] Antonio Soares de Castro and Alvaro de Souza Dutra. On the quantum Hamilton–Jacobi formalism. Foundations of Physics, 21(6): 649–663, June 1991. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00733275.

Oliver:1991:RPC

[1335] M. A. Oliver. The reconciliation of physics with cosmology. Foundations of Physics, 21(6):665–689, June 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00733276.

Ungar:1991:FDR

[1336] Abraham A. Ungar. Formalism to deal with Reichenbach’s special theory of relativity. Foundations of Physics, 21(6):691–726, June 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00733277.

Roman:1991:QLR

[1337] L. Román and B. Rumbos. Quantum logic revisited. Foundations of Physics, 21(6):727–734, June 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00733278.

Ablamowicz:1991:CR

[1338] Rafal Ablamowicz, Pertti Lounesto, and Johannes Maks. Conference report. Foundations of Physics, 21(6):735–748, June 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00733279.

Weitzner:1991:BR

[1339] Harold Weitzner and Daniel M. Greenberger. Book review. Foundations of Physics, 21(6):749–752, June 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00733280.

Anonymous:1991:Ca

[1340] Anonymous. Corrigendum. Foundations of Physics, 21(6):753–755, June 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00733281.

Frieden:1991:FIC

[1341] B. Roy Frieden. Fisher information and the complex nature of the Schrödinger wave equation. Foundations of Physics, 21(7):757–771, July 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00733343.

Foster:1991:SNG

[1342] Sara Foster and Andrew Elby. A SQUID no-go theorem without macrorealism: What SQUID’s really tell us about nature. Foundations of Physics, 21(7):773–785, July 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00733344.

Bartocci:1991:SAC

[1343] Umberto Bartocci and Marco Mamone Capria. Symmetries and asymmetries in classical and relativistic electrodynamics. Foundations of Physics, 21(7):787–801, July 1991. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00733345.

Zachar:1991:ICR

[1344] Oron Zachar and Orly Alter. Interpretation of the curious results of the new quantum formalism of pre- and post-selected systems. Foundations of Physics, 21(7):803–820, July 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00733346.

Dotson:1991:WDW

[1345] Allen C. Dotson. What determines whether a wave function is inherently necessary? Foundations of Physics, 21(7):821–829, July 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00733347.

Adomian:1991:ASS

[1346] G. Adomian. An analytical solution of the stochastic Navier–Stokes system. Foundations of Physics, 21(7):831–843, July 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00733348.

Boniolo:1991:QMI

[1347] Giovanni Boniolo and Francesco Gonella. Quantum mechanics and imagery: a hermeneutical approach. Foundations of Physics, 21(7): 845–854, July 1991. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00733349.

Trnkova:1991:AOL

[1348] Vera Trnková. Automorphisms of orthomodular lattices and symmetries of quantum logics. Foundations of Physics, 21(7):855–860, July 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00733350.

Rae:1991:BR

[1349] Alastair I. M. Rae. Book review. Foundations of Physics, 21(7): 861–863, July 1991. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00733351.

Ginsparg:1991:BR

[1350] Paul Ginsparg. Book review. Foundations of Physics, 21(7):865–867, July 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00733352.

Anonymous:1991:Cb

[1351] Anonymous. Corrigendum. Foundations of Physics, 21(7):869, July 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00733353.

Mladenov:1991:GQF

[1352] Ivailo M. Mladenov. Geometric quantization of the five-dimensional Kepler problem. Foundations of Physics, 21(8):871–888, August 1991. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00733213.

Busch:1991:POQ

[1353] Paul Busch and Peter Mittelstaedt. The problem of objectification in quantum mechanics. Foundations of Physics, 21(8):889–904, August 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00733214.

Stapleton:1991:AFE

[1354] David Stapleton. An application of functional equations to the analysis of the invariance identities of classical gauge field theory. Foundations of Physics, 21(8):905–929, August 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00733215.

Schulman:1991:SSQ

[1355] L. S. Schulman. “Special” states in quantum measurement apparatus: Structural requirements for the recovery of standard probabilities. Foundations of Physics, 21(8):931–945, August 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00733216.

Vaidman:1991:QTM

[1356] Lev Vaidman. A quantum time machine. Foundations of Physics, 21 (8):947–958, August 1991. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00733217.

Price:1991:ARR

[1357] Huw Price. The asymmetry of radiation: Reinterpreting the Wheeler– Feynman argument. Foundations of Physics, 21(8):959–975, August 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00733218.

Rosen:1991:SGU

[1358] Joe Rosen. Self-generating Universe and many worlds. Foundations of Physics, 21(8):977–981, August 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00733219.

Adomian:1991:LNS

[1359] G. Adomian and R. Rach. Linear and nonlinear Schrödinger equations. Foundations of Physics, 21(8):983–991, August 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00733220.

dEspagnat:1991:BR

[1360] Bernard d’Espagnat. Book review. Foundations of Physics, 21(8): 993–995, August 1991. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00733221.

Holman:1991:BR

[1361] Richard Holman. Book review. Foundations of Physics, 21(8):997– 999, August 1991. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00733222.

Getoor:1991:BR

[1362] Ronald Getoor. Book review. Foundations of Physics, 21(8):1001– 1002, August 1991. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00733223.

Anonymous:1991:FAa

[1363] Anonymous. First announcement. Foundations of Physics, 21(8): 1003, August 1991. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00733224.

Pavsic:1991:IRQ

[1364] Matej Pavsic. On the interpretation of the relativistic quantum mechanics with invariant evolution parameter. Foundations of Physics, 21(9): 1005–1019, September 1991. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00733384.

Kummer:1991:FQTa

[1365] Hans Kummer. The foundation of quantum theory and noncommutative spectral theory. Part I. Foundations of Physics, 21(9):1021– 1069, September 1991. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00733385.

Phipps:1991:PTS

[1366] Thomas E. Phipps, Jr. Proper time synchronization. Foundations of Physics, 21(9):1071–1087, September 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00733386.

Ishigaki:1991:FME

[1367] Toshio Ishigaki. Four mathematical expressions of the uncertainty relation. Foundations of Physics, 21(9):1089–1105, September 1991. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00733387.

Wan:1991:LST

[1368] K. Kong Wan. Local spatial transformations and local observables. Foundations of Physics, 21(9):1107–1116, September 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00733388.

Fujita:1991:BEC

[1369] S. Fujita, T. Kimura, and Y. Zheng. On the Bose–Einstein condensation of free relativistic bosons with or without mass. Foundations of Physics, 21(9):1117–1130, September 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00733389.

Klir:1991:BR

[1370] George J. Klir. Book review. Foundations of Physics, 21(9):1131– 1132, September 1991. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00733390.

Anonymous:1991:FAb

[1371] Anonymous. First announcement. Foundations of Physics, 21(9): 1133, September 1991. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00733391.

Sonego:1991:IHF

[1372] Sebastiano Sonego. Interpretation of the hydrodynamical formalism of quantum mechanics. Foundations of Physics, 21(10):1135–1181, October 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00734264.

Kummer:1991:FQTb

[1373] Hans Kummer. The foundation of quantum theory and noncommutative spectral theory: Part II. Foundations of Physics, 21(10):1183– 1236, October 1991. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00734265.

Israelit:1991:CEP

[1374] Mark Israelit and Nathan Rosen. Classical elementary particles in general relativity. Foundations of Physics, 21(10):1237–1247, October 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00734266.

Stuewer:1991:BR

[1375] Roger H. Stuewer. Book review. Foundations of Physics, 21(10): 1249–1251, October 1991. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00734267.

Puthoff:1991:BR

[1376] Harold Puthoff. Book review. Foundations of Physics, 21(10):1253– 1259, October 1991. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00734268.

Austern:1991:BR

[1377] Norman Austern. Book review. Foundations of Physics, 21(10):1261– 1262, October 1991. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00734269.

Anonymous:1991:FAc

[1378] Anonymous. First announcement. Foundations of Physics, 21(10): 1263, October 1991. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00734270.

Anandan:1991:GAQ

[1379] J. Anandan. A geometric approach to quantum mechanics. Foundations of Physics, 21(11):1265–1284, November 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00732829.

Brandt:1991:CGS

[1380] Howard E. Brandt. Connections and geodesics in the spacetime tangent bundle. Foundations of Physics, 21(11):1285–1295, November 1991. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00732830.

Rosen:1991:SFQ

[1381] Joe Rosen. Symmetry at the foundations of quantum theory. Foundations of Physics, 21(11):1297–1304, November 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00732831.

Gur:1991:PSP

[1382] H. Gür. Phase-space path integration of the relativistic particle equations. Foundations of Physics, 21(11):1305–1314, November 1991. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00732832.

Ferrero:1991:QOP

[1383] Miguel Ferrero and T. W. Marshall. Quantum optical predictions in Q representation for Bell’s type experiments. Foundations of Physics, 21(11):1315–1321, November 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00732833.

Dariescu:1991:GTC

[1384] C. Dariescu and Marina Dariescu. U(1) gauge theory for charged bosonic fields on R × S^{3} topology. Foundations of Physics, 21(11): 1323–1327, November 1991. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00732834.

Dariescu:1991:GTQ

[1385] C. Dariescu and Marina Dariescu. U(1) gauge theory of the quantum hall effect. Foundations of Physics, 21(11):1329–1333, November 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00732835.

Jammer:1991:BR

[1386] Max Jammer. Book review. Foundations of Physics, 21(11):1335– 1338, November 1991. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00732836.

Greenberg:1991:BR

[1387] O. W. Greenberg. Book review. Foundations of Physics, 21(11):1339– 1341, November 1991. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00732837.

Earman:1991:BR

[1388] John Earman. Book review. Foundations of Physics, 21(11):1343– 1346, November 1991. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00732838.

Wald:1991:BR

[1389] Robert M. Wald. Book review. Foundations of Physics, 21(11):1347– 1349, November 1991. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00732839.

Lewin:1991:BR

[1390] Walter H. G. Lewin. Book review. Foundations of Physics, 21(11): 1351–1352, November 1991. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00732840.

Jefferys:1991:BR

[1391] William H. Jefferys. Book review. Foundations of Physics, 21(11): 1353–1355, November 1991. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00732841.

Jammer:1991:SKP

[1392] Max Jammer. Sir Karl Popper and his philosophy of physics. Foundations of Physics, 21(12):1357–1368, December 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01889648.

Miller:1991:KPC

[1393] David Miller. Karl Popper’s contributions to logic and the philosophy of science. Foundations of Physics, 21(12):1369–1374, December 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01889649.

Selleri:1991:KPN

[1394] Franco Selleri and Alwyn van der Merwe. Karl Popper at ninety: Highlights of a lifelong intellectual quest. Foundations of Physics, 21(12): 1375–1386, December 1991. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889650.

Mugur-Schachter:1991:SQP

[1395] Mioara Mugur-Schächter. Spacetime quantum probabilities, relativized descriptions, and Popperian propensities. Part I: Spacetime quantum probabilities. Foundations of Physics, 21(12):1387–1449, December 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01889651.

Stapp:1991:QPB

[1396] Henry P. Stapp. Quantum propensities and the brain-mind connection. Foundations of Physics, 21(12):1451–1477, December 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01889652.

Suppes:1991:EHV

[1397] Patrick Suppes and Mario Zanotti. Existence of hidden variables having only upper probabilities. Foundations of Physics, 21(12):1479– 1499, December 1991. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889653.

Miller:1991:SCP

[1398] David Miller. Single-case probabilities. Foundations of Physics, 21(12): 1501–1516, December 1991. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889654.

Anonymous:1991:PKP

[1399] Anonymous. Photo of Karl Popper. Foundations of Physics, 21 (12):ii, December 1991. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889647.

Cufaro-Petroni:1992:SPT

[1400] Nicola Cufaro-Petroni and Jean-Pierre Vigier. Single-particle trajectories and interferences in quantum mechanics. Foundations of Physics, 22 (1):1–40, January 1992. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883379.

Lowdin:1992:SAO

[1401] Per-Olov Löwdin. Some aspects of objectivity and reality in modern science. Foundations of Physics, 22(1):41–105, January 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:

//link.springer.com/article/10.1007/BF01883380. In honor of Sir

Karl Popper.

Gudder:1992:RS

[1402] Stanley Gudder. Realistic spin. Foundations of Physics, 22(1):107– 120, January 1992. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883381.

deBeauregard:1992:TRT

[1403] O. Costa de Beauregard. Is there a reality out there? Foundations of Physics, 22(1):121–135, January 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01883382.

Barut:1992:HAQ

[1404] A. O. Barut. How to avoid “quantum paradoxes”. Foundations of Physics, 22(1):137–142, January 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01883383.

Prugovecki:1992:RPI

[1405] Eduard Prugovecki. Realism, positivism, instrumentalism, and quantum geometry. Foundations of Physics, 22(2):143–186, February 1992. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01893610.

Woodward:1992:LDP

[1406] James F. Woodward. Logic of discovery or psychology of invention? Foundations of Physics, 22(2):187–203, February 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01893611.

Mercier:1992:PIW

[1407] André Mercier. The problem of the imperfection of a world, itself created by a perfect god. Foundations of Physics, 22(2):205–219, February 1992.

CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01893612.

Leggett:1992:NRC

[1408] A. J. Leggett. On the nature of research in condensed-state physics. Foundations of Physics, 22(2):221–233, February 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01893613.

Mugur-Schachter:1992:SQP

[1409] M. Mugur-Schächter. Spacetime quantum probabilities II: Relativized descriptions and Popperian propensities. Foundations of Physics, 22(2): 235–312, February 1992. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01893614.

Hajek:1992:E

[1410] Alan Hájek and Jeffrey Bub. Epr. Foundations of Physics, 22(3): 313–332, March 1992. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883898.

Ballentine:1992:COD

[1411] L. E. Ballentine. Can one detect the state of an individual system? Foundations of Physics, 22(3):333–342, March 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01883899.

Vinduska:1992:ITB

[1412] Milan Vinduska. Information-theoretic bell inequalities and the relative measure of probability. Foundations of Physics, 22(3):343–355, March 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF01883900.

Peres:1992:RDE

[1413] Asher Peres. Recursive definition for elements of reality. Foundations of Physics, 22(3):357–361, March 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01883901.

Marshall:1992:HPP

[1414] T. W. Marshall. A historical perspective to the present-day locality debate. Foundations of Physics, 22(3):363–370, March 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01883902.

Santos:1992:PMQ

[1415] Emilio Santos. The physical meaning of quantization. Foundations of Physics, 22(3):371–379, March 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01883903.

Redhead:1992:PCM

[1416] Michael Redhead. Propensities, correlations, and metaphysics. Foundations of Physics, 22(3):381–394, March 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01883904.

Treder:1992:CDU

[1417] Hans-Jürgen Treder. Continuum and discretum-unified field theory and elementary constants. Foundations of Physics, 22(3):395–420, March 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF01883905.

Horwitz:1992:DES

[1418] L. P. Horwitz. On the definition and evolution of states in relativistic classical and quantum mechanics. Foundations of Physics, 22(3): 421–450, March 1992. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883906.

Namiki:1992:MHS

[1419] Mikio Namiki and Saverio Pascazio. Many-Hilbert-spaces approach to the wave-function collapse. Foundations of Physics, 22(3):451– 466, March 1992. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883907.

Hoekzema:1992:CQP

[1420] Dick J. Hoekzema. Contextual quantum process theory. Foundations of Physics, 22(4):467–486, April 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00732918.

Hoekzema:1992:QET

[1421] Dick J. Hoekzema. Quantum event theory: a Tetrode–Fokker version of quantum field theory. Foundations of Physics, 22(4):487–506, April 1992.

CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00732919.

Vargas:1992:GPTa

[1422] José G. Vargas. Geometrization of the physics with teleparallelism. I. The classical interactions. Foundations of Physics, 22(4):507–526, April 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00732920.

Vargas:1992:GPTb

[1423] José G. Vargas, Douglas G. Torr, and Alvaro Lecompte. Geometrization of the physics with teleparallelism. II. Towards a fully geometric Dirac equation. Foundations of Physics, 22(4):527–547, April 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00732921.

Israelit:1992:SCP

[1424] Mark Israelit and Nathan Rosen. The static character of prematter particles. Foundations of Physics, 22(4):549–554, April 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00732922.

Israelit:1992:WDG

[1425] Mark Israelit and Nathan Rosen. Weyl–Dirac geometry and dark matter. Foundations of Physics, 22(4):555–568, April 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00732923.

Castro:1992:WGR

[1426] Carlos Castro. On Weyl geometry, random processes, and geometric quantum mechanics. Foundations of Physics, 22(4):569–615, April 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00732924.

Braun:1992:DMM

[1427] M. A. Braun and K. Urbanowski. On the description of multiple measurements of an unstable state. Foundations of Physics, 22(4):617–630, April 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00732925.

Milonni:1992:BR

[1428] Peter W. Milonni. Book review. Foundations of Physics, 22(4): 631–633, April 1992. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00732926.

Blumel:1992:BR

[1429] R. Blümel. Book review. Foundations of Physics, 22(4):635–637, April 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00732927.

Henneaux:1992:BR

[1430] Marc Henneaux. Book review. Foundations of Physics, 22(4): 639–640, April 1992. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00732928.

Rueger:1992:BR

[1431] Alexander Rueger. Book review. Foundations of Physics, 22(4): 641–643, April 1992. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00732929.

Finley:1992:BR

[1432] J. D. Finley. Book review. Foundations of Physics, 22(4):645–648, April 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00732930.

Gregory:1992:BR

[1433] Bruce Gregory. Book review. Foundations of Physics, 22(4):649– 651, April 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00732931.

Cohen:1992:HMP

[1434] Leon Cohen and James L. Park. Henry Margenau: Physicist– philosopher. Foundations of Physics, 22(5):653–656, May 1992. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01889670.

Park:1992:PMQ

[1435] James L. Park and William Band. Preparation and measurement in quantum physics. Foundations of Physics, 22(5):657–668, May 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01889671.

Klein:1992:LSN

[1436] Lewis Klein. Line shapes in nonlinear spectroscopy. Foundations of Physics, 22(5):669–690, May 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01889672.

Cohen:1992:MWF

[1437] Leon Cohen. Multipart wave functions. Foundations of Physics, 22 (5):691–711, May 1992. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889673.

Herman:1992:COA

[1438] R. M. Herman. Classical origins of the Aharonov–Bohm effect. Foundations of Physics, 22(5):713–725, May 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01889674.

Royer:1992:ETR

[1439] Antoine Royer. Ehrenfest’s theorem reinterpreted and extended with Wigner’s function. Foundations of Physics, 22(5):727–736, May 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01889675.

Bub:1992:QMP

[1440] Jeffrey Bub. Quantum mechanics without the projection postulate. Foundations of Physics, 22(5):737–754, May 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01889676.

Anonymous:1992:PHMa

[1441] Anonymous. Photo of Henry Margenau. Foundations of Physics, 22(5):ii, May 1992. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889669.

Prugovecki:1992:FPQ

[1442] Eduard Prugovecki. Foundational problems in quantum gravity and quantum cosmology. Foundations of Physics, 22(6):755–806, June 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01883743.

Hellman:1992:BTI

[1443] Geoffrey Hellman. Bell-type inequalities in the nonideal case: Proof of a conjecture of Bell. Foundations of Physics, 22(6):807–817, June 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01883744.

Peres:1992:ELR

[1444] Asher Peres. Emergence of local realism in fuzzy observations of correlated quantum systems. Foundations of Physics, 22(6):819–828, June 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF01883745.

Pauling:1992:VRQ

[1445] Linus Pauling. The value of rough quantum mechanical calculations. Foundations of Physics, 22(6):829–838, June 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01883746.

Gudder:1992:QSM

[1446] Stanley Gudder. Quantum stochastic models. Foundations of Physics, 22(6):839–852, June 1992. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883747.

Grandy:1992:RT

[1447] W. T. Grandy, Jr. On randomness and thermodynamics. Foundations of Physics, 22(6):853–866, June 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01883748.

Chandler:1992:QPD

[1448] C. Chandler, L. Cohen, C. Lee, M. Scully, and K. Wódkiewicz. Quasiprobability distribution for spin-1/2 particles. Foundations of Physics, 22(7):867–878, July 1992. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889685.

Khalfin:1992:QCC

[1449] Leonid A. Khalfin and Boris S. Tsirelson. Quantum/ classical correspondence in the light of Bell’s inequalities. Foundations of Physics, 22(7):879–948, July 1992. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889686.

Busch:1992:WOJ

[1450] P. Busch, P. Lahti, and P. Mittelstaedt. Weak objectification, joint probabilities, and Bell inequalities in quantum mechanics. Foundations of Physics, 22(7):949–962, July 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01889687.

Mugur-Schachter:1992:TFI

[1451] Mioara Mugur-Schächter. Toward a factually induced space–time quantum logic. Foundations of Physics, 22(7):963–994, July 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01889688.

Anonymous:1992:PHMb

[1452] Anonymous. Photo of Henry Margenau. Foundations of Physics, 22(7):iv, July 1992. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889684.

Lee:1992:WTG

[1453] Hai-Woong Lee. Wigner trajectories of a Gaussian wave packet perturbed by a weak potential. Foundations of Physics, 22(8):995–1010, August 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00733392.

Cooperstock:1992:ELG

[1454] F. I. Cooperstock. Energy localization in general relativity: a new hypothesis. Foundations of Physics, 22(8):1011–1024, August 1992. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00733393.

Arensburg:1992:FOE

[1455] A. Arensburg and L. P. Horwitz. A first-order equation for spin in a manifestly relativistically covariant quantum theory. Foundations of Physics, 22(8):1025–1039, August 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00733394.

Drechsler:1992:QFD

[1456] W. Drechsler. Quantized fiber dynamics for extended elementary objects involving gravitation. Foundations of Physics, 22(8):1041–1077, August 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00733395.

vonBorzeszkowski:1992:JFT

[1457] Horst-Heino von Borzeszkowski and Sisir Roy. Jordan–Fock type uncertainty relations and cut-off lengths in quantum general relativity. Foundations of Physics, 22(8):1079–1087, August 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00733396.

Treder:1992:CIM

[1458] Hans-Jürgen Treder. Conform invariance and Mach’s Principle in cosmology. Foundations of Physics, 22(8):1089–1093, August 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00733397.

vonBorzeszkowski:1992:BR

[1459] Horst-Heino von Borzeszkowski and Michel Bawin. Book review. Foundations of Physics, 22(8):1095–1101, August 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00733398.

Srensen:1992:BR

[1460] Roy Srensen. Book review. Foundations of Physics, 22(8):1103– 1109, August 1992. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00733399.

Donald:1992:PPL

[1461] Matthew J. Donald. A priori probability and localized observers. Foundations of Physics, 22(9):1111–1172, September 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00732696.

Mensky:1992:CQM

[1462] Michael B. Mensky. Continuous quantum measurements and the action uncertainty principle. Foundations of Physics, 22(9):1173– 1193, September 1992. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00732697.

Brandt:1992:BR

[1463] Howard E. Brandt. Book review. Foundations of Physics, 22(9):1195– 1200, September 1992. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00732698.

Bell:1992:SPW

[1464] J. S. Bell. Six possible worlds of quantum mechanics. Foundations of Physics, 22(10):1201–1215, October 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01889711.

Dewdney:1992:WPD

[1465] C. Dewdney, G. Horton, M. M. Lam, Z. Malik, and M. Schmidt. Waveparticle dualism and the interpretation of quantum mechanics. Foundations of Physics, 22(10):1217–1265, October 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01889712.

Barut:1992:PLC

[1466] A. O. Barut and A. J. Bracken. Particle-like configurations of the electromagnetic field: an extension of de Broglie’s ideas. Foundations of Physics, 22(10):1267–1285, October 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01889713.

Holland:1992:DEB

[1467] P. R. Holland. The Dirac equation in the de Broglie–Bohm theory of motion. Foundations of Physics, 22(10):1287–1301, October 1992. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01889714.

Combourieu:1992:KRP

[1468] Marie-Christine Combourieu. Karl R. Popper, 1992: about the EPR controversy. Foundations of Physics, 22(10):1303–1323, October 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01889715.

Anonymous:1992:EN

[1469] Anonymous. Editorial note. Foundations of Physics, 22(10):ii, October 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF01889710.

Anonymous:1992:PLB

[1470] Anonymous. Photo of Louis de Broglie. Foundations of Physics, 22 (10):i, October 1992. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01889709.

Bozic:1992:BPD

[1471] M. Bozić, Z. Marić, and J. P. Vigier. De Broglian probabilities in the double-slit experiment. Foundations of Physics, 22(11):1325– 1344, November 1992. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883664.

Kaloyerou:1992:WZD

[1472] P. N. Kaloyerou. On the Wootters–Zurek development of Einstein’s twoslit experiment. Foundations of Physics, 22(11):1345–1377, November 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF01883665.

Cufaro-Petroni:1992:SQM

[1473] Nicola Cufaro-Petroni. On the structure of the quantum-mechanical probability models. Foundations of Physics, 22(11):1379–1401, November 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1007/
BF01883666.

Combourieu:1992:WPD

[1474] Marie-Christine Combourieu and Helmut Rauch. The wave-particle dualism in 1992: a summary. Foundations of Physics, 22(12):1403– 1434, December 1992. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883732.

Ghose:1992:WPD

[1475] Partha Ghose and Dipankar Home. Wave-particle duality of singlephoton states. Foundations of Physics, 22(12):1435–1447, December 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF01883733.

Kyprianidis:1992:PSF

[1476] A. Kyprianidis. The principles of a stochastic formulation of quantum theory. Foundations of Physics, 22(12):1449–1483, December 1992. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01883734.

deBeauregard:1992:EGI

[1477] O. Costa de Beauregard. Electromagnetic gauge as an integration condition: de Broglie’s argument revisited and expanded. Foundations of Physics, 22(12):1485–1494, December 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01883735.

dEspagnat:1992:HPR

[1478] B. d’Espagnat. Heisenberg picture and reality. Foundations of Physics, 22(12):1495–1504, December 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01883736.

Squires:1992:BR

[1479] Euan J. Squires. Book review. Foundations of Physics, 22(12):1505– 1507, December 1992. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883737.

Garstang:1992:BR

[1480] R. H. Garstang. Book review. Foundations of Physics, 22(12):1509– 1511, December 1992. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883738.

Weenink:1992:BR

[1481] M. P. H. Weenink. Book review. Foundations of Physics, 22(12):1513– 1515, December 1992. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883739.

Holton:1992:BR

[1482] Gerald Holton. Book review. Foundations of Physics, 22(12):1517– 1520, December 1992. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883740.

Baranger:1992:BR

[1483] Michel Baranger. Book review. Foundations of Physics, 22(12):1521– 1523, December 1992. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883741.

Anonymous:1992:A

[1484] Anonymous. Announcement. Foundations of Physics, 22(12):1525– 1527, December 1992. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883742.

Vigier:1993:DNE

[1485] J. P. Vigier. From Descartes and Newton to Einstein and de Broglie. Foundations of Physics, 23(1):1–4, January 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01883986.

Holland:1993:DJB

[1486] P. R. Holland and J. P. Vigier. David Joseph Bohm: 1917–1992. Foundations of Physics, 23(1):5–6, January 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01883987.

Rauch:1993:RNI

[1487] H. Rauch. Reality in neutron interference experiments. Foundations of Physics, 23(1):7–36, January 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01883988.

Mugur-Schachter:1993:QMU

[1488] Mioara Mugur-Schächter. From quantum mechanics to universal structures of conceptualization and feedback on quantum mechanics. Foundations of Physics, 23(1):37–122, January 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01883989.

Lochak:1993:LBC

[1489] Georges Lochak. Louis de Broglie’s conception of physics. Foundations of Physics, 23(1):123–131, January 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01883990.

Rae:1993:BR

[1490] Alastair I. M. Rae. Book review. Foundations of Physics, 23(1): 133–135, January 1993. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883991.

vonBorzeszkowski:1993:BR

[1491] Horst-Heino von Borzeszkowski and Doren A. Recker. Book review. Foundations of Physics, 23(1):137–143, January 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01883992.

Martin:1993:BR

[1492] André Martin. Book review. Foundations of Physics, 23(1):145– 147, January 1993. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883993.

Bunge:1993:BR

[1493] Mario Bunge. Book review. Foundations of Physics, 23(1):149– 151, January 1993. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883994.

McKinney:1993:BR

[1494] William J. McKinney. Book review. Foundations of Physics, 23(1): 153–155, January 1993. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883995.

Hall:1993:BR

[1495] Richard L. Hall. Book review. Foundations of Physics, 23(1):157– 159, January 1993. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883996.

Ginocchio:1993:BR

[1496] Joseph Ginocchio. Book review. Foundations of Physics, 23(1):161– 163, January 1993. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883997.

Goodstein:1993:BR

[1497] David L. Goodstein. Book review. Foundations of Physics, 23(1): 165–167, January 1993. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883998.

Anonymous:1993:SFM

[1498] Anonymous. Symposium on the foundations of modern physics 1993. Foundations of Physics, 23(1):169–171, January 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01883999.

Inomata:1993:AOB

[1499] Akira Inomata, Ray Wilson, and Alwyn van der Merwe. To Asim Orhan Barut on his sixty-fifth birthday. Foundations of Physics, 23(2): 172–176, February 1993. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883621.

Kien:1993:GCS

[1500] Fam Le Kien, M. O. Scully, and H. Walther. Generation of a coherent state of the micromaser field. Foundations of Physics, 23(2): 177–184, February 1993. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883622.

Solem:1993:UGP

[1501] J. C. Solem and L. C. Biedenharn. Understanding geometrical phases in quantum mechanics: An elementary example. Foundations of Physics, 23 (2):185–195, February 1993. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883623.

Moshinsky:1993:BEP

[1502] M. Moshinsky and G. Loyola. Barut equation for the particle-antiparticle system with a Dirac oscillator interaction. Foundations of Physics, 23 (2):197–210, February 1993. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883624.

Adelman:1993:GSS

[1503] M. Adelman, J. V. Corbett, and C. A. Hurst. The geometry of state space. Foundations of Physics, 23(2):211–223, February 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01883625.

Budinich:1993:EEU

[1504] Paolo Budinich and Ryszard Raczka. Eigenvibrations of the expanding universe. Foundations of Physics, 23(2):225–237, February 1993. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01883626.

Taylor:1993:TE

[1505] J. G. Taylor. On theories of everything. Foundations of Physics, 23 (2):239–243, February 1993. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883627.

Mukunda:1993:MML

[1506] N. Mukunda. Mechanical models for Lorentz group representations. Foundations of Physics, 23(2):245–260, February 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01883628.

Drechsler:1993:CVQ

[1507] Wolfgang Drechsler. Classical versus quantum gravity. Foundations of Physics, 23(2):261–276, February 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01883629.

Bracken:1993:CQS

[1508] A. J. Bracken and R. J. B. Fawcett. Compact quantum systems and the Pauli data problem. Foundations of Physics, 23(2):277–289, February 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF01883630.

Sawada:1993:BSN

[1509] Tetsuo Sawada. Bound states of the nucleon-monopole system. Foundations of Physics, 23(2):291–311, February 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01883631.

McEwan:1993:CFG

[1510] J. McEwan. A complex formulation of generalized Hamiltonian (Birkhoffian) theory. Foundations of Physics, 23(2):313–327, February 1993. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01883632.

Xu:1993:CFT

[1511] Bo wei Xu. Conformal field theories and critical phenomena. Foundations of Physics, 23(2):329–339, February 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01883633.

Ghirardi:1993:PDO

[1512] G. C. Ghirardi, R. Grassi, J. Butterfield, and G. N. Fleming. Parameter dependence and outcome dependence in dynamical models for state vector reduction. Foundations of Physics, 23(3):341–364, March 1993. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01883717.

Hestenes:1993:ZM

[1513] David Hestenes. Zitterbewegung modeling. Foundations of Physics, 23 (3):365–387, March 1993. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883718.

Benedict:1993:CSQ

[1514] M. G. Benedict and W. Schleich. On the correspondence of semiclassical and quantum phases in cyclic evolutions. Foundations of Physics, 23 (3):389–397, March 1993. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883719.

Anders:1993:UBF

[1515] T. B. Anders, R. von Mellenthin, B. Pfeil, and H. Salecker. Unitarity bounds for 4-fermion contact interactions. Foundations of Physics, 23 (3):399–410, March 1993. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883720.

Moffat:1993:QGO

[1516] J. W. Moffat. Quantum gravity, the origin of time and time’s arrow. Foundations of Physics, 23(3):411–437, March 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01883721.

Grandy:1993:NM

[1517] W. T. Grandy, Jr. A nonstandard model. Foundations of Physics, 23 (3):439–460, March 1993. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883722.

OConnell:1993:DES

[1518] R. F. O’Connell. Does the electron have a structure? Foundations of Physics, 23(3):461–464, March 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01883723.

Wyss:1993:CSL

[1519] Daniel W. Wyss and Walter Wyss. Coincidental spectral lines for the hydrogen atom. Foundations of Physics, 23(3):465–468, March 1993.

CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01883724.

Rodrigues:1993:GDN

[1520] Waldyr A. Rodrigues, Jr., Jayme Vaz, Jr., and Erasmo Recami. A generalization of Dirac nonlinear electrodynamics, and spinning charged particles. Foundations of Physics, 23(3):469–485, March 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01883725.

Fanchi:1993:RIT

[1521] J. R. Fanchi. Review of invariant time formulations of relativistic quantum theories. Foundations of Physics, 23(3):487–548, March 1993. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01883726.

Fronsdal:1993:CFQ

[1522] Christian Fronsdal. Classical foundations of quantum groups. Foundations of Physics, 23(4):551–569, April 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01883765.

Flato:1993:QDS

[1523] M. Flato, L. K. Hadjiivanov, and I. T. Todorov. Quantum deformations of singletons and of free zero-mass fields. Foundations of Physics, 23 (4):571–586, April 1993. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883766.

Flato:1993:WDQ

[1524] Moshé Flato, Zhi-Cheng Lu, and Daniel Sternheimer. From where do quantum groups come? Foundations of Physics, 23(4):587–598, April 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF01883767.

Budinich:1993:GPC

[1525] P. Budinich and R. Raczka. Global properties of conformally flat momentum space and their implications. Foundations of Physics, 23(4): 599–615, April 1993. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883768.

Zhu:1993:SAH

[1526] Chengjun Zhu and John R. Klauder. The self-adjointness of Hermitian Hamiltonians. Foundations of Physics, 23(4):617–631, April 1993. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01883769.

Grundland:1993:GTS

[1527] A. M. Grundland, J. A. Tuszyński, and P. Winternitz. Group theory and solutions of classical field theories with polynomial nonlinearities. Foundations of Physics, 23(4):633–665, April 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01883770.

Tchrakian:1993:NGF

[1528] D. H. Tchrakian. New gauge field models with non-Yang–Mills dynamics. Foundations of Physics, 23(4):667–673, April 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01883771.

Cordero:1993:AMN

[1529] Patricio Cordero and Sebastián Salamó. Algebraic methods for the Natanzon potentials. Foundations of Physics, 23(4):675–690, April 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01883772.

Iwao:1993:REH

[1530] Syurei Iwao. Radial excitation of hadronic states with null Pontryagin index. Foundations of Physics, 23(4):691–702, April 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01883773.

ORaifeartaigh:1993:ABE

[1531] L. O’Raifeartaigh, N. Straumann, and A. Wipf. Aharonov–Bohm effect in the presence of superconductors. Foundations of Physics, 23 (5):703–709, May 1993. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883804.

Berrondo:1993:PEN

[1532] M. Berrondo and J. F. Van Huele. The pole expansion in normalized QED. Foundations of Physics, 23(5):711–719, May 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01883805.

Durr:1993:GEF

[1533] Detlef Dürr, Sheldon Goldstein, and Nino Zanghí. A global equilibrium as the foundation of quantum randomness. Foundations of Physics, 23 (5):721–738, May 1993. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883806.

Takahashi:1993:SLT

[1534] Yasushi Takahashi. Surfaces and lines in terms of the delta and step functions. Foundations of Physics, 23(5):739–750, May 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01883807.

Bohm:1993:RND

[1535] A. Bohm. Relativistic and nonrelativistic dynamical groups. Foundations of Physics, 23(5):751–767, May 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01883808.

Kleinert:1993:GTO

[1536] H. Kleinert. Group theory and orbital fluctuations of the hydrogen atom. Foundations of Physics, 23(5):769–807, May 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01883809.

Duru:1993:CFB

[1537] I. H. Duru. Casimir force between two Aharonov–Bohm solenoids. Foundations of Physics, 23(5):809–818, May 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01883810.

Bozic:1993:SCM

[1538] M. Bozić and Z. Marić. A search for the classical model of spin. Foundations of Physics, 23(5):819–835, May 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01883811.

Aydm:1993:TBD

[1539] Z. Z. Aydm and A. U. Yilmazer. Two-body Dirac equation versus KDP equation. Foundations of Physics, 23(5):837–840, May 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01883812.

Salamin:1993:SEQ

[1540] Yousef I. Salamin. Self-energy quantum electrodynamics: Multipole radiation. Foundations of Physics, 23(5):841–849, May 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01883813.

Beker:1993:CES

[1541] Haluk Beker. Classification of exactly solvable potential problems. Foundations of Physics, 23(5):851–856, May 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01883814.

Zaugg:1993:BAM

[1542] T. Zaugg, M. Wilkens, and P. Meystre. Back-action in the measurement of “macroscopic quantum superpositions” in microwave cavities. Foundations of Physics, 23(6):857–871, June 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01891509.

Gilmore:1993:ADQ

[1543] R. Gilmore, H. G. Solari, and S. K. Kim. Algebraic description of the quantum defect. Foundations of Physics, 23(6):873–879, June 1993. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01891510.

Malin:1993:CQS

[1544] Shimon Malin. The collapse of quantum states: a new interpretation. Foundations of Physics, 23(6):881–893, June 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01891511.

Dowling:1993:SEC

[1545] Jonathan P. Dowling. Spontaneous emission in cavities: How much more classical can you get? Foundations of Physics, 23(6):895–905, June 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01891512.

Salamin:1993:BSF

[1546] Yousef I. Salamin. Bremsstrahlung in self-field QED. Foundations of Physics, 23(6):907–912, June 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01891513.

Collins:1993:TQU

[1547] Peter D. B. Collins and Euan J. Squires. Time in a quantum universe. Foundations of Physics, 23(6):913–921, June 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01891514.

Dreitlein:1993:UEC

[1548] Joseph Dreitlein. The unreasonable effectiveness of computer physics. Foundations of Physics, 23(6):923–930, June 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01891515.

Kundt:1993:MEP

[1549] Wolfgang Kundt. The most energetic processes in the universe. Foundations of Physics, 23(6):931–948, June 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01891516.

Budinich:1993:CCS

[1550] P. Budinich. Conformal compacifications from spinor geometry. Foundations of Physics, 23(6):949–963, June 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01891517.

Burt:1993:BCQ

[1551] Philip B. Burt. Boundary conditions in quantum field theories. Foundations of Physics, 23(6):965–968, June 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01891518.

Inomata:1993:TCQa

[1552] Akira Inomata, Georg Junker, and Raj Wilson. Topological charge quantization via path integration: An application of the Kustaanheimo–Stiefel transformation. Foundations of Physics, 23(6):969, June 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01891519.

Anonymous:1993:PB

[1553] Anonymous. Photo of a o bonus. Foundations of Physics, 23(6):i, June 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF01891508.

Elby:1993:WMT

[1554] Andrew Elby, Harvey R. Brown, and Sara Foster. What makes a theory physically “complete”? Foundations of Physics, 23(7):971–985, July 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00736011.

Elitzur:1993:QMI

[1555] Avshalom C. Elitzur and Lev Vaidman. Quantum mechanical interaction-free measurements. Foundations of Physics, 23(7):987–997, July 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1007/
BF00736012.

Hudson:1993:SLT

[1556] Robin L. Hudson and Sylvia Pulmannová. Sum logics and tensor products. Foundations of Physics, 23(7):999–1024, July 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00736013.

Bleyer:1993:ELH

[1557] Ulrich Bleyer. Energy levels of the hydrogen atom due to a generalized Dirac equation. Foundations of Physics, 23(7):1025–1048, July 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00736014.

Selleri:1993:BR

[1558] Franco Selleri and GianCarlo Ghirardi. Book review. Foundations of Physics, 23(7):1049–1055, July 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00736015.

MacKinnon:1993:BR

[1559] Edward M. MacKinnon. Book review. Foundations of Physics, 23(7): 1057–1059, July 1993. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00736016.

Gesztesy:1993:BR

[1560] F. Gesztesy. Book review. Foundations of Physics, 23(7):1061–1063, July 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00736017.

Devaney:1993:BR

[1561] Robert L. Devaney and Geoffrey Dixon. Book review. Foundations of Physics, 23(7):1065–1069, July 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00736018.

Inomata:1993:TCQb

[1562] Akira Inomata, Georg Junker, and Raj Wilson. Topological charge quantization via path integration: An application of the Kustaanheimo–Stiefel transformation. Foundations of Physics, 23(8):1073–1091, August 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00732414.

Parrott:1993:UPS

[1563] Stephen Parrott. Unphysical and physical(?) solutions of the Lorentz– Dirac equation. Foundations of Physics, 23(8):1093–1119, August 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00732415.

Comay:1993:RPM

[1564] E. Comay. Remarks on the physical meaning of the Lorentz–Dirac equation. Foundations of Physics, 23(8):1121–1136, August 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00732416.

Hemmo:1993:BR

[1565] Meir Hemmo. Book review. Foundations of Physics, 23(8):1137– 1145, August 1993. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00732417.

Goldberg:1993:BRa

[1566] Stanley Goldberg and Robert March. Book review. Foundations of Physics, 23(8):1147–1159, August 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00732418.

Eichhorn:1993:BR

[1567] Heinrich H. Eichhorn. Book review. Foundations of Physics, 23(8): 1161–1163, August 1993. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00732419.

Sevrin:1993:BR

[1568] Alexander Sevrin. Book review. Foundations of Physics, 23(8):1165– 1167, August 1993. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00732420.

Goldberg:1993:BRb

[1569] Joshua N. Goldberg. Book review. Foundations of Physics, 23(8): 1169–1171, August 1993. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF00732421.

Anonymous:1993:FFP

[1570] Anonymous. Frontiers of fundamental physics international conference. Foundations of Physics, 23(8):1173, August 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00732422.

Gull:1993:INR

[1571] Stephen Gull, Anthony Lasenby, and Chris Doran. Imaginary numbers are not real — the geometric algebra of spacetime. Foundations of Physics, 23(9):1175–1201, September 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01883676.

Lounesto:1993:CAH

[1572] Pertti Lounesto. Clifford algebras and Hestenes spinors. Foundations of Physics, 23(9):1203–1237, September 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01883677.

Doran:1993:SOS

[1573] Chris Doran, Anthony Lasenby, and Stephen Gull. States and operators in the spacetime algebra. Foundations of Physics, 23(9):1239– 1264, September 1993. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883678.

Kruger:1993:CLR

[1574] Heinz Krüger. Classical limit of real Dirac theory: Quantization of relativistic central field orbits. Foundations of Physics, 23(9):1265– 1288, September 1993. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883679.

Sobczyk:1993:DHE

[1575] Garret Sobczyk. David Hestenes: The early years. Foundations of Physics, 23(10):1290–1293, October 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01883780.

Lasenby:1993:MDA

[1576] Anthony Lasenby, Chris Doran, and Stephen Gull. A multivector derivative approach to Lagrangian field theory. Foundations of Physics, 23(10): 1295–1327, October 1993. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01883781.

Gull:1993:EPT

[1577] Stephen Gull, Anthony Lasenby, and Chris Doran. Electron paths, tunnelling, and diffraction in the spacetime algebra. Foundations of Physics, 23(10):1329–1356, October 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01883782.

Dress:1993:DGG

[1578] Andreas W. M. Dress and Timothy F. Havel. Distance geometry and geometric algebra. Foundations of Physics, 23(10):1357–1374, October 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF01883783.

Sobczyk:1993:LTU

[1579] Garret Sobczyk. Linear transformations in unitary geometric algebra. Foundations of Physics, 23(10):1375–1385, October 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01883784.

Boudet:1993:RCS

[1580] Roger Boudet. On the relativistic calculation of spontaneous emission. Foundations of Physics, 23(10):1387–1404, October 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01883785.

Jancewicz:1993:HSC

[1581] Bernard Jancewicz. A Hilbert space for the classical electromagnetic field. Foundations of Physics, 23(11):1405–1421, November 1993. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01243938.

Barut:1993:PBS

[1582] A. O. Barut and N. Ünal. On Poisson brackets and symplectic structures for the classical and quantum Zitterbewegung. Foundations of Physics, 23(11):1423–1429, November 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01243939.

Daviau:1993:LND

[1583] C. Daviau. Linear and nonlinear Dirac equation. Foundations of Physics, 23(11):1431–1443, November 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01243940.

Vroegindeweij:1993:EDH

[1584] P. G. Vroegindeweij. The equations of Dirac and the M_{2} (h)- representation of cl_{1,3}. Foundations of Physics, 23(11):1445–1463, November 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01243941.

Rodrigues:1993:CBN

[1585] Waldyr A. Rodrigues, Jr. and Quintino A. G. de Souza. The Clifford bundle and the nature of the gravitational field. Foundations of Physics, 23(11):1465–1490, November 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01243942.

Sommen:1993:IDO

[1586] F. Sommen and N. Van Acker. So ( m )-invariant differential operators on Clifford algebra-valued functions. Foundations of Physics, 23(11): 1491–1519, November 1993. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01243943.

Roberts:1993:VVR

[1587] D. E. Roberts. Vector-valued rational forms. Foundations of Physics, 23(11):1521–1533, November 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01243944.

Rodriguez-Romo:1993:AFI

[1588] Suemi Rodríguez-Romo. An analysis of Fierz identities, factorization and inversion theorems. Foundations of Physics, 23(11):1535– 1542, November 1993. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF01243945.

Semenov:1993:IHD

[1589] Pavel Semenov. On invertibility in high-dimensional Clifford algebras. Foundations of Physics, 23(11):1543–1546, November 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01243946.

Schroder:1993:HJT

[1590] Wilfried Schröder. Hans–Jürgen treder at 65. Foundations of Physics, 23(12):1549–1551, December 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00732364.

Saunders:1993:DRS

[1591] Simon Saunders. Decoherence, relative states, and evolutionary adaptation. Foundations of Physics, 23(12):1553–1585, December 1993. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00732365.

Beil:1993:ECC

[1592] R. G. Beil. The extended classical charged particle. II. Foundations of Physics, 23(12):1587–1600, December 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00732366.

Griffiths:1993:CCH

[1593] Robert B. Griffiths. The consistency of consistent histories: a reply to d’Espagnat. Foundations of Physics, 23(12):1601–1610, December 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00732367.

Page:1993:BR

[1594] Don Page, Richard Healey, and Richard Matzner. Book review. Foundations of Physics, 23(12):1611–1619, December 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00732368.

Folse:1993:BR

[1595] Henry J. Folse, Jr. and B. Greene. Book review. Foundations of Physics, 23(12):1621–1629, December 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF00732369.

Scholl:1993:BR

[1596] Eckehard Schöll and J. W. Darewych. Book review. Foundations of Physics, 23(12):1631–1634, December 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00732370.

Anonymous:1994:IA

[1597] Anonymous. Important announcement. Foundations of Physics, 24 (1):1, January 1994. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02053905.

Lam:1994:BACa

[1598] M. M. Lam and C. Dewdney. The Bohm approach to cavity quantum scalar field dynamics. Part I: The free field. Foundations of Physics, 24(1):3–27, January 1994. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02053906.

Lam:1994:BACb

[1599] M. M. Lam and C. Dewdney. The Bohm approach to cavity quantum scalar field dynamics. Part II: The interaction of the field with matter. Foundations of Physics, 24(1):29–60, January 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02053907.

Vigier:1994:PTR

[1600] J. P. Vigier. Possible test of the reality of superluminal phase waves and particle phase space motions in the Einstein–de Broglie–Bohm causal stochastic interpretation of quantum mechanics. Foundations of Physics, 24(1):61–83, January 1994. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02053908.

Scully:1994:SQD

[1601] M. O. Scully and K. Wódkiewicz. Spin quasi-distribution functions. Foundations of Physics, 24(1):85–107, January 1994. CODEN FNDPA4.

ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF02053909.

Hudson:1994:GTP

[1602] Robin Hudson and Sylvia Pulmannová. Graded tensor products of quantum logics. Foundations of Physics, 24(1):109–116, January 1994. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02053910.

Herbut:1994:DCE

[1603] Fedor Herbut. Delayed-choice experiments and retroactive apparent occurrence in the quantum theory of measurement. Foundations of Physics, 24(1):117–137, January 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02053911.

Hnilo:1994:TOL

[1604] A. A. Hnilo. On testing objective local theories by using Greenberger– Horne–Zeilinger states. Foundations of Physics, 24(1):139–162, January 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF02053912.

Yan:1994:TEL

[1605] C. C. Yan. Theoretical experimentation with the law of Biot– Savart. Foundations of Physics, 24(1):163–175, January 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02053913.

Costantini:1994:MCD

[1606] D. Costantini and U. Garibaldi. Microcanonical and canonical distributions and finite exchangeable random processes. Foundations of Physics, 24(1):177–202, January 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02053914.

Anonymous:1994:Aa

[1607] Anonymous. Announcement. Foundations of Physics, 24(1):203– 204, January 1994. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02053915.

Anonymous:1994:SFMa

[1608] Anonymous. Symposium on the foundations of modern physics 1994 13–16 June 1994, Helsinki, Finland. Foundations of Physics, 24(1):

205–206, January 1994. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02053916.

Anonymous:1994:Ab

[1609] Anonymous. Announcement. Foundations of Physics, 24(1):207, January 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF02053917.

Jammer:1994:FRH

[1610] Max Jammer. Fritz Rohrlich and his work — on the occasion of his retirement. Foundations of Physics, 24(2):209–216, February 1994. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02313122.

Haller:1994:QGE

[1611] K. Haller and E. Lim-Lombridas. Quantum gauge equivalence in QED. Foundations of Physics, 24(2):217–247, February 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02313123.

Grotch:1994:STH

[1612] Howard Grotch. Status of the theory of the hydrogen Lamb shift. Foundations of Physics, 24(2):249–272, February 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02313124.

Owen:1994:QET

[1613] David A. Owen. On quantum electrodynamics of two-particle bound states containing spinless particles. Foundations of Physics, 24(2): 273–296, February 1994. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02313125.

Crater:1994:SQM

[1614] H. W. Crater and P. Van Alstine. Structure of quantum-mechanical relativistic two-body interactions for spinning particles. Foundations of Physics, 24(2):297–328, February 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02313126.

Anonymous:1994:SFMb

[1615] Anonymous. Symposium on the foundations of modern physics 1994. Foundations of Physics, 24(2):329–330, February 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02313127.

Anonymous:1994:PACa

[1616] Anonymous. Preliminary announcement and call for papers: International Workshop on Quantum Communications and Measurement. Foundations of Physics, 24(2):331, February 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02313128.

Prugovecki:1994:FGC

[1617] Eduard Prugovecki. On foundational and geometric critical aspects of quantum electrodynamics. Foundations of Physics, 24(3):335– 362, March 1994. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02058096.

Nakano:1994:SRD

[1618] Akihiro Nakano and Philip Pearle. Statevector reduction in discrete time: a random walk in Hilbert space. Foundations of Physics, 24 (3):363–377, March 1994. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02058097.

Popescu:1994:QNA

[1619] Sandu Popescu and Daniel Rohrlich. Quantum nonlocality as an axiom. Foundations of Physics, 24(3):379–385, March 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02058098.

Coester:1994:VSH

[1620] F. Coester and W. Polyzou. Vacuum structures in Hamiltonian lightfront dynamics. Foundations of Physics, 24(3):387–400, March 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02058099.

Biedenharn:1994:ESW

[1621] L. C. Biedenharn and L. P. Horwitz. On the equivalence of the Skyrme– Witten model and large-N_{c} quark models. Foundations of Physics, 24 (3):401–417, March 1994. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02058100.

Islam:1994:CSN

[1622] M. M. Islam. Composite structure of the nucleon. Foundations of Physics, 24(3):419–425, March 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02058101.

Varadarajan:1994:BR

[1623] V. S. Varadarajan. Book review. Foundations of Physics, 24(3): 427–430, March 1994. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02058102.

Trimble:1994:BR

[1624] Virginia Trimble. Book review. Foundations of Physics, 24(3): 431–433, March 1994. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02058103.

Padmanabhan:1994:BR

[1625] T. Padmanabhan. Book review. Foundations of Physics, 24(3): 435–437, March 1994. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02058104.

Abarbanel:1994:BR

[1626] Henry D. I. Abarbanel. Book review. Foundations of Physics, 24(3): 439–442, March 1994. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02058105.

Wolfenstein:1994:BR

[1627] L. Wolfenstein. Book review. Foundations of Physics, 24(3):443– 445, March 1994. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02058106.

Landsberg:1994:BR

[1628] Peter T. Landsberg. Book review. Foundations of Physics, 24(3): 447–448, March 1994. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02058107.

Korepin:1994:BR

[1629] Vladimir Korepin. Book review. Foundations of Physics, 24(3): 449, March 1994. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02058108.

Anonymous:1994:PACb

[1630] Anonymous. Preliminary announcement and call for papers: International Workshop on Quantum Communications and Measurement. Foundations of Physics, 24(3):451, March 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02058109.

Balachandran:1994:TPP

[1631] A. P. Balachandran. Topology in physics — a perspective. Foundations of Physics, 24(4):455–466, April 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02058057.

Goldberg:1994:SDM

[1632] Joshua N. Goldberg. Self-dual Maxwell field on a null surface. II. Foundations of Physics, 24(4):467–476, April 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02058058.

Barut:1994:FES

[1633] A. O. Barut. On the formulation of electrodynamics from a single principle. Foundations of Physics, 24(4):477–485, April 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02058059.

Mannheim:1994:OQC

[1634] Philip D. Mannheim. Open questions in classical gravity. Foundations of Physics, 24(4):487–511, April 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02058060.

Biswas:1994:SRN

[1635] Tarun Biswas. Special relativistic Newtonian gravity. Foundations of Physics, 24(4):513–524, April 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02058061.

Droz-Vincent:1994:MBS

[1636] Philippe Droz-Vincent. The massive bosonic string in the center-of-mass gauge. Foundations of Physics, 24(4):525–541, April 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02058062.

Fanchi:1994:EVP

[1637] John R. Fanchi. Evaluating the validity of parametrized relativistic wave equations. Foundations of Physics, 24(4):543–562, April 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02058063.

Land:1994:SRD

[1638] M. C. Land, R. I. Arshansky, and L. P. Horwitz. Selection rules for dipole radiation from a relativistic bound state. Foundations of Physics, 24(4):563–578, April 1994. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02058064.

Anonymous:1994:LPF

[1639] Anonymous. List of publications by F. Rohrlich. Foundations of Physics, 24(4):579–587, April 1994. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02058065.

Anonymous:1994:SFMc

[1640] Anonymous. Symposium on the foundations of modern physics 1994. Foundations of Physics, 24(4):589–590, April 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF02058066.

Islam:1994:SEQ

[1641] Jamal Nazrul Islam. The Schrödinger equation in quantum field theory. Foundations of Physics, 24(5):593–630, May 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF02054667.

Cattaneo:1994:AUQ

[1642] Gianpiero Cattaneo and Federico Laudisa. Axiomatic unsharp quantum theory (from Mackey to Ludwig and Piron). Foundations of Physics, 24(5):631–683, May 1994. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02054668.

Belavkin:1994:NPQ

[1643] V. P. Belavkin. Nondemolition principle of quantum measurement theory. Foundations of Physics, 24(5):685–714, May 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF02054669.

Wan:1994:USP

[1644] K. Kong Wan and R. G. McLean. Unitary state preparation, local position measurements, and spin in quantum mechanics. Foundations of Physics, 24(5):715–737, May 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02054670.

Kummer:1994:HVB

[1645] H. Kummer and R. G. McLean. Hidden variables and Bell’s theorem in quantum mechanics. Foundations of Physics, 24(5):739–751, May 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02054671.

delaPena:1994:WPM

[1646] L. de la Peña and A. M. Cetto. The wave properties of matter and the zeropoint radiation field. Foundations of Physics, 24(5):753–781, May 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF02054672.

daCosta:1994:TQG

[1647] N. C. A. da Costa, F. A. Doria, A. F. Furtado do Amaral, and J. A. de Barros. Two questions on the geometry of gauge fields. Foundations of Physics, 24(5):783–800, May 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02054673.

Sheldon:1994:BR

[1648] Eric Sheldon. Book review. Foundations of Physics, 24(5):801–806, May 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF02054674.

Brown:1994:BR

[1649] H. R. Brown. Book review. Foundations of Physics, 24(5):807–810, May 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF02054675.

McGuire:1994:BR

[1650] J. B. McGuire. Book review. Foundations of Physics, 24(5):811–812, May 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF02054676.

Ritter:1994:BR

[1651] Rogers C. Ritter. Book review. Foundations of Physics, 24(5): 813–816, May 1994. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02054677.

Dubois:1994:BR

[1652] Robert Dubois. Book review. Foundations of Physics, 24(5):817– 819, May 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02054678.

Rabinowitz:1994:BR

[1653] Mario Rabinowitz. Book review. Foundations of Physics, 24(5): 821–823, May 1994. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02054679.

Meir:1994:BR

[1654] A. J. Meir. Book review. Foundations of Physics, 24(5):825–828, May 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF02054680.

Wan:1994:UMS

[1655] K. K. Wan and F. E. Harrison. Unitary models of single detector triggering and local position measurements. Foundations of Physics, 24 (6):831–853, June 1994. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02067650.

Baumgartner:1994:PTE

[1656] B. Baumgartner. Postulates for time evolution in quantum mechanics. Foundations of Physics, 24(6):855–872, June 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF02067651.

Yeh:1994:CBC

[1657] Leehwa Yeh and Y. S. Kim. Correspondence between the classical and quantum canonical transformation groups from an operator formulation of the Wigner function. Foundations of Physics, 24(6):873–884, June 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF02067652.

Capria:1994:MME

[1658] Marco Mamone Capria and Fernanda Pambianco. On the Michelson– Morley experiment. Foundations of Physics, 24(6):885–899, June 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02067653.

Israelit:1994:WDM

[1659] Mark Israelit and Nathan Rosen. Weylian dark matter and cosmology. Foundations of Physics, 24(6):901–915, June 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF02067654.

delaPena:1994:QPZ

[1660] L. de la Peña and A. M. Cetto. Quantum phenomena and the zeropoint radiation field. Foundations of Physics, 24(6):917–948, June 1994. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02067655.

vBorzeszkowski:1994:EEF

[1661] H.-H. v. Borzeszkowski and H.-J. Treder. Einstein equations and Fierz– Pauli equations with self-interaction in quantum gravity. Foundations of Physics, 24(6):949–962, June 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02067656.

Greenberger:1994:BR

[1662] Daniel Greenberger. Book review. Foundations of Physics, 24(6): 963–965, June 1994. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02067657.

Schieve:1994:BR

[1663] William C. Schieve. Book review. Foundations of Physics, 24(6): 967–969, June 1994. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02067658.

Blood:1994:BR

[1664] Casey Blood. Book review. Foundations of Physics, 24(6):971–974, June 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF02067659.

Franklin:1994:BR

[1665] Allan Franklin. Book review. Foundations of Physics, 24(6):975– 978, June 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02067660.

Price:1994:BR

[1666] David Long Price. Book review. Foundations of Physics, 24(6): 979–980, June 1994. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02067661.

Stedman:1994:BR

[1667] Geoff Stedman. Book review. Foundations of Physics, 24(6):981– 982, June 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02067662.

Hara:1994:BR

[1668] Takashi Hara. Book review. Foundations of Physics, 24(6):983–985, June 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF02067663.

Prugovecki:1994:GCS

[1669] Eduard Prugovecki. On the general covariance and strong equivalence principles in quantum general relativity. Foundations of Physics, 24 (7):989–1076, July 1994. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02054648.

vonBorzeszkowski:1994:QCE

[1670] H.-H. von Borzeszkowski and H.-J. Treder. Quantum cosmology and Eddington’s large cosmic numbers. Foundations of Physics, 24(7): 1077–1088, July 1994. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02054649.

Stulpe:1994:SRC

[1671] Werner Stulpe. Some remarks on classical representations of quantum mechanics. Foundations of Physics, 24(7):1089–1094, July 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02054650.

MacCallum:1994:BR

[1672] David MacCallum. Book review. Foundations of Physics, 24(7): 1095–1097, July 1994. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02054651.

Efinger:1994:BR

[1673] H. J. Efinger. Book review. Foundations of Physics, 24(7):1099– 1101, July 1994. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02054652.

Putnam:1994:BR

[1674] Hilary Putnam. Book review. Foundations of Physics, 24(7):1103– 1104, July 1994. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02054653.

Aerts:1994:CPS

[1675] Diederik Aerts. Constantin Piron at sixty-plus: Continuing a quest for the understanding of fundamental physical theories and the pursuit of their elaboration. Foundations of Physics, 24(8):1107–1111, August 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02057858.

Mielnik:1994:SP

[1676] Bogdan Mielnik. The screen problem. Foundations of Physics, 24(8): 1113–1129, August 1994. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02057859.

Cattaneo:1994:CQV

[1677] Gianpiero Cattaneo and Giuseppe Nisticò. Coexistence of questions versus Piron’s compatibility. Foundations of Physics, 24(8):1131– 1152, August 1994. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02057860.

Beltrametti:1994:BTI

[1678] Enrico G. Beltrametti and Maciej J. Maczynski. On Bell-type inequalities. Foundations of Physics, 24(8):1153–1159, August 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02057861.

DallaChiara:1994:PUQ

[1679] M. L. Dalla Chiara and R. Giuntini. Partial and unsharp quantum logics. Foundations of Physics, 24(8):1161–1177, August 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02057862.

Tasaki:1994:MTL

[1680] S. Tasaki, E. Eisenberg, and L. P. Horwitz. Measurement theory in the Lax–Phillips formalism. Foundations of Physics, 24(8):1179– 1194, August 1994. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02057863.

Dvurecenskij:1994:BTI

[1681] Anatolij Dvurecenskij and Helmut Länger. Bell-type inequalities in horizontal sums of Boolean algebras. Foundations of Physics, 24(8): 1195–1202, August 1994. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02057864.

Gudder:1994:TRQ

[1682] Stanley Gudder. Toward a rigorous quantum field theory. Foundations of Physics, 24(9):1205–1225, September 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02148565.

Aerts:1994:QSS

[1683] Diederik Aerts. Quantum structures, separated physical entities and probability. Foundations of Physics, 24(9):1227–1259, September 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02148566.

Bub:1994:SQP

[1684] Jeffrey Bub. On the structure of quantal proposition systems. Foundations of Physics, 24(9):1261–1279, September 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF02148567.

Enz:1994:HDQ

[1685] Charles P. Enz. Hamiltonian description and quantization of dissipative systems. Foundations of Physics, 24(9):1281–1292, September 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02148568.

Cattaneo:1994:ISP

[1686] Gianpiero Cattaneo, Tiziana Marsico, and Giuseppe NisticÒ. Individual samples, preparations, and states in Piron’s approach to quantum physics. Foundations of Physics, 24(9):1293–1303, September 1994. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02148569.

Pawlowski:1994:UCM

[1687] Marek Pawlowski and Ryszard Raczka. A unified conformal model for fundamental interactions without dynamical Higgs field. Foundations of Physics, 24(9):1305–1327, September 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02148570.

Foulis:1994:EAU

[1688] D. J. Foulis and M. K. Bennett. Effect algebras and unsharp quantum logics. Foundations of Physics, 24(10):1331–1352, October 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02283036.

Aerts:1994:QCI

[1689] Diederik Aerts and Thomas Durt. Quantum, classical and intermediate: an illustrative example. Foundations of Physics, 24(10):1353– 1369, October 1994. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02283037.

Martin:1994:SPR

[1690] Ph. A. Martin and M. Sassoli de Bianchi. Spin precession revisited. Foundations of Physics, 24(10):1371–1378, October 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF02283038.

Oziewicz:1994:CFT

[1691] Zbigniew Oziewicz. Classical field theory and analogy between Newton’s and Maxwell’s equations. Foundations of Physics, 24(10):1379– 1402, October 1994. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02283039.

Pulmannova:1994:QLH

[1692] Sylvia Pulmannová. Quantum logics and Hilbert space. Foundations of Physics, 24(10):1403–1414, October 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02283040.

Enderle:1994:ECQ

[1693] M. Enderle and H. Neumann. Embedding of the classical into the quantum description of photons. Foundations of Physics, 24(10):1415– 1424, October 1994. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02283041.

Mayet:1994:NOO

[1694] R. Mayet and S. Pulmannová. Nearly orthosymmetric ortholattices and Hilbert spaces. Foundations of Physics, 24(10):1425–1437, October 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02283042.

Antoniou:1994:LQS

[1695] I. Antoniou and Z. Suchanecki. The logic of quantum systems with diagonal singularities. Foundations of Physics, 24(10):1439–1457, October 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF02283043.

Barut:1994:CST

[1696] A. O. Barut, P. Budinich, J. Niederle, and R. Raçzka. Conformal space– times — the arenas of physics and cosmology. Foundations of Physics, 24(11):1461–1494, November 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02054779.

Pavsic:1994:EMI

[1697] Matej Pavsic. The embedding model of induced gravity with bosonic sources. Foundations of Physics, 24(11):1495–1518, November 1994. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02054780.

Evans:1994:MFR

[1698] M. W. Evans. The magnetic fields and rotation generators of free space electromagnetism. Foundations of Physics, 24(11):1519–1542, November 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF02054781.

Padmanabhan:1994:PIR

[1699] T. Padmanabhan. Path integral for the relativistic particle and harmonic oscillators. Foundations of Physics, 24(11):1543–1562, November 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02054782.

Rosen:1994:CKG

[1700] Nathan Rosen. A classical Klein–Gordon particle. Foundations of Physics, 24(11):1563–1569, November 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02054783.

Barut:1994:DFT

[1701] A. O. Barut. Debating the final theory. Foundations of Physics, 24(11): 1571–1576, November 1994. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02054784.

Dariescu:1994:GTB

[1702] Ciprian Dariescu and Marina-Aura Dariescu. SU(2) × U(1) gauge theory of bosonic and fermionic fields in S^{3} × R space–time. Foundations of Physics, 24(11):1577–1582, November 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02054785.

Caves:1994:BR

[1703] Carlton M. Caves. Book review. Foundations of Physics, 24(11):1583– 1585, November 1994. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02054786.

deMuynck:1994:IQM

[1704] W. M. de Muynck, W. De Baere, and H. Martens. Interpretations of quantum mechanics, joint measurement of incompatible observables, and counterfactual definiteness. Foundations of Physics, 24(12):1589– 1664, December 1994. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.

1007/BF02054787. See comments [1705].

Stapp:1994:CIQ

[1705] Henry P. Stapp. Comments on “Interpretations of quantum mechanics, joint measurement of incompatible observables, and counterfactual definiteness”. Foundations of Physics, 24(12):1665–1669, December 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF02054788. See

[1704].

Evans:1994:CRT

[1706] M. W. Evans. Classical relativistic theory of the longitudinal ghost fields of electromagnetism. Foundations of Physics, 24(12):1671– 1688, December 1994. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02054789.

Rosen:1994:CPP

[1707] N. Rosen. A classical Proca particle. Foundations of Physics, 24(12): 1689–1695, December 1994. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02054790.

Halpern:1994:MGU

[1708] Leopold Halpern. Matter and geometry in a unified theory. Foundations of Physics, 24(12):1697–1703, December 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02054791.

Brandt:1994:BR

[1709] Howard E. Brandt. Book review. Foundations of Physics, 24(12):1705– 1709, December 1994. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02054792.

Holland:1995:JPV

[1710] Peter Holland. Jean-Pierre Vigier at seventy-five: La lutte continue. Foundations of Physics, 25(1):1–4, January 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF02054654.

Ghirardi:1995:DMW

[1711] G. C. Ghirardi, R. Grassi, and F. Benatti. Describing the macroscopic world: Closing the circle within the dynamical reduction program. Foundations of Physics, 25(1):5–38, January 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02054655.

Horwitz:1995:USR

[1712] L. P. Horwitz. The unstable system in relativistic quantum mechanics. Foundations of Physics, 25(1):39–65, January 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF02054656.

Droz-Vincent:1995:QME

[1713] Philippe Droz-Vincent. Quantum mechanical evolution of relativistic particles. Foundations of Physics, 25(1):67–90, January 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02054657.

Eberhard:1995:BTB

[1714] Philippe H. Eberhard and Philippe Rosselet. Bell’s theorem based on a generalized EPR criterion of reality. Foundations of Physics, 25(1): 91–111, January 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02054658.

Leggett:1995:RQP

[1715] A. J. Leggett. Is “relative quantum phase” transitive? Foundations of Physics, 25(1):113–122, January 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02054659.

Redhead:1995:MAA

[1716] Michael Redhead. More ado about nothing. Foundations of Physics, 25 (1):123–137, January 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02054660.

Sudarshan:1995:TRS

[1717] E. C. G. Sudarshan and L. C. Biedenharn. Time reversal for systems with internal symmetry. Foundations of Physics, 25(1):139–143, January 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF02054661.

Vermaas:1995:MIQ

[1718] Pieter E. Vermaas and Dennis Dieks. The modal interpretation of quantum mechanics and its generalization to density operators. Foundations of Physics, 25(1):145–158, January 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02054662.

Bozic:1995:CSI

[1719] Mirjana Bozić and Zvonko Marić. Compatible statistical interpretation of a wave packet. Foundations of Physics, 25(1):159–173, January 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02054663.

Evans:1995:CQC

[1720] M. W. Evans. The charge quantization condition in O (3) vacuum electrodynamics. Foundations of Physics, 25(1):175–181, January 1995. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02054664.

Paty:1995:NEO

[1721] Michel Paty. The nature of Einstein’s objections to the Copenhagen interpretation of quantum mechanics. Foundations of Physics, 25(1): 183–204, January 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02054665.

Clifton:1995:BR

[1722] Rob Clifton. Book review. Foundations of Physics, 25(1):205– 209, January 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02054666.

Mandel:1995:IOP

[1723] L. Mandel. Indistinguishability in one-photon and two-photon interference. Foundations of Physics, 25(2):211–218, February 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02055204.

Agarwal:1995:ICS

[1724] G. S. Agarwal. Interference in complementary spaces. Foundations of Physics, 25(2):219–228, February 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02055205.

Leavens:1995:BTF

[1725] C. R. Leavens. Bohm trajectory and Feynman path approaches to the “tunneling time problem”. Foundations of Physics, 25(2):229– 268, February 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02055206.

Cushing:1995:QTT

[1726] James T. Cushing. Quantum tunneling times: a crucial test for the causal program? Foundations of Physics, 25(2):269–280, February 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02055207.

Pagonis:1995:UCD

[1727] Constantine Pagonis and Rob Clifton. Unremarkable contextualism: Dispositions in the Bohm theory. Foundations of Physics, 25(2):281– 296, February 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02055208.

Petroni:1995:QMS

[1728] Nicola Cufaro Petroni and Francesco Guerra. Quantum mechanical states as attractors for Nelson processes. Foundations of Physics, 25 (2):297–315, February 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02055209.

Rafii-Tabar:1995:SMQ

[1729] H. Rafii-Tabar. Simulating the motion of a quantum particle at constant temperature. Foundations of Physics, 25(2):317–328, February 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF02055210.

Brown:1995:BPT

[1730] H. R. Brown, C. Dewdney, and G. Horton. Bohm particles and their detection in the light of neutron interferometry. Foundations of Physics, 25 (2):329–347, February 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02055211.

Anandan:1995:RST

[1731] J. Anandan and H. R. Brown. On the reality of space–time geometry and the wavefunction. Foundations of Physics, 25(2):349–360, February 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF02055212.

Tourrenc:1995:GLV

[1732] P. Tourrenc and T. Melliti. Gravitational limitation on the verification of special relativity in the laboratory. Foundations of Physics, 25(2): 361–376, February 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02055213.

Barut:1995:QTS

[1733] A. O. Barut. Quantum theory of single events continued. Accelerating wavelets and the Stern–Gerlach experiment. Foundations of Physics, 25 (2):377–381, February 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02055214.

Evans:1995:FRR

[1734] M. W. Evans. The B^{(3)} field: Its role in the Rayleigh–Jeans law, Planck law, and Einstein coefficients. Foundations of Physics, 25(2): 383–389, February 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02055215.

Mackman:1995:LIR

[1735] Steve Mackman and Euan Squires. Lorentz invariance and the retarded Bohm model. Foundations of Physics, 25(2):391–397, February 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02055216.

Anonymous:1995:SWP

[1736] Anonymous. Some works and publications of Jean–Pierre Vigier. Foundations of Physics, 25(2):399–407, February 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF02055217.

Israel:1995:BR

[1737] Werner Israel. Book review. Foundations of Physics, 25(2):409– 410, February 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02055218.

Redei:1995:LIQ

[1738] Miklós Rédei. Logical independence in quantum logic. Foundations of Physics, 25(3):411–422, March 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02059228.

Dickson:1995:WPA

[1739] Michael Dickson. What is preferred about the preferred basis? Foundations of Physics, 25(3):423–440, March 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02059229.

Parrott:1995:NPA

[1740] Stephen Parrott and Daniel J. Endres. Nonuniqueness of the “physical” acceleration for the Lorentz–Dirac equation. Foundations of Physics, 25 (3):441–464, March 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02059230.

Cormier-Delanoue:1995:WCD

[1741] C. Cormier-Delanoue. Wave-corpuscle duality of light reconsidered. Foundations of Physics, 25(3):465–479, March 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF02059231.

Chiatti:1995:PIT

[1742] Leonardo Chiatti. Path integral and transactional interpretation. Foundations of Physics, 25(3):481–490, March 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02059232.

Yan:1995:CMM

[1743] C. C. Yan. Can the macroscopic Maxwell equations be obtained from the microscopic Maxwell–Lorentz equations by performing averages? Foundations of Physics, 25(3):491–502, March 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02059233.

Antonelli:1995:BR

[1744] Peter Antonelli. Book review. Foundations of Physics, 25(3):503– 506, March 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02059234.

Cushing:1995:BR

[1745] James T. Cushing. Book review. Foundations of Physics, 25(3): 507–510, March 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02059235.

Villain:1995:BR

[1746] Jacques Villain. Book review. Foundations of Physics, 25(3):511– 515, March 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02059236.

Parker:1995:BR

[1747] Eugene N. Parker. Book review. Foundations of Physics, 25(3): 517–519, March 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02059237.

Anile:1995:BR

[1748] Angelo Marcello Anile. Book review. Foundations of Physics, 25(3): 521–523, March 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02059238.

Riseborough:1995:BR

[1749] Peter S. Riseborough. Book review. Foundations of Physics, 25(3): 525–526, March 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02059239.

Donald:1995:MCP

[1750] Matthew J. Donald. A mathematical characterization of the physical structure of observers. Foundations of Physics, 25(4):529–571, April 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF02059006.

delaPena:1995:QPZ

[1751] L. de la Peña and A. M. Cetto. Quantum phenomena and the zeropoint radiation field. II. Foundations of Physics, 25(4):573–604, April 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02059007.

Omnes:1995:NIQ

[1752] Roland Omnès. A new interpretation of quantum mechanics and its consequences in epistemology. Foundations of Physics, 25(4):605–629, April 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF02059008.

Cooperstock:1995:EU

[1753] F. I. Cooperstock and M. Israelit. The energy of the Universe. Foundations of Physics, 25(4):631–635, April 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02059009.

Sudarshan:1995:GDM

[1754] E. C. G. Sudarshan. Groups as dynamical models. Foundations of Physics, 25(4):637–641, April 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02059010.

Wheeler:1995:HSS

[1755] James T. Wheeler. Horizons and singularities in static, spherically symmetric spacetimes. Foundations of Physics, 25(5):645–679, May 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02059122.

Hegstrom:1995:MQM

[1756] Roger A. Hegstrom and Fernando Sols. A model of quantum measurement in Josephson junctions. Foundations of Physics, 25(5): 681–700, May 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02059123.

Filho:1995:PPQ

[1757] J. Barretto Bastos Filho and F. Selleri. Propensity, probability, and quantum physics. Foundations of Physics, 25(5):701–716, May 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02059124.

Beil:1995:MFT

[1758] R. G. Beil. Moving frame transport and gauge transformations. Foundations of Physics, 25(5):717–742, May 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02059125.

Jadczyk:1995:QJE

[1759] A. Jadczyk. On quantum jumps, events, and spontaneous localization models. Foundations of Physics, 25(5):743–762, May 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02059126.

Israelit:1995:CDM

[1760] Mark Israelit and Nathan Rosen. Cosmic dark matter and Dirac gauge function. Foundations of Physics, 25(5):763–777, May 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02059127.

Capria:1995:MME

[1761] Marco Mamone Capria and Fernanda Pambianco. On the Michelson– Morley experiment. Foundations of Physics, 25(5):779, May 1995. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02059128.

Anonymous:1995:CPa

[1762] Anonymous. Call for papers. Foundations of Physics, 25(5):781, May 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF02059129.

Burakovsky:1995:ERM

[1763] L. Burakovsky and L. P. Horwitz. Equilibrium relativistic mass distribution for indistinguishable events. Foundations of Physics, 25(6): 785–818, June 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02080564.

Pavsic:1995:RBC

[1764] Matej Pavsic. Relativistic p-branes without constraints and their relation to the wiggly extended objects. Foundations of Physics, 25 (6):819–832, June 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02080565.

Bogdanowicz:1995:AML

[1765] Jerzy Bogdanowicz, Maciej Pindor, and Ryszard Raczka. An analysis of mean life and lifetime of unstable elementary particles. Foundations of Physics, 25(6):833–849, June 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02080566.

Formanek:1995:PSU

[1766] J. Formánek. On the problem of systems under the influence of imperfect instruments. Foundations of Physics, 25(6):851–870, June 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02080567.

Rodrigues:1995:MSN

[1767] Waldyr A. Rodrigues, Jr., Quintino A. G. de Souza, and Yuri Bozhkov. The mathematical structure of Newtonian spacetime: Classical dynamics and gravitation. Foundations of Physics, 25(6):871–924, June 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02080568.

Cereceda:1995:KST

[1768] José L. Cereceda. The Kochen–Specker theorem and Bell’s theorem: an algebraic approach. Foundations of Physics, 25(6):925–949, June 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02080569.

Asanov:1995:GFE

[1769] R. A. Asanov. Gravitational field of electrically charged mass in the Lobachevski space. Foundations of Physics, 25(6):951–957, June 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02080570.

Dariescu:1995:GTF

[1770] Marina-Aura Dariescu, C. Dariescu, and I. Gottlieb. Gauge theory of fermions on R × S^{3} spacetime. Foundations of Physics, 25(6): 959–963, June 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02080571.

Anonymous:1995:CPb

[1771] Anonymous. Call for papers. Foundations of Physics, 25(6):965, June 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF02080572.

Budinich:1995:CCH

[1772] P. Budinich. Conformally compactified homogeneous spaces. Possible observable consequences. Foundations of Physics, 25(7):969–993, July 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF02059522.

Cohen:1995:QHCa

[1773] Marcus S. Cohen. Quantization of helicity on a compact spacetime. Foundations of Physics, 25(7):995–1028, July 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02059523.

Carmeli:1995:CRS

[1774] M. Carmeli. Cosmological relativity: a special relativity for cosmology. Foundations of Physics, 25(7):1029–1040, July 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF02059524.

Gillespie:1995:ISE

[1775] Daniel T. Gillespie. Incompatibility of the Schrödinger equation with Langevin and Fokker–Planck equations. Foundations of Physics, 25 (7):1041–1053, July 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02059525.

Rylov:1995:PED

[1776] Y. A. Rylov. Pauli’s electron as a dynamic system. Foundations of Physics, 25(7):1055–1086, July 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02059526.

Misra:1995:TOC

[1777] B. Misra. From time operator to chronons. Foundations of Physics, 25 (7):1087–1104, July 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02059527.

Ghose:1995:AAA

[1778] Partha Ghose and Dipankar Home. An analysis of the Aharonov– Anandan–Vaidman model. Foundations of Physics, 25(7):1105–1109, July 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02059528.

Phipps:1995:BR

[1779] Thomas E. Phipps, Jr. Book review. Foundations of Physics, 25(7): 1111–1115, July 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02059529.

Welsch:1995:BR

[1780] Dirk-Gunnar Welsch. Book review. Foundations of Physics, 25(7): 1117–1119, July 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02059530.

Anonymous:1995:CPc

[1781] Anonymous. Call for papers. Foundations of Physics, 25(7):1121, July 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF02059531.

Anonymous:1995:A

[1782] Anonymous. Announcement. Foundations of Physics, 25(7):1123– 1124, July 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02059532.

Burakovsky:1995:RMD

[1783] L. Burakovsky and L. P. Horwitz. Relativistic mass distribution in eventanti-event system and “realistic” equation of state for hot hadronic matter. Foundations of Physics, 25(8):1127–1146, August 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02055255.

Cattaneo:1995:SRB

[1784] Gianpiero Cattaneo and Roberto Giuntini. Some results on BZ structures from Hilbertian unsharp quantum physics. Foundations of Physics, 25 (8):1147–1183, August 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02055256.

Coecke:1995:HMR

[1785] Bob Coecke. A hidden measurement representation for quantum entities described by finite-dimensional complex Hilbert spaces. Foundations of Physics, 25(8):1185–1208, August 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02055257.

Fujita:1995:BED

[1786] Shigeji Fujita, Salvador Godoy, and Diep Nguyen. Bloch electron dynamics. Foundations of Physics, 25(8):1209–1220, August 1995. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02055258.

Biedenharn:1995:QMT

[1787] L. C. Biedenharn and J. C. Solem. A quantum-mechanical treatment of Szilard’s engine: Implications for the entropy of information. Foundations of Physics, 25(8):1221–1229, August 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02055259.

Griffiths:1995:BR

[1788] Robert B. Griffiths. Book review. Foundations of Physics, 25(8): 1231–1236, August 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02055260.

Busch:1995:RMQ

[1789] Paul Busch, Marian Grabowski, and Pekka J. Lahti. Repeatable measurements in quantum theory: Their role and feasibility. Foundations of Physics, 25(9):1239–1266, September 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02055331.

Dove:1995:SVE

[1790] Chris Dove and Evan J. Squires. Symmetric versions of explicit wavefunction collapse models. Foundations of Physics, 25(9):1267– 1282, September 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02055332.

Chew:1995:SC

[1791] G. F. Chew. Standpoint cosmology. Foundations of Physics, 25(9): 1283–1333, September 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02055333.

Burakovsky:1995:GBE

[1792] L. Burakovsky and L. P. Horwitz. Generalized Boltzmann equation in a manifestly covariant relativistic statistical mechanics. Foundations of Physics, 25(9):1335–1358, September 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02055334.

Dotson:1995:CVP

[1793] Allen C. Dotson. On creating values of physical properties nonlocally. Foundations of Physics, 25(9):1359–1370, September 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02055335.

Miron:1995:RON

[1794] R. Miron and G. Zet. Relativistic optics of nondispersive media. Foundations of Physics, 25(9):1371–1382, September 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF02055336.

Eisenberg:1995:BRa

[1795] Judah M. Eisenberg. Book review. Foundations of Physics, 25(9):1383– 1385, September 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02055337.

Eisenberg:1995:BRb

[1796] Judah M. Eisenberg. Book review. Foundations of Physics, 25(9):1387– 1389, September 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02055338.

Eisenberg:1995:BRc

[1797] Judah M. Eisenberg. Book review. Foundations of Physics, 25(9):1391– 1393, September 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02055339.

Eisenberg:1995:BRd

[1798] Judah M. Eisenberg. Book review. Foundations of Physics, 25(9):1395– 1396, September 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02055340.

Collett:1995:CCM

[1799] Brian Collett, Philip Pearle, Frank Avignone, and Shmuel Nussinov. Constraint on collapse models by limit on spontaneous X-ray emission in Ge. Foundations of Physics, 25(10):1399–1412, October 1995. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02057460.

Kaloyerou:1995:CIM

[1800] P. N. Kaloyerou. Causal interpretation of the modified Klein–Gordon equation. Foundations of Physics, 25(10):1413–1460, October 1995. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02057461.

Vigier:1995:DIF

[1801] Jean-Pierre Vigier. Derivation of inertial forces from the Einstein–de Broglie–Bohm (E.d.B.B.) causal stochastic interpretation of quantum mechanics. Foundations of Physics, 25(10):1461–1494, October 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02057462.

Frastai:1995:SFP

[1802] J. Frastai and L. P. Horwitz. Off-shell fields and Pauli–Villars regularization. Foundations of Physics, 25(10):1495–1510, October 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02057463.

Chisholm:1995:USG

[1803] J. S. R. Chisholm and R. S. Farwell. Unified spin gauge model and the top quark mass. Foundations of Physics, 25(10):1511–1522, October 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF02057464.

Dariescu:1995:FKN

[1804] Marina-Aura Dariescu, C. Dariescu, and I. Gottlieb. Fermions in a Kerr– Newman space–time. Foundations of Physics, 25(10):1523–1528, October 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF02057465.

Sheldon:1995:BR

[1805] Eric Sheldon. Book review. Foundations of Physics, 25(10):1529– 1534, October 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02057466.

Ruijgrok:1995:BR

[1806] Th. W. Ruijgrok. Book review. Foundations of Physics, 25(10):1535– 1537, October 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02057467.

Cohen:1995:QHCb

[1807] Marcus S. Cohen. Quantization of helicity on a compact spacetime. Foundations of Physics, 25(10):1539, October 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02057468.

Svozil:1995:STP

[1808] K. Svozil. Set theory and physics. Foundations of Physics, 25(11): 1541–1560, November 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02055507.

Uchiyama:1995:LRC

[1809] Satoshi Uchiyama. Local reality: Can it exist in the EPR–Bohm gedanken experiment? Foundations of Physics, 25(11):1561–1575, November 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02055508.

Beil:1995:PTF

[1810] R. G. Beil. Poincaré transport of frames. Foundations of Physics, 25(11): 1577–1597, November 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02055509.

Dechoum:1995:NHSa

[1811] K. Dechoum and H. M. FranÇa. Non-Heisenberg states of the harmonic oscillator. Foundations of Physics, 25(11):1599–1620, November 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02055510.

Dechoum:1995:NHSb

[1812] K. Dechoum and H. M. FranÇa. Non-Heisenberg states of the harmonic oscillator. Foundations of Physics, 25(11):1620, November 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02055511.

Mould:1995:IOQ

[1813] Richard Mould. The inside observer in quantum mechanics. Foundations of Physics, 25(11):1621–1629, November 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02055512.

Hawkins:1995:BR

[1814] David Hawkins. Book review. Foundations of Physics, 25(11):1631– 1636, November 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02055513.

Henry:1995:BR

[1815] Richard C. Henry. Book review. Foundations of Physics, 25(11):1637– 1639, November 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02055514.

Boyd:1995:BR

[1816] Sir Robert Boyd, FRS. Book review. Foundations of Physics, 25(11): 1641–1643, November 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02055515.

Grassmann:1995:BR

[1817] Winfried Grassmann. Book review. Foundations of Physics, 25(11): 1645–1647, November 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02055516.

Rosen:1995:BR

[1818] S. P. Rosen. Book review. Foundations of Physics, 25(11):1649– 1651, November 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02055517.

Slemrod:1995:BR

[1819] Marshall Slemrod. Book review. Foundations of Physics, 25(11):1653– 1655, November 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02055518.

Vilgis:1995:BR

[1820] T. Vilgis. Book review. Foundations of Physics, 25(11):1657– 1659, November 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02055519.

Uzer:1995:BR

[1821] Turgay Uzer. Book review. Foundations of Physics, 25(11):1661– 1662, November 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02055520.

Olver:1995:BR

[1822] Peter Olver. Book review. Foundations of Physics, 25(11):1663– 1665, November 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02055521.

Cohen:1995:RQM

[1823] O. Cohen and B. J. Hiley. Retrodiction in quantum mechanics, preferred Lorentz frames, and nonlocal measurements. Foundations of Physics, 25(12):1669–1698, December 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02057882.

Bennett:1995:PSE

[1824] M. K. Bennett and D. J. Foulis. Phi-symmetric effect algebras. Foundations of Physics, 25(12):1699–1722, December 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF02057883.

Rybakov:1995:SMA

[1825] Yu. P. Rybakov and B. Saha. Soliton model of atom. Foundations of Physics, 25(12):1723–1731, December 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02057884.

Ben-Yaacov:1995:SMA

[1826] Uri Ben-Ya’acov. Statistical mechanics analysis of the “Twins Paradox”. Foundations of Physics, 25(12):1733–1740, December 1995. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02057885.

King:1995:EAL

[1827] B. Tom King. An Einstein addition law for nonparallel boosts using the geometric algebra of space–time. Foundations of Physics, 25(12): 1741–1755, December 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02057886.

Nistico:1995:ERB

[1828] Giuseppe Nisticò. Empirical relations between noncommuting observables. Foundations of Physics, 25(12):1757–1767, December 1995. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02057887.

Smida:1995:GDC

[1829] A. Smida, M. Hachemane, and M. Fellah. Geometro-differential conception of extended particles and their quantum theory in de Sitter space. Foundations of Physics, 25(12):1769–1795, December 1995. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02057888.

Coester:1995:BR

[1830] F. Coester. Book review. Foundations of Physics, 25(12):1797– 1798, December 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02057889.

Hansen:1995:BR

[1831] Uwe J. Hansen. Book review. Foundations of Physics, 25(12):1799– 1800, December 1995. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02057890.

Cohen:1996:ERL

[1832] O. Cohen and B. J. Hiley. Elements of reality, Lorentz invariance, and the product rule. Foundations of Physics, 26(1):1–15, January 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02058886.

Schray:1996:ORC

[1833] Jörg Schray and Corinne A. Manogue. Octonionic representations of Clifford algebras and triality. Foundations of Physics, 26(1):17–70, January 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF02058887.

Fahn:1996:MDE

[1834] Paul N. Fahn. Maxwell’s demon and the entropy cost of information. Foundations of Physics, 26(1):71–93, January 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF02058888.

Cormier-Delanoue:1996:SMW

[1835] Christian Cormier-Delanoue. Strangeness of matter waves. Foundations of Physics, 26(1):95–103, January 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02058889.

Dehnen:1996:DPE

[1836] H. Dehnen and D. Ebner. Derivation of the principle of equivalence for antimatter. Foundations of Physics, 26(1):105–115, January 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02058890.

Aharonov:1996:MPM

[1837] Yakir Aharonov, Jeeva Anandan, and Lev Vaidman. The meaning of protective measurements. Foundations of Physics, 26(1):117–126, January 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF02058891.

Kak:1996:IPC

[1838] Subhash C. Kak. Information, physics, and computation. Foundations of Physics, 26(1):127–137, January 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02058892.

Cushing:1996:BRa

[1839] James T. Cushing. Book review. Foundations of Physics, 26(1): 139–142, January 1996. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02058893.

Schlesinger:1996:BR

[1840] George N. Schlesinger. Book review. Foundations of Physics, 26(1): 143–145, January 1996. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02058894.

Silk:1996:BR

[1841] Joseph Silk. Book review. Foundations of Physics, 26(1):147– 149, January 1996. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02058895.

Anonymous:1996:IISa

[1842] Anonymous. II International Symposium on Fundamental Problems in Quantum Physics. Foundations of Physics, 26(1):155–156, January 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02058896.

Pavsic:1996:RTP

[1843] Matej Pavsic. On the resolution of time problem in quantum gravity induced from unconstrained membranes. Foundations of Physics, 26(2): 159–195, February 1996. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02058084.

Johnson:1996:DPD

[1844] Robert W. Johnson. Direct product and decomposition of certain physically important algebras. Foundations of Physics, 26(2):197– 222, February 1996. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02058085.

Piirainen:1996:MTC

[1845] Reijo Piirainen. Möbius transformation and conformal relativity. Foundations of Physics, 26(2):223–242, February 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF02058086.

Kocharovsky:1996:SCI

[1846] V. V. Kocharovsky and Vl. V. Kocharovsky. Self-consistent infrared and ultraviolet asymptotically free unitary renormalizable theory of quantum gravity and matter fields. Foundations of Physics, 26(2): 243–256, February 1996. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02058087.

Nikolov:1996:QPU

[1847] Blagowest A. Nikolov. Quantum probability and unified approach to quantization and dynamics. Foundations of Physics, 26(2):257– 269, February 1996. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02058088.

Assis:1996:NFI

[1848] André K. T. Assis and Peter Graneau. Nonlocal forces of inertia in cosmology. Foundations of Physics, 26(2):271–283, February 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02058089.

Anonymous:1996:IISb

[1849] Anonymous. II International Symposium on Fundamental Problems in Quantum Physics. Foundations of Physics, 26(2):285–286, February 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF02058090.

Peres:1996:NR

[1850] Asher Peres. Nathan Rosen, 1909–1995. Foundations of Physics, 26 (3):289–290, March 1996. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02069473.

Pearle:1996:GEC

[1851] Philip Pearle and Euan Squires. Gravity, energy conservation, and parameter values in collapse models. Foundations of Physics, 26(3): 291–305, March 1996. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02069474.

Batterman:1996:CAC

[1852] Robert W. Batterman and Homer White. Chaos and algorithmic complexity. Foundations of Physics, 26(3):307–336, March 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02069475.

Leaf:1996:PNF

[1853] Boris Leaf. Positive- and negative-frequency parts of D’Alembert’s equation with applications in electrodynamics. Foundations of Physics, 26 (3):337–368, March 1996. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02069476.

Nairz:1996:CMT

[1854] A. Nairz. A class of metric theories of gravitation on Minkowski spacetime. Foundations of Physics, 26(3):369–389, March 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02069477.

deBeauregard:1996:JPC

[1855] O. Costa de Beauregard. On the joint probability of correlated physical occurrences. Foundations of Physics, 26(3):391–399, March 1996. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02069478.

Ramachandran:1996:QPD

[1856] G. Ramachandran, A. R. Usha Devi, P. Devi, and Swarnamala Sirsi. Quasi-probability distributions for arbitrary spin- j particles. Foundations of Physics, 26(3):401–412, March 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02069479.

Carmeli:1996:CSR

[1857] M. Carmeli. Cosmological special relativity. Foundations of Physics, 26 (3):413–416, March 1996. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02069480.

Schroder:1996:GRV

[1858] Wilfried Schröder and Hans-Jürgen Treder. On general relativistic vortex-dynamics. Foundations of Physics, 26(3):417–422, March 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02069481.

Sklar:1996:BR

[1859] Lawrence Sklar. Book review. Foundations of Physics, 26(3):423– 426, March 1996. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02069482.

Cushing:1996:BRb

[1860] James T. Cushing. Book review. Foundations of Physics, 26(3): 427–429, March 1996. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02069483.

Gale:1996:BR

[1861] George Gale. Book review. Foundations of Physics, 26(3):431– 433, March 1996. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02069484.

Ding:1996:BR

[1862] Jiu Ding. Book review. Foundations of Physics, 26(3):435–436, March 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF02069485.

Anonymous:1996:IISc

[1863] Anonymous. II International symposium on fundamental problems in quantum physics. Foundations of Physics, 26(3):437–438, March 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02069486.

Kaveh:1996:MJH

[1864] Moshe Kaveh. For Max Jammer on his eightieth birthday. Foundations of Physics, 26(4):441–442, April 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02071213.

Rohrlich:1996:SRC

[1865] Fritz Rohrlich. Scientific realism: a challenge to physicists. Foundations of Physics, 26(4):443–451, April 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02071214.

Holton:1996:RTS

[1866] Gerald Holton. The role of themata in science. Foundations of Physics, 26(4):453–465, April 1996. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02071215.

Schlesinger:1996:PTE

[1867] George N. Schlesinger. The power of thought experiments. Foundations of Physics, 26(4):467–482, April 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02071216.

Agassi:1996:PMH

[1868] Joseph Agassi. The place of metaphysics in the historiography of science. Foundations of Physics, 26(4):483–499, April 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF02071217.

Hon:1996:DSD

[1869] Giora Hon. Disturbing, but not surprising: Did Gödel surprise Einstein with a rotating universe and time travel? Foundations of Physics, 26 (4):501–521, April 1996. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.

1007/BF02071218. See comments [2177].

Grunbaum:1996:TMC

[1870] Adolf Grünbaum. Theological misinterpretations of current physical cosmology. Foundations of Physics, 26(4):523–543, April 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02071219.

Beller:1996:CAH

[1871] Mara Beller. The conceptual and the anecdotal history of quantum mechanics. Foundations of Physics, 26(4):545–557, April 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02071220.

Bergmann:1996:BRa

[1872] Peter G. Bergmann. Book review. Foundations of Physics, 26(4): 559–561, April 1996. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02071221.

Murthy:1996:BR

[1873] Ganpathy Murthy. Book review. Foundations of Physics, 26(4): 563–566, April 1996. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02071222.

Anonymous:1996:Aa

[1874] Anonymous. Announcements. Foundations of Physics, 26(4):567– 572, April 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02071223.

Neeman:1996:PAG

[1875] Yuval Ne’eman. Plato alleges that God forever geometrizes. Foundations of Physics, 26(5):575–583, May 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02058232.

Israelit:1996:WDG

[1876] Mark Israelit and Nathan Rosen. A Weyl–Dirac geometric particle. Foundations of Physics, 26(5):585–594, May 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02058233.

Israel:1996:ISS

[1877] Werner Israel. Imploding stars, shifting continents, and the inconstancy of matter. Foundations of Physics, 26(5):595–616, May 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02058234.

Deser:1996:IFS

[1878] S. Deser. Inequivalence of first- and second-order formulations in d=2 gravity models. Foundations of Physics, 26(5):617–621, May 1996. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02058235.

Earman:1996:TSS

[1879] John Earman. Tolerance for spacetime singularities. Foundations of Physics, 26(5):623–640, May 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02058236.

Selleri:1996:NOW

[1880] F. Selleri. Noninvariant one-way velocity of light. Foundations of Physics, 26(5):641–664, May 1996. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02058237.

Ludu:1996:GKE

[1881] A. Ludu, R. A. Ionescu, and W. Greiner. Generalized KdV equation for fluid dynamics and quantum algebras. Foundations of Physics, 26 (5):665–678, May 1996. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02058238.

Kogan:1996:SWP

[1882] Eugene Kogan and Moshe Kaveh. Statistics of waves propagating in a random medium. Foundations of Physics, 26(5):679–690, May 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02058239.

Berkovits:1996:IED

[1883] Richard Berkovits. Interacting electrons in disordered potentials: The inverse compressibility. Foundations of Physics, 26(5):691–699, May 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02058240.

Freilikher:1996:TTD

[1884] V. Freilikher, M. Pustilnik, and I. Yurkevich. Tunneling through a disordered potential barrier. Foundations of Physics, 26(5):701–712, May 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF02058241.

Anonymous:1996:Ab

[1885] Anonymous. Announcements. Foundations of Physics, 26(5):713– 718, May 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02058242.

Bohm:1996:RHV

[1886] David Bohm. On the role of hidden variables in the fundamental structure of physics. Foundations of Physics, 26(6):719–786, June 1996. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02058632.

Bub:1996:STK

[1887] Jeffrey Bub. Schütte’s tautology and the Kochen–Specker theorem. Foundations of Physics, 26(6):787–806, June 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF02058633.

Peres:1996:GKS

[1888] Asher Peres. Generalized Kochen–Specker theorem. Foundations of Physics, 26(6):807–812, June 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02058634.

Gudder:1996:A

[1889] Stanley Gudder. D-algebras. Foundations of Physics, 26(6):813– 822, June 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02058635.

Bohm:1996:SMO

[1890] D. Bohm and B. J. Hiley. Statistical mechanics and the ontological interpretation. Foundations of Physics, 26(6):823–846, June 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02058636.

Revzen:1996:BIS

[1891] M. Revzen and A. Mann. Bell’s inequality for a single particle. Foundations of Physics, 26(6):847–850, June 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02058637.

Horwitz:1996:HQM

[1892] L. P. Horwitz. Hypercomplex quantum mechanics. Foundations of Physics, 26(6):851–862, June 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02058638.

Kutach:1996:BR

[1893] Douglas Kutach and Robert Weingard. Book review. Foundations of Physics, 26(6):863–866, June 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02058639.

Anonymous:1996:Ac

[1894] Anonymous. Announcements. Foundations of Physics, 26(6):867– 872, June 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02058640.

Busch:1996:SMQ

[1895] Paul Busch and Pekka J. Lahti. The standard model of quantum measurement theory: History and applications. Foundations of Physics, 26 (7):875–893, July 1996. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02148831.

Vaidman:1996:WME

[1896] Lev Vaidman. Weak-measurement elements of reality. Foundations of Physics, 26(7):895–906, July 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02148832.

Tarozzi:1996:QMM

[1897] G. Tarozzi. Quantum measurements and macrophysical reality: Epistemological implications of a proposed paradox. Foundations of Physics, 26(7):907–917, July 1996. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02148833.

dEspagnat:1996:IHR

[1898] Bernard d’Espagnat. Influences, histories, and reality. Foundations of Physics, 26(7):919–928, July 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02148834.

vonBorzeszkowski:1996:MED

[1899] H.-H. von Borzeszkowski and H.-J. Treder. Mach–Einstein doctrine and general relativity. Foundations of Physics, 26(7):929–942, July 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02148835.

Ghose:1996:TPE

[1900] Partha Ghose and Dipankar Home. The two-prism experiment and waveparticle duality of light. Foundations of Physics, 26(7):943–953, July 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF02148836.

Woodward:1996:BRa

[1901] James F. Woodward. Book reviews. Foundations of Physics, 26(7): 955–964, July 1996. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02148837.

Dixon:1996:BR

[1902] Geoffrey Dixon. Book reviews. Foundations of Physics, 26(7): 965–966, July 1996. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02148838.

Anonymous:1996:IISd

[1903] Anonymous. II International symposium on fundamental problems in quantum physics. Foundations of Physics, 26(7):967–968, July 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02148839.

Anonymous:1996:Ad

[1904] Anonymous. Announcements. Foundations of Physics, 26(7):969, July 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF02148840.

Anonymous:1996:BRB

[1905] Anonymous. Book review: Physical Interpretations of Relativity Theory, British Society for the Philosophy of Science 6–9 September 1996, London. Foundations of Physics, 26(7):970, July 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF02148841.

Giardina:1996:EIQ

[1906] Irene Giardina and Alberto Rimini. On the existence of inequivalent quasideterministic domains. Foundations of Physics, 26(8):973– 987, August 1996. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02061399.

Belnap:1996:BST

[1907] Nuel Belnap and László E. Szabó. Branching space–time analysis of the GHZ theorem. Foundations of Physics, 26(8):989–1002, August 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02061400.

Ingall:1996:NWW

[1908] J. E. M. Ingall. The Newton–Wigner and Wightman localization of the photon. Foundations of Physics, 26(8):1003–1031, August 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02061401.

Khrennikov:1996:UHS

[1909] Andrew Khrennikov. The ultrametric Hilbert-space description of quantum measurements with a finite exactness. Foundations of Physics, 26 (8):1033–1054, August 1996. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02061402.

Franca:1996:MET

[1910] H. M. França, A. Maia, Jr., and C. P. Malta. Maxwell electromagnetic theory, Planck’s radiation law, and Bose–Einstein statistics. Foundations of Physics, 26(8):1055–1068, August 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02061403.

Dariescu:1996:PSS

[1911] Ciprian Dariescu. Planary symmetric static worlds with massless scalar sources. Foundations of Physics, 26(8):1069–1080, August 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02061404.

Schroder:1996:HEC

[1912] W. Schröder and H.-J. Treder. Hans Ertel and cosmology. Foundations of Physics, 26(8):1081–1088, August 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02061405.

Piirainen:1996:C

[1913] Reijo Piirainen. Corrigendum. Foundations of Physics, 26(8):1089– 1090, August 1996. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02061406.

Stapp:1996:BR

[1914] Henry P. Stapp. Book review. Foundations of Physics, 26(8):1091– 1097, August 1996. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02061407.

Bergia:1996:BR

[1915] Silvio Bergia. Book review. Foundations of Physics, 26(8):1099– 1101, August 1996. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02061408.

Bergmann:1996:BRb

[1916] Peter G. Bergmann. Book review. Foundations of Physics, 26(8): 1103–1104, August 1996. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02061409.

Lichtenberg:1996:BR

[1917] Don Lichtenberg. Book review. Foundations of Physics, 26(8):1105– 1107, August 1996. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02061410.

Fulling:1996:BR

[1918] Stephen A. Fulling. Book review. Foundations of Physics, 26(8): 1109–1111, August 1996. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02061411.

Martin:1996:BR

[1919] Jürgen Martin and Richard N. Zare. Book review. Foundations of Physics, 26(8):1113–1115, August 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02061412.

Anonymous:1996:Ae

[1920] Anonymous. Announcements. Foundations of Physics, 26(8):1117– 1118, August 1996. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02061413.

Garola:1996:TAS

[1921] Claudio Garola and Luigi Solombrino. The theoretical apparatus of semantic realism: a new language for classical and quantum physics. Foundations of Physics, 26(9):1121–1164, September 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF02275624.

Wan:1996:QPS

[1922] K. Kong Wan and R. H. Fountain. Quantization by parts, self-adjoint extensions, and a novel derivation of the Josephson equation in superconductivity. Foundations of Physics, 26(9):1165–1199, September 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02275625.

Petrov:1996:EDS

[1923] A. N. Petrov and J. V. Narlikar. The energy distribution for a spherically symmetric isolated system in general relativity. Foundations of Physics, 26(9):1201–1229, September 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02275626.

Pitowsky:1996:LFM

[1924] Itamar Pitowsky and Noam Shoresh. Locality, factorizability, and the Maxwell Boltzmann distribution. Foundations of Physics, 26(9):1231– 1242, September 1996. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02275627.

Evans:1996:UGE

[1925] M. W. Evans. Unification of gravitation and electromagnetism with b^{(3)}. Foundations of Physics, 26(9):1243–1261, September 1996. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02275628.

Matzner:1996:BR

[1926] Richard Matzner. Book review. Foundations of Physics, 26(9):1263– 1266, September 1996. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02275629.

Greiner:1996:BR

[1927] Walter Greiner. Book review. Foundations of Physics, 26(9):1267– 1269, September 1996. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02275630.

Kulik:1996:BR

[1928] Andrzej S. Kulik and Yue Wu. Book review. Foundations of Physics, 26(9):1271–1273, September 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02275631.

Ungar:1996:MG

[1929] Abraham A. Ungar. Midpoints in gyrogroups. Foundations of Physics, 26(10):1277–1328, October 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02058271.

Garola:1996:SRV

[1930] Claudio Garola and Luigi Solombrino. Semantic realism versus EPR-like paradoxes: The Furry, Bohm–Aharonov, and Bell paradoxes. Foundations of Physics, 26(10):1329–1356, October 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF02058272.

Cassinello:1996:PPQ

[1931] Andrés Cassinello and José Luis Sánchez-Gómez. On the probabilistic postulate of quantum mechanics. Foundations of Physics, 26(10): 1357–1374, October 1996. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02058273.

Kak:1996:SCS

[1932] Subhash C. Kak. Speed of computation and simulation. Foundations of Physics, 26(10):1375–1386, October 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02058274.

Stoeckler:1996:BR

[1933] Manfred Stoeckler. Book review. Foundations of Physics, 26(10): 1387–1390, October 1996. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02058275.

Wald:1996:BR

[1934] Robert Wald. Book review. Foundations of Physics, 26(10):1391– 1393, October 1996. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02058276.

Sher:1996:BR

[1935] Marc Sher. Book review. Foundations of Physics, 26(10):1395– 1397, October 1996. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02058277.

Gilbert:1996:MPR

[1936] Barry C. Gilbert and Sue Sulcs. The measurement problem resolved and local realism preserved via a collapse-free photon detection model. Foundations of Physics, 26(11):1401–1439, November 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02272365.

Ghose:1996:RQM

[1937] Partha Ghose. Relativistic quantum mechanics of spin-0 and spin-1 bosons. Foundations of Physics, 26(11):1441–1455, November 1996. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02272366.

Rodriguez-Romo:1996:CRH

[1938] Suemi Rodríguez-Romo. On Clifford representation of Hopf algebras and Fierz identities. Foundations of Physics, 26(11):1457–1468, November 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF02272367.

Collins:1996:DPN

[1939] R. Eugene Collins. Differentiable probabilities: a new viewpoint on spin, gauge invariance, gauge fields, and relativistic quantum mechanics. Foundations of Physics, 26(11):1469–1527, November 1996. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02272368.

Johnson:1996:CTI

[1940] John L. Johnson. Coordinate transform invariance. Foundations of Physics, 26(11):1529–1557, November 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02272369.

Cole:1996:BR

[1941] Daniel C. Cole and Alfonso Rueda. Book review. Foundations of Physics, 26(11):1559–1562, November 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02272370.

Brooke:1996:BR

[1942] James A. Brooke. Book review. Foundations of Physics, 26(11):1563– 1566, November 1996. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02272371.

Phipps:1996:BR

[1943] Thomas E. Phipps, Jr. Book review. Foundations of Physics, 26(11): 1567–1571, November 1996. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02272372.

Schieve:1996:LHH

[1944] W. C. Schieve. For Lawrence Horwitz on his sixty-fifth birthday. Foundations of Physics, 26(12):1575–1578, December 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF02282120.

Adler:1996:NGP

[1945] Stephen L. Adler and Jeeva Anandan. Nonadiabatic geometric phase in quaternionic Hilbert space. Foundations of Physics, 26(12):1579– 1589, December 1996. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02282121.

Brumby:1996:GEQ

[1946] S. P. Brumby and G. C. Joshi. Global effects in quaternionic quantum field theory. Foundations of Physics, 26(12):1591–1599, December 1996.

CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02282122.

Dixon:1996:OXP

[1947] Geoffrey Dixon. Octonion x, y-product G_{2} variants. Foundations of Physics, 26(12):1601–1605, December 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02282123.

Neeman:1996:CQ

[1948] Yuval Ne’eman. CHN, QCD, and \overline{SA}(4, R). Foundations of Physics, 26(12):1607–1615, December 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02282124.

Rohrlich:1996:UEP

[1949] Fritz Rohrlich. The unreasonable effectiveness of physical intuition: Success while ignoring objections. Foundations of Physics, 26(12):1617– 1626, December 1996. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02282125.

Aerts:1996:RTW

[1950] Diederik Aerts. Relativity theory: What is reality? Foundations of Physics, 26(12):1627–1644, December 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02282126.

Prugovecki:1996:LQG

[1951] Eduard Prugovecki. On locality in quantum general relativity and quantum gravity. Foundations of Physics, 26(12):1645–1668, December 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02282127.

Blanchard:1996:RQE

[1952] Ph. Blanchard and A. Jadczyk. Relativistic quantum events. Foundations of Physics, 26(12):1669–1681, December 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02282128.

Mannheim:1996:LGG

[1953] Philip D. Mannheim. Local and global gravity. Foundations of Physics, 26(12):1683–1709, December 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02282129.

Anonymous:1996:BLP

[1954] Anonymous. Bibliography of Lawrence P. Horwitz. Foundations of Physics, 26(12):1711–1723, December 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02282130.

Woodward:1996:BRb

[1955] James F. Woodward. Book review. Foundations of Physics, 26(12): 1725–1731, December 1996. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02282131.

Coleman:1996:BR

[1956] Piers Coleman. Book review. Foundations of Physics, 26(12):1733– 1735, December 1996. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02282132.

Trump:1997:SPC

[1957] Matthew Trump and W. C. Schieve. The synchronization problem in covariant relativistic dynamics. Foundations of Physics, 27(1): 1–17, January 1997. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02550152.

Land:1997:PEC

[1958] M. C. Land. Particles and events in classical off-shell electrodynamics. Foundations of Physics, 27(1):19–41, January 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF02550153.

Hannibal:1997:RSP

[1959] Ludger Hannibal. Relativistic spin on the Poincaré group. Foundations of Physics, 27(1):43–56, January 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02550154.

Owen:1997:BSE

[1960] David A. Owen. The Bethe–Salpeter equation for spin-1 particles. Foundations of Physics, 27(1):57–66, January 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02550155.

VanAlstine:1997:TTE

[1961] Peter Van Alstine and Horace W. Crater. A tale of three equations: Breit, Eddington–Gaunt, and two-body Dirac. Foundations of Physics, 27(1):67–79, January 1997. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02550156.

Sastry:1997:DP

[1962] Ramchander R. Sastry and John R. Klauder. Discontinuous perturbations. Foundations of Physics, 27(1):81–91, January 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02550157.

Neuberger:1997:RCG

[1963] Herbert Neuberger. Regularization of chiral gauge theories. Foundations of Physics, 27(1):93–99, January 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02550158.

Kovner:1997:DSV

[1964] Alex Kovner. Dual superconductivity. Variations on a theme. Foundations of Physics, 27(1):101–111, January 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02550159.

Marinov:1997:CTD

[1965] M. S. Marinov and Bilha Segev. Causality and time dependence in quantum tunneling. Foundations of Physics, 27(1):113–132, January 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02550160.

Eisenberg:1997:QPD

[1966] E. Eisenberg, R. Avigur, and N. Shnerb. Quantization process for the driven well — where P fails to commute with P^{2}. Foundations of Physics, 27(2):135–151, February 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02550446.

Eisenberg:1997:LSA

[1967] Eli Eisenberg and Itzhack Dana. Limited sensitivity to analyticity: a manifestation of quantum chaos. Foundations of Physics, 27(2):

153–170, February 1997. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02550447.

Exner:1997:MLG

[1968] Pavel Exner. Magnetoresonances on a lasso graph. Foundations of Physics, 27(2):171–190, February 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02550448.

Eisenberg:1997:NTS

[1969] E. Eisenberg and Y. Ashkenazy. A new time-scale for tunneling. Foundations of Physics, 27(2):191–202, February 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02550449.

Levitan:1997:EDC

[1970] J. Levitan, M. Lewkowicz, and Y. Ashkenazy. Enhancement of decoherence by chaotic-like behavior. Foundations of Physics, 27(2): 203–214, February 1997. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02550450.

Marchand:1997:MPD

[1971] Jean-Paul Marchand. A minimal phenomenology for decay and resonance. Foundations of Physics, 27(2):215–226, February 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02550451.

Bach:1997:MTR

[1972] V. Bach, J. Fröhlich, and I. M. Sigal. Mathematical theory of radiation. Foundations of Physics, 27(2):227–237, February 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02550452.

Petrosky:1997:MNE

[1973] T. Petrosky and M. Rosenberg. Microscopic non-equilibrium structure and dynamical model of entropy flow. Foundations of Physics, 27(2): 239–259, February 1997. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02550453.

Leifer:1997:SEF

[1974] P. Leifer. Superrelativity as an element of a final theory. Foundations of Physics, 27(2):261–285, February 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02550454.

Gudder:1997:ETS

[1975] Stanley Gudder. Effect test spaces and effect algebras. Foundations of Physics, 27(2):287–304, February 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02550455.

Haller:1997:MEG

[1976] Kurt Haller. S-matrix elements for gauge theories with and without implemented constraints. Foundations of Physics, 27(2):305–313, February 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF02550456.

Burakovsky:1997:TRS

[1977] L. Burakovsky. Towards resolution of the scalar meson nonet enigma. Foundations of Physics, 27(2):315–330, February 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02550457.

Anonymous:1997:CPa

[1978] Anonymous. Call for papers. Foundations of Physics, 27(2):331– 332, February 1997. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02550458.

Antoniou:1997:FLC

[1979] I. Antoniou and Z. Suchanecki. The fuzzy logic of chaos and probabilistic inference. Foundations of Physics, 27(3):333–362, March 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02550161.

Droz-Vincent:1997:DIR

[1980] Philippe Droz-Vincent. Direct interactions in relativistic statistical mechanics. Foundations of Physics, 27(3):363–387, March 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02550162.

Trump:1997:MSD

[1981] Matthew A. Trump and William C. Schieve. Off-mass-shell dynamics in flat spacetime. Foundations of Physics, 27(3):389–414, March 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02550163.

Durt:1997:TIV

[1982] Thomas Durt. Three interpretations of the violation of Bell’s inequalities. Foundations of Physics, 27(3):415–434, March 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02550164.

Rodrigues:1997:EUP

[1983] Waldyr A. Rodrigues, Jr. and Jian-Yu Lu. On the existence of undistorted progressive waves (upws) of arbitrary speeds 0 ≤ θ < ∞ in nature. Foundations of Physics, 27(3):435–508, March 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02550165.

Granik:1997:WQM

[1984] A. Granik. Why quantum mechanics indeed? Foundations of Physics, 27(4):511–532, April 1997. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02550675.

Vargas:1997:EKK

[1985] José G. Vargas and Douglas G. Torr. The emergence of a Kaluza–Klein microgenometry from the invariants of optimally Euclidean Lorentzian spaces. Foundations of Physics, 27(4):533–558, April 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02550676.

Plaga:1997:PFE

[1986] R. Plaga. On a possibility to find experimental evidence for the manyworlds interpretation of quantum mechanics. Foundations of Physics, 27(4):559–577, April 1997. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02550677.

Hachemane:1997:IGD

[1987] Mahmoud Hachemane, Mohamed Abdelwahab Benbitour, and Abdallah Smida. Interaction in the geometro-differential conception of extended particles and the Galilei semigroup of trajectories. Foundations of Physics, 27(4):579–594, April 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02550678.

vBorzeszkowski:1997:MPH

[1988] H.-H. v. Borzeszkowski and H.-J. Treder. Mach’s Principle and hidden matter. Foundations of Physics, 27(4):595–603, April 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02550679.

Gudder:1997:PQR

[1989] Stanley Gudder. Perspectives on quantum reality: Non-relativistic, relativistic, and field theoretic. Foundations of Physics, 27(4):605–606, April 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF02550680.

Anonymous:1997:SFM

[1990] Anonymous. Symposium on the foundations of modern physics 1997. Foundations of Physics, 27(4):609–611, April 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF02550681.

Bohm:1997:JBE

[1991] A. Bohm, M. Loewe, P. Patuleanu, and C. Püntmann. Jordan blocks and exponentially decaying higher-order Gamow states. Foundations of Physics, 27(5):613–624, May 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02550171.

Santilli:1997:RHM

[1992] Ruggero Maria Santilli. Relativistic hadronic mechanics: Nonunitary, axiom-preserving completion of relativistic quantum mechanics. Foundations of Physics, 27(5):625–729, May 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02550172.

Unal:1997:SMC

[1993] Nuri Ünal. A simple model of the classical Zitterbewegung: Photon wave function. Foundations of Physics, 27(5):731–746, May 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02550173.

Unal:1997:KCZ

[1994] Nuri Ünal. Kernel of the classical Zitterbewegung. Foundations of Physics, 27(5):747–758, May 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02550174.

Nairz:1997:ECM

[1995] A. Nairz. Errata: a class of metric theories of gravitation of Minkowski spacetime. Foundations of Physics, 27(5):759–762, May 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02550175.

Ferrero:1997:ECS

[1996] Miguel Ferrero and Emilio Santos. Empirical consequences of the scientific construction: The program of local hidden-variables theories in quantum mechanics. Foundations of Physics, 27(6):765–800, June 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02550341.

Mancini:1997:CLD

[1997] Stefano Mancini, Vladimir I. Man’ko, and Paolo Tombest. Classicallike description of quantum dynamics by means of symplectic tomography. Foundations of Physics, 27(6):801–824, June 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02550342.

Vargas:1997:CTF

[1998] José G. Vargas and Douglas G. Torr. The construction of teleparallel Finsler connections and the emergence of an alternative concept of metric compatibility. Foundations of Physics, 27(6):825–843, June 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02550343.

Sutherland:1997:PSG

[1999] Roderick I. Sutherland. Phase space generalization of the de Broglie– Bohm model. Foundations of Physics, 27(6):845–863, June 1997. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02550344.

Meis:1997:PWP

[2000] C. Meis. Photon wave-particle duality and virtual electromagnetic waves. Foundations of Physics, 27(6):865–873, June 1997. CODEN FNDPA4.

ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF02550345.

Treder:1997:MCW

[2001] Hans-Jürgen Treder and Wilfried Schröder. Magnetohydrodynamics corresponding with wave mechanics. Foundations of Physics, 27(6): 875–879, June 1997. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02550346.

Ungar:1997:TPU

[2002] Abraham A. Ungar. Thomas precession: Its underlying gyrogroup axioms and their use in hyperbolic geometry and relativistic physics. Foundations of Physics, 27(6):881–951, June 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02550347.

Jones:1997:PCP

[2003] Roger Jones. Physics and chance: Philosophical issues in the foundations of statistical mechanics. Foundations of Physics, 27(6):953–955, June 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF02550348.

Sukhatme:1997:QMC

[2004] Uday Sukhatme. Quantum mechanics: Classical results, modern systems and visualized examples. Foundations of Physics, 27(6):957–958, June 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF02550349.

Kevan:1997:EPR

[2005] Larry Kevan. Electron paramagnetic resonance: Elementary theory and practical applications. Foundations of Physics, 27(6):959–960, June 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02550350.

Anonymous:1997:CPb

[2006] Anonymous. Call for papers. Foundations of Physics, 27(6):963– 964, June 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02550351.

Greiner:1997:LB

[2007] Walter Greiner. Lawrence Biedenharn. Foundations of Physics, 27 (7):963–967, July 1997. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02551147.

Bohm:1997:TRO

[2008] Arno Bohm and Sujeewa Wickramasekara. The time reversal operator for semigroup evolutions. Foundations of Physics, 27(7):969–993, July 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF02551148.

Danos:1997:WTI

[2009] Michael Danos. Ward–Takahashi identities and Noether’s theorem in quantum field theory. Foundations of Physics, 27(7):995–1009, July 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02551149.

Horwitz:1997:SAQ

[2010] L. P. Horwitz. Schwinger algebra for quaternionic quantum mechanics. Foundations of Physics, 27(7):1011–1034, July 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF02551150.

Asherova:1997:BLH

[2011] R. M. Asherova, J. P. Draayer, Yu. I. Kharitonov, and Yu. F. Smirnov. The Biedenharn–Louck–Hecht resolution of the outer multiplicity problem for the U(3) and U_{q}(3) groups. Foundations of Physics, 27(7): 1035–1046, July 1997. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02551151.

VanIsacker:1997:NMF

[2012] P. Van Isacker, O. Juillet, and B. K. Gjelsten. A nuclear mass formula based on SU(4) symmetry. Foundations of Physics, 27(7): 1047–1060, July 1997. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02551152.

Hess:1997:CSS

[2013] P. O. Hess. Comparative study of the Sp(2, R) and the Sp(6, R) models and an application to the Ba chain of isotopes. Foundations of Physics, 27(7):1061–1081, July 1997. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02551153.

Louck:1997:DSM

[2014] James D. Louck. Doubly stochastic matrices in quantum mechanics. Foundations of Physics, 27(8):1085–1104, August 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02551435.

Neeman:1997:WSC

[2015] Yuval Ne’eman and Djordje Sijacki. World spinors-construction and some applications. Foundations of Physics, 27(8):1105–1122, August 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02551436.

Ludu:1997:QCS

[2016] Andrei Ludu and Walter Greiner. Quasi-continuous symmetries of nonlie type. Foundations of Physics, 27(8):1123–1138, August 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02551437.

Moshinsky:1997:SRC

[2017] M. Moshinsky, A. del Sol Mesa, and V. Riquer. Supermultiplicity and the relativistic Coulomb problem with arbitrary spin. Foundations of Physics, 27(8):1139–1157, August 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02551438.

Santilli:1997:ILA

[2018] Ruggero Maria Santilli. Invariant Lie-admissible formulation of quantum deformations. Foundations of Physics, 27(8):1159–1177, August 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02551439.

Rowe:1997:RCC

[2019] D. J. Rowe and J. Repka. The representations and coupling coefficients of su(n); application to su(4). Foundations of Physics, 27(8): 1179–1209, August 1997. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02551440.

Eisenberg:1997:BRP

[2020] J. M. Eisenberg. Back-reaction in the presence of thermalizing collisions. Foundations of Physics, 27(9):1213–1220, September 1997. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02551524.

Obukhov:1997:CAN

[2021] Yuri N. Obukhov, Eckehard W. Mielke, Jan Budczies, and Friedrich W. Hehl. On the chiral anomaly in non-Riemannian spacetimes. Foundations of Physics, 27(9):1221–1236, September 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02551525.

Matinyan:1997:QFD

[2022] Sergei G. Matinyan and Berndt Müller. Quantum fluctuations and dynamical chaos: an effective potential approach. Foundations of Physics, 27(9):1237–1255, September 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02551526.

Schmid:1997:RMN

[2023] R. N. Schmid, E. Engel, and R. M. Dreizler. Relativistic models for nuclear structure calculations: Comparative study of mean-field and Hartree–Fock approximation for superheavy nuclei. Foundations of Physics, 27(9):1257–1274, September 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02551527.

Simkovic:1997:LHS

[2024] F. Simkovic, J. Schwieger, G. Pantis, and Amand Faessler. A large Hilbert space QRPA and RQRPA calculation of neutrinoless double beta decay. Foundations of Physics, 27(9):1275–1289, September 1997. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02551528.

Solem:1997:VKP

[2025] Johndale C. Solem. Variations on the Kepler problem. Foundations of Physics, 27(9):1291–1306, September 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02551529.

Redei:1997:RCC

[2026] Miklós Rédei. Reichenbach’s common cause principle and quantum field theory. Foundations of Physics, 27(10):1309–1321, October 1997. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02551514.

Cattaneo:1997:CPO

[2027] Gianpiero Cattaneo, Tiziana Marsico, Giuseppe Nisticò, and Guido Bacciagaluppi. A concrete procedure for obtaining sharp reconstructions of unsharp observables in finite-dimensional quantum mechanics. Foundations of Physics, 27(10):1323–1343, October 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF02551515.

Rice:1997:NCL

[2028] Dien A. Rice. Nonlocality and conservation laws in hidden variable theories. Foundations of Physics, 27(10):1345–1353, October 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02551516.

Bana:1997:PKC

[2029] Gergely Bana and Thomas Durt. Proof of Kolmogorovian censorship. Foundations of Physics, 27(10):1355–1373, October 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02551517.

Schroeck:1997:SQT

[2030] F. E. Schroeck, Jr. Symmetry in quantum theory: Implications for the convexity formalism, the measurement problem, and hidden variables. Foundations of Physics, 27(10):1375–1396, October 1997. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02551518.

Cocke:1997:IG

[2031] W. J. Cocke and B. Roy Frieden. Information and gravitation. Foundations of Physics, 27(10):1397–1412, October 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF02551519.

Vstovsky:1997:TIS

[2032] G. V. Vstovsky. Transform information: a symmetry breaking measure. Foundations of Physics, 27(10):1413–1444, October 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02551520.

Caluzi:1997:CAH

[2033] J. J. Caluzi and A. K. T. Assis. A critical analysis of Helmholtz’s argument against Weber’s electrodynamics. Foundations of Physics, 27(10):

1445–1452, October 1997. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02551521.

Schroder:1997:HMA

[2034] Wilfried Schröder and Hans-Jürgen Treder. On Hide’s magnetic analogue of Ertel’s vorticity theorem. Foundations of Physics, 27(10): 1453–1455, October 1997. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02551522.

Phipps:1997:NVE

[2035] Thomas E. Phipps, Jr. Newton versus Einstein: How matter interacts with matter. Foundations of Physics, 27(10):1457–1460, October 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02551523.

Inomata:1997:MNH

[2036] Akira Inomata, John R. Klauder, Ichiro Ohba, and Alwyn van der Merwe. For Mikio Namiki on his seventieth birthday. Foundations of Physics, 27(11):1463–1466, November 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02551493.

Klauder:1997:UQ

[2037] John R. Klauder. Understanding quantization. Foundations of Physics, 27(11):1467–1483, November 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF02551494.

Araki:1997:GKC

[2038] Huzihiro Araki and Takaaki Tabuchi. Generalization of Krinsky’s commutativity proof of transfer matrices with Hamiltonians. Foundations of Physics, 27(11):1485–1494, November 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02551495.

Ezawa:1997:CPE

[2039] Hiroshi Ezawa. Constructive proof of the existence of bound state in one dimension. Foundations of Physics, 27(11):1495–1509, November 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02551496.

Hida:1997:RFQ

[2040] Takeyuki Hida. Random fields and quantum dynamics. Foundations of Physics, 27(11):1511–1518, November 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02551497.

Sakita:1997:RMF

[2041] B. Sakita. Random matrices, fermions, collective fields, and universality. Foundations of Physics, 27(11):1519–1525, November 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02551498.

Selleri:1997:RPN

[2042] F. Selleri. The relativity principle and the nature of time. Foundations of Physics, 27(11):1527–1548, November 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02551499.

Khalfin:1997:UTF

[2043] Leonid A. Khalfin. Unconditional tests of fundamental discrete symmetries CP, T, CPT in rigorous quantum dynamics beyond the approximate Lee–Oehme–Yang theory. Foundations of Physics, 27(11): 1549–1570, November 1997. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02551500.

Li:1997:MWI

[2044] Miao Li and Yong-Shi Wu. Master Ward identity for nonlocal symmetries in D = 2 principal chiral models. Foundations of Physics, 27(11): 1571–1583, November 1997. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02551501.

Yabu:1997:DSV

[2045] Hiroyuki Yabu and Hiroshi Kuratsuji. Dynamics of a superfluid vortex and the Magnus force. Foundations of Physics, 27(11):1585– 1599, November 1997. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02551502.

Hama:1997:FHN

[2046] Yogiro Hama, Takeshi Kodama, and Samya Paiva. Fluctuations in hadronic and nuclear collisions. Foundations of Physics, 27(11):1601– 1621, November 1997. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02551503.

Schulman:1997:WIB

[2047] L. S. Schulman. Watching it boil: Continuous observation for the quantum Zeno effect. Foundations of Physics, 27(12):1623–1636, December 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF02551441.

Mensky:1997:DCM

[2048] Michael B. Mensky. Decoherence in continuous measurements: From models to phenomenology. Foundations of Physics, 27(12):1637– 1654, December 1997. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02551442.

Pascazio:1997:DOQ

[2049] Saverio Pascazio. Dynamical origin of the quantum Zeno effect. Foundations of Physics, 27(12):1655–1670, December 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF02551443.

Beige:1997:AQZ

[2050] Almut Beige, Gerhard C. Hegerfeldt, and Dirk G. Sondermann. Atomic quantum Zeno effect for ensembles and single systems. Foundations of Physics, 27(12):1671–1688, December 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02551444.

Rimini:1997:COP

[2051] Alberto Rimini. Compound objects as particles in quantum mechanics. Foundations of Physics, 27(12):1689–1699, December 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02551445.

Kirillov:1997:TMS

[2052] A. I. Kirillov. On the theory of metastable states. Foundations of Physics, 27(12):1701–1708, December 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF02551446.

Nakazato:1997:TDW

[2053] Hiromichi Nakazato. Time development of a wave packet and the time delay. Foundations of Physics, 27(12):1709–1723, December 1997. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02551447.

Ohba:1997:NMQ

[2054] Ichiro Ohba. A novel method to quantize systems of damped motion. Foundations of Physics, 27(12):1725–1738, December 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02551448.

Okano:1997:STC

[2055] K. Okano, L. Schülke, and B. Zheng. Short-time critical dynamics of statistical systems and field theory. Foundations of Physics, 27(12): 1739–1764, December 1997. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/BF02551449.

Home:1997:CIQ

[2056] Dipankar Home and Guruprasad Kar. Collapse-induced quantum nonlocal effect. Foundations of Physics, 27(12):1765–1770, December 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02551450.

Deotto:1998:BMR

[2057] E. Deotto and G. C. Ghirardi. Bohmian mechanics revisited. Foundations of Physics, 28(1):1–30, January 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1018752202576.

Kamefuchi:1998:SCQ

[2058] S. Kamefuchi. Some considerations on quantum mechanics-matter wave and probability wave. Foundations of Physics, 28(1):31–43, January 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018704319414.

Machida:1998:QMC

[2059] Shigeru Machida and Akio Motoyoshi. Quantum mechanics of the composite system and its subsystems. Foundations of Physics, 28(1):45–57, January 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018756303484.

Tonomura:1998:DOM

[2060] Akira Tonomura. Direct observation of magnetic vortices in superconductors using electron waves. Foundations of Physics, 28(1):59–86, January 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018708420323.

Kimball:1998:SST

[2061] J. C. Kimball. States on the Sierpinski triangle. Foundations of Physics, 28(1):87–105, January 1998. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1018760504393.

Inomata:1998:FAS

[2062] A. Inomata and O. Kizilkaya. Factorization, algebraization, and shape invariance. Foundations of Physics, 28(1):107–124, January 1998. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1018712621231.

Braginsky:1998:GPQ

[2063] V. B. Braginsky. From Galileo’s pendulum to a quantum one (a short review). Foundations of Physics, 28(1):125–130, January 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1018764605301.

Svetlichny:1998:QFS

[2064] George Svetlichny. Quantum formalism with state-collapse and superluminal communication. Foundations of Physics, 28(2):131–155, February 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018726717481.

Schafir:1998:NMW

[2065] R. L. Schafir. Nonlocality in the many-worlds and consistent-histories interpretations. Foundations of Physics, 28(2):157–166, February 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1018744601551.

Hushwater:1998:PQA

[2066] V. Hushwater. A path from the quantization of the action variable to quantum mechanical formalism. Foundations of Physics, 28(2):167–184, February 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018796618389.

Schon:1998:TPO

[2067] Martin Schön. Twin Paradox without one-way velocity assumptions. Foundations of Physics, 28(2):185–204, February 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1018748702460.

Israelit:1998:TWD

[2068] Mark Israelit. Torsional Weyl–Dirac electrodynamics. Foundations of Physics, 28(2):205–229, February 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1018700819298.

Esposito:1998:CMF

[2069] Salvatore Esposito. Covariant Majorana formulation of electrodynamics. Foundations of Physics, 28(2):231–244, February 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1018752803368.

Rylov:1998:QMD

[2070] Yuri A. Rylov. Quantum mechanics as a dynamic construction. Foundations of Physics, 28(2):245–272, February 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018704920207.

vBorzeszkowski:1998:DMV

[2071] H.-H. v. Borzeszkowski and H.-J. Treder. Dark matter versus Mach’s Principle. Foundations of Physics, 28(2):273–290, February 1998. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1018756904277.

Evans:1998:EVN

[2072] A. B. Evans. On expected values and “negative probability” in 4-space QED. Foundations of Physics, 28(2):291–306, February 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1018709021115.

Tavakol:1998:DAG

[2073] Reza Tavakol and Roustam Zalaletdinov. On the domain of applicability of general relativity. Foundations of Physics, 28(2):307–331, February 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018761005186.

Benci:1998:NVP

[2074] Vieri Benci and Donato Fortunato. A new variational principle for the fundamental equations of classical physics. Foundations of Physics, 28 (2):333–352, February 1998. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1018713122024.

Anonymous:1998:BPLa

[2075] Anonymous. Foundations of Physics Letters — list of contents: Volume 10, number 5, October 1997. Foundations of Physics, 28(2):353, February 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018765106094.

Scully:1998:AOB

[2076] Marlan O. Scully. Asim Orhan Barut: 1926–1994. Foundations of Physics, 28(3):355–356, March 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1018717308337.

Dowling:1998:ABP

[2077] Jonathan P. Dowling. Asim Barut: a personal tribute. Foundations of Physics, 28(3):357–359, March 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1018751625175.

Spreeuw:1998:CAE

[2078] Robert J. C. Spreeuw. A classical analogy of entanglement. Foundations of Physics, 28(3):361–374, March 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1018703709245.

Englert:1998:CAU

[2079] Berthold-Georg Englert. Classical analogs of unitarily equivalent Hamilton operators. Foundations of Physics, 28(3):375–384, March 1998. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1018755726084.

Steinberg:1998:SPN

[2080] Aephraim M. Steinberg. Single-particle nonlocality and conditional measurements. Foundations of Physics, 28(3):385–397, March 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1018707810154.

Scully:1998:OAQ

[2081] Marlan O. Scully and Herbert Walther. An operational analysis of quantum eraser and delayed choice. Foundations of Physics, 28(3):399–413, March 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018759926992.

Bozic:1998:QIQ

[2082] Mirjana Bozić and Zvonko Marić. Quantum interference, quantum theory of measurement, and (in)completeness of quantum mechanics. Foundations of Physics, 28(3):415–427, March 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018712011063.

Manko:1998:DST

[2083] V. I. Man’ko. Dynamical symmetries and tomography. Foundations of Physics, 28(3):429–438, March 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1018764027901.

Jaekel:1998:CSQ

[2084] Marc-Thierry Jaekel and Serge Reynaud. Conformal symmetry and quantum relativity. Foundations of Physics, 28(3):439–456, March 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1018716111971.

Crawford:1998:SMG

[2085] James P. Crawford. Spinor matter in a gravitational field: Covariant equations à la Heisenberg. Foundations of Physics, 28(3):457–470, March 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018768128810.

Barashenkov:1998:MSD

[2086] V. S. Barashenkov. Mechanics in six-dimensional spacetime. Foundations of Physics, 28(3):471–484, March 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1018720212880.

Ralph:1998:BGS

[2087] J. F. Ralph, T. D. Clark, R. J. Prance, H. Prance, and J. Diggins. Broken gauge symmetry in macroscopic quantum circuits. Foundations of Physics, 28(3):485–503, March 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1018772229718.

Melloy:1998:PBD

[2088] G. F. Melloy and A. J. Bracken. Probability backflow for a Dirac particle. Foundations of Physics, 28(3):505–514, March 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1018724313788.

Ranfagni:1998:TSP

[2089] A. Ranfagni, R. Ruggeri, and A. Agresti. Tunneling as a stochastic process. Foundations of Physics, 28(3):515–525, March 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1018776330627.

Ahluwalia:1998:BRB

[2090] D. V. Ahluwalia. Book review: Quantum Field Theory, Second Edition, by Lewis H. Ryder. Foundations of Physics, 28(3):527–529, March 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1018728414697.

Gea-Banacloche:1998:ECR

[2091] Julio Gea-Banacloche. Emergence of classical radiation fields through decoherence in the Scully–Lamb laser model. Foundations of Physics, 28(4):531–548, April 1998. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1018729904161.

Busch:1998:TCA

[2092] Thomas Busch, Berthold-Georg Englert, Kazimierz Rzazewski, and Martin Wilkens. Two cold atoms in a harmonic trap. Foundations of Physics, 28(4):549–559, April 1998. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1018705520999.

Kocharovskaya:1998:SEG

[2093] Olga Kocharovskaya and Y. V. Radeonychev. Spontaneous emission from the ground atomic state due to its crossing with the dynamic Stark level. Foundations of Physics, 28(4):561–584, April 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1018757505070.

Lee:1998:PLT

[2094] Hwang Lee and Marlan O. Scully. The physics of EIT and LWI in V - type configurations. Foundations of Physics, 28(4):585–600, April 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1018709621908.

Zakowicz:1998:SEA

[2095] Wladysław Zakowicz and Arkadiusz Orlowski. Spontaneously emitting atom in front of a two-slit interferometer. Foundations of Physics, 28 (4):601–609, April 1998. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1018761605978.

Zhu:1998:SNL

[2096] Shi-Yao Zhu, Dmitri E. Nikonov, and Marlan O. Scully. A scheme for noninversion lasing for short–wavelength lasers in heliumlike ions. Foundations of Physics, 28(4):611–619, April 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018713722817.

Welch:1998:OTE

[2097] George R. Welch, G. G. Padmabandu, Edward S. Fry, Mikhail D. Lukin, Dmitri E. Nikonov, Frank Sander, Marlan O. Scully, Antoin Weis, and Frank K. Tittel. Observation of V -type electromagnetically induced transparency in a sodium atomic beam. Foundations of Physics, 28 (4):621–638, April 1998. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1018765706887.

Clerk:1998:NRW

[2098] A. A. Clerk and J. E. Sipe. Nonlocality and the rotating wave approximation. Foundations of Physics, 28(4):639–651, April 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1018717823725.

Salamin:1998:NSF

[2099] Yousef I. Salamin. Nonrelativistic strong-field photoionization without making the dipole approximation. Foundations of Physics, 28(4):653– 665, April 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018769807796.

Rusek:1998:BLE

[2100] Marian Rusek and Arkadiusz Orlowski. Bands of localized electromagnetic waves in 3D random media. Foundations of Physics, 28(4):667– 681, April 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018721924634.

vanFraassen:1998:BRB

[2101] Bas C. van Fraassen. Book review: Interpreting the Quantum World by Jeffrey Bub. Foundations of Physics, 28(4):683–689, April 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1018773908704.

Moniz:1998:BRE

[2102] P. Vargas Moniz. Book review: Euclidean Quantum Gravity on Manifolds with Boundary by Giampiero Esposito, Alexander Yu. Kamenshchik, and Giuseppe Pollifrone. Foundations of Physics, 28(4):691–694, April 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1018726125542.

Wightman:1998:BRB

[2103] Arthur Wightman. Book review: A Tale of Two Continents: a Physicist’s Life in a Turbulent World by Abraham Pais. Foundations of Physics, 28(4):695–696, April 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1018778109613.

Anonymous:1998:BPLb

[2104] Anonymous. Foundations of Physics Letters — list of contents: Volume 11, number 1, February 1998. Foundations of Physics, 28(4):697, April 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018730226451.

Barut:1998:EMT

[2105] Asim O. Barut and Walter Wyss. The energy–momentum tensor for electromagnetic interactions. Foundations of Physics, 28(5):699–715, May 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018889518158.

Kapuscik:1998:SIE

[2106] Edward Kapuścik. Self-induced electrostatic fields. Foundations of Physics, 28(5):717–727, May 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1018841602228.

Inomata:1998:RMT

[2107] Akira Inomata, Georg Junker, and Claudia Rosch. Remarks on the magnetic top. Foundations of Physics, 28(5):729–739, May 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1018893619066.

Uzes:1998:SFE

[2108] C. A. Uzes and A. O. Barut. Spinor field as elementary excitations of a system of scalar fields. Foundations of Physics, 28(5):741–754, May 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1018845703137.

Unal:1998:PIQ

[2109] Nuri Ünal. Path integral quantization of a spinning particle. Foundations of Physics, 28(5):755–762, May 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1018897719975.

Salesi:1998:VFO

[2110] Giovanni Salesi and Erasmo Recami. A velocity field and operator for spinning particles in (nonrelativistic) quantum mechanics. Foundations of Physics, 28(5):763–773, May 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1018849804045.

Audretsch:1998:QPB

[2111] Jürgen Audretsch and Vladimir D. Skarzhinsky. Quantum processes beyond the Aharonov–Bohm effect. Foundations of Physics, 28(5):777– 788, May 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018801920883.

Rodrigues:1998:ERQ

[2112] W. A. Rodrigues, Jr. and J. Vaz, Jr. From electromagnetism to relativistic quantum mechanics. Foundations of Physics, 28(5):789–814, May 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018854004954.

Acikgoz:1998:VPS

[2113] I. Açikgöz and N. Ünal. Vacuum polarization in self-field quantum electrodynamics. Foundations of Physics, 28(5):815–828, May 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1018806121792.

Vairo:1998:BRS

[2114] Antonio Vairo. The Barbieri–Remiddi solution of the bound-state problem in QED. Foundations of Physics, 28(5):829–841, May 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1018858105862.

Carati:1998:PPC

[2115] A. Carati. Pair production in classical electrodynamics. Foundations of Physics, 28(5):843–853, May 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1018810222701.

Dowling:1998:CLS

[2116] Jonathan P. Dowling. The classical Lamb shift: Why Jackson is wrong! Foundations of Physics, 28(5):855–862, May 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1018862206771.

Anonymous:1998:BPLc

[2117] Anonymous. Foundations of Physics Letters — list of contents: Volume 11, number 1, February 1998. Foundations of Physics, 28(5): 863, May 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018814323609.

Nussinov:1998:REM

[2118] S. Nussinov. Realistic experiments for measuring the wave function of a single particle. Foundations of Physics, 28(6):865–880, June 1998. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1018891028578.

Holland:1998:NTI

[2119] P. R. Holland. New trajectory interpretation of quantum mechanics. Foundations of Physics, 28(6):881–911, June 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1018813112648.

Maroli:1998:MBC

[2120] C. Maroli and M. Cornelli. Microscopic behavior of the classical electron in the absence of external forces. Foundations of Physics, 28(6):913– 929, June 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018865129486.

Vargas:1998:TKC

[2121] Jose G. Vargas and Douglas G. Torr. Teleparallel Kähler calculus for spacetime. Foundations of Physics, 28(6):931–958, June 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1018817213557.

Paul:1998:RIF

[2122] Harry Paul and Mladen Pavicić. Realistic interaction–free detection of objects in a resonator. Foundations of Physics, 28(6):959–970, June 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1018869230395.

Krenn:1998:STQ

[2123] G. Krenn and K. Svozil. Stronger-than-quantum correlations. Foundations of Physics, 28(6):971–984, June 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018821314465.

Shirokov:1998:SSD

[2124] M. I. Shirokov. Spin state determination using a Stern–Gerlach device. Foundations of Physics, 28(6):985–997, June 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1018873331304.

Beker:1998:PWM

[2125] Haluk Beker. SU(1, 1) perturbations of the s-wave Morse potential problem. Foundations of Physics, 28(6):999–1004, June 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1018825415374.

Kak:1998:QID

[2126] Subhash C. Kak. Quantum information in a distributed apparatus. Foundations of Physics, 28(6):1005–1012, June 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018877532212.

Schroder:1998:ECP

[2127] Wilfried Schröder and Hans-Jürgen Treder. Einstein and cosmic physics. Foundations of Physics, 28(6):1013–1020, June 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1018829616282.

Lounesto:1998:BRB

[2128] Pertti Lounesto. Book review: Clifford Algebra: a Computational Tool for Physicists, by John Snygg. Foundations of Physics, 28(6):1021, June 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018881600353.

Anonymous:1998:Ea

[2129] Anonymous. Erratum. Foundations of Physics, 28(6):1023–1025, June 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018833717191.

Anonymous:1998:Eb

[2130] Anonymous. Erratum. Foundations of Physics, 28(6):1027, June 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1018885701261.

Anonymous:1998:BPLd

[2131] Anonymous. Foundations of Physics Letters — list of contents: Volume 11, number 1, February 1998. Foundations of Physics, 28(6):1029, June 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018837818100.

Rovelli:1998:ITI

[2132] Carlo Rovelli. “Incerto tempore, incertisque loci”: Can we compute the exact time at which a quantum measurement happens? Foundations of Physics, 28(7):1031–1043, July 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1018889802170.

Rohrlich:1998:ATE

[2133] Fritz Rohrlich. The arrow of time in the equations of motion. Foundations of Physics, 28(7):1045–1056, July 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1018841919008.

Rueda:1998:CIM

[2134] Alfonso Rueda and Bernhard Haisch. Contribution to inertial mass by reaction of the vacuum to accelerated motion. Foundations of Physics, 28(7):1057–1108, July 1998. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1018893903079.

Coecke:1998:RCQ

[2135] Bob Coecke. A representation for compound quantum systems as individual entities: Hard acts of creation and hidden correlations. Foundations of Physics, 28(7):1109–1135, July 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018846019917.

Good:1998:PSI

[2136] R. H. Good. Particle spectrum implied by the Dirac equation. Foundations of Physics, 28(7):1137–1156, July 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018898003987.

Sutherland:1998:DFQ

[2137] Roderick I. Sutherland. Density formalism for quantum theory. Foundations of Physics, 28(7):1157–1190, July 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018850120826.

Burns:1998:EVR

[2138] Jean E. Burns. Entropy and vacuum radiation. Foundations of Physics, 28(7):1191–1207, July 1998. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1018802204896.

Anonymous:1998:BPLe

[2139] Anonymous. Foundations of Physics Letters — list of contents: Volume 11, number 2, April 1998. Foundations of Physics, 28(7):1209, July 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1018854221734.

Trump:1998:CSC

[2140] M. A. Trump and W. C. Schieve. Classical scattering in the covariant two-body Coulomb potential. Foundations of Physics, 28(8):1211–1236, August 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018806308530.

Ridderbos:1998:SEE

[2141] T. M. Ridderbos and M. L. G. Redhead. The spin–echo experiments and the Second Law of Thermodynamics. Foundations of Physics, 28(8): 1237–1270, August 1998. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1018870725369.

Ralph:1998:STD

[2142] J. F. Ralph, T. D. Clark, H. Prance, R. J. Prance, A. Widom, and Y. N. Srivastava. Solutions of the time-dependent Schrödinger equation for a two-state system. Foundations of Physics, 28(8):1271–1282, August 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1018822809439.

Ungar:1998:PEH

[2143] Abraham A. Ungar. From Pythagoras to Einstein: The hyperbolic Pythagorean theorem. Foundations of Physics, 28(8):1283–1321, August 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018874826277.

Burgos:1998:WNP

[2144] M. E. Burgos. Which natural processes have the special status of measurements? Foundations of Physics, 28(8):1323–1346, August 1998. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1018826910348.

Coecke:1998:RSE

[2145] Bob Coecke. A representation for a spin-S entity as a compound system in R^{3} consisting of 2S individual spin-1/2 entities. Foundations of Physics, 28(8):1347–1365, August 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1018878927186.

Smida:1998:GDC

[2146] A. Smida, M. Hachemane, and A.-H. Hamici. Geometro-differential conception of extended particles and the semigroup of trajectories in Minkowski space–time. Foundations of Physics, 28(8):1367–1381, August 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018831011256.

Bar:1998:FPI

[2147] D. Bar. The Feynman path integrals and Everett’s universal wave function. Foundations of Physics, 28(8):1383–1391, August 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1018883028094.

Dariescu:1998:ZEM

[2148] Ciprian Dariescu and Marina-Aura Dariescu. Zero energy modes of massless fermions in S^{3}R spacetime. Foundations of Physics, 28(8):1393–1397, August 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018835112165.

Anonymous:1998:BPLf

[2149] Anonymous. Foundations of Physics Letters — list of contents: Volume 11, number 2, April 1998. Foundations of Physics, 28(8):1399, August 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018887129003.

Fanchi:1998:FCT

[2150] John R. Fanchi. Foreword: Changing times. Foundations of Physics, 28(9):1401–1405, September 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1018839325794.

Trump:1998:PPS

[2151] M. A. Trump and W. C. Schieve. Perihelion precession in the special relativistic two-body problem. Foundations of Physics, 28(9):1407–1416, September 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018848909864.

Pitts:1998:PGR

[2152] J. Brian Pitts and W. C. Schieve. On parametrized general relativity. Foundations of Physics, 28(9):1417–1424, September 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1018801126703.

Salisbury:1998:CCG

[2153] D. C. Salisbury. Classical canonical general coordinate and gauge symmetries. Foundations of Physics, 28(9):1425–1431, September 1998. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1018853110773.

Salisbury:1998:CQR

[2154] D. C. Salisbury. A classical and quantum relativistic interacting variablemass model. Foundations of Physics, 28(9):1433–1442, September 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1018805227611.

Pavsic:1998:FRT

[2155] Matej Pavsic. Formulation of a relativistic theory without constraints. Foundations of Physics, 28(9):1443–1451, September 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1018857211681.

Pavsic:1998:PFT

[2156] Matej Pavsic. Parametrized field theory. Foundations of Physics, 28(9): 1453–1464, September 1998. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1018809328520.

Pavsic:1998:QGI

[2157] Matej Pavsic. Quantum gravity induced from unconstrained membranes. Foundations of Physics, 28(9):1465–1477, September 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1018861312590.

Land:1998:PME

[2158] M. C. Land. Pre-Maxwell electrodynamics. Foundations of Physics, 28(9):1479–1487, September 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1018813429428.

Land:1998:CCP

[2159] M. C. Land. The classical Coulomb problem in pre-Maxwell electrodynamics. Foundations of Physics, 28(9):1489–1497, September 1998. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1018865413499.

Land:1998:PMQ

[2160] M. C. Land. Pre-Maxwell quantum electrodynamics. Foundations of Physics, 28(9):1499–1506, September 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018817530337.

Anonymous:1998:BPLg

[2161] Anonymous. Foundations of Physics Letters — list of contents: Volume 11, number 3, June 1998. Foundations of Physics, 28(9):1507, September 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018869514407.

Horwitz:1998:SQS

[2162] L. P. Horwitz and N. Shnerb. Second quantization of the Stueckelberg relativistic quantum theory and associated gauge fields. Foundations of Physics, 28(10):1509–1519, October 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018841000237.

Fanchi:1998:MON

[2163] John R. Fanchi. The mass operator and neutrino oscillations. Foundations of Physics, 28(10):1521–1528, October 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1018830117075.

Alvarez:1998:FPT

[2164] Edgardo T. Garcia Alvarez and Fabian H. Gaioli. Feynman’s proper time approach to QED. Foundations of Physics, 28(10):1529–1538, October 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018882101146.

Gaioli:1998:CQT

[2165] Fabián H. Gaioli and Edgardo T. Garcia Alvarez. Classical and quantum theories of spin. Foundations of Physics, 28(10):1539–1550, October 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018834217984.

Huang:1998:FSL

[2166] Young-Sea Huang. Formulation of Schrödinger-like relativistic wave equation of motion. Foundations of Physics, 28(10):1551–1559, October 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018886302054.

Gill:1998:CPT

[2167] Tepper L. Gill. Canonical proper-time Dirac theory. Foundations of Physics, 28(10):1561–1575, October 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1018838418892.

Burakovsky:1998:RSM

[2168] L. Burakovsky. Relativistic statistical mechanics and particle spectroscopy. Foundations of Physics, 28(10):1577–1594, October 1998. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1018890402963.

Burakovsky:1998:GSR

[2169] L. Burakovsky. Glueball spectroscopy in Regge phenomenology. Foundations of Physics, 28(10):1595–1605, October 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1018842519801.

Horwitz:1998:DUS

[2170] L. P. Horwitz and Y. Strauss. Description of unstable systems in relativistic quantum mechanics in the Lax–Phillips theory. Foundations of Physics, 28(10):1607–1616, October 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018894503871.

Anonymous:1998:BPLh

[2171] Anonymous. Foundations of Physics Letters — list of contents: Volume 11, number 3, June 1998. Foundations of Physics, 28(10):1617, October 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018846620710.

Wiseman:1998:EHM

[2172] H. M. Wiseman. Extending Heisenberg’s measurement-disturbance relation to the twin-slit case. Foundations of Physics, 28(11):1619–1631, November 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018889508782.

Shirai:1998:RQM

[2173] Hisato Shirai. Reinterpretation of quantum mechanics based on the statistical interpretation. Foundations of Physics, 28(11):1633–1662, November 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018841625620.

Rizzi:1998:SLR

[2174] G. Rizzi and A. Tartaglia. Speed of light on rotating platforms. Foundations of Physics, 28(11):1663–1683, November 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1018893609690.

Matolcsi:1998:SRF

[2175] T. Matolcsi. Spacetime without reference frames: an application to synchronizations on a rotating disk. Foundations of Physics, 28(11): 1685–1701, November 1998. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1018845726529.

Mould:1998:CQM

[2176] Richard A. Mould. Consciousness and quantum mechanics. Foundations of Physics, 28(11):1703–1718, November 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018897710599.

Yourgrau:1998:CDG

[2177] Palle Yourgrau. Comments on “Did Gödel Surprise Einstein with a Rotating Universe and Time Travel?” by Giora Hon. Foundations of Physics, 28(11):1719–1727, November 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018849827437. See [1869].

Gudder:1998:BRB

[2178] Stanley P. Gudder. Book review: Quantum Logic in Algebraic Approach, by Milklós Rédei. Foundations of Physics, 28(11):1729–1730, November 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018801911508.

Anonymous:1998:BPLi

[2179] Anonymous. Foundations of Physics Letters — list of contents: Volume 11, number 4, August 1998. Foundations of Physics, 28(11):1733, November 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018853928346.

Anonymous:1998:Aa

[2180] Anonymous. Announcement. Foundations of Physics, 28(11):1735–1736, November 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018806012416.

Anonymous:1998:IA

[2181] Anonymous. Important announcement. Foundations of Physics, 28 (12):1737, December 1998. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1018858002847.

Wan:1998:CSS

[2182] K. Kong Wan, Jason Bradshaw, Colin Trueman, and F. E. Harrison. Classical systems, standard quantum systems, and mixed quantum systems in Hilbert space. Foundations of Physics, 28(12):1739–1783, December 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018838919685.

Mehra:1998:JWG

[2183] Jagdish Mehra. Josiah Willard Gibbs and the foundations of statistical mechanics. Foundations of Physics, 28(12):1785–1815, December 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1018890903755.

Mugur-Schachter:1998:BRM

[2184] Mioara Mugur-Schächter. Book review: Mécanique quantique: une introduction philosophique, by Michel Bitbol. Foundations of Physics, 28(12): 1817–1824, December 1998. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1018843020594.

Anonymous:1998:BPLj

[2185] Anonymous. Foundations of Physics Letters — list of contents, volume 11, number 4, August 1998. Foundations of Physics, 28(12):1825, December 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018895104664.

Anonymous:1998:Ab

[2186] Anonymous. Announcement. Foundations of Physics, 28(12):1827–1828, December 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018847221502.

Benci:1999:QPC

[2187] Vieri Benci. Quantum phenomena in a classical model. Foundations of Physics, 29(1):1–28, January 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1018899202364.

Boudet:1999:RER

[2188] Roger Boudet. The relativistic expressions with retardation of the matrix elements used in the hydrogenic atomic transitions: application to the photoeffect and the Lamb shift. Foundations of Physics, 29(1):29–47, January 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018861019202.

Boudet:1999:RCM

[2189] Roger Boudet. A relativistic calculation of the matrix elements used in the photoeffect and the Lamb shift. Foundations of Physics, 29(1):49–66, January 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018813103272.

DeBaere:1999:LBT

[2190] W. De Baere, A. Mann, and M. Revzen. Locality and Bell’s theorem. Foundations of Physics, 29(1):67–77, January 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1018865120111.

Goldstein:1999:IOA

[2191] Stanisław Goldstein, Andrzej Luczak, and Ivan F. Wilde. Independence in operator algebras. Foundations of Physics, 29(1):79–89, January 1999.

CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1018817204181.

Mehra:1999:PHQ

[2192] Jagdish Mehra and Helmut Rechenberg. Planck’s half-quanta: a history of the concept of zero-point energy. Foundations of Physics, 29 (1):91–132, January 1999. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1018869221019.

Saslow:1999:BRB

[2193] Wayne M. Saslow. Book review: Classical Electrodynamics, by John David Jackson. Foundations of Physics, 29(1):133–135, January 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1018821305089.

Anonymous:1999:BPLa

[2194] Anonymous. Foundations of Physics Letters — list of contents: Volume 11, number 5, October 1998. Foundations of Physics, 29(1):137, January 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018873321928.

Anonymous:1999:Aa

[2195] Anonymous. Announcements. Foundations of Physics, 29(1):139–141, January 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018825405998.

Anonymous:1999:AIVa

[2196] Anonymous. Author index to volume 28. Foundations of Physics, 29 (1):143–144, January 1999. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1018877422836.

Vargas:1999:CEU

[2197] Jose G. Vargas and Douglas G. Torr. The Cartan–Einstein unification with teleparallelism and the discrepant measurements of Newton’s constant G. Foundations of Physics, 29(2):145–200, February 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1018840720961.

Fernandes:1999:GPS

[2198] M. C. B. Fernandes and J. D. M. Vianna. On the generalized phase space approach to Duffin–Kemmer–Petiau particles. Foundations of Physics, 29 (2):201–219, February 1999. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1018869505031.

Nistico:1999:CCP

[2199] Giuseppe Nisticò. Consistency conditions for probabilities of quantum histories. Foundations of Physics, 29(2):221–229, February 1999. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1018821621870.

Alonso:1999:TRQ

[2200] Vidal Alonso, Salvatore De Vincenzo, and Luigi Mondino. Tensorial relativistic quantum mechanics in (1 + 1) dimensions and boundary conditions. Foundations of Physics, 29(2):231–250, February 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1018873605940.

Lanciani:1999:MED

[2201] Paolo Lanciani. A model of the electron in a 6-dimensional spacetime. Foundations of Physics, 29(2):251–265, February 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1018825722778.

Kak:1999:IPQ

[2202] Subhash Kak. The initialization problem in quantum computing. Foundations of Physics, 29(2):267–279, February 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1018877706849.

Meessen:1999:SQE

[2203] A. Meessen. Spacetime quantization, elementary particles, and cosmology. Foundations of Physics, 29(2):281–316, February 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1018829823687.

Anonymous:1999:BPLb

[2204] Anonymous. Foundations of Physics Letters — list of contents: Volume 11, number 5, October 1998. Foundations of Physics, 29(2):317, February 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018881807757.

Anonymous:1999:Ab

[2205] Anonymous. Announcements. Foundations of Physics, 29(2):319–321, February 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018833924596.

Anonymous:1999:DGH

[2206] Anonymous. For Daniel Greenberger on his 65th birthday. Foundations of Physics, 29(3):323–324, March 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1018885913209.

Allman:1999:QPS

[2207] B. E. Allman, A. Cimmino, S. L. Griffin, and A. G. Klein. Quantum phase shift caused by spatial confinement. Foundations of Physics, 29 (3):325–332, March 1999. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1018858630047.

Anandan:1999:QMP

[2208] J. Anandan. The quantum measurement problem and the possible role of the gravitational field. Foundations of Physics, 29(3):333–348, March 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018810714118.

Calude:1999:EQU

[2209] Cristian S. Calude, Peter H. Hertling, and Karl Svozil. Embedding quantum Universes in classical ones. Foundations of Physics, 29(3):349–379, March 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018862730956.

Felber:1999:NTI

[2210] J. Felber, R. Gähler, R. Golub, P. Hank, V. Ignatovich, T. Keller, and U. Rauch. Neutron time interferometry. Foundations of Physics, 29 (3):381–396, March 1999. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1018814815026.

Greenberg:1999:RCP

[2211] O. W. Greenberg and Robert C. Hilborn. The relation of constraints on particle statistics for different species of particles. Foundations of Physics, 29(3):397–407, March 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1018866831865.

Gudder:1999:UGE

[2212] Stanley Gudder. Universal groups of effect spaces. Foundations of Physics, 29(3):409–422, March 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1018818915935.

Horne:1999:SIG

[2213] M. A. Horne. Schrödinger interferometry in a gravity field as seen from free fall. Foundations of Physics, 29(3):423–433, March 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1018870900005.

Jaeger:1999:EPN

[2214] Gregg Jaeger and Abner Shimony. An extremum principle for a neutron diffraction experiment. Foundations of Physics, 29(3):435–444, March 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018823116843.

Leggett:1999:STE

[2215] A. J. Leggett. Some thought–experiments involving macrosystems as illustrations of various interpretations of quantum mechanics. Foundations of Physics, 29(3):445–456, March 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018875100914.

Namiki:1999:DWC

[2216] Mikio Namiki. Decoherence and wavefunction collapse in quantum measurements. Foundations of Physics, 29(3):457–464, March 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1018827217752.

Pearle:1999:CCM

[2217] Philip Pearle, James Ring, Juan I. Collar, and Frank T. Avignone III. The CSL collapse model and spontaneous radiation: an update. Foundations of Physics, 29(3):465–480, March 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018879201822.

Peshkin:1999:FFI

[2218] Murray Peshkin. Force-free interactions and nondispersive phase shifts in interferometry. Foundations of Physics, 29(3):481–489, March 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1018831318661.

Tisza:1999:UUQ

[2219] Laszlo Tisza. Unreasonable and “unreasonable” in quantum mechanics? Foundations of Physics, 29(3):491–495, March 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1018883302731.

Anonymous:1999:BPLc

[2220] Anonymous. Foundations of Physics Letters — list of contents: Volume 11, number 6, December 1998. Foundations of Physics, 29(3):497, March 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018835419569.

Anonymous:1999:Ac

[2221] Anonymous. Announcements. Foundations of Physics, 29(3):499–501, March 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018887403640.

Bergou:1999:EFM

[2222] János A. Bergou. Entangled fields in multiple cavities as a testing ground for quantum mechanics. Foundations of Physics, 29(4):503–519, April 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018839507274.

Bernstein:1999:SVG

[2223] Herbert J. Bernstein. Simple version of the Greenberger–Horne–Zeilinger (GHZ) argument against local realism. Foundations of Physics, 29 (4):521–525, April 1999. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1018856024112.

Carteret:1999:ME

[2224] H. A. Carteret, N. Linden, S. Popescu, and A. Sudbery. Multiparticle entanglement. Foundations of Physics, 29(4):527–552, April 1999. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1018808108183.

Chiao:1999:RLW

[2225] Raymond Y. Chiao and John C. Garrison. Realism or locality: Which should we abandon? Foundations of Physics, 29(4):553–560, April 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1018860125021.

Fivel:1999:ME

[2226] Daniel I. Fivel. Multiparticle entanglement. Foundations of Physics, 29(4):561–570, April 1999. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1018812209091.

Mermin:1999:WDC

[2227] N. David Mermin. What do these correlations know about reality? Nonlocality and the absurd. Foundations of Physics, 29(4):571–587, April 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018864225930.

Peres:1999:ABI

[2228] Asher Peres. All the Bell inequalities. Foundations of Physics, 29(4):589– 614, April 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018816310000.

Vaidman:1999:VTG

[2229] Lev Vaidman. Variations on the theme of the Greenberger–Horne– Zeilinger proof. Foundations of Physics, 29(4):615–630, April 1999. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1018868326838.

Zeilinger:1999:FPQ

[2230] Anton Zeilinger. A foundational principle for quantum mechanics. Foundations of Physics, 29(4):631–643, April 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018820410908.

Anonymous:1999:SBD

[2231] Anonymous. Selected bibliography: Daniel Greenberger. Foundations of Physics, 29(4):645–647, April 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1018872427747.

Anonymous:1999:BPLd

[2232] Anonymous. Foundations of Physics Letters — list of contents: Volume 11, number 6, December 1998. Foundations of Physics, 29(4):649, April 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018824511817.

Anonymous:1999:Ad

[2233] Anonymous. Announcements. Foundations of Physics, 29(4):651–653, April 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018876528655.

Uffink:1999:TUR

[2234] Jos Uffink and Janneke van Lith. Thermodynamic uncertainty relations. Foundations of Physics, 29(5):655–692, May 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1018811305766.

Blanco:1999:HEA

[2235] R. Blanco. On a hypothetical explanation of the Aharonov–Bohm effect. Foundations of Physics, 29(5):693–720, May 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1018870322605.

Home:1999:IBA

[2236] Dipankar Home and A. S. Majumdar. On the importance of the Bohmian approach for interpreting CP-violation experiments. Foundations of Physics, 29(5):721–727, May 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1018822406675.

Assis:1999:EFO

[2237] A. K. T. Assis, W. A. Rodrigues, Jr., and A. J. Mania. The electric field outside a stationary resistive wire carrying a constant current. Foundations of Physics, 29(5):729–753, May 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018874523513.

Vaidman:1999:TSC

[2238] Lev Vaidman. Time-symmetrized counterfactuals in quantum theory. Foundations of Physics, 29(5):755–765, May 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1018826607583.

Mehra:1999:YJSa

[2239] Jagdish Mehra, Kimball A. Milton, and Peter Rembiesa. The young Julian Schwinger. I. a New York City childhood. Foundations of Physics, 29(5):767–786, May 1999. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1018878624422.

Mehra:1999:YJSb

[2240] Jagdish Mehra, Kimball A. Milton, and Peter Rembiesa. The young Julian Schwinger. II. Julian Schwinger at Columbia University. Foundations of Physics, 29(5):787–817, May 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018830708492.

Woodward:1999:BRB

[2241] James F. Woodward. Book review: Doubt and Certainty, by Tony Rothman and George Sudarshan. Foundations of Physics, 29(5):819– ??, May 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018882725330.

Englert:1999:BRB

[2242] Berthold-Georg Englert. Book review: Quantum Optics, by Marlan O. Scully and M. Suhail Zubairy. Foundations of Physics, 29(5):829– 831, May 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018834809401.

vBorzeszkowski:1999:BRB

[2243] Horst-Heino v. Borzeszkowski. Book review: Open Questions in Relativistic Physics, edited by Franco Selleri. Foundations of Physics, 29 (5):833–835, May 1999. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1018886826239.

Garcia-Colin:1999:BRB

[2244] L. S. García-Colín. Book review: Quantum Statistical Theory of Superconductivity, by Shigeji Fujita and Salvador Godoy. Foundations of Physics, 29(5):837–839, May 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1018838910309.

Garg:1999:BRB

[2245] Anupam Garg. Book review: Classical Dynamics: a Contemporary Approach, by Jorge V. José and Eugene J. Saletan. Foundations of Physics, 29(5):841–843, May 1999. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1018890927148.

Anonymous:1999:BPLe

[2246] Anonymous. Foundations of Physics Letters — list of contents: Volume 12, number 1, February 1999. Foundations of Physics, 29(5): 845, May 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018843011218.

Anonymous:1999:Ae

[2247] Anonymous. Announcements. Foundations of Physics, 29(5):847–849, May 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018895028056.

Kastner:1999:TBP

[2248] R. E. Kastner. The three-box “paradox” and other reasons to reject the counterfactual usage of the ABL rule. Foundations of Physics, 29 (6):851–863, June 1999. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1018865010735. See reply [2249].

Vaidman:1999:MER

[2249] Lev Vaidman. The meaning of elements of reality and quantum counterfactuals: Reply to Kastner. Foundations of Physics, 29(6):865–876, June 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018817227573. See [2248].

Gudder:1999:OSM

[2250] Stan Gudder. Observables and statistical maps. Foundations of Physics, 29(6):877–897, June 1999. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1018869211643.

Woodward:1999:WCI

[2251] James F. Woodward and Thomas Mahood. What is the cause of inertia? Foundations of Physics, 29(6):899–930, June 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1018821328482. See reply

[2309].

Mehra:1999:YJSc

[2252] Jagdish Mehra, Kimball A. Milton, and Peter Rembiesa. The young Julian Schwinger. III. Schwinger goes to Berkeley. Foundations of Physics, 29(6):931–966, June 1999. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1018873312552.

Mehra:1999:YJSd

[2253] Jagdish Mehra, Kimball A. Milton, and Peter Rembiesa. The young Julian Schwinger. IV. During the Second World War. Foundations of Physics, 29(6):967–1010, June 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1018825429390.

Barrett:1999:BRB

[2254] Jeffrey A. Barrett. Book review: Quantum Chance and Non-Locality, by Michael Dickson. Foundations of Physics, 29(6):1011–1018, June 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1018877413461.

Kafatos:1999:BRB

[2255] Menas Kafatos. Book review: Statistical Geometry and Applications to Microphysics and Cosmology, by Sisir Roy. Foundations of Physics, 29 (6):1017–1018, June 1999. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1018829530299.

Anonymous:1999:BPLf

[2256] Anonymous. Foundations of Physics Letters — list of contents, volume 12, number 2, April 1999. Foundations of Physics, 29(6):1019, June 1999.

CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1018881514369.

Anonymous:1999:Af

[2257] Anonymous. Announcement. Foundations of Physics, 29(6):1021–1022, June 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018833631207.

Drechsler:1999:BWI

[2258] W. Drechsler and H. Tann. Broken Weyl invariance and the origin of mass. Foundations of Physics, 29(7):1023–1064, July 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1012851715278.

Khrennikov:1999:CQM

[2259] Andrei Khrennikov. Classical and quantum mechanics on information spaces with applications to cognitive, psychological, social, and anomalous phenomena. Foundations of Physics, 29(7):1065–1098, July 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1018885632116.

deSabbata:1999:HFG

[2260] Venzo de Sabbata and Luca Ronchetti. A Hamiltonian formulation of gravitational theory that allows one to consider curvature and torsion as conjugate variables. Foundations of Physics, 29(7):1099–1117, July 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1018889716186.

Mehra:1999:YJSe

[2261] Jagdish Mehra, Kimball A. Milton, and Peter Rembiesa. The young Julian Schwinger. V. Winding up at the Radiation Lab, going to Harvard, and marriage. Foundations of Physics, 29(7):1119–1162, July 1999. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1018841800257.

Anonymous:1999:BPLg

[2262] Anonymous. Foundations of Physics Letters — list of contents, volume 12, number 2, April 1999. Foundations of Physics, 29(7):1163, July 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1018893817095.

Bacciagaluppi:1999:DMI

[2263] Guido Bacciagaluppi and Michael Dickson. Dynamics for modal interpretations. Foundations of Physics, 29(8):1165–1201, August 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1018803613886.

Kobe:1999:RSL

[2264] Donald H. Kobe. A relativistic Schrödinger-like equation for a photon and its second quantization. Foundations of Physics, 29(8):1203–1231, August 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018855630724.

DeLorenci:1999:RQV

[2265] V. A. De Lorenci and N. F. Svaiter. A rotating quantum vacuum. Foundations of Physics, 29(8):1233–1264, August 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1018807714794.

Norton:1999:QMS

[2266] John D. Norton. A quantum mechanical supertask. Foundations of Physics, 29(8):1265–1302, August 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1018859831633.

Israelit:1999:MCG

[2267] Mark Israelit. Matter creation by geometry in an integrable Weyl–Dirac theory. Foundations of Physics, 29(8):1303–1322, August 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1018811915703.

Anonymous:1999:BPLh

[2268] Anonymous. Foundations of Physics Letters — list of contents: Volume 12, number 4, April 1999. Foundations of Physics, 29(8):1323, August 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018863932541.

Held:1999:MAM

[2269] Alan Held. In memoriam: André Mercier, 1913–1999. Foundations of Physics, 29(9):1325–1326, September 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018816024789.

Drechsler:1999:MGW

[2270] Wolfgang Drechsler. Mass generation by Weyl symmetry breaking. Foundations of Physics, 29(9):1327–1369, September 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1018805108859.

Kudaka:1999:MBT

[2271] Shoju Kudaka. Modified Busch-type measurements and the universal conservation laws. Foundations of Physics, 29(9):1371–1388, September 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018857125698.

Pulmannova:1999:EAR

[2272] Sylvia Pulmannova. Effect algebras with the Riesz decomposition property and AF C∗-algebras. Foundations of Physics, 29(9):1389–1401, September 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018809209768.

Maroney:1999:QST

[2273] O. Maroney and B. J. Hiley. Quantum state teleportation understood through the Bohm interpretation. Foundations of Physics, 29(9):1403– 1415, September 1999. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1018861226606.

Petrosky:1999:TTCa

[2274] T. Petrosky. Transport theory and collective modes. I. The case of moderately dense gases. Foundations of Physics, 29(9):1417–1456, September 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018813310677.

Leaf:1999:DKG

[2275] Boris Leaf. Dissipation in the Klein–Gordon field. Foundations of Physics, 29(9):1457–1478, September 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018865327515.

Hachemane:1999:FRN

[2276] M. Hachemane, A. Smida, and R. Djelid. Functional realization and nonlinear induced representation in the geometrodifferential conception of extended particles. Foundations of Physics, 29(9):1479–1495, September 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018817411585.

Anonymous:1999:BPLi

[2277] Anonymous. Foundations of Physics Letters — list of contents: Volume 12, number 3, June 1999. Foundations of Physics, 29(9):1497, September 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018869428424.

Pauri:1999:CRU

[2278] M. Pauri and M. Vallisneri. Classical roots of the Unruh and Hawking effects. Foundations of Physics, 29(10):1499–1520, October 1999. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1018821619763.

Frieden:1999:IUV

[2279] B. Roy Frieden. F -information, a unitless variant of Fisher information. Foundations of Physics, 29(10):1521–1541, October 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1018806603833.

Vargas:1999:TAW

[2280] Jose G. Vargas and Douglas G. Torr. The theory of acceleration within its context of differential invariants: The root of the problem with cosmological models? Foundations of Physics, 29(10):1543–1580, October 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1018858620671.

Petrosky:1999:TTCb

[2281] T. Petrosky. Transport theory and collective modes II: Long-time tail and Green–Kubo formalism. Foundations of Physics, 29(10):1581–1605, October 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018810704742.

Cattaneo:1999:ASA

[2282] Gianpiero Cattaneo and Stanley Gudder. Algebraic structures arising in axiomatic unsharp quantum physics. Foundations of Physics, 29(10): 1607–1637, October 1999. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1018862721580.

Gudder:1999:BRB

[2283] Stanley P. Gudder. Book review: Quantum Computing and Quantum Communications, edited by Colin P. Williams. Foundations of Physics, 29(10):1639–1642, October 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1018814805650.

Anonymous:1999:BPLj

[2284] Anonymous. Foundations of Physics Letters — list of contents: Volume 12, number 3, June 1999. Foundations of Physics, 29(10):1643, October 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018866822489.

vanderMerwe:1999:MHM

[2285] Alwyn van der Merwe. In memoriam: Henry Margenau, 1903–1999. Foundations of Physics, 29(11):1645–1646, November 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1018872428366.

Anandan:1999:TDL

[2286] J. Anandan. Are there dynamical laws? Foundations of Physics, 29(11): 1647–1672, November 1999. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1018869712436.

Cole:1999:CTC

[2287] Daniel C. Cole. Cross-term conservation relationships for electromagnetic energy, linear momentum, and angular momentum. Foundations of Physics, 29(11):1673–1693, November 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018821829274.

Gron:1999:FPR

[2288] Øyvind Grøn and Kjell Vøyenli. On the foundation of the principle of relativity. Foundations of Physics, 29(11):1695–1733, November 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1018873813345.

Cardone:1999:BLI

[2289] Fabio Cardone and Roberto Mignani. Broken Lorentz invariance and metric description of interactions in a deformed Minkowski space.

Foundations of Physics, 29(11):1735–1783, November 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1018825930183.

Weigert:1999:GTD

[2290] Stefan Weigert. Gauge transformations for a driven quantum particle in an infinite square Well. Foundations of Physics, 29(11):1785–1805, November 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018878014253.

Schulman:1999:BRB

[2291] L. S. Schulman. Book review: Decoherence and Quantum Measurements, by Mikio Namiki. Foundations of Physics, 29(11):1807–1810, November 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018830131092.

Mizushima:1999:BRB

[2292] Masataka Mizushima. Book review: Quantum Mechanics, Third Edition, by Eugene Merzbacher. Foundations of Physics, 29(11):1811–1812, November 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018882115162.

Anonymous:1999:BPLk

[2293] Anonymous. Foundations of Physics Letters — list of contents: Volume 12, number 5, October 1999. Foundations of Physics, 29(11):1813, November 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018834232000.

Anonymous:1999:ACPa

[2294] Anonymous. Announcement: Call for papers. Foundations of Physics, 29(11):1815–1817, November 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1018886216070.

Cole:1999:CBT

[2295] Daniel C. Cole. Connections between the thermodynamics of classical electrodynamic systems and quantum mechanical systems for quasielectrostatic operations. Foundations of Physics, 29(12):1819–1847, December 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018838300141.

Kim:1999:ERP

[2296] Yoon-Ho Kim and Yanhua Shih. Experimental realization of Popper’s experiment: Violation of the uncertainty principle? Foundations of Physics, 29(12):1849–1861, December 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018890316979.

Appleby:1999:GBT

[2297] D. M. Appleby. Generic Bohmian trajectories of an isolated particle. Foundations of Physics, 29(12):1863–1883, December 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1018842401049.

Appleby:1999:BTP

[2298] D. M. Appleby. Bohmian trajectories post-decoherence. Foundations of Physics, 29(12):1885–1916, December 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018894417888.

Gour:1999:WSP

[2299] Gilad Gour and L. Sriramkumar. Will small particles exhibit Brownian motion in the quantum vacuum? Foundations of Physics, 29(12): 1917–1949, December 1999. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1018846501958.

Mould:1999:QC

[2300] Richard A. Mould. Quantum consciousness. Foundations of Physics, 29(12):1951–1961, December 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1018898518796.

Ben-Aryeh:1999:ESD

[2301] Y. Ben-Aryeh, A. Mann, and B. C. Sanders. Empirical state determination of entangled two-level systems and its relation to information theory. Foundations of Physics, 29(12):1963–1975, December 1999. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1018850602867.

Pitts:1999:FPG

[2302] J. Brian Pitts and W. C. Schieve. On the form of parametrized gravitation in flat spacetime. Foundations of Physics, 29(12):1977–1985, December 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018802719705.

Mehra:1999:BRB

[2303] Jagdish Mehra. Book review: Classical Electrodynamics, by Julian Schwinger. Foundations of Physics, 29(12):1987–1991, December 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1018854703775.

Anonymous:1999:AIVb

[2304] Anonymous. Author index to volume 29. Foundations of Physics, 29(12): 1993–1994, December 1999. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1018806920614.

Anonymous:1999:ACPb

[2305] Anonymous. Announcement: Call for papers. Foundations of Physics, 29(12):1995–1997, December 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1018858904684.

Anonymous:1999:BPLl

[2306] Anonymous. Foundations of Physics Letters — list of contents: Volume 12, number 5, October 1999. Foundations of Physics, 29(12):1999, December 1999. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1018811021522.

Belousek:2000:SSC

[2307] Darrin W. Belousek. Statistics, symmetry, and the conventionality of indistinguishability in quantum mechanics. Foundations of Physics, 30 (1):1–34, January 2000. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1003658107978.

Bell:2000:CBD

[2308] Sarah B. M. Bell, John P. Cullerne, and Bernard M. Diaz. Classical behavior of the Dirac bispinor. Foundations of Physics, 30(1):35–57, January 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1003686924816.

Dobyns:2000:CIV

[2309] York Dobyns, Alfonso Rueda, and Bernard Haisch. The case for inertia as a vacuum effect: a reply to Woodward and Mahood. Foundations of Physics, 30(1):59–80, January 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1003639008887. See [2251].

Costantini:2000:PPR

[2310] D. Costantini and U. Garibaldi. A purely probabilistic representation for the dynamics of a gas of particles. Foundations of Physics, 30(1):81–99, January 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1003691025725.

SantAnna:2000:QST

[2311] Adonai S. Sant’Anna and Alexandre M. S. Santos. Quasi-set-theoretical foundations of statistical mechanics: a research program. Foundations of Physics, 30(1):101–120, January 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1003643109795.

Bernui:2000:RRP

[2312] Armando Bernui. The radiation reaction problem in a simple coupled model. Foundations of Physics, 30(1):121–138, January 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1003695126634.

Avinash:2000:NAA

[2313] K. Avinash and V. L. Rvachev. Non-Archimedean algebra: applications to cosmology and gravitation. Foundations of Physics, 30(1):139–152, January 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1003647210704.

Belousek:2000:SSD

[2314] Darrin W. Belousek. Statistics, symmetry, and (in)distinguishability in Bohmian mechanics. Foundations of Physics, 30(1):153–164, January 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1003699327542.

Woodward:2000:BRBa

[2315] James F. Woodward. Book reviews: Time: a Traveler’s Guide. By Clifford A. Pickover. Oxford University Press, New York, 1998, xviii + 285 pp., 815.95 (softcover, 1999). ISBN 0-19-513096-0. Surfing Through Hyperspace: Understanding Higher Universes in Six Easy Lessons. By Clifford A. Pickover. Oxford University Press, New York, 1999, xxiv + 239 pp., 825.00 (hardcover). ISBN 0-19-513006-5. Foundations of Physics, 30 (1):165–170, January 2000. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1003651411612.

Phipps:2000:BRB

[2316] Thomas E. Phipps, Jr. Book review: Relational Mechanics. By Andre K. T. Assis. Apeiron, Montreal, Canada, 1999, 285 pp., $825.00 (softcover). (Available from C. Roy Keys, Inc., Montreal, Canada.) ISBN 0-9683689-2-1. Foundations of Physics, 30(1):171–173, January 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1003603528451.

Zimba:2000:QMU

[2317] Jason Zimba. Quantum-mechanical uncertainty and the stability of incompatibility. Foundations of Physics, 30(2):179–203, February 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1003660715847.

deMuynck:2000:PMT

[2318] Willem M. de Muynck. Preparation and measurement: Two independent sources of uncertainty in quantum mechanics. Foundations of Physics, 30 (2):205–225, February 2000. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1003640932686.

Caticha:2000:IRE

[2319] Ariel Caticha. Insufficient reason and entropy in quantum theory. Foundations of Physics, 30(2):227–251, February 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1003692916756.

Zareski:2000:EPW

[2320] David Zareski. Embedding of particle waves in a Schwarzschild metric background. Foundations of Physics, 30(2):253–285, February 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1003645000826.

Smida:2000:EIP

[2321] A. Smida, M. Hachemane, R. Djelid, and A.-H. Hamici. On external and internal properties of extended elementary objects. Foundations of Physics, 30(2):287–299, February 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1003697017665.

Gudder:2000:BPQ

[2322] Stanley Gudder. Basic properties of quantum automata. Foundations of Physics, 30(2):301–319, February 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1003649201735.

Garuccio:2000:NSF

[2323] Augusto Garuccio, Ramón Risco-Delgado, and Franco Selleri. Nonlocality of some factorable quantum mechanical state vectors. Foundations of Physics, 30(2):321–330, February 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1003601318573.

Ungar:2000:RCV

[2324] Abraham A. Ungar. The relativistic composite-velocity reciprocity principle. Foundations of Physics, 30(2):331–342, February 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1003653302643.

Anonymous:2000:AKH

[2325] Anonymous. In appreciation of Kurt Haller on his seventieth birthday. Foundations of Physics, 30(3):347–348, March 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1017209520390.

Gusynin:2000:PGP

[2326] V. P. Gusynin, V. A. Miransky, and I. A. Shovkovy. Physical gauge in the problem of dynamical chiral symmetry breaking in QED in a magnetic field. Foundations of Physics, 30(3):349–357, March 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1003605404461.

Swanson:2000:GGV

[2327] Mark S. Swanson. Of ghosts, gauge volumes, and Gauss’s law. Foundations of Physics, 30(3):359–370, March 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1003613621299.

Sucher:2000:GIT

[2328] J. Sucher. Gauge-independence and the two-body problem in QED. Foundations of Physics, 30(3):371–381, March 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1003665605369.

Greenberg:2000:SMQ

[2329] O. W. Greenberg. Study of a model of quantum electrodynamics. Foundations of Physics, 30(3):383–391, March 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1003617722208.

Bender:2000:EPS

[2330] Carl M. Bender and H. F. Jones. Effective potential for PT - Symmetric quantum field theories. Foundations of Physics, 30(3):393– 411, March 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1003669706278.

Holstein:2000:EEA

[2331] Barry R. Holstein. Evaluating the effective action. Foundations of Physics, 30(3):413–437, March 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1003621823116.

Elias:2000:VCA

[2332] Victor Elias, Kevin B. Sprague, and Ying Xue. Vacuum condensates and the anomalous magnetic moment of a Dirac fermion. Foundations of Physics, 30(3):439–461, March 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1003673807187.

Dunne:2000:IAV

[2333] Gerald V. Dunne. Instantons and asymmetric vacua. Foundations of Physics, 30(3):463–474, March 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1003625924025.

Antonuccio:2000:CLC

[2334] F. Antonuccio, S. Pinsky, and S. Tsujimaru. A comment on the lightcone vacuum in 1+1 dimensional super–Yang–Mills theory. Foundations of Physics, 30(3):475–486, March 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1003677908095.

Creutz:2000:TML

[2335] Michael Creutz. Transfer matrices and lattice fermions at finite density. Foundations of Physics, 30(3):487–492, March 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1003630124933.

Ralston:2000:ECW

[2336] John P. Ralston. Embedding coordinates for the well-dressed quark. Foundations of Physics, 30(4):493–518, April 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1003678210280.

Kim:2000:NQP

[2337] Victor T. Kim, Grigorii B. Pivovarov, and James P. Vary. A new qualitative prediction of the parton model for high-energy hadron collisions. Foundations of Physics, 30(4):519–527, April 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1003612727118.

Fried:2000:NTA

[2338] H. M. Fried. New thoughts about an old eikonal problem. Foundations of Physics, 30(4):529–532, April 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1003664711188.

Frankfurt:2000:PPW

[2339] L. Frankfurt, G. A. Miller, and M. Strikman. Perturbative pion wave function in coherent pion–nucleon di–jet production. Foundations of Physics, 30(4):533–542, April 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1003616828027.

Gamberg:2000:DIS

[2340] Leonard Gamberg, George R. Kalbfleisch, and Kimball A. Milton. Direct and indirect searches for low-mass magnetic monopoles. Foundations of Physics, 30(4):543–565, April 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1003668812097.

Chodos:2000:CSB

[2341] Alan Chodos. Concatenation of scales below 1 eV. Foundations of Physics, 30(4):567–576, April 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1003620928935.

Islam:2000:EFT

[2342] M. M. Islam and H. Weigel. An effective field theory model to describe nuclear matter in heavy-ion collisions. Foundations of Physics, 30(4):577–597, April 2000. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1003672913006.

Gelb:2000:ALJ

[2343] James M. Gelb and S. P. Rosen. Another look at just-so solar neutrino oscillations. Foundations of Physics, 30(4):599–606, April 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1003625029844.

Anonymous:2000:BRBa

[2344] Anonymous. Book review: Ultimate Zero and One: Computing at the Quantum Frontier. By Colin P. Williams and Scott H. Clearwater. Copernicus, Springer-Verlag, New York, New York, 2000, i–xiv, 250 pp., $27.00 (hardcover). Foundations of Physics, 30(4):607–609, April 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1017260613914.

Anonymous:2000:BRBb

[2345] Anonymous. Book review: Consciousness: The Quantum Mind and the Meaning of Life. By Evan Harris Walker. Helix–Perseus Books, Cambridge, Massachusetts, 1999, 320 pp., $27.50 (hardcover). ISBN 0-7382-0234-7. Foundations of Physics, 30(4):611–614, April 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1017229230753.

Anonymous:2000:BPLa

[2346] Anonymous. Foundations of Physics Letters. Foundations of Physics, 30(4):615, April 2000. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1017281214823.

Anonymous:2000:ACP

[2347] Anonymous. Announcement: Call for papers. Foundations of Physics, 30(4):617–619, April 2000. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1017233331661.

Rohrlich:2000:EPR

[2348] Fritz Rohrlich. The equivalence principle revisited. Foundations of Physics, 30(5):621–630, May 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1003703825166.

Moreau:2000:DGI

[2349] William Moreau. Dynamically generated inertia. Foundations of Physics, 30(5):631–651, May 2000. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1003728709237.

Strauss:2000:RRR

[2350] Y. Strauss and L. P. Horwitz. Representation of the resonance of a relativistic quantum field theoretical Lee–Friedrichs model in Lax–Phillips scattering theory. Foundations of Physics, 30(5):653–694, May 2000. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1003780726075.

West:2000:ASP

[2351] Geoffrey B. West. Asymptotic series and precocious scaling. Foundations of Physics, 30(5):695–704, May 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1003732910145.

Liboff:2000:GPD

[2352] Richard L. Liboff. Generalized partial differential equation and Fermat’s last theorem. Foundations of Physics, 30(5):705–708, May 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1003784926984.

Mannheim:2000:ARG

[2353] Philip D. Mannheim. Attractive and repulsive gravity. Foundations of Physics, 30(5):709–746, May 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1003737011054.

Mbonye:2000:GPR

[2354] Manasse R. Mbonye and Ronald L. Mallett. Gravitational perturbations of a radiating spacetime. Foundations of Physics, 30(5):747–774, May 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1003789027892.

Carlson:2000:CEA

[2355] Carl. E. Carlson and Ian J. Swanson. Casimir energy in astrophysics: Gamma-ray bursts from QED vacuum transitions. Foundations of Physics, 30(5):775–783, May 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1003741111963.

Davidson:2000:EEB

[2356] Aharon Davidson and Uzi Paz. Extensible embeddings of black-hole geometries. Foundations of Physics, 30(5):785–794, May 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1003793128801.

Ng:2000:MFS

[2357] Y. Jack Ng and H. van Dam. Measuring the foaminess of space–time with gravity-wave interferometers. Foundations of Physics, 30(5):795– 805, May 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1003745212871.

Bar:2000:ZEE

[2358] D. Bar. The Zeno effect in the EPR Paradox, in the teleportation process, and in Wheeler’s delayed-choice experiment. Foundations of Physics, 30 (6):813–838, June 2000. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1003692507147.

Urbanowski:2000:DLO

[2359] K. Urbanowski and J. Piskorski. On departures from the Lee, Oehme and Yang approximation. Foundations of Physics, 30(6):839–866, June 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1003698323985.

Schelb:2000:CFM

[2360] Udo Schelb. Characterizability of free motion in special relativity. Foundations of Physics, 30(6):867–892, June 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1003650408056.

Boyer:2000:DAB

[2361] Timothy H. Boyer. Does the Aharonov–Bohm effect exist? Foundations of Physics, 30(6):893–905, June 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1003602524894.

Boyer:2000:CEA

[2362] Timothy H. Boyer. Classical electromagnetism and the Aharonov–Bohm phase shift. Foundations of Physics, 30(6):907–932, June 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1003654508964.

Strunin:2000:SAA

[2363] D. V. Strunin. Stable autosoliton of the action function as a particle-type structure. Foundations of Physics, 30(6):933–949, June 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1003606725803.

Teller:2000:IQM

[2364] Paul Teller and Michael Redhead. Is indistinguishability in quantum mechanics conventional? Foundations of Physics, 30(6):951–957, June 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1003658709873.

Woodward:2000:BRBb

[2365] James F. Woodward. Book review: Concepts of Mass in Contemporary Physics and Philosophy, by Max Jammer. Foundations of Physics, 30 (6):959–964, June 2000. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1003610826711.

Anonymous:2000:ASSa

[2366] Anonymous. Announcement: Summer School on Theoretical Physics. Foundations of Physics, 30(6):965–966, June 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1017295410781.

Anonymous:2000:AWOa

[2367] Anonymous. Announcement: The 9th Workshop on Open Systems and Information Dynamics. Foundations of Physics, 30(6):967–968, June 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1017214927620.

Anonymous:2000:FPLa

[2368] Anonymous. Foundations of Physics Letters: List of contents: Volume 13, number 3, June 2000. Foundations of Physics, 30(6):969, June 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1017201423019.

Zisis:2000:AHV

[2369] M. Zisis. Approximate hidden variables. Foundations of Physics, 30 (7):971–1000, July 2000. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1003662802913.

Svozil:2000:RR

[2370] K. Svozil. Relativizing relativity. Foundations of Physics, 30(7):1001– 1016, July 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1003600519752.

Wheeler:2000:SS

[2371] James T. Wheeler. String without strings. Foundations of Physics, 30 (7):1017–1091, July 2000. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1003652503822.

Singleton:2000:OFA

[2372] D. Singleton and V. Dzhunushaliev. Orbital and field angular momentum in the nucleon. Foundations of Physics, 30(7):1093–1105, July 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1003604620660.

Assis:2000:PES

[2373] A. K. T. Assis. On the propagation of electromagnetic signals in wires and coaxial cables according to Weber’s electrodynamics. Foundations of Physics, 30(7):1107–1121, July 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1003656604731.

Anastasovski:2000:ODG

[2374] P. K. Anastasovski, T. E. Bearden, C. Ciubotariu, W. T. Coffey, W. T. Coffey, L. B. Crowell, G. J. Evans, M. W. Evans, R. Flower, A. Labounsky, B. Lehnert, P. R. Molnár, S. Roy, and J. P. Vigier. Operator derivation of the gauge-invariant Proca and Lehnert equations; elimination of the Lorenz condition. Foundations of Physics, 30(7):1123– 1129, July 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1003608721569.

Anonymous:2000:BRBc

[2375] Anonymous. Book review: The Quantum Theory of Fields, Volume I II: Supersymmetry. Steven Weinberg. Cambridge University Press, New York, New York, 2000, vii–xxii + 1-419, $849.95 (hardcover). Foundations of Physics, 30(7):1131–1133, July 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1017244005639.

Anonymous:2000:BRBd

[2376] Anonymous. Book review: Quantum Chaos — An Introduction. Hans-Jürgen Stöckmann. Cambridge University Press, Cambridge, United Kingdom, 1999, xi + 368 pp., 885.00 (hardcover). ISBN 0-521-59284-4. Foundations of Physics, 30(7):1135–1137, July 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1017212822478.

Anonymous:2000:ASSb

[2377] Anonymous. Announcement: Summer School on Theoretical Physics, July 31–August 5, 2000, Escuela de Fisica, Zacatecas Autonomous University, Mexico. Foundations of Physics, 30(7):1139–1140, July 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1017264806548.

Anonymous:2000:AWOb

[2378] Anonymous. Announcement: The 9th Workshop on Open Systems and Information Dynamics, September 19–21, 2000, Torun, Poland. Foundations of Physics, 30(7):1141–1142, July 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1017216923386.

Anonymous:2000:BPLb

[2379] Anonymous. Foundations of Physics Letters: List of contents: Volume 13, number 3, June 2000. Foundations of Physics, 30(7):1143, July 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1017268907456.

Pearle:2000:WCC

[2380] Philip Pearle. Wavefunction collapse and conservation laws. Foundations of Physics, 30(8):1145–1160, August 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1003677103804.

Fanchi:2000:QPR

[2381] John R. Fanchi. Quantum potential in relativistic dynamics. Foundations of Physics, 30(8):1161–1189, August 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1003680320642.

Poirier:2000:ASC

[2382] Bill Poirier. Algebraically self-consistent quasiclassical approximation on phase space. Foundations of Physics, 30(8):1191–1226, August 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1003632404712.

Sheehan:2000:SSW

[2383] D. P. Sheehan, J. Glick, and J. D. Means. Steady-state work by an asymmetrically inelastic gravitator in a gas: a Second Law paradox. Foundations of Physics, 30(8):1227–1256, August 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1003684421550.

Ungar:2000:BAH

[2384] Abraham A. Ungar. The bifurcation approach to hyperbolic geometry. Foundations of Physics, 30(8):1257–1282, August 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1003636505621.

Boudet:2000:ERC

[2385] Roger Boudet and Bruno Blaive. Exact relativistic calculation with retardation of the matrix elements used in the photoeffect of hydrogenic atoms. Foundations of Physics, 30(8):1283–1300, August 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1003688522459.

Castro:2000:SUR

[2386] Carlos Castro. The string uncertainty relations follow from the new relativity principle. Foundations of Physics, 30(8):1301–1316, August 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1003640606529.

Anonymous:2000:BRBe

[2387] Anonymous. Book review: Black Hole Physics. Basic Concepts and New Developments. By Valeri P. Frolov and Igor D. Novikov. Fundamental Theories of Physics, Vol. 96. Foundations of Physics, 30(8):1317–1320, August 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1017228323368.

Anonymous:2000:BRBf

[2388] Anonymous. Book review: The Genius of Science: a Portrait Gallery of 20th Century Physicists. Abraham Pais. Oxford University Press, New York, 2000, 356 pp. Foundations of Physics, 30(8):1321–1323, August 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1017244707438.

Anonymous:2000:BRBg

[2389] Anonymous. Book review: Principles of Optics. By Max Born and Emil Wolf. Cambridge University Press, Cambridge, United Kingdom, 1999, 985 pp. Foundations of Physics, 30(8):1325–1326, August 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1017296724276.

Anonymous:2000:AWOc

[2390] Anonymous. Announcement: The 9th Workshop on Open Systems and Information Dynamics. Foundations of Physics, 30(8):1327–1328, August 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1017200925185.

Anonymous:2000:FPLb

[2391] Anonymous. Foundations of Physics Letters: List of contents: Volume 13, number 3, June 2000. Foundations of Physics, 30(8):1329, August 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1017248808347.

Anonymous:2000:EPI

[2392] Anonymous. Editors’ preface: Interdisciplinary studies of quantum structures. Foundations of Physics, 30(9):1331–1335, September 2000. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1026420315635.

Ghirardi:2000:LMN

[2393] GianCarlo Ghirardi. Local measurements of nonlocal observables and the relativistic reduction process. Foundations of Physics, 30(9):1337– 1385, September 2000. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1026497632474.

Aerts:2000:VBI

[2394] Diederik Aerts, Sven Aerts, Jan Broekaert, and Liane Gabora. The violation of Bell inequalities in the macroworld. Foundations of Physics, 30(9): 1387–1414, September 2000. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1026449716544.

Beltrametti:2000:RTS

[2395] E. G. Beltrametti and S. Bugajski. Remarks on two-slit probabilities. Foundations of Physics, 30(9):1415–1429, September 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1026401800614.

Bacciagaluppi:2000:DPM

[2396] Guido Bacciagaluppi. Delocalized properties in the modal interpretation of a continuous model of decoherence. Foundations of Physics, 30(9): 1431–1444, September 2000. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1026453817453.

Peruzzi:2000:CIB

[2397] Giulio Peruzzi and Alberto Rimini. Compoundation invariance and Bohmian mechanics. Foundations of Physics, 30(9):1445–1472, September 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1026405901523.

Aerts:2000:WDQ

[2398] Diederik Aerts, Ellie D’Hondt, and Liane Gabora. Why the disjunction in quantum logic is not classical. Foundations of Physics, 30(9): 1473–1480, September 2000. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1026457918361.

Broekaert:2000:MQL

[2399] Jan Broekaert and Bart D’Hooghe. A model with quantum logic, but non-quantum probability: The product test issue. Foundations of Physics, 30(9):1481–1501, September 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1026410002431.

Pykacz:2000:LOF

[2400] Jarosław Pykacz. Lukasiewicz operations in fuzzy set and many-valued representations of quantum logics. Foundations of Physics, 30(9):1503– 1524, September 2000. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1026462019270.

Laudisa:2000:TAH

[2401] Federico Laudisa. On time asymmetry and history in an Everett quantum world. Foundations of Physics, 30(9):1525–1538, September 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1026414103340.

Garola:2000:OVN

[2402] Claudio Garola. Objectivity versus nonobjectivity in quantum mechanics. Foundations of Physics, 30(9):1539–1565, September 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1026466120178.

Anonymous:2000:BPLc

[2403] Anonymous. Foundations of Physics Letters. Foundations of Physics, 30 (9):1567, September 2000. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1026418204249.

Suppes:2000:ISM

[2404] Patrick Suppes. Invariance, symmetry and meaning. Foundations of Physics, 30(10):1569–1585, October 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1026437914611.

Castellani:2000:LPP

[2405] Elena Castellani and Peter Mittelstaedt. Leibniz’s principle, physics, and the language of physics. Foundations of Physics, 30(10):1587–1604, October 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1026489931449.

Dorato:2000:SRS

[2406] Mauro Dorato. Substantivalism, relationism, and structural spacetime realism. Foundations of Physics, 30(10):1605–1628, October 2000. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1026442015519.

Boniolo:2000:PFM

[2407] Giovanni Boniolo and Fernando de Felice. On the philosophical foundations of measurements in general relativity. Foundations of Physics, 30(10):1629–1641, October 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1026494032358.

Ghins:2000:EVT

[2408] Michel Ghins. Empirical versus theoretical existence and truth. Foundations of Physics, 30(10):1643–1654, October 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1026446116428.

vanFraassen:2000:MGE

[2409] Bas C. van Fraassen. Michel Ghins on the empirical versus the theoretical. Foundations of Physics, 30(10):1655–1661, October 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1026498100498.

Gudder:2000:WFP

[2410] S. Gudder. What is fuzzy probability theory? Foundations of Physics, 30(10):1663–1678, October 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1026450217337.

Riecan:2000:JDO

[2411] Beloslav Riecan. On the joint distribution of observables. Foundations of Physics, 30(10):1679–1686, October 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1026402301407.

Foulis:2000:MHE

[2412] D. J. Foulis. MV and Heyting effect algebras. Foundations of Physics, 30(10):1687–1706, October 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1026454318245.

Isham:2000:SPR

[2413] C. J. Isham and J. Butterfield. Some possible roles for topos theory in quantum theory and quantum gravity. Foundations of Physics, 30(10): 1707–1735, October 2000. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1026406502316.

Chevalier:2000:PBG

[2414] Georges Chevalier, Anatolij Dvurecenskij, and Karl Svozil. Piron’s and Bell’s geometric lemmas and Gleason’s theorem. Foundations of Physics, 30(10):1737–1755, October 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1026458519154.

Cassinelli:2000:TLR

[2415] G. Cassinelli, E. De Vito, P. Lahti, and A. Levrero. A theorem of Ludwig revisited. Foundations of Physics, 30(10):1757–1763, October 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1026410603224.

Cattaneo:2000:PBD

[2416] G. Cattaneo, R. Giuntini, and S. Pulmannovà. Pre-BZ and degenerate BZ posets: applications to fuzzy sets and unsharp quantum theories. Foundations of Physics, 30(10):1765–1799, October 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1026462620062.

Cattaneo:2000:MPS

[2417] G. Cattaneo, J. Hamhalter, and P. Pták. On the de Morgan property of the standard Brouwer–Zadeh poset. Foundations of Physics, 30(10): 1801–1805, October 2000. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1026414704133.

Neggers:2000:FFA

[2418] J. Neggers, A. Dvurecenskij, and Hee Sik Kim. On d-fuzzy functions in d-algebras. Foundations of Physics, 30(10):1807–1816, October 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1026466720971.

Anonymous:2000:FPLc

[2419] Anonymous. Foundations of Physics Letters: List of contents: Volume 13, number 3, June 2000. Foundations of Physics, 30(10):1817, October 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1026418805041.

Svetlichny:2000:STO

[2420] George Svetlichny. The space–time origin of quantum mechanics: Covering law. Foundations of Physics, 30(11):1819–1847, November 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1003797204134.

Cole:2000:TBR

[2421] Daniel C. Cole. Thermodynamics of blackbody radiation via classical physics for arbitrarily shaped cavities with perfectly conducting walls. Foundations of Physics, 30(11):1849–1867, November 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1003706320972.

Kruger:2000:TDU

[2422] Thomas Krüger. Towards a deeper understanding of the Einstein– Podolsky–Rosen problem. Foundations of Physics, 30(11):1869–1890, November 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1003758305043.

Szabo:2000:FRE

[2423] László E. Szabó. On Fine’s resolution of the EPR–Bell Problem. Foundations of Physics, 30(11):1891–1909, November 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1003710421881.

Shaarawi:2000:CAF

[2424] Amr M. Shaarawi. Clifford algebra formulation of an electromagnetic charge–current wave theory. Foundations of Physics, 30(11):1911–1941, November 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1003762405951.

Jin:2000:TIN

[2425] Wei Min Jin. From time inversion to nonlinear QED. Foundations of Physics, 30(11):1943–1973, November 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1003714522790.

Schroder:2000:GRCa

[2426] Wilfried Schröder and Hans-Jürgen Treder. The general-relativistic and covariant form of the first Helmholtz vorticity theorem. Foundations of Physics, 30(11):1975–1982, November 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1003766506860.

Schroder:2000:GRCb

[2427] Wilfried Schröder and Hans-Jürgen Treder. On the general-relativistic and covariant form of integration of the Helmholtz theorem. Foundations of Physics, 30(11):1983–1986, November 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1003718623698.

Schroder:2000:BFH

[2428] Wilfried Schröder and Hans-Jürgen Treder. Boltzmann’s formulation of the Helmholtz vorticity theorem. Foundations of Physics, 30(11): 1987–1989, November 2000. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1003770607768.

vonBorzeszkowski:2000:BRB

[2429] H.-H. von Borzeszkowski. Book review: Quantum Measurements and Decoherence: Models and Phenomenology. By Michael B. Mensky. Kluwer Academic Publishers, Dordrecht, Boston, London, 2000, xvi + 226 pages, 8118 (hardcover). Foundations of Physics, 30(11):1991–1993, November 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1003722724607.

Anonymous:2000:FPLd

[2430] Anonymous. Foundations of Physics Letters: List of contents: Volume 13, number 4, August 2000. Foundations of Physics, 30(11):1995, November 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1017215408677.

Anonymous:2000:MCG

[2431] Anonymous. To Martin C. Gutzwiller on his seventy-fifth birthday. Foundations of Physics, 30(12):1997–1998, December 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1017270320625.

Fazekas:2000:SOO

[2432] P. Fazekas. Spin and orbital order in itinerant ferromagnets. Foundations of Physics, 30(12):1999–2009, December 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1003724522132.

Bunemann:2000:MBG

[2433] Jörg Bünemann, Florian Gebhard, and Werner Weber. Multi-band Gutzwiller wave functions for itinerant ferromagnetism. Foundations of Physics, 30(12):2011–2031, December 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1003733306203.

Baeriswyl:2000:VSM

[2434] D. Baeriswyl. Variational scheme for the Mott transition. Foundations of Physics, 30(12):2033–2048, December 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1003785323041.

Fulde:2000:QCT

[2435] Peter Fulde and Hermann Stoll. Quantum chemical treatment of strong electron correlations. Foundations of Physics, 30(12):2049–2060, December 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1003737407111.

Hasegawa:2000:SBM

[2436] Hideo Hasegawa. Slave-boson mean–field theory of spin- and orbital- ordered states in the degenerate Hubbard model. Foundations of Physics, 30(12):2061–2078, December 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1003789423950.

Kimura:2000:MDA

[2437] N. Kimura and Y. Kakehashi. Molecular-dynamics approach to the magnetic structure of competing magnetic alloys: Fe–Cr alloys. Foundations of Physics, 30(12):2079–2100, December 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1003741508020.

Metzner:2000:STD

[2438] Walter Metzner. Superconductivity in the two-dimensional Hubbard model. Foundations of Physics, 30(12):2101–2112, December 2000. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1003793524858.

Ruckenstein:2000:BCB

[2439] Andrei E. Ruckenstein. Bose condensation without broken symmetries. Foundations of Physics, 30(12):2113–2124, December 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1003745608929.

Wolfle:2000:QTP

[2440] Peter Wölfle. Quasiclassical theory of phase relaxation by gauge field fluctuations. Foundations of Physics, 30(12):2125–2133, December 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1003797625767.

Galperin:2000:GNA

[2441] Y. M. Galperin and K. A. Chao. Giant nonlinear absorption by an ensemble of metallic grains. Foundations of Physics, 30(12):2135–2150, December 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1003749709837.

Anonymous:2000:AIV

[2442] Anonymous. Author index to volume 30. Foundations of Physics, 30(12): 2151–2153, December 2000. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1017389126783.

Anonymous:2000:FPLe

[2443] Anonymous. Foundations of Physics Letters: List of contents: Volume 13, number 4, August 2000. Foundations of Physics, 30(12):2155, December 2000. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1017240126676.

Anonymous:2001:MCGa

[2444] Anonymous. To Martin C. Gutzwiller on his seventy-fifth birthday. Foundations of Physics, 31(1):1–5, January 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1017376202635.

Zaitsev:2001:QBO

[2445] Oleg Zaitsev, R. Narevich, and R. E. Prange. Quasiclassical Born– Oppenheimer approximations. Foundations of Physics, 31(1):7–26, January 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1004153019473.

Scalapino:2001:NRH

[2446] Douglas J. Scalapino and S. R. White. Numerical results for the Hubbard model: Implications for the high T_{c} pairing mechanism. Foundations of Physics, 31(1):27–39, January 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1004147703543.

Devreese:2001:PIS

[2447] J. T. Devreese, F. Brosens, and L. F. Lemmens. Path integrals and statistics of identical particles. Foundations of Physics, 31(1):41–55, January 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1004199720382.

Klauder:2001:CSP

[2448] John R. Klauder. Coherent state path integrals without resolutions of unity. Foundations of Physics, 31(1):57–67, January 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1004151804452.

Carati:2001:TDS

[2449] A. Carati and L. Galgani. Theory of dynamical systems and the relations between classical and quantum mechanics. Foundations of Physics, 31 (1):69–87, January 2001. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1004103921290.

Contopoulos:2001:DND

[2450] G. Contopoulos. The development of nonlinear dynamics in astronomy. Foundations of Physics, 31(1):89–114, January 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1004155905361.

Rahav:2001:SSR

[2451] Saar Rahav and Shmuel Fishman. Spectral statistics of the rectangular billiard with a flux line. Foundations of Physics, 31(1):115–146, January 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1004108022199.

Borondo:2001:SMV

[2452] F. Borondo, J. C. Losada, and R. M. Benito. Scars in molecular vibrations and spectra of LiCN. Foundations of Physics, 31(1):147–163, January 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1004160006269.

Papadopoulos:2001:PIT

[2453] G. J. Papadopoulos and P. Melas. Photon induced tunneling oscillations in a double quantum Well. Foundations of Physics, 31(1):165–175, January 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1004112123108.

Takahashi:2001:CDS

[2454] Kin’ya Takahashi and Kensuke S. Ikeda. Complex-domain semiclassical theory: application to time-dependent barrier tunneling problems. Foundations of Physics, 31(1):177–201, January 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1004164107178.

Anonymous:2001:BPLa

[2455] Anonymous. Foundations of Physics Letters. Foundations of Physics, 31(1):203, January 2001. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1017395624016.

Anonymous:2001:MCGb

[2456] Anonymous. To Martin C. Gutzwiller on his seventy-fifth birthday. Foundations of Physics, 31(2):205–208, February 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1017515617678.

Brack:2001:BCS

[2457] Matthias Brack. Bifurcation cascades and self-similarity of periodic orbits with analytical scaling constants in Hénon–Heiles type potentials. Foundations of Physics, 31(2):209–232, February 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1017582218587.

Berggren:2001:CQD

[2458] K.-F. Berggren and T. Ouchterlony. Chaos in a quantum dot with spinorbit coupling. Foundations of Physics, 31(2):233–242, February 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1017586402657.

Reichl:2001:SSL

[2459] L. E. Reichl and G. Akguc. Scattering from spatially localized chaotic and disordered systems. Foundations of Physics, 31(2):243–267, February 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1017538519495.

Blumel:2001:RSB

[2460] R. Blümel, P. M. Koch, and L. Sirko. Ray-splitting billiards. Foundations of Physics, 31(2):269–281, February 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1017590503566.

Berglund:2001:ADH

[2461] Nils Berglund and Turgay Uzer. The averaged dynamics of the hydrogen atom in crossed electric and magnetic fields as a perturbed Kepler problem. Foundations of Physics, 31(2):283–326, February 2001. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1017542620404.

Richter:2001:TTF

[2462] A. Richter. Test of trace formulas for spectra of superconducting microwave billiards. Foundations of Physics, 31(2):327–354, February 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1017594604474.

Monteiro:2001:QWT

[2463] T. S. Monteiro and D. S. Saraga. Quantum Wells in tilted fields: Semiclassical amplitudes and phase coherence times. Foundations of Physics, 31(2):355–370, February 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1017546721313.

Ramey:2001:HCH

[2464] R. Dwayne Ramey and N. L. Balazs. Hamiltonian chaos in homogeneous ADM cosmologies? Foundations of Physics, 31(2):371–398, February 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1017598705383.

Frustaglia:2001:BQT

[2465] Diego Frustaglia and Klaus Richter. Ballistic quantum transport: Effect of geometrical phases. Foundations of Physics, 31(2):399–421, February 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1017550822221.

Bolte:2001:SEV

[2466] Jens Bolte. Semiclassical expectation values for relativistic particles with spin 1/2. Foundations of Physics, 31(2):423–444, February 2001. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1017502906292.

Anonymous:2001:BPLb

[2467] Anonymous. Foundations of Physics Letters. Foundations of Physics, 31(2):445, February 2001. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1017554923130.

Main:2001:POQ

[2468] Jörg Main and Günter Wunner. Periodic orbit quantization: How to make semiclassical trace formulae convergent. Foundations of Physics, 31(3):447–474, March 2001. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1017587012035.

Schafer:2001:SWR

[2469] R. Schäfer, U. Kuhl, M. Barth, and H.-J. Stöckmann. Spectra and wavefunctions in a ray–splitting Sinai microwave billiard and their semiclassical interpretation. Foundations of Physics, 31(3):475–487, March 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1017517628874.

Bohigas:2001:SSC

[2470] O. Bohigas, P. Leboeuf, and M. J. Sánchez. Spectral spacing correlations for chaotic and disordered systems. Foundations of Physics, 31 (3):489–517, March 2001. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1017569612944.

Hejhal:2001:QCM

[2471] Dennis A. Hejhal and Andreas Strömbergsson. On quantum chaos and Maass waveforms of CM-type. Foundations of Physics, 31(3):519–533, March 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1017521729782.

Pattard:2001:TFU

[2472] Thomas Pattard and Jan M. Rost. Threshold fragmentation under dipole forces. Foundations of Physics, 31(3):535–541, March 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1017573713853.

Eckhardt:2001:SQA

[2473] Bruno Eckhardt and Uzy Smilansky. Stroboscopic quantization of autonomous systems. Foundations of Physics, 31(3):543–556, March 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1017525830691.

Greenberger:2001:BRBa

[2474] Daniel Greenberger. Book review: New Developments on Fundamental Problems in Quantum Physics. By Miguel Ferrero and Alwyn van der Merwe, eds. Kluwer Academic Publishers, Dordrecht, The Netherlands, 1997. Foundations of Physics, 31(3):557–559, March 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1015742314761.

Taylor:2001:BRB

[2475] John R. Taylor. Book review: Introduction to Quantum Mechanics. By David J. Griffiths. Prentice Hall, New York, New York, 1995. Foundations of Physics, 31(3):561–563, March 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1015729931600.

Anonymous:2001:BPLc

[2476] Anonymous. Foundations of Physics Letters. Foundations of Physics, 31(3):565, March 2001. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1017577815670.

Anonymous:2001:SII

[2477] Anonymous. This special issue of invited papers is dedicated to Martin C. Gutzwiller. Foundations of Physics, 31(4):563–565, April 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1017556707359.

Anonymous:2001:MGH

[2478] Anonymous. To Martin Gutzwiller on his seventy-fifth birthday. Foundations of Physics, 31(4):567, April 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1017508824198.

Aurich:2001:OSR

[2479] R. Aurich and F. Steiner. Orbit sum rules for the quantum wave functions of the strongly chaotic Hadamard billiard in arbitrary dimensions. Foundations of Physics, 31(4):569–592, April 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1017560808268.

Kleppner:2001:BQM

[2480] Daniel Kleppner and John B. Delos. Beyond quantum mechanics: Insights from the work of Martin Gutzwiller. Foundations of Physics, 31 (4):593–612, April 2001. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1017512925106.

Braun:2001:LDU

[2481] Peter Braun, Sven Gnutzmann, Fritz Haake, Marek Kuś, and Karol Życzkowski. Level dynamics and universality of spectral fluctuations. Foundations of Physics, 31(4):613–622, April 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1017564909177.

Choquard:2001:HHJ

[2482] Philippe Choquard. The homogeneous Hamilton–Jacobi and Bernoulli equations revisited. Foundations of Physics, 31(4):623–640, April 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1017517026015.

Palla:2001:NCS

[2483] Gergely Palla, Gábor Vattay, André Voros, Niels Søndergaard, and Carl Philip Dettmann. Noise corrections to stochastic trace formulas. Foundations of Physics, 31(4):641–657, April 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1017569010085.

Berry:2001:KZQ

[2484] Michael Berry. Knotted zeros in the quantum states of hydrogen. Foundations of Physics, 31(4):659–667, April 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1017521126923.

Connors:2001:QEL

[2485] R. D. Connors and J. P. Keating. Quantum energy levels and classical periodic orbits: Discreteness and statistical duality. Foundations of Physics, 31(4):669–692, April 2001. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1017573110994.

Briggs:2001:DTD

[2486] John S. Briggs and Jan M. Rost. On the derivation of the timedependent equation of Schrödinger. Foundations of Physics, 31(4):693– 712, April 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1017525227832.

Schulman:2001:CGE

[2487] L. S. Schulman and Bernard Gaveau. Coarse grains: The emergence of space and order. Foundations of Physics, 31(4):713–731, April 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1017577211902.

Anonymous:2001:BPLd

[2488] Anonymous. Foundations of Physics Letters. Foundations of Physics, 31(4):733, April 2001. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1017529428741.

Bub:2001:QBC

[2489] Jeffrey Bub. The quantum bit commitment theorem. Foundations of Physics, 31(5):735–756, May 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1017597528026.

Ulfbeck:2001:GFW

[2490] Ole Ulfbeck and Aage Bohr. Genuine fortuitousness. Where did that click come from? Foundations of Physics, 31(5):757–774, May 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1017596312096.

Capria:2001:CSS

[2491] Marco Mamone Capria. On the conventionality of simultaneity in special relativity. Foundations of Physics, 31(5):775–818, May 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1017548428934.

Woodward:2001:GIQ

[2492] James F. Woodward. Gravity, inertia, and quantum vacuum zero point fields. Foundations of Physics, 31(5):819–835, May 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1017500513005.

deBeauregard:2001:RQM

[2493] O. Costa de Beauregard. Relativistic quantum mechanics as a telegraph. Foundations of Physics, 31(5):837–848, May 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1017552529843.

Horwitz:2001:BRB

[2494] Lawrence P. Horwitz. Book review: Classical Relativistic Many-Body Dynamics. By M. A. Trump and W. C. Schieve. Kluwer Academic Publishers, Fundamental Theories of Physics, Dordrecht, The Netherlands, 1999, 365 pp., $186.00 (hardcover). ISBN 0-7923-5737-X. Foundations of Physics, 31(5):849–854, May 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1017504613913.

Greenberger:2001:BRBb

[2495] Daniel Greenberger. Book review: Conceptual Foundations of Quantum Physics: an Overview from Modern Perspectives. By Dipankar Home. Plenum Publishing Corporation, New York, New York, 1997, xvii + 386 pp., $167.00 (hardcover). ISBN 0-306-45660-5. Foundations of Physics, 31(5):855–857, May 2001. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1017556630751.

Stachel:2001:BRB

[2496] John J. Stachel. Book review: Quantum Reprogramming. Ensembles and Single Systems: a Two-Tier Approach to Quantum Mechanics. By Evert Jan Post. Kluwer Academic Publishers, Dordrecht/Boston/London, 1995, xiv + 320 pp., $137.00 (hardcover). ISBN 0-7923-3565-1. Foundations of Physics, 31(5):859–861, May 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1017508814822.

Gudder:2001:BRBa

[2497] Stanley P. Gudder. Book review: New Trends in Quantum Structures. Anatolij Dvurecenskij and Sylvia Pulmannová. Kluwer Academic Publishers, Dordrecht, 2000, i-xvi, 1-541, $185 (hardcover). Foundations of Physics, 31(5):863–865, May 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1017560831660.

Sklar:2001:BRB

[2498] Lawrence Sklar. Book review: The Philosophy of Physics. By Roberto Torretti. Cambridge University Press, Cambridge, United Kingdom, 1999, xvi + 512 pp., $23.95 (softcover). Foundations of Physics, 31 (5):867–868, May 2001. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1017512915730.

Greenberg:2001:BRB

[2499] O. W. Greenberg. Book review: A Tale of Two Continents. By Abraham Pais. Princeton University Press, Princeton, New Jersey, 1997, vii– xvi + 1–511, $35.00 (hardcover). Foundations of Physics, 31(5):869– 870, May 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1017564932569.

Anonymous:2001:BPLe

[2500] Anonymous. Foundations of Physics Letters. Foundations of Physics, 31(5):871, May 2001. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1017517016639.

Anonymous:2001:P

[2501] Anonymous. Preface. Foundations of Physics, 31(6):873–874, June 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1017586916359.

Gyulassy:2001:CNE

[2502] Miklos Gyulassy and Dénes Molnár. Covariant non-equilibrium transport theory solutions for RHIC. Foundations of Physics, 31(6):875– 894, June 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1017555800429.

Bekenstein:2001:EKT

[2503] Jacob D. Bekenstein and Asaf Oron. Extended Kelvin theorem in relativistic magnetohydrodynamics. Foundations of Physics, 31(6):895– 907, June 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1017507917267.

Sklarz:2001:RMC

[2504] S. Sklarz and L. P. Horwitz. Relativistic mechanics of continuous media. Foundations of Physics, 31(6):909–934, June 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1017559901338.

Harpaz:2001:REC

[2505] Amos Harpaz and Noam Soker. Radiation from an electric charge. Foundations of Physics, 31(6):935–949, June 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1017512018176.

Oron:2001:RRC

[2506] O. Oron and L. P. Horwitz. Radiation reaction of the classical offshell relativistic charged particle. Foundations of Physics, 31(6):951– 966, June 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1017564002246.

Land:2001:CSE

[2507] M. C. Land and L. P. Horwitz. The covariant Stark effect. Foundations of Physics, 31(6):967–991, June 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1017516119084.

Segev:2001:LFA

[2508] Bilha Segev and J. C. Wells. A light-fronts approach to a two-center time-dependent Dirac equation. Foundations of Physics, 31(6):993–1015, June 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1017568103155.

Anonymous:2001:BPLf

[2509] Anonymous. Foundations of Physics Letters. Foundations of Physics, 31(6):1017, June 2001. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1017520219993.

Guendelman:2001:VES

[2510] E. I. Guendelman. The volume element of space–time and scale invariance. Foundations of Physics, 31(7):1019–1037, July 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1017572522313.

Balakin:2001:RDV

[2511] A. Balakin and V. Kurbanova. Relativistic dynamics of vector bosons in the field of gravitational radiation. Foundations of Physics, 31(7): 1039–1049, July 2001. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1017522206384.

Rezzolla:2001:SEF

[2512] L. Rezzolla, B. J. Ahmedov, and J. C. Miller. Stationary electromagnetic fields of a slowly rotating magnetized neutron star in general relativity. Foundations of Physics, 31(7):1051–1065, July 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1017574223222.

Cooperstock:2001:EAM

[2513] F. I. Cooperstock. Energy and angular momentum of systems in general relativity. Foundations of Physics, 31(7):1067–1082, July 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1017526307292.

Pitts:2001:FSG

[2514] J. B. Pitts and W. C. Schieve. Flat spacetime gravitation with a preferred foliation. Foundations of Physics, 31(7):1083–1104, July 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1017578424131.

Salisbury:2001:QGI

[2515] D. C. Salisbury. Quantum general invariance and loop gravity. Foundations of Physics, 31(7):1105–1118, July 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1017530508201.

Recami:2001:SMB

[2516] Erasmo Recami. Superluminal motions? A bird’s-eye view of the experimental situation. Foundations of Physics, 31(7):1119–1135, July 2001.

CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1017582525039.

Anonymous:2001:BPLg

[2517] Anonymous. Foundations of Physics Letters. Foundations of Physics, 31(7):1137, July 2001. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1017534609109.

Ivezic:2001:TTR

[2518] Tomislav Ivezić. “True transformations relativity” and electrodynamics. Foundations of Physics, 31(8):1139–1183, August 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1017547720033.

Pavsic:2001:CAB

[2519] Matej Pavsic. Clifford-algebra based polydimensional relativity and relativistic dynamics. Foundations of Physics, 31(8):1185–1209, August 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1017599804103.

Gersten:2001:MEO

[2520] Alexander Gersten. Maxwell equations — the one-photon quantum equation. Foundations of Physics, 31(8):1211–1231, August 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1017551920941.

Lieu:2001:MRB

[2521] Richard Lieu. Microscopic relativity: The basic theory. Foundations of Physics, 31(8):1233–1250, August 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1017504005011.

Piazzese:2001:PMP

[2522] F. I. Piazzese. Pythagorean metric in Particle’s relativistic dynamics. Foundations of Physics, 31(8):1251–1263, August 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1017556021850.

Anonymous:2001:BPLh

[2523] Anonymous. Foundations of Physics Letters. Foundations of Physics, 31(8):1265, August 2001. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1017508105920.

Fanchi:2001:NRD

[2524] John R. Fanchi. Nonlocality in relativistic dynamics. Foundations of Physics, 31(9):1267–1285, September 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1012206526801.

Huang:2001:SLR

[2525] Y.-S. Huang. Schrödinger-like relativistic wave equation of motion for the Lorentz-scalar potential. Foundations of Physics, 31(9):1287–1298, September 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1012270110871.

Gill:2001:CEP

[2526] Tepper L. Gill, W. W. Zachary, and J. Lindesay. The classical electron problem. Foundations of Physics, 31(9):1299–1355, September 2001. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1012222227710.

Montanus:2001:PTF

[2527] J. M. C. Montanus. Proper-time formulation of relativistic dynamics. Foundations of Physics, 31(9):1357–1400, September 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1012274211780.

Anonymous:2001:BPLi

[2528] Anonymous. Foundations of Physics Letters. Foundations of Physics, 31 (9):1401, September 2001. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1017347828618.

Spekkens:2001:NOC

[2529] R. W. Spekkens and J. E. Sipe. Non-orthogonal core projectors for modal interpretations of quantum mechanics. Foundations of Physics, 31(10): 1403–1430, October 2001. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1012630512689.

Spekkens:2001:MIQ

[2530] R. W. Spekkens and J. E. Sipe. A modal interpretation of quantum mechanics based on a principle of entropy minimization. Foundations of Physics, 31(10):1431–1464, October 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1012630429527.

Stapp:2001:QTR

[2531] Henry P. Stapp. Quantum theory and the role of mind in nature. Foundations of Physics, 31(10):1465–1499, October 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1012682413597.

Assis:2001:SCC

[2532] A. K. T. Assis, J. A. Hernandes, and J. E. Lamesa. Surface charges in conductor plates carrying constant currents. Foundations of Physics, 31(10):1501–1511, October 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1012634530435.

Anonymous:2001:BPLj

[2533] Anonymous. Foundations of Physics Letters. Foundations of Physics, 31(10):1513, October 2001. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1017378314506.

Foulis:2001:OCE

[2534] David J. Foulis and Stanley P. Gudder. Observables, calibration, and effect algebras. Foundations of Physics, 31(11):1515–1544, November 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1012686531344.

Anastopoulos:2001:QTH

[2535] C. Anastopoulos. Quantum theory without Hilbert spaces. Foundations of Physics, 31(11):1545–1580, November 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1012690715414.

Poirier:2001:PSO

[2536] Bill Poirier. Phase space optimization of quantum representations: Non-Cartesian coordinate spaces. Foundations of Physics, 31(11):1581–1610, November 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1012642832253.

Chen:2001:GSC

[2537] Jingling Chen and Abraham A. Ungar. From the group SL(2, C) to gyrogroups and gyrovector spaces and hyperbolic geometry. Foundations of Physics, 31(11):1611–1639, November 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1012694816323.

Antoniou:2001:NLE

[2538] I. Antoniou, E. Karpov, and G. Pronko. Non-locality in electrodynamics. Foundations of Physics, 31(11):1641–1655, November 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1012646900393.

Lamb:2001:BRB

[2539] Willis E. Lamb, Jr. Book review: Climbing the Mountain: The Scientific Biography of Julian Schwinger. By Jagdish Mehra and Kimball A. Milton. Oxford University Press, New York, New York, 2000, xii + 677 pp., $49.95 (hardcover). ISBN 0-19-850658-9. Foundations of Physics, 31(11): 1657–1660, November 2001. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1012698917232.

Post:2001:BRB

[2540] E. J. Post. Book review: Quantum Dialogue: The Making of a Revolution. By Mara Beller. University of Chicago Press, Chicago, Illinois, 1999, 365 pp., $35.00/$20.00 (hardcover/softcover). Foundations of Physics, 31(11):1661–1664, November 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1012651001302.

Gudder:2001:BRBb

[2541] Stanley P. Gudder. Book review: Quantum Computation and Quantum Information. By Michael A. Nielsen and Isaac L. Chuang. Cambridge University Press, Cambridge, United Kingdom, 2000, i-xxv + 676 pp., $42.00 (hardcover). Foundations of Physics, 31(11):1665–1667, November 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1012603118140.

Anonymous:2001:BPLk

[2542] Anonymous. Foundations of Physics Letters. Foundations of Physics, 31 (11):1669, November 2001. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1017338602210.

Lutz:2001:CMB

[2543] M. F. M. Lutz and E. E. Kolomeitsev. Covariant meson–baryon scattering with chiral and large Nc constraints. Foundations of Physics, 31(12): 1671–1702, December 2001. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1012655116247.

delaPena:2001:QTL

[2544] L. de la Peña and A. M. Cetto. Quantum theory and linear stochastic electrodynamics. Foundations of Physics, 31(12):1703–1731, December 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1012670800317.

Friedman:2001:NAS

[2545] Yaakov Friedman and Bernard Russo. A new approach to spinors and some representations of the Lorentz group on them. Foundations of Physics, 31(12):1733–1766, December 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1012622917155.

Rodrigues:2001:RFS

[2546] W. A. Rodrigues, Jr. and M. Sharif. Rotating frames in SRT: The Sagnac effect and related issues. Foundations of Physics, 31(12):1767–1783, December 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1012675001225.

Rodrigues:2001:EPP

[2547] W. A. Rodrigues, Jr. and M. Sharif. Equivalence principle and the principle of local Lorentz invariance. Foundations of Physics, 31(12): 1785–1806, December 2001. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1012627118064. See corrigenda [2591].

Anonymous:2001:AIV

[2548] Anonymous. Author index to volume 31. Foundations of Physics, 31(12): 1807–1808, December 2001. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1017453211350.

Anonymous:2001:BPLl

[2549] Anonymous. Foundations of Physics Letters. Foundations of Physics, 31 (12):1809, December 2001. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1017387902134.

Boyer:2002:CEI

[2550] Timothy H. Boyer. Classical electromagnetic interaction of a point charge and a magnetic moment: Considerations related to the Aharonov–Bohm phase shift. Foundations of Physics, 32(1):1–39, January 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1013803229638.

Boyer:2002:SEM

[2551] Timothy H. Boyer. Semiclassical explanation of the Matteucci–Pozzi and Aharonov–Bohm phase shifts. Foundations of Physics, 32(1):41–49, January 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1013896713708.

Batchelor:2002:SMV

[2552] David Batchelor. Semiclassical models for virtual antiparticle pairs, the unit of charge e, and the QCD coupling α_{s}. Foundations of Physics, 32(1):51–76, January 2002. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1013848830547. See erratum [2568].

Bernal:2002:FUL

[2553] A. N. Bernal, M. P. López, and M. Sánchez. Fundamental units of length and time. Foundations of Physics, 32(1):77–108, January 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1013800914617.

Grandpeix:2002:PDZa

[2554] Jean-Yves Grandpeix and François Lurçat. Particle description of zeroenergy vacuum I: Virtual particles. Foundations of Physics, 32(1): 109–131, January 2002. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1013852931455.

Grandpeix:2002:PDZb

[2555] Jean-Yves Grandpeix and François Lurçat. Particle description of zeroenergy vacuum II: Basic vacuum systems. Foundations of Physics, 32 (1):133–158, January 2002. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1013805015526.

Bailey:2002:ULC

[2556] C. D. Bailey. The unifying laws of classical mechanics. Foundations of Physics, 32(1):159–176, January 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1013857032364.

Schroder:2002:PNC

[2557] W. Schröder and H.-J. Treder. Post-Newtonian corrections in the dynamics in the Earth–Moon system and their importance for the relativistic theories of gravitation: a historical case study. Foundations of Physics, 32 (1):177–186, January 2002. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1013809116434.

Kragh:2002:BRB

[2558] Helge Kragh. Book review: The Historical Development of Quantum Theory, Volume 6: The Completion of Quantum Mechanics 1926–1941. By Jagdish Mehra and Helmut Rechenberg. Springer, New York, 2000, xxxvi + 1612 pp., $149.00 (hardcover). Foundations of Physics, 32(1): 187–189, January 2002. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1013861100505.

Anonymous:2002:BPLa

[2559] Anonymous. Foundations of Physics Letters. Foundations of Physics, 32(1):191, January 2002. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1017482617343.

Pulmannova:2002:HVB

[2560] Sylvia Pulmannová. Hidden variables and Bell inequalities on quantum logics. Foundations of Physics, 32(2):193–216, February 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1014424425657.

Mohrhoff:2002:WAQ

[2561] Ulrich Mohrhoff. The world according to quantum mechanics (or the 18 errors of Henry P. Stapp). Foundations of Physics, 32(2):217–254, February 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1014409209728.

Stapp:2002:FW

[2562] Henry P. Stapp. The 18-fold way. Foundations of Physics, 32(2):255–266, February 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1014461226566.

Falsaperla:2002:QLE

[2563] P. Falsaperla, G. Fonte, and G. Salesi. Quantum Lyapunov exponents. Foundations of Physics, 32(2):267–294, February 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1014413310636.

Israelit:2002:PMC

[2564] Mark Israelit. Primary matter creation in a Weyl–Dirac cosmological model. Foundations of Physics, 32(2):295–321, February 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1014465327475.

Hut:2002:BRB

[2565] Piet Hut. Book review: Structure and Interpretation of Classical Mechanics. By Gerald J. Sussman and Jack Wisdom, with Meinhard E. Mayer. The MIT Press, Cambridge, Massachusetts, 2001, xvii + 534 pp., $60.00 (hardcover). ISBN 0-262-19455-4. Foundations of Physics, 32 (2):323–326, February 2002. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1014417411545.

Walter:2002:BRB

[2566] Scott Walter. Book review: Beyond Einstein’s Velocity Addition Law. By Abraham A. Ungar. Fundamental Theories of Physics 117. Kluwer Academic, Dordrecht, The Netherlands, 2001, xlii + 413 pp., $138.00 (hardcover). Foundations of Physics, 32(2):327–330, February 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1014469428383.

Anonymous:2002:BPLb

[2567] Anonymous. Foundations of Physics Letters. Foundations of Physics, 32(2):331, February 2002. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1017457112453.

Batchelor:2002:ESM

[2568] David Batchelor. Erratum: “Semiclassical Models for Virtual Antiparticle Pairs, the Unit of Charge e, and the QCD Couplings α_{s}”. Foundations of Physics, 32(2):333, February 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1014421501766. See [2552].

Redei:2002:LPC

[2569] Miklós Rédei and Stephen J. Summers. Local primitive causality and the common cause principle in quantum field theory. Foundations of Physics, 32(3):335–355, March 2002. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1014869211488.

Megill:2002:DOO

[2570] Normal D. Megill and Mladen Pavicić. Deduction, ordering, and operations in quantum logic. Foundations of Physics, 32(3):357–378, March 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1014857228326.

Atmanspacher:2002:WQT

[2571] Harald Atmanspacher, Hartmann Römer, and Harald Walach. Weak quantum theory: Complementarity and entanglement in physics and beyond. Foundations of Physics, 32(3):379–406, March 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1014809312397.

Parrott:2002:RUA

[2572] Stephen Parrott. Radiation from a uniformly accelerated charge and the equivalence principle. Foundations of Physics, 32(3):407–440, March 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1014861329235.

Sheehan:2002:PSP

[2573] D. P. Sheehan, J. Glick, T. Duncan, J. A. Langton, M. J. Gagliardi, and R. Tobe. Phase space portraits of an unresolved gravitational Maxwell demon. Foundations of Physics, 32(3):441–462, March 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1014813413305.

Horton:2002:BPW

[2574] George Horton, Chris Dewdney, and Ulrike Ne’eman. de Broglie’s pilotwave theory for the Klein–Gordon equation and its space–time pathologies. Foundations of Physics, 32(3):463–476, March 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1014865430144.

Anonymous:2002:BPLc

[2575] Anonymous. Foundations of Physics Letters. Foundations of Physics, 32(3):477, March 2002. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1017404914214.

Svozil:2002:CRT

[2576] Karl Svozil. Conventions in relativity theory and quantum mechanics. Foundations of Physics, 32(4):479–502, April 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1015017831247.

Melloy:2002:GRP

[2577] G. F. Melloy. The generalized representation of particle localization in quantum mechanics. Foundations of Physics, 32(4):503–530, April 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1015080215317.

Chen:2002:BG

[2578] Jing-Ling Chen and Abraham A. Ungar. The Bloch gyrovector. Foundations of Physics, 32(4):531–565, April 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1015032332156.

Grove:2002:CPC

[2579] Peter G. Grove. Can the past be changed? Foundations of Physics, 32(4):567–587, April 2002. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1015084316226.

Genovese:2002:CTL

[2580] Marco Genovese, Carlo Novero, and Enrico Predazzi. On the conclusive tests of local realism and pseudoscalar mesons. Foundations of Physics, 32(4):589–605, April 2002. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1015036400296.

Mir-Kasimov:2002:RBR

[2581] R. M. Mir-Kasimov. Relation between relativistic and non-relativistic quantum mechanics as integral transformation. Foundations of Physics, 32(4):607–626, April 2002. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1015088417135.

Antoniou:2002:CFT

[2582] Ioannis E. Antoniou. Carathéodory and the foundations of thermodynamics and statistical physics. Foundations of Physics, 32(4):627–641, April 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1015040501205.

Anonymous:2002:BPLd

[2583] Anonymous. Foundations of Physics Letters. Foundations of Physics, 32(4):643, April 2002. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1017402118043.

Bene:2002:PVM

[2584] Gyula Bene and Dennis Dieks. A perspectival version of the modal interpretation of quantum mechanics and the origin of macroscopic behavior. Foundations of Physics, 32(5):645–671, May 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1016014008418.

Stefanovich:2002:MST

[2585] Eugene V. Stefanovich. Is Minkowski space–time compatible with quantum mechanics? Foundations of Physics, 32(5):673–703, May 2002. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1016052825257.

Rupp:2002:EDR

[2586] S. Rupp, S. Hunzinger, and M. Sorg. Exchange degeneracy of relativistic two-particle quantum states. Foundations of Physics, 32(5):705– 750, May 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1016004909327.

Frenkel:2002:TEK

[2587] Andor Frenkel. A tentative expression of the Károlyházy uncertainty of the space–time structure through vacuum spreads in quantum gravity.

Foundations of Physics, 32(5):751–771, May 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1016057026165.

Faraoni:2002:EPS

[2588] Valerio Faraoni and Donovan M. Faraoni. Elimination of the potential from the Schrödinger and Klein–Gordon equations by means of conformal transformations. Foundations of Physics, 32(5):773–788, May 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1016522910236.

Sancho:2002:PER

[2589] Pedro Sancho. Popper’s experiment revisited. Foundations of Physics, 32(5):789–805, May 2002. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1016009127074.

Anselmino:2002:BRB

[2590] Mauro Anselmino. Book review: Spin in Particle Physics. By Elliot Leader. Cambridge University Press, Cambridge, United Kingdom, 2001, xxi + 500 pp., $130.00 (hardcover). ISBN 0-521-35281-9. Foundations of Physics, 32(5):807–809, May 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1016013211144.

Rodrigues:2002:CEP

[2591] W. A. Rodrigues, Jr. and M. Sharif. Corrigenda: Equivalence Principle and the Principle of Local Lorentz Invariance. Foundations of Physics, 32(5):811–812, May 2002. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1016065227983. See [2547].

Anonymous:2002:BPLe

[2592] Anonymous. Foundations of Physics Letters. Foundations of Physics, 32(5):813, May 2002. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1016017312053.

Gadella:2002:UMF

[2593] M. Gadella and F. Gómez. A unified mathematical formalism for the Dirac formulation of quantum mechanics. Foundations of Physics, 32 (6):815–869, June 2002. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1016069311589.

Ghose:2002:CTBa

[2594] Partha Ghose. A continuous transition between quantum and classical mechanics. I. Foundations of Physics, 32(6):871–892, June 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1016055128428.

Ghose:2002:CTBb

[2595] Partha Ghose and Manoj K. Samal. A continuous transition between quantum and classical mechanics. II. Foundations of Physics, 32(6):893– 906, June 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1016007212498.

Gour:2002:QPP

[2596] Gilad Gour. The quantum phase problem: Steps toward a resolution. Foundations of Physics, 32(6):907–926, June 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1016059229336.

Puthoff:2002:PVP

[2597] H. E. Puthoff. Polarizable-vacuum (PV) approach to general relativity. Foundations of Physics, 32(6):927–943, June 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1016011413407.

Israelit:2002:QDM

[2598] Mark Israelit. Quintessence and dark matter created by Weyl–Dirac geometry. Foundations of Physics, 32(6):945–961, June 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1016063430245.

Horwitz:2002:RNM

[2599] L. P. Horwitz and Ari Belenkiy. Relativistic notion of mass and a resolution of a conflict between Schopenhauer and Hegel. Foundations of Physics, 32(6):963–979, June 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1016015514315.

Emch:2002:BRB

[2600] Gérard G. Emch. Book review: John von Neumann and the Foundations of Quantum Physics. Edited by Miklós Rédei and Michael Stöltzner.

Kluwer Academic Publishers, Dordrecht, Boston, London, 2001, ix + 371 pages, ISBN 0-7923-6812-6, 134.00 EUR /119.00 USD /89.00 GBP (hardcover). Foundations of Physics, 32(6):981–985, June 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1016067531153.

Anonymous:2002:BPLf

[2601] Anonymous. Foundations of Physics Letters. Foundations of Physics, 32(6):987, June 2002. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1016019615224.

Benioff:2002:TCT

[2602] Paul Benioff. Towards a coherent theory of physics and mathematics. Foundations of Physics, 32(7):989–1029, July 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1016561108807.

VanFlandern:2002:ERS

[2603] Tom Van Flandern and Jean-Pierre Vigier. Experimental repeal of the speed limit for gravitational, electrodynamic, and quantum field interactions. Foundations of Physics, 32(7):1031–1068, July 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1016530625645.

Dzhunushaliev:2002:PSF

[2604] V. Dzhunushaliev. Polarized spacetime foam. Foundations of Physics, 32(7):1069–1090, July 2002. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1016582609715.

Marto:2002:NLI

[2605] J. Marto and J. R. Croca. Non-local interferometry: a causal explanation by means of local wavelet analysis. Foundations of Physics, 32 (7):1091–1109, July 2002. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1016534726554.

Santilli:2002:NPA

[2606] Ruggero Maria Santilli. New problematic aspects of current string theories and their invariant resolution. Foundations of Physics, 32(7):1111– 1140, July 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1016586710624.

Cirkovic:2002:URS

[2607] Milan M. Ćirković. Is the Universe really that simple? Foundations of Physics, 32(7):1141–1157, July 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1016538827462.

Khrennikov:2002:FAE

[2608] Andrei Khrennikov. Frequency analysis of the EPR–Bell argumentation. Foundations of Physics, 32(7):1159–1174, July 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1016590811533.

Dunning-Davies:2002:BRB

[2609] Jeremy Dunning-Davies. Book review: Foundations of Hadronic Chemistry with Applications to New Clean Energies and Fuels. By R. M. Santilli. Kluwer Academic Publishers, Boston/Dordrecht/London, 2001, liv + 397 pp., $138/£95 (hardcover). ISBN 1-4020-0087-1. Foundations of Physics, 32(7):1175–1178, July 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1016542928371.

Anonymous:2002:BPLg

[2610] Anonymous. Foundations of Physics Letters. Foundations of Physics, 32(7):1179, July 2002. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1016594912441.

Klapdor-Kleingrothaus:2002:NDB

[2611] H. V. Klapdor-Kleingrothaus, A. Dietz, and I. V. Krivosheina. Neutrinoless double Beta decay: Status of evidence. Foundations of Physics, 32 (8):1181–1223, August 2002. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1019722802931. See corrigenda [2688].

Choquard:2002:HHJ

[2612] Philippe Choquard and Joël Wagner. The homogeneous Hamilton–Jacobi and Bernoulli equations revisited, II. Foundations of Physics, 32(8): 1225–1249, August 2002. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1019719219769.

Sulcs:2002:IFO

[2613] S. Sulcs, B. C. Gilbert, and C. F. Osborne. On the interference of fullerenes and other massive particles. Foundations of Physics, 32(8): 1251–1271, August 2002. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1019771303840.

Rosenblum:2002:OQE

[2614] Bruce Rosenblum and Fred Kuttner. The observer in the quantum experiment. Foundations of Physics, 32(8):1273–1293, August 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1019723420678.

Mohrhoff:2002:MSW

[2615] Ulrich Mohrhoff. Making sense of a world of clicks. Foundations of Physics, 32(8):1295–1311, August 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1019775404748.

Mohrhoff:2002:WLP

[2616] Ulrich Mohrhoff. Why the laws of physics are just so. Foundations of Physics, 32(8):1313–1324, August 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1019727521587. See remarks [2762, 2763].

Sklar:2002:BRB

[2617] Lawrence Sklar. Book review: The Discovery of Dynamics. By Julian B. Barbour. Oxford University Press, New York, New York, xxiv + 746 pp., $45.00 (softcover). ISBN 0-19-513202-5. Foundations of Physics, 32 (8):1325–1326, August 2002. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1019779505657.

Anonymous:2002:BPLh

[2618] Anonymous. Foundations of Physics Letters. Foundations of Physics, 32(8):1327, August 2002. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1019731622495.

Vuskovic:2002:NCB

[2619] Lepsa Vusković, Dusan Arsenović, and Mirjana Bozić. Non-classical behavior of atoms in an interferometer. Foundations of Physics, 32(9):

1329–1346, September 2002. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1020365405788.

Budinich:2002:GPS

[2620] Paolo Budinich. From the geometry of pure spinors with their division algebras to fermion physics. Foundations of Physics, 32(9):1347–1398, September 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1020307822626.

Duvenhage:2002:NIQ

[2621] Rocco Duvenhage. The nature of information in quantum mechanics. Foundations of Physics, 32(9):1399–1417, September 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1020359806696.

Grotch:2002:BSQ

[2622] H. Grotch and D. A. Owen. Bound states in quantum electrodynamics: Theory and application. Foundations of Physics, 32(9):1419–1457, September 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1020311923535.

Esposito:2002:OMF

[2623] Giampiero Esposito. On the occurrence of mass in field theory. Foundations of Physics, 32(9):1459–1483, September 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1020363907605. See comment

[3132].

Woodward:2002:BRB

[2624] James F. Woodward. Book review: The Future of Spacetime. By Stephen W. Hawking, Kip S. Thorne, Igor Novikov, Timothy Ferris, Alan Lightman, and Richard Price. W. W. Norton, New York and London, 2002, 220 pp., $25.95 (hardcover). ISBN 0-393-02022-3. Foundations of Physics, 32(9):1485–1491, September 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1020316024443.

Anonymous:2002:BPLi

[2625] Anonymous. Foundations of Physics Letters. Foundations of Physics, 32 (9):1493, September 2002. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1020368108514.

Rubin:2002:LEI

[2626] Mark A. Rubin. Locality in the Everett interpretation of quantum field theory. Foundations of Physics, 32(10):1495–1523, October 2002. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1020477902039.

Rizzi:2002:SGR

[2627] Guido Rizzi and Matteo Luca Ruggiero. Space geometry of rotating platforms: an operational approach. Foundations of Physics, 32(10): 1525–1556, October 2002. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1020427318877.

Sheehan:2002:SSM

[2628] D. P. Sheehan, A. R. Putnam, and J. H. Wright. A solid-state Maxwell demon. Foundations of Physics, 32(10):1557–1595, October 2002. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1020479302947.

Garola:2002:SMO

[2629] Claudio Garola. A simple model for an objective interpretation of quantum mechanics. Foundations of Physics, 32(10):1597–1615, October 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1020431419786.

Roa-Neri:2002:AAC

[2630] J. A. E. Roa-Neri and J. L. Jiménez. An alternative approach to the classical dynamics of an extended charged particle. Foundations of Physics, 32(10):1617–1634, October 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1020483403856.

Introzzi:2002:BRB

[2631] Gianluca Introzzi. Book review: Foundations and Interpretation of Quantum Mechanics. By Gennaro Auletta. World Scientific, Singapore, 2000, xxxii + 981 pp., $168,00 (hardcover). ISBN 981-02-4039-2. Foundations of Physics, 32(10):1635–1637, October 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1020435620694.

Anonymous:2002:BPLj

[2632] Anonymous. Foundations of Physics Letters. Foundations of Physics, 32(10):1639, October 2002. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1020487604764.

Blume-Kohout:2002:CMS

[2633] Robin Blume-Kohout, Carlton M. Caves, and Ivan H. Deutsch. Climbing Mount Scalable: Physical resource requirements for a scalable quantum computer. Foundations of Physics, 32(11):1641–1670, November 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1021471621587.

Ungar:2002:HGS

[2634] Abraham A. Ungar. The hyperbolic geometric structure of the density matrix for mixed state qubits. Foundations of Physics, 32(11): 1671–1699, November 2002. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1021446605657.

vBorzeszkowski:2002:LNL

[2635] H.-H. v. Borzeszkowski, B. K. Datta, V. de Sabbata, L. Ronchetti, and H.-J. Treder. Local and non-local aspects of quantum gravity. Foundations of Physics, 32(11):1701–1716, November 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1021498622495.

Friedman:2002:RLS

[2636] Yaakov Friedman and Yuriy Gofman. Relativistic linear spacetime transformations based on symmetry. Foundations of Physics, 32(11):1717– 1736, November 2002. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1021450706566.

Fuss:2002:LGF

[2637] Thomas Fuß. SU(3) local gauge field theory as effective dynamics of composite gluons. Foundations of Physics, 32(11):1737–1755, November 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1021402823404.

Luo:2002:MSE

[2638] Shunlong Luo. Maximum Shannon entropy, minimum Fisher information, and an elementary game. Foundations of Physics, 32(11):1757– 1772, November 2002. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1021454807474.

Myrvold:2002:MIR

[2639] Wayne C. Myrvold. Modal interpretations and relativity. Foundations of Physics, 32(11):1773–1784, November 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1021406924313.

daCosta:2002:TT

[2640] Newton C. A. da Costa and Adonai S. Sant’Anna. Time in thermodynamics. Foundations of Physics, 32(11):1785–1796, November 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1021458908383.

Cirkovic:2002:BRB

[2641] Milan M. Ćirković. Book review: Anthropic Bias: Observation Selection Effects in Science and Philosophy. By Nick Bostrom. Routledge, New York and London, 2002, xiii + 224 pp., $70 (hardcover). ISBN 0-415-93858-9. Foundations of Physics, 32(11):1797–1801, November 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1021411025221.

Anonymous:2002:C

[2642] Anonymous. Corrigenda. Foundations of Physics, 32(11):1803, November 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1021463009292.

Anonymous:2002:BPLk

[2643] Anonymous. Foundations of Physics Letters. Foundations of Physics, 32 (11):1805, November 2002. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1021415126130.

Horwitz:2002:P

[2644] L. P. Horwitz. Preface. Foundations of Physics, 32(12):1807–1808, December 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1022339031738.

Makela:2002:BHA

[2645] Jarmo Mäkelä. Black holes as atoms. Foundations of Physics, 32(12): 1809–1849, December 2002. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1022362515809.

Barvinsky:2002:DSC

[2646] Andrei Barvinsky, Saurya Das, and Gabor Kunstatter. Discrete spectra of charged black holes. Foundations of Physics, 32(12):1851–1862, December 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1022314632647.

Staruszkiewicz:2002:QMS

[2647] Andrzej Staruszkiewicz. Quantum mechanics of the supplementary series. Foundations of Physics, 32(12):1863–1876, December 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1022366616717.

Davies:2002:HFC

[2648] P. C. W. Davies and Tamara M. Davis. How far can the Generalized Second Law be generalized? Foundations of Physics, 32(12):1877–1889, December 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1022318700787.

Gibbons:2002:MTP

[2649] G. W. Gibbons. The maximum tension principle in general relativity. Foundations of Physics, 32(12):1891–1901, December 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1022370717626.

Comer:2002:DNS

[2650] G. L. Comer. Do neutron star gravitational waves carry superfluid imprints? Foundations of Physics, 32(12):1903–1942, December 2002. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1022322801696.

Anonymous:2002:AIV

[2651] Anonymous. Author index to volume 32. Foundations of Physics, 32(12): 1943–1944, December 2002. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1022342114922.

Anonymous:2002:BPLl

[2652] Anonymous. Foundations of Physics Letters. Foundations of Physics, 32 (12):1945, December 2002. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1022374818534.

Zaslavskii:2003:RSC

[2653] O. B. Zaslavskii. Regular self-consistent geometries with infinite quantum backreaction in 2D dilaton gravity and black hole thermodynamics: Unfamiliar features of familiar models. Foundations of Physics, 33 (1):1–35, January 2003. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1022837108018.

Sinha:2003:BHF

[2654] Sukanya Sinha, Alpan Raval, and B. L. Hu. Black hole fluctuations and backreaction in stochastic gravity. Foundations of Physics, 33(1):37–64, January 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1022815724856.

Mann:2003:EAG

[2655] R. B. Mann. Expanding the area of gravitational entropy. Foundations of Physics, 33(1):65–86, January 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1022867708927.

Fulling:2003:IBF

[2656] S. A. Fulling, B.-G. Englert, and M. D. Pilloff. Interacting bosons at finite temperature: How Bogolubov visited a black hole and came home again. Foundations of Physics, 33(1):87–110, January 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1022819825765.

Winstanley:2003:ECC

[2657] Elizabeth Winstanley. On the existence of conformally coupled scalar field hair for black holes in (anti-)de Sitter space. Foundations of Physics, 33(1):111–143, January 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1022871809835.

Barbon:2003:RBH

[2658] J. L. F. Barbón and E. Rabinovici. Remarks on black hole instabilities and closed string tachyons. Foundations of Physics, 33(1):145–165, January 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1022823926674.

Reznik:2003:EV

[2659] Benni Reznik. Entanglement from the vacuum. Foundations of Physics, 33(1):167–176, January 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1022875910744.

Anonymous:2003:BPL

[2660] Anonymous. Foundations of Physics Letters. Foundations of Physics, 33(1):177, January 2003. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1022828027582.

Gerlach:2003:RBE

[2661] Ulrich H. Gerlach. Radiation from bodies with extreme acceleration II: Kinematics. Foundations of Physics, 33(2):179–221, February 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1023712604772.

Katz:2003:TSG

[2662] Joseph Katz. Thermodynamics of self-gravitating systems. Foundations of Physics, 33(2):223–269, February 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1023776921610.

Mukhanov:2003:OBH

[2663] V. F. Mukhanov. On the origin of black-hole entropy. Foundations of Physics, 33(2):271–277, February 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1023729005681.

Dou:2003:BHE

[2664] Djamel Dou and Rafael D. Sorkin. Black-hole entropy as causal links. Foundations of Physics, 33(2):279–296, February 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1023781022519.

Bousso:2003:UES

[2665] Raphael Bousso, Oliver DeWolfe, and Robert C. Myers. Unbounded entropy in spacetimes with positive cosmological constant. Foundations of Physics, 33(2):297–321, February 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1023733106589.

Jacobson:2003:HE

[2666] Ted Jacobson and Renaud Parentani. Horizon entropy. Foundations of Physics, 33(2):323–348, February 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1023785123428.

Volovik:2003:WCQ

[2667] G. E. Volovik. What can the quantum liquid say on the brane black hole, the entropy of an extremal black hole, and the vacuum energy? Foundations of Physics, 33(2):349–368, February 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023762013553.

Anonymous:2003:PJD

[2668] Anonymous. Publications of Jacob D. Bekenstein. Foundations of Physics, 33(2):369–375, February 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1023737207498.

Anonymous:2003:FPLa

[2669] Anonymous. Foundations of Physics Letters: List of contents: Volume 16, number 1, February 2003. Foundations of Physics, 33(2):377, February 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1023789224336.

Rubin:2003:RFP

[2670] Mark A. Rubin. Relative frequency and probability in the Everett interpretation of Heisenberg-picture quantum mechanics. Foundations of Physics, 33(3):379–405, March 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1023765114419.

Kryukov:2003:CFA

[2671] Alexey A. Kryukov. Coordinate formalism on abstract Hilbert space: Kinematics of a quantum measurement. Foundations of Physics, 33 (3):407–443, March 2003. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1023711631257.

Castro:2003:ESR

[2672] Carlos Castro and Alex Granik. Extended scale relativity, p-loop harmonic oscillator, and logarithmic corrections to the black hole entropy. Foundations of Physics, 33(3):445–466, March 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1023763615328.

Cirkovic:2003:TAT

[2673] Milan M. Ćirković. The thermodynamical arrow of time: Reinterpreting the Boltzmann–Schuetz argument. Foundations of Physics, 33(3):467– 490, March 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1023715732166.

Vaidman:2003:MIF

[2674] Lev Vaidman. The meaning of the interaction-free measurements. Foundations of Physics, 33(3):491–510, March 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023767716236.

Lanyi:2003:TEB

[2675] G. Lanyi. Thermal equilibrium between radiation and matter. Foundations of Physics, 33(3):511–528, March 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1023719800307.

vBorzeszkowski:2003:TTD

[2676] H.-H. v. Borzeszkowski and T. Chrobok. Are there thermodynamical degrees of freedom of gravitation? Foundations of Physics, 33(3):529–539, March 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1023771817145.

Mashhoon:2003:BRB

[2677] Bahram Mashhoon. Book review: Relativity: Special, General, and Cosmological. By Wolfgang Rindler. Oxford University Press, New York, New York, 2001, xiii + 428 pp., $95.00/$40.00 (hardcover/softcover). ISBN 0 19 850835 2/ISBN 0 19 850836 0. Foundations of Physics, 33 (3):541–542, March 2003. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1023723901215.

Anonymous:2003:FPLb

[2678] Anonymous. Foundations of Physics Letters: List of contents: Volume 16, number 1, February 2003. Foundations of Physics, 33(3):543, March 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1023775918054.

Aldrovandi:2003:EPR

[2679] R. Aldrovandi, P. B. Barros, and J. G. Pereira. The equivalence principle revisited. Foundations of Physics, 33(4):545–575, April 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1023728002453.

Matteucci:2003:ABP

[2680] G. Matteucci, D. Iencinella, and C. Beeli. The Aharonov–Bohm phase shift and Boyer’s critical considerations: New experimental result but still an open subject? Foundations of Physics, 33(4):577–590, April 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1023766519291.

Mould:2003:QBS

[2681] Richard A. Mould. Quantum brain states. Foundations of Physics, 33 (4):591–612, April 2003. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1023718603362.

Aldrovandi:2003:KSI

[2682] R. Aldrovandi, A. L. Barbosa, M. Calçada, and J. G. Pereira. Kinematics of a spacetime with an infinite cosmological constant. Foundations of Physics, 33(4):613–624, April 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1023770620200.

Ahmedov:2003:CMG

[2683] B. J. Ahmedov and N. I. Rakhmatov. Concerning measurement of gravitomagnetism in electromagnetic systems. Foundations of Physics, 33 (4):625–639, April 2003. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1023722704270.

Svetlichny:2003:CIF

[2684] George Svetlichny. Causality implies formal state collapse. Foundations of Physics, 33(4):641–655, April 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1023774721109.

Garuccio:2003:VCH

[2685] A. Garuccio and V. Berardi. On the validity of Clauser and Horne factorizability. Foundations of Physics, 33(4):657–664, April 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1023726805179.

Ferrero:2003:IIC

[2686] Miguel Ferrero. The information interpretation and the conceptual problems of quantum mechanics. Foundations of Physics, 33(4):665–676, April 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1023778822017.

Ungar:2003:BRB

[2687] Abraham A. Ungar. Book review: Theory of K-Loops. By Hubert Kiechle. Springer, Berlin, 2002, x + 186 pp., $36.80 (softcover). ISBN 3-540-43262-0. Foundations of Physics, 33(4):677–678, April 2003. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1023730906088.

Klapdor-Kleingrothaus:2003:CND

[2688] H. V. Klapdor-Kleingrothaus, A. Dietz, and I. V. Krivosheina. Corrigenda: “Neutrinoless Double Beta Decay: Status of Evidence,” Found. Phys. 32(8), 1181–1223 (2002). Foundations of Physics, 33(4):679–684, April 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1023782922926. See [2611].

Anonymous:2003:FPLc

[2689] Anonymous. Foundations of Physics Letters: List of contents: Volume 16, number 1, February 2003. Foundations of Physics, 33(4):685, April 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1023735006996.

Ahluwalia:2003:P

[2690] D. V. Ahluwalia and M. Kirchbach. Preface. Foundations of Physics, 33(5):687–688, May 2003. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1025647104639.

Kaku:2003:HSS

[2691] Michio Kaku. Highlights of supersymmetry, superstrings, and M-theory. Foundations of Physics, 33(5):689–706, May 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1025688621477.

Sudarshan:2003:HSF

[2692] E. C. G. Sudarshan. Higher spin fields and non-holonomic constraints. Foundations of Physics, 33(5):707–717, May 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1025640705548.

Castro:2003:UPG

[2693] Gabriel López Castro and Alejandro Mariano. Unstable particles, gauge invariance and the Δ^{++} resonance parameters. Foundations of Physics, 33(5):719–734, May 2003. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1025692722386.

Villanueva:2003:NME

[2694] V. M. Villanueva, J. A. Nieto, and O. Obregón. Non-minimal electromagnetic coupling for Spin-3/ 2 fields. Foundations of Physics, 33 (5):735–740, May 2003. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1025644806456.

Napsuciale:2003:PIE

[2695] Mauro Napsuciale. “Principle of Indistinguishability” and equations of motion for particles with spin. Foundations of Physics, 33(5):741– 768, May 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1025696823295.

Kato:2003:FAM

[2696] A. Kato, G. Muñoz, D. Singleton, J. Dryzek, and V. Dzhunushaliev. Field angular momentum. Foundations of Physics, 33(5):769–780, May 2003.

CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1025649007365.

Kirchbach:2003:QSI

[2697] M. Kirchbach. Quantum states of indefinite spins: From baryons to massive gravitino. Foundations of Physics, 33(5):781–812, May 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1025601124203.

Klapdor-Kleingrothaus:2003:FEN

[2698] H. V. Klapdor-Kleingrothaus. First evidence for neutrinoless double Beta decay. Foundations of Physics, 33(5):813–829, May 2003. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1025653108273.

Klapdor-Kleingrothaus:2003:SCD

[2699] H. V. Klapdor-Kleingrothaus and I. V. Krivosheina. Search for cold dark matter and solar neutrinos with GENIUS and GENIUS–TF. Foundations of Physics, 33(5):831–837, May 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1025760925112.

Gonzalez-Robles:2003:AGB

[2700] V. M. González-Robles. About a generalization of Bell’s inequality. Foundations of Physics, 33(5):839–853, May 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1025605209182.

Gallegos:2003:EFT

[2701] A. Gallegos and J. L. M. Lucio. An effective field theory approach to f_{0}(980) − a_{0}(980) mixing. Foundations of Physics, 33(5):855–862, May 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1025609326020.

Espinoza:2003:MFK

[2702] Augusto Espinoza and Andrew Chubykalo. Mathematical foundation of Kapitsa’s hypothesis about the origin and structure of ball lightning. Foundations of Physics, 33(5):863–873, May 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1025661310091.

Anonymous:2003:FPLd

[2703] Anonymous. Foundations of Physics Letters: List of contents: Volume 16, number 1, February 2003. Foundations of Physics, 33(5): 875, May 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1025613426929.

Castagnino:2003:GAT

[2704] Mario Castagnino, Olimpia Lombardi, and Luis Lara. The global arrow of time as a geometrical property of the Universe. Foundations of Physics, 33(6):877–912, June 2003. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1025665410999.

Verschl:2003:RST

[2705] M. Verschl and M. Sorg. Relativistic Schrödinger theory and the Hartree– Fock approach. Foundations of Physics, 33(6):913–954, June 2003. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1025617527838.

Tarzi:2003:EPR

[2706] S. Tarzi. Exclusion principles as restricted permutation symmetries. Foundations of Physics, 33(6):955–979, June 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1025669511908.

Ghosal:2003:RSE

[2707] S. K. Ghosal, Biplab Raychaudhuri, Anjan Kumar Chowdhury, and Minakshi Sarker. Relativistic Sagnac effect and Ehrenfest paradox. Foundations of Physics, 33(6):981–1001, June 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1025621628746.

Dieks:2003:BRB

[2708] Dennis Dieks. Book review: Foundations of Quantum Mechanics, an Empiricist Approach. By Willem M. de Muynck. Kluwer Academic Publishers, Dordrecht/Boston/London, 2002, xxiv + 680 pp., $219.00 (hardcover). ISBN 1-4020-0932-1. Foundations of Physics, 33(6):1003–1006, June 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1025673612816.

Anonymous:2003:FPLe

[2709] Anonymous. Foundations of Physics Letters: List of contents: Volume 16, number 1, February 2003. Foundations of Physics, 33(6):1007, June 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1025625729655.

Hrasko:2003:TCT

[2710] P. Hraskó. Time correlation in tunneling of photons. Foundations of Physics, 33(7):1009–1031, July 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1025677700523.

Cooperstock:2003:EDW

[2711] F. I. Cooperstock and S. Tieu. The energy of a dynamical wave-emitting system in general relativity. Foundations of Physics, 33(7):1033–1059, July 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1025633817361.

Nicrosini:2003:RSL

[2712] Oreste Nicrosini and Alberto Rimini. Relativistic spontaneous localization: a proposal. Foundations of Physics, 33(7):1061–1084, July 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1025685801431.

Hyllus:2003:PIA

[2713] Philipp Hyllus and Erik Sjöqvist. Precession and interference in the Aharonov–Casher and scalar Aharonov–Bohm effects. Foundations of Physics, 33(7):1085–1105, July 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1025637918269. See comment [3157].

Beil:2003:FGR

[2714] R. G. Beil. Finsler geometry and relativistic field theory. Foundations of Physics, 33(7):1107–1127, July 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1025689902340.

Hanson:2003:WWC

[2715] Robin Hanson. When worlds collide: Quantum probability from observer selection? Foundations of Physics, 33(7):1129–1150, July 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1025642019178.

Anonymous:2003:FPLf

[2716] Anonymous. Foundations of Physics Letters: List of contents: Volume 16, number 2, April 2003. Foundations of Physics, 33(7):1151, July 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1025694003248.

Horwitz:2003:P

[2717] L. P. Horwitz. Preface. Foundations of Physics, 33(8):1153–1156, August 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1025646122716.

Land:2003:HOK

[2718] Martin Land. Higher-order Kinetic term for controlling photon mass in off–shell electrodynamics. Foundations of Physics, 33(8):1157–1175, August 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1025670806787.

Oron:2003:CMA

[2719] O. Oron and L. P. Horwitz. The conformal metric associated with the U(1) gauge of the Stueckelberg–Schrödinger equation. Foundations of Physics, 33(8):1177–1187, August 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1025622923625.

Fanchi:2003:RDT

[2720] John R. Fanchi. Relativistic dynamical theory of particle decay and application to K-mesons. Foundations of Physics, 33(8):1189–1205, August 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1025674907695.

Harpaz:2003:EME

[2721] Amos Harpaz and Noam Soker. Equation of motion of an electric charge. Foundations of Physics, 33(8):1207–1221, August 2003. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1025627024534.

Piazzese:2003:ERA

[2722] F. I. Piazzese. On the energy radiation from an accelerated charge. Foundations of Physics, 33(8):1223–1236, August 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1025679008604.

Gersten:2003:ESR

[2723] Alexander Gersten. Euclidean special relativity. Foundations of Physics, 33(8):1237–1251, August 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1025631125442.

Alfred:2003:TCT

[2724] Marcus Alfred and James Lindesay. A test of the calculability of a threebody relativistic, cluster decomposable, unitary, covariant scattering theory. Foundations of Physics, 33(8):1253–1264, August 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1025683109512.

Anonymous:2003:FPLg

[2725] Anonymous. Foundations of Physics Letters: List of contents: Volume 16, number 2, April 2003. Foundations of Physics, 33(8):1265, August 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1025635226351.

Baugh:2003:TD

[2726] James Baugh, David Ritz Finkelstein, Andrei Galiautdinov, and Mohsen Shiri-Garakani. Transquantum dynamics. Foundations of Physics, 33(9): 1267–1275, September 2003. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1025661209919.

Pavsic:2003:CSA

[2727] Matej Pavsic. Clifford space as the arena for physics. Foundations of Physics, 33(9):1277–1306, September 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1025637126758.

Mallett:2003:GFC

[2728] Ronald L. Mallett. The gravitational field of a circulating light beam. Foundations of Physics, 33(9):1307–1314, September 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1025689110828.

Pitts:2003:NFR

[2729] J. Brian Pitts and W. C. Schieve. Nonsingularity of flat Robertson– Walker models in the special relativistic approach to Einstein’s equations. Foundations of Physics, 33(9):1315–1321, September 2003. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1025641327666.

Oron:2003:EAW

[2730] O. Oron and L. P. Horwitz. Eikonal approximation to 5D wave equations and the 4D space–time metric. Foundations of Physics, 33(9): 1323–1338, September 2003. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1025693311737.

Ivezic:2003:PST

[2731] Tomislav Ivezić. The proof that the standard transformations of E and B are not the Lorentz transformations. Foundations of Physics, 33(9): 1339–1347, September 2003. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1025645428575.

OHara:2003:RIS

[2732] Paul O’Hara. Rotational invariance and the spin-statistics theorem. Foundations of Physics, 33(9):1349–1368, September 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1025697412645.

Ben-Amots:2003:BAR

[2733] N. Ben-Amots. Basic aspect of relativistic rotation: Franklin rotation of a sphere. Foundations of Physics, 33(9):1369–1372, September 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1025649529483.

Santilli:2003:EIG

[2734] Ruggero Maria Santilli. Elements of iso-, geno-, hyper-mathematics for matter, their isoduals for antimatter, and their applications in physics, chemistry, and biology. Foundations of Physics, 33(9):1373– 1416, September 2003. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1025601613554.

Anonymous:2003:FPLh

[2735] Anonymous. Foundations of Physics Letters: List of contents: Volume 16, number 2, April 2003. Foundations of Physics, 33(9):1417, September 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1025653630392.

Greenberger:2003:PDM

[2736] Daniel Greenberger and Abner Shimony. The presence of David Mermin. Foundations of Physics, 33(10):1419–1422, October 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1026033228750.

Griffiths:2003:PQR

[2737] Robert B. Griffiths. Probabilities and quantum reality: are there correlata? Foundations of Physics, 33(10):1423–1459, October 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1026092212820.

Schack:2003:QTF

[2738] Rüdiger Schack. Quantum theory from four of Hardy’s axioms. Foundations of Physics, 33(10):1461–1468, October 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1026044329659.

Leggett:2003:NHV

[2739] A. J. Leggett. Nonlocal hidden-variable theories and quantum mechanics: an incompatibility theorem. Foundations of Physics, 33(10):1469–1493, October 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1026096313729.

Collett:2003:WCR

[2740] Brian Collett and Philip Pearle. Wavefunction collapse and random walk. Foundations of Physics, 33(10):1495–1541, October 2003. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1026048530567.

Peres:2003:WWO

[2741] Asher Peres. What’s wrong with these observables? Foundations of Physics, 33(10):1543–1547, October 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1026000614638.

Wootters:2003:WTF

[2742] William K. Wootters. Why things fall. Foundations of Physics, 33(10): 1549–1557, October 2003. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1026052631476.

Anonymous:2003:FPLi

[2743] Anonymous. Foundations of Physics Letters: List of contents: Volume 16, number 3, June 2003. Foundations of Physics, 33(10):1559, October 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1026004715546.

Clifton:2003:CQT

[2744] Rob Clifton, Jeffrey Bub, and Hans Halvorson. Characterizing quantum theory in terms of information-theoretic constraints. Foundations of Physics, 33(11):1561–1591, November 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1026056716397.

Brassard:2003:QCC

[2745] Gilles Brassard. Quantum communication complexity. Foundations of Physics, 33(11):1593–1616, November 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1026009100467.

Bruss:2003:ESQ

[2746] Dagmar Bruß and Chiara Macchiavello. On the entanglement structure in quantum cloning. Foundations of Physics, 33(11):1617–1628, November 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1026061117305.

DiVincenzo:2003:HQD

[2747] David P. DiVincenzo, Patrick Hayden, and Barbara M. Terhal. Hiding quantum data. Foundations of Physics, 33(11):1629–1647, November 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1026013201376.

Plotnitsky:2003:MMC

[2748] Arkady Plotnitsky. Mysteries without mysticism and correlations without correlata: On quantum knowledge and knowledge in general. Foundations of Physics, 33(11):1649–1689, November 2003. CODEN FNDPA4.

ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1026065218214.

Elser:2003:MFP

[2749] Veit Elser. The Mermin fixed point. Foundations of Physics, 33(11): 1691–1698, November 2003. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1026017302284.

Shimony:2003:FPD

[2750] Abner Shimony. Four poems dedicated to David Mermin. Foundations of Physics, 33(11):1699–1700, November 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1026069319123.

Anonymous:2003:FPLj

[2751] Anonymous. Foundations of Physics Letters: List of contents: Volume 16, number 3, June 2003. Foundations of Physics, 33(11):1701, November 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1026021403193.

Lifshitz:2003:QMD

[2752] Ron Lifshitz. Quasicrystals: a matter of definition. Foundations of Physics, 33(12):1703–1711, December 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1026247120031.

Saslow:2003:MTB

[2753] Wayne M. Saslow. Multicarrier transport: Batteries, semiconductors, mixed ionic–electronic conductors, and biology. Foundations of Physics, 33(12):1713–1734, December 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1026225504102.

Rehr:2003:FQP

[2754] J. J. Rehr. Failure of the quasiparticle Picture of X-ray absorption? Foundations of Physics, 33(12):1735–1742, December 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1026277520940.

Marder:2003:SEL

[2755] M. Marder. The shape of the edge of a leaf. Foundations of Physics, 33(12):1743–1768, December 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1026229605010.

Rabson:2003:CA

[2756] David A. Rabson, John F. Huesman, and Benji N. Fisher. Cohomology for anyone. Foundations of Physics, 33(12):1769–1796, December 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1023/A%3A1026281621848.

Anonymous:2003:PDM

[2757] Anonymous. Publications of N. David Mermin. Foundations of Physics, 33(12):1797–1809, December 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/A%3A1026233805919.

Anonymous:2003:AIV

[2758] Anonymous. Author index to volume 33. Foundations of Physics, 33(12): 1811–1812, December 2003. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/A%3A1026208419070.

Anonymous:2003:FPLk

[2759] Anonymous. Foundations of Physics Letters: List of contents: Volume 16, number 3, June 2003. Foundations of Physics, 33(12):1813–1814, December 2003. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/A%
3A1026285822757.

Grandy:2004:TEMa

[2760] W. T. Grandy,, Jr. Time evolution in macroscopic systems. I. Equations of motion. Foundations of Physics, 34(1):1–20, January 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/B%
3AFOOP.0000012007.06843.ed.

Grandy:2004:TEMb

[2761] W. T. Grandy, Jr. Time evolution in macroscopic systems. II. The entropy. Foundations of Physics, 34(1):21–57, January 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://

link.springer.com/article/10.1023/B%3AFOOP.0000012008.36856.
c1.

Marchildon:2004:RMI

[2762] Louis Marchildon. Remarks on Mohrhoff’s interpretation of quantum mechanics. Foundations of Physics, 34(1):59–73, January 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/B%
3AFOOP.0000012009.79192.ff. See [2616, 2763].

Mohrhoff:2004:DQS

[2763] Ulrich Mohrhoff. Do quantum states evolve? Apropos of Marchildon’s remarks. Foundations of Physics, 34(1):75–97, January 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://

link.springer.com/article/10.1023/B%3AFOOP.0000012010.89014.

f4. See [2616, 2762].

Schust:2004:QER

[2764] P. Schust, M. Mattes, and M. Sorg. Quantum entanglement in relativistic three-particle systems. Foundations of Physics, 34(1):99–167, January 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/B%
3AFOOP.0000012011.48629.5a.

Lindesay:2004:CPT

[2765] James Lindesay and Tepper Gill. Canonical proper time formulation for physical systems. Foundations of Physics, 34(1):169–182, January 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/B%
3AFOOP.0000012012.27956.d3.

Vacchini:2004:BRB

[2766] Bassano Vacchini. Book review: The Theory of Open Quantum Systems. By H.-P. Breuer and F. Petruccione. Oxford University Press, New York, New York, 2002, xxi + 625 pp., $55.00 (hardcover). ISBN 0-19-852063-8. Foundations of Physics, 34(1):183–186, January 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/B%
3AFOOP.0000012127.36628.af.

Greenberg:2004:BRB

[2767] O. W. Greenberg. Book review: Quantum Field Theory in a Nutshell. By A. Zee, Princeton University Press, Princeton, NJ, 2003, ISBN: 0-691-01019-6, xv + 518 pp. $49.50 (hardcover). Foundations of Physics, 34 (1):187–188, January 2004. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/B%3AFOOP.0000012128.27150.89.

Esposito:2004:E

[2768] G. Esposito. Erratum. Foundations of Physics, 34(1):189, January 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/B%
3AFOOP.0000012013.82420.12.

Anonymous:2004:FPLa

[2769] Anonymous. Foundations of Physics Letters: List of contents: Volume 16, number 4, August 2003. Foundations of Physics, 34(1):191, January 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/B%
3AFOOP.0000012129.00640.e1.

Caves:2004:GTD

[2770] Carlton M. Caves, Christopher A. Fuchs, Kiran K. Manne, and Joseph M. Renes. Gleason-type derivations of the quantum probability rule for generalized measurements. Foundations of Physics, 34(2):193–209, February 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/B%
3AFOOP.0000019581.00318.a5.

Pitts:2004:NCE

[2771] J. Brian Pitts and W. C. Schieve. Null cones and Einstein’s equations in Minkowski spacetime. Foundations of Physics, 34(2):211–238, February 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/B%
3AFOOP.0000019582.44548.6a.

Cirkovic:2004:FNE

[2772] Milan M. Ćirković. Forecast for the next eon: applied cosmology and the long-term fate of intelligent beings. Foundations of Physics, 34(2): 239–261, February 2004. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/B%3AFOOP.0000019583.67831.60.

Puccini:2004:AFG

[2773] G. Puccini and H. Vucetich. Axiomatic foundations of Galilean quantum field theories. Foundations of Physics, 34(2):263–295, February 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/B%
3AFOOP.0000019584.39535.28.

Bell:2004:QDT

[2774] Sarah B. M. Bell, John P. Cullerne, and Bernard M. Diaz. QED derived from the two-body interaction. Foundations of Physics, 34(2): 297–333, February 2004. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/B%3AFOOP.0000019585.77172.33.

Bell:2004:TBI

[2775] Sarah B. M. Bell, John P. Cullerne, and Bernard M. Diaz. The twobody interaction with a circle in time. Foundations of Physics, 34(2): 335–358, February 2004. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/B%3AFOOP.0000019586.03478.18.

Anonymous:2004:FPLb

[2776] Anonymous. Foundations of Physics Letters: List of contents: Volume 16, number 4, August 2003. Foundations of Physics, 34(2):359, February 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/B%
3AFOOP.0000019692.81570.87.

Krol:2004:BIQ

[2777] Jerzy Król. Background independence in quantum gravity and forcing constructions. Foundations of Physics, 34(3):361–403, March 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/B%
3AFOOP.0000019620.04821.a2.

Bohr:2004:PUQ

[2778] Aage Bohr, Ben R. Mottelson, and Ole Ulfbeck. The principle underlying quantum mechanics. Foundations of Physics, 34(3):405–417, March 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/B%
3AFOOP.0000019621.02554.7e.

Gambetta:2004:MDP

[2779] Jay Gambetta and H. M. Wiseman. Modal dynamics for positive operator measures. Foundations of Physics, 34(3):419–448, March 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/B%
3AFOOP.0000019622.81881.f8.

Garola:2004:LMW

[2780] Claudio Garola and Jarosław Pykacz. Locality and measurements within the SR model for an objective interpretation of quantum mechanics. Foundations of Physics, 34(3):449–475, March 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/B%3AFOOP.0000019623.41496.fc.

Maclay:2004:GSO

[2781] G. Jordan Maclay and Robert L. Forward. A Gedanken spacecraft that operates using the quantum vacuum (dynamic Casimir effect). Foundations of Physics, 34(3):477–500, March 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/B%3AFOOP.0000019624.51662.50.

Strocchi:2004:RQM

[2782] F. Strocchi. Relativistic quantum mechanics and field theory. Foundations of Physics, 34(3):501–527, March 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/B%3AFOOP.0000019625.30165.35.

DallaChiara:2004:BRQ

[2783] Maria Luisa Dalla Chiara. Book review: Quantum Mechanics, Mathematics, Cognition and Action: Proposals for a Formalized Epistemology. Mioara Mugur–Schächter and Alwyn van der Merwe, eds., Kluwer Academic, Dordrecht, The Netherlands, 2002, xviii + 493 pp., $191.00 (hardcover). ISBN 1-4020-1120-2. Foundations of Physics, 34(3):529– 532, March 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/B%
3AFOOP.0000019696.81352.ce.

Grandy:2004:BRBa

[2784] W. T. Grandy, Jr. Book review: Probability Theory: The Logic of Science. By Edwin T. Jaynes, edited by G. Larry Bretthorst. Cambridge University Press, Cambridge, United Kingdom, 2003, xxix + 727 pp., S60.00 (hardcover). ISBN 0-521-59271-2. Foundations of Physics, 34 (3):533–536, March 2004. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/B%3AFOOP.0000019697.60679.4e.

Obata:2004:BRB

[2785] Tsunehiro Obata. Book review: Theory of High Temperature Superconductivity. By Shigeji Fujita and Salvador Godoy. Fundamental Theories of Physics, Vol. 121, Kluwer Academic Publishers, Dordrecht, The Netherlands, 2002, xix + 367, $139.50 (hardcover). ISBN 1-4020-0149-5. Foundations of Physics, 34(3):537–540, March 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/B%3AFOOP.0000019698.75239.16.

Cirkovic:2004:BRB

[2786] Milan M. Ćirković. Book review: Modern Cosmology. By Scott Dodelson. Academic Press, London, San Diego, California, 2003. xiii + 440 pp., $70 (hardcover). ISBN 0-12-219141-2. Foundations of Physics, 34 (3):541–544, March 2004. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/B%3AFOOP.0000019699.88151.ed.

Anonymous:2004:FPLc

[2787] Anonymous. Foundations of Physics Letters: List of contents: Volume 16, number 6, December 2003. Foundations of Physics, 34(3):545, March 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/B%
3AFOOP.0000019700.53468.2a.

Roa-Neri:2004:RRN

[2788] J. A. E. Roa-Neri and J. L. Jiménez. Radiation reaction of a nonrelativistic quantum charged particle. Foundations of Physics, 34(4):547– 580, April 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/B%
3AFOOP.0000019626.21305.c8.

Alfred:2004:NFP

[2789] Marcus Alfred, Petero Kwizera, James V. Lindesay, and H. Pierre Noyes. A nonperturbative, finite particle number approach to relativistic scattering theory. Foundations of Physics, 34(4):581–616, April 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/B%
3AFOOP.0000019627.19038.50.

Cardone:2004:KSGa

[2790] Fabio Cardone, Alessio Marrani, and Roberto Mignani. Killing symmetries of generalized Minkowski spaces. I. Algebraic–infinitesimal structure of spacetime rotation groups. Foundations of Physics, 34(4):617– 641, April 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/B%
3AFOOP.0000019628.97334.f0.

Puccini:2004:STA

[2791] Gabriel D. Puccini and Héctor Vucetich. Steps towards the axiomatic foundations of the relativistic quantum field theory: Spin-statistics, commutation relations, and CPT theorems. Foundations of Physics, 34 (4):643–667, April 2004. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/B%3AFOOP.0000019629.12926.2b.

Lavis:2004:SES

[2792] D. A. Lavis. The spin–echo system reconsidered. Foundations of Physics, 34(4):669–688, April 2004. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/B%3AFOOP.0000019630.61758.b6.

Khrennikov:2004:RBP

[2793] Andrei Y. Khrennikov and Elena R. Loubenets. On relations between probabilities under quantum and classical measurements. Foundations of Physics, 34(4):689–704, April 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/B%3AFOOP.0000019631.84010.a6.

Selleri:2004:BRB

[2794] F. Selleri. Book review: Towards a Nonlinear Quantum Physics. By J. R. Croca. World Scientific, Singapore, 2003, xi + 215 pp. (hardcover). Foundations of Physics, 34(4):705–710, April 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/B%3AFOOP.0000019701.29636.5b.

Fujita:2004:BRB

[2795] S. Fujita. Book review: Methods of Statistical Physics. By Tomoyasu Tanaka. Cambridge University Press, Cambridge, United Kingdom, 2002, xv + 293 pp., $110.00/$40.00 (hardcover/softcover). ISBN 0-521-58056-0/0-521-58958-4. Foundations of Physics, 34(4):711–713, April 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/B%
3AFOOP.0000019702.00446.27.

Anonymous:2004:FPLd

[2796] Anonymous. Foundations of Physics Letters: List of contents: Volume 17, number 1, February 2004. Foundations of Physics, 34(4):715, April 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/B%
3AFOOP.0000019703.01751.c9.

deMuynck:2004:TNC

[2797] Willem M. de Muynck. Towards a neo-Copenhagen interpretation of quantum mechanics. Foundations of Physics, 34(5):717–770, May 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/B%
3AFOOP.0000022186.59877.0c.

Grandy:2004:TEMc

[2798] W. T. Grandy, Jr. Time evolution in macroscopic systems. III: Selected applications. Foundations of Physics, 34(5):771–813, May 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/B%
3AFOOP.0000022187.45866.81.

Armour:2004:SMF

[2799] Rollin S. Armour, Jr. Spin-1/2 Maxwell fields. Foundations of Physics, 34(5):815–842, May 2004. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/B%3AFOOP.0000022188.90097.10.

Krol:2004:ESN

[2800] Jerzy Król. Exotic smoothness and noncommutative spaces. The model-theoretical approach. Foundations of Physics, 34(5):843–869, May 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/B%
3AFOOP.0000022189.71690.34.

Drago:2004:FWQ

[2801] A. Drago and S. Esposito. Following Weyl on quantum mechanics: The contribution of Ettore Majorana. Foundations of Physics, 34(5):871– 887, May 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/B%
3AFOOP.0000022190.47446.f3.

Anonymous:2004:FPLe

[2802] Anonymous. Foundations of Physics Letters: List of contents: Volume 17, number 2, April 2004. Foundations of Physics, 34(5):889, May 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/B%
3AFOOP.0000022290.46764.37.

Olavo:2004:FQM

[2803] L. S. F. Olavo. Foundations of quantum mechanics: The connection between QM and the Central Limit Theorem. Foundations of Physics, 34(6):891–935, June 2004. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/B%3AFOOP.0000034222.65544.26.

Raju:2004:EBP

[2804] C. K. Raju. The electrodynamic 2-body problem and the origin of quantum mechanics. Foundations of Physics, 34(6):937–962, June 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/B%
3AFOOP.0000034223.58332.d4.

Struyve:2004:PSO

[2805] W. Struyve, W. De Baere, J. De Neve, and S. De Weirdt. On Peres’ statement “Opposite Momenta Lead to Opposite Directions,” decaying systems and optical imaging. Foundations of Physics, 34(6):963– 985, June 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/B%
3AFOOP.0000034224.37659.96.

Byun:2004:LID

[2806] Yanghyun Byun. Lorentz invariant decompositions of the state vector spaces and the basis problem. Foundations of Physics, 34(6):987– 1003, June 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/B%
3AFOOP.0000034225.69458.a0.

Herrmann:2004:SAD

[2807] Heiko Herrmann, W. Muschik, G. Rückner, and H.-H. von Borzeszkowski. Spin axioms in different geometries of relativistic continuum physics. Foundations of Physics, 34(6):1005–1021, June 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/B%3AFOOP.0000034226.19527.0e.

Rockwood:2004:BRB

[2808] Alyn P. Rockwood. Book review: Geometric Algebra for Physicists, Chris Doran and Anthony Lasenby, Cambridge University Press, Cambridge, U.K., 2003, xiv + 578 pp., $95.00 (hardcover). ISBN 0-521-48022-1. Foundations of Physics, 34(6):1023–1026, June 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/B%3AFOOP.0000034227.17261.10.

Anonymous:2004:FPLf

[2809] Anonymous. Foundations of Physics Letters: List of contents: Volume 17, number 3, June 2004. Foundations of Physics, 34(6):1027, June 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/B%
3AFOOP.0000034228.10942.8e.

Wheeler:2004:RCG

[2810] John C. Wheeler. Resolution of a classical gravitational Second-Law Paradox. Foundations of Physics, 34(7):1029–1062, July 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://

link.springer.com/article/10.1023/B%3AFOOP.0000037622.22389.
20.

Zafiris:2004:QES

[2811] Elias Zafiris. Quantum event structures from the perspective of Grothendieck topoi. Foundations of Physics, 34(7):1063–1090, July 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/B%
3AFOOP.0000037623.08379.df.

Castro:2004:CMS

[2812] Carlos Castro. The charge–mass–spin relation of Clifford polyparticles, Kerr–Newman black holes and the fine structure constant. Foundations of Physics, 34(7):1091–1113, July 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1023/B%3AFOOP.0000037624.96564.25.

Anonymous:2004:FPLg

[2813] Anonymous. Foundations of Physics Letters: List of contents: Volume 17, number 4, August 2004. Foundations of Physics, 34(7):1115, July 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/B%
3AFOOP.0000037625.75892.53.

Cordes:2004:SDH

[2814] H. O. Cordes. The split of the Dirac Hamiltonian into precisely predictable energy components. Foundations of Physics, 34(8):1117–1153, August 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/B%
3AFOOP.0000041287.25713.fe.

Cardone:2004:KSGb

[2815] Fabio Cardone, Alessio Marrani, and Roberto Mignani. Killing symmetries of generalized Minkowski spaces. Part 2: Finite structure of space– time rotation groups in four dimensions. Foundations of Physics, 34(8): 1155–1201, August 2004. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/B%3AFOOP.0000041288.74545.c1.

Bar:2004:IWS

[2816] D. Bar. Internet Websites statistics expressed in the framework of the Ursell–Mayer cluster formalism. Foundations of Physics, 34(8): 1203–1223, August 2004. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/B%3AFOOP.0000041289.82512.8f.

Kryukov:2004:PEC

[2817] Alexey A. Kryukov. On the problem of emergence of classical space– time: The quantum-mechanical approach. Foundations of Physics, 34 (8):1225–1248, August 2004. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/B%3AFOOP.0000041290.67197.a7.

Garola:2004:SAC

[2818] Claudio Garola and Sandro Sozzo. A semantic approach to the completeness problem in quantum mechanics. Foundations of Physics, 34 (8):1249–1266, August 2004. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/B%3AFOOP.0000041291.04835.e9.

Woodward:2004:BRB

[2819] James F. Woodward. Book review: The Fabric of the Cosmos. By Brian Greene, Alfred A. Knopf, New York, U.S.A., 2004, xii + 569 pp., $28.95 (hardcover). Foundations of Physics, 34(8):1267–1273, August 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/B%
3AFOOP.0000041439.22726.8b.

Spector:2004:BRB

[2820] Donald Spector. Book review: Essential Mathematical Methods for Physicists. By Hans Weber and George Arfken, Academic Press, San Diego, California, U.S.A., 2003, xxii + 932 pp., $89.95 (hardcover). ISBN 0-12-059877-9. Foundations of Physics, 34(8):1275–1280, August 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/B%
3AFOOP.0000041440.46009.eb.

Lusanna:2004:BRB

[2821] Luca Lusanna. Book review: Relativity in Rotating Frames. Relativistic Physics in Rotating Reference Frames. Edited by G. Rizzi and M.

L. Ruggiero, (Fundamental Theories of Physics 135), 452 pp., $193.00. ISBN 1-4020-1805-3. Foundations of Physics, 34(8):1281–1282, August 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/B%
3AFOOP.0000041441.83646.60.

Anonymous:2004:FPLh

[2822] Anonymous. Foundations of Physics Letters: List of contents: Volume 17, number 5, October 2004. Foundations of Physics, 34(8):1283, August 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/B%
3AFOOP.0000041442.21284.f4.

Gyenis:2004:WCS

[2823] Balázs Gyenis and Miklós Rédei. When can statistical theories be causally closed? Foundations of Physics, 34(9):1285–1303, September 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/B%
3AFOOP.0000044094.09861.12.

Whitaker:2004:EPB

[2824] M. A. B. Whitaker. The EPR paper and Bohr’s response: a re-assessment. Foundations of Physics, 34(9):1305–1340, September 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/B%
3AFOOP.0000044095.69270.31;   http://www.springerlink.com/content/
j032495h61374575/.

Falmagne:2004:MOI

[2825] Jean-Claude Falmagne. Meaningfulness and order-invariance: Two fundamental principles for scientific laws. Foundations of Physics, 34(9): 1341–1384, September 2004. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/B%3AFOOP.0000044096.50863.8e.

Bailey:2004:HLV

[2826] C. D. Bailey. Hamilton and the law of varying action revisited. Foundations of Physics, 34(9):1385–1406, September 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/B%3AFOOP.0000044097.65629.20.

Cardone:2004:KSGc

[2827] Fabio Cardone, Alessio Marrani, and Roberto Mignani. Killing symmetries of generalized Minkowski spaces, 3: Spacetime translations in four dimensions. Foundations of Physics, 34(9):1407–1429, September 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/B%
3AFOOP.0000044098.47155.38.

DiGrezia:2004:FMS

[2828] E. Di Grezia and S. Esposito. Fermi, Majorana and the statistical model of atoms. Foundations of Physics, 34(9):1431–1450, September 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1023/B:FOOP.
0000044099.18859.1d.

Anonymous:2004:FPLi

[2829] Anonymous. Foundations of Physics Letters: List of contents: Volume 17, number 6, November 2004. Foundations of Physics, 34(9): 1451, September 2004. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/B%3AFOOP.0000044208.96364.4e.

Marchildon:2004:WSW

[2830] Louis Marchildon. Why should we interpret quantum mechanics? Foundations of Physics, 34(10):1453–1466, October 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/B%3AFOOP.0000044100.95918.b2.

Jones:2004:CWC

[2831] Gordon Jones, Philip Pearle, and James Ring. Consequence for wavefunction collapse model of the Sudbury Neutrino Observatory Experiment. Foundations of Physics, 34(10):1467–1474, October 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/B%
3AFOOP.0000044101.51344.93.

Woodward:2004:FCO

[2832] James F. Woodward. Flux capacitors and the origin of inertia. Foundations of Physics, 34(10):1475–1514, October 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/B%3AFOOP.0000044102.03268.46.

Erber:2004:RP

[2833] T. Erber. Reflections on parity. Foundations of Physics, 34(10):1515– 1540, October 2004. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1023/B%3AFOOP.0000044103.06633.b8.

vonBorzeszkowski:2004:MMU

[2834] Horst-Heino von Borzeszkowski and Hans-Jürgen Treder. On metric and matter in unconnected, connected, and metrically connected manifolds. Foundations of Physics, 34(10):1541–1572, October 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/B%
3AFOOP.0000044104.36647.ef.

Lindesay:2004:NUQ

[2835] James Lindesay and H. Pierre Noyes. Nonperturbative, unitary quantumparticle scattering amplitudes from three-particle equations. Foundations of Physics, 34(10):1573–1606, October 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/B%3AFOOP.0000044105.74284.04.

Grandy:2004:BRBb

[2836] W. T. Grandy, Jr. Book review: Equilibrium and Non-Equilibrium Statistical Thermodynamics. By Michel Le Bellac, Fabrice Mortessagne and G. George Batrouni. Cambridge University Press, Cambridge, United Kingdom, 2004, xvi + 632 pp., $75 (hardcover). ISBN 0-521-82143-6. Foundations of Physics, 34(10):1607–1609, October 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/B%3AFOOP.0000044209.48220.60.

Anonymous:2004:FPLj

[2837] Anonymous. Foundations of Physics Letters: List of contents: Volume 17, number 7, December 2004. Foundations of Physics, 34(10):1611, October 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/B%
3AFOOP.0000044210.21091.f3.

Tarozzi:2004:FSH

[2838] Gino Tarozzi and Alwyn van der Merwe. For Franco Selleri on his seventieth birthday. Foundations of Physics, 34(11):1613–1615, November 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-004-1306-1.

Jammer:2004:SSC

[2839] Max Jammer. The strange story of the concept which inaugurated modern theoretical physics. Foundations of Physics, 34(11):1617– 1641, November 2004. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-004-1307-0.

Santos:2004:FPL

[2840] Emilio Santos. The failure to perform a loophole-free test of Bell’s inequality supports local realism. Foundations of Physics, 34(11):1643– 1673, November 2004. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-004-1308-z.

Auletta:2004:PRQ

[2841] Gennaro Auletta and Gino Tarozzi. On the physical reality of quantum waves. Foundations of Physics, 34(11):1675–1694, November 2004. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-004-1309-y.

deBeauregard:2004:BBPa

[2842] O. Costa de Beauregard. To believe or not believe in the A potential, That’s a question. Flux quantization in autistic magnets. Prediction of a new effect. Foundations of Physics, 34(11):1695–1702, November 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-004-1310-5.

Arminjon:2004:GAT

[2843] Mayeul Arminjon. Gravity as Archimedes’ thrust and a bifurcation in that theory. Foundations of Physics, 34(11):1703–1724, November 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-004-1312-3.

Hatch:2004:CEP

[2844] Ronald R. Hatch. Clocks and the equivalence principle. Foundations of Physics, 34(11):1725–1739, November 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-004-1313-2.

Freire:2004:HRF

[2845] Olival Freire, Jr. The historical roots of “foundations of quantum physics” as a field of research (1950–1970). Foundations of Physics, 34(11):1741–1760, November 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-004-1314-1.

Bergia:2004:WWW

[2846] Silvio Bergia. The way we were: Bubble chamber pictures, pion– nucleon interactions and polology. Foundations of Physics, 34(11):1761– 1776, November 2004. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-004-1315-0.

Wegener:2004:ICT

[2847] Mogens Wegener. The idea of a cosmic time. Foundations of Physics, 34(11):1777–1799, November 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-004-1316-z.

Adamenko:2004:EAS

[2848] S. V. Adamenko and V. I. Vysotskii. Evolution of annular self-controlled electron–nucleus collapse in condensed targets. Foundations of Physics, 34(11):1801–1831, November 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-004-1317-y.

Rizzi:2004:SGP

[2849] Guido Rizzi, Matteo Luca Ruggiero, and Alessio Serafini. Synchronization gauges and the principles of special relativity. Foundations of Physics, 34(12):1835–1887, December 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-004-1624-3.

Nimtz:2004:SSV

[2850] Günter Nimtz. Superluminal signal velocity and causality. Foundations of Physics, 34(12):1889–1903, December 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-004-1625-2.

Levy:2004:ERC

[2851] Joseph Levy. Experimental and real coordinates in space–time transformations. Foundations of Physics, 34(12):1905–1922, December 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-004-1626-1.

deBeauregard:2004:BBPb

[2852] O. Costa de Beauregard. To believe or not believe in the 4-potential, That’s a question. The electric Helmholtz–Mikhailov effect and its magnetic analog. Foundations of Physics, 34(12):1923–1928, December 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-004-1627-0.

Croca:2004:BTL

[2853] J. R. Croca. De Broglie tired light model and the reality of the quantum waves. Foundations of Physics, 34(12):1929–1954, December 2004. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-004-1628-z.

Ranada:2004:PAA

[2854] Antonio F. Rañada. The pioneer anomaly as acceleration of the clocks. Foundations of Physics, 34(12):1955–1971, December 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-004-1629-y.

Barone:2004:VEL

[2855] M. Barone. The vacuum as ether in the last century. Foundations of Physics, 34(12):1973–1982, December 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-004-1630-5.

Garuccio:2004:IPC

[2856] Augusto Garuccio. Interferometry with phase conjugate mirrors and measure of one-way velocity of light. Foundations of Physics, 34(12): 1983–1992, December 2004. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-004-1631-4.

Ferrero:2004:EVQ

[2857] M. Ferrero, D. Salgado, and J. L. Sánchez-Gómez. Is the epistemic view of quantum mechanics incomplete? Foundations of Physics, 34(12): 1993–2003, December 2004. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-004-1632-3.

Fano:2004:NMN

[2858] Vincenzo Fano. Non-materiality of non-locality. Foundations of Physics, 34(12):2005–2013, December 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-004-1633-2.

Smirnov-Rueda:2005:EIH

[2859] R. Smirnov-Rueda. On essential incompleteness of Hertz’s experiments on propagation of electromagnetic interactions. Foundations of Physics, 35(1):1–31, January 2005. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-004-1911-z.

DeBaere:2005:CRG

[2860] W. De Baere. On the consequences of retaining the general validity of locality in physical theory. Foundations of Physics, 35(1):33–56, January 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
004-1912-y.

Bitsakis:2005:STO

[2861] Eftichios Bitsakis. Space and time: The ongoing quest. Foundations of Physics, 35(1):57–83, January 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-004-1913-x.

Oliynyk:2005:HMH

[2862] Todd A. Oliynyk. Hidden measurements, hidden variables and the volume representation of transition probabilities. Foundations of Physics, 35(1):85–107, January 2005. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-004-1921-x.

Vianna:2005:GCC

[2863] J. D. M. Vianna, M. C. B. Fernandes, and A. E. Santana. Galileancovariant Clifford algebras in the phase-space representation. Foundations of Physics, 35(1):109–129, January 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-004-1926-5.

Carroll:2005:IQM

[2864] Robert Carroll. Information, quantum mechanics, and gravity. Foundations of Physics, 35(1):131–154, January 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-004-1928-3.

Auletta:2005:QII

[2865] Gennaro Auletta. Quantum information and inferential reasoning. Foundations of Physics, 35(1):155–156, January 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-004-1929-2.

Anonymous:2005:FPLa

[2866] Anonymous. Foundations of Physics Letters. Foundations of Physics, 35 (1):171–??, January 2005. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-004-1930-9.

Dickson:2005:TJC

[2867] Michael Dickson and Antony Valentini. A tribute to James T. Cushing: 1937–2002. Foundations of Physics, 35(2):173–176, February 2005. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-004-1939-0.

Holland:2005:WWE

[2868] Peter Holland. What’s wrong with Einstein’s 1927 hidden-variable interpretation of quantum mechanics? Foundations of Physics, 35(2): 177–196, February 2005. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-004-1940-7.

Schlosshauer:2005:ZDB

[2869] Maximilian Schlosshauer and Arthur Fine. On Zurek’s derivation of the Born rule. Foundations of Physics, 35(2):197–213, February 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-004-1941-6.

Shimony:2005:AEB

[2870] Abner Shimony. An analysis of ensembles that are both pre- and postselected. Foundations of Physics, 35(2):215–232, February 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-004-1942-5.

Butterfield:2005:PP

[2871] J. Butterfield. On the persistence of particles. Foundations of Physics, 35 (2):233–269, February 2005. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-004-1943-4.

Isham:2005:QC

[2872] C. J. Isham. Quantising on a category. Foundations of Physics, 35 (2):271–297, February 2005. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-004-1944-3.

Vaidman:2005:RBQ

[2873] Lev Vaidman. The reality in Bohmian quantum mechanics or can you kill with an empty wave bullet? Foundations of Physics, 35(2): 299–312, February 2005. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-004-1945-2.

Timpson:2005:NIF

[2874] C. G. Timpson. Nonlocality and information flow: The approach of Deutsch and Hayden. Foundations of Physics, 35(2):313–343, February 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
004-1946-1.

Anonymous:2005:FPLb

[2875] Anonymous. Foundations of Physics Letters. Foundations of Physics, 35(2):345, February 2005. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-004-1947-0.

Bokulich:2005:NBG

[2876] Peter Bokulich and Alisa Bokulich. Niels Bohr’s generalization of classical mechanics. Foundations of Physics, 35(3):347–371, March 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-004-1979-5.

Berkovitz:2005:MIQ

[2877] Joseph Berkovitz and Meir Hemmo. Modal interpretations of quantum mechanics and relativity: a reconsideration. Foundations of Physics, 35 (3):373–397, March 2005. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-004-1980-z.

Dieks:2005:QMI

[2878] Dennis Dieks. Quantum mechanics: an intelligible description of objective reality? Foundations of Physics, 35(3):399–415, March 2005. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-004-1981-y.

Saunders:2005:CSR

[2879] Simon Saunders. Complementarity and scientific rationality. Foundations of Physics, 35(3):417–447, March 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-004-1982-x.

Durr:2005:RDM

[2880] Detlef Dürr, Sheldon Goldstein, Roderich Tumulka, and Nino Zanghí. On the role of density matrices in Bohmian mechanics. Foundations of Physics, 35(3):449–467, March 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-004-1983-9.

Leavens:2005:WMP

[2881] C. R. Leavens. Weak measurements from the point of view of Bohmian mechanics. Foundations of Physics, 35(3):469–491, March 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-004-1984-8.

Maroney:2005:DMB

[2882] O. J. E. Maroney. The density matrix in the de Broglie–Bohm approach. Foundations of Physics, 35(3):493–510, March 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-004-1985-7.

Peres:2005:EPR

[2883] A. Peres. Einstein, Podolsky, Rosen, and Shannon. Foundations of Physics, 35(3):511–514, March 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-004-1986-6.

Brown:2005:SMP

[2884] Harvey R. Brown and David Wallace. Solving the measurement problem: de Broglie–Bohm loses out to Everett. Foundations of Physics, 35(4):517–540, April 2005. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-004-2009-3.

Bub:2005:QMA

[2885] Jeffrey Bub. Quantum mechanics is about quantum information. Foundations of Physics, 35(4):541–560, April 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-004-2010-x.

Redhead:2005:BBR

[2886] Michael Redhead. Broken bootstraps — the rise and fall of a research programme. Foundations of Physics, 35(4):561–575, April 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-004-2011-9.

Pinto-Neto:2005:BIQ

[2887] Nelson Pinto-Neto. The Bohm interpretation of quantum cosmology. Foundations of Physics, 35(4):577–603, April 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-004-2012-8.

Bowman:2005:CLB

[2888] Gary E. Bowman. On the classical limit in Bohm’s theory. Foundations of Physics, 35(4):605–625, April 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-004-2013-7.

Appleby:2005:FVQ

[2889] D. M. Appleby. Facts, values and Quanta. Foundations of Physics, 35(4):627–668, April 2005. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-004-2014-6.

Belousek:2005:URT

[2890] Darrin W. Belousek. Underdetermination, realism, and theory appraisal: an epistemological reflection on quantum mechanics. Foundations of Physics, 35(4):669–695, April 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-004-2015-5.

Anonymous:2005:LCV

[2891] Anonymous. List of contents: Volume 18, number 2, April 2005. Foundations of Physics, 35(4):697, April 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-004-2016-4.

Lindesay:2005:CNP

[2892] James Lindesay and H. Pierre. Noyes. Construction of non-perturbative, unitary particle–antiparticle amplitudes for finite particle number scattering formalisms. Foundations of Physics, 35(5):699–741, May 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-005-4563-8.

Collins:2005:MBP

[2893] R. Eugene Collins. The mathematical basis for physical laws. Foundations of Physics, 35(5):743–785, May 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-005-4564-7.

Auletta:2005:QIG

[2894] Gennaro Auletta. Quantum information as a general paradigm. Foundations of Physics, 35(5):787–815, May 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-005-4565-6.

Cirkovic:2005:PVS

[2895] Milan M. Ćirković. Physics versus semantics: a puzzling case of the missing quantum theory. Foundations of Physics, 35(5):817–838, May 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-005-4566-5.

Broekaert:2005:MLT

[2896] Jan Broekaert. A modified Lorentz-transformation-based gravity model confirming basic GRT experiments. Foundations of Physics, 35(5): 839–864, May 2005. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-005-4567-4.

Gsponer:2005:CLD

[2897] Andre Gsponer and Jean-Pierre Hurni. Cornelius Lanczos’s derivation of the usual action integral of classical electrodynamics. Foundations of Physics, 35(5):865–880, May 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-005-4568-3.

Abernethy:2005:DBC

[2898] Deanna Abernethy and John R. Klauder. The distance between classical and quantum systems. Foundations of Physics, 35(5):881–895, May 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-005-4569-2.

Klauder:2005:BRB

[2899] John R. Klauder. Book review: Time, Quantum and Information. First Edition, Corrected 2nd Printing. By Lutz Castell and Otfried Ischebeck, Springer, Berlin, Germany, 2004, xiii + 456 pp., $ 69.95. Foundations of Physics, 35(5):897–899, May 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-005-4570-9.

Hestenes:2005:GTG

[2900] David Hestenes. Gauge theory gravity with geometric calculus. Foundations of Physics, 35(6):903–970, June 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-005-5828-y.

Castro:2005:ERT

[2901] Carlos Castro. The extended relativity theory in Born–Clifford phase spaces with a lower and upper length scales and Clifford group geometric unification. Foundations of Physics, 35(6):971–1041, June 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-005-5829-x.

Schust:2005:SEA

[2902] P. Schust, F. Stary, M. Mattes, and M. Sorg. Self-energy and action principle in relativistic Schrödinger theory. Foundations of Physics, 35 (6):1043–1105, June 2005. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-005-5830-4.

Lakshminarayanan:2005:BRB

[2903] V. Lakshminarayanan. Book review: Quantum Mechanics: Theory and Applications. By Ajoy K. Ghatak and S. Lokanathan, Kluwer Academic Publishers, Dordrecht, The Netherlands, 2004, xlviii + 858 pp., $303.00/$99.00 (hardcover/softcover). ISBN 1-4020-1850-9. Foundations of Physics, 35(6):1107–1109, June 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-005-5831-3.

Horwitz:2005:P

[2904] L. P. Horwitz, J. R. Fanchi, and M. Land. Preface. Foundations of Physics, 35(7):1113–1115, July 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-005-6403-2.

Johnson:2005:UAC

[2905] Philip R. Johnson and B. L. Hu. Uniformly accelerated charge in a quantum field: From radiation reaction to Unruh effect. Foundations of Physics, 35(7):1117–1147, July 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-005-6404-1.

Higuchi:2005:CQR

[2906] Atsushi Higuchi and Giles D. R. Martin. Classical and quantum radiation reaction for linear acceleration. Foundations of Physics, 35(7):1149–1179, July 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
005-6405-0.

Oron:2005:RBM

[2907] O. Oron and L. P. Horwitz. Relativistic Brownian motion and gravity as an eikonal approximation to a quantum evolution equation. Foundations of Physics, 35(7):1181–1203, July 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-005-6406-z.

Rapoport:2005:UGR

[2908] Diego L. Rapoport. On the unification of geometric and random structures through torsion fields: Brownian motions, viscous and magnetofluid-dynamics. Foundations of Physics, 35(7):1205–1244, July 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-005-6407-y.

Land:2005:DSE

[2909] Martin Land. Duality in off–shell electromagnetism. Foundations of Physics, 35(7):1245–1262, July 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-005-6408-x.

Land:2005:DSS

[2910] Martin Land. Discrete symmetries of off–shell electromagnetism. Foundations of Physics, 35(7):1263–1288, July 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-005-6409-9.

Kim:2005:SWL

[2911] Y. S. Kim and Marilyn E. Noz. Standing waves in the Lorentz-covariant world. Foundations of Physics, 35(7):1289–1305, July 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-005-6410-3.

Shovkovy:2005:TLC

[2912] Igor A. Shovkovy. Two lectures on color superconductivity*. Foundations of Physics, 35(8):1309–1358, August 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-005-6440-x.

Schieve:2005:CRS

[2913] Wm. C. Schieve. Covariant relativistic statistical mechanics of many particles. Foundations of Physics, 35(8):1359–1381, August 2005. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-005-6441-9.

Rapoport:2005:CWD

[2914] Diego L. Rapoport. Cartan–Weyl Dirac and Laplacian operators, Brownian motions: The quantum potential and scalar curvature, Maxwell’s and Dirac–Hestenes equations, and supersymmetric systems. Foundations of Physics, 35(8):1383–1431, August 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-005-6443-7.

Gersten:2005:FAC

[2915] A. Gersten. Filed approach to classical mechanics. Foundations of Physics, 35(8):1433–1443, August 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-005-6446-4.

Gersten:2005:ETW

[2916] Alexander Gersten. Experiment to test whether we live in a fourdimensional physical space–time. Foundations of Physics, 35(8):1445– 1452, August 2005. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-005-6447-3.

Fanchi:2005:IRQ

[2917] John R. Fanchi. Introducing relativistic quantum mechanics to energy students. Foundations of Physics, 35(8):1453–1474, August 2005. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-005-6449-1.

Anderson:2005:ELC

[2918] Matthew Anderson and Richard A. Matzner. Extended lifetime in computational evolution of isolated black holes. Foundations of Physics, 35(9):1477–1495, September 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-005-6477-x.

Cooperstock:2005:CTC

[2919] F. I. Cooperstock and S. Tieu. Closed timelike curves and time travel: Dispelling the myth. Foundations of Physics, 35(9):1497–1509, September 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
005-6478-9.

Culetu:2005:HP

[2920] Hristu Culetu. Λ and the Heisenberg principle. Foundations of Physics, 35(9):1511–1519, September 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-005-6479-8.

Harpaz:2005:CO

[2921] Amos Harpaz and Noam Soker. A coaccelerated observer. Foundations of Physics, 35(9):1521–1531, September 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-005-6480-2.

Ben-Amots:2005:DRL

[2922] N. Ben-Amots. Detection of rotation by a local elevator-like Gedanken experiment. Foundations of Physics, 35(9):1533–1542, September 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-005-6481-1.

Montanus:2005:FSG

[2923] J. M. C. Montanus. Flat space gravitation. Foundations of Physics, 35(9): 1543–1562, September 2005. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-005-6482-0.

OHara:2005:QMM

[2924] Paul O’Hara. Quantum mechanics and the metrics of general relativity. Foundations of Physics, 35(9):1563–1584, September 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-005-6483-z.

Ivezic:2005:PME

[2925] Tomislav Ivezić. The proof that Maxwell equations with the 3D E and B are not covariant upon the Lorentz transformations but upon the standard transformations: The new Lorentz invariant field equations. Foundations of Physics, 35(9):1585–1615, September 2005. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-005-6484-y.

Pavsic:2005:CSG

[2926] Matej Pavsic. Clifford space as a generalization of spacetime: Prospects for QFT of point particles and strings. Foundations of Physics, 35(9): 1617–1642, September 2005. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-005-6485-x.

Piron:2005:NEG

[2927] C. Piron. New Einstein gravitation*. Foundations of Physics, 35(9): 1643–1647, September 2005. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.

1007/s10701-005-6486-9. See clarification [3070].

Piron:2005:C

[2928] C. Piron. Complement. Foundations of Physics, 35(9):1649–1651, September 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-005-6487-8.

Khrennikov:2005:PSC

[2929] Andrei Khrennikov. The principle of supplementarity: a contextual probabilistic viewpoint to complementarity, the interference of probabilities and incompatibility of variables in quantum mechanics. Foundations of Physics, 35(10):1655–1693, October 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-005-6511-z.

Smirnov-Rueda:2005:TCT

[2930] R. Smirnov-Rueda. On two complementary types of total time derivative in classical field theories and Maxwell’s equations. Foundations of Physics, 35(10):1695–1723, October 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-005-6515-8.

Israelit:2005:WIM

[2931] Mark Israelit. Wesson’s induced matter theory with a Weylian bulk. Foundations of Physics, 35(10):1725–1748, October 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-005-6518-5.

Hess:2005:BTS

[2932] Karl Hess and Walter Philipp. The Bell theorem as a special case of a theorem of Bass. Foundations of Physics, 35(10):1749–1767, October 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-005-6520-y.

Israelit:2005:MSW

[2933] Mark Israelit. On measuring standards in Weyl’s geometry. Foundations of Physics, 35(10):1769–1782, October 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-005-6523-8.

Fuchs:2005:P

[2934] Christopher Fuchs. Preface. Foundations of Physics, 35(11):1785– 1786, November 2005. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-005-7348-1.

Bechmann-Pasquinucci:2005:QST

[2935] H. Bechmann-Pasquinucci. From quantum state targeting to Bell inequalities. Foundations of Physics, 35(11):1787–1804, November 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-005-7349-0.

Bekenstein:2005:HDE

[2936] Jacob D. Bekenstein. How does the Entropy/ information bound work? Foundations of Physics, 35(11):1805–1823, November 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-005-7350-7.

Benioff:2005:TCT

[2937] Paul Benioff. Towards a coherent theory of physics and mathematics: The theory–experiment connection. Foundations of Physics, 35(11): 1825–1856, November 2005. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-005-7351-6.

Blume-Kohout:2005:SEQ

[2938] Robin Blume-Kohout and Wojciech H. Zurek. A simple example of “Quantum Darwinism”: Redundant information storage in many-spin environments. Foundations of Physics, 35(11):1857–1876, November 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
005-7352-5.

Brassard:2005:QPT

[2939] Gilles Brassard, Anne Broadbent, and Alain Tapp. Quantum pseudotelepathy. Foundations of Physics, 35(11):1877–1907, November 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-005-7353-4.

Brukner:2005:CID

[2940] Caslav Brukner, Markus Aspelmeyer, and Anton Zeilinger. Complementarity and information in “delayed-choice for entanglement swapping”. Foundations of Physics, 35(11):1909–1919, November 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-005-7355-2.

Bruss:2005:HFP

[2941] Dagmar Bruß and Chiara Macchiavello. How the first partial transpose was written. Foundations of Physics, 35(11):1921–1926, November 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-005-7357-0.

Cabello:2005:BTI

[2942] Adán Cabello. Bell’s theorem without inequalities and without unspeakable information. Foundations of Physics, 35(11):1927–1934, November 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
005-7358-z.

Deser:2005:NST

[2943] S. Deser. A note on stress-tensors, conservation and equations of motion. Foundations of Physics, 35(11):1935–1940, November 2005. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-005-7360-5.

dEspagnat:2005:CWF

[2944] Bernard d’Espagnat. Consciousness and the Wigner’s friend problem. Foundations of Physics, 35(12):1943–1966, December 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-005-8656-1.

DiVincenzo:2005:FLO

[2945] David P. DiVincenzo and Barbara M. Terhal. Fermionic linear optics revisited. Foundations of Physics, 35(12):1967–1984, December 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-005-8657-0.

Flammia:2005:MIC

[2946] Steven T. Flammia, Andrew Silberfarb, and Carlton M. Caves. Minimal informationally complete measurements for pure states. Foundations of Physics, 35(12):1985–2006, December 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-005-8658-z.

Hehl:2005:CEF

[2947] Friedrich W. Hehl and Yuri N. Obukhov. To consider the electromagnetic field as fundamental, and the metric only as a subsidiary field. Foundations of Physics, 35(12):2007–2025, December 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-005-8659-y.

Horodecki:2005:ITA

[2948] Karol Horodecki, Michal Horodecki, Pawel Horodecki, and Jonathan Oppenheim. Information theories with adversaries, intrinsic information, and entanglement. Foundations of Physics, 35(12):2027–2040, December 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
005-8660-5.

Horodecki:2005:CON

[2949] Michal Horodecki, Ryszard Horodecki, Aditi Sen(De), and Ujjwal Sen. Common origin of no-cloning and no-deleting principles conservation of information. Foundations of Physics, 35(12):2041–2049, December 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-005-8661-4.

Loubenets:2005:LRB

[2950] Elena R. Loubenets. “Local realism”, Bell’s theorem and quantum “locally realistic” inequalities. Foundations of Physics, 35(12):2051– 2072, December 2005. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-005-8662-3.

Mermin:2005:WWC

[2951] N. David Mermin. What’s wrong with this criticism. Foundations of Physics, 35(12):2073–2077, December 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-005-8663-2.

Peres:2006:CWW

[2952] Asher Peres. I am the cat who walks by himself. Foundations of Physics, 36(1):1–18, January 2006. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-005-9010-3.

Peshkin:2006:SZP

[2953] Murray Peshkin. Spin-zero particles must be bosons: a new proof within nonrelativistic quantum mechanics. Foundations of Physics, 36 (1):19–29, January 2006. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-005-9011-2.

Plotnitsky:2006:NBN

[2954] Arkady Plotnitsky. A new book of numbers: On the precise definition of quantum variables and the relationships between mathematics and physics in quantum theory. Foundations of Physics, 36(1):30–60, January 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
005-9004-1.

Shimony:2006:ASB

[2955] Abner Shimony. An analysis of Stapp’s “A Bell-Type Theorem without Hidden Variables”. Foundations of Physics, 36(1):61–72, January 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-

005-9002-3. See [2956].

Stapp:2006:CSA

[2956] Henry P. Stapp. Comments on Shimony’s “An Analysis of Stapp’s ‘A Bell-Type Theorem without Hidden Variables’ ”. Foundations of Physics, 36(1):73–82, January 2006. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-005-9005-0. See [2955].

Butterley:2006:CSS

[2957] Paul Butterley, Anthony Sudbery, and Jason Szulc. Compatibility of subsystem states. Foundations of Physics, 36(1):83–101, January 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-005-9006-z.

Terno:2006:IQC

[2958] Daniel R. Terno. Inconsistency of quantum–classical dynamics, and what it implies. Foundations of Physics, 36(1):102–111, January 2006. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-005-9007-y.

Wootters:2006:QMF

[2959] W. K. Wootters. Quantum measurements and finite geometry. Foundations of Physics, 36(1):112–126, January 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-005-9008-x.

Ziman:2006:RSI

[2960] Mário Ziman, Martin Plesch, and Vladimír Buz zek. Reconstruction of superoperators from incomplete measurements. Foundations of Physics, 36(1):127–156, January 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-005-9009-9.

Anonymous:2006:APL

[2961] Anonymous. Asher Peres — list of publications. Foundations of Physics, 36(1):157–173, January 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-005-9003-2.

Kryukov:2006:QMH

[2962] Alexey A. Kryukov. Quantum mechanics on Hilbert manifolds: The principle of functional relativity. Foundations of Physics, 36(2):175– 226, February 2006. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-005-9012-1.

Ellis:2006:PRW

[2963] George F. R. Ellis. Physics and the real world. Foundations of Physics, 36 (2):227–262, February 2006. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-005-9016-x.

Cardone:2006:NER

[2964] Fabio Cardone, Roberto Mignani, and Renato Scrimaglio. New experimental results on the lower limits of local Lorentz invariance. Foundations of Physics, 36(2):263–290, February 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-005-9014-z.

beimGraben:2006:CCD

[2965] Peter beim Graben and Harald Atmanspacher. Complementarity in classical dynamical systems. Foundations of Physics, 36(2):291–306, February 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
005-9013-0.

Faria:2006:TCE

[2966] A. J. Faria, H. M. França, and R. C. Sponchiado. Tunneling as a classical escape rate induced by the vacuum zero-point radiation. Foundations of Physics, 36(2):307–320, February 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-005-9017-9.

Woodward:2006:BRB

[2967] James F. Woodward. Book review: World Without Time: The Forgotten Legacy of Gödel and Einstein. By Palle Yourgrau, Basic Books, New York, New York, USA, 2005, viii + 210 pp., $24 (hard cover). ISBN 0-465-09293-4. Foundations of Physics, 36(2):321–325, February 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-005-9018-8.

Ferrero:2006:FEE

[2968] Migual Ferrero and Alwyn Van der Merwe. Faith, endeavour and expectations of an unfinished quest: a tribute to Emilio Santos on his seventieth birthday. Foundations of Physics, 36(3):327–332, March 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-005-9024-x.

Marshall:2006:AWP

[2969] Trevor W. Marshall. Are atoms waves or particles? Foundations of Physics, 36(3):333–349, March 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-005-9023-y.

delaPena:2006:FLS

[2970] L. de la Peña and A. M. Cetto. The foundations of linear stochastic electrodynamics. Foundations of Physics, 36(3):350–368, March 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-005-9020-1.

Holland:2006:HVC

[2971] Peter Holland. Hidden variables as computational tools: The construction of a relativistic spinor field. Foundations of Physics, 36(3):369–384, March 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
005-9022-z.

Boya:2006:UW

[2972] Luis J. Boya. The uniqueness of the world. Foundations of Physics, 36 (3):385–395, March 2006. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-005-9019-7.

Rosales:2006:PAM

[2973] J. L. Rosales. The pioneer anomaly: The measure of a topological phase defect of light in cosmology. Foundations of Physics, 36(3):396–406, March 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
005-9027-7.

Castagnino:2006:TRI

[2974] M. Castagnino, M. Gadella, and O. Lombardi. Time-reversal, irreversibility and arrow of time in quantum mechanics. Foundations of Physics, 36(3):407–426, March 2006. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-005-9021-0.

Ranada:2006:TQM

[2975] Antonio F. Rañada and José Luis Trueba. Topological quantization of the magnetic flux. Foundations of Physics, 36(3):427–436, March 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-005-9026-8.

Pykacz:2006:SEP

[2976] Jarosław Pykacz. “Solution” of the EPR Paradox: Negative, or rather fuzzy probabilities? Foundations of Physics, 36(3):437–442, March 2006.

CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-005-9025-9.

Selleri:2006:SSR

[2977] F. Selleri. Superluminal signals and the resolution of the causal paradox. Foundations of Physics, 36(3):443–463, March 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-005-9028-6.

Porras:2006:PSS

[2978] D. Porras and J. I. Cirac. Phonon superfluids in sets of trapped ions. Foundations of Physics, 36(4):465–476, April 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-005-9036-6.

Hulpke:2006:UPE

[2979] Florian Hulpke, Uffe V. Poulsen, Anna Sanpera, Aditi Sen(De), Ujjwal Sen, and Maciej Lewenstein. Unitarity as preservation of entropy and entanglement in quantum systems. Foundations of Physics, 36(4):477–499, April 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
005-9035-7.

Facchi:2006:ZSC

[2980] Paolo Facchi, Rosario Fazio, Giuseppe Florio, Saverio Pascazio, and Tetsuya Yoneda. Zeno subspaces for coupled superconducting qubits. Foundations of Physics, 36(4):500–511, April 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-005-9033-9.

Cabello:2006:CCP

[2981] Adán Cabello. Communication complexity as a principle of quantum mechanics. Foundations of Physics, 36(4):512–525, April 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-005-9029-5.

Salgado:2006:MCQ

[2982] D. Salgado, J. L. Sánchez-Gómez, and M. Ferrero. A mathematical characterization of quantum Gaussian stochastic evolution schemes. Foundations of Physics, 36(4):526–540, April 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-005-9034-8.

Zukowski:2006:SQS

[2983] Marek Zukowski. Separability of quantum states vs. original Bell (1964) inequalities. Foundations of Physics, 36(4):541–545, April 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-005-9038-4.

Revzen:2006:WFD

[2984] M. Revzen. The Wigner function as distribution function. Foundations of Physics, 36(4):546–562, April 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-005-9037-5.

Bramon:2006:BIT

[2985] A. Bramon, R. Escribano, and G. Garbarino. Bell’s inequality tests with meson–antimeson pairs. Foundations of Physics, 36(4):563–584, April 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
005-9030-z.

Stairs:2006:LRC

[2986] Allen Stairs and Jeffrey Bub. Local realism and conditional probability. Foundations of Physics, 36(4):585–601, April 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-005-9031-y.

vanderMerwe:2006:PES

[2987] Alwyn van der Merwe. Publications of Emilio Santos. Foundations of Physics, 36(4):602–612, April 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-005-9032-x.

Nobbenhuis:2006:CDA

[2988] Stefan Nobbenhuis. Categorizing different approaches to the cosmological constant problem. Foundations of Physics, 36(5):613–680, May 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-005-9042-8.

Madarasz:2006:TPL

[2989] Judit X. Madarász, István Németi, and Gergely Székely. Twin Paradox and the logical foundation of relativity theory. Foundations of Physics, 36(5):681–714, May 2006. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-005-9041-9.

Kholmetskii:2006:GRC

[2990] Alexander L. Kholmetskii. On “gauge renormalization” in classical electrodynamics. Foundations of Physics, 36(5):715–744, May 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-005-9039-3.

Hofer-Szabo:2006:RCC

[2991] Gábor Hofer-Szabó and Miklós Rédei. Reichenbachian common cause systems of arbitrary finite size exist. Foundations of Physics, 36(5): 745–756, May 2006. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-005-9040-x.

Treder:2006:CQP

[2992] H.-J. Treder and H.-H. von Borzeszkowski. Covariance and quantum principles–censors of the space–time structure. Foundations of Physics, 36(5):757–763, May 2006. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-005-9043-7.

Hut:2006:MMM

[2993] Piet Hut, Mark Alford, and Max Tegmark. On math, matter and mind. Foundations of Physics, 36(6):765–794, June 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-006-9048-x.

Manousakis:2006:FQT

[2994] Efstratios Manousakis. Founding quantum theory on the basis of consciousness. Foundations of Physics, 36(6):795–838, June 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-006-9049-9.

Hartnett:2006:DMD

[2995] John G. Hartnett. The distance modulus determined from Carmeli’s cosmology fits the accelerating Universe data of the high-redshift Type Ia supernovae without dark matter. Foundations of Physics, 36(6):839–861, June 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
006-9047-y.

Garola:2006:QMS

[2996] Claudio Garola, Jarosław Pykacz, and Sandro Sozzo. Quantum machine and semantic realism approach: a unified model. Foundations of Physics, 36(6):862–882, June 2006. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-006-9046-z.

Schroder:2006:AME

[2997] W. Schröder and H.-J. Treder. Aspects of the Mach–Einstein doctrine and geophysical application (a historical review). Foundations of Physics, 36(6):883–901, June 2006. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-006-9050-3.

Eringen:2006:MET

[2998] A. Cemal Eringen. Micromorphic electromagnetic theory and waves. Foundations of Physics, 36(6):902–919, June 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-006-9044-1.

Castagnino:2006:PCL

[2999] Mario Castagnino and Manuel Gadella. The problem of the classical limit of quantum mechanics and the role of self-induced decoherence. Foundations of Physics, 36(6):920–952, June 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-006-9074-8.

Agashe:2006:EEL

[3000] S. D. Agashe. Einstein’s “Zur Elektrodynamik … ” (1905) revisited, with some consequences. Foundations of Physics, 36(7):955–1011, July 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-

006-9055-y. See addendum [3049].

Myers:2006:CPD

[3001] John M. Myers. Conditional probabilities and density operators in quantum modeling. Foundations of Physics, 36(7):1012–1035, July 2006. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-006-9053-0.

Low:2006:RRN

[3002] Stephen G. Low. Reciprocal relativity of noninertial frames and the quaplectic group. Foundations of Physics, 36(7):1036–1069, July 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-006-9051-2.

Krol:2006:MSR

[3003] Jerzy Król. A model for spacetime: The role of interpretation in some Grothendieck topoi. Foundations of Physics, 36(7):1070–1098, July 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-006-9052-1.

Raju:2006:TTR

[3004] C. K. Raju. Time travel and the reality of spontaneity. Foundations of Physics, 36(7):1099–1113, July 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-006-9056-x.

Rubin:2006:SDF

[3005] Mark A. Rubin. Spatial degrees of freedom in Everett quantum mechanics. Foundations of Physics, 36(8):1115–1159, August 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-006-9062-z.

Gudder:2006:QMF

[3006] Stan Gudder. Quantum mechanics on finite groups. Foundations of Physics, 36(8):1160–1192, August 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-006-9060-1.

Meschini:2006:ESP

[3007] Diego Meschini and Markku Lehto. Is empty spacetime a physical thing? Foundations of Physics, 36(8):1193–1216, August 2006. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-006-9058-8.

Long:2006:DMQ

[3008] Gui Lu Long, Yi-Fan Zhou, Jia-Qi Jin, Yang Sun, and Hai-Woong Lee. Density matrix in quantum mechanics and distinctness of ensembles having the same compressed density matrix. Foundations of Physics, 36(8): 1217–1243, August 2006. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-006-9057-9.

Kopeikin:2006:AFS

[3009] Sergei M. Kopeikin and Edward B. Fomalont. Aberration and the fundamental speed of gravity in the Jovian deflection experiment. Foundations of Physics, 36(8):1244–1285, August 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-006-9059-7.

Plotnitsky:2006:BRB

[3010] Arkady Plotnitsky. Book review: It’s About Time: Understanding Einstein’s Relativity. By N. David Mermin, Princeton University Press, Princeton, NJ, USA, 2005, xv, 192 pp. (Hardcover). Foundations of Physics, 36(8):1286–1290, August 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-006-9063-y.

Sidharth:2006:BRB

[3011] B. G. Sidharth. Book review: A Different Universe. By R. B. Laughlin, Basic Books, New York, 2005, xviii, 254 pp. (Hardcover). Foundations of Physics, 36(8):1291–1294, August 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-006-9061-0.

Hagar:2006:EUW

[3012] Amit Hagar and Meir Hemmo. Explaining the unobserved — why quantum mechanics ain’t only about information. Foundations of Physics, 36(9):1295–1324, September 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-006-9065-9.

Stary:2006:OPH

[3013] F. Stary and M. Sorg. Ortho- and para-helium in relativistic Schrödinger theory. Foundations of Physics, 36(9):1325–1403, September 2006. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-006-9066-8.

Blume:2006:NPA

[3014] Frank Blume. A nontemporal probabilistic approach to special and general relativity. Foundations of Physics, 36(9):1404–1440, September 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-006-9064-x.

Conway:2006:FWT

[3015] John Conway and Simon Kochen. The Free Will Theorem. Foundations of Physics, 36(10):1441–1473, October 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-006-9068-6. See comments [3043, 3112,

deParga:2006:PDE

[3016] G. Ares de Parga. A physical deduction of an equivalent Landau–Lifshitz equation of motion in classical electrodynamics. A new expression for the large distance radiation rate of energy. Foundations of Physics, 36(10): 1474–1510, October 2006. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-006-9072-x.

Ivezic:2006:FDG

[3017] Tomislav Ivezić. Four-dimensional geometric quantities versus the usual three-dimensional quantities: The resolution of Jackson’s Paradox. Foundations of Physics, 36(10):1511–1534, October 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-006-9071-y.

Greenberg:2006:WF

[3018] O. W. Greenberg. Why is CPT fundamental? Foundations of Physics, 36(10):1535–1553, October 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-006-9070-z.

Pucci:2006:MAM

[3019] R. Pucci and G. G. N. Angilella. Majorana: From atomic and molecular, to nuclear physics. Foundations of Physics, 36(10):1554–1572, October 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
006-9067-7.

Seevinck:2006:QWB

[3020] Michael Seevinck. The quantum world is not built up from correlations. Foundations of Physics, 36(10):1573–1586, October 2006. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-006-9073-9.

Cornejo-Perez:2006:TWS

[3021] O. Cornejo-Pérez, J. Negro, L. M. Nieto, and H. C. Rosu. Travelingwave solutions for Korteweg–de Vries–Burgers equations through factorizations. Foundations of Physics, 36(10):1587–1599, October 2006. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-006-9069-5.

Anastopoulos:2006:CVQ

[3022] Charis Anastopoulos. Classical versus quantum probability in sequential measurements. Foundations of Physics, 36(11):1601–1661, November 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
006-9077-5.

Planat:2006:SFA

[3023] Michel Planat, Haret C. Rosu, and Serge Perrine. A survey of finite algebraic geometrical structures underlying mutually unbiased quantum measurements. Foundations of Physics, 36(11):1662–1680, November 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-006-9079-3.

Kato:2006:DSI

[3024] Goro Kato and Tsunefumi Tanaka. Double-slit interference and temporal topos. Foundations of Physics, 36(11):1681–1700, November 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-006-9078-4.

Christov:2006:MIF

[3025] Christo I. Christov. On the material invariant formulation of Maxwell’s displacement current. Foundations of Physics, 36(11):1701–1717, November 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
006-9075-7.

Soler:2006:RFR

[3026] D. Soler. Reference frames and rigid motions in relativity: Applications. Foundations of Physics, 36(11):1718–1735, November 2006. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-006-9080-x.

Aldrovandi:2006:ASF

[3027] R. Aldrovandi, R. R. Cuzinatto, and L. G. Medeiros. Analytic solutions for the Λ-FRW model. Foundations of Physics, 36(11):1736– 1752, November 2006. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-006-9076-6.

Bishop:2006:CED

[3028] Robert C. Bishop and Harald Atmanspacher. Contextual emergence in the description of properties. Foundations of Physics, 36(12):1753– 1777, December 2006. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-006-9082-8.

Krol:2006:MSI

[3029] Jerzy Król. A model for spacetime II. The emergence of higher dimensions and field Theory/ strings dualities. Foundations of Physics, 36(12): 1778–1800, December 2006. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-006-9087-3.

Rothman:2006:CGD

[3030] Tony Rothman and Stephen Boughn. Can gravitons be detected? Foundations of Physics, 36(12):1801–1825, December 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-006-9081-9.

Guerra:2006:CBA

[3031] Vasco Guerra and Rodrigo de Abreu. On the consistency between the assumption of a special system of reference and special relativity. Foundations of Physics, 36(12):1826–1845, December 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-006-9085-5.

Guendelman:2006:UAE

[3032] E. Guendelman and H. Ruchvarger. The Universe accelerated expansion using extra-dimensions with metric components found by a new equivalence principle. Foundations of Physics, 36(12):1846–1868, December 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
006-9084-6.

Hiley:2006:WEQ

[3033] B. J. Hiley and R. E. Callaghan. What is erased in the quantum erasure? Foundations of Physics, 36(12):1869–1883, December 2006. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-006-9086-4.

Gu:2006:ETE

[3034] Xiao-Yan Gu. Energy for two-electron quantum dots: The quantization rule approach. Foundations of Physics, 36(12):1884–1892, December 2006. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
006-9083-7.

tHooft:2007:Ea

[3035] Gerardus ’t Hooft. Editorial. Foundations of Physics, 37(1):1–2, January 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10701-006-9096-2.pdf.

Kryukov:2007:MPT

[3036] Alexey A. Kryukov. On the measurement problem for a two-level quantum system. Foundations of Physics, 37(1):3–39, January 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-006-9093-5.

Pavsic:2007:RPS

[3037] Matej Pavsic. Rigid particle and its spin revisited. Foundations of Physics, 37(1):40–79, January 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-006-9094-4.

Hnilo:2007:LDD

[3038] Alejandro A. Hnilo, Marcelo G. Kovalsky, and Guillermo Santiago. Low dimension dynamics in the EPRB experiment with random variable analyzers. Foundations of Physics, 37(1):80–102, January 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-006-9091-7.

Fulp:2007:BEG

[3039] Ronald Fulp. BRST extension of geometric quantization. Foundations of Physics, 37(1):103–124, January 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-006-9090-8.

Kalev:2007:LHV

[3040] A. Kalev, A. Mann, and M. Revzen. Local hidden variables underpinning of entanglement and teleportation. Foundations of Physics, 37(1): 125–143, January 2007. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-006-9092-6.

Ainsworth:2007:UE

[3041] Peter M. Ainsworth. Untangling entanglement. Foundations of Physics, 37(1):144–158, January 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-006-9095-3.

Wharton:2007:TSQ

[3042] K. B. Wharton. Time-symmetric quantum mechanics. Foundations of Physics, 37(1):159–168, January 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-006-9089-1.

Bassi:2007:CKA

[3043] Angelo Bassi and GianCarlo Ghirardi. The Conway–Kochen argument and relativistic GRW models. Foundations of Physics, 37(2): 169–185, February 2007. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-006-9097-1. See reply [3112].

Tumulka:2007:CFW

[3044] Roderich Tumulka. Comment on “The Free Will Theorem”. Foundations of Physics, 37(2):186–197, February 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-006-9098-0. See [3015, 3043, 3112].

Klauber:2007:RRC

[3045] Robert D. Klauber. Relativistic rotation: a comparison of theories. Foundations of Physics, 37(2):198–252, February 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-006-9099-z.

Muller:2007:BDC

[3046] Thomas Müller. Branch dependence in the “consistent histories” approach to quantum mechanics. Foundations of Physics, 37(2):253– 276, February 2007. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-006-9100-x.

Recami:2007:DSW

[3047] Erasmo Recami and Giovanni Salesi. Deriving spin within a discrete-time theory. Foundations of Physics, 37(2):277–294, February 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-006-9101-9.

Afshar:2007:PWP

[3048] Shahriar S. Afshar, Eduardo Flores, Keith F. McDonald, and Ernst Knoesel. Paradox in wave–particle duality. Foundations of Physics, 37 (2):295–305, February 2007. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-006-9102-8.

Agashe:2007:AEE

[3049] S. D. Agashe. Addendum to “Einstein’s ‘Zur Electrodynamik … ’ (1905) Revisited, with some Consequences”^{(1)} by S. D. Agashe. Foundations of Physics, 37(2):306–309, February 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-006-9103-7. See [3000].

Norsen:2007:AR

[3050] Travis Norsen. Against ‘realism’. Foundations of Physics, 37(3):311–340, March 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
007-9104-1.

Garecki:2007:TAR

[3051] Janusz Garecki. The tensors of the averaged relative energy–momentum and angular momentum in general relativity and some of their applications. Foundations of Physics, 37(3):341–365, March 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-007-9107-y.

Castro:2007:DEW

[3052] Carlos Castro. On dark energy, Weyl’s geometry, different derivations of the vacuum energy density and the pioneer anomaly. Foundations of Physics, 37(3):366–409, March 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-007-9106-z.

Jimenez:2007:MMA

[3053] J. L. Jiménez, J. A. E. Roa-Neri, and P. Vargas. A mechanical model for analyzing the runaway solutions in the radiation reaction problem. Foundations of Physics, 37(3):410–426, March 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-007-9109-9.

Smerlak:2007:RE

[3054] Matteo Smerlak and Carlo Rovelli. Relational EPR. Foundations of Physics, 37(3):427–445, March 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-007-9105-0.

Hartnett:2007:LDA

[3055] John G. Hartnett and Firmin J. Oliveira. Luminosity distance, angular size and surface brightness in cosmological general relativity. Foundations of Physics, 37(3):446–454, March 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-007-9108-x.

vanderMerwe:2007:EN

[3056] Alwyn van der Merwe. Editorial note. Foundations of Physics, 37 (4–5):455, May 2007. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-007-9143-7.

Horwitz:2007:P

[3057] L. P. Horwitz. Preface. Foundations of Physics, 37(4–5):456–459, May 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
007-9125-9.

Galley:2007:GSF

[3058] Chad R. Galley. Gravitational self-force from quantized linear metric perturbations in curved space. Foundations of Physics, 37(4–5):460–479, May 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
007-9111-2.

Lin:2007:NIU

[3059] Shih-Yuin Lin and B. L. Hu. New insights into uniformly accelerated detector in a quantum field. Foundations of Physics, 37(4–5):480–490, May 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
007-9120-1.

Lindesay:2007:CCP

[3060] James Lindesay. Consequences of a cosmological phase transition at the TeV scale. Foundations of Physics, 37(4–5):491–531, May 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-007-9115-y.

Mannheim:2007:SGP

[3061] Philip D. Mannheim. Solution to the ghost problem in fourth order derivative theories. Foundations of Physics, 37(4–5):532–571, May 2007.

CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-007-9119-7.

Seidewitz:2007:UES

[3062] Ed Seidewitz. The Universe as an eigenstate: Spacetime paths and decoherence. Foundations of Physics, 37(4–5):572–596, May 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-007-9123-y.

Land:2007:CPL

[3063] Martin Land. Coulomb potential from Lorentz invariance in N dimensions. Foundations of Physics, 37(4–5):597–631, May 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-007-9117-9.

Silverman:2007:CCQ

[3064] Mark P. Silverman. Condensates in the cosmos: Quantum stabilization of the collapse of relativistic degenerate stars to black holes. Foundations of Physics, 37(4–5):632–669, May 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-007-9121-0.

Santilli:2007:ENH

[3065] Ruggero Maria Santilli. The etherino and/or the neutrino hypothesis. Foundations of Physics, 37(4–5):670–711, May 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-007-9122-z.

Ansoldi:2007:SKP

[3066] Stefano Ansoldi and Eduardo I. Guendelman. Solitons as key parts to produce a Universe in the Laboratory. Foundations of Physics, 37 (4–5):712–722, May 2007. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-007-9113-0.

Cox:2007:GFE

[3067] David Eric Cox, James G. O’Brien, Ronald L. Mallett, and Chandra Roychoudhuri. Gravitational Faraday effect produced by a ring laser. Foundations of Physics, 37(4–5):723–733, May 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-007-9114-z.

Horwitz:2007:QIT

[3068] Lawrence P. Horwitz. Quantum interference in time. Foundations of Physics, 37(4–5):734–746, May 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-007-9127-7.

Ivezic:2007:TNP

[3069] Tomislav Ivezić. Trouton–Noble Paradox revisited. Foundations of Physics, 37(4–5):747–760, May 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-007-9116-x.

Piron:2007:FCN

[3070] Constantin Piron. Further clarification of “New Einstein Gravitation”. Foundations of Physics, 37(4–5):761–762, May 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-007-9128-6. See [2927].

Harpaz:2007:EFG

[3071] Amos Harpaz. Electric field in a gravitational field. Foundations of Physics, 37(4–5):763–772, May 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-007-9118-8.

Ben-Amots:2007:REG

[3072] N. Ben-Amots. Relativistic exponential gravitation and exponential potential of electric charge. Foundations of Physics, 37(4–5):773–787, May 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
007-9112-1.

Whitney:2007:RDB

[3073] Cynthia Kolb Whitney. Relativistic dynamics in basic chemistry. Foundations of Physics, 37(4–5):788–812, May 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-007-9124-x.

Rapoport:2007:TFC

[3074] Diego L. Rapoport. Torsion fields, Cartan–Weyl space–time and state– space quantum geometries, their Brownian motions, and the time variables. Foundations of Physics, 37(4–5):813–854, May 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-007-9126-8.

Domenech:2007:DPN

[3075] Graciela Domenech and Federico Holik. A discussion on particle number and quantum indistinguishability. Foundations of Physics, 37(6): 855–878, June 2007. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-007-9129-5.

Korbicz:2007:RGT

[3076] J. K. Korbicz and M. Lewenstein. Remark on a group-theoretical formalism for quantum mechanics and the quantum-to-classical transition. Foundations of Physics, 37(6):879–896, June 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-007-9130-z.

Iorio:2007:CPA

[3077] Lorenzo Iorio. Can the Pioneer anomaly be of gravitational origin? A phenomenological answer. Foundations of Physics, 37(6):897–918, June 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
007-9132-x.

Cirilo-Lombardo:2007:NCGa

[3078] Diego Julio Cirilo-Lombardo. Non-compact groups, coherent states, relativistic wave equations and the harmonic oscillator. Foundations of Physics, 37(6):919–950, June 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-007-9133-9.

Wallace:2007:NLGa

[3079] David Wallace and Christopher G. Timpson. Non-locality and gauge freedom in Deutsch and Hayden’s formulation of quantum mechanics. Foundations of Physics, 37(6):951–955, June 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-007-9135-7.

Esposito:2007:HTQa

[3080] Salvatore Esposito. Hole theory and quantum electrodynamics in an unknown manuscript in French by Ettore Majorana. Foundations of Physics, 37(6):956–976, June 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-007-9136-6.

Bock:2007:RQR

[3081] Robert D. Bock. Radial quantization in rotating space–times. Foundations of Physics, 37(6):977–988, June 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-007-9137-5.

Whitaker:2007:SAH

[3082] M. A. B. Whitaker. Solomon’s argument on hidden variables in quantum theory. Foundations of Physics, 37(6):989–997, June 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-007-9138-4.

Boyer:2007:CBR

[3083] Timothy H. Boyer. Connecting blackbody radiation, relativity, and discrete charge in classical electrodynamics. Foundations of Physics, 37 (7):999–1026, July 2007. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-007-9139-3.

Tyc:2007:NACa

[3084] Tomás Tyc, Brett Hamilton, Barry C. Sanders, and William D. Oliver. No approximate complex fermion coherent states. Foundations of Physics, 37(7):1027–1048, July 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-007-9140-x.

Esposito:2007:HTQb

[3085] S. Esposito. Hole theory and quantum electrodynamics in an unknown manuscript in French by Ettore Majorana. Foundations of Physics, 37 (7):1049–1068, July 2007. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-007-9141-9.

Wallace:2007:NLGb

[3086] David Wallace and Christopher G. Timpson. Non-locality and gauge freedom in Deutsch and Hayden’s formulation of quantum mechanics. Foundations of Physics, 37(7):1069–1073, July 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-007-9142-8.

Gambini:2007:RPR

[3087] Rodolfo Gambini and Jorge Pullin. Relational physics with real rods and clocks and the measurement problem of quantum mechanics. Foundations of Physics, 37(7):1074–1092, July 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-007-9144-6.

Beck:2007:TTP

[3088] T. Beck and M. Sorg. Two- and three-particle systems in relativistic Schrödinger theory. Foundations of Physics, 37(7):1093–1147, July 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-007-9145-5.

Cirilo-Lombardo:2007:NCGb

[3089] Diego Julio Cirilo-Lombardo. Non-compact groups, coherent states, relativistic wave equations and the harmonic oscillator. Foundations of Physics, 37(8):1149–1180, August 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-007-9110-3.

Lindesay:2007:CNS

[3090] James Lindesay. Coordinates with non-singular curvature for a time dependent black hole horizon. Foundations of Physics, 37(8):1181– 1196, August 2007. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-007-9146-4.

Pavsic:2007:UTG

[3091] M. Pavsic. On a unified theory of generalized branes coupled to gauge fields, including the gravitational and Kalb–Ramond fields. Foundations of Physics, 37(8):1197–1242, August 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-007-9147-3.

delaTorre:2007:ONV

[3092] Alberto C. de la Torre. Observables have no value: a no-go theorem for position and momentum observables. Foundations of Physics, 37(8): 1243–1252, August 2007. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-007-9148-2.

Binicioglu:2007:ESS

[3093] Sinem Binicioğlu, M. Ali Can, Alexander A. Klyachko, and Alexander S. Shumovsky. Entanglement of a single spin-1 object: an example of ubiquitous entanglement. Foundations of Physics, 37(8):1253–1277, August 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
007-9149-1.

Scardigli:2007:QLD

[3094] Fabio Scardigli. A quantum-like description of the planetary systems. Foundations of Physics, 37(8):1278–1295, August 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-007-9151-7.

Faria:2007:VEF

[3095] A. J. Faria, H. M. França, G. G. Gomes, and R. C. Sponchiado. The vacuum electromagnetic fields and the Schrödinger equation. Foundations of Physics, 37(8):1296–1305, August 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-007-9152-6.

Entov:2007:AGP

[3096] Michael Entov, Leonid Polterovich, and Frol Zapolsky. An “anti-Gleason” phenomenon and simultaneous measurements in classical mechanics. Foundations of Physics, 37(8):1306–1316, August 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-007-9158-0.

Kundt:2007:FP

[3097] Wolfgang Kundt. Fundamental physics. Foundations of Physics, 37(9): 1317–1369, September 2007. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-007-9150-8.

Steuernagel:2007:AED

[3098] Ole Steuernagel. Afshar’s experiment does not show a violation of complementarity. Foundations of Physics, 37(9):1370–1385, September 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-

007-9153-5. See reply [3173].

Floyd:2007:IRA

[3099] Edward R. Floyd. Interference, reduced action, and trajectories. Foundations of Physics, 37(9):1386–1402, September 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-007-9154-4.

Floyd:2007:WWT

[3100] Edward R. Floyd. Welcher Weg? A trajectory representation of a quantum Young’s diffraction experiment. Foundations of Physics, 37(9):1403– 1420, September 2007. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-007-9155-3.

Foulis:2007:EOS

[3101] David J. Foulis. Effects, observables, states, and symmetries in physics. Foundations of Physics, 37(10):1421–1446, October 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-007-9170-4.

vanEnk:2007:TMQ

[3102] S. J. van Enk. A toy model for quantum mechanics. Foundations of Physics, 37(10):1447–1460, October 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-007-9171-3.

Lewis:2007:TLH

[3103] Peter J. Lewis. Towards a local hidden variable theory. Foundations of Physics, 37(10):1461–1469, October 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-007-9172-2.

Ishak:2007:RFC

[3104] Mustapha Ishak. Remarks on the formulation of the cosmological Constant/ dark energy problems. Foundations of Physics, 37(10):1470– 1498, October 2007. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-007-9175-z.

Petkov:2007:RMS

[3105] Vesselin Petkov. On the reality of Minkowski space. Foundations of Physics, 37(10):1499–1502, October 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10701-007-9178-9.pdf.

Aldrovandi:2007:NEG

[3106] R. Aldrovandi, J. G. Pereira, and K. H. Vu. The nonlinear essence of gravitational waves. Foundations of Physics, 37(10):1503–1517, October 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
007-9180-2.

Tyc:2007:NACb

[3107] Tomás Tyc, Brett Hamilton, Barry C. Sanders, and William D. Oliver. No approximate complex fermion coherent states. Foundations of Physics, 37(11):1519–1539, November 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-007-9131-y.

Galvan:2007:TVP

[3108] Bruno Galvan. Typicality vs. probability in trajectory-based formulations of quantum mechanics. Foundations of Physics, 37(11):1540– 1562, November 2007. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-007-9174-0.

Nikolic:2007:QMM

[3109] Hrvoje Nikolić. Quantum mechanics: Myths and facts. Foundations of Physics, 37(11):1563–1611, November 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-007-9176-y.

Nauenberg:2007:CQE

[3110] Michael Nauenberg. Critique of “quantum enigma: Physics encounters consciousness”. Foundations of Physics, 37(11):1612–1627, November 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-

007-9179-8. See [3137].

Israelit:2007:CNF

[3111] Mark Israelit. Creation of neutral fundamental particles in the Weyl– Dirac version of Wesson’s IMT. Foundations of Physics, 37(11):1628– 1642, November 2007. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-007-9181-1.

Conway:2007:RCB

[3112] John Conway and Simon Kochen. Reply to comments of Bassi, Ghirardi, and Tumulka on the Free Will Theorem. Foundations of Physics, 37(11): 1643–1647, November 2007. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-007-9173-1. See [3015, 3043, 3044].

Chubykalo:2007:RCH

[3113] A. Chubykalo, A. Espinoza, R. Alvarado Flores, and A. Gutierrez Rodriguez. Reply to ‘Comment on “Helmholtz Theorem and the V -Gauge in the Problem of Superluminal and Instantaneous Signals in Classical Electrodynamics” by A. Chubykalo et al.’ by J. A. Heras [Found. Phys. Lett. vol. 19(6) p. 579 (2006)]. Foundations of Physics, 37(11):1648– 1652, November 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10701-007-9183-z.pdf.

Sheehan:2007:SLT

[3114] D. P. Sheehan. The Second Law of Thermodynamics: Foundations and status. Foundations of Physics, 37(12):1653–1658, December 2007. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-007-9164-2.

tHooft:2007:Eb

[3115] Gerard ’t Hooft. Editorial. Foundations of Physics, 37(12):1659, December 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10701-007-9184-y.pdf.

Miller:2007:ISL

[3116] S. L. Miller. Insights into the Second Law of Thermodynamics from anisotropic gas–surface interactions. Foundations of Physics, 37(12): 1660–1684, December 2007. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-007-9167-z.

Denur:2007:SDW

[3117] Jack Denur. Speed-dependent weighting of the Maxwellian distribution in rarefied gases: a Second-Law Paradox? Foundations of Physics, 37(12): 1685–1706, December 2007. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-007-9160-6.

Crosignani:2007:RFD

[3118] Bruno Crosignani and Paolo Di Porto. Random fluctuations of diathermal and adiabatic pistons. Foundations of Physics, 37(12):1707– 1715, December 2007. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-007-9165-1.

Schulman:2007:DIE

[3119] L. S. Schulman. Decoherence induced equilibration. Foundations of Physics, 37(12):1716–1726, December 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-007-9169-x.

Burns:2007:VRE

[3120] Jean E. Burns. Vacuum radiation, entropy, and molecular chaos. Foundations of Physics, 37(12):1727–1737, December 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-007-9161-5.

Berger:2007:NST

[3121] Jorge Berger. A nonconventional scenario for thermal equilibrium. Foundations of Physics, 37(12):1738–1743, December 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-007-9162-4.

Leff:2007:ELI

[3122] Harvey S. Leff. Entropy, its language, and interpretation. Foundations of Physics, 37(12):1744–1766, December 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-007-9163-3.

Duncan:2007:ILF

[3123] T. L. Duncan and J. S. Semura. Information loss as a foundational principle for the Second Law of Thermodynamics. Foundations of Physics, 37(12):1767–1773, December 2007. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-007-9159-z.

Sheehan:2007:TL

[3124] D. P. Sheehan. Thermosynthetic life. Foundations of Physics, 37(12): 1774–1797, December 2007. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-007-9168-y.

tHooft:2008:EN

[3125] Gerard ’t Hooft. Editorial note. Foundations of Physics, 38(1):1–2, January 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
007-9187-8.

Bruhn:2008:NLI

[3126] Gerhard W. Bruhn. On the non-Lorentz-invariance of M. W. Evans’ O(3)-symmetry law. Foundations of Physics, 38(1):3–6, January 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-007-9189-6.

Hehl:2008:AEUa

[3127] Friedrich W. Hehl. An assessment of Evans’ unified field theory I. Foundations of Physics, 38(1):7–37, January 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-007-9190-0.

Hehl:2008:AEUb

[3128] Friedrich W. Hehl and Yuri N. Obukhov. An assessment of Evans’ unified field theory II. Foundations of Physics, 38(1):38–46, January 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-007-9188-7.

Mashburn:2008:OMI

[3129] Joe Mashburn. An order model for infinite classical states. Foundations of Physics, 38(1):47–75, January 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-007-9182-0.

Kosyakov:2008:CRC

[3130] B. P. Kosyakov. Is classical reality completely deterministic? Foundations of Physics, 38(1):76–88, January 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-007-9185-x.

Sidharth:2008:DRL

[3131] B. G. Sidharth. Different routes to Lorentz symmetry violations. Foundations of Physics, 38(1):89–95, January 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-007-9193-x.

Esposito:2008:CAO

[3132] Giampiero Esposito. Comment and addendum to “On the Occurrence of Mass in Field Theory”. Foundations of Physics, 38(1):96–98, January 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-

007-9191-z. See [2623].

Cirilo-Lombardo:2008:NCG

[3133] Diego Julio Cirilo-Lombardo. Non-compact groups, coherent states, relativistic wave equations and the harmonic oscillator. Foundations of Physics, 38(1):99, January 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-007-9192-y.

Tegmark:2008:MU

[3134] Max Tegmark. The mathematical Universe. Foundations of Physics, 38 (2):101–150, February 2008. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-007-9186-9. See comments [3220].

Huang:2008:EER

[3135] Young-Sea Huang and Kang-Hao Lu. Exact expression for radiation of an accelerated charge in classical electrodynamics. Foundations of Physics, 38(2):151–159, February 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-007-9194-9. See comment [3623].

Doran:2008:ISB

[3136] Rosa Doran, Francisco S. N. Lobo, and Paulo Crawford. Interior of a Schwarzschild black hole revisited. Foundations of Physics, 38(2): 160–187, February 2008. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-007-9197-6.

Kuttner:2008:RNC

[3137] Fred Kuttner. Response to Nauenberg’s “critique of quantum enigma: Physics encounters consciousness”. Foundations of Physics, 38(2):188– 190, February 2008. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-007-9195-8. See [3110].

Barcelo:2008:RLF

[3138] Carlos Barceló and Gil Jannes. A real Lorentz–FitzGerald contraction. Foundations of Physics, 38(2):191–199, February 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-007-9196-7.

Hartnett:2008:ERD

[3139] John G. Hartnett. Extending the redshift-distance relation in cosmological general relativity to higher redshifts. Foundations of Physics, 38 (3):201–215, March 2008. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-007-9198-5.

Bohmer:2008:PDE

[3140] C. G. Böhmer and T. Harko. Physics of dark energy particles. Foundations of Physics, 38(3):216–227, March 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-007-9199-4.

DiGrezia:2008:MQS

[3141] E. Di Grezia and S. Esposito. Majorana and the quasi-stationary states in nuclear physics. Foundations of Physics, 38(3):228–240, March 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-007-9200-2.

Niestegge:2008:AQM

[3142] Gerd Niestegge. An approach to quantum mechanics via conditional probabilities. Foundations of Physics, 38(3):241–256, March 2008. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-007-9201-1.

Aiello:2008:ATU

[3143] Matías Aiello, Mario Castagnino, and Olimpia Lombardi. The arrow of time: From Universe time–asymmetry to local irreversible processes. Foundations of Physics, 38(3):257–292, March 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-007-9202-0.

Marsh:2008:CGEa

[3144] Gerald E. Marsh. Charge, geometry, and effective mass. Foundations of Physics, 38(3):293–300, March 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-008-9209-1.

Sen:2008:PMC

[3145] R. N. Sen. Physics and the measurement of continuous variables. Foundations of Physics, 38(4):301–316, April 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-007-9203-z.

Manchak:2008:PPG

[3146] John Byron Manchak. Is prediction possible in general relativity? Foundations of Physics, 38(4):317–321, April 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-008-9204-6.

Zhao:2008:EES

[3147] Shuang Zhao, Hans De Raedt, and Kristel Michielsen. Event-by-event simulation of Einstein–Podolsky–Rosen–Bohm experiments. Foundations of Physics, 38(4):322–347, April 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10701-008-9205-5.pdf.

Stuckey:2008:RSQ

[3148] W. M. Stuckey, Michael Silberstein, and Michael Cifone. Reconciling spacetime and the quantum: Relational blockworld and the Quantum Liar Paradox. Foundations of Physics, 38(4):348–383, April 2008. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-008-9206-4.

Dasgupta:2008:MDQ

[3149] Ratan Dasgupta and Sisir Roy. Multinomial distribution, quantum statistics and Einstein–Podolsky–Rosen like phenomena. Foundations of Physics, 38(4):384–394, April 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-008-9207-3.

Hurley:2008:UFR

[3150] D. J. Hurley and M. A. Vandyck. A unified framework for relativity and curvilinear–time Newtonian mechanics. Foundations of Physics, 38 (4):395–408, April 2008. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-008-9208-2.

Broekaert:2008:SVG

[3151] Jan Broekaert. A spatially-VSL gravity model with 1−P N limit of GRT. Foundations of Physics, 38(5):409–435, May 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-008-9210-8.

Whitaker:2008:CSI

[3152] M. A. B. Whitaker. Can the statistical interpretation of quantum mechanics be inferred from the Schrödinger equation? — Bell and Gottfried. Foundations of Physics, 38(5):436–447, May 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-008-9211-7.

VandenNest:2008:MBQ

[3153] Maarten Van den Nest and Hans J. Briegel. Measurement-based quantum computation and undecidable logic. Foundations of Physics, 38 (5):448–457, May 2008. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-008-9212-6.

Ranada:2008:TCP

[3154] Antonio F. Rañada and A. Tiemblo. Time, clocks and parametric invariance. Foundations of Physics, 38(5):458–469, May 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-008-9214-4.

Catren:2008:CQO

[3155] Gabriel Catren. On classical and quantum objectivity. Foundations of Physics, 38(5):470–487, May 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-008-9216-2.

Yahalom:2008:GMT

[3156] Asher Yahalom. The geometrical meaning of time. Foundations of Physics, 38(6):489–497, June 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-008-9215-3.

Boyer:2008:CER

[3157] Timothy H. Boyer. Comment on experiments related to the Aharonov– Bohm phase shift. Foundations of Physics, 38(6):498–505, June 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-

008-9217-1. See [2680].

Anders:2008:HMO

[3158] Janet Anders, Michal Hajdusek, Damian Markham, and Vlatko Vedral. How much of one-way computation is just thermodynamics? Foundations of Physics, 38(6):506–522, June 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-008-9218-0.

Vankov:2008:RGG

[3159] Anatoli Andrei Vankov. On relativistic generalization of gravitational force. Foundations of Physics, 38(6):523–545, June 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-008-9219-z.

Marochnik:2008:CAV

[3160] Leonid Marochnik, Daniel Usikov, and Grigory Vereshkov. Cosmological acceleration from virtual gravitons. Foundations of Physics, 38 (6):546–555, June 2008. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-008-9220-6.

Rosen:2008:QGP

[3161] Steven M. Rosen. Quantum gravity and phenomenological philosophy. Foundations of Physics, 38(6):556–582, June 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-008-9221-5.

Suarez:2008:NRL

[3162] Antoine Suarez. Nonlocal “realistic” Leggett models can be considered refuted by the before–before experiment. Foundations of Physics, 38 (6):583–589, June 2008. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-008-9228-y. See [3192, 3206].

Higashi:2008:LCC

[3163] Katsuaki Higashi. The limits of common cause approach to EPR correlation. Foundations of Physics, 38(7):591–609, July 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-008-9222-4.

Vargas:2008:FQM

[3164] Jose G. Vargas. The foundations of quantum mechanics and the evolution of the Cartan–Kähler calculus. Foundations of Physics, 38(7):610–647, July 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
008-9223-3.

Pegg:2008:RQM

[3165] David T. Pegg. Retrocausality and quantum measurement. Foundations of Physics, 38(7):648–658, July 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-008-9224-2.

Catillon:2008:SBP

[3166] P. Catillon, N. Cue, M. J. Gaillard, R. Genre, M. Gouanère, R. G. Kirsch, J.-C. Poizat, J. Remillieux, L. Roussel, and M. Spighel. A search for the de Broglie particle internal clock by means of electron channeling. Foundations of Physics, 38(7):659–664, July 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-008-9225-1.

Nyambuya:2008:NCS

[3167] G. G. Nyambuya. New curved spacetime Dirac equations. Foundations of Physics, 38(7):665–677, July 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-008-9226-0.

Kosyakov:2008:BHI

[3168] B. P. Kosyakov. Black holes: Interfacing the classical and the quantum. Foundations of Physics, 38(7):678–694, July 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-008-9227-z.

Sidharth:2008:LSR

[3169] B. G. Sidharth. The limits of special relativity. Foundations of Physics, 38(8):695–706, August 2008. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-008-9229-x.

Held:2008:AQM

[3170] Carsten Held. Axiomatic quantum mechanics and completeness. Foundations of Physics, 38(8):707–732, August 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10701-008-9230-4.pdf.

tHooft:2008:LFM

[3171] Gerard ’t Hooft. A locally finite model for gravity. Foundations of Physics, 38(8):733–757, August 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10701-008-9231-3.pdf.

Matsutani:2008:GOG

[3172] Shigeki Matsutani. Gauss optics and Gauss sum on an optical phenomena. Foundations of Physics, 38(8):758–777, August 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-008-9233-1.

Flores:2008:RCS

[3173] Eduardo V. Flores. Reply to comments of Steuernagel on the Afshar’s experiment. Foundations of Physics, 38(8):778–781, August 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-

008-9234-0. See [3098].

Niestegge:2008:RQM

[3174] Gerd Niestegge. A representation of quantum measurement in order-unit spaces. Foundations of Physics, 38(9):783–795, September 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-008-9236-y.

Schlosshauer:2008:CEI

[3175] M. Schlosshauer. Classicality, the ensemble interpretation, and decoherence: Resolving the Hyperion dispute. Foundations of Physics, 38(9): 796–803, September 2008. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-008-9237-x. See reply [3182].

Marchildon:2008:RER

[3176] Louis Marchildon. On relativistic elements of reality. Foundations of Physics, 38(9):804–817, September 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-008-9238-9.

Helland:2008:QMF

[3177] Inge S. Helland. Quantum mechanics from focusing and symmetry. Foundations of Physics, 38(9):818–842, September 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-008-9239-8.

Bachtold:2008:IQM

[3178] Manuel Bächtold. Interpreting quantum mechanics according to a pragmatist approach. Foundations of Physics, 38(9):843–868, September 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
008-9240-2.

Nikolic:2008:PRB

[3179] Hrvoje Nikolić. Probability in relativistic Bohmian mechanics of particles and strings. Foundations of Physics, 38(9):869–881, September 2008.

CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-008-9241-1.

Burgos:2008:TCT

[3180] M. E. Burgos. Transitions to the continuum: Three different approaches. Foundations of Physics, 38(10):883–907, October 2008. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-008-9213-5.

Westman:2008:EOG

[3181] Hans Westman and Sebastiano Sonego. Events and observables in generally invariant spacetime theories. Foundations of Physics, 38(10): 908–915, October 2008. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-008-9235-z.

Ballentine:2008:CDR

[3182] Leslie Ballentine. Classicality without decoherence: a reply to Schlosshauer. Foundations of Physics, 38(10):916–922, October 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s10701-008-9242-0.pdf. See [3175].

Dieks:2008:IQP

[3183] Dennis Dieks and Marijn A. M. Versteegh. Identical quantum particles and weak discernibility. Foundations of Physics, 38(10):923–934, October 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10701-008-9243-z.pdf.

Sestito:2008:DIP

[3184] A. Sestito. Detection of incompatible properties in a double-slit experiment. Foundations of Physics, 38(10):935–958, October 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-008-9244-y.

Marsh:2008:CGEb

[3185] Gerald E. Marsh. Charge, geometry, and effective mass in the Kerr– Newman solution to the Einstein field equations. Foundations of Physics, 38(10):959–968, October 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-008-9245-x.

Domenech:2008:QSF

[3186] Graciela Domenech, Federico Holik, and Décio Krause. Q-spaces and the foundations of quantum mechanics. Foundations of Physics, 38(11): 969–994, November 2008. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-008-9246-9.

Delbourgo:2008:BRP

[3187] R. Delbourgo and D. Lashmar. Born reciprocity and the 1/r potential. Foundations of Physics, 38(11):995–1010, November 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-008-9247-8.

Ghirardi:2008:RMP

[3188] GianCarlo Ghirardi. Reconsidering Mermin’s “in praise of measurement”. Foundations of Physics, 38(11):1011–1019, November 2008. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-008-9248-7.

Arminjon:2008:DTE

[3189] Mayeul Arminjon. Dirac-type equations in a gravitational field, with vector wave function. Foundations of Physics, 38(11):1020–1045, November 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
008-9249-6.

Herbut:2008:ETE

[3190] F. Herbut. On EPR-type entanglement in the experiments of Scully et al. I. The micromaser case and delayed-choice quantum erasure. Foundations of Physics, 38(11):1046–1064, November 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-008-9251-z.

Monroe:2008:CVU

[3191] Hunter Monroe. Are causality violations undesirable? Foundations of Physics, 38(11):1065–1069, November 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-008-9254-9.

Zukowski:2008:CNR

[3192] Marek Zukowski. Comment on: Nonlocal “realistic” Leggett models can be considered refuted by the before–before experiment. Foundations of Physics, 38(11):1070–1071, November 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-008-9250-0. See [3206, 3162].

Rave:2008:IQI

[3193] M. J. Rave. Interpreting quantum interference using a Berry’s phaselike quantity. Foundations of Physics, 38(12):1073–1081, December 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-008-9252-y.

Slobodov:2008:UCT

[3194] Sergei Slobodov. Unwrapping closed timelike curves. Foundations of Physics, 38(12):1082–1109, December 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-008-9253-x.

Laudisa:2008:NLR

[3195] Federico Laudisa. Non-local realistic theories and the Scope of the Bell theorem. Foundations of Physics, 38(12):1110–1132, December 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-008-9255-8.

Heinosaari:2008:NJM

[3196] Teiko Heinosaari, Daniel Reitzner, and Peter Stano. Notes on joint measurability of quantum observables. Foundations of Physics, 38(12): 1133–1147, December 2008. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-008-9256-7.

Alexander:2008:QGF

[3197] Stephon H. S. Alexander and Gianluca Calcagni. Quantum gravity as a Fermi liquid. Foundations of Physics, 38(12):1148–1184, December 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-008-9257-6.

Aldrovandi:2009:SRN

[3198] R. Aldrovandi and J. G. Pereira. De Sitter relativity: a new road to quantum gravity? Foundations of Physics, 39(1):1–19, January 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-008-9258-5.

Allori:2009:CLQ

[3199] Valia Allori and Nino Zanghì. On the classical limit of quantum mechanics. Foundations of Physics, 39(1):20–32, January 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-008-9259-4.

Atiq:2009:NWE

[3200] Mahdi Atiq, Mozafar Karamian, and Mahdi Golshani. A new way for the extension of quantum theory: Non-Bohmian quantum potentials. Foundations of Physics, 39(1):33–44, January 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-008-9260-y.

Scholz:2009:CSB

[3201] Erhard Scholz. Cosmological spacetimes balanced by a Weyl geometric scale covariant scalar field. Foundations of Physics, 39(1):45–72, January 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
008-9261-x.

Schmelzer:2009:CMI

[3202] I. Schmelzer. A condensed matter interpretation of SM fermions and gauge fields. Foundations of Physics, 39(1):73–107, January 2009. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-008-9262-9.

Angelo:2009:IET

[3203] Renato M. Angelo. On the interpretative essence of the term “interactionfree measurement”: The role of entanglement. Foundations of Physics, 39 (2):109–119, February 2009. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-008-9263-8.

Niestegge:2009:RQM

[3204] Gerd Niestegge. A representation of quantum measurement in nonassociative algebras. Foundations of Physics, 39(2):120–136, February 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-008-9264-7.

Mohrhoff:2009:OPQ

[3205] U. Mohrhoff. Objective probability and quantum fuzziness. Foundations of Physics, 39(2):137–155, February 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-008-9266-5.

Suarez:2009:BSS

[3206] Antoine Suarez. On Bell, Suarez–Scarani, and Leggett experiments: Reply to a comment by Marek Zukowski in [Found. Phys. 38:1070, 2008]. Foundations of Physics, 39(2):156–159, February 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-

008-9267-4. See [3162, 3192].

Laura:2009:TTQ

[3207] Roberto Laura and Leonardo Vanni. Time translation of quantum properties. Foundations of Physics, 39(2):160–173, February 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-008-9268-3.

Hudson:2009:MSR

[3208] Robert Hudson. The methodological strategy of robustness in the context of experimental WIMP research. Foundations of Physics, 39(2): 174–193, February 2009. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-009-9271-3.

deGosson:2009:SCU

[3209] Maurice A. de Gosson. The symplectic camel and the uncertainty principle: The tip of an iceberg? Foundations of Physics, 39(2): 194–214, February 2009. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-009-9272-2.

Lampe:2009:DTM

[3210] Bodo Lampe. Development of the tetron model. Foundations of Physics, 39(3):215–236, March 2009. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-009-9278-9.

Foulis:2009:SFG

[3211] David J. Foulis and Sylvia Pulmannová. Spin factors as generalized Hermitian algebras. Foundations of Physics, 39(3):237–255, March 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-009-9279-8.

Orefice:2009:CHD

[3212] A. Orefice, R. Giovanelli, and D. Ditto. Complete Hamiltonian description of wave-like features in classical and quantum physics. Foundations of Physics, 39(3):256–272, March 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-009-9280-2.

Norsen:2009:LCC

[3213] Travis Norsen. Local causality and completeness: Bell vs. Jarrett. Foundations of Physics, 39(3):273–294, March 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-009-9281-1.

Caprez:2009:FRE

[3214] Adam Caprez and Herman Batelaan. Feynman’s relativistic electrodynamics paradox and the Aharonov–Bohm effect. Foundations of Physics, 39(3):295–306, March 2009. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-009-9283-z.

Poplawski:2009:GEC

[3215] Nikodem J. Poplawski. Gravitation, electromagnetism and cosmological constant in purely affine gravity. Foundations of Physics, 39(3):307–330, March 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
009-9284-y.

Boughn:2009:NG

[3216] Stephen Boughn. Nonquantum gravity. Foundations of Physics, 39 (4):331–351, April 2009. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-009-9282-0.

Alicki:2009:NBT

[3217] Robert Alicki. On von Neumann and Bell theorems applied to quantumness tests. Foundations of Physics, 39(4):352–360, April 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-009-9285-x.

Sancho:2009:CBI

[3218] Pedro Sancho. Correlations between the interacting fragments of decaying processes. Foundations of Physics, 39(4):361–372, April 2009. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-009-9288-7.

Cirilo-Lombardo:2009:NCG

[3219] Diego Julio Cirilo-Lombardo. Non-compact groups, coherent states, relativistic wave equations and the harmonic oscillator II: Physical and geometrical considerations. Foundations of Physics, 39(4):373–396, April 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
009-9289-6.

Jannes:2009:SCM

[3220] Gil Jannes. Some comments on “The Mathematical Universe”. Foundations of Physics, 39(4):397–406, April 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-009-9286-9. See [3134].

deHaro:2009:BRJ

[3221] Sebastian de Haro and Thomas van Lier. Book review: J. R. Leibowitz: Hidden Harmony. The Connected Worlds of Physics and Art. Foundations of Physics, 39(4):407–410, April 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10701-009-9291-z.pdf.

Hajicek:2009:IPQ

[3222] P. Hájícek and J. Tolar. Intrinsic properties of quantum systems. Foundations of Physics, 39(5):411–432, May 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-009-9296-7.

Hurley:2009:DHS

[3223] D. J. Hurley and M. A. Vandyck. D-differentiation in Hilbert space and the structure of quantum mechanics. Foundations of Physics, 39 (5):433–473, May 2009. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-009-9297-6.

Corbin:2009:SCL

[3224] Vincent Corbin and Neil J. Cornish. Semi-classical limit and minimum decoherence in the conditional probability interpretation of quantum mechanics. Foundations of Physics, 39(5):474–485, May 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-009-9298-5.

Schmelzer:2009:WHO

[3225] I. Schmelzer. Why the Hamilton operator alone is not enough. Foundations of Physics, 39(5):486–498, May 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-009-9299-4.

Sela:2009:CSR

[3226] Or Sela, Boaz Tamir, Shahar Dolev, and Avshalom C. Elitzur. Can special relativity be derived from Galilean mechanics alone? Foundations of Physics, 39(5):499–509, May 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-009-9301-1.

Tanaka:2009:KRS

[3227] Sho Tanaka. Kinematical reduction of spatial degrees of freedom and holographic relation in Yang’s quantized space–time algebra. Foundations of Physics, 39(5):510–518, May 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-008-9270-9.

Busch:2009:PJL

[3228] Paul Busch, Dennis Dieks, and Gerardus ’t Hooft. Pekka Johannes Lahti — 60th birthday. Foundations of Physics, 39(6):519– 520, June 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10701-009-9309-6.pdf.

Carmeli:2009:GOR

[3229] C. Carmeli, G. Cassinelli, and F. Zizzi. Generalized orthogonality relations and SU(1, 1)-quantum tomography. Foundations of Physics, 39 (6):521–549, June 2009. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-009-9290-0.

Buhagiar:2009:GTG

[3230] David Buhagiar, Emmanuel Chetcuti, and Anatolij Dvurecenskij. On Gleason’s theorem without Gleason. Foundations of Physics, 39(6): 550–558, June 2009. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-008-9265-6.

DallaChiara:2009:ASA

[3231] Maria Luisa Dalla Chiara, Roberto Giuntini, Hector Freytes, Antonio Ledda, and Giuseppe Sergioli. The algebraic structure of an approximately universal system of quantum computational gates. Foundations of Physics, 39(6):559–572, June 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-009-9302-0.

Gudder:2009:TEM

[3232] Stan Gudder. Transition effect matrices and quantum Markov chains. Foundations of Physics, 39(6):573–592, June 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-008-9269-2.

Ali:2009:CPF

[3233] S. Twareque Ali, Claudio Carmeli, Teiko Heinosaari, and Alessandro Toigo. Commutative POVMs and fuzzy observables. Foundations of Physics, 39(6):593–612, June 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-009-9292-y.

Jencova:2009:CCP

[3234] Anna Jencová and Sylvia Pulmannová. Characterizations of commutative POV measures. Foundations of Physics, 39(6):613–624, June 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-009-9273-1.

Holevo:2009:ECO

[3235] Alexander S. Holevo and Juha-Pekka Pellonpää. Extreme covariant observables for type I symmetry groups. Foundations of Physics, 39 (6):625–641, June 2009. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-009-9274-0.

Stenholm:2009:EPS

[3236] Stig Stenholm. Entanglement of pure states. Foundations of Physics, 39(6):642–655, June 2009. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-009-9276-y.

Ylinen:2009:QIR

[3237] Kari Ylinen. Quantum instruments and related transformation valued functions. Foundations of Physics, 39(6):656–675, June 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-009-9294-9.

Brukner:2009:IIQ

[3238] Caslav Brukner and Anton Zeilinger. Information invariance and quantum probabilities. Foundations of Physics, 39(7):677–689, July 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-009-9316-7.

Bub:2009:CNN

[3239] Jeffrey Bub and Allen Stairs. Contextuality and nonlocality in ‘no signaling’ theories. Foundations of Physics, 39(7):690–711, July 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-009-9307-8.

Busch:2009:SBQ

[3240] Paul Busch. On the sharpness and bias of quantum effects. Foundations of Physics, 39(7):712–730, July 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-009-9287-8.

Caspers:2009:IQL

[3241] Martijn Caspers, Chris Heunen, Nicolaas P. Landsman, and Bas Spitters. Intuitionistic quantum logic of an n-level system. Foundations of Physics, 39(7):731–759, July 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10701-009-9308-7.pdf.

Dieks:2009:OPR

[3242] Dennis Dieks. Objectivity in perspective: Relationism in the interpretation of quantum mechanics. Foundations of Physics, 39(7):760– 775, July 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10701-009-9293-x.pdf.

Ghirardi:2009:CNC

[3243] Gian Carlo Ghirardi and Karl Wienand. Contextuality, nonlocality and counterfactual arguments. Foundations of Physics, 39(7):776–789, July 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
009-9303-z.

Ann:2009:FTD

[3244] Kevin Ann and Gregg Jaeger. Finite-time destruction of entanglement and non-locality by environmental influences. Foundations of Physics, 39(7):790–828, July 2009. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-009-9295-8.

Kiukas:2009:TTC

[3245] J. Kiukas, A. Ruschhaupt, and R. F. Werner. Tunneling times with covariant measurements. Foundations of Physics, 39(7):829–846, July 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-009-9275-z.

Mittelstaedt:2009:CVC

[3246] Peter Mittelstaedt. Cognition versus constitution of objects: From Kant to modern physics. Foundations of Physics, 39(7):847–859, July 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-009-9277-x.

Anonymous:2009:PJL

[3247] Anonymous. Pekka Johannes Lahti — publications until 2009. Foundations of Physics, 39(7):860–868, July 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-009-9314-9.

Hewitt-Horsman:2009:IMW

[3248] Clare Hewitt-Horsman. An introduction to many worlds in quantum computation. Foundations of Physics, 39(8):869–902, August 2009. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-009-9300-2.

Matzkin:2009:BMQ

[3249] A. Matzkin. Bohmian mechanics, the quantum–classical correspondence and the classical limit: The case of the square billiard. Foundations of Physics, 39(8):903–920, August 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-009-9304-y.

Weinstein:2009:NN

[3250] Steven Weinstein. Nonlocality without nonlocality. Foundations of Physics, 39(8):921–936, August 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-009-9305-x.

Atkinson:2009:NEL

[3251] David Atkinson and Porter Johnson. Nonconservation of energy and loss of determinism I. Infinitely many colliding balls. Foundations of Physics, 39(8):937–957, August 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10701-009-9306-9.pdf.

Schurmann:2009:CLU

[3252] Thomas Schürmann and Ingo Hoffmann. A closer look at the uncertainty relation of position and momentum. Foundations of Physics, 39 (8):958–963, August 2009. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-009-9310-0.

Klein:2009:SEG

[3253] U. Klein. Schrödinger’s equation with gauge coupling derived from a continuity equation. Foundations of Physics, 39(8):964–995, August 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-009-9311-z.

Khrennikov:2009:DMB

[3254] Andrei Khrennikov. Detection model based on representation of quantum particles by classical random fields: Born’s rule and beyond. Foundations of Physics, 39(9):997–1022, September 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-009-9312-y.

Ardenghi:2009:QMM

[3255] Juan Sebastian Ardenghi, Mario Castagnino, and Olimpia Lombardi. Quantum mechanics: Modal interpretation and Galilean transformations. Foundations of Physics, 39(9):1023–1045, September 2009. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-009-9313-x.

Preti:2009:SRB

[3256] Giovanni Preti. Schwarzschild radius before general relativity: Why does Michell–Laplace argument provide the correct answer? Foundations of Physics, 39(9):1046–1054, September 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-009-9315-8.

Timko:2009:SDB

[3257] J. A. Timko and E. R. Vrscay. Spin-dependent Bohmian electronic trajectories for helium. Foundations of Physics, 39(9):1055–1071, September 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
009-9317-6.

Hajicek:2009:QMC

[3258] P. Hájícek. Quantum model of classical mechanics: Maximum entropy packets. Foundations of Physics, 39(9):1072–1096, September 2009. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-009-9318-5.

Page:2009:HG

[3259] D. N. Page. The height of a giraffe. Foundations of Physics, 39(10): 1097–1108, October 2009. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-009-9322-9.

Nikolic:2009:BFU

[3260] Hrvoje Nikolić. Boson–fermion unification, superstrings, and Bohmian mechanics. Foundations of Physics, 39(10):1109–1138, October 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-009-9323-8.

Kastner:2009:VAT

[3261] R. E. Kastner. On visibility in the Afshar two-slit experiment. Foundations of Physics, 39(10):1139–1144, October 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-009-9329-2.

Lewin:2009:WFC

[3262] K. Lewin. The wave function collapse as an effect of field quantization. Foundations of Physics, 39(10):1145–1160, October 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-009-9330-9.

Capozziello:2009:DED

[3263] S. Capozziello, M. De Laurentis, M. Francaviglia, and S. Mercadante. From dark energy & dark matter to dark metric. Foundations of Physics, 39(10):1161–1176, October 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-009-9332-7.

Franca:2009:SGP

[3264] Humberto M. França. The Stern–Gerlach phenomenon according to classical electrodynamics. Foundations of Physics, 39(10):1177–1190, October 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
009-9338-1.

Jelinek:2009:ETE

[3265] Karel Jelínek, Jirí Pavlu, Jaromír Havlica, and Jan Wild. Experimental test of the Evans’ B (3)-field: Measuring the interaction with free electrons. Foundations of Physics, 39(10):1191–1196, October 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-009-9339-0.

Page:2009:QMS

[3266] Don N. Page. Quantum mechanics as a simple generalization of classical mechanics. Foundations of Physics, 39(11):1197–1204, November 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-009-9324-7.

Aguilar:2009:ICC

[3267] José Edgar Madriz Aguilar and Carlos Romero. Inducing the cosmological constant from five-dimensional Weyl space. Foundations of Physics, 39(11):1205–1216, November 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-009-9340-7.

VanWilligenburg:2009:ESL

[3268] L. G. Van Willigenburg and W. L. De Koning. Emergence of the Second Law out of reversible dynamics. Foundations of Physics, 39(11):1217– 1239, November 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10701-009-9341-6.pdf.

delaPena:2009:QME

[3269] L. de la Peña, A. Valdés-Hernández, and A. M. Cetto. Quantum mechanics as an emergent property of ergodic systems embedded in the zeropoint radiation field. Foundations of Physics, 39(11):1240–1272, November 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
009-9348-z.

Maia:2009:QLQ

[3270] M. D. Maia, S. S. e Almeida Silva, and F. S. Carvalho. Quaternion-loop quantum gravity. Foundations of Physics, 39(11):1273–1279, November 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
009-9350-5.

Coll:2009:FCC

[3271] Bartolomé Coll, Joan Josep Ferrando, and Juan Antonio Morales-Lladosa. Four causal classes of Newtonian frames. Foundations of Physics, 39(11):1280–1295, November 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-009-9353-2.

Banerjee:2009:TNG

[3272] Rabin Banerjee, Biswajit Chakraborty, Subir Ghosh, Pradip Mukherjee, and Saurav Samanta. Topics in noncommutative geometry inspired physics. Foundations of Physics, 39(12):1297–1345, December 2009. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-009-9349-y.

Nimtz:2009:VPP

[3273] Günter Nimtz. On virtual phonons, photons, and electrons. Foundations of Physics, 39(12):1346–1355, December 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-009-9356-z.

Hamada:2009:RQG

[3274] Ken ji Hamada. Renormalizable 4D quantum gravity as a perturbed theory from CFT. Foundations of Physics, 39(12):1356–1393, December 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
009-9358-x.

Fiorentin:2009:RIC

[3275] Stefano Re Fiorentin. A re-interpretation of the concept of mass and of the relativistic mass–energy relation. Foundations of Physics, 39(12): 1394–1406, December 2009. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-009-9359-9.

Demir:2009:VEO

[3276] Durmus A. Demir. Vacuum energy as the origin of the gravitational constant. Foundations of Physics, 39(12):1407–1425, December 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-009-9364-z.

Hestenes:2010:ZQM

[3277] David Hestenes. Zitterbewegung in quantum mechanics. Foundations of Physics, 40(1):1–54, January 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-009-9360-3.

Bahrami:2010:PPO

[3278] M. Bahrami and A. Shafiee. Postponing the past: an operational analysis of delayed-choice experiments. Foundations of Physics, 40(1):55–92, January 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
009-9375-9.

Gambini:2010:UPO

[3279] Rodolfo Gambini, Luis Pedro García Pintos, and Jorge Pullin. Undecidability and the problem of outcomes in quantum measurements. Foundations of Physics, 40(1):93–115, January 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-009-9376-8.

Finster:2010:QOC

[3280] Felix Finster and Christian Hainzl. Quantum oscillations can prevent the big bang singularity in an Einstein–Dirac cosmology. Foundations of Physics, 40(1):116–124, January 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-009-9380-z.

Harrigan:2010:EIE

[3281] Nicholas Harrigan and Robert W. Spekkens. Einstein, incompleteness, and the epistemic view of quantum states. Foundations of Physics, 40 (2):125–157, February 2010. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-009-9347-0.

Rubin:2010:CPT

[3282] Jacques L. Rubin. Conformal proper times according to the Woodhouse causal axiomatics of relativistic spacetimes. Foundations of Physics, 40 (2):158–178, February 2010. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-009-9379-5.

Atkinson:2010:NEL

[3283] David Atkinson and Porter Johnson. Nonconservation of energy and loss of determinism II. Colliding with an open set. Foundations of Physics, 40(2):179–189, February 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10701-009-9384-8.pdf.

Afonin:2010:LLN

[3284] V. V. Afonin and V. Y. Petrov. Is the Luttinger liquid a new state of matter? Foundations of Physics, 40(2):190–204, February 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-009-9385-7.

Maroney:2010:DCA

[3285] Owen J. E. Maroney. Does a computer have an arrow of time? Foundations of Physics, 40(2):205–238, February 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-009-9386-6.

Piazza:2010:GPG

[3286] Federico Piazza. Glimmers of a pre-geometric perspective. Foundations of Physics, 40(3):239–266, March 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-009-9387-5.

Perez-Madrid:2010:SMT

[3287] A. Pérez-Madrid and I. Santamaría-Holek. Statistical mechanical theory of a closed oscillating Universe. Foundations of Physics, 40(3):267–275, March 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
009-9388-4.

Manchak:2010:PSG

[3288] John Byron Manchak. On the possibility of supertasks in general relativity. Foundations of Physics, 40(3):276–288, March 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-009-9390-x.

Schmelzer:2010:OPW

[3289] I. Schmelzer. Overlaps in pilot wave field theories. Foundations of Physics, 40(3):289–300, March 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-009-9394-6.

Herbut:2010:ETE

[3290] F. Herbut. On EPR-type entanglement in the experiments of Scully et al. II. Insight in the real random delayed-choice erasure experiment. Foundations of Physics, 40(3):301–312, March 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-009-9395-5.

Wharton:2010:NIK

[3291] K. B. Wharton. A novel interpretation of the Klein–Gordon equation. Foundations of Physics, 40(3):313–332, March 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-009-9398-2.

Rynasiewicz:2010:EI

[3292] R. A. Rynasiewicz. Editorial introduction. Foundations of Physics, 40(4): 333–334, April 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10701-010-9412-8.pdf.

Goldstein:2010:BMQ

[3293] Sheldon Goldstein. Bohmian mechanics and quantum information. Foundations of Physics, 40(4):335–355, April 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-009-9319-4.

Hiley:2010:RBW

[3294] B. J. Hiley. On the relationship between the Wigner–Moyal and Bohm approaches to quantum mechanics: a step to a more general theory? Foundations of Physics, 40(4):356–367, April 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-009-9320-y.

Demopoulos:2010:EP

[3295] William Demopoulos. Effects and propositions. Foundations of Physics, 40(4):368–389, April 2010. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-009-9321-x.

vanFraassen:2010:RW

[3296] Bas C. van Fraassen. Rovelli’s world. Foundations of Physics, 40 (4):390–417, April 2010. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-009-9326-5.

Kiessling:2010:MSA

[3297] Michael K.-H. Kiessling. Misleading signposts along the de Broglie–Bohm road to quantum mechanics. Foundations of Physics, 40(4):418–429, April 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
009-9327-4.

Conway:2010:TSC

[3298] John Conway and Simon Kochen. Thou shalt not clone one bit! Foundations of Physics, 40(4):430–433, April 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-010-9409-3.

Wilce:2010:FIQ

[3299] Alexander Wilce. Formalism and interpretation in quantum theory. Foundations of Physics, 40(4):434–462, April 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-010-9410-x.

Brassard:2010:CQM

[3300] Gilles Brassard and André Allan Méthot. Can quantum-mechanical description of physical reality be considered correct? Foundations of Physics, 40(4):463–468, April 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-010-9411-9.

Bogolubov:2010:ALH

[3301] Nikolai N. Bogolubov, Jr. and Anatoliy K. Prykarpatsky. The analysis of Lagrangian and Hamiltonian properties of the classical relativistic electrodynamics models and their quantization. Foundations of Physics, 40(5):469–493, May 2010. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-009-9399-1.

Fehr:2010:QC

[3302] Serge Fehr. Quantum cryptography. Foundations of Physics, 40(5): 494–531, May 2010. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-010-9408-4.

Hernandez-Zapata:2010:CNR

[3303] Sergio Hernández-Zapata and Ernesto Hernández-Zapata. Classical and non-relativistic limits of a Lorentz-invariant Bohmian model for a system of spinless particles. Foundations of Physics, 40(5):532–544, May 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-010-9413-7.

Kononets:2010:CCK

[3304] Y. V. Kononets. Charge conservation, Klein’s Paradox and the concept of paulions in the Dirac electron theory. Foundations of Physics, 40 (5):545–572, May 2010. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-010-9414-6.

Lampe:2010:EDT

[3305] Bodo Lampe. Erratum to: Development of the tetron model. Foundations of Physics, 40(5):573–574, May 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10701-010-9416-4.pdf.

Christov:2010:ERM

[3306] C. I. Christov. The effect of the relative motion of atoms on the frequency of the emitted light and the reinterpretation of the Ives–Stilwell experiment. Foundations of Physics, 40(6):575–584, June 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s10701-010-9418-2.pdf.

Surya:2010:QCQ

[3307] Sumati Surya and Petros Wallden. Quantum covers in quantum measure theory. Foundations of Physics, 40(6):585–606, June 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-010-9419-1.

Corbett:2010:SLQ

[3308] John Corbett and Thomas Durt. Spatial localization in quantum theory based on qr-numbers. Foundations of Physics, 40(6):607–628, June 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-010-9424-4.

McCabe:2010:NUU

[3309] Gordon McCabe. The non-unique Universe. Foundations of Physics, 40(6):629–637, June 2010. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-010-9425-3.

Cadoni:2010:HEE

[3310] Mariano Cadoni and Maurizio Melis. Holographic entanglement entropy of the BTZ black hole. Foundations of Physics, 40(6):638–657, June 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-010-9430-6.

Horton:2010:RHV

[3311] George Horton and Chris Dewdney. A relativistic hidden-variable interpretation for the massive vector field based on energy–momentum flows. Foundations of Physics, 40(6):658–678, June 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-010-9456-9.

Stachel:2010:LE

[3312] John Stachel. Letter to the Editor. Foundations of Physics, 40(6): 679, June 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10701-010-9420-8.pdf.

Milotti:2010:SIS

[3313] Edoardo Milotti, C. Curceanu (Petrascu), Stephen L. Adler, and Michael Berry. Special issue on spin statistics. Foundations of Physics, 40(7): 681–683, July 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10701-010-9421-7.pdf.

Solomon:2010:NOA

[3314] Allan I. Solomon. Normal ordering and abnormal nonsense. Foundations of Physics, 40(7):684–691, July 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-009-9325-6.

Balachandran:2010:CSG

[3315] A. P. Balachandran, Anosh Joseph, and Pramod Padmanabhan. Causality and statistics on the Groenewold–Moyal plane. Foundations of Physics, 40(7):692–702, July 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-009-9335-4.

Barabash:2010:ETP

[3316] A. S. Barabash. Experimental test of the Pauli Exclusion Principle. Foundations of Physics, 40(7):703–718, July 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-009-9336-3.

Doplicher:2010:SSF

[3317] Sergio Doplicher. Spin and statistics and first principles. Foundations of Physics, 40(7):719–732, July 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-009-9337-2.

Kienle:2010:OEC

[3318] Paul Kienle. Orbital electron capture rate and time modulation of H^{−} and He-like ions in a storage-ring. Foundations of Physics, 40(7):733–745, July 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
009-9343-4.

Marchetti:2010:SST

[3319] P. A. Marchetti. Spin-statistics transmutation in quantum field theory. Foundations of Physics, 40(7):746–764, July 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-009-9345-2.

Bartalucci:2010:VEL

[3320] S. Bartalucci, S. Bertolucci, M. Bragadireanu, M. Cargnelli, C. Curceanu (Petrascu), S. Di Matteo, J.-P. Egger, C. Guaraldo, M. Iliescu, T. Ishiwatari, M. Laubenstein, J. Marton, E. Milotti, D. Pietreanu, T. Ponta, A. Romero Vidal, D. L. Sirghi, F. Sirghi, L. Sperandio, O. Vazquez Doce, E. Widmann, and J. Zmeskal. The VIP experimental limit on the Pauli Exclusion Principle violation by electrons. Foundations of Physics, 40(7):765–775, July 2010. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-009-9346-1.

Jabs:2010:CSS

[3321] Arthur Jabs. Connecting spin and statistics in quantum mechanics. Foundations of Physics, 40(7):776–792, July 2010. CODEN FNDPA4.

ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-009-9351-4. See addendum

[3322].

Jabs:2010:ACS

[3322] Arthur Jabs. Addendum to: Connecting Spin and Statistics in Quantum Mechanics. DOI 10.1007/s10701-009-9351-4. Foundations of Physics, 40 (7):793–794, July 2010. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-009-9377-7. See [3321].

Gozzi:2010:CQM

[3323] E. Gozzi. Classical and quantum mechanics via supermetrics in time. Foundations of Physics, 40(7):795–806, July 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-009-9355-0.

Bernabei:2010:NPN

[3324] R. Bernabei, P. Belli, F. Cappella, R. Cerulli, C. J. Dai, A. d’Angelo, H. L. He, A. Incicchitti, H. H. Kuang, X. H. Ma, F. Montecchia, F. Nozzoli, D. Prosperi, X. D. Sheng, and Z. P. Ye. Non-Paulian nuclear processes in highly radiopure NaI(Tl): Status and perspectives. Foundations of Physics, 40(7):807–813, July 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-009-9362-1.

Lavagno:2010:TDB

[3325] A. Lavagno and P. Narayana Swamy. Thermostatistics of deformed bosons and fermions. Foundations of Physics, 40(7):814–828, July 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-009-9363-0.

Papadopoulos:2010:GBR

[3326] Nikolaos Papadopoulos and Andrés F. Reyes-Lega. On the geometry of the Berry–Robbins approach to spin-statistics. Foundations of Physics, 40(7):829–851, July 2010. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-009-9365-y.

DiDomenico:2010:CSQ

[3327] Antonio Di Domenico. CPT symmetry and quantum mechanics tests in the neutral kaon system at KLOE. Foundations of Physics, 40 (7):852–866, July 2010. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-009-9366-x.

Capozziello:2010:RAI

[3328] S. Capozziello and M. De Laurentis. A review about invariance induced gravity: Gravity and spin from local conformal–affine symmetry. Foundations of Physics, 40(7):867–899, July 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-009-9367-9.

Bernabei:2010:RDL

[3329] R. Bernabei, P. Belli, F. Cappella, R. Cerulli, C. J. Dai, A. d’Angelo, H. L. He, A. Incicchitti, H. H. Kuang, X. H. Ma, F. Montecchia, F. Nozzoli, D. Prosperi, X. D. Sheng, and Z. P. Ye. Results from DAMA/LIBRA at Gran Sasso. Foundations of Physics, 40(7):900–916, July 2010. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-009-9368-8.

Mavromatos:2010:DCV

[3330] Nick E. Mavromatos. Decoherence and CPT violation in a stringy model of space–time foam. Foundations of Physics, 40(7):917–960, July 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-009-9372-z.

Galehouse:2010:PEP

[3331] Daniel C. Galehouse. Pauli’s exclusion principle in spinor coordinate space. Foundations of Physics, 40(7):961–977, July 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-009-9374-x.

Sarkar:2010:IST

[3332] Sarben Sarkar. Implications of space–time foam for entanglement correlations of neutral kaons. Foundations of Physics, 40(7):978–1003, July 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
009-9389-3.

Reyes-Lega:2010:RCS

[3333] Andrés F. Reyes-Lega and Carlos Benavides. Remarks on the configuration space approach to spin-statistics. Foundations of Physics, 40 (7):1004–1029, July 2010. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-009-9397-3.

Bressan:2010:SSN

[3334] Andrea Bressan. The spin structure of the nucleon; a COMPASS perspective. Foundations of Physics, 40(7):1030–1041, July 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-009-9405-7.

Selyugin:2010:GPE

[3335] O. V. Selyugin and O. V. Teryaev. A gravitational potential with extra-dimensions and spin effects in hadronic reactions. Foundations of Physics, 40(7):1042–1050, July 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-010-9437-z.

Khrennikov:2010:DCQ

[3336] Andrei Khrennikov. Description of composite quantum systems by means of classical random fields. Foundations of Physics, 40(8):1051–1064, August 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
009-9392-8.

Bohata:2010:BCS

[3337] Martin Bohata and Jan Hamhalter. Bell’s correlations and spin systems. Foundations of Physics, 40(8):1065–1075, August 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-009-9401-y.

Jacobson:2010:BHT

[3338] Ted Jacobson and Aron C. Wall. Black hole thermodynamics and Lorentz symmetry. Foundations of Physics, 40(8):1076–1080, August 2010. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-010-9423-5.

Buric:2010:PTC

[3339] Nikola Burić. Probability in theories with complex dynamics and Hardy’s fifth axiom. Foundations of Physics, 40(8):1081–1087, August 2010. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-010-9426-2.

Alhaidari:2010:DEC

[3340] A. D. Alhaidari. Dirac equation with coupling to 1/r singular vector potentials for all angular momenta. Foundations of Physics, 40(8): 1088–1095, August 2010. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-010-9431-5.

Gould:2010:IBP

[3341] Elizabeth Gould and P. K. Aravind. Isomorphism between the Peres and Penrose proofs of the BKS theorem in three dimensions. Foundations of Physics, 40(8):1096–1101, August 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-010-9434-2.

Boyer:2010:BRS

[3342] Timothy H. Boyer. Blackbody radiation and the scaling symmetry of relativistic classical electron theory with classical electromagnetic zeropoint radiation. Foundations of Physics, 40(8):1102–1116, August 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-010-9436-0.

Budiyono:2010:QCT

[3343] Agung Budiyono. On quantum–classical transition of a single particle. Foundations of Physics, 40(8):1117–1133, August 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-010-9438-y.

Sidharth:2010:FWP

[3344] B. G. Sidharth. The Feynman–Wheeler perfect absorber theory in a new light. Foundations of Physics, 40(8):1134–1140, August 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-010-9444-0.

Vedral:2010:ESQ

[3345] Vlatko Vedral. The elusive source of quantum speedup. Foundations of Physics, 40(8):1141–1154, August 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-010-9452-0.

Wronski:2010:OCR

[3346] Leszek Wroński and Michal Marczyk. Only countable Reichenbachian common cause systems exist. Foundations of Physics, 40(8):1155– 1160, August 2010. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-010-9457-8.

Busch:2010:BPP

[3347] Paul Busch. Between physics and philosophy — Festschrift for Peter Mittelstaedt on his 80th birthday. Foundations of Physics, 40(9–10):1161– 1162, October 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10701-010-9468-5.pdf.

Busch:2010:QMS

[3348] Paul Busch, Joachim Pfarr, Manfred L. Ristig, and Ernst-Walther Stachow. Quantum-matter-spacetime: Peter Mittelstaedt’s contributions to physics and its foundations. Foundations of Physics, 40(9–10):1163– 1170, October 2010. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-010-9478-3.

Falkenburg:2010:LR

[3349] Brigitte Falkenburg. Language and reality. Foundations of Physics, 40(9–10):1171–1188, October 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-010-9449-8.

Busch:2010:PML

[3350] Paul Busch. Peter Mittelstaedt: List of publications until 2010. Foundations of Physics, 40(9–10):1189–1199, October 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-010-9466-7.

Clark:2010:AEQ

[3351] John W. Clark, Hessam Habibian, Aikaterini D. Mandilara, and Manfred L. Ristig. Aspects of entanglement in quantum many-body systems. Foundations of Physics, 40(9–10):1200–1220, October 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-010-9467-6.

Nimtz:2010:EVP

[3352] Günter Nimtz. Erratum to: On virtual phonons, photons, and electrons. Foundations of Physics, 40(9–10):1221–1230, October 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s10701-010-9503-6.pdf.

Nitsch:2010:HFM

[3353] Jürgen Nitsch and Sergey Tkachenko. High-frequency multiconductor transmission–line theory. Foundations of Physics, 40(9–10):1231– 1252, October 2010. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-010-9443-1.

Ristig:2010:SPS

[3354] M. L. Ristig and K. A. Gernoth. Statistical properties of strongly correlated quantum liquids. Foundations of Physics, 40(9–10):1253–1262, October 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
009-9370-1.

Barbour:2010:DMP

[3355] Julian Barbour. The definition of Mach’s Principle. Foundations of Physics, 40(9–10):1263–1284, October 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-010-9490-7.

Grunbaum:2010:DMC

[3356] Adolf Grünbaum. David Malament and the conventionality of simultaneity: a reply. Foundations of Physics, 40(9–10):1285–1297, October 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-009-9328-3.

Lazar:2010:CSS

[3357] Markus Lazar and Friedrich W. Hehl. Cartan’s spiral staircase in physics and, in particular, in the gauge theory of dislocations. Foundations of Physics, 40(9–10):1298–1325, October 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-010-9440-4.

Pfarr:2010:CTC

[3358] Joachim Pfarr. Closed timelike curves — time and again. Foundations of Physics, 40(9–10):1326–1332, October 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-010-9448-9.

Bub:2010:NNH

[3359] Jeffrey Bub. Von Neumann’s ‘no hidden variables’ proof: a re-appraisal. Foundations of Physics, 40(9–10):1333–1340, October 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-010-9480-9.

Busch:2010:UQR

[3360] Paul Busch and Gregg Jaeger. Unsharp quantum reality. Foundations of Physics, 40(9–10):1341–1367, October 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-010-9497-0.

Brunetti:2010:TQP

[3361] Romeo Brunetti, Klaus Fredenhagen, and Marc Hoge. Time in quantum physics: From an external parameter to an intrinsic observable. Foundations of Physics, 40(9–10):1368–1378, October 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-009-9400-z.

Ghirardi:2010:DQN

[3362] GianCarlo Ghirardi. Does quantum nonlocality irremediably conflict with special relativity? Foundations of Physics, 40(9–10):1379–1395, October 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
009-9391-9.

Jaeger:2010:IQM

[3363] Gregg Jaeger. Individuation in quantum mechanics and space–time. Foundations of Physics, 40(9–10):1396–1409, October 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-010-9489-0.

Kiefer:2010:CQT

[3364] Claus Kiefer. Can quantum theory be applied to the Universe as a whole? Foundations of Physics, 40(9–10):1410–1418, October 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-010-9441-3.

Lahti:2010:CQO

[3365] Pekka Lahti and Juha-Pekka Pellonpää. On the complementarity of the quadrature observables. Foundations of Physics, 40(9–10):1419– 1428, October 2010. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-009-9373-y.

Lyre:2010:WQT

[3366] Holger Lyre. Why quantum theory is possibly wrong. Foundations of Physics, 40(9–10):1429–1438, October 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-010-9463-x.

Redei:2010:OIO

[3367] Miklós Rédei. Operational independence and operational separability in algebraic quantum mechanics. Foundations of Physics, 40(9–10): 1439–1449, October 2010. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-010-9447-x.

Stachow:2010:OVL

[3368] Ernst-Walther Stachow. Objectivity vs. locality in quantum physics. Foundations of Physics, 40(9–10):1450–1475, October 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-009-9393-7.

Zeh:2010:QDI

[3369] H. Dieter Zeh. Quantum discreteness is an illusion. Foundations of Physics, 40(9–10):1476–1493, October 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-009-9383-9.

DallaChiara:2010:ESR

[3370] Maria Luisa Dalla Chiara, Roberto Giuntini, Antonio Ledda, Roberto Leporini, and Giuseppe Sergioli. Entanglement as a semantic resource. Foundations of Physics, 40(9–10):1494–1518, October 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-010-9407-5.

Dvurecenskij:2010:SPB

[3371] A. Dvurecenskij, R. Giuntini, and T. Kowalski. On the structure of pseudo BL-algebras and pseudo hoops in quantum logics. Foundations of Physics, 40(9–10):1519–1542, October 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-009-9342-5.

Foulis:2010:TDE

[3372] David J. Foulis and Sylvia Pulmannová. Type-decomposition of an effect algebra. Foundations of Physics, 40(9–10):1543–1565, October 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-009-9344-3.

Gudder:2010:EAA

[3373] Stan Gudder. Effect algebras are not adequate models for quantum mechanics. Foundations of Physics, 40(9–10):1566–1577, October 2010. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-009-9369-7.

Weingartner:2010:BLA

[3374] Paul Weingartner. Basis logic for application in physics and its intuitionistic alternative. Foundations of Physics, 40(9–10):1578–1596, October 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
009-9406-6.

Esfeld:2010:PC

[3375] Michael Esfeld. Physics and causation. Foundations of Physics, 40(9–10): 1597–1610, October 2010. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-009-9357-y.

Jaeger:2010:CLN

[3376] Lydia Jaeger. The contingency of laws of nature in science and theology. Foundations of Physics, 40(9–10):1611–1624, October 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-010-9470-y.

Kuhlmann:2010:WCR

[3377] Meinard Kuhlmann. Why conceptual rigour matters to philosophy: on the ontological significance of algebraic quantum field theory. Foundations of Physics, 40(9–10):1625–1637, October 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-010-9474-7.

Mainzer:2010:ETS

[3378] Klaus Mainzer. The emergence of temporal structures in dynamical systems. Foundations of Physics, 40(9–10):1638–1650, October 2010. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-010-9451-1.

Vollmer:2010:IO

[3379] Gerhard Vollmer. Invariance and objectivity. Foundations of Physics, 40(9–10):1651–1667, October 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-010-9471-x.

DHooghe:2010:POM

[3380] Bart D’Hooghe. A possible operational motivation for the orthocomplementation in quantum structures. Foundations of Physics, 40(11): 1669–1680, November 2010. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-010-9464-9.

Brannen:2010:SPI

[3381] Carl Brannen. Spin path integrals and generations. Foundations of Physics, 40(11):1681–1699, November 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-010-9465-8.

Struyve:2010:ETM

[3382] Ward Struyve. On Epstein’s trajectory model of non-relativistic quantum mechanics. Foundations of Physics, 40(11):1700–1711, November 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-010-9475-6.

Castro:2010:NQM

[3383] Carlos Castro. On nonlinear quantum mechanics, noncommutative phase spaces, fractal-scale calculus and vacuum energy. Foundations of Physics, 40(11):1712–1730, November 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-010-9476-5.

Flores:2010:CPS

[3384] E. V. Flores and J. M. De Tata. Complementarity Paradox solved: Surprising consequences. Foundations of Physics, 40(11):1731–1743, November 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
010-9477-4.

Machicote:2010:TGC

[3385] Juan Eduardo Reluz Machicote. Time as a geometric concept involving angular relations in classical mechanics and quantum mechanics. Foundations of Physics, 40(11):1744–1778, November 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-010-9479-2.

Ribeiro:2010:WSG

[3386] José Edmar Arantes Ribeiro. Was the Stern–Gerlach phenomenon classically described? Foundations of Physics, 40(11):1779–1782, November 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
010-9486-3.

Kleemans:2010:BRK

[3387] Machiel Kleemans. Book review: Kristian Camilleri: Heisenberg and the Interpretation of Quantum Mechanics — The Physicist as Philosopher. Foundations of Physics, 40(11):1783–1787, November 2010. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-010-9459-6.

Singh:2010:SNG

[3388] Dinesh Singh, Nader Mobed, and Pierre-Philippe Ouimet. Signatures of noncommutative geometry in muon decay for nonsymmetric gravity. Foundations of Physics, 40(12):1789–1799, December 2010. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-010-9487-2.

Schroer:2010:CLY

[3389] Bert Schroer. A critical look at 50 years particle theory from the perspective of the crossing property. Foundations of Physics, 40(12):1800– 1857, December 2010. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-010-9492-5.

Norsen:2010:TEL

[3390] Travis Norsen. The theory of (exclusively) local beables. Foundations of Physics, 40(12):1858–1884, December 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-010-9495-2.

Falciano:2010:GRQ

[3391] F. T. Falciano, M. Novello, and J. M. Salim. Geometrizing relativistic quantum mechanics. Foundations of Physics, 40(12):1885–1901, December 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
010-9496-1.

Laraudogoitia:2010:FAA

[3392] Jon Perez Laraudogoitia. A flawed argument against actual infinity in physics. Foundations of Physics, 40(12):1902–1910, December 2010. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-010-9498-z.

Horwitz:2011:PIP

[3393] L. P. Horwitz. Preface IARD 2008 Proceedings. Foundations of Physics, 41(1):1–3, January 2011. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s10701-010-9491-6.pdf.

Fanchi:2011:MCQ

[3394] John R. Fanchi. Manifestly covariant quantum theory with invariant evolution parameter in relativistic dynamics. Foundations of Physics, 41(1):4–32, January 2011. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-009-9371-0.

Rapoport:2011:SCC

[3395] Diego L. Rapoport. Surmounting the Cartesian cut through philosophy, physics, logic, cybernetics, and geometry: Self-reference, torsion, the Klein bottle, the time operator, multivalued logics and quantum mechanics. Foundations of Physics, 41(1):33–76, January 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-009-9334-5.

Hsiang:2011:BMC

[3396] Jen-Tsung Hsiang, Tai-Hung Wu, and Da-Shin Lee. Brownian motion of a charged particle in electromagnetic fluctuations at finite temperature. Foundations of Physics, 41(1):77–87, January 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-009-9333-6.

Gersten:2011:TLL

[3397] Alexander Gersten. Tensor Lagrangians, Lagrangians equivalent to the Hamilton–Jacobi equation and relativistic dynamics. Foundations of Physics, 41(1):88–98, January 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-009-9352-3.

Gill:2011:TME

[3398] Tepper L. Gill and Woodford W. Zachary. Two mathematically equivalent versions of Maxwell’s equations. Foundations of Physics, 41(1): 99–128, January 2011. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-009-9331-8.

Kryukov:2011:GUQ

[3399] A. Kryukov. Geometry of the unification of quantum mechanics and relativity of a single particle. Foundations of Physics, 41(1):129–140, January 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
009-9354-1.

Horwitz:2011:HMC

[3400] Lawrence Horwitz, Avi Gershon, and Marcelo Schiffer. Hamiltonian map to conformal modification of spacetime metric: Kaluza–Klein and TeVeS. Foundations of Physics, 41(1):141–157, January 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-010-9483-6.

Schmelzer:2011:PQI

[3401] I. Schmelzer. Pure quantum interpretations are not viable. Foundations of Physics, 41(2):159–177, February 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-010-9484-5.

Dolce:2011:CTD

[3402] Donatello Dolce. Compact time and determinism for bosons: Foundations. Foundations of Physics, 41(2):178–203, February 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-010-9485-4.

Schroer:2011:BMS

[3403] Bert Schroer. Bondi–Metzner–Sachs symmetry, holography on nullsurfaces and area proportionality of “light-slice” entropy. Foundations of Physics, 41(2):204–241, February 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-010-9494-3.

Babin:2011:EBC

[3404] Anatoli Babin and Alexander Figotin. Electrodynamics of balanced charges. Foundations of Physics, 41(2):242–260, February 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s10701-010-9502-7.pdf.

Opatrny:2011:EMD

[3405] Tomás Opatrný and Zuzana Misáková. Exorcising Maxwell’s demon from Liboff’s three-channel conundrum. Foundations of Physics, 41(2):

261–269, February 2011. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-010-9505-4.

Faizal:2011:BAB

[3406] Mir Faizal. BRST and anti–BRST symmetries in perturbative quantum gravity. Foundations of Physics, 41(2):270–277, February 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-010-9511-6.

Kennefick:2011:REU

[3407] Daniel Kennefick. Review of Einstein’s Unification by Jeroen van Dongen. Foundations of Physics, 41(2):278–280, February 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-010-9532-1.

Bengtsson:2011:P

[3408] Ingemar Bengtsson and Andrei Khrennikov. Preface. Foundations of Physics, 41(3):281, March 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10701-010-9524-1.pdf.

Curceanu:2011:NEL

[3409] C. Curceanu (Petrascu), S. Bartalucci, S. Bertolucci, M. Bragadireanu, M. Cargnelli, S. Di Matteo, J.-P. Egger, C. Guaraldo, M. Iliescu, T. Ishiwatari, M. Laubenstein, J. Marton, E. Milotti, D. Pietreanu, T. Ponta, A. Romero Vidal, D. L. Sirghi, F. Sirghi, L. Sperandio, O. Vazquez Doce, E. Widmann, and J. Zmeskal. New experimental limit on the Pauli Exclusion Principle violation by electrons — the VIP experiment. Foundations of Physics, 41(3):282–287, March 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-009-9378-6.

Uhlmann:2011:TPF

[3410] Armin Uhlmann. Transition probability (fidelity) and its relatives. Foundations of Physics, 41(3):288–298, March 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-009-9381-y.

Jaeger:2011:IQM

[3411] Gregg Jaeger. Individuation in quantum mechanics. Foundations of Physics, 41(3):299–304, March 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-009-9382-x.

Brida:2011:DBC

[3412] Giorgio Brida, Maria Bondani, Ivo P. Degiovanni, Marco Genovese, Matteo G. A. Paris, Ivano Ruo Berchera, and Valentina Schettini. On the discrimination between classical and quantum states. Foundations of Physics, 41(3):305–316, March 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-009-9396-4.

Khrennikov:2011:PCS

[3413] Andrei Khrennikov. Prequantum classical statistical field theory: Schrödinger dynamics of entangled systems as a classical stochastic process. Foundations of Physics, 41(3):317–329, March 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-009-9402-x.

Manko:2011:PDE

[3414] Margarita A. Man’ko and Vladimir I. Man’ko. Probability description and entropy of classical and quantum systems. Foundations of Physics, 41(3):330–344, March 2011. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-009-9403-9.

Fuchs:2011:QBR

[3415] Christopher A. Fuchs and Rüdiger Schack. A quantum-Bayesian route to quantum-state space. Foundations of Physics, 41(3):345–356, March 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
009-9404-8.

Marchildon:2011:CEI

[3416] Louis Marchildon. Can Everett be interpreted without extravaganza? Foundations of Physics, 41(3):357–362, March 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-010-9415-5.

DeMartini:2011:EQS

[3417] Francesco De Martini. Entanglement and quantum superposition of a macroscopic–macroscopic system. Foundations of Physics, 41(3): 363–370, March 2011. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-010-9417-3.

Craig:2011:CHQ

[3418] David Craig and Parampreet Singh. Consistent histories in quantum cosmology. Foundations of Physics, 41(3):371–379, March 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-010-9422-6.

Rau:2011:MBQ

[3419] Jochen Rau. Measurement-based quantum foundations. Foundations of Physics, 41(3):380–388, March 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-010-9427-1.

Baladron:2011:SPD

[3420] C. Baladrón. Study on a possible Darwinian origin of quantum mechanics. Foundations of Physics, 41(3):389–395, March 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-010-9428-0.

Ududec:2011:TSE

[3421] Cozmin Ududec, Howard Barnum, and Joseph Emerson. Three slit experiments and the structure of quantum theory. Foundations of Physics, 41(3):396–405, March 2011. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-010-9429-z.

Iriyama:2011:QME

[3422] Satoshi Iriyama and Masanori Ohya. Quantum mutual entropy defined by liftings. Foundations of Physics, 41(3):406–413, March 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-010-9432-4.

Badziag:2011:PP

[3423] Piotr Badziag, Ingemar Bengtsson, Adán Cabello, Helena Granström, and Jan-Åke Larsson. Pentagrams and paradoxes. Foundations of Physics, 41(3):414–423, March 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-010-9433-3.

Garola:2011:GOB

[3424] Claudio Garola and Sandro Sozzo. Generalized observables, Bell’s inequalities and mixtures in the ESR model for QM. Foundations of Physics, 41(3):424–449, March 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-010-9435-1.

Gustafson:2011:TQS

[3425] Karl Gustafson. Trigonometry of quantum states. Foundations of Physics, 41(3):450–465, March 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-010-9439-x.

Plotnitsky:2011:RUE

[3426] Arkady Plotnitsky. On the reasonable and unreasonable effectiveness of mathematics in classical and quantum physics. Foundations of Physics, 41(3):466–491, March 2011. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-010-9442-2.

DeMartini:2011:GHR

[3427] Francesco De Martini, Fabio Sciarrino, Nicolò Spagnolo, and Chiara Vitelli. Generation of highly resilient to decoherence macroscopic quantum superpositions via phase-covariant quantum cloning. Foundations of Physics, 41(3):492–508, March 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-010-9445-z.

Gevorkyan:2011:NQM

[3428] Ashot S. Gevorkyan. Nonrelativistic quantum mechanics with fundamental environment. Foundations of Physics, 41(3):509–515, March 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-010-9446-y.

Volovich:2011:RCM

[3429] Igor V. Volovich. Randomness in classical mechanics and quantum mechanics. Foundations of Physics, 41(3):516–528, March 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-010-9450-2.

Haven:2011:ILQ

[3430] Emmanuel Haven. Itô’s lemma with quantum calculus (q-calculus): Some implications. Foundations of Physics, 41(3):529–537, March 2011. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-010-9453-z.

Asano:2011:QLM

[3431] Masanari Asano, Masanori Ohya, and Andrei Khrennikov. Quantumlike model for decision making process in two players game. Foundations of Physics, 41(3):538–548, March 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-010-9454-y.

Watanabe:2011:NEQ

[3432] Noboru Watanabe. Note on entropies of quantum dynamical systems. Foundations of Physics, 41(3):549–563, March 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-010-9455-x.

Appleby:2011:PQS

[3433] D. M. Appleby, Åsa Ericsson, and Christopher A. Fuchs. Properties of QBist state spaces. Foundations of Physics, 41(3):564–579, March 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-010-9458-7.

Nieuwenhuizen:2011:CLF

[3434] T. M. Nieuwenhuizen. Is the contextuality loophole fatal for the derivation of Bell inequalities? Foundations of Physics, 41(3):580–591, March 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
010-9461-z;   http://link.springer.com/content/pdf/10.1007/
s10701-010-9461-z.pdf.

Ozawa:2011:QRM

[3435] Masanao Ozawa. Quantum reality and measurement: a quantum logical approach. Foundations of Physics, 41(3):592–607, March 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-010-9462-y.

Nielsen:2011:IPA

[3436] H. B. Nielsen. Imaginary part of action, future functioning as hidden variables. Foundations of Physics, 41(3):608–635, March 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-010-9469-4.

Ghirardi:2011:VVD

[3437] GianCarlo Ghirardi. Vedral, V.: Decoding reality: the Universe as quantum information. Foundations of Physics, 41(4):637–639, April 2011.

CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-011-9541-8.

Hajicek:2011:QMP

[3438] P. Hájícek. The quantum measurement problem and cluster separability. Foundations of Physics, 41(4):640–666, April 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-010-9506-3.

Hurley:2011:DHS

[3439] D. J. Hurley and M. A. Vandyck. D-differentiation in Hilbert space and the structure of quantum mechanics Part II: Accelerated observers and fictitious forces. Foundations of Physics, 41(4):667–685, April 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-010-9509-0.

Bedingham:2011:RSR

[3440] Daniel J. Bedingham. Relativistic state reduction dynamics. Foundations of Physics, 41(4):686–704, April 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-010-9510-7.

Griffiths:2011:QL

[3441] Robert B. Griffiths. Quantum locality. Foundations of Physics, 41 (4):705–733, April 2011. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-010-9512-5.

Paraoanu:2011:RSQ

[3442] G. S. Paraoanu. Realism and single–quanta nonlocality. Foundations of Physics, 41(4):734–743, April 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-010-9513-4.

Lee:2011:QME

[3443] Jae-Weon Lee. Quantum mechanics emerges from information theory applied to causal horizons. Foundations of Physics, 41(4):744–753, April 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
010-9514-3.

Hofer:2011:UAO

[3444] Werner A. Hofer. Unconventional approach to orbital-free density functional theory derived from a model of extended electrons. Foundations of Physics, 41(4):754–791, April 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-010-9517-0.

Liu:2011:TEM

[3445] Cheng-Shi Liu. Trial equation method based on symmetry and applications to nonlinear equations arising in mathematical physics. Foundations of Physics, 41(5):793–804, May 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-010-9521-4.

Kauffmann:2011:UQM

[3446] Steven Kenneth Kauffmann. Unambiguous quantization from the maximum classical correspondence that is self-consistent: The slightly stronger canonical commutation rule Dirac missed. Foundations of Physics, 41(5):805–819, May 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-010-9523-2.

Kurkov:2011:LFR

[3447] M. A. Kurkov and V. A. Franke. Local fields without restrictions on the spectrum of 4-momentum operator and relativistic Lindblad equation. Foundations of Physics, 41(5):820–842, May 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-010-9525-0.

Valdes-Hernandez:2011:BEI

[3448] A. Valdés-Hernández, L. de la Peña, and A. M. Cetto. Bipartite entanglement induced by a common background (zero-point) radiation field. Foundations of Physics, 41(5):843–862, May 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-010-9527-y.

Hamada:2011:VOQ

[3449] Ken ji Hamada. Vertex operators in 4D quantum gravity formulated as CFT. Foundations of Physics, 41(5):863–882, May 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-010-9533-0.

Waegell:2011:PPB

[3450] Mordecai Waegell, P. K. Aravind, Norman D. Megill, and Mladen Pavicić. Parity proofs of the Bell–Kochen–Specker Theorem based on the 600-cell. Foundations of Physics, 41(5):883–904, May 2011. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-011-9534-7.

Heller:2011:FPU

[3451] Michael Heller, Leszek Pysiak, and Wiesław Sasin. Fundamental problems in the unification of physics. Foundations of Physics, 41(5):905–918, May 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
011-9535-6.

Anonymous:2011:ERE

[3452] Anonymous. Editorial: Reduction, emergence, and physics. Foundations of Physics, 41(6):919, June 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10701-011-9560-5.pdf.

Butterfield:2011:ERS

[3453] J. Butterfield. Emergence, reduction and supervenience: a varied landscape. Foundations of Physics, 41(6):920–959, June 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-011-9549-0.

Callender:2011:HHM

[3454] Craig Callender. Hot and heavy matters in the foundations of statistical mechanics. Foundations of Physics, 41(6):960–981, June 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s10701-010-9518-z.pdf.

Hartle:2011:QRQ

[3455] James B. Hartle. The quasiclassical realms of this quantum Universe. Foundations of Physics, 41(6):982–1006, June 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10701-010-9460-0.pdf.

Healey:2011:REB

[3456] Richard Healey. Reduction and emergence in Bose–Einstein condensates. Foundations of Physics, 41(6):1007–1030, June 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-010-9481-8.

Batterman:2011:ESS

[3457] Robert W. Batterman. Emergence, singularities, and symmetry breaking. Foundations of Physics, 41(6):1031–1050, June 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-010-9493-4.

Dieks:2011:HCP

[3458] Dennis Dieks and Andrea Lubberdink. How classical particles emerge from the quantum world. Foundations of Physics, 41(6):1051–1064, June 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10701-010-9515-2.pdf.

Butterfield:2011:LDE

[3459] J. Butterfield. Less is different: Emergence and reduction reconciled. Foundations of Physics, 41(6):1065–1135, June 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-010-9516-1.

Kastner:2011:QNE

[3460] Ruth E. Kastner. Quantum nonlocality: Not eliminated by the Heisenberg Picture. Foundations of Physics, 41(7):1137–1142, July 2011. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-011-9536-5.

Dvurecenskij:2011:SPE

[3461] Anatolij Dvurecenskij. States on pseudo effect algebras and integrals. Foundations of Physics, 41(7):1143–1162, July 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-011-9537-4.

Seidewitz:2011:CHS

[3462] Ed Seidewitz. Consistent histories of systems and measurements in spacetime. Foundations of Physics, 41(7):1163–1192, July 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-011-9538-3.

Nimtz:2011:TCS

[3463] Günter Nimtz. Tunneling confronts Special Relativity. Foundations of Physics, 41(7):1193–1199, July 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-011-9539-2;   https://springerlink3.
metapress.com/content/n67286542k5r4186.

Rosado:2011:RQS

[3464] José Ignacio Rosado. Representation of quantum states as points in a probability simplex associated to a SIC–POVM. Foundations of Physics, 41(7):1200–1213, July 2011. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-011-9540-9.

Paraoanu:2011:PMR

[3465] G. S. Paraoanu. Partial measurements and the realization of quantummechanical counterfactuals. Foundations of Physics, 41(7):1214–1235, July 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
011-9542-7.

Rubin:2011:OLE

[3466] Mark A. Rubin. Observers and locality in Everett quantum field theory. Foundations of Physics, 41(7):1236–1262, July 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-011-9543-6.

Nistico:2011:QMC

[3467] Giuseppe Nisticò and Angela Sestito. Quantum mechanics, can it be consistent with locality? Foundations of Physics, 41(7):1263–1278, July 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
011-9547-2.

Akhmedov:2011:NOE

[3468] E. K. Akhmedov and A. Y. Smirnov. Neutrino oscillations: Entanglement, energy–momentum conservation and QFT. Foundations of Physics, 41(8):1279–1306, August 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-011-9545-4.

Benedictus:2011:JRL

[3469] Fedde Benedictus. Joe Rosen: Lawless Universe. Foundations of Physics, 41(8):1307–1308, August 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10701-011-9546-3.pdf.

Ghirardi:2011:RPN

[3470] Gian Carlo Ghirardi. On a recent proof of nonlocality without inequalities. Foundations of Physics, 41(8):1309–1317, August 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-011-9548-1.

Rave:2011:QIM

[3471] M. J. Rave. Quantum interference and many worlds: a new family of classical analogies. Foundations of Physics, 41(8):1318–1330, August 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-011-9550-7.

Romannikov:2011:EPR

[3472] A. Romannikov. Ehrenfest’s Paradox and radial electric field in quasineutral tokamak plasma. Foundations of Physics, 41(8):1331–1337, August 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
011-9551-6.

Arenhart:2011:DFQ

[3473] Jonas Rafael Becker Arenhart. A discussion on finite quasi-cardinals in quasi-set theory. Foundations of Physics, 41(8):1338–1354, August 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-011-9552-5.

Zahiri-Abyaneh:2011:ZEM

[3474] Mehran Zahiri-Abyaneh and Mehrdad Farhoudi. Zitterbewegung in external magnetic field: Classic versus quantum approach. Foundations of Physics, 41(8):1355–1374, August 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-011-9553-4.

Capria:2011:SDA

[3475] Marco Mamone Capria. Spatial directions, anisotropy and special relativity. Foundations of Physics, 41(8):1375–1397, August 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-011-9554-3.

Hofer-Szabo:2011:BID

[3476] Gábor Hofer-Szabó. Bell (δ) inequalities derived from separate common causal explanation of almost perfect EPR anticorrelations. Foundations of Physics, 41(8):1398–1413, August 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-011-9555-2.

deGosson:2011:IQW

[3477] Maurice A. de Gosson and Basil J. Hiley. Imprints of the quantum world in classical mechanics. Foundations of Physics, 41(9):1415–1436, September 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
011-9544-5.

Grossing:2011:CEQ

[3478] Gerhard Grössing, Johannes Mesa Pascasio, and Herbert Schwabl. A classical explanation of quantization. Foundations of Physics, 41(9): 1437–1453, September 2011. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-011-9556-1.

Mitra:2011:MFF

[3479] Abhas Mitra. Macroscopic form of the first law of thermodynamics for an adiabatically evolving non-singular self-gravitating fluid. Foundations of Physics, 41(9):1454–1461, September 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-011-9557-0.

Gaveau:2011:HPI

[3480] B. Gaveau, A. M. Nounou, and L. S. Schulman. Homotopy and path integrals in the time-dependent Aharonov–Bohm effect. Foundations of Physics, 41(9):1462–1474, September 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-011-9559-y.

Rovelli:2011:FT

[3481] Carlo Rovelli. “Forget time”. Foundations of Physics, 41(9):1475– 1490, September 2011. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-011-9561-4.

Benedictus:2011:BRD

[3482] Fedde Benedictus. Book review: D. E. Neuenschwander: Emmy Noether’s Wonderful Theorem. Foundations of Physics, 41(9):1491–1492, September 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10701-011-9562-3.pdf.

Fritz:2011:QAH

[3483] Tobias Fritz. Quantum analogues of Hardy’s nonlocality paradox. Foundations of Physics, 41(9):1493–1501, September 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-011-9563-2.

Chuprikov:2011:CSD

[3484] Nikolay L. Chuprikov. From a 1D completed scattering and double slit diffraction to the quantum–classical problem for isolated systems. Foundations of Physics, 41(9):1502–1520, September 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-011-9564-1.

Hossenfelder:2011:TSD

[3485] Sabine Hossenfelder. Testing super–deterministic hidden variables theories. Foundations of Physics, 41(9):1521–1531, September 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-011-9565-0.

Barcelo:2011:QNG

[3486] C. Barceló, L. J. Garay, and G. Jannes. Quantum non-gravity and stellar collapse. Foundations of Physics, 41(9):1532–1541, September 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-011-9577-9.

Schroer:2011:AGT

[3487] Bert Schroer. An alternative to the gauge theoretic setting. Foundations of Physics, 41(10):1543–1568, October 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-011-9567-y.

Vaccaro:2011:VUT

[3488] Joan A. Vaccaro. T violation and the unidirectionality of time. Foundations of Physics, 41(10):1569–1596, October 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-011-9568-x.

Mantz:2011:RCS

[3489] Christiaan L. M. Mantz and Tomislav Prokopec. Resolving curvature singularities in holomorphic gravity. Foundations of Physics, 41(10):1597– 1633, October 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10701-011-9570-3.pdf.

Paseka:2011:CSL

[3490] Jan Paseka and Zdenka Riecanová. Considerable sets of linear operators in Hilbert spaces as operator generalized effect algebras. Foundations of Physics, 41(10):1634–1647, October 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-011-9573-0.

Simpson:2011:LSR

[3491] Michael Simpson. Larc: a state reduction theory of quantum measurement. Foundations of Physics, 41(10):1648–1663, October 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-011-9574-z.

Arecchi:2011:F

[3492] F. Tito Arecchi. Foreword. Foundations of Physics, 41(11):1665– 1666, November 2011. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-011-9572-1.

Mittelstaedt:2011:PIM

[3493] Peter Mittelstaedt. The problem of interpretation of modern physics. Foundations of Physics, 41(11):1667–1676, November 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-011-9571-2.

Pauri:2011:EPV

[3494] Massimo Pauri. Epistemic primacy vs. ontological elusiveness of spatial extension: Is there an evolutionary role for the quantum? Foundations of Physics, 41(11):1677–1702, November 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-011-9581-0.

dEspagnat:2011:QPR

[3495] Bernard d’Espagnat. Quantum physics and reality. Foundations of Physics, 41(11):1703–1716, November 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-011-9582-z.

Torretti:2011:HPC

[3496] Roberto Torretti. How percepts and concepts engage the future. Foundations of Physics, 41(11):1717–1728, November 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-011-9584-x.

Omnes:2011:WUE

[3497] Roland Omnès. Wigner’s “Unreasonable Effectiveness of Mathematics”, revisited. Foundations of Physics, 41(11):1729–1739, November 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-011-9587-7.

Calosi:2011:QOE

[3498] Claudio Calosi, Vincenzo Fano, and Gino Tarozzi. Quantum ontology and extensional mereology. Foundations of Physics, 41(11):1740– 1755, November 2011. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-011-9590-z.

Gillis:2011:CME

[3499] Edward J. Gillis. Causality, measurement, and elementary interactions. Foundations of Physics, 41(12):1757–1785, December 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-011-9576-x.

Waegell:2011:PPK

[3500] Mordecai Waegell and P. K. Aravind. Parity proofs of the Kochen– Specker theorem based on the 24 rays of Peres. Foundations of Physics, 41(12):1786–1799, December 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-011-9578-8.

Rebilas:2011:TPB

[3501] Krzysztof Rebilas. Thomas precession and the Bargmann–Michel– Telegdi equation. Foundations of Physics, 41(12):1800–1809, December 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10701-011-9579-7.pdf.

Lee:2011:INT

[3502] Chunghyoung Lee. Infinity and Newton’s three laws of motion. Foundations of Physics, 41(12):1810–1828, December 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-011-9580-1.

tHooft:2011:CEP

[3503] Gerard ’t Hooft. A class of elementary particle models without any adjustable real parameters. Foundations of Physics, 41(12):1829–1856, December 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10701-011-9586-8.pdf.

Omnes:2011:DWF

[3504] Roland Omnès. Decoherence and wave function collapse. Foundations of Physics, 41(12):1857–1880, December 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-011-9588-6.

Aspelmeyer:2012:E

[3505] Markus Aspelmeyer, Caslav Brukner, and Anton Zeilinger. Editorial. Foundations of Physics, 42(1):1–3, January 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-011-9621-9.

Pearle:2012:CC

[3506] Philip Pearle. Cosmogenesis and collapse. Foundations of Physics, 42 (1):4–18, January 2012. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-010-9482-7.

Wootters:2012:ESR

[3507] William K. Wootters. Entanglement sharing in real-vector-space quantum theory. Foundations of Physics, 42(1):19–28, January 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-010-9488-1.

Hasegawa:2012:EBD

[3508] Yuji Hasegawa. Entanglement between degrees of freedom in a single– particle system revealed in neutron interferometry. Foundations of Physics, 42(1):29–45, January 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-010-9499-y.

Greenberger:2012:TTT

[3509] Daniel M. Greenberger. The tic-tac-toe theory of gravity. Foundations of Physics, 42(1):46–52, January 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-010-9500-9.

Laloe:2012:QPS

[3510] F. Laloë and W. J. Mullin. Quantum properties of a single beam splitter. Foundations of Physics, 42(1):53–67, January 2012. CODEN FNDPA4.

ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-010-9501-8.

Cabello:2012:RBM

[3511] Adán Cabello. The role of bounded memory in the foundations of quantum mechanics. Foundations of Physics, 42(1):68–79, January 2012. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-010-9507-2.

Gisin:2012:NRD

[3512] Nicolas Gisin. Non-realism: Deep thought or a soft option? Foundations of Physics, 42(1):80–85, January 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-010-9508-1.

Stenholm:2012:CQR

[3513] Stig Stenholm. The construction of quantum reality. Foundations of Physics, 42(1):86–97, January 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-010-9519-y.

Juffmann:2012:NPB

[3514] Thomas Juffmann, Stefan Nimmrichter, Markus Arndt, Herbert Gleiter, and Klaus Hornberger. New prospects for de Broglie interferometry. Foundations of Physics, 42(1):98–110, January 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-010-9520-5.

Tamma:2012:PND

[3515] V. Tamma, C. O. Alley, W. P. Schleich, and Y. H. Shih. Prime number decomposition, the hyperbolic function and multi-path Michelson interferometers. Foundations of Physics, 42(1):111–121, January 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-010-9522-3.

Werner:2012:OBG

[3516] Sam Werner. Observation of Berry’s geometric phase by neutron interferometry. Foundations of Physics, 42(1):122–139, January 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-010-9526-z.

Horne:2012:NCF

[3517] Michael Horne. On non-conservative forces and topological quantum phases. Foundations of Physics, 42(1):140–146, January 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-010-9528-x.

Klein:2012:ANI

[3518] A. G. Klein. Adventures in neutron interferometry. Foundations of Physics, 42(1):147–152, January 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-010-9529-9.

Rauch:2012:QPN

[3519] Helmut Rauch. Quantum physics with neutrons: From spinor symmetry to Kochen–Specker phenomena. Foundations of Physics, 42(1): 153–172, January 2012. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-010-9530-3.

Chiao:2012:QIF

[3520] R. Y. Chiao, S. J. Minter, K. Wegter-McNelly, and L. A. Martinez. Quantum incompressibility of a falling Rydberg atom, and a gravitationallyinduced charge separation effect in superconducting systems. Foundations of Physics, 42(1):173–191, January 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10701-010-9531-2.pdf.

Hiley:2012:CAD

[3521] B. J. Hiley and R. E. Callaghan. Clifford algebras and the Dirac– Bohm quantum Hamilton–Jacobi equation. Foundations of Physics, 42 (1):192–208, January 2012. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-011-9558-z.

Weatherall:2012:BRE

[3522] James Owen Weatherall. A brief remark on energy conditions and the Geroch–Jang Theorem. Foundations of Physics, 42(2):209–214, February 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
011-9583-y.

Louis-Martinez:2012:RAD

[3523] Domingo J. Louis-Martinez. Relativistic action at a distance and fields. Foundations of Physics, 42(2):215–223, February 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-011-9589-5.

Romero:2012:CFS

[3524] C. Romero, J. B. Fonseca-Neto, and M. L. Pucheu. Conformally flat spacetimes and Weyl frames. Foundations of Physics, 42(2):224– 240, February 2012. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-011-9593-9.

Hofer-Szabo:2012:RCC

[3525] Gábor Hofer-Szabó and Péter Vecsernyés. Reichenbach’s common cause principle in algebraic quantum field theory with locally finite degrees of freedom. Foundations of Physics, 42(2):241–255, February 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-011-9594-8.

Wilde:2012:ACL

[3526] Mark M. Wilde and Ari Mizel. Addressing the clumsiness loophole in a Leggett–Garg test of macrorealism. Foundations of Physics, 42(2): 256–265, February 2012. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-011-9598-4.

Das:2012:VGS

[3527] Arundhati Das, Surajit Chattopadhyay, and Ujjal Debnath. Validity of the generalized Second Law of Thermodynamics in the logamediate and intermediate scenarios of the Universe. Foundations of Physics, 42 (2):266–283, February 2012. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-011-9600-1.

Tammaro:2012:MNC

[3528] Elliott Tammaro. Mechanics: Non-classical, non-quantum. Foundations of Physics, 42(2):284–290, February 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-011-9602-z.

Fivel:2012:DRQ

[3529] Daniel I. Fivel. Derivation of the rules of quantum mechanics from information-theoretic axioms. Foundations of Physics, 42(2):291– 318, February 2012. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-011-9603-y.

Kuic:2012:MTE

[3530] Domagoj Kuić, Pasko Zupanović, and Davor Juretić. Macroscopic time evolution and MaxEnt inference for closed systems with Hamiltonian dynamics. Foundations of Physics, 42(2):319–339, February 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-011-9604-x.

Brun:2012:PSD

[3531] Todd A. Brun and Mark M. Wilde. Perfect state distinguishability and computational speedups with postselected closed timelike curves. Foundations of Physics, 42(3):341–361, March 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-011-9601-0.

Perez:2012:RSQ

[3532] B. Carballo Pérez and H. García-Compeán. Rarita–Schwinger quantum free field via deformation quantization. Foundations of Physics, 42 (3):362–368, March 2012. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-011-9605-9.

DAbramo:2012:NSS

[3533] Germano D’Abramo. A note on solid-state Maxwell demon. Foundations of Physics, 42(3):369–376, March 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-011-9607-7.

Itto:2012:GNW

[3534] Yuichi Itto and Sumiyoshi Abe. On the geodesic nature of Wegner’s flow. Foundations of Physics, 42(3):377–387, March 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-011-9606-8.

Mannheim:2012:MCC

[3535] Philip D. Mannheim. Making the case for conformal gravity. Foundations of Physics, 42(3):388–420, March 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-011-9608-6.

Kent:2012:RWI

[3536] Adrian Kent. Real world interpretations of quantum theory. Foundations of Physics, 42(3):421–435, March 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-011-9610-z.

Barletta:2012:GFM

[3537] Elisabetta Barletta and Sorin Dragomir. Gravity as a Finslerian metric phenomenon. Foundations of Physics, 42(3):436–453, March 2012. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-011-9614-8.

Hardy:2012:LHR

[3538] Lucien Hardy and William K. Wootters. Limited holism and real-vector-space quantum theory. Foundations of Physics, 42(3):454–473, March 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
011-9616-6.

Ozawa:2012:RBR

[3539] Masanao Ozawa and Yuichiro Kitajima. Reconstructing Bohr’s reply to EPR in algebraic quantum theory. Foundations of Physics, 42 (4):475–487, April 2012. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-011-9615-7.

Chamseddine:2012:VFS

[3540] Riad Chamseddine. Vectorial form of the successive Lorentz transformations. Application: Thomas rotation. Foundations of Physics, 42 (4):488–511, April 2012. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-011-9617-5.

Mazzola:2012:RCC

[3541] Claudio Mazzola. Reichenbachian common cause systems revisited. Foundations of Physics, 42(4):512–523, April 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-011-9622-8.

Mazharimousavi:2012:GSH

[3542] S. Habib Mazharimousavi and M. Halilsoy. Ground state H-atom in Born–Infeld theory. Foundations of Physics, 42(4):524–530, April 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-011-9623-7.

Tamir:2012:WMW

[3543] Boaz Tamir and Sergei Masis. Weak measurement and weak information. Foundations of Physics, 42(4):531–543, April 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-012-9624-1.

Budroni:2012:BIC

[3544] Costantino Budroni and Giovanni Morchio. Bell inequalities as constraints on unmeasurable correlations. Foundations of Physics, 42(4): 544–554, April 2012. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-012-9625-0.

Hajicek:2012:CSS

[3545] P. Hájícek. Changes of separation status during registration and scattering. Foundations of Physics, 42(4):555–581, April 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-012-9626-z.

Swendsen:2012:CDE

[3546] Robert H. Swendsen. Choosing a definition of entropy that works. Foundations of Physics, 42(4):582–593, April 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-012-9627-y.

Boyer:2012:BRS

[3547] Timothy H. Boyer. The blackbody radiation spectrum follows from zeropoint radiation and the structure of relativistic spacetime in classical physics. Foundations of Physics, 42(5):595–614, May 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-012-9628-x.

Grinbaum:2012:WFT

[3548] Alexei Grinbaum. Which fine-tuning arguments are fine? Foundations of Physics, 42(5):615–631, May 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-012-9629-9.

Nikolic:2012:HVN

[3549] Hrvoje Nikolić. Hidden variables with nonlocal time. Foundations of Physics, 42(5):632–646, May 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-012-9631-2.

Stapp:2012:QL

[3550] Henry P. Stapp. Quantum locality? Foundations of Physics, 42 (5):647–655, May 2012. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-012-9632-1.

Wallman:2012:RCC

[3551] Joel J. Wallman and Stephen D. Bartlett. Revisiting consistency conditions for quantum states of systems on closed timelike curves: an epistemic perspective. Foundations of Physics, 42(5):656–673, May 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-012-9635-y.

Griffiths:2012:QCL

[3552] Robert B. Griffiths. Quantum counterfactuals and locality. Foundations of Physics, 42(5):674–684, May 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-012-9637-9.

tHooft:2012:PQG

[3553] Gerard ’t Hooft. The pursuit of quantum gravity. Foundations of Physics, 42(5):685–687, May 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10701-012-9638-8.pdf.

Pusey:2012:SNS

[3554] Matthew F. Pusey. Stabilizer notation for Spekkens’ toy theory. Foundations of Physics, 42(5):688–708, May 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-012-9639-7.

Mansfield:2012:HNL

[3555] Shane Mansfield and Tobias Fritz. Hardy’s Non-locality Paradox and possibilistic conditions for non-locality. Foundations of Physics, 42 (5):709–719, May 2012. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-012-9640-1.

Khrennikov:2012:PSI

[3556] Andrei Khrennikov and Gregor Weihs. Preface of the special issue quantum foundations: Theory and experiment. Foundations of Physics, 42(6): 721–724, June 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10701-012-9644-x.pdf.

Khrennikov:2012:TFM

[3557] Andrei Khrennikov. Towards a field model of prequantum reality. Foundations of Physics, 42(6):725–741, June 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-011-9611-y.

Sollner:2012:TBR

[3558] Immo Söllner, Benjamin Gschösser, Patrick Mai, Benedikt Pressl, and Gregor Weihs Zoltán Vörös. Testing Born’s rule in quantum mechanics for three mutually exclusive events. Foundations of Physics, 42 (6):742–751, June 2012. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-011-9597-5.

Jaeger:2012:GQP

[3559] Gregg Jaeger. Generalized quantum probability and entanglement enhancement witnessing. Foundations of Physics, 42(6):752–759, June 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
012-9633-0.

Rauch:2012:NMW

[3560] Helmut Rauch. Neutron matter wave quantum optics. Foundations of Physics, 42(6):760–777, June 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10701-011-9569-9.pdf.

DiDomenico:2012:HUR

[3561] Antonio Di Domenico, Andreas Gabriel, Beatrix C. Hiesmayr, Florian Hipp, Marcus Huber, Gerd Krizek, Karoline Mühlbacher, Sasa Radic, Christoph Spengler, and Lukas Theussl. Heisenberg’s uncertainty relation and Bell inequalities in high energy physics. Foundations of Physics, 42(6):778–802, June 2012. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-011-9575-y.

Spence:2012:EED

[3562] S. E. Spence and A. D. Parks. Experimental evidence for a dynamical non-locality induced effect in quantum interference using weak values. Foundations of Physics, 42(6):803–815, June 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-011-9596-6.

Coecke:2012:P

[3563] Bob Coecke, Prakash Panangaden, and Peter Selinger. Preface. Foundations of Physics, 42(7):817–818, July 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10701-012-9661-9.pdf.

Baez:2012:DAQ

[3564] John C. Baez. Division algebras and quantum theory. Foundations of Physics, 42(7):819–855, July 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-011-9566-z.

Heunen:2012:CCQ

[3565] Chris Heunen. Complementarity in categorical quantum mechanics. Foundations of Physics, 42(7):856–873, July 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-011-9585-9.

Jacobs:2012:ICM

[3566] Bart Jacobs. Involutive categories and monoids, with a GNScorrespondence. Foundations of Physics, 42(7):874–895, July 2012. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/content/pdf/10.1007/s10701-011-9595-
7.pdf.

Spitters:2012:SMO

[3567] Bas Spitters. The space of measurement outcomes as a spectral invariant for non-commutative algebras. Foundations of Physics, 42(7): 896–908, July 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10701-011-9619-3.pdf.

Crane:2012:HEM

[3568] Louis Crane. Holography in the EPRL model. Foundations of Physics, 42(7):909–917, July 2012. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-012-9630-3.

Schumacher:2012:MQT

[3569] Benjamin Schumacher and Michael D. Westmoreland. Modal quantum theory. Foundations of Physics, 42(7):918–925, July 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-012-9650-z.

Schumacher:2012:IIF

[3570] Benjamin Schumacher and Michael D. Westmoreland. Isolation and information flow in quantum dynamics. Foundations of Physics, 42 (7):926–931, July 2012. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-012-9651-y.

Jacobs:2012:CAQ

[3571] Bart Jacobs and Jorik Mandemaker. Coreflections in algebraic quantum logic. Foundations of Physics, 42(7):932–958, July 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s10701-012-9654-8.pdf.

Feng:2012:DUC

[3572] Johnny Feng. A domain of unital channels. Foundations of Physics, 42(7):959–975, July 2012. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-012-9656-6.

Crowder:2012:ITR

[3573] Tanner Crowder and Keye Martin. Information theoretic representations of qubit channels. Foundations of Physics, 42(7):976–983, July 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-012-9657-5.

Navascues:2012:PAT

[3574] M. Navascués, T. Cooney, D. Pérez-García, and N. Villanueva. A physical approach to Tsirelson’s problem. Foundations of Physics, 42(8): 985–995, August 2012. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-012-9641-0.

Babin:2012:RDA

[3575] Anatoli Babin and Alexander Figotin. Relativistic dynamics of accelerating particles derived from field equations. Foundations of Physics, 42 (8):996–1014, August 2012. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-012-9642-z.

Elliott:2012:ILP

[3576] S. R. Elliott, B. H. LaRoque, V. M. Gehman, M. F. Kidd, and M. Chen. An improved limit on Pauli-exclusion-principle forbidden atomic transitions. Foundations of Physics, 42(8):1015–1030, August 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-012-9643-y.

Castro:2012:BRG

[3577] Carlos Castro. Born’s reciprocal gravity in curved phase-spaces and the cosmological constant. Foundations of Physics, 42(8):1031–1055, August 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
012-9645-9.

Lamport:2012:BP

[3578] Leslie Lamport. Buridan’s principle. Foundations of Physics, 42(8): 1056–1066, August 2012. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-012-9647-7.

Sucu:2012:SIC

[3579] Yusuf Sucu and Nuri Ünal. Symmetry and integrability in the classical model of Zitterbewegung. Foundations of Physics, 42(8):1067–1077, August 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
012-9648-6.

Dvurecenskij:2012:AEA

[3580] Anatolij Dvurecenskij and Yongjian Xie. Atomic effect algebras with the Riesz decomposition property. Foundations of Physics, 42(8):1078– 1093, August 2012. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-012-9655-7.

Kastner:2012:PTI

[3581] Ruth E. Kastner. The possibilist transactional interpretation and relativity. Foundations of Physics, 42(8):1094–1113, August 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-012-9658-4.

Besnard:2012:SMS

[3582] Fabien Besnard. Simultaneity in Minkowski spacetime: From uniqueness to arbitrariness. Foundations of Physics, 42(9):1115–1134, September 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
012-9649-5.

Castro:2012:SPE

[3583] Carlos Castro. On superluminal particles and the extended relativity theories. Foundations of Physics, 42(9):1135–1152, September 2012. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-012-9659-3.

Lee:2012:OEG

[3584] Jae-Weon Lee. On the origin of entropic gravity and inertia. Foundations of Physics, 42(9):1153–1164, September 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-012-9660-x.

Kupervasser:2012:UAT

[3585] Oleg Kupervasser, Hrvoje Nikolić, and Vinko Zlatić. The universal arrow of time. Foundations of Physics, 42(9):1165–1185, September 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-012-9662-8.

Nussinov:2012:ESR

[3586] Shmuel Nussinov and Jeff Tollaksen. Extreme sub-radiance: Can quantum effects generate dramatically longer atomic lifetimes? Foundations of Physics, 42(9):1186–1199, September 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-012-9663-7.

Kennerly:2012:ID

[3587] Sam Kennerly. Illusory decoherence. Foundations of Physics, 42(9): 1200–1209, September 2012. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-012-9664-6.

Gryb:2012:RTR

[3588] Sean Gryb and Karim Thébault. The role of time in relational quantum theories. Foundations of Physics, 42(9):1210–1238, September 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-012-9665-5.

Smolin:2012:REI

[3589] Lee Smolin. A real ensemble interpretation of quantum mechanics. Foundations of Physics, 42(10):1239–1261, October 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-012-9666-4.

Giacosa:2012:NED

[3590] Francesco Giacosa. Non-exponential decay in quantum field theory and in quantum mechanics: The case of two (or more) decay channels. Foundations of Physics, 42(10):1262–1299, October 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-012-9667-3.

Pollock:2012:VFP

[3591] M. D. Pollock. On vacuum fluctuations and particle masses. Foundations of Physics, 42(10):1300–1328, October 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-012-9668-2.

Cavalcanti:2012:BNS

[3592] Eric G. Cavalcanti and Howard M. Wiseman. Bell nonlocality, signal locality and unpredictability (or what Bohr could have told Einstein at Solvay had He known about Bell experiments). Foundations of Physics, 42(10):1329–1338, October 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-012-9669-1.

Arcos:2012:SFH

[3593] H. I. Arcos, C. S. O. Mayor, G. Otalora, and J. G. Pereira. Spin-2 fields and helicity. Foundations of Physics, 42(10):1339–1349, October 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-012-9670-8.

Hernandez-Coronado:2012:BSR

[3594] H. Hernandez-Coronado. From Bargmann’s superselection rule to quantum Newtonian spacetime. Foundations of Physics, 42(10):1350– 1364, October 2012. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-012-9673-5.

Mamone-Capria:2012:SIE

[3595] Marco Mamone-Capria. Simultaneity as an invariant equivalence relation. Foundations of Physics, 42(11):1365–1383, November 2012. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-012-9674-4.

Wetterich:2012:PT

[3596] C. Wetterich. Probabilistic time. Foundations of Physics, 42(11):1384– 1443, November 2012. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-012-9675-3.

Briscese:2012:TNR

[3597] Fabio Briscese, Yeinzon Rodríguez, and Guillermo A. González. On the true nature of renormalizability in Horava–Lifshitz gravity. Foundations of Physics, 42(11):1444–1451, November 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-012-9677-1.

Hossenfelder:2012:QSS

[3598] Sabine Hossenfelder. Quantum superpositions of the speed of light. Foundations of Physics, 42(11):1452–1468, November 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-012-9678-0.

Dvornikov:2012:CQM

[3599] Maxim Dvornikov. Canonical quantization of a massive Weyl field. Foundations of Physics, 42(11):1469–1479, November 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-012-9679-z.

Schroer:2012:CDR

[3600] Bert Schroer. Causality and dispersion relations and the role of the Smatrix in the ongoing research. Foundations of Physics, 42(12):1481– 1522, December 2012. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-012-9676-2.

Peil:2012:PTR

[3601] Steven Peil. Proposed test of relative phase as hidden variable in quantum mechanics. Foundations of Physics, 42(12):1523–1533, December 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-012-9680-6.

Healey:2012:QDP

[3602] Richard Healey. Quantum decoherence in a pragmatist view: Dispelling Feynman’s mystery. Foundations of Physics, 42(12):1534–1555, December 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
012-9681-5.

Lochan:2012:STN

[3603] Kinjalk Lochan, Seema Satin, and Tejinder P. Singh. Statistical thermodynamics for a non-commutative special relativity: Emergence of a generalized quantum dynamics. Foundations of Physics, 42(12):1556– 1572, December 2012. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-012-9683-3.

Lindesay:2012:SCS

[3604] James Lindesay. Self-consistent solutions of canonical proper selfgravitating quantum systems. Foundations of Physics, 42(12):1573– 1585, December 2012. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-012-9684-2.

Esposito:2012:MST

[3605] S. Esposito and A. Naddeo. Majorana solutions to the two-electron problem. Foundations of Physics, 42(12):1586–1608, December 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-012-9685-1.

deHaro:2013:FYS

[3606] Sebastian de Haro, Dennis Dieks, Gerard ’t Hooft, and Erik Verlinde. Forty years of string theory reflecting on the foundations. Foundations of Physics, 43(1):1–7, January 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10701-012-9691-3.pdf.

Rovelli:2013:CLS

[3607] Carlo Rovelli. A critical look at strings. Foundations of Physics, 43 (1):8–20, January 2013. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-011-9599-3.

Smolin:2013:PLP

[3608] Lee Smolin. A perspective on the landscape problem. Foundations of Physics, 43(1):21–45, January 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-012-9652-x.

tHooft:2013:FST

[3609] Gerard ’t Hooft. On the foundations of superstring theory. Foundations of Physics, 43(1):46–53, January 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10701-012-9682-4.pdf.

Rickles:2013:MSO

[3610] Dean Rickles. Mirror symmetry and other miracles in superstring theory. Foundations of Physics, 43(1):54–80, January 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-010-9504-5.

Dawid:2013:TAF

[3611] Richard Dawid. Theory assessment and final theory claim in string theory. Foundations of Physics, 43(1):81–100, January 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-011-9592-x.

Balasubramanian:2013:WWD

[3612] Vijay Balasubramanian. What we don’t know about time. Foundations of Physics, 43(1):101–114, January 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-011-9591-y.

Giddings:2013:STT

[3613] Steven B. Giddings. Is string theory a theory of quantum gravity? Foundations of Physics, 43(1):115–139, January 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-011-9612-x.

Gubser:2013:GSD

[3614] Steven S. Gubser. The gauge–string duality and heavy ion collisions. Foundations of Physics, 43(1):140–155, January 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-011-9613-9.

Martinec:2013:ENG

[3615] Emil J. Martinec. Evolving notions of geometry in string theory. Foundations of Physics, 43(1):156–173, January 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-011-9609-5.

Susskind:2013:ST

[3616] Leonard Susskind. String theory. Foundations of Physics, 43(1): 174–181, January 2013. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-011-9620-x.

Duff:2013:SMT

[3617] M. J. Duff. String and M-theory: answering the critics. Foundations of Physics, 43(1):182–200, January 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-011-9618-4. See erratum [3625].

Hammond:2013:ERR

[3618] Richard T. Hammond. Electrodynamics and radiation reaction. Foundations of Physics, 43(2):201–209, February 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-012-9687-z.

Dvurecenskij:2013:SOS

[3619] Anatolij Dvurecenskij. Smearing of observables and spectral measures on quantum structures. Foundations of Physics, 43(2):210–224, February 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
012-9689-x.

Hassanabadi:2013:ESD

[3620] H. Hassanabadi, Z. Molaee, and A. Boumali. Exact solutions of the Duffin–Kemmer–Petiau equation with a pseudoharmonic potential in the presence of a magnetic field in (1 + 2) dimensions. Foundations of Physics, 43(2):225–235, February 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-012-9690-4.

Dantas:2013:ALQ

[3621] Christine C. Dantas. An approach to loop quantum cosmology through integrable discrete Heisenberg spin chains. Foundations of Physics, 43 (2):236–242, February 2013. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-012-9692-2.

Guendelman:2013:MLP

[3622] Eduardo Guendelman and Roee Steiner. Mach like principle from conserved charges. Foundations of Physics, 43(2):243–266, February 2013.

CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-012-9693-1.

Singal:2013:CEE

[3623] Ashok K. Singal. Comment on “Exact Expression for Radiation of an Accelerated Charge in Classical Electrodynamics”. Foundations of Physics, 43(2):267–270, February 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-012-9694-0. See [3135].

Sestito:2013:SER

[3624] Angela Sestito. Simultaneous elements of reality for incompatible properties by exploiting locality. Foundations of Physics, 43(2):271–283, February 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
012-9697-x.

Duff:2013:ESM

[3625] M. J. Duff. Erratum to: String and M-theory: answering the critics. Foundations of Physics, 43(2):284, February 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10701-012-9696-y.pdf. See

[3617].

Xi:2013:QCC

[3626] Zhengjun Xi and Yongming Li. Quantum and classical correlations in quantum measurement. Foundations of Physics, 43(3):285–293, March 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
012-9686-0.

Boisvert:2013:ATI

[3627] Jean-Sébastien Boisvert and Louis Marchildon. Absorbers in the transactional interpretation of quantum mechanics. Foundations of Physics, 43(3):294–309, March 2013. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-012-9695-z.

Schroeren:2013:DHS

[3628] David P. B. Schroeren. Decoherent histories of spin networks. Foundations of Physics, 43(3):310–328, March 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-013-9698-4.

Schroer:2013:MLF

[3629] Bert Schroer. Modular localization and the foundational origin of integrability. Foundations of Physics, 43(3):329–372, March 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-013-9699-3.

Landsman:2013:FSC

[3630] N. P. (Klaas) Landsman and Robin Reuvers. A flea on Schrödinger’s cat. Foundations of Physics, 43(3):373–407, March 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-013-9700-1.

Dewdney:2013:I

[3631] Chris Dewdney and Paavo Pylkkänen. Introduction. Foundations of Physics, 43(4):409–411, April 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10701-013-9705-9.pdf.

Atmanspacher:2013:AHR

[3632] Harald Atmanspacher. Appreciating a Hiley respected colleague. Foundations of Physics, 43(4):412–414, April 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10701-013-9706-8.pdf.

Anonymous:2013:BHP

[3633] Anonymous. Basil Hiley: Publications 1959–2013. Foundations of Physics, 43(4):415–423, April 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-013-9704-x.

Binz:2013:CAS

[3634] Ernst Binz, Maurice A. de Gosson, and Basil J. Hiley. Clifford algebras in symplectic geometry and quantum mechanics. Foundations of Physics, 43(4):424–439, April 2013. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-012-9634-z.

deGosson:2013:QB

[3635] Maurice A. de Gosson. Quantum blobs. Foundations of Physics, 43(4): 440–457, April 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10701-012-9636-x.pdf.

Coecke:2013:CCR

[3636] Bob Coecke and Raymond Lal. Causal categories: Relativistically interacting processes. Foundations of Physics, 43(4):458–501, April 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-012-9646-8.

Silberstein:2013:BBU

[3637] Michael Silberstein, W. M. Stuckey, and Timothy McDevitt. Being, becoming and the undivided Universe: a dialogue between relational blockworld and the implicate order concerning the unification of relativity and quantum theory. Foundations of Physics, 43(4):502–532, April 2013. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-012-9653-9.

Filk:2013:TNL

[3638] Thomas Filk. Temporal non-locality. Foundations of Physics, 43(4): 533–547, April 2013. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-012-9671-7.

Seager:2013:CLR

[3639] William Seager. Classical levels, Russellian monism and the implicate order. Foundations of Physics, 43(4):548–567, April 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-012-9672-6.

Wikman:2013:NOM

[3640] Georg Wikman. The notion of order in mathematics and physics. Similarity, difference and indistinguishability. Foundations of Physics, 43(4): 568–596, April 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10701-012-9688-y.pdf.

tHooft:2013:DBD

[3641] Gerard ’t Hooft. Duality between a deterministic cellular automaton and a bosonic quantum field theory in 1 + 1 dimensions. Foundations of Physics, 43(5):597–614, May 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-013-9702-z.

Pollock:2013:ESS

[3642] M. D. Pollock. On the entropy of Schwarzschild space–time. Foundations of Physics, 43(5):615–630, May 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-013-9701-0.

Santamato:2013:DDE

[3643] Enrico Santamato and Francesco De Martini. Derivation of the Dirac equation by conformal differential geometry. Foundations of Physics, 43(5):631–641, May 2013. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-013-9703-y.

Brasil:2013:HMT

[3644] Carlos Alexandre Brasil, L. A. de Castro, and R. D. J. Napolitano. How much time does a measurement take? Foundations of Physics, 43(5):642–655, May 2013. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-013-9707-7.

Graydon:2013:QQD

[3645] Matthew A. Graydon. Quaternionic quantum dynamics on complex Hilbert spaces. Foundations of Physics, 43(5):656–664, May 2013. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-013-9708-6.

Drory:2013:SRC

[3646] Alon Drory. Special relativity cannot be derived from Galilean mechanics alone. Foundations of Physics, 43(5):665–684, May 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-013-9709-5.

DiLorenzo:2013:RLI

[3647] Antonio Di Lorenzo. Reassessment of Leggett inequality. Foundations of Physics, 43(5):685–698, May 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-013-9710-z.

Afriat:2013:WGA

[3648] Alexander Afriat. Weyl’s gauge argument. Foundations of Physics, 43 (5):699–705, May 2013. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-013-9712-x.

Kent:2013:MQI

[3649] Adrian Kent. Might quantum-induced deviations from the Einstein equations detectably affect gravitational wave propagation? Foundations of Physics, 43(6):707–718, June 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-013-9716-6.

Vacaru:2013:SLE

[3650] Sergiu I. Vacaru. Super-luminal effects for Finsler branes as a way to preserve the paradigm of relativity theories. Foundations of Physics, 43(6):719–732, June 2013. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-013-9711-y.

Heaney:2013:SIK

[3651] Michael B. Heaney. A symmetrical interpretation of the Klein–Gordon equation. Foundations of Physics, 43(6):733–746, June 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-013-9713-9.

Pearle:2013:CCC

[3652] Philip Pearle. Chronogenesis, cosmogenesis and collapse. Foundations of Physics, 43(6):747–768, June 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-013-9714-8.

Vervoort:2013:BTT

[3653] Louis Vervoort. Bell’s theorem: Two neglected solutions. Foundations of Physics, 43(6):769–791, June 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-013-9715-7.

Sozzo:2013:QHO

[3654] Sandro Sozzo. The quantum harmonic oscillator in the ESR model. Foundations of Physics, 43(6):792–804, June 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-013-9717-5.

Niestegge:2013:TSE

[3655] Gerd Niestegge. Three-slit experiments and quantum nonlocality. Foundations of Physics, 43(6):805–812, June 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-013-9719-3.

Donadi:2013:CMT

[3656] Sandro Donadi, Angelo Bassi, Catalina Curceanu, Antonio Di Domenico, and Beatrix C. Hiesmayr. Are collapse models testable via flavor oscillations? Foundations of Physics, 43(7):813–844, July 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-013-9720-x.

daCosta:2013:PLQ

[3657] N. da Costa and C. de Ronde. The paraconsistent logic of quantum superpositions. Foundations of Physics, 43(7):845–858, July 2013. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-013-9721-9.

John:2013:CSM

[3658] Moncy V. John and Kiran Mathew. Coherent states and modified de Broglie–Bohm complex quantum trajectories. Foundations of Physics, 43(7):859–871, July 2013. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-013-9722-8.

Egg:2013:FSL

[3659] Matthias Egg. The foundational significance of Leggett’s non-local hidden-variable theories. Foundations of Physics, 43(7):872–880, July 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
013-9723-7.

Ghirardi:2013:APE

[3660] GianCarlo Ghirardi and Raffaele Romano. About possible extensions of quantum theory. Foundations of Physics, 43(7):881–894, July 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-013-9724-6.

Asano:2013:NKA

[3661] Masanari Asano, Irina Basieva, Andrei Khrennikov, Masanori Ohya, and Ichiro Yamato. Non-Kolmogorovian approach to the context-dependent systems breaking the classical probability law. Foundations of Physics, 43(7):895–911, July 2013. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-013-9725-5.

Polyakov:2013:PIT

[3662] Sergey V. Polyakov, Fabrizio Piacentini, Paolo Traina, Ivo P. Degiovanni, Alan Migdall, Giorgio Brida, and Marco Genovese. Practical implementation of a test of event–based corpuscular model as an alternative to quantum mechanics. Foundations of Physics, 43(8):913–922, August 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
013-9718-4.

Boyer:2013:CCQ

[3663] Timothy H. Boyer. Contrasting classical and quantum vacuum states in non-inertial frames. Foundations of Physics, 43(8):923–947, August 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-013-9726-4.

Foulis:2013:TDS

[3664] David J. Foulis and Sylvia Pulmannová. Type-decomposition of a synaptic algebra. Foundations of Physics, 43(8):948–968, August 2013. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-013-9727-3.

Arageorgis:2013:PPP

[3665] Aristidis Arageorgis and Chrysovalantis Stergiou. On particle phenomenology without particle ontology: How much local is almost local? Foundations of Physics, 43(8):969–977, August 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-013-9728-2.

Nomura:2013:QMS

[3666] Yasunori Nomura. Quantum mechanics, spacetime locality, and gravity. Foundations of Physics, 43(8):978–1007, August 2013. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-013-9729-1;
 http://link.springer.com/content/pdf/10.1007/s10701-013-9729-
1.pdf.

Henson:2013:NSD

[3667] Joe Henson. Non-separability does not relieve the problem of Bell’s Theorem. Foundations of Physics, 43(8):1008–1038, August 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-013-9730-8.

Xie:2013:SQM

[3668] Yongjian Xie, Aili Yang, and Fang Ren. Super quantum measures on finite spaces. Foundations of Physics, 43(9):1039–1065, September 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-013-9731-7.

Donadi:2013:ESC

[3669] Sandro Donadi, Angelo Bassi, Luca Ferialdi, and Catalina Curceanu. The effect of spontaneous collapses on neutrino oscillations. Foundations of Physics, 43(9):1066–1089, September 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-013-9732-6.

Zafiris:2013:CSR

[3670] Elias Zafiris and Vassilios Karakostas. A categorial semantic representation of quantum event structures. Foundations of Physics, 43(9):1090– 1123, September 2013. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-013-9733-5.

Egg:2013:DCE

[3671] Matthias Egg. Delayed-choice experiments and the metaphysics of entanglement. Foundations of Physics, 43(9):1124–1135, September 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-013-9734-4.

Dvurecenskij:2013:BFP

[3672] Anatolij Dvurecenskij and Jirí Janda. On bilinear forms from the point of view of generalized effect algebras. Foundations of Physics, 43(9): 1136–1152, September 2013. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-013-9736-2.

Weatherall:2013:SGB

[3673] James Owen Weatherall. The scope and generality of Bell’s Theorem. Foundations of Physics, 43(9):1153–1169, September 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-013-9737-1.

Barakat:2013:GDE

[3674] T. Barakat and H. A. Alhendi. Generalized Dirac equation with induced energy-dependent potential via simple similarity transformation and asymptotic iteration methods. Foundations of Physics, 43(10): 1171–1181, October 2013. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-013-9738-0.

Kitajima:2013:ESB

[3675] Yuichiro Kitajima. EPR states and Bell correlated states in algebraic quantum field theory. Foundations of Physics, 43(10):1182–1192, October 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
013-9739-z.

Svensson:2013:WQM

[3676] Bengt E. Y. Svensson. What is a quantum-mechanical “weak value” the value of? Foundations of Physics, 43(10):1193–1205, October 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
013-9740-6;   http://link.springer.com/content/pdf/10.1007/
s10701-013-9740-6.pdf.

Oeckl:2013:PFQ

[3677] Robert Oeckl. A positive formalism for quantum theory in the general boundary formulation. Foundations of Physics, 43(10):1206–1232, October 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
013-9741-5.

Kaplan:2013:PEP

[3678] I. G. Kaplan. The Pauli Exclusion Principle. Can it be proved? Foundations of Physics, 43(10):1233–1251, October 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-013-9742-4.

Halabi:2013:CSS

[3679] T. Halabi. A conservative solution to the stochastic dynamical reduction problem. Foundations of Physics, 43(10):1252–1256, October 2013. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-013-9743-3.

Xu:2013:QTT

[3680] Dai-Yu Xu, Towe Wang, and Xun Xue. Quantum tunneling time: Relativistic extensions. Foundations of Physics, 43(11):1257–1274, November 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
013-9744-2.

Civitarese:2013:CEQ

[3681] O. Civitarese and M. Gadella. On the concept of entropy for quantum decaying systems. Foundations of Physics, 43(11):1275–1294, November 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
013-9746-0.

Haag:2013:SLI

[3682] Rudolf Haag. On the sharpness of localization of individual events in space and time. Foundations of Physics, 43(11):1295–1313, November 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
013-9747-z.

Dvurecenskij:2013:KPE

[3683] Anatolij Dvurecenskij. Kite pseudo effect algebras. Foundations of Physics, 43(11):1314–1338, November 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-013-9748-y.

Omnes:2013:LPE

[3684] Roland Omnès. Local properties of entanglement and application to collapse. Foundations of Physics, 43(11):1339–1368, November 2013. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-013-9750-4.

Fletcher:2013:LCC

[3685] Samuel C. Fletcher. Light clocks and the clock hypothesis. Foundations of Physics, 43(11):1369–1383, November 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-013-9751-3.

Norton:2013:BCT

[3686] John D. Norton. Brownian computation is thermodynamically irreversible. Foundations of Physics, 43(11):1384–1410, November 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-013-9753-1.

Barnum:2013:ESW

[3687] Howard Barnum, Carl Philipp Gaebler, and Alexander Wilce. Ensemble steering, weak self-duality, and the structure of probabilistic theories. Foundations of Physics, 43(12):1411–1427, December 2013. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-013-9752-2.

Gomes:2013:FAQ

[3688] Henrique Gomes and Tim Koslowski. Frequently asked questions about shape dynamics. Foundations of Physics, 43(12):1428–1458, December 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
013-9754-0.

Buric:2013:HFS

[3689] N. Burić, D. B. Popović, M. Radonjić, and S. Prvanović. Hamiltonian formulation of statistical ensembles and mixed states of quantum and hybrid systems. Foundations of Physics, 43(12):1459–1477, December 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-013-9755-z.

Bohmer:2013:NAM

[3690] Christian G. Böhmer and Nicola Tamanini. A new approach to modifying theories of gravity. Foundations of Physics, 43(12):1478–1488, December 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
013-9756-y.

Franklin:2013:RBM

[3691] Jerrold Franklin. Rigid body motion in Special Relativity. Foundations of Physics, 43(12):1489–1501, December 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-013-9757-x.

Wallden:2013:DIS

[3692] Petros Wallden. Distinguishing initial state-vectors from each other in histories formulations and the PBR argument. Foundations of Physics, 43(12):1502–1525, December 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-013-9759-8.

Vassallo:2014:PBO

[3693] Antonio Vassallo and Michael Esfeld. A proposal for a Bohmian ontology of quantum gravity. Foundations of Physics, 44(1):1–18, January 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-013-9745-1.

Okon:2014:CCH

[3694] Elias Okon and Daniel Sudarsky. On the consistency of the consistent histories approach to quantum mechanics. Foundations of Physics, 44 (1):19–33, January 2014. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-013-9760-2.

Heinosaari:2014:SIQ

[3695] Teiko Heinosaari, Takayuki Miyadera, and Daniel Reitzner. Strongly incompatible quantum devices. Foundations of Physics, 44(1):34–57, January 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
013-9761-1.

Ryff:2014:NRD

[3696] Luiz Carlos Ryff. Null-result detection and Einstein–Podolsky–Rosen correlations. Foundations of Physics, 44(1):58–70, January 2014. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-013-9762-0.

Pellonpaa:2014:CMQ

[3697] Juha-Pekka Pellonpää. Complete measurements of quantum observables. Foundations of Physics, 44(1):71–90, January 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-013-9764-y.

Rovelli:2014:WG

[3698] Carlo Rovelli. Why gauge? Foundations of Physics, 44(1):91–104, January 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
013-9768-7.

Batouli:2014:CID

[3699] J. Batouli and M. El Baz. Classical interpretation of a deformed quantum oscillator. Foundations of Physics, 44(2):105–113, February 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-013-9766-9.

Okon:2014:BOC

[3700] Elias Okon and Daniel Sudarsky. Benefits of objective collapse models for cosmology and quantum gravity. Foundations of Physics, 44(2): 114–143, February 2014. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-014-9772-6.

Davidson:2014:TVM

[3701] Mark Davidson. Theories of variable mass particles and low energy nuclear phenomena. Foundations of Physics, 44(2):144–174, February 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-014-9774-4.

Abe:2014:FPT

[3702] Sumiyoshi Abe. Fokker–Planck theory of nonequilibrium systems governed by hierarchical dynamics. Foundations of Physics, 44(2):175– 182, February 2014. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-014-9775-3.

Eingorn:2014:EQE

[3703] Maxim V. Eingorn and Vitaliy D. Rusov. Emergent quantum Euler equation and Bose–Einstein condensates. Foundations of Physics, 44 (2):183–191, February 2014. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-014-9782-4.

Barnum:2014:LTJ

[3704] Howard Barnum and Alexander Wilce. Local tomography and the Jordan structure of quantum theory. Foundations of Physics, 44(2): 192–212, February 2014. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-014-9777-1.

Woodard:2014:NMC

[3705] R. P. Woodard. Nonlocal models of cosmic acceleration. Foundations of Physics, 44(2):213–233, February 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-014-9780-6.

Sheehan:2014:ETT

[3706] D. P. Sheehan, D. J. Mallin, J. T. Garamella, and W. F. Sheehan. Experimental test of a thermodynamic paradox. Foundations of Physics, 44(3):235–247, March 2014. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-014-9781-5.

Dzhafarov:2014:NFN

[3707] Ehtibar N. Dzhafarov and Janne V. Kujala. No-forcing and no-matching theorems for classical probability applied to quantum mechanics. Foundations of Physics, 44(3):248–265, March 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-014-9783-3.

Muralidhar:2014:CVF

[3708] K. Muralidhar. Complex vector formalism of harmonic oscillator in geometric algebra: Particle mass, spin and dynamics in complex vector space. Foundations of Physics, 44(3):266–295, March 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-014-9784-2.

Laudisa:2014:LTR

[3709] Federico Laudisa. On Leggett theories: a reply. Foundations of Physics, 44(3):296–304, March 2014. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-014-9787-z.

Gryb:2014:SEQ

[3710] Sean Gryb and Karim Thébault. Symmetry and evolution in quantum gravity. Foundations of Physics, 44(3):305–348, March 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-014-9789-x.

Shanahan:2014:CLR

[3711] Daniel Shanahan. A case for Lorentzian relativity. Foundations of Physics, 44(4):349–367, April 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10701-013-9765-x.pdf.

Pollock:2014:GES

[3712] M. D. Pollock. On gravitational effects in the Schrödinger equation. Foundations of Physics, 44(4):368–388, April 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-014-9776-2.

Khrennikov:2014:PFD

[3713] Andrei Khrennikov, Börje Nilsson, Sven Nordebo, and Igor Volovich. Photon flux and distance from the source: Consequences for quantum communication. Foundations of Physics, 44(4):389–405, April 2014. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-014-9786-0.

tHooft:2014:RQM

[3714] Gerard ’t Hooft. Relating the quantum mechanics of discrete systems to standard canonical quantum mechanics. Foundations of Physics, 44 (4):406–425, April 2014. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-014-9788-y.

Fortin:2014:PTD

[3715] Sebastian Fortin and Olimpia Lombardi. Partial traces in decoherence and in interpretation: What do reduced states refer to? Foundations of Physics, 44(4):426–446, April 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-014-9791-3.

Anonymous:2014:POP

[3716] Anonymous. Preface of the organizers to the Proceedings of “Horizons of Quantum Physics”. Foundations of Physics, 44(5):447–448, May 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-014-9798-9.

Zeilinger:2014:IPO

[3717] Anton Zeilinger. Introduction to the Proceedings of “Horizons of Quantum Physics” 2012. Foundations of Physics, 44(5):449–451, May 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-014-9799-8.

Bekenstein:2014:CQG

[3718] Jacob D. Bekenstein. Can quantum gravity be exposed in the laboratory? Foundations of Physics, 44(5):452–462, May 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-014-9779-z.

tHooft:2014:SFQ

[3719] Gerard ’t Hooft. Superstrings and the foundations of quantum mechanics. Foundations of Physics, 44(5):463–471, May 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-014-9790-4.

Kempf:2014:QGQ

[3720] Achim Kempf. Quantum gravity on a quantum computer? Foundations of Physics, 44(5):472–482, May 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-013-9735-3.

Diosi:2014:GRW

[3721] Lajos Diósi. Gravity-related wave function collapse. Foundations of Physics, 44(5):483–491, May 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-013-9767-8.

Mann:2014:QT

[3722] Robert B. Mann and Eduardo Martín-Martínez. Quantum thermometry. Foundations of Physics, 44(5):492–511, May 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-014-9797-x.

Jizba:2014:STG

[3723] Petr Jizba and Fabio Scardigli. Space–time grains: Roots of special and doubly special relativity. Foundations of Physics, 44(5):512–522, May 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
013-9758-9.

Scarani:2014:SCM

[3724] Valerio Scarani, Jean-Daniel Bancal, Antoine Suarez, and Nicolas Gisin. Strong constraints on models that explain the violation of Bell inequalities with hidden superluminal influences. Foundations of Physics, 44 (5):523–531, May 2014. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-014-9785-1.

Unruh:2014:HRB

[3725] W. G. Unruh. Has Hawking radiation been measured? Foundations of Physics, 44(5):532–545, May 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-014-9778-0.

Kleinert:2014:QFT

[3726] H. Kleinert. Quantum field theory of black-swan events. Foundations of Physics, 44(5):546–556, May 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-013-9749-x.

Penrose:2014:GQMa

[3727] Roger Penrose. On the gravitization of quantum mechanics 1: Quantum state reduction. Foundations of Physics, 44(5):557–575, May 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s10701-013-9770-0.pdf.

Celeri:2014:QCF

[3728] Lucas C. Céleri, Rafael M. Gomes, Radu Ionicioiu, Thomas Jennewein, Robert B. Mann, and Daniel R. Terno. Quantum control in foundational experiments. Foundations of Physics, 44(5):576–587, May 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-014-9792-2.

Benedictus:2014:STS

[3729] Fedde Benedictus. String theory & the scientific method. Foundations of Physics, 44(6):589–593, June 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-014-9794-0.

Wuthrich:2014:LA

[3730] Adrian Wüthrich. Local acausality. Foundations of Physics, 44(6): 594–609, June 2014. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-014-9796-y.

Griffiths:2014:NQL

[3731] Robert B. Griffiths. The new quantum logic. Foundations of Physics, 44(6):610–640, June 2014. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-014-9802-4.

Mohrhoff:2014:MQW

[3732] Ulrich Mohrhoff. Manifesting the quantum world. Foundations of Physics, 44(6):641–677, June 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-014-9803-3.

Aichmann:2014:TTB

[3733] Horst Aichmann and Günter Nimtz. On the traversal time of barriers. Foundations of Physics, 44(6):678–688, June 2014. CODEN FNDPA4.

ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-014-9804-2.

Pena:2014:PMU

[3734] Igor Peña and Daniel Sudarsky. On the possibility of measuring the Unruh effect. Foundations of Physics, 44(6):689–708, June 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-014-9806-0.

Jaeger:2014:IPC

[3735] Gregg Jaeger. On the identification of the parts of compound quantum objects. Foundations of Physics, 44(7):709–724, July 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-014-9795-z.

Arsenovic:2014:LFS

[3736] D. Arsenović, N. Burić, D. M. Davidović, and S. Prvanović. Lagrangian form of Schrödinger equation. Foundations of Physics, 44(7):725–735, July 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
014-9810-4.

Bierhorst:2014:RAC

[3737] Peter Bierhorst. A rigorous analysis of the Clauser–Horne–Shimony–Holt inequality experiment when trials need not be independent. Foundations of Physics, 44(7):736–761, July 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-014-9811-3.

Watson:2014:QPS

[3738] P. Watson and A. J. Bracken. Quantum phase space from Schwinger’s measurement algebra. Foundations of Physics, 44(7):762–780, July 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-014-9813-1.

Cator:2014:CDB

[3739] Eric Cator and Klaas Landsman. Constraints on determinism: Bell versus Conway–Kochen. Foundations of Physics, 44(7):781–791, July 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-014-9815-z.

Chajda:2014:HPT

[3740] Ivan Chajda, Jirí Janda, and Jan Paseka. How to produce S-tense operators on lattice effect algebras. Foundations of Physics, 44(7):792–811, July 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
014-9818-9.

Ohmori:2014:OEQ

[3741] Kenji Ohmori. Optically engineered quantum states in ultrafast and ultracold systems. Foundations of Physics, 44(8):813–818, August 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-014-9773-5.

VanMeter:2014:QCC

[3742] Rodney Van Meter. Quantum computing’s classical problem, classical computing’s quantum problem. Foundations of Physics, 44(8): 819–828, August 2014. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-014-9807-z.

Muller:2014:HIT

[3743] Markus P. Müller, Jonathan Oppenheim, and Oscar C. O. Dahlsten. Hiding information in theories beyond quantum mechanics, and it’s application to the black hole information problem. Foundations of Physics, 44(8):829–842, August 2014. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-014-9809-x.

Crespi:2014:FQP

[3744] Andrea Crespi, Roberto Osellame, Linda Sansoni, Paolo Mataloni, Fabio Sciarrino, and Roberta Ramponi. Fabrication of quantum photonic integrated circuits by means of femtosecond laser pulses. Foundations of Physics, 44(8):843–855, August 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-014-9800-6.

Beausoleil:2014:LSI

[3745] Raymond G. Beausoleil. Large scale integrated photonics for twentyfirst century information technologies. Foundations of Physics, 44(8): 856–872, August 2014. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-013-9771-z.

Penrose:2014:GQMb

[3746] Roger Penrose. On the gravitization of quantum mechanics 2: Conformal cyclic cosmology. Foundations of Physics, 44(8):873–890, August 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10701-013-9763-z.pdf.

Scardigli:2014:PPH

[3747] Fabio Scardigli. Posters presented at Horizons Workshop. Foundations of Physics, 44(8):891–904, August 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-013-9769-6.

Dolce:2014:RQR

[3748] Donatello Dolce and Andrea Perali. The role of quantum recurrence in superconductivity, carbon nanotubes and related gauge symmetry breaking. Foundations of Physics, 44(9):905–922, September 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-014-9816-y.

Gonzalez:2014:NDP

[3749] Diego González, Sergio Davis, and Gonzalo Gutiérrez. Newtonian dynamics from the Principle of Maximum Caliber. Foundations of Physics, 44(9):923–931, September 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-014-9819-8.

Odagiri:2014:SMG

[3750] Kosuke Odagiri. Standard model gauge couplings from gauge-dilatation symmetry breaking. Foundations of Physics, 44(9):932–952, September 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
014-9820-2.

Cheng:2014:CPC

[3751] Jian-Yuan Cheng. A complete proof of the confinement limit of one-dimensional Dirac particles. Foundations of Physics, 44(9):953– 959, September 2014. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-014-9821-1.

Kohler:2014:HOT

[3752] H. S. Köhler. Harmonic oscillator trap and the phase-shift approximation. Foundations of Physics, 44(9):960–972, September 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-014-9822-0.

Tomazelli:2014:MOA

[3753] J. L. Tomazelli and G. A. M. A. Fernandes. Majorana–Oppenheimer approach to Proca field equations. Foundations of Physics, 44(9):973– 989, September 2014. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-014-9824-y.

Castro:2014:CSR

[3754] Carlos Castro. On Clifford space relativity, black hole entropy, rainbow metrics, generalized dispersion and uncertainty relations. Foundations of Physics, 44(9):990–1008, September 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-014-9825-x.

Yang:2014:QMF

[3755] Aili Yang and Yongjian Xie. Quantum measures on finite effect algebras with the Riesz decomposition properties. Foundations of Physics, 44(10): 1009–1037, October 2014. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-014-9826-9.

Hermens:2014:CKF

[3756] Ronnie Hermens. Conway–Kochen and the finite precision loophole. Foundations of Physics, 44(10):1038–1048, October 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-014-9827-8.

Alhaidari:2014:RSA

[3757] A. D. Alhaidari. Renormalization of the strongly attractive inverse square potential: Taming the singularity. Foundations of Physics, 44(10): 1049–1058, October 2014. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-014-9828-7.

Cour:2014:LDD

[3758] Brian R. La Cour. A locally deterministic, detector-based model of quantum measurement. Foundations of Physics, 44(10):1059–1084, October 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10701-014-9829-6.pdf.

Waegell:2014:PPK

[3759] Mordecai Waegell and P. K. Aravind. Parity proofs of the Kochen– Specker theorem based on the 120-cell. Foundations of Physics, 44(10): 1085–1095, October 2014. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-014-9830-0.

deGosson:2014:BJQ

[3760] Maurice A. de Gosson. Born–Jordan quantization and the equivalence of the Schrödinger and Heisenberg pictures. Foundations of Physics, 44(10): 1096–1106, October 2014. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s10701-014-9831-z.pdf.

Baumgartner:2014:CES

[3761] Bernhard Baumgartner. Characterizing entropy in statistical physics and in quantum information theory. Foundations of Physics, 44(10): 1107–1123, October 2014. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-014-9832-y.

Spekkens:2014:SDP

[3762] Robert W. Spekkens. The status of determinism in proofs of the impossibility of a noncontextual model of quantum theory. Foundations of Physics, 44(11):1125–1155, November 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-014-9833-x.

Bachar:2014:LIB

[3763] Yossi Bachar, Rafael I. Arshansky, Lawrence P. Horwitz, and Igal Aharonovich. Lorentz invariant Berry phase for a perturbed relativistic four dimensional harmonic oscillator. Foundations of Physics, 44(11): 1156–1167, November 2014. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-014-9834-9.

Hurley:2014:MFN

[3764] D. J. Hurley and M. A. Vandyck. A minimal framework for noncommutative quantum mechanics. Foundations of Physics, 44(11):1168– 1187, November 2014. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-014-9835-8.

Hassanabadi:2014:RFG

[3765] H. Hassanabadi, A. Armat, and L. Naderi. Relativistic Fermi-gas model for nucleus. Foundations of Physics, 44(11):1188–1194, November 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-014-9836-7.

Wallden:2014:CIC

[3766] Petros Wallden. Contrary inferences in consistent histories and a set selection criterion. Foundations of Physics, 44(11):1195–1215, November 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
014-9837-6.

Niestegge:2014:GQT

[3767] Gerd Niestegge. A generalized quantum theory. Foundations of Physics, 44(11):1216–1229, November 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-014-9838-5.

Banik:2014:OMP

[3768] Manik Banik, Some Sankar Bhattacharya, Sujit K. Choudhary, Amit Mukherjee, and Arup Roy. Ontological models, preparation contextuality and nonlocality. Foundations of Physics, 44(11):1230–1244, November 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
014-9839-4.

Dieks:2014:PSI

[3769] Dennis Dieks, Décio Krause, and Christian de Ronde. Preface special issue foundations of physics. Foundations of Physics, 44(12):1245, December 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10701-014-9853-6.pdf.

daCosta:2014:QMO

[3770] Newton da Costa and Olimpia Lombardi. Quantum mechanics: Ontology without individuals. Foundations of Physics, 44(12):1246–1257, December 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
014-9793-1.

Fortin:2014:QDL

[3771] Sebastian Fortin and Leonardo Vanni. Quantum decoherence: a logical perspective. Foundations of Physics, 44(12):1258–1268, December 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-014-9805-1.

Krause:2014:SNI

[3772] Décio Krause and Jonas R. B. Arenhart. Separability and nonindividuality: Is it possible to conciliate (at least a form of) Einstein’s realism with quantum mechanics? Foundations of Physics, 44(12): 1269–1288, December 2014. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-014-9808-y.

Page:2014:TGL

[3773] Julien Page and Gabriel Catren. Towards a Galoisian interpretation of Heisenberg indeterminacy principle. Foundations of Physics, 44(12): 1289–1301, December 2014. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-014-9812-2.

Dieks:2014:LID

[3774] Dennis Dieks. The logic of identity: Distinguishability and indistinguishability in classical and quantum physics. Foundations of Physics, 44(12):1302–1316, December 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-014-9814-0.

Catren:2014:RBG

[3775] Gabriel Catren. On the relation between gauge and phase symmetries. Foundations of Physics, 44(12):1317–1335, December 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-014-9817-x.

Roldan-Charria:2014:ICO

[3776] Jairo Roldán-Charria. Indivisibility, complementarity and ontology: a Bohrian interpretation of quantum mechanics. Foundations of Physics, 44(12):1336–1356, December 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-014-9823-z.

Freytes:2014:PPM

[3777] H. Freytes, G. Domenech, and C. de Ronde. Physical properties as modal operators in the topos approach to quantum mechanics. Foundations of Physics, 44(12):1357–1368, December 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-014-9842-9.

daCosta:2014:NRL

[3778] N. C. A. da Costa and C. de Ronde. Non-reflexive logical foundation for quantum mechanics. Foundations of Physics, 44(12):1369–1380, December 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
014-9848-3.

Smolin:2015:USD

[3779] Lee Smolin. Unification of the state with the dynamical law. Foundations of Physics, 45(1):1–10, January 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-014-9855-4.

DellAntonio:2015:TCC

[3780] G. F. Dell’Antonio. On tracks in a cloud chamber. Foundations of Physics, 45(1):11–21, January 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-014-9850-9.

Consoli:2015:PVP

[3781] Maurizio Consoli. Probing the vacuum of particle physics with precise laser interferometry. Foundations of Physics, 45(1):22–43, January 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-014-9849-2.

Oikonomou:2015:LFS

[3782] V. K. Oikonomou. Localized fermions on superconducting domain walls and extended supersymmetry with non-trivial topological charges. Foundations of Physics, 45(1):44–61, January 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-014-9844-7.

Kitajima:2015:ICO

[3783] Yuichiro Kitajima. Imperfect cloning operations in algebraic quantum theory. Foundations of Physics, 45(1):62–74, January 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-014-9843-8.

Alsid:2015:UGR

[3784] Scott Alsid and Mario Serna. Unifying geometrical representations of gauge theory. Foundations of Physics, 45(1):75–103, January 2015. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-014-9841-x.

Fiscaletti:2015:PNO

[3785] Davide Fiscaletti and Amrit Sorli. Perspectives of the numerical order of material changes in timeless approaches in physics. Foundations of Physics, 45(2):105–133, February 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-014-9840-y.

Holland:2015:QPE

[3786] Peter Holland. Quantum potential energy as concealed motion. Foundations of Physics, 45(2):134–141, February 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-014-9852-7.

Vatsya:2015:FST

[3787] S. R. Vatsya. Formulation of spinors in terms of gauge fields. Foundations of Physics, 45(2):142–157, February 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-014-9854-5.

Johal:2015:RHE

[3788] Ramandeep S. Johal, Renuka Rai, and Günter Mahler. Reversible heat engines: Bounds on estimated efficiency from inference. Foundations of Physics, 45(2):158–170, February 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-014-9856-3.

DiCasola:2015:BQC

[3789] Eolo Di Casola, Stefano Liberati, and Sebastiano Sonego. Between quantum and classical gravity: Is there a mesoscopic spacetime? Foundations of Physics, 45(2):171–176, February 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-014-9859-0.

Gadella:2015:DPG

[3790] M. Gadella. A discussion on the properties of Gamow states. Foundations of Physics, 45(2):177–197, February 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-014-9860-7.

Li:2015:EIT

[3791] Haichao Li, Guoqin Ge, Lingmin Liao, and Shunbin Feng. Electromagnetically induced transparency and Autler–Townes splitting in a superconducting quantum circuit with a four-level V -type energy spectrum. Foundations of Physics, 45(2):198–210, February 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-014-9861-6.

Kent:2015:DIM

[3792] Adrian Kent. Does it make sense to speak of self-locating uncertainty in the universal wave function? Remarks on Sebens and Carroll. Foundations of Physics, 45(2):211–217, February 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-014-9862-5.

Schroer:2015:HSS

[3793] Bert Schroer. A Hilbert space setting for interacting higher spin fields and the Higgs issue. Foundations of Physics, 45(3):219–252, March 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-015-9867-8.

Buric:2015:RBD

[3794] Nikola Burić. Relations between different notions of degrees of freedom of a quantum system and its classical model. Foundations of Physics, 45 (3):253–278, March 2015. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-014-9858-1.

Jannes:2015:CML

[3795] Gil Jannes. Condensed matter lessons about the origin of time. Foundations of Physics, 45(3):279–294, March 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-014-9864-3.

Garcia-Morales:2015:QMP

[3796] Vladimir Garcia-Morales. Quantum mechanics and the principle of least radix economy. Foundations of Physics, 45(3):295–332, March 2015. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-015-9865-x.

Huang:2015:DES

[3797] Wayne Cheng-Wei Huang and Herman Batelaan. Discrete excitation spectrum of a classical harmonic oscillator in zero–point radiation. Foundations of Physics, 45(3):333–353, March 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-015-9866-9.

Guo:2015:STT

[3798] Zhihua Guo, Huaixin Cao, and Shixian Qu. Structures of three types of local quantum channels based on quantum correlations. Foundations of Physics, 45(4):355–369, April 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-015-9874-9.

Bennett:2015:SSC

[3799] A. F. Bennett. Spin-statistics connection for relativistic quantum mechanics. Foundations of Physics, 45(4):370–381, April 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-015-9869-6.

Moyer:2015:TQT

[3800] Curt A. Moyer. Timelines and quantum time operators. Foundations of Physics, 45(4):382–403, April 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-015-9870-0.

Uffink:2015:LTE

[3801] Jos Uffink and Giovanni Valente. Lanford’s theorem and the emergence of irreversibility. Foundations of Physics, 45(4):404–438, April 2015. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-015-9871-z.

Drory:2015:FUJ

[3802] Alon Drory. Failure and uses of Jaynes’ principle of transformation groups. Foundations of Physics, 45(4):439–460, April 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-015-9876-7.

Okon:2015:BHI

[3803] Elias Okon and Daniel Sudarsky. The black hole information paradox and the collapse of the wave function. Foundations of Physics, 45(4):461–470, April 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
015-9877-6.

Esfeld:2015:BTI

[3804] Michael Esfeld. Bell’s theorem and the issue of determinism and indeterminism. Foundations of Physics, 45(5):471–482, May 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-015-9883-8.

Busch:2015:PPM

[3805] Paul Busch. Philosophical problems of modern physics: Peter Mittelstaedt 1929–2014. Foundations of Physics, 45(5):483–495, May 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-015-9872-y.

Laraudogoitia:2015:AJH

[3806] Jon Pérez Laraudogoitia. On the Atkinson–Johnson homogeneous solution for infinite systems. Foundations of Physics, 45(5):496–506, May 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
015-9878-5.

Falek:2015:KPB

[3807] M. Falek, M. Merad, and M. Moumni. Klein paradox for the bosonic equation in the presence of minimal length. Foundations of Physics, 45(5):507–524, May 2015. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-015-9880-y.

Niestegge:2015:DCG

[3808] Gerd Niestegge. Dynamical correspondence in a generalized quantum theory. Foundations of Physics, 45(5):525–534, May 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-015-9881-x.

Feyereisen:2015:HWV

[3809] M. R. Feyereisen. How the weak variance of momentum can turn out to be negative. Foundations of Physics, 45(5):535–556, May 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s10701-015-9885-6.pdf.

Kochen:2015:RQM

[3810] Simon Kochen. A reconstruction of quantum mechanics. Foundations of Physics, 45(5):557–590, May 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-015-9886-5.

Jaaskelainen:2015:WFM

[3811] Markku Jääskeläinen. The wave function as matter density: Ontological assumptions and experimental consequences. Foundations of Physics, 45(6):591–610, June 2015. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-015-9884-7.

Pollock:2015:GGM

[3812] M. D. Pollock. On the gyromagnetic and gyrogravito–magnetic ratios of the electron. Foundations of Physics, 45(6):611–643, June 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-015-9887-4.

Chuprikov:2015:SDE

[3813] Nikolay L. Chuprikov. The stationary Dirac equation as a generalized Pauli equation for two quasiparticles. Foundations of Physics, 45 (6):644–656, June 2015. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-015-9888-3.

deHaro:2015:BRC

[3814] Sebastian de Haro, Jans Henke, and Darrell Tang. Book review: A. Cappelli, E. Castellani, F. Colomo and P. Di Vecchia (Eds.): The Birth of String Theory. Foundations of Physics, 45(6):657–660, June 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-015-9900-y.

Dreisigmeyer:2015:NLM

[3815] David W. Dreisigmeyer and Peter M. Young. Nonconservative Lagrangian mechanics: Purely causal equations of motion. Foundations of Physics, 45(6):661–672, June 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-015-9892-7.

Englman:2015:OSD

[3816] Robert Englman and Asher Yahalom. Open systems’ density matrix properties in a time coarsened formalism. Foundations of Physics, 45 (6):673–690, June 2015. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-015-9894-5.

Vaccaro:2015:VUT

[3817] Joan A. Vaccaro. T violation and the unidirectionality of time: Further details of the interference. Foundations of Physics, 45(6):691–706, June 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
015-9896-3.

Khrennikov:2015:PSI

[3818] Andrei Khrennikov, Hans de Raedt, Arkady Plotnitsky, and Sergey Polyakov. Preface of the special issue probing the limits of quantum mechanics: Theory and experiment, volume 1. Foundations of Physics, 45(7):707–710, July 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10701-015-9911-8.pdf.

Khrennikov:2015:CIQ

[3819] Andrei Khrennikov. CHSH inequality: Quantum probabilities as classical conditional probabilities. Foundations of Physics, 45(7):711–725, July 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
014-9851-8.

Peters:2015:EBC

[3820] Joffrey K. Peters, Jingyun Fan, Alan L. Migdall, and Sergey V. Polyakov. Experimental bounds on classical random field theories. Foundations of Physics, 45(7):726–734, July 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-014-9857-2.

Kupczynski:2015:BIE

[3821] Marian Kupczynski. Bell inequalities, experimental protocols and contextuality. Foundations of Physics, 45(7):735–753, July 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-014-9863-4.

Marchildon:2015:WQ

[3822] Louis Marchildon. Why I am not a QBist. Foundations of Physics, 45(7): 754–761, July 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10701-015-9875-8.pdf.

Dzhafarov:2015:CTT

[3823] Ehtibar N. Dzhafarov, Janne V. Kujala, and Jan-Åke Larsson. Contextuality in three types of quantum-mechanical systems. Foundations of Physics, 45(7):762–782, July 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-015-9882-9.

Chernega:2015:DEI

[3824] Vladimir N. Chernega, Olga V. Man’ko, and Vladimir I. Man’ko. Deformed entropy and information relations for composite and noncomposite systems. Foundations of Physics, 45(7):783–798, July 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-015-9890-9.

Rosinger:2015:FDL

[3825] Elemér E. Rosinger. Five departures in logic, mathematics, and thuswhether we like it, or not-in physics as well …. Foundations of Physics, 45(7):799–805, July 2015. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-015-9891-8.

Jaeger:2015:MFP

[3826] Gregg Jaeger. Measurement and fundamental processes in quantum mechanics. Foundations of Physics, 45(7):806–819, July 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-015-9893-6.

Khots:2015:LCM

[3827] Boris Khots and Dmitriy Khots. Lagrangian in classical mechanics and in special relativity from observer’s mathematics point of view. Foundations of Physics, 45(7):820–826, July 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-015-9895-4.

Roknizadeh:2015:CMG

[3828] R. Roknizadeh and H. Heydari. Complexifier method for generation of coherent states of nonlinear harmonic oscillator. Foundations of Physics, 45(7):827–839, July 2015. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-015-9901-x.

Loubenets:2015:CIL

[3829] Elena R. Loubenets. Context-invariant and local quasi hidden variable (qHV) modelling versus contextual and nonlocal HV modelling. Foundations of Physics, 45(7):840–850, July 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-015-9903-8.

Heydari:2015:GSQ

[3830] Hoshang Heydari. Geometry and structure of quantum phase space. Foundations of Physics, 45(7):851–857, July 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-015-9907-4.

Santamato:2015:PSS

[3831] Enrico Santamato and Francesco De Martini. Proof of the spin–statistics theorem. Foundations of Physics, 45(7):858–873, July 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-015-9912-7.

Eingorn:2015:IDQ

[3832] Maxim V. Eingorn and Vitaliy D. Rusov. Inflation due to quantum potential. Foundations of Physics, 45(8):875–882, August 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-015-9897-2.

Gray:2015:MVQ

[3833] John E. Gray and Andrew Vogt. Mean and variance in quantum theory. Foundations of Physics, 45(8):883–888, August 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-015-9898-1.

Kremnizer:2015:III

[3834] Kobi Kremnizer and André Ranchin. Integrated information-induced quantum collapse. Foundations of Physics, 45(8):889–899, August 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-015-9905-6.

Pykacz:2015:BTI

[3835] Jarosław Pykacz, L’ubica Valásková, and Ol’ga Nánásiová. Bell-type inequalities for bivariate maps on orthomodular lattices. Foundations of Physics, 45(8):900–913, August 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10701-015-9906-5.pdf.

Majima:2015:IEC

[3836] Hiroki Majima and Akira Suzuki. Identities for entropy change associated with the time-evolution of an open system. Foundations of Physics, 45 (8):914–922, August 2015. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-015-9908-3.

Rastegin:2015:URG

[3837] Alexey E. Rastegin. Uncertainty relations for general canonically conjugate observables in terms of unified entropies. Foundations of Physics, 45 (8):923–942, August 2015. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-015-9909-2.

Zaopo:2015:ITC

[3838] Marco Zaopo. Information theoretic characterization of physical theories with projective state space. Foundations of Physics, 45(8):943–958, August 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
015-9910-9.

Doyen:2015:CQM

[3839] G. Doyen and D. Drakova. Copenhagen quantum mechanics emerges from a deterministic Schrödinger theory in 11 dimensional spacetime including weak field gravitation. Foundations of Physics, 45(8):959– 999, August 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10701-015-9914-5.pdf.

Toader:2015:ENP

[3840] Iulian D. Toader. Editorial: New perspectives in the philosophy of physics. Foundations of Physics, 45(9):1001, September 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
015-9924-3;   http://link.springer.com/content/pdf/10.1007/
s10701-015-9924-3.pdf.

Held:2015:EBI

[3841] Carsten Held. Einstein’s boxes: Incompleteness of quantum mechanics without a separation principle. Foundations of Physics, 45(9):1002– 1018, September 2015. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-014-9845-6.

Crull:2015:LIM

[3842] Elise M. Crull. Less interpretation and more decoherence in quantum gravity and inflationary cosmology. Foundations of Physics, 45(9):1019– 1045, September 2015. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-014-9847-4. See reply [3872].

Wolff:2015:OVQ

[3843] Johanna Wolff. Observability, visualizability and the question of metaphysical neutrality. Foundations of Physics, 45(9):1046–1062, September 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
015-9904-7.

Shech:2015:TAF

[3844] Elay Shech. Two approaches to fractional statistics in the quantum Hall effect: Idealizations and the curious case of the anyon. Foundations of Physics, 45(9):1063–1100, September 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-015-9899-0.

Dawid:2015:TND

[3845] Richard Dawid. Turning Norton’s dome against material induction. Foundations of Physics, 45(9):1101–1109, September 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-015-9943-0.

Hofer-Szabo:2015:RBL

[3846] Gábor Hofer-Szabó. Relating Bell’s local causality to the causal Markov condition. Foundations of Physics, 45(9):1110–1136, September 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-015-9868-7.

Bisio:2015:FQF

[3847] Alessandro Bisio, Giacomo Mauro D’Ariano, and Paolo Perinotti. Free quantum field theory from quantum cellular automata. Foundations of Physics, 45(10):1137–1152, October 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-015-9934-1.

Sanz:2015:IPA

[3848] A. S. Sanz. Investigating puzzling aspects of the quantum theory by means of its hydrodynamic formulation. Foundations of Physics, 45(10):

1153–1165, October 2015. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-015-9917-2.

Gooding:2015:BTD

[3849] Cisco Gooding and William G. Unruh. Bootstrapping time dilation decoherence. Foundations of Physics, 45(10):1166–1178, October 2015. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-015-9939-9.

Khrennikov:2015:UQC

[3850] Andrei Khrennikov. Unconditional quantum correlations do not violate Bell’s inequality. Foundations of Physics, 45(10):1179–1189, October 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
015-9930-5.

Nieuwenhuizen:2015:SHG

[3851] Theodorus M. Nieuwenhuizen and Matthew T. P. Liska. Simulation of the hydrogen ground state in stochastic electrodynamics-2: Inclusion of relativistic corrections. Foundations of Physics, 45(10):1190–1202, October 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
015-9919-0.

Bisio:2015:WDM

[3852] Alessandro Bisio, Giacomo Mauro D’Ariano, and Paolo Perinotti. Weyl, Dirac and Maxwell quantum cellular automata. Foundations of Physics, 45(10):1203–1221, October 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-015-9927-0.

Plotnitsky:2015:MPP

[3853] Arkady Plotnitsky. A matter of principle: The principles of quantum theory, Dirac’s equation, and quantum information. Foundations of Physics, 45(10):1222–1268, October 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-015-9928-z.

Plotnitsky:2015:RRO

[3854] Arkady Plotnitsky and Andrei Khrennikov. Reality without realism: On the ontological and epistemological architecture of quantum mechanics. Foundations of Physics, 45(10):1269–1300, October 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-015-9942-1.

Pombo:2015:DSP

[3855] Claudia Pombo. Differentiation with stratification: a principle of theoretical physics in the tradition of the memory art. Foundations of Physics, 45(10):1301–1310, October 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-015-9936-z.

Watanabe:2015:EQC

[3856] Noboru Watanabe. On entropy of quantum compound systems. Foundations of Physics, 45(10):1311–1329, October 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-015-9925-2.

Dickson:2015:RRQ

[3857] Michael Dickson. Reconstruction and reinvention in quantum theory. Foundations of Physics, 45(10):1330–1340, October 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-015-9946-x.

Grinbaum:2015:QTC

[3858] Alexei Grinbaum. Quantum theory as a critical regime of language dynamics. Foundations of Physics, 45(10):1341–1350, October 2015. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-015-9937-y.

Stairs:2015:QLQ

[3859] Allen Stairs. Quantum logic and quantum reconstruction. Foundations of Physics, 45(10):1351–1361, October 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-015-9879-4.

Asano:2015:QIB

[3860] Masanari Asano, Irina Basieva, Andrei Khrennikov, and Masanori Ohya. Quantum information biology: From information interpretation of quantum mechanics to applications in molecular biology and cognitive psychology. Foundations of Physics, 45(10):1362–1378, October 2015. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-015-9929-y.

Basieva:2015:PCO

[3861] Irina Basieva and Andrei Khrennikov. On the possibility to combine the order effect with sequential reproducibility for quantum measurements. Foundations of Physics, 45(10):1379–1393, October 2015. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-015-9932-3.

Haven:2015:PFC

[3862] Emmanuel Haven. Potential functions and the characterization of economics-based information. Foundations of Physics, 45(10):1394– 1406, October 2015. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-015-9931-4.

Angelo:2015:WPD

[3863] R. M. Angelo and A. D. Ribeiro. Wave–particle duality: an informationbased approach. Foundations of Physics, 45(11):1407–1420, November 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
015-9913-6.

Faizal:2015:SDD

[3864] Mir Faizal and Tsou Sheung Tsun. Supersymmetric duality in deformed superloop space. Foundations of Physics, 45(11):1421–1432, November 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
015-9915-4.

Garcia-Escartin:2015:SCT

[3865] Juan Carlos Garcia-Escartin and Pedro Chamorro-Posada. Secure communication in the twin paradox. Foundations of Physics, 45(11):1433– 1453, November 2015. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-015-9916-3.

Sutherland:2015:LDP

[3866] Roderick I. Sutherland. Lagrangian description for particle interpretations of quantum mechanics: Single–particle case. Foundations of Physics, 45(11):1454–1464, November 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-015-9918-1.

Biro:2015:SST

[3867] T. S. Biró and P. Ván. Splitting the source term for the Einstein equation to classical and quantum parts. Foundations of Physics, 45(11): 1465–1482, November 2015. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-015-9920-7.

Olavo:2015:FQM

[3868] L. S. F. Olavo. Foundations of quantum mechanics: On rotations by 4π for half-integral spin particles. Foundations of Physics, 45(11):1483– 1494, November 2015. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-015-9921-6.

Dariescu:2015:QAR

[3869] Ciprian Dariescu and Marina-Aura Dariescu. Quantum analysis of k = −1 Robertson–Walker universe. Foundations of Physics, 45(11): 1495–1513, November 2015. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-015-9922-5.

Heifetz:2015:TTH

[3870] Eyal Heifetz and Eliahu Cohen. Toward a thermo-hydrodynamic like description of Schrödinger equation via the Madelung formulation and Fisher information. Foundations of Physics, 45(11):1514–1525, November 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
015-9926-1.

Bersons:2015:NME

[3871] I. Bersons and R. Veilande. Nonlinear models of electric charge and magnetic moment. Foundations of Physics, 45(11):1526–1532, November 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
015-9940-3.

Vassallo:2015:IIQ

[3872] Antonio Vassallo and Michael Esfeld. On the importance of interpretation in quantum physics: a reply to Elise Crull. Foundations of Physics, 45(12):1533–1536, December 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-015-9902-9. See [3842, 4039, 4040].

Bera:2015:CBM

[3873] Sayantani Bera, Sandro Donadi, Kinjalk Lochan, and Tejinder P. Singh. A comparison between models of gravity induced decoherence. Foundations of Physics, 45(12):1537–1560, December 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-015-9933-2.

Frenkel:2015:DTR

[3874] Andor Frenkel. Dependence of the time-reading process of the Salecker– Wigner quantum clock on the size of the clock. Foundations of Physics, 45(12):1561–1573, December 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-015-9938-x.

Krishnan:2015:GG

[3875] Chethan Krishnan. A generalization of gravity. Foundations of Physics, 45(12):1574–1585, December 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-015-9941-2.

Ajaib:2015:FFS

[3876] Muhammad Adeel Ajaib. A fundamental form of the Schrödinger equation. Foundations of Physics, 45(12):1586–1598, December 2015. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-015-9944-z.

Stergiou:2015:ECP

[3877] Chrysovalantis Stergiou. Explaining correlations by partitions. Foundations of Physics, 45(12):1599–1612, December 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-015-9945-y.

Sych:2015:QU

[3878] Denis Sych and Gerd Leuchs. Quantum uniqueness. Foundations of Physics, 45(12):1613–1619, December 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-015-9947-9.

Gould:2015:NLR

[3879] Elizabeth S. Gould and Niayesh Afshordi. A non-local reality: Is there a phase uncertainty in quantum mechanics? Foundations of Physics, 45(12):1620–1644, December 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-015-9948-8.

Svensson:2015:NRQ

[3880] B. E. Y. Svensson. Non-representative quantum mechanical weak values. Foundations of Physics, 45(12):1645–1656, December 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://

link.springer.com/article/10.1007/s10701-015-9951-0. See com-

ment [3993] and response [4031].

tHooft:2016:E

[3881] Gerard ’t Hooft, Carlo Rovelli, and Angela Lahee. Editorial. Foundations of Physics, 46(1):1–2, January 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10701-015-9975-5.pdf.

Niehaus:2016:PMS

[3882] Arend Niehaus. A probabilistic model of spin and spin measurements. Foundations of Physics, 46(1):3–13, January 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-015-9953-y.

Perelman:2016:NRP

[3883] Carlos Castro Perelman. Novel remarks on point mass sources, firewalls, null singularities and gravitational entropy. Foundations of Physics, 46 (1):14–27, January 2016. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-015-9954-x.

Drieschner:2016:PRF

[3884] Michael Drieschner. Probability and relative frequency. Foundations of Physics, 46(1):28–43, January 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-015-9955-9.

Fearn:2016:DCQ

[3885] H. Fearn. A delayed choice quantum eraser explained by the transactional interpretation of quantum mechanics. Foundations of Physics, 46(1):44– 69, January 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10701-015-9956-8.pdf.

Backens:2016:CGC

[3886] Miriam Backens and Ali Nabi Duman. A complete graphical calculus for Spekkens’ toy bit theory. Foundations of Physics, 46(1):70–103, January 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
015-9957-7.

Pienaar:2016:SDL

[3887] Jacques Pienaar. A separable, dynamically local ontological model of quantum mechanics. Foundations of Physics, 46(1):104–119, January 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
015-9958-6.

Auffeves:2016:CSM

[3888] Alexia Auffèves and Philippe Grangier. Contexts, systems and modalities: a new ontology for quantum mechanics. Foundations of Physics, 46 (2):121–137, February 2016. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-015-9952-z.

Kaloyerou:2016:CQO

[3889] P. N. Kaloyerou. Critique of quantum optical experimental refutations of Bohr’s principle of complementarity, of the Wootters–Zurek principle of complementarity, and of the particle–wave duality relation. Foundations of Physics, 46(2):138–175, February 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-015-9959-5.

Scholz:2016:MLA

[3890] Erhard Scholz. MOND-like acceleration in integrable Weyl geometric gravity. Foundations of Physics, 46(2):176–208, February 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-015-9960-z.

Read:2016:IST

[3891] James Read. The interpretation of string-theoretic dualities. Foundations of Physics, 46(2):209–235, February 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-015-9961-y.

Kholmetskii:2016:PTR

[3892] Alexander Kholmetskii, Oleg Missevitch, and Tolga Yarman. Poynting theorem, relativistic transformation of total energy–momentum and electromagnetic energy–momentum tensor. Foundations of Physics, 46 (2):236–261, February 2016. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-015-9963-9.

Curceanu:2016:SEX

[3893] C. Curceanu, S. Bartalucci, A. Bassi, M. Bazzi, S. Bertolucci, C. Berucci, A. M. Bragadireanu, M. Cargnelli, A. Clozza, L. De Paolis, S. Di Matteo, S. Donadi, A. D’Uffizi, J.-P. Egger, C. Guaraldo, and et al. Spontaneously emitted X-rays: an experimental signature of the dynamical reduction models. Foundations of Physics, 46(3):263–268, March 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-015-9923-4.

DAriano:2016:QTI

[3894] Giacomo M. D’Ariano and Paolo Perinotti. Quantum theory is an information theory. Foundations of Physics, 46(3):269–281, March 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-015-9935-0.

Kujala:2016:PCC

[3895] Janne V. Kujala and Ehtibar N. Dzhafarov. Proof of a conjecture on contextuality in cyclic systems with binary variables. Foundations of Physics, 46(3):282–299, March 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-015-9964-8.

Kowalski:2016:TPA

[3896] F. V. Kowalski and R. S. Browne. Two-particle asynchronous quantum correlation: Wavefunction collapse acting as a beamsplitter. Foundations of Physics, 46(3):300–329, March 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-015-9965-7.

Hamhalter:2016:BCE

[3897] Jan Hamhalter and Veronika Sobotíková. Bell correlated and EPR states in the framework of Jordan algebras. Foundations of Physics, 46(3): 330–349, March 2016. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-015-9966-6.

Davis:2016:BIF

[3898] Sergio Davis, Joaquín Peralta, Yasmín Navarrete, Diego González, and Gonzalo Gutiérrez. A Bayesian interpretation of first-order phase transitions. Foundations of Physics, 46(3):350–359, March 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-015-9967-5.

Castagnoli:2016:HMQ

[3899] Giuseppe Castagnoli. Highlighting the mechanism of the quantum speedup by time-symmetric and relational quantum mechanics. Foundations of Physics, 46(3):360–381, March 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-015-9968-4.

Grigorenko:2016:PFC

[3900] A. N. Grigorenko. Particles, fields and a canonical distance form. Foundations of Physics, 46(3):382–392, March 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10701-015-9982-6.pdf.

Mauri:2016:PMR

[3901] Roberto Mauri. The principle of minimal resistance in non-equilibrium thermodynamics. Foundations of Physics, 46(4):393–408, April 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-015-9969-3.

Katore:2016:UDB

[3902] S. D. Katore, S. P. Hatkar, and R. J. Baxi. Unified description of Bianchi Type-I universe in f(R) gravity. Foundations of Physics, 46(4):409–427, April 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
015-9970-x.

Chamseddine:2016:GRS

[3903] Riad Chamseddine. Graphical representations for the successive Lorentz transformations. application: Lorentz contraction and its dependence on Thomas rotation. Foundations of Physics, 46(4):428–457, April 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-015-9971-9.

Vervoort:2016:NGT

[3904] Louis Vervoort. No-go theorems face background-based theories for quantum mechanics. Foundations of Physics, 46(4):458–472, April 2016. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-015-9973-7.

Giardino:2016:QPR

[3905] Sergio Giardino. Quaternionic particle in a relativistic box. Foundations of Physics, 46(4):473–483, April 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-015-9974-6.

Lopez:2016:LIQ

[3906] Carlos Lopez. A local interpretation of quantum mechanics. Foundations of Physics, 46(4):484–504, April 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-015-9976-4.

Hossenfelder:2016:RMD

[3907] S. Hossenfelder. The remote Maxwell demon as energy down-converter. Foundations of Physics, 46(4):505–516, April 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-015-9981-7.

Zak:2016:NNA

[3908] Michail Zak. Non-Newtonian aspects of artificial intelligence. Foundations of Physics, 46(5):517–553, May 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-015-9977-3.

Singal:2016:CLF

[3909] Ashok K. Singal. Compatibility of Larmor’s formula with radiation reaction for an accelerated charge. Foundations of Physics, 46(5):554–574, May 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
015-9978-2.

Avanzini:2016:PWQ

[3910] Francesco Avanzini, Barbara Fresch, and Giorgio J. Moro. Pilot-wave quantum theory with a single Bohm’s trajectory. Foundations of Physics, 46(5):575–605, May 2016. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-015-9979-1.

Sbitnev:2016:HPVa

[3911] Valeriy I. Sbitnev. Hydrodynamics of the physical vacuum: I. Scalar quantum sector. Foundations of Physics, 46(5):606–619, May 2016. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-015-9980-8.

Klobus:2016:CSC

[3912] Waldemar Klobus, Michal Oszmaniec, and Remigiusz Augusiak. Communication strength of correlations violating monogamy relations. Foundations of Physics, 46(5):620–634, May 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-015-9983-5.

Zwirn:2016:MPD

[3913] Hervé Zwirn. The measurement problem: Decoherence and convivial solipsism. Foundations of Physics, 46(6):635–667, June 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-016-9999-5.

Khlestkov:2016:ESE

[3914] Yu. A. Khlestkov and L. A. Sukhanova. Exact solutions to the Einstein– Maxwell equations describing wormholes and handles. Foundations of Physics, 46(6):668–688, June 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-016-9998-6.

Torres-Vega:2016:URL

[3915] Gabino Torres-Vega. Universal raising and lowering operators for a discrete energy spectrum. Foundations of Physics, 46(6):689–701, June 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
016-9997-7.

Boddy:2016:SSD

[3916] Kimberly K. Boddy, Sean M. Carroll, and Jason Pollack. De Sitter space without dynamical quantum fluctuations. Foundations of Physics, 46(6):702–735, June 2016. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-016-9996-8.

Smolin:2016:QMP

[3917] Lee Smolin. Quantum mechanics and the principle of maximal variety. Foundations of Physics, 46(6):736–758, June 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-016-9994-x.

Myrvold:2016:QMN

[3918] Wayne C. Myrvold. Quantum mechanics and narratability. Foundations of Physics, 46(7):759–775, July 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-016-0015-x.

Luo:2016:FRY

[3919] Shao-Kai Luo, Jin-Man He, Yan-Li Xu, and Xiao-Tian Zhang. Fractional relativistic Yamaleev oscillator model and its dynamical behaviors. Foundations of Physics, 46(7):776–786, July 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-015-9984-4.

deCordoba:2016:HQ

[3920] P. Fernández de Córdoba, J. M. Isidro, and J. Vazquez Molina. The holographic quantum. Foundations of Physics, 46(7):787–803, July 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-015-9986-2.

Ber:2016:RSP

[3921] Ran Ber and Erez Zohar. Remote state preparation for quantum fields. Foundations of Physics, 46(7):804–814, July 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-016-0001-3.

Heifetz:2016:EPM

[3922] Eyal Heifetz, Roumen Tsekov, Eliahu Cohen, and Zohar Nussinov. On entropy production in the Madelung fluid and the role of Bohm’s potential in classical diffusion. Foundations of Physics, 46(7):815–824, July 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
016-0003-1.

Sverdlov:2016:HMC

[3923] Roman Sverdlov. How Mensky’s continuous measurement can emerge from GRW on larger time scales. Foundations of Physics, 46(7):825–835, July 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
016-0004-0.

Khrennikov:2016:RZB

[3924] Andrei Khrennikov. Reflections on Zeilinger–Brukner information interpretation of quantum mechanics. Foundations of Physics, 46(7):836–844, July 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
016-0005-z.

Ilyin:2016:BRT

[3925] Aleksey V. Ilyin. The Born rule and time-reversal symmetry of quantum equations of motion. Foundations of Physics, 46(7):845–851, July 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-016-0006-y.

Okon:2016:LDM

[3926] Elias Okon and Daniel Sudarsky. Less decoherence and more coherence in quantum gravity, inflationary cosmology and elsewhere. Foundations of Physics, 46(7):852–879, July 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-016-0007-x.

Boyer:2016:CZP

[3927] Timothy H. Boyer. Classical zero-point radiation and relativity: The problem of atomic collapse revisited. Foundations of Physics, 46(7): 880–890, July 2016. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-016-0008-9.

Kuic:2016:PSM

[3928] Domagoj Kuić. Predictive statistical mechanics and macroscopic time evolution: Hydrodynamics and entropy production. Foundations of Physics, 46(7):891–914, July 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-016-0009-8.

Cheng:2016:BEC

[3929] Ze Cheng. Bose–Einstein condensation of nonideal Cooper pairs in the Hartree–Fock–Popov theory. Foundations of Physics, 46(8):915–942, August 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
016-9993-y.

Vassallo:2016:CIS

[3930] Antonio Vassallo and Pui Him Ip. On the conceptual issues surrounding the notion of relational Bohmian dynamics. Foundations of Physics, 46 (8):943–972, August 2016. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-016-9992-z.

Ballentine:2016:PPQ

[3931] Leslie E. Ballentine. Propensity, probability, and quantum theory. Foundations of Physics, 46(8):973–1005, August 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-016-9991-0.

Lush:2016:DBQ

[3932] David C. Lush. Does Bohm’s quantum force have a classical origin? Foundations of Physics, 46(8):1006–1021, August 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-016-9990-1.

Tessarotto:2016:GLP

[3933] Massimo Tessarotto and Claudio Cremaschini. Generalized Lagrangianpath representation of non-relativistic quantum mechanics. Foundations of Physics, 46(8):1022–1061, August 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-016-9989-7.

Fakhri:2016:URO

[3934] H. Fakhri and M. Sayyah-Fard. An uncertainty relation for the orbital angular momentum operator. Foundations of Physics, 46(8): 1062–1073, August 2016. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-016-9988-8.

Slagter:2016:NFW

[3935] Reinoud Jan Slagter and Supriya Pan. A new fate of a warped 5D FLRW model with a U(1) scalar gauge field. Foundations of Physics, 46(9): 1075–1089, September 2016. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-016-0002-2.

Bennett:2016:DKP

[3936] A. F. Bennett. Duffin–Kemmer–Petiau particles are bosons. Foundations of Physics, 46(9):1090–1108, September 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-016-0010-2.

Gondran:2016:RSS

[3937] Michel Gondran and Alexandre Gondran. Replacing the singlet spinor of the EPR-B experiment in the configuration space with two singleparticle spinors in physical space. Foundations of Physics, 46(9):1109– 1126, September 2016. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-016-0011-1.

Tessarotto:2016:HSS

[3938] Massimo Tessarotto, Michael Mond, and Davide Batic. Hamiltonian structure of the Schrödinger classical dynamical system. Foundations of Physics, 46(9):1127–1167, September 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-016-0012-0.

Bigaj:2016:STM

[3939] Tomasz Bigaj. On some troubles with the metaphysics of fermionic compositions. Foundations of Physics, 46(9):1168–1184, September 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s10701-016-0013-z.pdf.

tHooft:2016:BHU

[3940] Gerard ’t Hooft. Black hole unitarity and antipodal entanglement. Foundations of Physics, 46(9):1185–1198, September 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10701-016-0014-y.pdf.

Zhang:2016:FEC

[3941] Anwei Zhang. The formalism for energy changing rate of an accelerated atom coupled with electromagnetic vacuum fluctuations. Foundations of Physics, 46(9):1199–1207, September 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-016-0016-9.

Walleczek:2016:NQI

[3942] Jan Walleczek and Gerhard Grössing. Nonlocal quantum information transfer without superluminal signalling and communication. Foundations of Physics, 46(9):1208–1228, September 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10701-016-9987-9.pdf.

Rovelli:2016:AAR

[3943] Carlo Rovelli. An argument against the realistic interpretation of the wave function. Foundations of Physics, 46(10):1229–1237, October 2016.

CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-016-0032-9.

Sbitnev:2016:HPVb

[3944] Valeriy I. Sbitnev. Hydrodynamics of the physical vacuum: II. Vorticity dynamics. Foundations of Physics, 46(10):1238–1252, October 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-015-9985-3.

Bolotin:2016:CUW

[3945] Arkady Bolotin. Constructibility of the universal wave function. Foundations of Physics, 46(10):1253–1268, October 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-016-0018-7.

Dantas:2016:IST

[3946] Christine C. Dantas. An inhomogeneous space–time patching model based on a nonlocal and nonlinear Schrödinger equation. Foundations of Physics, 46(10):1269–1292, October 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-016-0019-6.

DeZela:2016:GTT

[3947] F. De Zela. Gleason-type theorem for projective measurements, including qubits: The Born rule beyond quantum physics. Foundations of Physics, 46(10):1293–1306, October 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-016-0020-0.

Fiscaletti:2016:ADE

[3948] Davide Fiscaletti. About dark energy and dark matter in a threedimensional quantum vacuum model. Foundations of Physics, 46(10): 1307–1340, October 2016. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-016-0021-z.

Cassinelli:2016:ABQ

[3949] Gianni Cassinelli and Pekka Lahti. An axiomatic basis for quantum mechanics. Foundations of Physics, 46(10):1341–1373, October 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-016-0022-y.

Borghi:2016:PTT

[3950] Claudio Borghi. Physical time and thermal clocks. Foundations of Physics, 46(10):1374–1379, October 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-016-0030-y.

DeHaro:2016:CAG

[3951] Sebastian De Haro, Daniel R. Mayerson, and Jeremy N. Butterfield. Conceptual aspects of gauge/gravity duality. Foundations of Physics, 46(11):1381–1425, November 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10701-016-0037-4.pdf.

Prosperi:2016:ICL

[3952] G. M. Prosperi. Introduction of a classical level in quantum theory. Foundations of Physics, 46(11):1426–1460, November 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-016-0023-x.

Rfifi:2016:ECQ

[3953] Saad Rfifi and Fatimazahra Siyouri. Effect of cavity QED on entanglement. Foundations of Physics, 46(11):1461–1470, November 2016. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-016-0024-9.

Schulman:2016:SSD

[3954] L. S. Schulman. Special states demand a force for the observer. Foundations of Physics, 46(11):1471–1494, November 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10701-016-0025-8.pdf.

Schulman:2016:LSC

[3955] L. S. Schulman and M. G. E. da Luz. Looking for the source of change. Foundations of Physics, 46(11):1495–1501, November 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s10701-016-0031-x.pdf.

Gluza:2016:PWQ

[3956] Janusz Gluza and Jerzy Kosek. Pilot-wave quantum theory in discrete space and time and the principle of least action. Foundations of Physics, 46(11):1502–1521, November 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10701-016-0026-7.pdf.

Miyadera:2016:ETU

[3957] Takayuki Miyadera. Energy-time uncertainty relations in quantum measurements. Foundations of Physics, 46(11):1522–1550, November 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-016-0027-6.

Oliynyk:2016:CQL

[3958] Todd A. Oliynyk. Classical-quantum limits. Foundations of Physics, 46(12):1551–1572, December 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-016-0028-5.

Burde:2016:SRK

[3959] Georgy I. Burde. Special relativity kinematics with anisotropic propagation of light and correspondence principle. Foundations of Physics, 46(12):1573–1597, December 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-016-0029-4.

Kononets:2016:NPC

[3960] Yu. V. Kononets. Novel principles and the charge-symmetric design of Dirac’s quantum mechanics: I. Enhanced Eriksen’s Theorem and the universal charge–index formalism for Dirac’s equation in (strong) external static fields. Foundations of Physics, 46(12):1598–1633, December 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-016-0033-8.

Naeij:2016:DSI

[3961] Hamid Reza Naeij and Afshin Shafiee. Double-slit interference pattern for a macroscopic quantum system. Foundations of Physics, 46(12): 1634–1648, December 2016. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-016-0034-7.

Das:2016:PTK

[3962] Praloy Das and Subir Ghosh. Particle on a torus knot: a Hamiltonian analysis. Foundations of Physics, 46(12):1649–1665, December 2016. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-016-0035-6.

Tkachuk:2016:GLT

[3963] V. M. Tkachuk. Galilean and Lorentz transformations in a space with generalized uncertainty principle. Foundations of Physics, 46(12):1666– 1679, December 2016. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-016-0036-5.

Mamone-Capria:2016:FTT

[3964] Marco Mamone-Capria. On the fundamental theorem of the theory of relativity. Foundations of Physics, 46(12):1680–1712, December 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-016-0038-3.

Risueno:2016:IPC

[3965] Pablo García Risueño, Apostolos Syropoulos, and Natàlia Vergés. Ideograms for physics and chemistry. Foundations of Physics, 46(12): 1713–1721, December 2016. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-016-0044-5.

Helein:2017:CST

[3966] Frédéric Hélein and Dimitri Vey. Curved space–times by crystallization of liquid fiber bundles. Foundations of Physics, 47(1):1–41, January 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-016-0039-2.

Floyd:2017:NON

[3967] Edward R. Floyd. Neutrino oscillations with nil mass. Foundations of Physics, 47(1):42–60, January 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-016-0040-9.

deGosson:2017:AMD

[3968] Maurice A. de Gosson. The angular momentum dilemma and Born– Jordan quantization. Foundations of Physics, 47(1):61–70, January 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s10701-016-0041-8.pdf.

Ferraro:2017:FFS

[3969] Rafael Ferraro. The frame of fixed stars in relational mechanics. Foundations of Physics, 47(1):71–88, January 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-016-0042-7.

Lee:2017:HOI

[3970] Ciarán M. Lee and John H. Selby. Higher-order interference in extensions of quantum theory. Foundations of Physics, 47(1):89–112, January 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10701-016-0045-4.pdf.

Schmelzer:2017:ANL

[3971] I. Schmelzer. About a “nonlocal” local model considered by L. Vervoort, and the necessity to distinguish locality from Einstein locality. Foundations of Physics, 47(1):113–116, January 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-016-0046-3.

Schmelzer:2017:ACL

[3972] I. Schmelzer. About a “contextuality loophole” in Bell’s Theorem claimed to exist by nieuwenhuizen. Foundations of Physics, 47(1):117–119, January 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
016-0047-2.

Okon:2017:BHI

[3973] Elias Okon and Daniel Sudarsky. Black holes, information loss and the measurement problem. Foundations of Physics, 47(1):120–131, January 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
016-0048-1.

Kellner:2017:PKM

[3974] Jakob Kellner. Pitowsky’s Kolmogorovian models and super-determinism. Foundations of Physics, 47(1):132–148, January 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s10701-016-0049-0.pdf.

Smilga:2017:TCF

[3975] Walter Smilga. Towards a constructive foundation of quantum mechanics. Foundations of Physics, 47(1):149–159, January 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-016-0050-7.

Healey:2017:QSO

[3976] Richard Healey. Quantum states as objective informational bridges. Foundations of Physics, 47(2):161–173, February 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-015-9949-7.

Sutherland:2017:LDP

[3977] Roderick I. Sutherland. Lagrangian description for particle interpretations of quantum mechanics: Entangled many-particle case. Foundations of Physics, 47(2):174–207, February 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-016-0043-6.

Cabral:2017:ESG

[3978] Francisco Cabral and Francisco S. N. Lobo. Electrodynamics and spacetime geometry: Foundations. Foundations of Physics, 47(2):208– 228, February 2017. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-016-0051-6.

Goswami:2017:URI

[3979] Ashutosh K. Goswami and Prasanta K. Panigrahi. Uncertainty relation and inseparability criterion. Foundations of Physics, 47(2):229– 235, February 2017. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-016-0052-5.

Fukuda:2017:WVQ

[3980] Kazuki Fukuda, Jaeha Lee, and Izumi Tsutsui. Weak value, quasiprobability and Bohmian mechanics. Foundations of Physics, 47(2):236– 255, February 2017. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-016-0054-3.

Paillet:2017:SRS

[3981] J. L. Paillet and A. Meulenberg. Special relativity, the source of electron deep orbits. Foundations of Physics, 47(2):256–264, February 2017. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-016-0055-2.

Bandyopadhyay:2017:ROM

[3982] Somshubhro Bandyopadhyay, Manik Banik, Some Sankar Bhattacharya, Sibasish Ghosh, Guruprasad Kar, Amit Mukherjee, and Arup Roy. Reciprocal ontological models show indeterminism comparable to quantum theory. Foundations of Physics, 47(2):265–273, February 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-016-0058-z.

Niestegge:2017:QTG

[3983] Gerd Niestegge. Quantum teleportation and Grover’s algorithm without the wavefunction. Foundations of Physics, 47(2):274–293, February 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-016-0060-5.

Feintzeig:2017:NNK

[3984] Benjamin H. Feintzeig and Samuel C. Fletcher. On noncontextual, non-Kolmogorovian hidden variable theories. Foundations of Physics, 47(2): 294–315, February 2017. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-017-0061-z.

Nieuwenhuizen:2017:CLF

[3985] Theodorus M. Nieuwenhuizen and Marian Kupczynski. The contextuality loophole is fatal for the derivation of Bell inequalities: Reply to a comment by I. Schmelzer. Foundations of Physics, 47(2):316–319, February 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
017-0062-y.

Lambare:2017:CFB

[3986] Justo Pastor Lambare. On the CHSH form of Bell’s inequalities. Foundations of Physics, 47(3):321–326, March 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/s10701-017-0063-x.

Davidson:2017:MER

[3987] Mark Davidson. On the Mössbauer effect and the rigid recoil question. Foundations of Physics, 47(3):327–354, March 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-017-0064-9.

Seidewitz:2017:AHT

[3988] Ed Seidewitz. Avoiding Haag’s theorem with parameterized quantum field theory. Foundations of Physics, 47(3):355–374, March 2017. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-017-0065-8.

Brun:2017:SCT

[3989] Todd A. Brun and Mark M. Wilde. Simulations of closed timelike curves. Foundations of Physics, 47(3):375–391, March 2017. CODEN FNDPA4.

ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-017-0066-7.

Floyd:2017:AQE

[3990] Edward R. Floyd. Action quantization, energy quantization, and time parametrization. Foundations of Physics, 47(3):392–429, March 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-017-0067-6.

Svensson:2017:QWV

[3991] Bengt E. Y. Svensson. Quantum weak values and logic: an uneasy couple. Foundations of Physics, 47(3):430–452, March 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s10701-017-0068-5.pdf.

Egg:2017:PCI

[3992] Matthias Egg, Vincent Lam, and Andrea Oldofredi. Particles, cutoffs and inequivalent representations. Foundations of Physics, 47(3):453–466, March 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10701-
017-0069-4.

Ben-Israel:2017:CNR

[3993] Alon Ben-Israel and Lev Vaidman. Comment on “Non-representative Quantum Mechanical Weak Values”. Foundations of Physics, 47(4): 467–470, April 2017. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.

1007/s10701-017-0071-x. See [3880] and response [4031].

Ardourel:2017:IDB

[3994] Vincent Ardourel. Irreversibility in the derivation of the Boltzmann equation. Foundations of Physics, 47(4):471–489, April 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-017-0072-9.

Bopp:2017:TSQ

[3995] Fritz W. Bopp. Time symmetric quantum mechanics and causal classical physics? Foundations of Physics, 47(4):490–504, April 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/s10701-017-0074-7.

Koks:2017:SRD

[3996] Don Koks. Simultaneity on the rotating disk. Foundations of Physics, 47(4):505–531, April 2017. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-017-0075-6.

Muralidhar:2017:TSS

[3997] Kundeti Muralidhar. Theory of stochastic Schrödinger equation in complex vector space. Foundations of Physics, 47(4):532–552, April 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-017-0076-5.

Eckart:2017:MWS

[3998] Andreas Eckart, Andreas Hüttemann, Claus Kiefer, Silke Britzen, Michal Zajacek, Claus Lämmerzahl, Manfred Stöckler, Monica Valencia-S, Vladimir Karas, and Macarena García-Marín. The Milky Way’s supermassive black hole: How good a case is it? Foundations of Physics, 47(5):553–624, May 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Koslowski:2017:QIC

[3999] Tim Koslowski. Quantum inflation of classical shapes. Foundations of Physics, 47(5):625–639, May 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Boughn:2017:MSB

[4000] Stephen Boughn. Making sense of Bell’s Theorem and quantum nonlocality. Foundations of Physics, 47(5):640–657, May 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Sudbery:2017:SWT

[4001] Anthony Sudbery. Single-world theory of the extended Wigner’s friend experiment. Foundations of Physics, 47(5):658–669, May 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s10701-017-0082-7.pdf.

Bedingham:2017:TRS

[4002] D. J. Bedingham and O. J. E. Maroney. Time reversal symmetry and collapse models. Foundations of Physics, 47(5):670–696, May 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s10701-017-0084-5.pdf.

Kastner:2017:DWM

[4003] R. E. Kastner. Demystifying weak measurements. Foundations of Physics, 47(5):697–707, May 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). See reply [4032].

Capozziello:2017:ASS

[4004] Salvatore Capozziello, Tomislav Prokopec, and Alessandro D. A. M. Spallicci. Aims and scopes of the special issue: Foundations of astrophysics and cosmology. Foundations of Physics, 47(6):709–710, June 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10701-017-0081-8.pdf.

Lopez-Corredoira:2017:TPS

[4005] Martín López-Corredoira. Tests and problems of the standard model in cosmology. Foundations of Physics, 47(6):711–768, June 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Akarsu:2017:ILT

[4006] Özgür Akarsu, Metin Arik, and Nihan Katirci. Inflation and late time acceleration designed by Stueckelberg massive photon. Foundations of Physics, 47(6):769–796, June 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Brandenberger:2017:BCP

[4007] Robert Brandenberger and Patrick Peter. Bouncing cosmologies: Progress and problems. Foundations of Physics, 47(6):797–850, June 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Bonnet-Bidaud:2017:DLU

[4008] Jean-Marc Bonnet-Bidaud. The diffuse light of the universe. Foundations of Physics, 47(6):851–869, June 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Uffink:2017:MOT

[4009] Jos Uffink. Masanes and Oppenheim on the Third Law of Thermodynamics. Foundations of Physics, 47(7):871–872, July 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Mathew:2017:IQT

[4010] Kiran Mathew and Moncy V. John. Interfering quantum trajectories without which-way information. Foundations of Physics, 47(7):873–886, July 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Asenjo:2017:CES

[4011] Felipe A. Asenjo and Sergio A. Hojman. Class of exact solutions for a cosmological model of unified gravitational and quintessence fields. Foundations of Physics, 47(7):887–896, July 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Bera:2017:SFS

[4012] Sayantani Bera, Priyanka Giri, and Tejinder P. Singh. Spacetime fluctuations and a stochastic Schrödinger–Newton equation. Foundations of Physics, 47(7):897–910, July 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Kujala:2017:MDA

[4013] Janne V. Kujala. Minimal distance to approximating noncontextual system as a measure of contextuality. Foundations of Physics, 47(7):911– 932, July 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Borghesi:2017:EQE

[4014] Christian Borghesi. Equivalent quantum equations in a system inspired by bouncing droplets experiments. Foundations of Physics, 47(7):933– 958, July 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Knoll:2017:QMS

[4015] Yehonatan Knoll. Quantum mechanics as a statistical description of classical electrodynamics. Foundations of Physics, 47(7):959–990, July 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Benavoli:2017:GTT

[4016] Alessio Benavoli, Alessandro Facchini, and Marco Zaffalon. A Gleasontype theorem for any dimension based on a gambling formulation of quantum mechanics. Foundations of Physics, 47(7):991–1002, July 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Khrennikov:2017:ASS

[4017] Andrei Khrennikov and Blake C. Stacey. Aims and scope of the special issue, ”Quantum Foundations: Informational Perspective”. Foundations of Physics, 47(8):1003–1008, August 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10701-017-0106-3.pdf.

DeBrota:2017:NBW

[4018] John B. DeBrota and Christopher A. Fuchs. Negativity bounds for Weyl– Heisenberg quasiprobability representations. Foundations of Physics, 47 (8):1009–1030, August 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Bengtsson:2017:NBS

[4019] Ingemar Bengtsson. The number behind the simplest SIC–POVM. Foundations of Physics, 47(8):1031–1041, August 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10701-017-0078-3.pdf.

Appleby:2017:SAN

[4020] Marcus Appleby, Steven Flammia, Gary McConnell, and Jon Yard. SICs and algebraic number theory. Foundations of Physics, 47(8):1042–1059, August 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Stacey:2017:SSN

[4021] Blake C. Stacey. Sporadic SICs and the normed division algebras. Foundations of Physics, 47(8):1060–1064, August 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Bisio:2017:QWW

[4022] Alessandro Bisio, Giacomo Mauro D’Ariano, and Paolo Perinotti. Quantum walks, Weyl equation and the Lorentz group. Foundations of Physics, 47(8):1065–1076, August 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Khrennikov:2017:PSQ

[4023] Andrei Khrennikov. The present situation in quantum theory and its merging with general relativity. Foundations of Physics, 47(8):1077– 1099, August 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10701-017-0089-0.pdf.

Loubenets:2017:BNG

[4024] Elena R. Loubenets. Bell’s nonlocality in a general nonsignaling case: Quantitatively and conceptually. Foundations of Physics, 47(8):1100– 1114, August 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Plotnitsky:2017:CQL

[4025] Arkady Plotnitsky. On the character of quantum law: Complementarity, entanglement, and information. Foundations of Physics, 47(8): 1115–1154, August 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Gomes:2017:SCL

[4026] Henrique Gomes. Semi-classical locality for the non-relativistic path integral in configuration space. Foundations of Physics, 47(9):1155–1184, September 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Josset:2017:SLT

[4027] Thibaut Josset. The Second Law of Thermodynamics at the microscopic scale. Foundations of Physics, 47(9):1185–1190, September 2017. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Selby:2017:DDH

[4028] John H. Selby and Bob Coecke. A diagrammatic derivation of the Hermitian adjoint. Foundations of Physics, 47(9):1191–1207, September 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s10701-017-0102-7.pdf.

Paturel:2017:HLM

[4029] Georges Paturel, Pekka Teerikorpi, and Yurij Baryshev. Hubble Law: Measure and interpretation. Foundations of Physics, 47(9):1208–1228, September 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Mannheim:2017:PSC

[4030] Philip D. Mannheim. PT symmetry, conformal symmetry, and the metrication of electromagnetism. Foundations of Physics, 47(9):1229–1257, September 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Svensson:2017:RCN

[4031] B. E. Y. Svensson. Response to comment on ‘Non-representative Quantum Mechanical Weak Values’ by Ben–Israel and Vaidman. Foundations of Physics, 47(9):1258–1260, September 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). See [3880, 3993].

Cohen:2017:WWM

[4032] Eliahu Cohen. What weak measurements and weak values really mean: Reply to Kastner. Foundations of Physics, 47(10):1261–1266, October 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10701-017-0107-2.pdf. See [4003].

Todd:2017:SCS

[4033] Scott L. Todd and Nicolas C. Menicucci. Sound clocks and sonic relativity. Foundations of Physics, 47(10):1267–1293, October 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s10701-017-0109-0.pdf.

deBarros:2017:CWF

[4034] J. Acacio de Barros and Gary Oas. Can we falsify the consciousnesscauses-collapse hypothesis in quantum mechanics? Foundations of Physics, 47(10):1294–1308, October 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Perlman:2017:QMI

[4035] H. S. Perlman. Quantum mechanics is incomplete but it is consistent with locality. Foundations of Physics, 47(10):1309–1316, October 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Reddiger:2017:MPF

[4036] Maik Reddiger. The Madelung picture as a foundation of geometric quantum theory. Foundations of Physics, 47(10):1317–1367, October 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Chibbaro:2017:CLN

[4037] Sergio Chibbaro and Angelo Vulpiani. Compressibility, laws of nature, initial conditions and complexity. Foundations of Physics, 47(10):1368– 1386, October 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Castagnino:2017:IQT

[4038] Mario Castagnino, Sebastian Fortin, Roberto Laura, and Daniel Sudarsky. Interpretations of quantum theory in the light of modern cosmology. Foundations of Physics, 47(11):1387–1422, November 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Fortin:2017:IDC

[4039] Sebastian Fortin and Olimpia Lombardi. Interpretation and decoherence: a contribution to the debate Vassallo & Esfeld versus Crull. Foundations of Physics, 47(11):1423–1427, November 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). See [3872, 4040].

Crull:2017:YMD

[4040] Elise M. Crull. Yes, more decoherence: a reply to critics. Foundations of Physics, 47(11):1428–1463, November 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). See [3872, 4039].

DeHaro:2017:ID

[4041] Sebastian De Haro. The invisibility of diffeomorphisms. Foundations of Physics, 47(11):1464–1497, November 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10701-017-0117-0.pdf.

Reed:2017:BRL

[4042] Donald Reed. Book review: Lawrence P. Horwitz: Relativistic Quantum Mechanics. Foundations of Physics, 47(11):1498–1502, November 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Hooft:2017:FTB

[4043] Gerard ’t Hooft. The firewall transformation for black holes and some of its implications. Foundations of Physics, 47(12):1503–1542, December 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10701-017-0122-3.pdf.

Mazzola:2017:DRC

[4044] Claudio Mazzola and Peter W. Evans. Do Reichenbachian common cause systems of arbitrary finite size exist? Foundations of Physics, 47(12): 1543–1558, December 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Van:2017:TSS

[4045] Thuan Vo Van. A time–space symmetry based cylindrical model for quantum mechanical interpretations. Foundations of Physics, 47(12): 1559–1581, December 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Lienert:2017:MTW

[4046] Matthias Lienert, Sören Petrat, and Roderich Tumulka. Multi-time wave functions versus multiple timelike dimensions. Foundations of Physics, 47(12):1582–1590, December 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Lambare:2017:NTC

[4047] Justo Pastor Lambare. On Nieuwenhuizen’s treatment of contextuality in Bell’s theorem. Foundations of Physics, 47(12):1591–1596, December 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Leon:2017:PO

[4048] Juan Leon and Lorenzo Maccone. The Pauli objection. Foundations of Physics, 47(12):1597–1608, December 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Santamato:2017:PSS

[4049] Enrico Santamato and Francesco De Martini. Proof of the spin statistics connection 2: Relativistic theory. Foundations of Physics, 47(12):1609– 1625, December 2017. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Aharonov:2018:IFE

[4050] Yakir Aharonov, Eliahu Cohen, Avshalom C. Elitzur, and Lee Smolin. Interaction-free effects between distant atoms. Foundations of Physics, 48(1):1–16, January 2018. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s10701-017-0127-y.pdf.

Luongo:2018:SAU

[4051] Orlando Luongo and Hernando Quevedo. Self-accelerated universe induced by repulsive effects as an alternative to dark energy and modified gravities. Foundations of Physics, 48(1):17–26, January 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Sanchez-Kuntz:2018:QLR

[4052] Natalia Sánchez-Kuntz and Eduardo Nahmad-Achar. Quantum locality, rings a bell?: Bell’s inequality meets local reality and true determinism. Foundations of Physics, 48(1):27–47, January 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Bryan:2018:RCU

[4053] K. L. H. Bryan and A. J. M. Medved. Realistic clocks for a universe without time. Foundations of Physics, 48(1):48–59, January 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Sciarretta:2018:LRM

[4054] Antonio Sciarretta. A local-realistic model of quantum mechanics based on a discrete spacetime. Foundations of Physics, 48(1):60–91, January 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Srikanth:2018:QBC

[4055] R. Srikanth. Quantum bit commitment and the reality of the quantum state. Foundations of Physics, 48(1):92–109, January 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Calmet:2018:NMC

[4056] Xavier Calmet, Iberê Kuntz, and Ian G. Moss. Non-minimal coupling of the Higgs Boson to curvature in an inflationary universe. Foundations of Physics, 48(1):110–120, January 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10701-017-0131-2.pdf.

Smolin:2018:DD

[4057] Lee Smolin. The dynamics of difference. Foundations of Physics, 48(2): 121–134, February 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Loveridge:2018:SRF

[4058] Leon Loveridge, Takayuki Miyadera, and Paul Busch. Symmetry, reference frames, and relational quantities in quantum mechanics. Foundations of Physics, 48(2):135–198, February 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10701-018-0138-3.pdf.

Atiyah:2018:RGD

[4059] M. F. Atiyah and J. Malkoun. The relativistic geometry and dynamics of electrons. Foundations of Physics, 48(2):199–208, February 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Vink:2018:PTQ

[4060] Jeroen C. Vink. Particle trajectories for quantum field theory. Foundations of Physics, 48(2):209–236, February 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Doboszewski:2018:HEN

[4061] Juliusz Doboszewski and Niels Linnemann. How not to establish the non-renormalizability of gravity. Foundations of Physics, 48(2):237–252, February 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Avalos:2018:NPP

[4062] R. Avalos, F. Dahia, and C. Romero. A note on the problem of proper time in Weyl space–time. Foundations of Physics, 48(2):253–270, February 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Tessarotto:2018:BGL

[4063] Massimo Tessarotto, Claudio Cremaschini, Michael Mond, Claudio Asci, Alessandro Soranzo, and Gino Tironi. On the Boltzmann–grad limit for smooth hard-sphere systems. Foundations of Physics, 48(3):271–294, March 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Friedberg:2018:WQM

[4064] R. Friedberg and P. C. Hohenberg. What is quantum mechanics? A minimal formulation. Foundations of Physics, 48(3):295–332, March 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Castagnoli:2018:CPR

[4065] Giuseppe Castagnoli. Completing the physical representation of quantum algorithms provides a quantitative explanation of their computational speedup. Foundations of Physics, 48(3):333–354, March 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Rosaler:2018:GER

[4066] Joshua Rosaler. Generalized Ehrenfest relations, deformation quantization, and the geometry of inter-model reduction. Foundations of Physics, 48(3):355–385, March 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Cooperstock:2018:EUG

[4067] F. I. Cooperstock and M. J. Dupre. Energy and uncertainty in general relativity. Foundations of Physics, 48(4):387–394, April 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Montambaux:2018:GSB

[4068] Gilles Montambaux. Generalized Stefan–Boltzmann Law. Foundations of Physics, 48(4):395–410, April 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Okon:2018:LSB

[4069] Elias Okon and Daniel Sudarsky. Losing stuff down a black hole. Foundations of Physics, 48(4):411–428, April 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Sparenberg:2018:DDO

[4070] Jean-Marc Sparenberg and David Gaspard. Decoherence and determinism in a one-dimensional cloud-chamber model. Foundations of Physics, 48(4):429–439, April 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Ruyant:2018:CWM

[4071] Quentin Ruyant. Can we make sense of relational quantum mechanics? Foundations of Physics, 48(4):440–455, April 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Bleybel:2018:GTT

[4072] Ali Bleybel and Abdallah Zaiour. A general theorem on temporal foliations of causal sets. Foundations of Physics, 48(4):456–478, April 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Redei:2018:E

[4073] Miklos Redei, Klaas Landsman, and Harvey Brown. Editorial. Foundations of Physics, 48(5):479–480, May 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10701-018-0171-2.pdf.

Rovelli:2018:PNP

[4074] Carlo Rovelli. Physics needs philosophy. Philosophy needs physics. Foundations of Physics, 48(5):481–491, May 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Dawid:2018:DU

[4075] Richard Dawid. Delimiting the unconceived. Foundations of Physics, 48 (5):492–506, May 2018. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s10701-017-0132-1.pdf.

Bangu:2018:MAC

[4076] Sorin Bangu and Robert H. C. Moir. The ‘miracle’ of applicability? The curious case of the simple harmonic oscillator. Foundations of Physics, 48(5):507–525, May 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Palacios:2018:WWE

[4077] Patricia Palacios. Had we but world enough, and time … but we don’t!: Justifying the thermodynamic and infinite-time limits in statistical mechanics. Foundations of Physics, 48(5):526–541, May 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Kay:2018:MGE

[4078] Bernard S. Kay. The matter–gravity entanglement hypothesis. Foundations of Physics, 48(5):542–557, May 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10701-018-0150-7.pdf.

Dewar:2018:GEN

[4079] Neil Dewar and James Owen Weatherall. On gravitational energy in Newtonian theories. Foundations of Physics, 48(5):558–578, May 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). See comments [4201].

Pitts:2018:ETC

[4080] J. Brian Pitts. Equivalent theories and changing Hamiltonian observables in general relativity. Foundations of Physics, 48(5):579– 590, May 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10701-018-0148-1.pdf.

Martens:2018:ALR

[4081] Niels C. M. Martens. Against Laplacian reduction of Newtonian mass to spatiotemporal quantities. Foundations of Physics, 48(5):591–609, May 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Manchak:2018:IPR

[4082] J. B. Manchak and James Owen Weatherall. (Information) paradox regained? A brief comment on Maudlin on black hole information loss. Foundations of Physics, 48(6):611–627, June 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Dente:2018:CCL

[4083] Gregory C. Dente. Causality in the classical limit for quantum electrodynamics. Foundations of Physics, 48(6):628–635, June 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Vanchurin:2018:IGF

[4084] Vitaly Vanchurin. Information graph flow: a geometric approximation of quantum and statistical systems. Foundations of Physics, 48(6):636– 653, June 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Li:2018:OQM

[4085] Sheng-Wen Li, C. Y. Cai, X. F. Liu, and C. P. Sun. Objectivity in quantum measurement. Foundations of Physics, 48(6):654–667, June 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Gomes:2018:TCS

[4086] Henrique Gomes. Timeless configuration space and the emergence of classical behavior. Foundations of Physics, 48(6):668–715, June 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Schurmann:2018:UPD

[4087] Thomas Schürmann. Uncertainty principle on 3-dimensional manifolds of constant curvature. Foundations of Physics, 48(6):716–725, June 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Lezcano:2018:SVN

[4088] Alfredo González Lezcano and Alejandro Cabo Montes de Oca. A stochastic version of the Noether theorem. Foundations of Physics, 48 (6):726–746, June 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Rovelli:2018:BRS

[4089] Carlo Rovelli. Book review: Shan Gao: The Meaning of the Wave Function. In Search of the Ontology of Quantum Mechanics. Foundations of Physics, 48(6):747–749, June 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Mandolesi:2018:AWP

[4090] André L. G. Mandolesi. Analysis of Wallace’s proof of the Born rule in Everettian quantum mechanics: Formal aspects. Foundations of Physics, 48(7):751–782, July 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Laino:2018:KPR

[4091] Luigi Laino. Is knowledge of physical reality still Kantian? Some remarks about the transcendental character of loop quantum gravity. Foundations of Physics, 48(7):783–802, July 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Vervoort:2018:HVT

[4092] Louis Vervoort. Are hidden-variable theories for pilot-wave systems possible? Foundations of Physics, 48(7):803–826, July 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Page:2018:NOP

[4093] Don N. Page. Normalized observational probabilities from unnormalizable quantum states or phase-space distributions. Foundations of Physics, 48(7):827–836, July 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Li:2018:TQF

[4094] Benliang Li, Daniel W. Hewak, and Qi Jie Wang. The transition from quantum field theory to one-particle quantum mechanics and a proposed interpretation of Aharonov–Bohm effect. Foundations of Physics, 48(7): 837–852, July 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Barnum:2018:IQI

[4095] Howard Barnum, Stephanie Wehner, and Alexander Wilce. Introduction: Quantum information theory and quantum foundations. Foundations of Physics, 48(8):853–856, August 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10701-018-0188-6.pdf.

Palazuelos:2018:RCB

[4096] Carlos Palazuelos. Random constructions in Bell inequalities: a survey. Foundations of Physics, 48(8):857–885, August 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Garner:2018:ICB

[4097] Andrew J. P. Garner. Interferometric computation beyond quantum theory. Foundations of Physics, 48(8):886–909, August 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Li:2018:SEE

[4098] Ke Li and Andreas Winter. Squashed entanglement, k-extendibility, quantum Markov chains, and recovery maps. Foundations of Physics, 48 (8):910–924, August 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Sainz:2018:MCC

[4099] Ana Belén Sainz and Elie Wolfe. Multipartite composition of contextuality scenarios. Foundations of Physics, 48(8):925–953, August 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Barnum:2018:OQL

[4100] Howard Barnum, Ciarán M. Lee, and John H. Selby. Oracles and query lower bounds in generalised probabilistic theories. Foundations of Physics, 48(8):954–981, August 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10701-018-0198-4.pdf.

Branford:2018:DHB

[4101] Dominic Branford, Oscar C. O. Dahlsten, and Andrew J. P. Garner. On defining the Hamiltonian beyond quantum theory. Foundations of Physics, 48(8):982–1006, August 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Mermin:2018:HNN

[4102] N. David Mermin and Rüdiger Schack. Homer nodded: Von Neumann’s surprising oversight. Foundations of Physics, 48(9):1007–1020, September 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10701-018-0197-5.pdf.

Kent:2018:QQ

[4103] Adrian Kent. Quanta and qualia. Foundations of Physics, 48(9):1021– 1037, September 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10701-018-0193-9.pdf.

Bonder:2018:EGQ

[4104] Yuri Bonder, Chryssomalis Chryssomalakos, and Daniel Sudarsky. Extracting geometry from quantum spacetime. Foundations of Physics, 48 (9):1038–1060, September 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Zalamea:2018:TRO

[4105] Federico Zalamea. The twofold role of observables in classical and quantum kinematics. Foundations of Physics, 48(9):1061–1091, September 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Trassinelli:2018:RQM

[4106] M. Trassinelli. Relational quantum mechanics and probability. Foundations of Physics, 48(9):1092–1111, September 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Caprara:2018:LLJ

[4107] Sergio Caprara and Angelo Vulpiani. Law without law or “just” limit theorems? Foundations of Physics, 48(9):1112–1127, September 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Gudder:2018:PB

[4108] Stan Gudder, Pekka Lahti, and Leon Loveridge. Paul Busch 1955–2018. Foundations of Physics, 48(9):1128–1130, September 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s10701-018-0206-8.pdf.

Bellucci:2018:BHG

[4109] Stefano Bellucci, Alfio Bonanno, S. J. Gabriele Gionti, and Fabio Scardigli. Black holes, gravitational waves and space time singularities. Foundations of Physics, 48(10):1131–1133, October 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s10701-018-0200-1.pdf.

tHooft:2018:VBH

[4110] Gerard ’t Hooft. Virtual black holes and space–time structure. Foundations of Physics, 48(10):1134–1149, October 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10701-017-0133-0.pdf.

Meissner:2018:NLE

[4111] Krzysztof A. Meissner and Hermann Nicolai. Non-local effects of conformal anomaly. Foundations of Physics, 48(10):1150–1158, October 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Kamenshchik:2018:SCT

[4112] A. Yu. Kamenshchik. Singularity crossing, transformation of matter properties and the problem of parametrization in field theories. Foundations of Physics, 48(10):1159–1176, October 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Penrose:2018:BBD

[4113] Roger Penrose. The Big Bang and its dark-matter content: Whence, whither, and wherefore. Foundations of Physics, 48(10):1177–1190, October 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10701-018-0162-3.pdf.

Kempf:2018:QGI

[4114] Achim Kempf. Quantum gravity, information theory and the CMB. Foundations of Physics, 48(10):1191–1203, October 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Casadio:2018:HQM

[4115] Roberto Casadio, Andrea Giugno, Andrea Giusti, and Octavian Micu. Horizon quantum mechanics: Spherically symmetric and rotating sources. Foundations of Physics, 48(10):1204–1218, October 2018. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Dvali:2018:QLB

[4116] Gia Dvali. On quantum life of black holes. Foundations of Physics, 48(10):1219–1225, October 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Ellis:2018:SCM

[4117] George Ellis. The standard cosmological model: achievements and issues. Foundations of Physics, 48(10):1226–1245, October 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Linde:2018:PIC

[4118] Andrei Linde. On the problem of initial conditions for inflation. Foundations of Physics, 48(10):1246–1260, October 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Turner:2018:CMM

[4119] Michael S. Turner. Λ CDM: Much more than we expected, but now less than what we want. Foundations of Physics, 48(10):1261–1278, October 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Kallosh:2018:SFP

[4120] Renata Kallosh. From the sky to the fundamental physics. Foundations of Physics, 48(10):1279–1290, October 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Platania:2018:FRG

[4121] Alessia Platania and Frank Saueressig. Functional renormalization group flows on Friedman–Lemaître–Robertson–Walker backgrounds. Foundations of Physics, 48(10):1291–1304, October 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10701-018-0181-0.pdf.

Silk:2018:TLC

[4122] Joseph Silk. Towards the limits of cosmology. Foundations of Physics, 48(10):1305–1332, October 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10701-018-0183-y.pdf.

Kragh:2018:GLP

[4123] Helge Kragh. Georges Lemaître, pioneer of modern theoretical cosmology. Foundations of Physics, 48(10):1333–1348, October 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Bini:2018:BHP

[4124] Donato Bini and Andrea Geralico. Black hole perturbations: a review of recent analytical results. Foundations of Physics, 48(10):1349–1363, October 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Percacci:2018:UQG

[4125] R. Percacci. Unimodular quantum gravity and the cosmological constant. Foundations of Physics, 48(10):1364–1379, October 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Vidotto:2018:MLB

[4126] Francesca Vidotto. Measuring the last burst of non-singular black holes. Foundations of Physics, 48(10):1380–1392, October 2018. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Bonanno:2018:GCQ

[4127] Alfio Bonanno, Benjamin Koch, and Alessia Platania. Gravitational collapse in quantum Einstein gravity. Foundations of Physics, 48(10): 1393–1406, October 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Eichhorn:2018:SAS

[4128] Astrid Eichhorn. Status of the asymptotic safety paradigm for quantum gravity and matter. Foundations of Physics, 48(10):1407–1429, October 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Manchak:2018:SNH

[4129] J. B. Manchak. Some “no hole” spacetime properties are unstable. Foundations of Physics, 48(11):1539–1545, November 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Fedullo:2018:HUR

[4130] Aniello Fedullo. Heisenberg uncertainty relations as statistical invariants. Foundations of Physics, 48(11):1546–1556, November 2018. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Adler:2018:CDM

[4131] Stephen L. Adler. Connecting the dots: Mott for emulsions, collapse models, colored noise, frame dependence of measurements, evasion of the “free will theorem”. Foundations of Physics, 48(11):1557–1567, November 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Healey:2018:QTL

[4132] Richard Healey. Quantum theory and the limits of objectivity. Foundations of Physics, 48(11):1568–1589, November 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). See comments [4181] and reply [4186].

Davidson:2018:BTK

[4133] Mark Davidson. Bohmian trajectories for Kerr–Newman particles in complex space–time. Foundations of Physics, 48(11):1590–1616, November 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Holman:2018:HPN

[4134] M. Holman. How problematic is the near–Euclidean spatial geometry of the large-scale universe? Foundations of Physics, 48(11):1617–1647, November 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Gazeau:2018:CQM

[4135] Jean-Pierre Gazeau. From classical to quantum models: The regularising rôle of integrals, symmetry and probabilities. Foundations of Physics, 48 (11):1648–1667, November 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Szangolies:2018:EHF

[4136] Jochen Szangolies. Epistemic horizons and the foundations of quantum mechanics. Foundations of Physics, 48(12):1669–1697, December 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Waegell:2018:ONL

[4137] Mordecai Waegell. An ontology of nature with local causality, parallel lives, and many relative worlds. Foundations of Physics, 48(12):1698– 1730, December 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Bartley:2018:IPL

[4138] David L. Bartley. Individual particle localization per relativistic de Broglie–Bohm. Foundations of Physics, 48(12):1731–1752, December 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Tilloy:2018:BQM

[4139] Antoine Tilloy. Binding quantum matter and space–time, without romanticism. Foundations of Physics, 48(12):1753–1769, December 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10701-018-0224-6.pdf.

Rizzi:2018:DPT

[4140] Anthony Rizzi. Does the PBR theorem rule out a statistical understanding of QM? Foundations of Physics, 48(12):1770–1793, December 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Kandhadai:2019:PQR

[4141] Adithya Kandhadai and Antony Valentini. Perturbations and quantum relaxation. Foundations of Physics, 49(1):1–23, January 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Mandolesi:2019:AWP

[4142] André L. G. Mandolesi. Analysis of Wallace’s proof of the Born rule in Everettian quantum mechanics II: Concepts and axioms. Foundations of Physics, 49(1):24–52, January 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Bopp:2019:BDB

[4143] Fritz W. Bopp. A bi-directional Big Bang/Crunch Universe within a twostate-vector quantum mechanics? Foundations of Physics, 49(1):53–62, January 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Choudhuri:2019:CDP

[4144] Amitava Choudhuri and Aritra Ganguly. Cosmological density perturbations in Newtonian- and MONDian gravity scenario: a symmetry-based approach. Foundations of Physics, 49(1):63–82, January 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Sen:2019:LER

[4145] Indrajit Sen. A local ψ-epistemic retrocausal hidden-variable model of Bell correlations with wavefunctions in physical space. Foundations of Physics, 49(2):83–95, February 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Martin-Dussaud:2019:NLR

[4146] P. Martin-Dussaud, C. Rovelli, and F. Zalamea. The notion of locality in relational quantum mechanics. Foundations of Physics, 49(2):96–106, February 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). See comment [4217].

Sbitnev:2019:QAS

[4147] Valeriy I. Sbitnev. Quaternion algebra on 4D superfluid quantum space– time: Gravitomagnetism. Foundations of Physics, 49(2):107–143, February 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Errehymy:2019:MAC

[4148] Abdelghani Errehymy and Mohammed Daoud. Model astrophysical configurations with the equation of state of Chaplygin gas. Foundations of Physics, 49(2):144–175, February 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Giddings:2019:QFG

[4149] Steven B. Giddings. Quantum-first gravity. Foundations of Physics, 49 (3):177–190, March 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Kuntz:2019:NSG

[4150] Iberê Kuntz. Non-smooth gravity and parity violation. Foundations of Physics, 49(3):191–201, March 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10701-019-00240-8.pdf.

Araujo:2019:PTD

[4151] Mateus Araújo. Probability in two deterministic universes. Foundations of Physics, 49(3):202–231, March 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Matarese:2019:LQG

[4152] Vera Matarese. Loop quantum gravity: a new threat to Humeanism? Part I: The problem of spacetime. Foundations of Physics, 49(3):232– 259, March 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Hermens:2019:OPS

[4153] Ronnie Hermens. An operationalist perspective on setting dependence. Foundations of Physics, 49(3):260–282, March 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10701-019-00243-5.pdf.

Filatov:2019:IPO

[4154] Stanislav Filatov and Marcis Auzinsh. On interchangeability of probe– object roles in quantum–quantum interaction–free measurement. Foundations of Physics, 49(3):283–297, March 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Matzkin:2019:WVQ

[4155] A. Matzkin. Weak values and quantum properties. Foundations of Physics, 49(3):298–316, March 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

John:2019:PMI

[4156] Moncy V. John and Kiran Mathew. Position measurement-induced collapse: a unified quantum description of Fraunhofer and Fresnel diffractions. Foundations of Physics, 49(4):317–329, April 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Gratus:2019:EBH

[4157] Jonathan Gratus, Paul Kinsler, and Martin W. McCall. Evaporating black holes, wormholes, and vacuum polarisation: Must they always conserve charge? Foundations of Physics, 49(4):330–350, April 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10701-019-00251-5.pdf.

Chaitanya:2019:QHK

[4158] K. V. S. Shiv Chaitanya. Quantum hydrodynamics: Kirchhoff equations. Foundations of Physics, 49(4):351–364, April 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Sebens:2019:EQP

[4159] Charles T. Sebens. Electromagnetism as quantum physics. Foundations of Physics, 49(4):365–389, April 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Hassan:2019:NPH

[4160] Syed Moeez Hassan. A non-perturbative Hamiltonian approach to the cosmological constant problem. Foundations of Physics, 49(5):391–427, May 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Wells:2019:FCI

[4161] James D. Wells. Finetuned cancellations and improbable theories. Foundations of Physics, 49(5):428–443, May 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Brandenburger:2019:ABD

[4162] Adam Brandenburger and Kai Steverson. Axioms for the Boltzmann distribution. Foundations of Physics, 49(5):444–456, May 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s10701-019-00257-z.pdf.

Gudder:2019:PBH

[4163] Stan Gudder and Pekka Lahti. Paul Busch: at the heart of quantum mechanics. Foundations of Physics, 49(6):457–459, June 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s10701-019-00270-2.pdf.

Werner:2019:UHT

[4164] Reinhard F. Werner and Terry Farrelly. Uncertainty from Heisenberg to today. Foundations of Physics, 49(6):460–491, June 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Carmeli:2019:NDR

[4165] Claudio Carmeli, Teiko Heinosaari, Takayuki Miyadera, and Alessandro Toigo. Noise-disturbance relation and the Galois connection of quantum measurements. Foundations of Physics, 49(6):492–505, June 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s10701-019-00255-1.pdf.

Kiukas:2019:COQ

[4166] Jukka Kiukas, Pekka Lahti, Juha-Pekka Pellonpää, and Kari Ylinen. Complementary observables in quantum mechanics. Foundations of Physics, 49(6):506–531, June 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10701-019-00261-3.pdf.

Carmeli:2019:CEC

[4167] Claudio Carmeli, Gianni Cassinelli, and Alessandro Toigo. Constructing extremal compatible quantum observables by means of two mutually unbiased bases. Foundations of Physics, 49(6):532–548, June 2019. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Loveridge:2019:RQT

[4168] Leon Loveridge and Takayuki Miyadera. Relative quantum time. Foundations of Physics, 49(6):549–560, June 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Beneduci:2019:SLP

[4169] Roberto Beneduci and Franklin Schroeck, Jr. Space localization of the photon. Foundations of Physics, 49(6):561–576, June 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Schmidt:2019:REF

[4170] Heinz-Jürgen Schmidt. Reduction and extremality of finite observables. Foundations of Physics, 49(6):577–593, June 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Wright:2019:GTT

[4171] Victoria J. Wright and Stefan Weigert. Gleason-type theorems from Cauchy’s functional equation. Foundations of Physics, 49(6):594– 606, June 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10701-019-00275-x.pdf. See correction [4245].

Dvurecenskij:2019:PEA

[4172] Anatolij Dvurecenskij. Perfect effect algebras and spectral resolutions of observables. Foundations of Physics, 49(6):607–628, June 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Dieks:2019:QRP

[4173] Dennis Dieks. Quantum reality, perspectivalism and covariance. Foundations of Physics, 49(6):629–646, June 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10701-019-00264-0.pdf.

Gudder:2019:CQM

[4174] Stanley Gudder. Contexts in quantum measurement theory. Foundations of Physics, 49(6):647–662, June 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Jaeger:2019:QUP

[4175] Gregg Jaeger. Quantum unsharpness, potentiality, and reality. Foundations of Physics, 49(6):663–676, June 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Stulpe:2019:AQC

[4176] Werner Stulpe. Aspects of the quantum–classical connection based on statistical maps. Foundations of Physics, 49(6):677–692, June 2019. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Rovelli:2019:NDD

[4177] Carlo Rovelli and Václav Zatloukal. Natural discrete differential calculus in physics. Foundations of Physics, 49(7):693–699, July 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Benitez:2019:SRF

[4178] Federico Benitez. Selective realism and the framework/interaction distinction: a taxonomy of fundamental physical theories. Foundations of Physics, 49(7):700–716, July 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Kastner:2019:DCQ

[4179] R. E. Kastner. The ‘delayed choice quantum eraser’ neither erases nor delays. Foundations of Physics, 49(7):717–727, July 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Smilga:2019:SPI

[4180] Walter Smilga. Structural properties of irreducible two-particle representations of the Poincaré group. Foundations of Physics, 49(7):728– 740, July 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Baumann:2019:CHQ

[4181] Veronika Baumann, Flavio Del Santo, and Caslav Brukner. Comment on Healey’s “Quantum Theory and the Limits of Objectivity”. Foundations of Physics, 49(7):741–749, July 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10701-019-00276-w.pdf. See [4132] and reply

[4186].

Khokonov:2019:EBS

[4182] M. Kh. Khokonov and J. U. Andersen. Equivalence between self-energy and self-mass in classical electron model. Foundations of Physics, 49(7): 750–782, July 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Ballentine:2019:MW

[4183] Leslie E. Ballentine. A meeting with Wigner. Foundations of Physics, 49 (8):783–785, August 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Pitts:2019:WOH

[4184] J. Brian Pitts. What are observables in Hamiltonian Einstein– Maxwell theory? Foundations of Physics, 49(8):786–796, August 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s10701-019-00284-w.pdf.

Maclay:2019:TQI

[4185] G. Jordan Maclay and Eric W. Davis. Testing a quantum inequality with a meta-analysis of data for squeezed light. Foundations of Physics, 49 (8):797–815, August 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Healey:2019:RCQ

[4186] Richard Healey. Reply to a comment on “Quantum Theory and the Limits of Objectivity”. Foundations of Physics, 49(8):816–819, August 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). See [4132, 4181].

Bennett:2019:CAH

[4187] Charles H. Bennett, Robin Hanson, and C. Jess Riedel. Comment on ‘The Aestivation Hypothesis for Resolving Fermi’s Paradox’. Foundations of Physics, 49(8):820–829, August 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). See [4691].

Cerdonio:2019:AMB

[4188] Massimo Cerdonio. About the measure of the bare cosmological constant. Foundations of Physics, 49(8):830–836, August 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Brox:2019:MPN

[4189] Daniel Brox. Modern physics and number theory. Foundations of Physics, 49(8):837–853, August 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Rosaler:2019:P

[4190] Joshua Rosaler, Robert Harlander, Gregor Schiemann, and Miguel Ángel Carretero Sahuquillo. Preface. Foundations of Physics, 49(9):855– 859, September 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10701-019-00288-6.pdf.

Borrelli:2019:PNH

[4191] Arianna Borrelli and Elena Castellani. The practice of naturalness: a historical–philosophical perspective. Foundations of Physics, 49(9):860– 878, September 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Harlander:2019:HNR

[4192] Robert Harlander and Joshua Rosaler. Higgs naturalness and renormalized parameters. Foundations of Physics, 49(9):879–897, September 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Bain:2019:WN

[4193] Jonathan Bain. Why be natural? Foundations of Physics, 49(9):898–914, September 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Jegerlehner:2019:HPC

[4194] Fred Jegerlehner. The hierarchy problem and the cosmological constant problem revisited. Foundations of Physics, 49(9):915–971, September 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Gubitosi:2019:SHA

[4195] Giulia Gubitosi, Chris Ripken, and Frank Saueressig. Scales and hierachies in asymptotically safe quantum gravity: a review. Foundations of Physics, 49(9):972–990, September 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10701-019-00263-1.pdf.

Wells:2019:NEE

[4196] James D. Wells. Naturalness, extra-empirical theory assessments, and the implications of skepticism. Foundations of Physics, 49(9):991–1010, September 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Friederich:2019:NFT

[4197] Simon Friederich. A new fine-tuning argument for the multiverse. Foundations of Physics, 49(9):1011–1021, September 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10701-019-00246-2.pdf.

Williams:2019:TNN

[4198] Porter Williams. Two notions of naturalness. Foundations of Physics, 49(9):1022–1050, September 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Pas:2019:PBM

[4199] Heinrich Päs. Physics beyond the multiverse: Naturalness and the quest for a fundamental theory. Foundations of Physics, 49(9):1051–1065, September 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Hatkar:2019:DES

[4200] S. P. Hatkar, P. S. Dudhe, and S. D. Katore. Dark energy scenario in metric f(R) formalism. Foundations of Physics, 49(10):1067– 1085, October 2019. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-019-00302-x.

Duerr:2019:GEN

[4201] Patrick M. Duerr and James Read. Gravitational energy in Newtonian gravity: a response to Dewar and Weatherall. Foundations of Physics, 49(10):1086–1110, October 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-019-00301-y;   https://
link.springer.com/content/pdf/10.1007/s10701-019-00301-y.pdf.

See [4079].

Rosaler:2019:GRC

[4202] Joshua Rosaler. The geometry of reduction: Compound reduction and overlapping state space domains. Foundations of Physics, 49(10): 1111–1142, October 2019. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-019-00299-3.

Boge:2019:QIV

[4203] Florian J. Boge. Quantum information versus epistemic logic: an analysis of the Frauchiger–Renner Theorem. Foundations of Physics, 49(10): 1143–1165, October 2019. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-019-00298-4.

Drezet:2019:LIR

[4204] Aurélien Drezet. Lorentz-invariant, retrocausal, and deterministic hidden variables. Foundations of Physics, 49(10):1166–1199, October 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-019-00297-5.

Castagnoli:2019:RTS

[4205] G. Castagnoli, E. Cohen, A. K. Ekert, and A. C. Elitzur. A relational time-symmetric framework for analyzing the quantum computational speedup. Foundations of Physics, 49(10):1200–1230, October 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-019-00300-z.

Stefanon:2019:LAT

[4206] Mario Stefanon. Looking at the Arrow of Time and Loschmidt’s Paradox through the magnifying glass of mathematical-billiard. Foundations of Physics, 49(10):1231–1251, October 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-019-00295-7.

Ellis:2019:HDC

[4207] George Ellis and Barbara Drossel. How downwards causation occurs in digital computers. Foundations of Physics, 49(11):1253–1277, November 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-019-00307-6.

Roncaglia:2019:CIQ

[4208] Marco Roncaglia. On the conservation of information in quantum physics. Foundations of Physics, 49(11):1278–1286, November 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-019-00304-9.

Viaggiu:2019:CCP

[4209] S. Viaggiu. The cosmological constant from Planckian fluctuations and the averaging procedure. Foundations of Physics, 49(11):1287– 1305, November 2019. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-019-00308-5.

Yang:2019:CDQ

[4210] Jianhao M. Yang. Consistent descriptions of quantum measurement. Foundations of Physics, 49(11):1306–1324, November 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

https://link.springer.com/article/10.1007/s10701-019-00305-
8;   https://link.springer.com/content/pdf/10.1007/s10701-019-
00305-8.pdf.

Rovelli:2019:NPN

[4211] Carlo Rovelli. Neither presentism nor eternalism. Foundations of Physics, 49(12):1325–1335, December 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-019-00312-9.

Thyssen:2019:CR

[4212] Pieter Thyssen. Conventionality and reality. Foundations of Physics, 49(12):1336–1354, December 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-019-00294-8;   https://
link.springer.com/content/pdf/10.1007/s10701-019-00294-8.pdf.

Ben-Yami:2019:ADS

[4213] Hanoch Ben-Yami. Absolute distant simultaneity in special relativity. Foundations of Physics, 49(12):1355–1364, December 2019. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

https://link.springer.com/article/10.1007/s10701-019-00306-
7;   https://link.springer.com/content/pdf/10.1007/s10701-019-
00306-7.pdf.

Hnilo:2019:SEC

[4214] Alejandro A. Hnilo. Simple explanation of the classical limit. Foundations of Physics, 49(12):1365–1371, December 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-019-00310-x.

Fortin:2019:CPU

[4215] Sebastian Fortin and Olimpia Lombardi. The correspondence principle and the understanding of decoherence. Foundations of Physics, 49(12): 1372–1393, December 2019. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-019-00309-4.

Maccone:2019:FPQ

[4216] Lorenzo Maccone. A fundamental problem in quantizing general relativity. Foundations of Physics, 49(12):1394–1403, December 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-019-00311-w.

Pienaar:2019:CNL

[4217] Jacques Pienaar. Comment on “The Notion of Locality in Relational Quantum Mechanics”. Foundations of Physics, 49(12):1404–1414, December 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-019-00303-w. See [4146].

Zwirn:2020:NVM

[4218] Hervé Zwirn. Nonlocality versus modified realism. Foundations of Physics, 50(1):1–26, January 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-019-00314-7.

Cetto:2020:SPO

[4219] Ana María Cetto, Andrea Valdés-Hernández, and Luis de la Peña. On the spin projection operator and the probabilistic meaning of the bipartite correlation function. Foundations of Physics, 50(1):27–39, January 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-019-00313-8.

Vanchurin:2020:EMT

[4220] Vitaly Vanchurin. Entropic mechanics: Towards a stochastic description of quantum mechanics. Foundations of Physics, 50(1):40–53, January 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-019-00315-6.

Gaveau:2020:RKQ

[4221] B. Gaveau and L. S. Schulman. Reconciling kinetic and quantum theory. Foundations of Physics, 50(2):55–60, February 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10701-019-00317-4.pdf.

Landsman:2020:RWR

[4222] Klaas Landsman. Randomness? What randomness? Foundations of Physics, 50(2):61–104, February 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10701-020-00318-8.pdf.

Engelbert:2020:HPD

[4223] Nicholas G. Engelbert and Renato M. Angelo. Hardy’s paradox as a demonstration of quantum irrealism. Foundations of Physics, 50(2):105– 119, February 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Hetzroni:2020:REH

[4224] Guy Hetzroni. Relativity and equivalence in Hilbert space: a principle– theory approach to the Aharonov–Bohm effect. Foundations of Physics, 50(2):120–135, February 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

deOliveira:2020:NVA

[4225] César R. de Oliveira and Renan G. Romano. A new version of the Aharonov–Bohm effect. Foundations of Physics, 50(3):137–146, March 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Searight:2020:GKK

[4226] Trevor P. Searight. f(R) gravity in a Kaluza–Klein theory with degenerate metric. Foundations of Physics, 50(3):147–160, March 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Ellis:2020:ET

[4227] George F. R. Ellis and Barbara Drossel. Emergence of time. Foundations of Physics, 50(3):161–190, March 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Al-Rashid:2020:RES

[4228] Saeed Naif Turki Al-Rashid, Mohammed A. Z. Habeeb, and Tugdual S. LeBohec. Riccati equations as a scale-relativistic gateway to quantum mechanics. Foundations of Physics, 50(3):191–203, March 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Rizzi:2020:HNI

[4229] Anthony Rizzi. How the natural interpretation of QM avoids the recent no-go theorem. Foundations of Physics, 50(3):204–215, March 2020. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Roberts:2020:NPH

[4230] Bryan W. Roberts and James Owen Weatherall. New perspectives on the hole argument. Foundations of Physics, 50(4):217–227, April 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s10701-020-00334-8.pdf.

Fletcher:2020:RCA

[4231] Samuel C. Fletcher. On representational capacities, with an application to general relativity. Foundations of Physics, 50(4):228–249, April 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Roberts:2020:RLE

[4232] Bryan W. Roberts. Regarding ‘Leibniz Equivalence’. Foundations of Physics, 50(4):250–269, April 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10701-020-00325-9.pdf.

Bradley:2020:RRS

[4233] Clara Bradley and James Owen Weatherall. On representational redundancy, surplus structure, and the hole argument. Foundations of Physics, 50(4):270–293, April 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Dewar:2020:GRC

[4234] Neil Dewar. General-relativistic covariance. Foundations of Physics, 50 (4):294–318, April 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Ladyman:2020:HAH

[4235] James Ladyman and Stuart Presnell. The hole argument in homotopy type theory. Foundations of Physics, 50(4):319–329, April 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s10701-019-00293-9.pdf.

Dougherty:2020:HAT

[4236] John Dougherty. The hole argument, take n. Foundations of Physics, 50(4):330–347, April 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Brighouse:2020:CCS

[4237] Carolyn Brighouse. Confessions of a (cheap) sophisticated substantivalist. Foundations of Physics, 50(4):348–359, April 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Norton:2020:HAA

[4238] Joshua Norton. The hole argument against everything. Foundations of Physics, 50(4):360–378, April 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Doboszewski:2020:SON

[4239] Juliusz Doboszewski. Some other ”no hole” spacetimes properties are unstable too. Foundations of Physics, 50(5):379–384, May 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s10701-019-00316-5.pdf.

Ling:2020:BBC

[4240] Eric Ling. The Big Bang is a coordinate singularity for k = −1 inflationary FLRW spacetimes. Foundations of Physics, 50(5):385– 428, May 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10701-020-00335-7.pdf.

Dragowski:2020:DEA

[4241] Michał Dragowski and Marta Włodarczyk. The Doppler effect and the anisotropy of the speed of light. Foundations of Physics, 50(5):429– 440, May 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10701-020-00337-5.pdf.

Kastner:2020:UOQ

[4242] R. E. Kastner. Unitary-only quantum theory cannot consistently describe the use of itself: On the Frauchiger–Renner Paradox. Foundations of Physics, 50(5):441–456, May 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Mainland:2020:PVF

[4243] G. B. Mainland and Bernard Mulligan. Polarization of vacuum fluctuations: Source of the vacuum permittivity and speed of light. Foundations of Physics, 50(5):457–480, May 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10701-020-00339-3.pdf.

Bishop:2020:CEP

[4244] Robert C. Bishop and George F. R. Ellis. Contextual emergence of physical properties. Foundations of Physics, 50(5):481–510, May 2020. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Wright:2020:CGT

[4245] Victoria J. Wright and Stefan Weigert. Correction to: Gleason-type theorems from Cauchy’s functional equation. Foundations of Physics, 50(5): 511–514, May 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10701-020-00340-w.pdf. See [4171].

Hofer-Szabo:2020:MP

[4246] Gábor Hofer-Szabó, Joanna Luc, and Tomasz Placek. Modality in physics. Foundations of Physics, 50(6):515–521, June 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:

//link.springer.com/content/pdf/10.1007/s10701-020-00344-6.
pdf.

Maudlin:2020:MFL

[4247] Tim Maudlin. A modal free lunch. Foundations of Physics, 50(6):522– 529, June 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Gomori:2020:VID

[4248] Márton Gömöri. On the very idea of distant correlations. Foundations of Physics, 50(6):530–554, June 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Szabo:2020:IEC

[4249] László E. Szabó. Intrinsic, extrinsic, and the constitutive a priori. Foundations of Physics, 50(6):555–567, June 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10701-019-00281-z.pdf.

Gyenis:2020:DPP

[4250] Balázs Gyenis. Determinism, physical possibility, and laws of nature. Foundations of Physics, 50(6):568–581, June 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10701-020-00320-0.pdf.

Fletcher:2020:WWR

[4251] Samuel C. Fletcher. Which worldlines represent possible particle histories? Foundations of Physics, 50(6):582–599, June 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Placek:2020:MAC

[4252] Tomasz Placek. On the modal aspects of causal sets. Foundations of Physics, 50(6):600–620, June 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10701-019-00280-0.pdf.

Luc:2020:GMB

[4253] Joanna Luc. Generalised manifolds as basic objects of general relativity. Foundations of Physics, 50(6):621–643, June 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10701-019-00292-w.pdf.

Muller:2020:DRP

[4254] Thomas Müller. Defining a relativity-proof notion of the present via spatio-temporal indeterminism. Foundations of Physics, 50(6):644– 664, June 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10701-019-00267-x.pdf.

Arve:2020:EMP

[4255] Per Arve. Everett’s missing postulate and the Born rule. Foundations of Physics, 50(7):665–692, July 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Barzegar:2020:QSD

[4256] Ali Barzegar. QBism is not so simply dismissed. Foundations of Physics, 50(7):693–707, July 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Uzan:2020:CTC

[4257] Pierre Uzan. The CHSH test can be used for all quantum-describable systems. Foundations of Physics, 50(7):708–714, July 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

He:2020:CIC

[4258] Guang Ping He. Copenhagen interpretation can survive the upgraded Schrödinger’s cat Gedankenexperiment. Foundations of Physics, 50(7): 715–726, July 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Buchholz:2020:PID

[4259] Detlev Buchholz and Klaus Fredenhagen. From path integrals to dynamical algebras: a macroscopic view of quantum physics. Foundations of Physics, 50(7):727–734, July 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10701-020-00345-5.pdf.

Vieira:2020:SAI

[4260] S. L. R. Vieira and K. Bakke. Semiclassical analysis of the interaction of the magnetic quadrupole moment of a neutral particle with axial electric fields in a uniformly rotating frame. Foundations of Physics, 50(7):735– 748, July 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Solis-Labastida:2020:AMP

[4261] Aldo F. G. Solis-Labastida and Jorge G. Hirsch. Algorithmic measurement procedures. Foundations of Physics, 50(8):749–763, August 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Barroso:2020:QFT

[4262] Filipe C. R. Barroso and Orfeu Bertolami. A quantum field theory view of interaction free measurements. Foundations of Physics, 50(8):764–771, August 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Renziehausen:2020:CBB

[4263] Klaus Renziehausen and Ingo Barth. The connection between Bohmian mechanics and many-particle quantum hydrodynamics. Foundations of Physics, 50(8):772–798, August 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10701-020-00349-1.pdf.

Forgione:2020:PIH

[4264] Marco Forgione. Path integrals and holism. Foundations of Physics, 50 (8):799–827, August 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Roza:2020:SDM

[4265] Engel Roza. On the second dipole moment of Dirac’s particle. Foundations of Physics, 50(8):828–849, August 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Filippov:2020:ORG

[4266] Sergey N. Filippov, Stan Gudder, Teiko Heinosaari, and Leevi Leppäjärvi. Operational restrictions in general probabilistic theories. Foundations of Physics, 50(8):850–876, August 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10701-020-00352-6.pdf.

Stergiou:2020:EUP

[4267] Chrysovalantis Stergiou. Empirical underdetermination for physical theories in C^{∗} algebraic setting: Comments to an Arageorgis’s argument. Foundations of Physics, 50(9):877–892, September 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-020-00358-0.

Capozziello:2020:ACH

[4268] Salvatore Capozziello, Micol Benetti, and Alessandro D. A. M. Spallicci. Addressing the cosmological H_{0} tension by the Heisenberg uncertainty. Foundations of Physics, 50(9):893–899, September 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-020-00356-2.

Poplawski:2020:NMT

[4269] Nikodem Popławski. Noncommutative momentum and torsional regularization. Foundations of Physics, 50(9):900–923, September 2020. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

https://link.springer.com/article/10.1007/s10701-020-00357-
1.

Poinat:2020:QMI

[4270] Sébastien Poinat. Quantum mechanics and its interpretations: a defense of the quantum principles. Foundations of Physics, 50(9):924– 941, September 2020. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-020-00360-6.

Samar:2020:KPS

[4271] M. I. Samar and V. M. Tkachuk. Kepler problem in space with deformed Lorentz-covariant Poisson brackets. Foundations of Physics, 50(9):942– 959, September 2020. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-020-00359-z.

Yang:2020:TIC

[4272] Ciann-Dong Yang and Shiang-Yi Han. Trajectory interpretation of correspondence principle: Solution of nodal issue. Foundations of Physics, 50(9):960–976, September 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.com/

article/10.1007/s10701-020-00363-3.

Nistico:2020:GTD

[4273] Giuseppe Nisticò. Group theoretical derivation of consistent free particle theories. Foundations of Physics, 50(9):977–1007, September 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-020-00364-2.

Glick:2020:MNM

[4274] Jennifer R. Glick and Christoph Adami. Markovian and non-Markovian quantum measurements. Foundations of Physics, 50(9):1008–1055, September 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL https://link.springer.com/article/10.
1007/s10701-020-00362-4.

Ellis:2020:CCP

[4275] George F. R. Ellis. The causal closure of physics in real world contexts. Foundations of Physics, 50(10):1057–1097, October 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-020-00366-0.

Ellis:2020:ESS

[4276] George F. R. Ellis. Emergence in solid state physics and biology. Foundations of Physics, 50(10):1098–1139, October 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-020-00367-z. See response

[4277].

Luu:2020:MEF

[4277] Thomas Luu and Ulf-G. Meißner. Misconceptions on effective field theories and spontaneous symmetry breaking: Response to Ellis’ article. Foundations of Physics, 50(10):1140–1151, October 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-020-00368-y. See

[4276].

Jebeile:2020:KRA

[4278] Julie Jebeile. The Kac ring or the art of making idealisations. Foundations of Physics, 50(10):1152–1170, October 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-020-00373-1.

Bigaj:2020:HHB

[4279] Tomasz Bigaj and Antonio Vassallo. How Humean is Bohumianism? Foundations of Physics, 50(10):1171–1188, October 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-020-00361-5.

Barandes:2020:MQD

[4280] Jacob A. Barandes and David Kagan. Measurement and quantum dynamics in the minimal modal interpretation of quantum theory. Foundations of Physics, 50(10):1189–1218, October 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-020-00374-0.

Myrvold:2020:S

[4281] Wayne C. Myrvold. The science of ΘΔ^{cs}. Foundations of Physics, 50 (10):??, October 2020. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-020-00371-3.

Trzetrzelewski:2020:EPR

[4282] Maciej Trzetrzelewski. On the equivalence principle and relativistic quantum mechanics. Foundations of Physics, 50(11):1253–1269, November 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-020-00388-8.

Rosa:2020:CRB

[4283] Luiz Pinguelli Rosa, Elaine Andrade, and Jean Faber. Constructivism and realism in Boltzmann’s thermodynamics’ atomism. Foundations of Physics, 50(11):1270–1293, November 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-020-00372-2.

Elouard:2020:IFQ

[4284] Cyril Elouard, Mordecai Waegell, and Andrew N. Jordan. An interactionfree quantum measurement-driven engine. Foundations of Physics, 50 (11):1294–1314, November 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.com/

article/10.1007/s10701-020-00381-1.

Oldofredi:2020:CBP

[4285] Andrea Oldofredi and Cristian López. On the classification between ψ-ontic and ψ-epistemic ontological models. Foundations of Physics, 50(11):1315–1345, November 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-020-00377-x.

Makukov:2020:TPE

[4286] Maxim Makukov and Eduard Mychelkin. Triple path to the exponential metric. Foundations of Physics, 50(11):1346–1355, November 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-020-00384-y.

Colombo:2020:ESD

[4287] Fabrizio Colombo and Giovanni Valente. Evolution of superoscillations in the Dirac field. Foundations of Physics, 50(11):1356–1375, November 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL https://link.springer.com/article/10.
1007/s10701-020-00382-0.

Carneiro:2020:ECN

[4288] Saulo Carneiro. Elementary charge and neutrino’s mass from Planck length. Foundations of Physics, 50(11):1376–1381, November 2020. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

https://link.springer.com/article/10.1007/s10701-020-00383-
z.

Higashi:2020:HRC

[4289] Katsuaki Higashi. Hardy relations and common cause. Foundations of Physics, 50(11):1382–1397, November 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-020-00392-y.

Slavov:2020:EPR

[4290] Matias Slavov. Eternalism and perspectival realism about the ‘now’. Foundations of Physics, 50(11):1398–1410, November 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-020-00385-x.

Binder:2020:IP

[4291] P.-M. Binder. Intuitionist physics. Foundations of Physics, 50(11):1411– 1417, November 2020. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-020-00365-1.

Dewar:2020:CIN

[4292] Neil Dewar and James Read. Conformal invariance of the Newtonian Weyl tensor. Foundations of Physics, 50(11):1418–1425, November 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-020-00386-w.

Goyal:2020:DCM

[4293] Philip Goyal. Derivation of classical mechanics in an energetic framework via conservation and relativity. Foundations of Physics, 50(11): 1426–1479, November 2020. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-020-00376-y.

Schmidt:2020:CAQ

[4294] Heinz-Jürgen Schmidt. Conditional action and quantum versions of Maxwell’s demon. Foundations of Physics, 50(11):1480–1508, November 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/

s10701-020-00387-9. See correction [4413].

Braak:2020:FGR

[4295] D. Braak and J. Mannhart. Fermi’s Golden Rule and the Second Law of Thermodynamics. Foundations of Physics, 50(11):1509–1540, November 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL https://link.springer.com/article/10.
1007/s10701-020-00380-2.

Gao:2020:PFO

[4296] Shan Gao. A puzzle for the field ontologists. Foundations of Physics, 50(11):1541–1553, November 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-020-00390-0.

Gavassino:2020:ZLT

[4297] L. Gavassino. The zeroth law of thermodynamics in special relativity. Foundations of Physics, 50(11):1554–1586, November 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-020-00393-x.

Santos:2020:LME

[4298] Emilio Santos. Local model of entangled photon experiments compatible with quantum predictions based on the reality of the vacuum fields. Foundations of Physics, 50(11):1587–1607, November 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-020-00395-9.

Varieschi:2020:NFD

[4299] Gabriele U. Varieschi. Newtonian fractional-dimension gravity and MOND. Foundations of Physics, 50(11):1608–1644, November 2020. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

https://link.springer.com/article/10.1007/s10701-020-00389-

7. See correction [4356].

Luc:2020:QSU

[4300] Joanna Luc. Quantumness of states and unitary operations. Foundations of Physics, 50(11):1645–1685, November 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-020-00391-z.

Kholmetskii:2020:PPB

[4301] A. L. Kholmetskii, O. V. Missevitch, and R. Smirnov-Rueda. Propagation properties of bound electromagnetic field: Classical and quantum viewpoints. Foundations of Physics, 50(11):1686–1722, November 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-020-00396-8.

Knight:2020:NPW

[4302] Andrew Knight. No paradox in wave–particle duality. Foundations of Physics, 50(11):1723–1727, November 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-020-00379-9.

Vaidman:2020:TNN

[4303] Lev Vaidman. There is no new problem for quantum mechanics. Foundations of Physics, 50(11):1728–1734, November 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-020-00394-w.

Khrennikov:2020:PSI

[4304] Andrei Khrennikov, Hans de Raedt, and Sergey Polyakov. Preface of the special issue probing the limits of quantum mechanics: Theory and experiment, volume 2. Foundations of Physics, 50(11):1735– 1738, November 2020. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-015-9950-1.

Cooperman:2020:MCC

[4305] Joshua H. Cooperman. Making the case for causal dynamical triangulations. Foundations of Physics, 50(11):1739–1755, November 2020. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

https://link.springer.com/article/10.1007/s10701-015-9972-8.

Fuchs:2020:PSI

[4306] Christopher A. Fuchs and Andrei Khrennikov. Preface to special issue: Quantum information revolution: Impact to foundations. Foundations of Physics, 50(12):1757–1761, December 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-020-00401-0.

Khrennikov:2020:QVC

[4307] Andrei Khrennikov. Quantum versus classical entanglement: Eliminating the issue of quantum nonlocality. Foundations of Physics, 50(12): 1762–1780, December 2020. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-020-00319-7.

Auffeves:2020:DBR

[4308] Alexia Auffèves and Philippe Grangier. Deriving Born’s rule from an inference to the best explanation. Foundations of Physics, 50(12):1781– 1793, December 2020. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-020-00326-8.

Bengtsson:2020:SSE

[4309] Ingemar Bengtsson. SICs: Some explanations. Foundations of Physics, 50(12):1794–1808, December 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-020-00341-9.

Hofmann:2020:QCR

[4310] Holger F. Hofmann. Quantum causality relations and the emergence of reality from coherent superpositions. Foundations of Physics, 50(12): 1809–1823, December 2020. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-020-00346-4.

Plotnitsky:2020:UIB

[4311] Arkady Plotnitsky. “The unavoidable interaction between the object and the measuring instruments”: Reality, probability, and nonlocality in quantum physics. Foundations of Physics, 50(12):1824–1858, December 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-020-00353-5.

DeBrota:2020:ROF

[4312] John B. DeBrota, Christopher A. Fuchs, and Rüdiger Schack. Respecting one’s fellow: QBism’s analysis of Wigner’s friend. Foundations of Physics, 50(12):1859–1874, December 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-020-00369-x.

Elze:2020:QSS

[4313] Hans-Thomas Elze. Are quantum spins but small perturbations of ontological Ising spins? Foundations of Physics, 50(12):1875–1893, December 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-020-00370-4.

Pienaar:2020:EAQ

[4314] Jacques Pienaar. Extending the agent in QBism. Foundations of Physics, 50(12):1894–1920, December 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-020-00375-z.

DAriano:2020:NPO

[4315] Giacomo Mauro D’Ariano. No purification ontology, no quantum paradoxes. Foundations of Physics, 50(12):1921–1933, December 2020. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

https://link.springer.com/article/10.1007/s10701-020-00398-
6.

Naus:2021:QMM

[4316] H. W. L. Naus. On the quantum mechanical measurement process. Foundations of Physics, 51(1):??, February 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-021-00404-5.

Luo:2021:RFA

[4317] M. J. Luo. Ricci flow approach to the cosmological constant problem. Foundations of Physics, 51(1):??, February 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-021-00405-4.

Chirco:2021:SMC

[4318] Goffredo Chirco and Thibaut Josset. Statistical mechanics of covariant systems with multi-fingered time. Foundations of Physics, 51 (1):??, February 2021. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-021-00406-3.

Singal:2021:CPE

[4319] Ashok K. Singal. Contribution of pressure to the energy-momentum density in a moving perfect fluid: a physical perspective. Foundations of Physics, 51(1):??, February 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-021-00403-6. See comment [4462].

Stergiou:2021:CPA

[4320] Chrysovalantis Stergiou. Causal processes in C∗-algebraic setting. Foundations of Physics, 51(1):??, February 2021. CODEN FNDPA4.

ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-021-00411-6.

Saldanha:2021:LDA

[4321] Pablo L. Saldanha. Local description of the Aharonov–Bohm effect with a quantum electromagnetic field. Foundations of Physics, 51 (1):??, February 2021. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-021-00414-3.

Barandes:2021:GIC

[4322] Jacob A. Barandes. Gauge invariance for classical massless particles with spin. Foundations of Physics, 51(1):??, February 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-021-00415-2. See

correction [4385].

Kozlova:2021:POR

[4323] Svetlana G. Kozlova and Denis P. Pishchur. On the possibility to observe relations between quantum measurements and the entropy of phase transitions in Zn 2 (BDC) 2 (DABCO). Foundations of Physics, 51 (1):??, February 2021. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-021-00416-1.

Sahin:2021:CPL

[4324] Inanç Sahin. Can parallel lives provide a solution to Hardy’s Paradox? Foundations of Physics, 51(1):??, February 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-021-00418-z.

Gokler:2021:QBC

[4325] Can Gokler. Quantum behavior of a classical particle subject to a random force. Foundations of Physics, 51(1):??, February 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-021-00422-3.

Searight:2021:MMU

[4326] Trevor P. Searight. Mirror matter from a unified field theory. Foundations of Physics, 51(1):??, February 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-021-00421-4.

Hance:2021:HQQ

[4327] Jonte R. Hance, James Ladyman, and John Rarity. How quantum is quantum counterfactual communication? Foundations of Physics, 51 (1):??, February 2021. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-021-00412-5. See correction [4355].

Jaeger:2021:EFP

[4328] Gregg Jaeger. Exchange forces in particle physics. Foundations of Physics, 51(1):??, February 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.com/

article/10.1007/s10701-021-00425-0.

Aliakbarzadeh:2021:CAI

[4329] Mojtaba Aliakbarzadeh and Kirsty Kitto. Is contextuality about the identity of random variables? Foundations of Physics, 51(1):??, February 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-021-00402-7.

Bopp:2021:IQS

[4330] Fritz W. Bopp. An intricate quantum statistical effect and the foundation of quantum mechanics. Foundations of Physics, 51(1):??, February 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-021-00408-1.

Khrennikov:2021:QPV

[4331] Andrei Khrennikov. Quantum postulate vs. quantum nonlocality: on the role of the Planck constant in Bell’s argument. Foundations of Physics, 51(1):??, February 2021. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-021-00430-3.

Meehan:2021:CNP

[4332] Alexander Meehan. Clarifying the new problem for quantum mechanics: Reply to Vaidman. Foundations of Physics, 51(1):??, February 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-021-00438-9.

Oldofredi:2021:BTA

[4333] Andrea Oldofredi. The bundle theory approach to relational quantum mechanics. Foundations of Physics, 51(1):??, February 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-021-00407-2.

Yeter:2021:ATC

[4334] U. Yeter, K. Sogut, and M. Salti. Anisotropic teleparallel cosmology via thermodynamics. Foundations of Physics, 51(1):??, February 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-021-00435-y.

Koberinski:2021:RVE

[4335] Adam Koberinski. Regularizing (away) vacuum energy. Foundations of Physics, 51(1):??, February 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-021-00442-z.

Roselli:2021:DAP

[4336] Carlo Roselli and Bruno Raffaele Stella. The dead–alive physicist experiment: a case-study against the hypothesis that consciousness causes the wave-function collapse in the quantum mechanical measurement process. Foundations of Physics, 51(1):??, February 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-021-00427-y. See correction

[4366].

Barrow:2021:CPP

[4337] John D. Barrow and João Magueijo. A contextual Planck parameter and the classical limit in quantum cosmology. Foundations of Physics, 51(1):??, February 2021. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-021-00433-0.

Oks:2021:RTD

[4338] E. Oks. Role of a time delay in the gravitational two-body problem. Foundations of Physics, 51(1):??, February 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-021-00443-y.

Martin-Dussaud:2021:RSF

[4339] Pierre Martin-Dussaud. Relational structures of fundamental theories. Foundations of Physics, 51(1):??, February 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-021-00444-x.

Borzou:2021:NGQ

[4340] Ahmad Borzou. A non-geometrodynamic quantum Yang–Mills theory of gravity based on the homogeneous Lorentz group. Foundations of Physics, 51(1):??, February 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.com/

article/10.1007/s10701-021-00410-7.

Perlman:2021:WQM

[4341] H. S. Perlman. Why quantum measurements yield single values. Foundations of Physics, 51(1):??, February 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-021-00440-1.

Cosgrove:2021:EPE

[4342] Joseph K. Cosgrove. Einstein’s principle of equivalence and the heuristic significance of general covariance. Foundations of Physics, 51(1): ??, February 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL https://link.springer.com/article/10.
1007/s10701-021-00434-z.

Cuffaro:2021:ERT

[4343] Michael E. Cuffaro and Emerson P. Doyle. Essay review of tanya and Jeffrey Bub’s totally random: Why nobody understands quantum mechanics: a serious comic on entanglement. Foundations of Physics, 51 (1):??, February 2021. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-021-00423-2.

Romano:2021:AES

[4344] Davide Romano. On the alleged extra-structures of quantum mechanics. Foundations of Physics, 51(1):??, February 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-021-00426-z.

DiBiagio:2021:SFR

[4345] Andrea Di Biagio and Carlo Rovelli. Stable facts, relative facts. Foundations of Physics, 51(1):??, February 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-021-00429-w.

Metcalf:2021:FNF

[4346] Thomas Metcalf. On Friederich’s new fine-tuning argument. Foundations of Physics, 51(2):??, April 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-021-00449-6.

Reichert:2021:ITC

[4347] Paula Reichert. Investigating total collisions of the Newtonian Nbody problem on shape space. Foundations of Physics, 51(2):??, April 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-021-00428-x.

Held:2021:NCL

[4348] Carsten Held. Non-contextual and local hidden-variable model for the Peres–Mermin and Greenberger–Horne–Zeilinger systems. Foundations of Physics, 51(2):??, April 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.com/

article/10.1007/s10701-021-00409-0.

Tanaka:2021:ATT

[4349] Shigenori Tanaka. Appearance of thermal time. Foundations of Physics, 51(2):??, April 2021. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-021-00445-w.

Padmanabhan:2021:WLP

[4350] T. Padmanabhan. World-line path integral for the propagator expressed as an ordinary integral: Concept and applications. Foundations of Physics, 51(2):??, April 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.com/

article/10.1007/s10701-021-00447-8.

Bub:2021:UFR

[4351] Jeffrey Bub. Understanding the frauchiger-renner argument. Foundations of Physics, 51(2):??, April 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-021-00420-5.

Masoumi:2021:CGN

[4352] Saeed Masoumi. On the continuity of geometrized Newtonian gravitation and general relativity. Foundations of Physics, 51(2):??, April 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-021-00419-y.

Hermens:2021:HRQ

[4353] R. Hermens. How real are quantum states in ψ-ontic models? Foundations of Physics, 51(2):??, April 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-021-00448-7.

Cavalcanti:2021:VWB

[4354] Eric G. Cavalcanti. The view from a Wigner bubble. Foundations of Physics, 51(2):??, April 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.com/

article/10.1007/s10701-021-00417-0.

Hance:2021:CHQ

[4355] Jonte R. Hance, James Ladyman, and John Rarity. Correction to: How quantum is quantum counterfactual communication? Foundations of Physics, 51(2):??, April 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-021-00450-z. See [4327].

Varieschi:2021:CNF

[4356] Gabriele U. Varieschi. Correction to: Newtonian Fractional-Dimension Gravity and MOND. Foundations of Physics, 51(2):??, April 2021. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

https://link.springer.com/article/10.1007/s10701-021-00439-

8. See [4299].

Landsman:2021:SBH

[4357] Klaas Landsman. Singularities, black holes, and cosmic censorship: a tribute to Roger Penrose. Foundations of Physics, 51(2):??, April 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-021-00432-1.

Lorenzetti:2021:RPA

[4358] Lorenzo Lorenzetti. A refined propensity account for GRW theory. Foundations of Physics, 51(2):??, April 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-021-00455-8.

Capellmann:2021:STQ

[4359] H. Capellmann. Space-time in quantum theory. Foundations of Physics, 51(2):??, April 2021. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-021-00441-0.

Waaijer:2021:RAF

[4360] Marijn Waaijer and Jan van Neerven. Relational analysis of the frauchiger-renner paradox and interaction-free detection of records from the past. Foundations of Physics, 51(2):??, April 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-021-00413-4.

Tavernelli:2021:GQD

[4361] Ivano Tavernelli. Gravitational quantum dynamics: a geometrical perspective. Foundations of Physics, 51(2):??, April 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-021-00436-x.

Schonfeld:2021:FDC

[4362] Jonathan F. Schonfeld. The first droplet in a cloud chamber track. Foundations of Physics, 51(2):??, April 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-021-00452-x.

Mulder:2021:GUS

[4363] Ruward A. Mulder. Gauge-underdetermination and shades of locality in the Aharonov–Bohm effect. Foundations of Physics, 51(2):??, April 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-021-00446-9.

Feintzeig:2021:LPC

[4364] Benjamin H. Feintzeig, Jonah Librande, and Rory Soiffer. Localizable particles in the classical limit of quantum field theory. Foundations of Physics, 51(2):??, April 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.com/

article/10.1007/s10701-021-00458-5.

Sharma:2021:WWF

[4365] Umesh Kumar Sharma and Ambuj Kumar Mishra. Wormholes within the framework of f(R, T) = R + αR^{2} + λT gravity. Foundations of Physics, 51(2):??, April 2021. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-021-00457-6.

Roselli:2021:CDA

[4366] Carlo Roselli and Bruno Raffaele Stella. Correction to: The Dead– Alive Physicist Experiment: A Case-Study Against the Hypothesis that Consciousness Causes the Wave-Function Collapse in the Quantum Mechanical Measurement Process. Foundations of Physics, 51(2): ??, April 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-021-00451-y. See [4336].

Mauri:2021:NLP

[4367] Roberto Mauri. A non-local phase field model of Bohm’s quantum potential. Foundations of Physics, 51(2):??, April 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-021-00454-9.

Linnemann:2021:SNG

[4368] Niels Linnemann and James Read. On the status of Newtonian gravitational radiation. Foundations of Physics, 51(2):??, April 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-021-00453-w.

Aurell:2021:QBH

[4369] Erik Aurell, Michał Eckstein, and Paweł Horodecki. Quantum black holes as solvents. Foundations of Physics, 51(2):??, April 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-021-00456-7.

Kaszlikowski:2021:LBC

[4370] Dagomir Kaszlikowski and Paweł Kurzyński. A little bit of classical magic to achieve (super-)quantum speedup. Foundations of Physics, 51(3):??, June 2021. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-021-00461-w.

Campanelli:2021:CNM

[4371] Leonardo Campanelli. A conjecture on the neutrality of matter. Foundations of Physics, 51(3):??, June 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-021-00462-9.

Zwirn:2021:PD

[4372] Herve Zwirn. Is the past determined? Foundations of Physics, 51(3):??, June 2021. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-021-00459-4.

Shaikh:2021:PBH

[4373] A. Y. Shaikh and K. S. Wankhade. Panorama behaviors of holographic dark energy models in modified gravity. Foundations of Physics, 51(3):??, June 2021. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-021-00463-8.

Weaver:2021:PCE

[4374] Christopher Gregory Weaver. In praise of Clausius entropy: Reassessing the foundations of Boltzmannian statistical mechanics. Foundations of Physics, 51(3):??, June 2021. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-021-00437-w.

deGosson:2021:QPD

[4375] Maurice A. de Gosson. Quantum polar duality and the symplectic camel: a new geometric approach to quantization. Foundations of Physics, 51(3):??, June 2021. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-021-00465-6.

Cabello:2021:BNL

[4376] Adán Cabello. Bell non-locality and Kochen–Specker contextuality: How are they connected? Foundations of Physics, 51(3):??, June 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-021-00466-5.

Kent:2021:CMC

[4377] Adrian Kent. Collapse and measures of consciousness. Foundations of Physics, 51(3):??, June 2021. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-021-00467-4.

Hooft:2021:FVF

[4378] Gerard ’t Hooft. Fast vacuum fluctuations and the emergence of quantum mechanics. Foundations of Physics, 51(3):??, June 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-021-00464-7.

Gradenigo:2021:SQI

[4379] Giacomo Gradenigo. Symplectic quantization II: Dynamics of space– time quantum fluctuations and the cosmological constant. Foundations of Physics, 51(3):??, June 2021. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-021-00468-3.

Anastopoulos:2021:MBI

[4380] Charis Anastopoulos. Mind-body interaction and modern physics. Foundations of Physics, 51(3):??, June 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-021-00472-7.

Gradenigo:2021:SQD

[4381] Giacomo Gradenigo and Roberto Livi. Symplectic quantization I: Dynamics of quantum fluctuations in a relativistic field theory. Foundations of Physics, 51(3):??, June 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-021-00470-9.

Voracek:2021:GKS

[4382] Václav Vorácek and Mirko Navara. Generalised Kochen–Specker theorem in three dimensions. Foundations of Physics, 51(3):??, June 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-021-00476-3.

Newman:2021:RPT

[4383] Andrew Newman. The rates of the passing of time, presentism, and the issue of co-existence in special relativity. Foundations of Physics, 51(3):??, June 2021. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-021-00469-2.

Stein:2021:PPM

[4384] Howard Stein. Physics and philosophy meet: the strange case of Poincaré. Foundations of Physics, 51(3):??, June 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-021-00460-x.

Barandes:2021:CGI

[4385] Jacob A. Barandes. Correction to: Gauge invariance for classical massless particles with spin. Foundations of Physics, 51(3):??, June 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-021-00471-8. See

[4322].

Rastogi:2021:TTS

[4386] Pallavi Rastogi and Shripad P. Mahulikar. Thermodynamic theory for simple and complex dissipative structures. Foundations of Physics, 51(3):??, June 2021. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-021-00473-6.

Cheng:2021:WSP

[4387] Bryan Cheng and James Read. Why not a sound postulate? Foundations of Physics, 51(3):??, June 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-021-00479-0.

Acuna:2021:NTR

[4388] Pablo Acuña. von Neumann’s theorem revisited. Foundations of Physics, 51(3):??, June 2021. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-021-00474-5.

Ni:2021:QFT

[4389] Jun Ni. Quantum field theory formulated as a Markov process determined by local configuration. Foundations of Physics, 51(3):??, June 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-021-00481-6.

Sebens:2021:ECD

[4390] Charles T. Sebens. Electron charge density: a clue from quantum chemistry for quantum foundations. Foundations of Physics, 51(4):??, August 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-021-00480-7.

Grangier:2021:CQF

[4391] Philippe Grangier. Completing the quantum formalism in a contextually objective framework. Foundations of Physics, 51(4):??, August 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-021-00424-1.

Jackson:2021:MPO

[4392] Peter A. Jackson and John S. Minkowski. The measurement problem, an ontological solution. Foundations of Physics, 51(4):??, August 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-021-00475-4.

Bolotin:2021:WPD

[4393] Arkady Bolotin. Wave–particle duality and the objectiveness of “True” and “False”. Foundations of Physics, 51(4):??, August 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-021-00478-1.

Barandes:2021:MFW

[4394] Jacob A. Barandes. On magnetic forces and work. Foundations of Physics, 51(4):??, August 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.com/

article/10.1007/s10701-021-00483-4.

Maiezza:2021:HTR

[4395] Alessio Maiezza and Juan Carlos Vasquez. On Haag’s theorem and renormalization ambiguities. Foundations of Physics, 51(4):??, August 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-021-00484-3.

Singal:2021:INZ

[4396] Ashok K. Singal. Implications of a non-zero Poynting flux at infinity sans radiation reaction for a uniformly accelerated charge. Foundations of Physics, 51(4):??, August 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-021-00486-1.

Oldofredi:2021:RQM

[4397] Andrea Oldofredi and Caludio Calosi. Relational quantum mechanics and the PBR Theorem: a peaceful coexistence. Foundations of Physics, 51(4):??, August 2021. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-021-00485-2.

Carroll:2021:ENC

[4398] Sean M. Carroll and Jackie Lodman. Energy non-conservation in quantum mechanics. Foundations of Physics, 51(4):??, August 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-021-00490-5.

Lambare:2021:NBT

[4399] Justo Pastor Lambare and Rodney Franco. A note on Bell’s Theorem logical consistency. Foundations of Physics, 51(4):??, August 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-021-00488-z.

Buniy:2021:MSS

[4400] Roman V. Buniy and Stephen D. H. Hsu. Macroscopic superposition states in isolated quantum systems. Foundations of Physics, 51 (4):??, August 2021. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-021-00477-2.

Weststeijn:2021:WFR

[4401] Nikki Weststeijn. Wigner’s friend and relational quantum mechanics: a reply to Laudisa. Foundations of Physics, 51(4):??, August 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-021-00487-0.

Wilson-Ewing:2021:PCW

[4402] Edward Wilson-Ewing. Potential consequences of wormhole-mediated entanglement. Foundations of Physics, 51(4):??, August 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-021-00489-y.

Park:2021:IIE

[4403] Jeong-Wan Park. Imperfect instantaneous Energy Conservation in Wheeler–Feynman absorber theory. Foundations of Physics, 51(4):??, August 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-021-00491-4.

Kastner:2021:UID

[4404] R. E. Kastner. Unitary interactions do not yield outcomes: Attempting to model “Wigner’s friend”. Foundations of Physics, 51(4):??, August 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-021-00492-3.

Lavis:2021:BLB

[4405] David A. Lavis, Reimer Kühn, and Roman Frigg. Becoming large, becoming infinite: The anatomy of thermal physics and phase transitions in finite systems. Foundations of Physics, 51(5):??, October 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Azhar:2021:FTM

[4406] Feraz Azhar and Abraham Loeb. Finely tuned models sacrifice explanatory depth. Foundations of Physics, 51(5):??, October 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Haji:2021:SRI

[4407] Amir Hossein Haji. Sequential rearrangement of information in formation of patterns. Foundations of Physics, 51(5):??, October 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Tolksdorf:2021:DCC

[4408] Jürgen Tolksdorf and Rainer Verch. The D-CTC condition is generically fulfilled in classical (non-quantum) statistical systems. Foundations of Physics, 51(5):??, October 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Katsnelson:2021:EQN

[4409] Mikhail I. Katsnelson and Vitaly Vanchurin. Emergent quantumness in neural networks. Foundations of Physics, 51(5):??, October 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Benavoli:2021:WTO

[4410] Alessio Benavoli, Alessandro Facchini, and Marco Zaffalon. The Weirdness Theorem and the origin of quantum paradoxes. Foundations of Physics, 51(5):??, October 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Pienaar:2021:QRQ

[4411] Jacques Pienaar. QBism and relational quantum mechanics compared. Foundations of Physics, 51(5):??, October 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Pienaar:2021:QQF

[4412] Jacques Pienaar. A quintet of quandaries: Five no-go theorems for relational quantum mechanics. Foundations of Physics, 51(5):??, October 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Schmidt:2021:CCA

[4413] Heinz-Jürgen Schmidt. Correction to: Conditional Action and Quantum Versions of Maxwell’s Demon. Foundations of Physics, 51(5):??, October 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). See [4294].

Perelman:2021:FIT

[4414] Carlos Castro Perelman. Further insights into thermal relativity theory and black hole thermodynamics. Foundations of Physics, 51(5):??, October 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Snoke:2021:MSC

[4415] D. W. Snoke. A model of spontaneous collapse with energy conservation. Foundations of Physics, 51(5):??, October 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Safranek:2021:BIO

[4416] Dominik Safránek, Anthony Aguirre, and J. M. Deutsch. A brief introduction to observational entropy. Foundations of Physics, 51(5):??, October 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Trompetter:2021:EAM

[4417] W. J. Trompetter. Electromagnetic angular momentum of an orbiting charge. Foundations of Physics, 51(5):??, October 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Gao:2021:NGR

[4418] Shan Gao. A no-go result for QBism. Foundations of Physics, 51(5):??, October 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Ferber:2021:NBA

[4419] R. Ferber. A numbers-based approach to a free particle’s proper spacetime. Foundations of Physics, 51(5):??, October 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). See correction [4444].

Price:2021:ESA

[4420] Huw Price and Ken Wharton. Entanglement swapping and action at a distance. Foundations of Physics, 51(6):??, December 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-021-00511-3.

Adlam:2021:CFT

[4421] Emily Adlam. Contextuality, fine-tuning and teleological explanation. Foundations of Physics, 51(6):??, December 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-021-00516-y.

Mouchet:2021:PIM

[4422] Amaury Mouchet. Path integrals in a multiply-connected configuration space (50 years after). Foundations of Physics, 51(6):??, December 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-021-00497-y.

Resende:2021:ATP

[4423] Pedro Resende. An abstract theory of physical measurements. Foundations of Physics, 51(6):??, December 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-021-00513-1.

Pitts:2021:CHG

[4424] J. Brian Pitts. Change in Hamiltonian general relativity with spinors. Foundations of Physics, 51(6):??, December 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-021-00509-x.

Sbisa:2021:DVH

[4425] Fulvio Sbisà. The divergence of Van Hove’s model and its consequences. Foundations of Physics, 51(6):??, December 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-021-00517-x.

Glowacki:2021:IPB

[4426] Jan Głowacki. Inevitability of the Poisson bracket structure of the relativistic constraints. Foundations of Physics, 51(6):??, December 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-021-00514-0.

Nistico:2021:GTD

[4427] Giuseppe Nisticò. Group theoretical derivation of consistent massless particle theories. Foundations of Physics, 51(6):??, December 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-021-00494-1.

Hanson:2022:CDE

[4428] Jason Hanson. Coupling the Dirac and Einstein equations through geometry. Foundations of Physics, 52(1):??, February 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-021-00519-9.

Rosaler:2022:DEF

[4429] Joshua Rosaler. Dogmas of effective field theory: Scheme dependence, fundamental parameters, and the many faces of the Higgs naturalness principle. Foundations of Physics, 52(1):??, February 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-021-00510-4.

Sobolev:2022:FTG

[4430] Alexander P. Sobolev. Foundations of a theory of gravity with a constraint and its canonical quantization. Foundations of Physics, 52 (1):??, February 2022. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-021-00521-1.

Svozil:2022:GGH

[4431] Karl Svozil. Generalized Greenberger–Horne–Zeilinger arguments from quantum logical analysis. Foundations of Physics, 52(1):??, February 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-021-00515-z.

Caponio:2022:NSE

[4432] Erasmo Caponio and Antonio Masiello. A note on the Sagnac effect in general relativity as a Finslerian effect. Foundations of Physics, 52 (1):??, February 2022. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-021-00523-z.

Kay:2022:QEG

[4433] Bernard S. Kay. Quantum electrostatics, Gauss’s Law, and a product picture for quantum electrodynamics; or, the temporal gauge revised. Foundations of Physics, 52(1):??, February 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-021-00512-2.

Schlatter:2022:FST

[4434] Andreas Schlatter. On the foundation of space and time by quantumevents. Foundations of Physics, 52(1):??, February 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-021-00526-w.

Bonilla-Licea:2022:BTB

[4435] Moise Bonilla-Licea and Dieter Schuch. Bohmian trajectories as borders of regions of constant probability. Foundations of Physics, 52 (1):??, February 2022. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-021-00525-x.

Tyagi:2022:SPI

[4436] Narayani Tyagi and Ken Wharton. Spacetime path integrals for entangled states. Foundations of Physics, 52(1):??, February 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-021-00520-2.

Fortin:2022:RET

[4437] Sebastian Fortin, Olimpia Lombardi, and Matías Pasqualini. Relational event-time in quantum mechanics. Foundations of Physics, 52 (1):??, February 2022. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-021-00528-8.

Gavassino:2022:PLI

[4438] L. Gavassino. Proving the Lorentz invariance of the entropy and the covariance of thermodynamics. Foundations of Physics, 52(1):??, February 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-021-00518-w.

Kitajima:2022:NMT

[4439] Yuichiro Kitajima. Negations and meets in topos quantum theory. Foundations of Physics, 52(1):??, February 2022. CODEN FNDPA4.

ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-021-00529-7.

Kujala:2022:CDR

[4440] Janne V. Kujala and Ehtibar N. Dzhafarov. Contextuality and dichotomizations of random variables. Foundations of Physics, 52(1): ??, February 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL https://link.springer.com/article/10.
1007/s10701-021-00527-9.

Lawrence:2022:OQM

[4441] Jay Lawrence. Observing a quantum measurement. Foundations of Physics, 52(1):??, February 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.com/

article/10.1007/s10701-021-00522-0.

Benitez:2022:RIE

[4442] Federico Benitez, Diego Romero-Maltrana, and Pablo Razeto-Barry. (re)interpreting E = mc^{2}. Foundations of Physics, 52(1):??, February 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-021-00530-0.

Martens:2022:DMR

[4443] Niels C. M. Martens. Dark matter realism. Foundations of Physics, 52(1):??, February 2022. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-021-00524-y.

Ferber:2022:CNB

[4444] R. Ferber. Correction to: A Numbers-Based Approach to a Free Particle’s Proper Spacetime. Foundations of Physics, 52(1):??, February 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-021-00532-y. See [4419].

Napoletani:2022:IOT

[4445] Domenico Napoletani and Daniele C. Struppa. Isolated objects and their evolution: A derivation of the propagator’s path integral for spinless elementary particles. Foundations of Physics, 52(1):??, February 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-021-00535-9.

Sbitnev:2022:QAS

[4446] Valeriy I. Sbitnev. Quaternion algebra on 4D superfluid quantum space-time. Dirac’s ghost fermion fields. Foundations of Physics, 52 (1):??, February 2022. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-021-00533-x.

Ipsen:2022:AWV

[4447] Asger C. Ipsen. Anomalous weak values are caused by disturbance. Foundations of Physics, 52(1):??, February 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-021-00534-w.

Blum:2022:JJA

[4448] Alexander S. Blum. Jácome (Jay) Armas (ed.): Review of “Conversations on Quantum Gravity”. Foundations of Physics, 52(1):??, February 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-021-00537-7.

Harrison:2022:RTQ

[4449] Alan K. Harrison. A realistic theory of quantum measurement. Foundations of Physics, 52(1):??, February 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-021-00536-8.

Spallicci:2022:HLC

[4450] Alessandro D. A. M. Spallicci, Micol Benetti, and Salvatore Capozziello. The Heisenberg limit at cosmological scales. Foundations of Physics, 52(1):??, February 2022. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-021-00531-z.

Gao:2022:RQM

[4451] Shan Gao. Is retrocausal quantum mechanics consistent with special relativity? Foundations of Physics, 52(1):??, February 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-022-00543-3.

Wani:2022:QIA

[4452] Salman Sajad Wani, James Q. Quach, and Behnam Pourhassan. A quantum informational approach to the problem of time. Foundations of Physics, 52(1):??, February 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-022-00540-6.

Paganini:2022:NEC

[4453] Claudio F. Paganini. No events on closed causal curves. Foundations of Physics, 52(1):??, February 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-022-00542-4.

Beesley:2022:ECP

[4454] Jonathan J. Beesley. On the equivalence of causal propagators of the Dirac equation in vacuum-destabilising external fields. Foundations of Physics, 52(1):??, February 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.com/

article/10.1007/s10701-022-00538-0.

Adlam:2022:LNC

[4455] Emily Adlam. Laws of nature as constraints. Foundations of Physics, 52(1):??, February 2022. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-022-00546-0.

Poveda:2022:NRA

[4456] Luis A. Poveda, Luis Grave de Peralta, and Bill Poirier. A non-relativistic approach to relativistic quantum mechanics: The case of the harmonic oscillator. Foundations of Physics, 52(1):??, February 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-022-00541-5.

Tessarotto:2022:CLQa

[4457] Massimo Tessarotto and Claudio Cremaschini. The common logic of quantum universe — Part I: The case of non-relativistic quantum mechanics. Foundations of Physics, 52(1):??, February 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-022-00547-z.

Abelson:2022:FTV

[4458] Shannon Sylvie Abelson. The fate of tensor-vector-scalar modified gravity. Foundations of Physics, 52(1):??, February 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-022-00545-1.

Mairhofer:2022:RRB

[4459] Lukas Mairhofer and Oliver Passon. Reconsidering the relation between “matter wave interference” and “wave-particle duality”. Foundations of Physics, 52(2):??, April 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.com/

article/10.1007/s10701-022-00544-2.

Fiorentin:2022:GES

[4460] Stefano Re Fiorentin. Gravity expands space. Foundations of Physics, 52(2):??, April 2022. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-022-00553-1.

Sheykin:2022:UCM

[4461] A. A. Sheykin. Universal constants as manifestations of relativity. Foundations of Physics, 52(2):??, April 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-022-00552-2.

Onoochin:2022:CPK

[4462] Vladimir Onoochin. Comment on the paper of A. K. Singal ‘Contribution of pressure to the energy-momentum density in a moving perfect fluid: A physical perspective’. Foundations of Physics, 52(2):??, April 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-022-00550-4. See [4319].

Glampedakis:2022:MRQ

[4463] Kostas Glampedakis. A Machian reformulation of quantum mechanics. Foundations of Physics, 52(2):??, April 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-022-00551-3.

Muir:2022:EFS

[4464] Heather A. Muir. Emergence of the fused spacetime from a continuum computing construct of reality. Foundations of Physics, 52(2): ??, April 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL https://link.springer.com/article/10.
1007/s10701-022-00549-x.

Shnerb:2022:TIT

[4465] Nadav M. Shnerb. The time–identity tradeoff. Foundations of Physics, 52(2):??, April 2022. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-022-00557-x.

Tessarotto:2022:CLQb

[4466] Massimo Tessarotto and Claudio Cremaschini. The common logic of quantum universe — Part II: The case of quantum gravity. Foundations of Physics, 52(2):??, April 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-022-00548-y.

Carcassi:2022:RPL

[4467] Gabriele Carcassi and Christine A. Aidala. Reverse physics: From laws to physical assumptions. Foundations of Physics, 52(2):??, April 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-022-00555-z.

Niehaus:2022:TAA

[4468] Arend Niehaus. Trying an alternative ansatz to quantum physics. Foundations of Physics, 52(2):??, April 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-022-00556-y.

Corda:2022:EBR

[4469] Christian Corda. On the equivalence between rotation and gravity: “gravitational” and “cosmological” redshifts in the laboratory. Foundations of Physics, 52(2):??, April 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-022-00558-w.

Zafiris:2022:EASb

[4470] Elias Zafiris. The equiareal Archimedean synchronization method of the quantum symplectic phase space: II. circle-valued moment map, integrality, and symplectic Abelian shadows. Foundations of Physics, 52(2):??, April 2022. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-022-00559-9.

Zafiris:2022:EASa

[4471] Elias Zafiris. The equiareal Archimedean synchronization method of the quantum symplectic phase space: I. Spinorial amplitudes, transition probability, and areal measure of time. Foundations of Physics, 52(2):??, April 2022. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-022-00560-2.

Weingarten:2022:MRQ

[4472] Don Weingarten. Macroscopic reality from quantum complexity. Foundations of Physics, 52(2):??, April 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-022-00554-0.

Zhang:2022:PE

[4473] Fan Zhang. Particulate exotica. Foundations of Physics, 52(2):??, April 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-022-00566-w.

Williams:2022:ECC

[4474] Porter Williams. Entanglement, complexity, and causal asymmetry in quantum theories. Foundations of Physics, 52(2):??, April 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-022-00562-0.

Margoni:2022:CTP

[4475] Emilia Margoni. Can there be a process without time? Processualism within timeless physics. Foundations of Physics, 52(2):??, April 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-022-00564-y.

Roselli:2022:PIA

[4476] Carlo Roselli. The physicist inside the ambiguous room: an argument against the need of consciousness in the quantum mechanical measurement process. Foundations of Physics, 52(2):??, April 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-022-00563-z.

Grinwald:2022:SEC

[4477] P. M. Grinwald. Schrödinger’s equation as a consequence of the Central Limit Theorem without assuming prior physical laws. Foundations of Physics, 52(2):??, April 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.com/

article/10.1007/s10701-022-00561-1.

Fiscaletti:2022:BVQ

[4478] Davide Fiscaletti. From Bohm’s vision of quantum processes to quantum field theory … to the transactional approach. Variations on the theme. Foundations of Physics, 52(3):??, June 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-022-00569-7.

Simeonov:2022:MMM

[4479] Lachezar S. Simeonov. Mechanical model of Maxwell’s equations and of Lorentz transformations. Foundations of Physics, 52(3):??, June 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-022-00567-9.

Bricmont:2022:ESP

[4480] Jean Bricmont, Sheldon Goldstein, and Douglas Hemmick. From EPR– Schrödinger Paradox to nonlocality based on perfect correlations. Foundations of Physics, 52(3):??, June 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-022-00568-8.

Kazemi:2022:EUA

[4481] M. J. Kazemi and S. Y. Rokni. Epistemic uncertainty from an averaged Hamilton–Jacobi formalism. Foundations of Physics, 52(3):??, June 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-022-00571-z.

Helland:2022:BEL

[4482] Inge S. Helland. The Bell experiment and the limitations of actors. Foundations of Physics, 52(3):??, June 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-022-00572-y.

Aharonov:2022:CVC

[4483] Yakir Aharonov and Tomer Shushi. Complex-valued classical behavior from the correspondence limit of quantum mechanics with two boundary conditions. Foundations of Physics, 52(3):??, June 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-022-00576-8.

Deledicque:2022:DEE

[4484] Vincent Deledicque. Dark energy explained by a bias in the measurements. Foundations of Physics, 52(3):??, June 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-022-00575-9.

Cabaret:2022:SWF

[4485] D-M. Cabaret, T. Grandou, and E. Perrier. Status of the wave function of quantum mechanics, or, what is quantum mechanics trying to tell us? Foundations of Physics, 52(3):??, June 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-022-00574-w.

Rovelli:2022:PBM

[4486] Carlo Rovelli. Preparation in Bohmian mechanics. Foundations of Physics, 52(3):??, June 2022. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-022-00565-x.

Kumar:2022:TQO

[4487] B. Shravan Kumar and S. Balakrishnan. Two-qubit operators in nosplitting theorems. Foundations of Physics, 52(3):??, June 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-022-00577-7.

Wang:2022:DBE

[4488] Lihong V. Wang. Derivation from Bloch equation to von Neumann equation to Schrödinger–Pauli equation. Foundations of Physics, 52 (3):??, June 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL https://link.springer.com/article/10.
1007/s10701-022-00578-6.

DiBiagio:2022:RQM

[4489] Andrea Di Biagio and Carlo Rovelli. Relational quantum mechanics is about facts, not states: a reply to Pienaar and Brukner. Foundations of Physics, 52(3):??, June 2022. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-022-00579-5.

Farshi:2022:EAT

[4490] Farhang Hadad Farshi and Silvia DeBianchi. An epistemic analysis of time phenomenon. Foundations of Physics, 52(3):??, June 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-022-00583-9.

Lombardi:2022:HDI

[4491] Olimpia Lombardi and Juan Sebastián Ardenghi. How different interpretations of quantum mechanics can enrich each other: The case of the relational quantum mechanics and the modal-Hamiltonian interpretation. Foundations of Physics, 52(3):??, June 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-022-00580-y.

Poletti:2022:SQM

[4492] Marcello Poletti. On the strangeness of quantum mechanics. Foundations of Physics, 52(3):??, June 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-022-00582-w.

Dorato:2022:WOR

[4493] Mauro Dorato and Matteo Morganti. What ontology for relational quantum mechanics? Foundations of Physics, 52(3):??, June 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-022-00581-x.

FagundesdosSantos:2022:ZPV

[4494] Leonardo Sioufi Fagundes dos Santos. Zeno’s paradoxes and the viscous friction force. Foundations of Physics, 52(3):??, June 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-022-00589-3.

Leegwater:2022:WGH

[4495] Gijs Leegwater. When Greenberger, Horne and Zeilinger meet Wigner’s friend. Foundations of Physics, 52(4):??, August 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-022-00586-6.

Gergely:2022:SAD

[4496] Bret Gergely and Herman Batelaan. Simulation of Afshar’s double slit experiment. Foundations of Physics, 52(4):??, August 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-022-00585-7.

Pasqualini:2022:TSB

[4497] Matías Pasqualini and Sebastian Fortin. Trans-statistical behavior of a multiparticle system in an ontology of properties. Foundations of Physics, 52(4):??, August 2022. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-022-00590-w.

Jun:2022:TID

[4498] Hyun-Su Jun. Theoretical investigation of deceleration parameterdependent gravitation in a complex spacetime manifold. Foundations of Physics, 52(4):??, August 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-022-00588-4.

Islami:2022:UEM

[4499] Arezoo Islami. The unreasonable effectiveness of mathematics: From Hamming to Wigner and back again. Foundations of Physics, 52 (4):??, August 2022. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-022-00592-8.

Landsman:2022:BMD

[4500] Klaas Landsman. Bohmian mechanics is not deterministic. Foundations of Physics, 52(4):??, August 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-022-00591-9.

Gogberashvili:2022:TID

[4501] Merab Gogberashvili. Towards an information description of space– time. Foundations of Physics, 52(4):??, August 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-022-00594-6.

Sutherland:2022:PCW

[4502] Roderick I. Sutherland. Probabilities and certainties within a causally symmetric model. Foundations of Physics, 52(4):??, August 2022. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

https://link.springer.com/article/10.1007/s10701-022-00573-
x.

Schneider:2022:SCP

[4503] Mike D. Schneider. A (strictly) contemporary perspective on trans-Planckian censorship. Foundations of Physics, 52(4):??, August 2022.

CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

https://link.springer.com/article/10.1007/s10701-022-00593-
7.

Garrison:2022:QSI

[4504] J. C. Garrison. Quantum statistics of identical particles. Foundations of Physics, 52(4):??, August 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-022-00596-4.

Mousavi:2022:SBS

[4505] S. V. Mousavi and S. Miret-Artés. Stochastic Bohmian and scaled trajectories. Foundations of Physics, 52(4):??, August 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-022-00598-2.

Allori:2022:WIL

[4506] Valia Allori. What is it like to be a relativistic GRW theory? or: Quantum mechanics and relativity, still in conflict after all these years. Foundations of Physics, 52(4):??, August 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-022-00595-5.

Nikolic:2022:RQB

[4507] Hrvoje Nikolić. Relativistic QFT from a Bohmian perspective: a proof of concept. Foundations of Physics, 52(4):??, August 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-022-00600-x.

Hance:2022:SVB

[4508] Jonte R. Hance, Sabine Hossenfelder, and Tim N. Palmer. Supermeasured: Violating Bell-statistical independence without violating physical statistical independence. Foundations of Physics, 52(4):??, August 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-022-00602-9.

Deutsch:2022:SSC

[4509] J. M. Deutsch and Anthony Aguirre. State-to-state cosmology: A new view on the cosmological arrow of time and the past hypothesis. Foundations of Physics, 52(4):??, August 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-022-00597-3.

Matarese:2022:OME

[4510] Vera Matarese. ‘Oumuamua and meta-empirical confirmation. Foundations of Physics, 52(4):??, August 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-022-00587-5.

Cetto:2022:REV

[4511] Ana María Cetto and Luis de la Peña. Role of the electromagnetic vacuum in the transition from classical to quantum mechanics. Foundations of Physics, 52(4):??, August 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-022-00605-6.

Tzemos:2022:BCM

[4512] Athanasios C. Tzemos and George Contopoulos. Bohmian chaos in multinodal bound states. Foundations of Physics, 52(4):??, August 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-022-00599-1.

Silk:2022:SR

[4513] Joseph Silk and Michel Cassé. Swampland revisited. Foundations of Physics, 52(4):??, August 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.com/

article/10.1007/s10701-022-00601-w.

Bolotin:2022:PCR

[4514] Arkady Bolotin. The paradox of classical reasoning. Foundations of Physics, 52(4):??, August 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.com/

article/10.1007/s10701-022-00604-7.

Healey:2022:SOR

[4515] Richard A. Healey. Securing the objectivity of relative facts in the quantum world. Foundations of Physics, 52(4):??, August 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-022-00603-8.

Gao:2022:BET

[4516] Shan Gao. On Bell’s Everett (?) theory. Foundations of Physics, 52(4):??, August 2022. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-022-00609-2.

Gill:2022:DRQ

[4517] Tepper L. Gill, Gonzalo Ares de Parga, and Mamadou Wade. Dual relativistic quantum mechanics I. Foundations of Physics, 52(4):??, August 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-022-00607-4.

Merriam:2022:PFQ

[4518] Paul Merriam. Presentist fragmentalism and quantum mechanics. Foundations of Physics, 52(4):??, August 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-022-00606-5.

Jia:2022:WSO

[4519] Ding Jia. What should be the ontology for the standard model? Foundations of Physics, 52(4):??, August 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-022-00610-9.

Sheikhahmadi:2022:SBH

[4520] Haidar Sheikhahmadi. Schwarzschild black hole perturbed by a force-free magnetic field. Foundations of Physics, 52(4):??, August 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-022-00612-7.

Maltrana:2022:HPT

[4521] Diego Maltrana and Nicolás Sepúlveda-Quiroz. The heuristic power of theory classification, the case of general relativity. Foundations of Physics, 52(4):??, August 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.com/

article/10.1007/s10701-022-00614-5.

Dieks:2022:PQR

[4522] Dennis Dieks. Perspectival quantum realism. Foundations of Physics, 52(4):??, August 2022. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-022-00611-8.

Szabados:2022:NPS

[4523] László B. Szabados. A note on Penrose’s spin-geometry theorem and the geometry of ‘empirical quantum angles’. Foundations of Physics, 52(4):??, August 2022. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-022-00616-3.

Vink:2022:SCE

[4524] Jeroen C. Vink. Spin and contextuality in extended de Broglie– Bohm–Bell quantum mechanics. Foundations of Physics, 52(5):??, October 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-022-00584-8.

Lambare:2022:MLR

[4525] Justo Pastor Lambare. On the meaning of local realism. Foundations of Physics, 52(5):??, October 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-022-00618-1.

Adlam:2022:WCI

[4526] Emily Adlam. Watching the clocks: Interpreting the Page–Wootters formalism and the internal quantum reference frame programme. Foundations of Physics, 52(5):??, October 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-022-00620-7.

Ellerman:2022:FMM

[4527] David Ellerman. Follow the math!: The mathematics of quantum mechanics as the mathematics of set partitions linearized to (Hilbert) vector spaces. Foundations of Physics, 52(5):??, October 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-022-00608-3.

Doner:2022:GEE

[4528] M. Kemal Döner and André Großardt. Is gravitational entanglement evidence for the quantization of spacetime? Foundations of Physics, 52(5):??, October 2022. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-022-00619-0.

Szabados:2022:TSQ

[4529] László B. Szabados. Three-space from quantum mechanics. Foundations of Physics, 52(5):??, October 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-022-00617-2.

Norsen:2022:PWA

[4530] Travis Norsen. A pilot-wave approach to the many-body problem: Beyond the small entanglement approximation. Foundations of Physics, 52(5):??, October 2022. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-022-00621-6.

Alkhateeb:2022:RBT

[4531] M. Alkhateeb and A. Matzkin. Relativistic Bohmian trajectories and Klein–Gordon currents for spin-0 particles. Foundations of Physics, 52(5):??, October 2022. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-022-00625-2.

Rizzuti:2022:TRL

[4532] Bruno F. Rizzuti, Luca M. Gaio, and Lucas T. Cardoso. Is time the real line? Foundations of Physics, 52(5):??, October 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-022-00623-4.

Tokasi:2022:SRC

[4533] Sahand Tokasi and Peter Pickl. Symplectic reduction of classical mechanics on shape space. Foundations of Physics, 52(5):??, October 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-022-00613-6.

Tappenden:2022:PWT

[4534] Paul Tappenden. Pilot-wave theory without nonlocality. Foundations of Physics, 52(5):??, October 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-022-00627-0.

Buchholz:2022:CLP

[4535] Detlev Buchholz, Roberto Longo, and Karl-Henning Rehren. Causal Lie products of free fields and the emergence of quantum field theory. Foundations of Physics, 52(5):??, October 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-022-00629-y.

Riek:2022:LTU

[4536] Roland Riek. Lorentz transformation under a discrete dynamical time and continuous space. Foundations of Physics, 52(5):??, October 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-022-00622-5.

Tumulka:2022:LGM

[4537] Roderich Tumulka. Limitations to genuine measurements in ontological models of quantum mechanics. Foundations of Physics, 52 (5):??, October 2022. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-022-00633-2.

Allori:2022:GIP

[4538] Valia Allori. On the Galilean invariance of the pilot-wave theory. Foundations of Physics, 52(5):??, October 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-022-00631-4.

Lopez:2022:SCT

[4539] Cristian López. Spontaneous collapse theories and temporal primitivism about time’s direction. Foundations of Physics, 52(5):??, October 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-022-00632-3.

Sebens:2022:DRP

[4540] Charles T. Sebens. The disappearance and reappearance of potential energy in classical and quantum electrodynamics. Foundations of Physics, 52(5):??, October 2022. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-022-00630-5.

Gao:2022:UTR

[4541] Shan Gao. Understanding time reversal in quantum mechanics: a new derivation. Foundations of Physics, 52(5):??, October 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-022-00634-1.

Adlam:2022:TEQ

[4542] Emily Adlam. Tabletop experiments for quantum gravity are also tests of the interpretation of quantum mechanics. Foundations of Physics, 52(5):??, October 2022. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-022-00636-z.

Fabbri:2022:BBF

[4543] Luca Fabbri. de Broglie–Bohm formulation of Dirac fields. Foundations of Physics, 52(6):??, December 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-022-00641-2.

Bloch:2022:IMQ

[4544] Yakov Bloch and Eliahu Cohen. The identification of mean quantum potential with Fisher information leads to a strong uncertainty relation. Foundations of Physics, 52(6):??, December 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-022-00638-x.

Dzhunushaliev:2022:QEN

[4545] Vladimir Dzhunushaliev and Vladimir Folomeev. QCD effects in non– QCD theories. Foundations of Physics, 52(6):??, December 2022. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

https://link.springer.com/article/10.1007/s10701-022-00642-
1.

Kent:2022:H

[4546] Adrian Kent. Hodology. Foundations of Physics, 52(6):??, December 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-022-00637-y.

Lazarovici:2022:PPO

[4547] Dustin Lazarovici and Paula Reichert. The point of primitive ontology. Foundations of Physics, 52(6):??, December 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-022-00639-w.

Arroyo:2022:WDR

[4548] Raoni Wohnrath Arroyo and Jonas R. Becker Arenhart. Whence deep realism for Everettian quantum mechanics? Foundations of Physics, 52(6):??, December 2022. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-022-00643-0.

Diaz-Pachon:2023:SSD

[4549] Daniel Andrés Díaz-Pachón, Ola Hössjer, and Robert J. Marks II. Sometimes size does not matter. Foundations of Physics, 53(1):??, February 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-022-00650-1.

Durt:2023:TBD

[4550] T. Durt. Testing de Broglie’s double solution in the mesoscopic regime. Foundations of Physics, 53(1):??, February 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-022-00626-1.

Wolf:2023:EMD

[4551] William J. Wolf, James Read, and Nicholas J. Teh. Edge modes and dressing fields for the Newton–Cartan quantum Hall effect. Foundations of Physics, 53(1):??, February 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-022-00615-4.

DelSanto:2023:OPI

[4552] Flavio Del Santo and Nicolas Gisin. The open past in an indeterministic physics. Foundations of Physics, 53(1):??, February 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-022-00645-y.

Lu:2023:TWP

[4553] Jianbo Lu, Shining Yang, Jing Guo, Yan Liu, Mou Xu, and Jie Wang. Traversable-wormhole physics in GBD theory of modified gravity. Foundations of Physics, 53(1):??, February 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-022-00644-z.

Valentini:2023:BBR

[4554] Antony Valentini. Beyond the Born rule in quantum gravity. Foundations of Physics, 53(1):??, February 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-022-00635-0.

Aristarhov:2023:HUP

[4555] Serj Aristarhov. Heisenberg’s uncertainty principle and particle trajectories. Foundations of Physics, 53(1):??, February 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-022-00646-x.

Bielinska:2023:TSO

[4556] Marta Bielińska and James Read. Testing spacetime orientability. Foundations of Physics, 53(1):??, February 2023. CODEN FNDPA4.

ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-022-00649-8.

Holland:2023:EWQ

[4557] Peter Holland. Eliminating the wavefunction from quantum dynamics: The bi-Hamilton–Jacobi theory, trajectories and time reversal. Foundations of Physics, 53(1):??, February 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-022-00640-3.

Oldofredi:2023:RDQ

[4558] Andrea Oldofredi. The relational dissolution of the quantum measurement problems. Foundations of Physics, 53(1):??, February 2023. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

https://link.springer.com/article/10.1007/s10701-022-00652-
z.

Gao:2023:CPQ

[4559] Shan Gao. Can pragmatist quantum realism explain protective measurements? Foundations of Physics, 53(1):??, February 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-022-00654-x.

Helland:2023:SQM

[4560] Inge S. Helland. A simple quantum model linked to decisions. Foundations of Physics, 53(1):??, February 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-022-00658-7.

Durt:2023:DIE

[4561] T. Durt. Do(es the influence of) empty waves survive in configuration space? Foundations of Physics, 53(1):??, February 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-022-00624-3.

Croca:2023:PEC

[4562] J. R. Croca, P. Castro, M. Gatta, and R. N. Moreira. Proposed experiments to clarify the real nature of the quantum waves. Foundations of Physics, 53(1):??, February 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-022-00656-9.

Bengtsson:2023:ENG

[4563] Ingemar Bengtsson and Tobias Eklund. Energy in Newtonian gravity. Foundations of Physics, 53(1):??, February 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-022-00660-z.

Hubert:2023:SIQ

[4564] Mario Hubert. Is the statistical interpretation of quantum mechanics ψ-ontic or ψ-epistemic? Foundations of Physics, 53(1):??, February 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-022-00651-0.

Nikolaev:2023:ASM

[4565] Vitaly Nikolaev and Louis Vervoort. Aspects of superdeterminism made intuitive. Foundations of Physics, 53(1):??, February 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-022-00648-9.

Matarese:2023:BBT

[4566] Vera Matarese. De Broglie–Bohm theory, Quo Vadis? Foundations of Physics, 53(1):??, February 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-022-00647-w.

Englman:2023:LIA

[4567] R. Englman and A. Yahalom. Lindbladian-induced alignment in quantum measurements. Foundations of Physics, 53(1):??, February 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-022-00659-6.

Dagan:2023:RHI

[4568] Yuval Dagan. Relativistic hydrodynamic interpretation of de Broglie matter waves. Foundations of Physics, 53(1):??, February 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-022-00657-8.

Saxena:2023:CVC

[4569] Ashwin Saxena and Anirban Pathak. Continuous variable controlled quantum conference. Foundations of Physics, 53(1):??, February 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-022-00661-y.

Luc:2023:SVC

[4570] Joanna Luc. States vs. changes of states: a reformulation of the ontic vs. epistemic distinction in quantum mechanics. Foundations of Physics, 53(1):??, February 2023. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-022-00662-x.

Wharton:2023:EPI

[4571] Ken Wharton and Raylor Liu. Entanglement and the path integral. Foundations of Physics, 53(1):??, February 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-022-00664-9.

Dewdney:2023:RBB

[4572] Christopher Dewdney. Rekindling of de Broglie–Bohm pilot wave theory in the Late Twentieth Century: a personal account. Foundations of Physics, 53(1):??, February 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-022-00655-w.

Gutierrez:2023:UCM

[4573] S. Gutiérrez, A. Camacho, and H. Ríos. Ultra-cold many-body systems and phenomenology of gravity theories with compact dimensions. Foundations of Physics, 53(1):??, February 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-022-00665-8.

Martin-Dussaud:2023:FNT

[4574] Pierre Martin-Dussaud, Titouan Carette, Jan Głowacki, Vaclav Zatloukal, and Federico Zalamea. Fact-nets: Towards a mathematical framework for relational quantum mechanics. Foundations of Physics, 53 (1):??, February 2023. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-022-00653-y.

Vink:2023:QES

[4575] Jeroen C. Vink. Quantum equilibrium in stochastic de Broglie–Bohm– Bell quantum mechanics. Foundations of Physics, 53(1):??, February 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-022-00668-5.

Verrall:2023:GSQ

[4576] Steven C. Verrall, Micah Atkins, Andrew Kaminsky, Emily Friederick, Andrew Otto, Kelly S. Verrall, and Peter Lynch. Ground state quantum vortex proton model. Foundations of Physics, 53(1):??, February 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-023-00669-y.

Ntelis:2023:FA

[4577] Pierros Ntelis and Adam Morris. Functors of actions. Foundations of Physics, 53(1):??, February 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-022-00628-z.

Drezet:2023:CBR

[4578] Aurélien Drezet. Can a Bohmian be a Rovellian for all practical purposes? Foundations of Physics, 53(1):??, February 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-023-00670-5.

Drezet:2023:QSN

[4579] Aurélien Drezet. Quantum solitodynamics: Non-linear wave mechanics and pilot-wave theory. Foundations of Physics, 53(1):??, February 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-023-00671-4.

Ali:2023:QUD

[4580] Md. Manirul Ali. Quantum uncertainty dynamics. Foundations of Physics, 53(1):??, February 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.com/

article/10.1007/s10701-023-00672-3.

Shaghaghi:2023:PFQ

[4581] Mehran Shaghaghi. The physical foundation of quantum theory. Foundations of Physics, 53(1):??, February 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-023-00673-2.

Snoke:2023:MFN

[4582] D. W. Snoke. Mathematical formalism for nonlocal spontaneous collapse in quantum field theory. Foundations of Physics, 53(2):??, April 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-023-00674-1.

Heifetz:2023:BNM

[4583] Eyal Heifetz, Anirban Guha, and Leo Maas. de Broglie normal modes in the Madelung fluid. Foundations of Physics, 53(2):??, April 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-023-00676-z.

Karlsson:2023:BSE

[4584] Anna Karlsson. Building spacetime from effective interactions between quantum fluctuations. Foundations of Physics, 53(2):??, April 2023. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

https://link.springer.com/article/10.1007/s10701-022-00667-
6.

Muslih:2023:PEW

[4585] Sami I. Muslih. Propagation of electromagnetic waves in fractional space time dimensions. Foundations of Physics, 53(2):??, April 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-023-00677-y.

Qureshi:2023:UMT

[4586] Tabish Qureshi. Understanding modified two-slit experiments using path markers. Foundations of Physics, 53(2):??, April 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-023-00684-z.

Toader:2023:ECC

[4587] Iulian D. Toader. Einstein completeness as categoricity. Foundations of Physics, 53(2):??, April 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.com/

article/10.1007/s10701-023-00675-0.

Lee:2023:VVS

[4588] Seokcheon Lee. A viable varying speed of light model in the RW metric. Foundations of Physics, 53(2):??, April 2023. CODEN FNDPA4.

ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-023-00682-1.

Romano:2023:DBA

[4589] Davide Romano. A decoherence-based approach to the classical limit in Bohm’s theory. Foundations of Physics, 53(2):??, April 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-023-00679-w.

Husain:2023:ICC

[4590] Viqar Husain and Vladimir Tasić. Indeterminacy in classical cosmology with dark matter. Foundations of Physics, 53(2):??, April 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-023-00683-0.

deGosson:2023:PSC

[4591] Maurice de Gosson and Charlyne de Gosson. Pointillisme à la Signac and construction of a quantum fiber bundle over convex bodies. Foundations of Physics, 53(2):??, April 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-023-00681-2.

Adlam:2023:DWV

[4592] Emily Adlam. Do we have any viable solution to the measurement problem? Foundations of Physics, 53(2):??, April 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-023-00686-x.

Bossche:2023:CUC

[4593] Mathias Van Den Bossche and Philippe Grangier. Contextual unification of classical and quantum physics. Foundations of Physics, 53(2):??, April 2023. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-023-00678-x.

Bloch:2023:SSR

[4594] Yakov Bloch. Spacetime superoscillations and the relativistic quantum potential. Foundations of Physics, 53(2):??, April 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-023-00680-3.

Sarkar:2023:ONC

[4595] Shubhayan Sarkar. An operational notion of classicality based on physical principles. Foundations of Physics, 53(2):??, April 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-023-00687-w.

Rodriguez-Warnier:2023:TGC

[4596] Pascal Rodríguez-Warnier. Temporal global correlations in timesymmetric collapse models. Foundations of Physics, 53(3):??, June 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

https://link.springer.com/article/10.1007/s10701-023-00688-
9.

Sontz:2023:NOQ

[4597] Stephen Bruce Sontz. A new organization of quantum theory based on quantum probability. Foundations of Physics, 53(3):??, June 2023. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

https://link.springer.com/article/10.1007/s10701-023-00691-
0.

Dvorak:2023:ABR

[4598] Marc Dvorak. An alternative to the Born rule: Spectral quantization. Foundations of Physics, 53(3):??, June 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-023-00692-z.

Nikulov:2023:PTG

[4599] Alexey Nikulov. Physical thinking and the GHZ theorem. Foundations of Physics, 53(3):??, June 2023. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-023-00693-y.

Torrome:2023:EOI

[4600] Ricardo Gallego Torromé, J. M. Isidro, and Pedro Fernández de Córdoba. On the emergent origin of the inertial mass. Foundations of Physics, 53(3):??, June 2023. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-023-00689-8.

Good:2023:IAR

[4601] Michael R. R. Good and Paul C. W. Davies. Infrared acceleration radiation. Foundations of Physics, 53(3):??, June 2023. CODEN FNDPA4.

ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-023-00694-x.

Fabbri:2023:DTH

[4602] Luca Fabbri. Dirac theory in hydrodynamic form. Foundations of Physics, 53(3):??, June 2023. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-023-00695-w.

Shanahan:2023:LTF

[4603] Daniel Shanahan. The Lorentz transformation in a fishbowl: A comment on Cheng and Read’s “Why Not a Sound Postulate?”. Foundations of Physics, 53(3):??, June 2023. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-023-00698-7.

Boykin:2023:PMC

[4604] Timothy B. Boykin. The position–momentum commutator as a generalized function: Resolution of the apparent discrepancy between continuous and discrete bases. Foundations of Physics, 53(3):??, June 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-023-00697-8.

Beck:2023:NCR

[4605] James L. Beck. Neo-classical relativistic mechanics theory for electrons that exhibits spin, Zitterbewegung, dipole moments, wavefunctions and Dirac’s wave equation. Foundations of Physics, 53(3):??, June 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-023-00696-9.

Feldmann:2023:ITI

[4606] Michel Feldmann. Information-theoretic interpretation of quantum formalism. Foundations of Physics, 53(3):??, June 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-023-00690-1.

Sorella:2023:RBO

[4607] S. P. Sorella. On the representations of Bell’s operators in quantum mechanics. Foundations of Physics, 53(3):??, June 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-023-00699-6.

Jovanovic:2023:AGR

[4608] Bozidar Jovanović. Affine geometry and relativity. Foundations of Physics, 53(3):??, June 2023. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-023-00700-2.

Cevik:2023:CFW

[4609] Ahmet Çevik and Zeki Seskir. On the cardinality of future worldlines in discrete spacetime structures. Foundations of Physics, 53(3): ??, June 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-023-00701-1.

Drezet:2023:FSI

[4610] Aurélien Drezet. Forewords for the special issue ‘Pilot-wave and Beyond: Louis de Broglie and David Bohm’s Quest for a Quantum Ontology’. Foundations of Physics, 53(3):??, June 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-023-00685-y.

Bernardini:2023:QPP

[4611] A. E. Bernardini and O. Bertolami. Quantum prey–predator dynamics: a Gaussian ensemble analysis. Foundations of Physics, 53(3):??, June 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-023-00703-z.

Cabbolet:2023:NPT

[4612] Marcoen J. T. F. Cabbolet. On the nonreality of the PBR Theorem. Foundations of Physics, 53(3):??, June 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-023-00704-y.

Born:1959:POE

[4613] Max Born and Emil Wolf. Principles of optics: electromagnetic theory of propagation, interference, and diffraction of light. Pergamon, New York, NY, USA, 1959. 803 pp. LCCN QC355 .B63.

Jackson:1962:CE

[4614] John David Jackson. Classical Electrodynamics. John Wiley, New York, NY, USA, 1962. 641 pp. LCCN QC631 .J3.

Jackson:1975:CE

[4615] John David Jackson. Classical Electrodynamics. John Wiley, New York, NY, USA, second edition, 1975. ISBN 0-471-43132-X. xxii + 848 pp. LCCN QC631 .J3 1975. URL http://www.

loc.gov/catdir/enhancements/fy0607/75009962-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0607/75009962-t.html.

Griffiths:1995:IQM

[4616] David J. (David Jeffery) Griffiths. Introduction to Quantum Mechanics. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1995. ISBN 0-13-124405-1. ix + 394 pp. LCCN QC174.12 .G75 1995. US$40.00.

Post:1995:QRE

[4617] Evert Jan Post. Quantum Reprogramming: Ensembles and Single Systems: a Two-Tier Approach to Quantum Mechanics, volume 181 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1995. ISBN 0-7923-3565-1 (hardcover), 94-015-8410-9. ISSN 0068-0346. xiv + 320 pp. LCCN Q174 .B67 vol. 181; Q174; Q174.12; QC174.12 P675 1995. URL http://books.google.com/

books?id=lK3aAAAAMAAJ;   http://catalog.hathitrust.org/api/
volumes/oclc/32469387.html;   http://catdir.loc.gov/catdir/
enhancements/fy0823/95017463d.html;   http://catdir.loc.gov/
catdir/enhancements/fy0823/95017463t.html;   https://link.springer.
com/book/10.1007/978-94-015-8410-4.

Fujita:1996:QST

[4618] Shigeji Fujita and Salvador Godoy. Quantum statistical theory of superconductivity. Selected topics in superconductivity; The language of science. Plenum Press, New York, NY, USA; London, UK, 1996. ISBN 0-306-45363-0. xix + 338 pp. LCCN QC611.92 .F84 1996. URL http:

//www.loc.gov/catdir/enhancements/fy0820/96022326-d.html;
http://www.loc.gov/catdir/enhancements/fy0820/96022326-t.html.

Bub:1997:IQW

[4619] Jeffrey Bub. Interpreting the quantum world. Cambridge University Press, Cambridge, UK, 1997. ISBN 0-521-56082-9 (hardback). xiv + 298 pp. LCCN QC174.12 .B83 1997. URL

http://www.loc.gov/catdir/description/cam028/96031563.html;
 http://www.loc.gov/catdir/samples/cam031/96031563.html;   http:
//www.loc.gov/catdir/toc/cam021/96031563.html.

Esposito:1997:EQG

[4620] Giampiero Esposito, Alexander Yu Kamenshchik, and Giuseppe Pollifrone. Euclidean quantum gravity on manifolds with boundary, volume 85 of Fundamental theories of physics. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1997. ISBN 0-7923-4472-3. xiii + 319 pp. LCCN QC178 .E85 1997.

Home:1997:CFQ

[4621] Dipankar Home. Conceptual foundations of quantum physics: an overview from modern perspectives. The language of science. Plenum Press, New York, NY, USA; London, UK, 1997. ISBN 0-306-45660-5 (hardcover). xvii + 386 pp. LCCN QC174.12; QC174.12 .H65

1997. URL http://catdir.loc.gov/catdir/enhancements/fy0813/
97031703-d.html;   http://catdir.loc.gov/catdir/enhancements/
fy0813/97031703-t.html;   http://www.gbv.de/dms/hebisdarmstadt/
toc/59664223.pdf;   http://www.zentralblattmath.org/zmath/en/
search/?an=0972.81001.

Namiki:1997:DQM

[4622] Mikio Namiki, Saverio Pascazio, and Hiromichi Nakazato. Decoherence and quantum measurements. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1997. ISBN 981-02-3077-X. ix + 228 pp. LCCN QC174.12 .N36 1997.

Pais:1997:TTC

[4623] Abraham Pais. A tale of two continents: a physicist’s life in a turbulent world. Princeton University Press, Princeton, NJ, USA, 1997. ISBN 0-691-01243-1. xvi + 511 pp. LCCN QC16.P24 A3 1997.

URL http://www.loc.gov/catdir/description/prin031/96039313.
html;   http://www.loc.gov/catdir/toc/prin031/96039313.html.

Dickson:1998:QCN

[4624] William Michael Dickson. Quantum chance and non-locality: probability and non-locality in the interpretations of quantum mechanics. Cambridge University Press, Cambridge, UK, 1998. ISBN 0-521-58127-3. xix + 244 pp. LCCN QC174.12 .D53 1998. URL

http://www.loc.gov/catdir/description/cam028/97008813.html;
 http://www.loc.gov/catdir/toc/cam021/97008813.html.

Pickover:1998:TTG

[4625] Clifford A. Pickover. Time: a traveler’s guide. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1998. ISBN 0-19-512042-6, 0-19-513096-0. xviii + 285 pp. LCCN ????

Redei:1998:QLA

[4626] Miklós Rédei. Quantum logic in algebraic approach, volume 91 of Fundamental theories of physics. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1998. ISBN 0-7923-4903-2 (hardcover). x + 238 pp. LCCN QC174.17.M35 R44 1998. URL http://www.

loc.gov/catdir/enhancements/fy0813/97046597-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0813/97046597-t.html.

Rothman:1998:DCC

[4627] Tony Rothman and E. C. G. Sudarshan. Doubt and certainty: the celebrated academy: debates on science, mysticism, reality, in general on the knowable and unknowable, with particular forays into such esoteric matters as the mind fluid, the behavior of the stock market, and the disposition of a quantum mechanical sphinx, to name a few. Helix books. Perseus Publishers, Cambridge, MA, USA, 1998. ISBN 0-7382-0006-9. xvi + 320 pp. LCCN Q175 .R56494 1998. US$25.00. URL http://www.

loc.gov/catdir/enhancements/fy0830/98087634-d.html.

Roy:1998:SGA

[4628] Sisir Roy. Statistical geometry and applications to microphysics and cosmology, volume 92 of Fundamental theories of physics. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1998. ISBN 0-7923-4907-5. xi + 246 pp. LCCN QC173.4.M5 R69 1998.

URL http://www.loc.gov/catdir/enhancements/fy0822/97050114-
d.html;   http://www.loc.gov/catdir/enhancements/fy0822/97050114-
t.html.

Assis:1999:RME

[4629] André Koch Torres Assis. Relational Mechanics. Apeiron, Montréal, Québec, Canada, 1999. ISBN 0-9683689-2-1. 285 pp. LCCN QC125.2

A88 1999. US$20.00. URL http://www.gbv.de/dms/bowker/toc/
9780968368923.pdf.

Auletta:1999:FIQ

[4630] Gennaro Auletta. Foundations and interpretation of quantum mechanics: in the lights of a critical-historical analysis of the problems and of a synthesis of the results. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1999. ISBN 981-02-4039-

2. xxxii + 981 pp. LCCN ???? URL http://www.zentralblatt-math.
org/zmath/en/search/?an=0965.81001.

Jackson:1999:CE

[4631] John David Jackson. Classical Electrodynamics. John Wiley, New York, NY, USA, third edition, 1999. ISBN 0-471-30932-X (hardcover). xxi + 808 pp. LCCN QC631 .J3 1999. URL http://www.

loc.gov/catdir/description/wiley033/97046873.html;   http:/
/www.loc.gov/catdir/toc/onix04/97046873.html.

Pickover:1999:STH

[4632] Clifford A. Pickover. Surfing through hyperspace: understanding higher universes in six easy lessons. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1999. ISBN 0-19-513006-5, 0-19-535155-X, 1-280-47238-3, 1-4237-3671-0 (e-book), 1-60256-371-3 (ebook). xxiv + 239 pp. LCCN QB981 .P625 1999. URL http:/

/catdir.loc.gov/catdir/enhancements/fy0604/98048660-d.html;
 http://catdir.loc.gov/catdir/enhancements/fy0723/98048660-
b.html;   http://site.ebrary.com/id/10085395;   http://www.
loc.gov/catdir/enhancements/fy0604/98048660-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0723/98048660-b.html.

Pickover:1999:TTG

[4633] Clifford A. Pickover. Time: a traveler’s guide. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1999. ISBN 0-19-513096-0 (paperback), 0-19-512042-6. xviii + 285 pp. LCCN QC173.59.S65 P53 1999.

Stockmann:1999:QCI

[4634] Hans-Jürgen Stöckmann. Quantum Chaos: an Introduction, volume 3 of Cambridge nonlinear science series. Cambridge University Press, Cambridge, UK, 1999. ISBN 0-521-59284-4 (hardcover). xi + 368 pp. LCCN QC174.17.C45 S84 1999. URL http:

//catdir.loc.gov/catdir/description/cam029/98045454.html;
http://catdir.loc.gov/catdir/toc/cam021/98045454.html;   http:
//site.ebrary.com/id/10447340;   http://www.loc.gov/catdir/
description/cam029/98045454.html;   http://www.loc.gov/catdir/
toc/cam021/98045454.html.

Torretti:1999:PP

[4635] Roberto Torretti. The philosophy of physics. The evolution of modern philosophy. Cambridge University Press, Cambridge, UK, 1999. ISBN 0-521-56259-7 (hardcover). xvi + 512 pp. LCCN

QC6 .T656 1999. URL http://www.loc.gov/catdir/description/
cam0210/99042504.html;   http://www.loc.gov/catdir/samples/
cam032/99042504.html;   http://www.loc.gov/catdir/toc/cam025/
99042504.html.

Trump:1999:CRM

[4636] Matthew A. Trump and William Carl Schieve. Classical relativistic many-body dynamics, volume 103 of Fundamental theories of physics. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1999. ISBN 0-7923-5737-X (hardcover). xv + 365 pp. LCCN QC174.45 .T7 1999. URL http:/

/catdir.loc.gov/catdir/enhancements/fy0814/99025676-d.html;
 http://catdir.loc.gov/catdir/enhancements/fy0814/99025676-
t.html;   http://www.loc.gov/catdir/enhancements/fy0814/99025676-
d.html;   http://www.loc.gov/catdir/enhancements/fy0814/99025676-
t.html.

Walker:1999:PCQ

[4637] Evan Harris Walker. The physics of consciousness: quantum minds and the meaning of life. Perseus Publishers, Cambridge, MA, USA, 1999. ISBN 0-7382-0234-7. ix + 368 pp. LCCN ????

Auletta:2000:FIQ

[4638] Gennaro Auletta. Foundations and interpretation of quantum mechanics: in the light of a critical-historical analysis of the problems and of a synthesis of the results. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2000. ISBN 981-02-4039-

2. xxxii + 981 pp. LCCN QC174.12 .A94 2000. URL http://www.

zentralblattmath.org/zmath/en/search/?an=0965.81001. With a

foreword by Giorgio Parisi.

Dvurecenskij:2000:NTQ

[4639] Anatolij Dvurečenskij and Sylvia Pulmannová. New trends in quantum structures, volume 516 of Mathematics and its applications. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 0-7923-6471-6. xvi + 541 pp. LCCN QC174.13 .D88 2000. URL http://www.

loc.gov/catdir/enhancements/fy0814/00042414-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0814/00042414-t.html.

Mehra:2000:CMS

[4640] Jagdish Mehra and Kimball A. Milton. Climbing the mountain: the scientific biography of Julian Schwinger. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2000. ISBN 0-19-850658-9 (hardcover), 0-19-852745-4 (paperback). xii + 677 pp. LCCN QC16.S29 M45 2000.

URL http://www.loc.gov/catdir/enhancements/fy0606/99087323-
d.html;   http://www.loc.gov/catdir/enhancements/fy0606/99087323-
t.html.

Stockmann:2000:QCI

[4641] Hans-Jürgen Stöckmann. Quantum chaos: an introduction. Cambridge University Press, Cambridge, UK, reprint edition, 2000. ISBN 0-521-59284-4, 0-521-02715-2. xi + 368 pp. LCCN QC174.17 C45S84 1999.

URL http://www.gbv.de/dms/bowker/toc/9780521592840.pdf;
http://www.zentralblatt-math.org/zmath/en/search/?an=0940.81019.

Walker:2000:PCQ

[4642] Evan Harris Walker. The physics of consciousness: the quantum minds and the meaning of life. Perseus Publishers, Cambridge, MA, USA, 2000. ISBN 0-7382-0234-7 (hardcover), 0-7382-0436-6 (paperback). ix + 368 pp. LCCN BF311 .W266 2000; QP411 .W35 2000. URL http://catdir.

loc.gov/catdir/enhancements/fy0832/99068944-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0832/99068944-d.html.

Auletta:2001:FIQ

[4643] Gennaro Auletta. Foundations and interpretation of quantum mechanics: in the light of a critical-historical analysis of the problems and of a synthesis of the results. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, revised edition, 2001. ISBN 981-02-4039-2, 981-02-4614-5 (paperback). xxxii + 981 pp. LCCN QC174.12 .A85 2001. With a foreword by Giorgio Parisi.

Barbour:2001:DD

[4644] Julian B. Barbour. The discovery of dynamics: a study from a Machian point of view of the discovery and the structure of dynamical theories. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2001. ISBN 0-19-535110-X (e-book), 0-19-513202-5 (paperback). xxiv + 746 pp. LCCN QC133 .B36 2001. URL http:/

/catdir.loc.gov/catdir/enhancements/fy0610/00062340-d.html;
 http://catdir.loc.gov/catdir/enhancements/fy0610/00062340-
t.html;   http://site.ebrary.com/id/10269116;   http://www.
loc.gov/catdir/enhancements/fy0610/00062340-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0610/00062340-t.html.

Fujita:2001:THT

[4645] Shigeji Fujita and Salvador Godoy. Theory of High Temperature Superconductivity, volume 121 of Fundamental theories of physics. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2001. ISBN 1-4020-0149-5 (paperback), 0-306-48216-9 (e-book). viii + 367 pp. LCCN QC611.98.H54 F84 2001.

URL http://ebooks.springerlink.com/summary.asp?id=99510;

http://site.ebrary.com/id/10067344;   http://www.gbv.de/dms/
bowker/toc/9781402001499.pdf;   http://www.loc.gov/catdir/enhancements/
fy0822/2001052594-d.html;   http://www.loc.gov/catdir/enhancements/
fy0822/2001052594-t.html;   http://www.zentralblattmath.org/
zmath/en/search/?an=1052.82034.

Leader:2001:SPP

[4646] Elliot Leader. Spin in particle physics, volume 15 of Cambridge monographs on particle physics, nuclear physics, and cosmology. Cambridge University Press, Cambridge, UK, 2001. ISBN 0-511-52445-5 (e-book), 0-521-35281-9 (hardback), 0-521-02077-8 (paperback). xxi + 500 pp. LCCN QC793.3.S6 L43 2001. URL

http://www.loc.gov/catdir/description/cam021/00058523.html;
 http://www.loc.gov/catdir/enhancements/fy0731/00058523-b.html;
 http://www.loc.gov/catdir/toc/cam022/00058523.html.

Pickover:2001:STH

[4647] Clifford A. Pickover. Surfing through hyperspace: understanding higher universes in six easy lessons. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2001. ISBN 0-19-514241-1 (paperback), 0-19-513006-5. xxiv + 239 pp. LCCN ????

Rindler:2001:RSG

[4648] Wolfgang Rindler. Relativity: special, general, and cosmological. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2001. ISBN 0-19-850835-2 (cased), 0-19-850836-0 (paperback). xiii + 428 pp. LCCN QC173.55 .R563 2001. URL http:/

/catdir.loc.gov/catdir/enhancements/fy0610/00067605-d.html;
 http://catdir.loc.gov/catdir/enhancements/fy0610/00067605-
t.html;   http://www.gbv.de/dms/goettingen/323206352.pdf;
http://www.loc.gov/catdir/enhancements/fy0610/00067605-d.html;
 http://www.loc.gov/catdir/enhancements/fy0610/00067605-t.html.

Santilli:2001:FHC

[4649] Ruggero Maria Santilli. Foundations of hadronic chemistry: with applications to the new clean energies and fuels, volume 120 of Fundamental theories of physics. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2001. ISBN 1-4020-0087-1. ISSN 0168-1222. liv + 397 pp. LCCN QC793.5.H32 S36 2001. URL http://catdir.loc.

gov/catdir/enhancements/fy0813/2001050388-d.html;   http://
catdir.loc.gov/catdir/enhancements/fy0813/2001050388-t.html;
 http://www.gbv.de/dms/bowker/toc/9781402000874.pdf;   http:
//www.zentralblattmath.org/zmath/en/search/?an=1025.81051.

Sussman:2001:SIC

[4650] Gerald Jay Sussman, Jack Wisdom, and Meinhard E. (Meinhard Edwin) Mayer. Structure and interpretation of classical mechanics. MIT Press, Cambridge, MA, USA, 2001. ISBN 0-262-19455-4. xxi + 534 pp. LCCN QC125.2 .S895 2001.

Breuer:2002:TOQ

[4651] Heinz-Peter Breuer and F. (Francesco) Petruccione. The theory of open quantum systems. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2002. ISBN 0-19-852063-8 (hardcover). xxi + 625 pp.

LCCN QC174.85.O6 B74 2002. URL http://catdir.loc.gov/catdir/
enhancements/fy0613/2002075713-d.html;   http://catdir.loc.
gov/catdir/enhancements/fy0613/2002075713-t.html;   http://
www.loc.gov/catdir/enhancements/fy0613/2002075713-d.html;
http://www.loc.gov/catdir/enhancements/fy0613/2002075713-t.
html.

deMuynck:2002:FQM

[4652] Willem M. de Muynck. Foundations of quantum mechanics: an empiricist approach, volume 127 of Fundamental theories of physics. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2002. ISBN 1-4020-0932-1 (hardcover), 0-306-48047-6 (e-book). xxiv + 680 pp. LCCN QC174.12 .M89 2002. URL http://catdir.loc.

gov/catdir/enhancements/fy0822/2002034361-d.html;   http://
catdir.loc.gov/catdir/enhancements/fy0822/2002034361-t.html;
 http://ebooks.springerlink.com/summary.asp?id=99349;   http:
//site.ebrary.com/id/10067332;   http://www.loc.gov/catdir/
enhancements/fy0822/2002034361-d.html;   http://www.loc.gov/
catdir/enhancements/fy0822/2002034361-t.html.

Kiechle:2002:TL

[4653] Hubert Kiechle. Theory of K-loops, volume 1778 of Lecture notes in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 3-540-43262-0 (paperback), 3-540-45817-4. ISSN 0075-8434. x + 186 pp. LCCN QA3

L4 V.1778. URL http://catdir.loc.gov/catdir/enhancements/
fy0817/2002022101-d.html;   http://catdir.loc.gov/catdir/enhancements/
fy0817/2002022101-t.html;   http://www.springerlink.com/openurl.
asp?genre=book\%26isbn=978-3-540-43262-3;   http://www.springerlink.
com/openurl.asp?genre=issue\%26issn=0075-8434\%26volume=1778.

Mugur-Schachter:2002:QMM

[4654] Mioara Mugur-Schächter and Alwyn Van der Merwe. Quantum mechanics, mathematics, cognition, and action: proposals for a formalized epistemology, volume 129 of Fundamental theories of physics. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2002. ISBN 0-306-48144-8 (e-book), 1-4020-1120-2 (paperback). xviii + 493 pp. LCCN QC6 .Q36 2002. URL http://catdir.loc.

gov/catdir/enhancements/fy0822/2002044870-d.html;   http://
catdir.loc.gov/catdir/enhancements/fy0822/2002044870d.html;
 http://catdir.loc.gov/catdir/toc/fy037/2002044870.html;   http:
//ebooks.springerlink.com/summary.asp?id=99443;   http://site.
ebrary.com/id/10067365;   http://www.gbv.de/dms/bowker/toc/
9781402011207.pdf;   http://www.loc.gov/catdir/enhancements/
fy0822/2002044870-d.html;   http://www.loc.gov/catdir/toc/fy037/
2002044870.html.

Tanaka:2002:MSP

[4655] Tomoyasu Tanaka. Methods of Statistical Physics. Cambridge University Press, Cambridge, UK, 2002. ISBN 0-511-07748-3 (e-book), 0-511-75567-8 (e-book), 0-521-58958-4 (paperback), 0-521-58056-0, 0-511-15693-6, 0-511-07591-X. xv + 293 pp. LCCN QC174.8 .T36 2002. URL http://

catdir.loc.gov/catdir/description/cam021/2001035650.html;
http://catdir.loc.gov/catdir/samples/cam033/2001035650.html;
 http://catdir.loc.gov/catdir/toc/cam021/2001035650.html;
http://site.ebrary.com/id/10005055;   http://www.loc.gov/catdir/
description/cam021/2001035650.html;   http://www.loc.gov/catdir/
samples/cam033/2001035650.html;   http://www.loc.gov/catdir/
toc/cam021/2001035650.html.

Auletta:2003:FIQ

[4656] Gennaro Auletta. Foundations and interpretation of quantum mechanics: in the light of a critical-historical analysis of the problems and of a synthesis of the results. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, revised edition, 2003. ISBN 981-02-4039-2, 981-02-4614-5. xxxii + 981 pp. LCCN QC174.12 .A85 2001.

Breuer:2003:TOQ

[4657] Heinz-Peter Breuer and Francesco Petruccione. The theory of open quantum systems. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2003. ISBN 0-19-852063-8. xxi + 625 pp. LCCN QC174.85.O6.

Castell:2003:TQI

[4658] Lutz Castell and Otfried Ischebeck. Time, quantum, and information. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. ISBN 3-540-44033-X. xiii + 456 pp. LCCN QC780 .T56 2003. URL http://www.loc.gov/

catdir/enhancements/fy0817/2003053941-d.html;   http://www.
loc.gov/catdir/enhancements/fy0817/2003053941-t.html.

Dodelson:2003:MCA

[4659] Scott Dodelson. Modern Cosmology: anisotropies and Inhomogeneities in the Universe. Academic Press, New York, USA, 2003. ISBN 0-12-219141-2 (hardcover), 0-08-051197-X (e-book), 1-281-03624-2. xiii + 440 + 4 pp. LCCN QB981 .D63 2003. URL http://catdir.

loc.gov/catdir/description/els051/2002117793.html;   http:/
/catdir.loc.gov/catdir/toc/els051/2002117793.html;   http:/
/site.ebrary.com/id/10190024;   http://www.gbv.de/dms/bowker/
toc/9780122191411.pdf;   http://www.loc.gov/catdir/description/
els051/2002117793.html;   http://www.loc.gov/catdir/toc/els051/
2002117793.html;   http://www.sciencedirect.com/science/book/
9780122191411.

Doran:2003:GAP

[4660] Chris (Chris J. L.) Doran and A. N. (Anthony N.) Lasenby. Geometric Algebra for Physicists. Cambridge University Press, Cambridge, UK, 2003. ISBN 0-521-48022-1. xiv + 578 pp. LCCN QC20.7.C55

D67 2003. URL http://www.loc.gov/catdir/description/cam031/
2002035182.html;   http://www.loc.gov/catdir/toc/cam031/2002035182.
html.

Jaynes:2003:PTL

[4661] E. T. (Edwin T.) Jaynes and G. Larry Bretthorst. Probability theory: the logic of science. Cambridge University Press, Cambridge, UK, 2003. ISBN 0-521-59271-2. xxix + 727 pp. LCCN QA273

.J36 2003eb.  URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/;   http://www.loc.gov/catdir/description/cam031/
2002071486.html;   http://www.loc.gov/catdir/samples/cam033/
2002071486.html;   http://www.loc.gov/catdir/toc/cam031/2002071486.
html.

Mugur-Schachter:2003:QMM

[4662] Mioara Mugur-Schächter and Alwyn Van der Merwe, editors. Quantum mechanics, mathematics, cognition and action: proposals for a formalized epistemology, volume 129 of Fundamental theories of physics. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2003. ISBN 1-4020-1120-2. xviii + 493 pp. LCCN ????

Rindler:2003:RSG

[4663] Wolfgang Rindler. Relativity: special, general, and cosmological. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2003. ISBN 0-19-850836-0 (paperback), 0-19-850835-2 (hardcover). xiii + 428 pp. LCCN ????

Weber:2003:EMM

[4664] Hans J. (Hans Jürgen) Weber and George B. (George Brown) Arfken. Essential Mathematical Methods for Physicists. Elsevier Academic Press, Amsterdam, The Netherlands, 2003. ISBN 0-12-059877-9 (hardcover). xxii + 932 pp. LCCN ????

Zee:2003:QFT

[4665] A. Zee. Quantum field theory in a nutshell. Princeton University Press, Princeton, NJ, USA, 2003. ISBN 0-691-01019-6. xv + 518 pp. LCCN QC174.45 .Z44 2003. URL http://www.

loc.gov/catdir/description/prin031/2002031743.html;   http:
//www.loc.gov/catdir/samples/prin031/2002031743.html;   http:
//www.loc.gov/catdir/toc/prin032/2002031743.html.

Ghatak:2004:QMH

[4666] Ajoy Kumar Ghatak and S. Lokanathan. Quantum mechanics: theory and applications, volume 137 of Fundamental theories of physics. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2004. ISBN 1-4020-1850-9 (hardcover), 1-4020-2129-1 (paperback). ISSN 0168-1222. xlviii + 858 pp. LCCN QC174.12 .G485

2004. URL http://catdir.loc.gov/catdir/enhancements/fy0822/
2004041246d.html;   http://catdir.loc.gov/catdir/enhancements/
fy0822/2004041246t.html;   http://www.gbv.de/dms/bowker/toc/
9781402018503.pdf;   http://www.zentralblattmath.org/zmath/
en/search/?an=1059.81001.

LeBellac:2004:ENE

[4667] Michel Le Bellac, Fabrice Mortessagne, and Ghassan George Batrouni. Equilibrium and non-equilibrium statistical thermodynamics. Cambridge University Press, Cambridge, UK, 2004. ISBN 0-511-19228-2 (e-book), 0-511-64817-0 (e-book), 0-511-60657-5 (e-book), 0-511-19370-X, 0-521-82143-6 (hardcover), 0-511-19444-7. xvi + 616 pp. LCCN QC311.5 L398 2004; QC318.T47 L43 2004. URL http://catdir.loc.gov/

catdir/description/cam041/2003055894.html;   http://catdir.

loc.gov/catdir/enhancements/fy0732/2003055894b.html;   http:
//catdir.loc.gov/catdir/toc/cam041/2003055894.html;   http:
//site.ebrary.com/id/10130372;   http://www.gbv.de/dms/bowker/
toc/9780521821438.pdf;   http://www.loc.gov/catdir/description/
cam041/2003055894.html;   http://www.loc.gov/catdir/toc/cam041/
2003055894.html;   http://www.zentralblattmath.org/zmath/en/
search/?an=1103.82001.

Rindler:2004:RSG

[4668] Wolfgang Rindler. Relativity: special, general, and cosmological. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2004. ISBN 0-19-850835-2 (hardcover), 0-19-850836-0 (paperback). xiii + 428

pp. LCCN QC173.55. URL http://www.gbv.de/dms/bowker/toc/
9780198508366.pdf.

Weber:2004:EMM

[4669] Hans-Jürgen Weber and George B. (George Brown) Arfken. Essential mathematical methods for physicists. Elsevier Academic Press, Amsterdam, The Netherlands, 2004. ISBN 0-08-046985-X (e-book), 0-12-059877-9. xxii + 932 pp. LCCN QC20 W43 2004; QA37.2 .A74

2004. URL http://catdir.loc.gov/catdir/description/els041/
2003277216.html;   http://catdir.loc.gov/catdir/toc/els041/
2003277216.html;   http://site.ebrary.com/id/10167088;   http:
//www.loc.gov/catdir/description/els041/2003277216.html;   http:
//www.loc.gov/catdir/toc/els041/2003277216.html.

Greene:2005:FCS

[4670] Brian Greene. The Fabric of the Cosmos: Space, Time, and the Texture of Reality. Vintage Books, New York, NY, USA, 2005. ISBN 0-375-72720-5. xii + 569 pp. LCCN QB982 .G74 2004.

Griffiths:2005:IQM

[4671] David J. (David Jeffery) Griffiths. Introduction to Quantum Mechanics. Pearson Prentice Hall, Upper Saddle River, NJ 07458, USA, second edition, 2005. ISBN 0-13-111892-7. ix + 468 pp. LCCN QC174.12 .G75 2005.

Leader:2005:SPP

[4672] Elliot Leader. Spin in particle physics, volume 15 of Cambridge monographs on particle physics, nuclear physics and. Cambridge University Press, Cambridge, UK, 2005. ISBN 0-521-02077-8, 0-521-35281-9 (hardcover). xxi + 526 pp. LCCN ????

Yourgrau:2005:WTF

[4673] Palle Yourgrau. A World Without Time: the Forgotten Legacy of Gödel and Einstein. Basic Books, New York, NY, USA, 2005. ISBN 0-465-09293-4. viii + 210 + 4 pp. LCCN BD638 .Y73 2005.

Breuer:2006:TOQ

[4674] Heinz-Peter Breuer and Francesco Petruccione. The theory of open quantum systems. Clarendon Press, New York, NY, USA, 2006. ISBN 0-19-921390-9 (paperback), 0-19-852063-8 (hardcover). xxi + 625 pp. LCCN QC174.85.O6.

Dodelson:2006:MC

[4675] Scott Dodelson. Modern cosmology. Academic Press, New York, USA, 4. [print.]. edition, 2006. ISBN 0-12-219141-2. xiii + 440 pp. LCCN ????

LeBellac:2006:ENEb

[4676] Michel Le Bellac, Fabrice Mortessagne, and Ghassan George Batrouni. Equilibrium and non-equilibrium statistical thermodynamics. Cambridge University Press, Cambridge, UK, reprint edition, 2006. ISBN 0-521-82143-6 (hardcover). xvi + 616 pp. LCCN ????

Stockmann:2006:QCI

[4677] Hans-Jürgen Stöckmann. Quantum chaos: an introduction. Cambridge University Press, Cambridge, UK, 2006. ISBN 0-521-59284-4 (hardcover), 0-521-02715-2. xi + 368 pp. LCCN ???? URL http://www.

gbv.de/dms/bowker/toc/9780521592840.pdf.

Breuer:2007:TOQ

[4678] Heinz-Peter Breuer and Francesco Petruccione. The Theory of Open Quantum Systems. Clarendon Press, New York, NY, USA, 2007. ISBN 0-19-921390-9 (paperback), 0-19-852063-8 (hardcover). xxi + 613 pp. LCCN ????

Dodelson:2007:MC

[4679] Scott Dodelson. Modern cosmology. Academic Press, New York, USA, 2007. ISBN 0-12-219141-2. xiii + 440 pp. LCCN QB981.

Doran:2007:GAP

[4680] Chris (Chris J. L.) Doran and A. N. (Anthony N.) Lasenby. Geometric Algebra for Physicists. Cambridge University Press, Cambridge, UK, 2007. ISBN 0-521-71595-4 (paperback). xiv + 578 pp. LCCN QC20.7.C55 D67 2007.

Breuer:2010:TOQ

[4681] Heinz-Peter Breuer and Francesco Petruccione. The Theory of Open Quantum Systems. Clarendon Press, New York, NY, USA, 2010. ISBN 0-19-852063-8, 0-19-921390-9 (paperback). xxi + 613 pp. LCCN ????

LeBellac:2010:ENEb

[4682] Michel Le Bellac, Fabrice Mortessagne, and Ghassan George Batrouni. Equilibrium and Non-Equilibrium Statistical Thermodynamics. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-82143-6, 0-521-52895-X (paperback). xvi + 616 pp. LCCN QC311.5 L398 2010.

vanDongen:2010:EU

[4683] Jeroen van Dongen. Einstein’s unification. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-88346-6. x + 213 pp. LCCN

QC173.6 .D66 2010. URL http://assets.cambridge.org/97805218/
83467/cover/9780521883467.jpg;   http://www.loc.gov/catdir/
enhancements/fy1009/2010017820-b.html;   http://www.loc.gov/
catdir/enhancements/fy1009/2010017820-d.html;   http://www.
loc.gov/catdir/enhancements/fy1009/2010017820-t.html.

Neuenschwander:2011:ENW

[4684] Dwight E. Neuenschwander. Emmy Noether’s wonderful theorem. The Johns Hopkins University Press, Baltimore, MD, USA, 2011. ISBN 0-8018-9694-0, 0-8018-9693-2 (hardcover). xviii + 243 pp.

LCCN QC174.17.S9 N48 2011. URL http://www.loc.gov/catdir/
enhancements/fy1112/2010923134-b.html;   http://www.loc.gov/
catdir/enhancements/fy1112/2010923134-d.html.

Redei:2001:AGR

[4685] Miklós Rédei and Michael Stöltzner, editors. John von Neumann and the foundations of quantum physics (Budapest, 1999), volume 8 of Vienna Circle Institute yearbook. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2001. ISBN 0-7923-6812-6. ISSN 0929-6328. ix + 371 pp. LCCN ????

Redei:2001:JNF

[4686] Miklós Rédei and Michael Stöltzner, editors. John von Neumann and the foundations of quantum physics, volume 8 of Vienna Circle Institute Yearbook. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2001. ISBN 0-7923-6812-6. ix + 371 pp. LCCN QC173.98 .J65 2001.

Bostrom:2002:ABO

[4687] Nick Bostrom. Anthropic bias: observation selection effects in science and philosophy. Studies in philosophy. Routledge, London, UK, 2002. ISBN 0-415-93858-9. xiii + 224 pp. LCCN BD241 .B657 2002. URL

http://www.loc.gov/catdir/enhancements/fy0647/2001058887-d.
html.

Hawking:2002:FS

[4688] S. W. (Stephen W.) Hawking, Kip S. Thorne, Igor D. Novikov, Timothy Ferris, Alan Lightman, and Richard Price, editors. The future of spacetime. W. W. Norton & Co., New York, NY, USA, 2002. ISBN 0-393-02022-3. 220 pp. LCCN QC173.59.S65 F87 2002.

Rizzi:2004:RRF

[4689] Guido Rizzi and Matteo Luca Ruggiero, editors. Relativity in rotating frames: relativistic physics in rotating reference frames, volume 135 of Fundamental theories of physics. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2004. ISBN 1-4020-1805-3 (hardcover). ISSN 0168-1222. xxiii + 452 pp. LCCN QC173.585 .R45 2004. URL http://

catdir.loc.gov/catdir/enhancements/fy0822/2003064119d.html;
 http://www.gbv.de/dms/bowker/toc/9781402018053.pdf;   http:
//www.loc.gov/catdir/enhancements/fy0822/2003064119-d.html;
 http://www.zentralblattmath.org/zmath/en/search/?an=1050.83005.

Camilleri:2009:HIQ

[4690] Kristian Camilleri. Heisenberg and the interpretation of quantum mechanics: the physicist as philosopher. Cambridge University Press, Cambridge, UK, 2009. ISBN 0-521-88484-5 (hardcover). xii + 199 pp.

LCCN QC16.H395 .C34 2009. URL http://www.loc.gov/catdir/toc/
ecip0825/2008034113.html.

Sandberg:2016:DWE

[4691] Anders Sandberg, Stuart Armstrong, and Milan M. Ćirkoviç. That is not dead which eternal lie: The aestivation hypothesis for resolving Fermi’s Paradox. Journal of the British Interplanetary Society, 69:406–415, 2016. CODEN JBISAW. ISSN 0007-084X. URL https://arxiv.org/abs/

1705.03394;   https://www.jbis.org.uk/paper/2016.69.406. See

comment [4187].

Sebens:2023:EES

[4692] Charles T. Sebens. Eliminating electron self-repulsion. Foundations of Physics, 53(4):??, August 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-023-00702-0.

Lawrence:2023:PQM

[4693] Jay Lawrence. Pointers for quantum measurement theory. Foundations of Physics, 53(4):??, August 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-023-00707-9.

Joshi:2023:EP

[4694] Mahendra Joshi. Evolution via projection. Foundations of Physics, 53(4):??, August 2023. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-023-00706-w.

Acharya:2023:OCI

[4695] Arnab Acharya, Pratik Jeware, and Soumitro Banerjee. Objective collapse induced by a macroscopic object. Foundations of Physics, 53 (4):??, August 2023. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-023-00709-7.

Jha:2023:CFT

[4696] Aditya Jha, Douglas Campbell, Clemency Montelle, and Phillip L. Wilson. On the continuum fallacy: Is temperature a continuous function? Foundations of Physics, 53(4):??, August 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-023-00713-x.

Lyons:2023:RST

[4697] Tony Lyons. Relational space–time and de Broglie waves. Foundations of Physics, 53(4):??, August 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-023-00715-9.

Boyer:2023:CEI

[4698] Timothy H. Boyer. Classical electromagnetic interaction of a charge with a solenoid or toroid. Foundations of Physics, 53(4):??, August 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-023-00712-y.

Drezet:2023:TSS

[4699] Aurélien Drezet. A time–symmetric soliton dynamics à la de Broglie. Foundations of Physics, 53(4):??, August 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-023-00711-z.

Torres:2023:NIQ

[4700] Fernanda Torres, Sujoy K. Modak, and Alfredo Aranda. New insights on the quantum–classical division in light of collapse models. Foundations of Physics, 53(4):??, August 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-023-00716-8.

Park:2023:BHE

[4701] I. Y. Park. Black hole entropy from non-dirichlet sectors, and a bounce solution. Foundations of Physics, 53(4):??, August 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-023-00719-5.

Zhong:2023:QEA

[4702] Shengyang Zhong. Quantum entanglement: An analysis via the orthogonality relation. Foundations of Physics, 53(4):??, August 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-023-00710-0.

Linford:2023:BC

[4703] Daniel Linford. On the boundary of the Cosmos. Foundations of Physics, 53(4):??, August 2023. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-023-00718-6.

Schneider:2023:EDE

[4704] Mike D. Schneider. On efforts to decouple early universe cosmology and quantum gravity phenomenology. Foundations of Physics, 53(4):??, August 2023. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-023-00720-y.

Bersons:2023:MMP

[4705] Imants Bersons, Rita Veilande, and Ojars Balcers. Mathematical models of photons. Foundations of Physics, 53(4):??, August 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-023-00708-8.

vandeVen:2023:EPN

[4706] Christiaan J. F. van de Ven. Emergent phenomena in nature: a paradox with theory? Foundations of Physics, 53(5):??, October 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-023-00721-x.

Smith:2023:EWT

[4707] Leslie Smith. Electron wave trajectories within Schrödinger’s hydrogen atom, and relativistic consequences. Foundations of Physics, 53 (5):??, October 2023. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-023-00722-w.

Golovko:2023:EMM

[4708] V. A. Golovko. Is the electron magnetic moment unique? Foundations of Physics, 53(5):??, October 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-023-00723-9.

Gurwich:2023:NVI

[4709] Ioseph Gurwich. The non-vanishing imprint of gravitational waves as the result of its nonlinear evolution in space. Foundations of Physics, 53(5):??, October 2023. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-023-00714-w.

Ruggiero:2023:SFT

[4710] Matteo Luca Ruggiero. Synchronization and fundamental time: a connection between relativity and quantum mechanics. Foundations of Physics, 53(5):??, October 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.com/

article/10.1007/s10701-023-00724-8.

Khrennikov:2023:EOQ

[4711] Andrei Khrennikov and Irina Basieva. Entanglement of observables: Quantum conditional probability approach. Foundations of Physics, 53(5):??, October 2023. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-023-00725-7.

Buchholz:2023:ATQ

[4712] Detlev Buchholz and Klaus Fredenhagen. Arrow of time and quantum physics. Foundations of Physics, 53(5):??, October 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-023-00728-4.

Handel:2023:CDM

[4713] Peter H. Handel and Klara E. Splett. Calculation of dark matter as a feature of space–time. Foundations of Physics, 53(5):??, October 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-023-00705-x.

Gao:2023:SSP

[4714] Shan Gao. Is superluminal signaling possible in collapse theories of quantum mechanics? Foundations of Physics, 53(5):??, October 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-023-00729-3.

Ravelonjato:2023:QRC

[4715] Rivo Herivola Manjakamanana Ravelonjato, Ravo Tokiniaina Ranaivoson, Raoelina Andriambololona, Roland Raboanary, Hanitriarivo Rakotoson, and Naivo Rabesiranana. Quantum and relativistic corrections to Maxwell–Boltzmann ideal gas model from a quantum phase space approach. Foundations of Physics, 53(5):??, October 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-023-00727-5.

Hardel:2023:RQE

[4716] Vincent Hardel, Paul-Antoine Hervieux, and Giovanni Manfredi. Relaxation to quantum equilibrium and the Born rule in Nelson’s stochastic dynamics. Foundations of Physics, 53(6):??, December 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-023-00730-w.

Carroll:2023:CDF

[4717] Sean M. Carroll. Completely discretized, finite quantum mechanics. Foundations of Physics, 53(6):??, December 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-023-00726-6.

Gao:2023:COB

[4718] Shan Gao. Can the ontology of Bohmian mechanics consists only in particles? The PBR theorem says no. Foundations of Physics, 53 (6):??, December 2023. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-023-00731-9.

Silagadze:2023:EQM

[4719] Zurab K. Silagadze. Evading quantum mechanics à la Sudarshan: Quantum-mechanics-free subsystem as a realization of Koopman– von Neumann mechanics. Foundations of Physics, 53(6):??, December 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-023-00734-6.

Helbig:2023:LMF

[4720] Phillip Helbig. Life, the multiverse, and fine-tuning. Foundations of Physics, 53(6):??, December 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-023-00732-8.

Ellis:2024:PTH

[4721] George F. R. Ellis. Physical time and human time. Foundations of Physics, 54(1):??, February 2024. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.com/

article/10.1007/s10701-023-00738-2.

Lopez-Castillo:2024:QSA

[4722] Alejandro López-Castillo. On the quasi-separability of atoms and molecules. Foundations of Physics, 54(1):??, February 2024. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-023-00736-4.

Helland:2024:AFQ

[4723] Inge S. Helland. An alternative foundation of quantum theory. Foundations of Physics, 54(1):??, February 2024. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-023-00735-5.

Piccione:2024:PNK

[4724] Nicolò Piccione. A proposal for a new kind of spontaneous collapse model. Foundations of Physics, 54(1):??, February 2024. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-023-00739-1.

Drezet:2024:CAR

[4725] Aurélien Drezet. A critical analysis of Relative Facts Do Not Exist: Relational Quantum Mechanics Is Incompatible with Quantum Mechanics by Jay Lawrence, Marcin Markiewicz and Marek Źukowski. Foundations of Physics, 54(1):??, February 2024. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-023-00743-5.

Menon:2024:SRR

[4726] Tushar Menon and James Read. Some remarks on recent formalist responses to the hole argument. Foundations of Physics, 54(1): ??, February 2024. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL https://link.springer.com/article/10.
1007/s10701-023-00746-2.

Gao:2024:UMS

[4727] Shan Gao. Is the universe in a mixed state? Foundations of Physics, 54(1):??, February 2024. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-023-00742-6.

Scharnhorst:2024:MPH

[4728] Jordan Scharnhorst and Anthony Aguirre. Modeling the past hypothesis: a mechanical cosmology. Foundations of Physics, 54(1):??, February 2024. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-023-00745-3.

Niestegge:2024:NST

[4729] Gerd Niestegge. Non-symmetric transition probability in generalized qubit models. Foundations of Physics, 54(1):??, February 2024. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-023-00744-4.

Etesi:2024:STA

[4730] Gábor Etesi. A set-theoretic analysis of the black hole entropy puzzle. Foundations of Physics, 54(1):??, February 2024. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-023-00737-3.

Soulas:2024:DHD

[4731] Antoine Soulas. Decoherence as a high-dimensional geometrical phenomenon. Foundations of Physics, 54(1):??, February 2024. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https:

//link.springer.com/article/10.1007/s10701-023-00740-8.

Bettini:2024:LP

[4732] Alessandro Bettini. Learning from paradoxes. Foundations of Physics, 54(1):??, February 2024. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-023-00733-7.

Adlam:2024:WDN

[4733] Emily Adlam. What does ‘(Non)-absoluteness of observed events’ mean? Foundations of Physics, 54(1):??, February 2024. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-023-00747-1.

Carcassi:2024:RQS

[4734] Gabriele Carcassi, Andrea Oldofredi, and Christine A. Aidala. On the reality of the quantum state once again: a no-go theorem for ψ-ontic models. Foundations of Physics, 54(1):??, February 2024. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-023-00748-0.

Vigneron:2024:NRR

[4735] Quentin Vigneron. Non-relativistic regime and topology: Topological term in the Einstein equation. Foundations of Physics, 54(1):??, February 2024. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-023-00749-z.

Novello:2024:FUG

[4736] M. Novello and V. Antunes. Fundamental units in gravitational, electromagnetic and weak (Fermi) interactions. Foundations of Physics, 54(1):??, February 2024. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-023-00741-7.

Maccone:2024:SCQ

[4737] Lorenzo Maccone. Schrödinger cats and quantum complementarity. Foundations of Physics, 54(1):??, February 2024. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-023-00750-6.

Elkhateeb:2024:RGC

[4738] Esraa Ali Elkhateeb. Reconstruction of f(R) gravity from cosmological unified dark fluid model. Foundations of Physics, 54(1):??, February 2024. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.1007/
s10701-023-00751-5.

Gao:2024:WGP

[4739] Shan Gao. Why the global phase is not real. Foundations of Physics, 54(2):??, April 2024. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-024-00754-w.